A Complete Bibliography of the Journal of the Royal Statistical Society, Series B family

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 October 2024 Version 1.27

Title word cross-reference

1/ω [2140]. 2 [632, 275]. 2^{K} [3321]. 2^{m} × 3 [1084]. 2^{n−m} [1208]. 2 × 2 [937, 676, 2024, 2637, 138]. 2 × 2 × 2 ⋯ [47]. 3 [3621]. ^{2} [3231]. ^{n} [826, 938]. _{k} [363, 644]. λ [774]. _{O} [2606]. A_{n} [2088]. α [3282, 3003]. AR(p) [3064]. b < v [1432]. b_{2} [1323]. β [3658]. c [969, 3520]. χ^{2} [307, 1147, 960, 230, 70, 13, 699]. D [2201, 2362, 1565, 1226, 1496, 1065, 1193, 1157]. E [1446, 1447, 1552, 3905, 3729]. E(r_{s}) [237]. E(s^{2}) [2600]. ϵ [1944]. F [791, 1194, 1147, 1253, 1277]. G [3568, 2124]. h [3320]. k [1892, 1916, 1293, 1613]. k^{n} [886, 1117]. L [1906, 2506, 3520]. L_{1} [2960, 3452, 3876]. L_{2} [1625]. λ [1615]. M [3348, 770, 1815, 1696, 2161, 1397, 1459, 1572, 1521]. R^{n} [3424]. N [380, 381, 878, 3327]. N = 3 [246]. NEAR(2) [1845, 1770]. NLAR(2) [1845]. P [2524, 3238, 3441, 884, 860, 3961, 3220, 3327]. P − pp′ [2193]. πP S [1179, 832, 929, 904]. q^{n−m} [2206]. R^{d} [3052]. ρ [3412]. r × s [968]. S [3748]. √b_{1} [1323]. T [1237, 2813, 1070, 2704, 3155, 1147, 3437, 2689, 1094]. T^{2} [629]. τ [3412]. U [2368, 2431, 3986, 376]. V [515]. φ_{k} [3520]. X^{2} [885]. Z [3769].

-Approximation [699]. -Associate [1459, 1696]. -ball [3876]. -Component [860]. -Concurrence [1572]. -Contaminations [1944]. -designs [2206]. -Dimensional [1397, 1521]. -Distribution [2704, 2689]. -Estimator [2506, 3348].

-Estimators [1815, 2161]. -filter [3441]. -Independence [1194]. -Investing [3003, 3282]. -likelihood [3320]. -Mask [515]. -Method [1237]. -model [3658]. -Moments [1906]. -Noise [2140]. -optimal [3520, 1552, 1226, 1496, 2124, 2600]. -Optimality [2201, 1446, 1447]. -Optimum [2362, 1565, 2813, 1065]. -Out [1892, 1916]. -Plots [2524]. -Regularization [2960]. -Sample [1293, 969, 275]. -shaped [3748]. -splines [3238]. -Statistic [1157, 3155]. -Statistics [2368, 2431, 3986]. -Test [1094, 1253]. -Tests [1277]. -Type [791]. -value [3220, 3769]. -values [3961, 3905, 3729]. -Variates [1070]. -year [3437].

/1 [644, 774]. /M/1 [363].

1 [1500, 465, 1853, 363, 774, 1046]. 15th [80]. 18 [382]. 19 [3759]. 1935 [32]. 1936 [50]. 1938 [80]. 1945 [111].

23rd [4]. 26th [50]. 28th [32].

30th [89].

4th [111].

611 [3654].

78 [3654].

= [2756].

A. [3953]. Aalen [3173]. ABC [3638]. aberrant [3639]. Aberration [2441]. absence [3396]. Absolute [2840]. Absorbing [530]. abundance [3508]. Accelerated [1842, 2623]. Acceleration [520, 2346]. Acceptance [1431, 1181]. Accident [103, 264, 798, 757]. Accommodate [2536]. Accommodating [2735]. Accommodation [2981].

Accounting [42, 2263, 3991]. Accuracy [988, 1896, 2607, 167, 20, 3155, 3317]. Accurate [1947, 2261, 3355]. Achievable [2485]. Achieve [1428]. achieved [3139]. Achieving [3189]. Acid [2207]. Acknowledgement [2357, 2358, 850, 2431, 2824, 2825, 798]. Acknowledgment [2824, 163]. Acquired [2190, 1966]. across [3268, 2793, 2873, 3451, 3392]. active [3264]. activity [3843]. Acyclic [2751, 3108, 3306]. Adams [3922]. adapt [3512, 3493]. Adaptation [2217]. Adaptive [2767, 573, 3163, 2821, 2683, 3942, 3599, 3184, 2466, 2518, 2808, 2761, 2522, 2687, 2472, 2647, 3433, 3544, 2382, 3277, 3377, 3639, 3769, 3164, 3518, 3596, 3882, 3588, 3814, 3609, 3193]. Adaptively [3086]. Addenda [929]. Addendum [2864, 816, 502, 1486, 2865]. Adding [2294]. Addition [75, 1225]. Additional [734]. Additions [2924]. Additive [2046, 1352, 2778, 2462, 3421, 3113, 2760, 3021, 3095, 2397, 1460, 2302, 3008, 2286, 519, 3312, 3366, 3166, 3378, 3173, 3274]. Additivity [614]. Adequacy [2656]. Adequate [1832]. Adjacency [2848]. Adjusted [2144, 2898, 2092, 2605, 2170, 2262, 1767, 2149, 2065, 2495, 1923, 2705, 3196, 2558, 3792, 3533, 3629]. Adjusting [1834, 3230, 3298]. Adjustment [755, 816, 2149, 2792, 1962, 1921, 2352, 3697, 3569, 3580, 3776]. Adjustments [1675, 1781, 1818, 3216, 3903]. Admissibility [1650, 1223, 1101, 3422, 1467]. Admitting [313]. adoption [3689]. Adversarial [2122, 3679]. Affine [1812, 1872, 2152, 2153, 3204]. After [755, 816, 1996, 421, 501, 2741, 2971, 3739]. Against [1878, 1372, 940, 2894, 2574, 3403]. Age [2306, 204]. Age-Dependent [2306]. Age-Distribution [204]. agency [3943]. Agent [2026, 2011]. aggregate [3538].

Aggregated [3640]. aggregation [3423]. Agricultural [4, 3, 80, 2484, 296, 91, 61, 50, 67, 25, 9, 6]. Agriculture [59]. Ahead [2790, 2325]. AIC [3193]. aid [3147]. Aids [140, 1010, 2048]. Air [115]. airstrikes [3771]. Aitkin [1813]. Akaike [2395, 2411, 1288, 1408]. Akilagun [3920]. al [2256, 2375, 2378, 3673, 3677, 3671]. al. [3670, 3669, 3672, 3674, 3675, 3676]. Alessandro [3979]. Alexander [3826]. Algebra [2409]. Algebraic [342, 1423]. Algebras [1471]. Algorithm [2453, 1578, 1286, 2072, 424, 502, 2273, 2346, 2220, 1549, 1873, 2345, 2496, 2467, 2960, 1708, 2164, 3544, 3071, 3175, 3176, 3231, 3191, 3764, 3518, 3504, 3605, 2637]. Algorithms [2831, 1399, 2336, 2957, 3497]. Ali [1532]. Alias [1279]. Aliasing [1268]. aligned [3573]. all-or-nothing [3310]. alleviation [3219]. Allocation [1513, 296, 1393, 1461, 741, 806, 2030, 1179, 922, 1677]. Allowing [708, 760, 864, 3227]. Almost [773, 783, 2557, 3149]. Alpha [3628, 3603]. also [2097]. Alternative [192, 991, 585, 358, 940, 1561, 1031, 394, 756, 2894, 3105, 935, 1380, 1685, 1164, 2385, 1933, 3779, 2459]. Alternatives [1093, 2098, 528, 1388, 1764, 1643, 2574, 1264, 1795]. Always [2301, 2862]. Alzheimer [3987]. Amalgamation [541]. AMF [3640]. Among [1440, 1688, 820, 1174, 949]. Amount [338, 2051]. Amplified [1913]. Amplitude [681]. Analogue [977]. Analogues [1099]. Analyses [756, 1891, 1994, 1277, 3780, 3792, 3662]. Analysing [2697, 3004, 2487, 1567]. Analysis [533, 1542, 937, 2042, 3852, 192, 1659, 2856, 310, 404, 285, 2639, 1456, 15, 36, 120, 144, 621, 1347, 1485, 1212, 1924, 1172, 1775, 2484, 441, 3828, 1003, 666, 22, 1336, 1818, 1479, 3117, 3829, 791, 3823, 3830, 2402, 1289, 1306, 1402, 2532, 2716, 2790, 3116, 3658, 3831, 2340, 429, 2377, 1230, 35, 54, 1294, 2787, 172, 415, 418, 470, 455, 482, 394, 2582, 2568, 2980, 1964, 2437, 2280, 1340, 2333, 2503, 1143, 1320, 1433, 234, 2162, 2478, 824, 114, 582, 2538, 629, 1674, 1688, 2109, 970, 1043, 1083, 1178, 424, 502, 1957, 2229, 1715, 233, 2082, 2581]. Analysis [2675, 855, 2874, 1015, 2260, 1261, 1328, 3827, 2461, 639, 680, 710, 3821, 1906, 2979, 2820, 14, 845, 2594, 369, 1966, 2293, 314, 628, 1355, 119, 185, 2159, 2912, 3079, 717, 3832, 2926, 1713, 1702, 2113, 685, 2893, 2554, 2796, 3833, 447, 1795, 1895, 2493, 3052, 1634, 2762, 414, 1944, 1638, 1466, 2417, 3834, 2794, 1676, 3835, 3822, 1019, 3836, 486, 2448, 569, 589, 1121, 1976, 2287, 947, 361, 750, 2361, 2418, 1136, 2508, 2942, 299, 2187, 392, 1080, 1400, 2241, 1998, 138, 186, 1721, 653, 2519, 2125, 187, 2477, 241, 1913, 1499, 3826, 3838, 2003, 2383, 3062]. Analysis [3824, 3839, 2304, 268, 1591, 820, 1777, 359, 37, 1029, 3009, 2869, 76, 84, 387, 3840, 2904, 3785, 2667, 3457, 3289, 3205, 3544, 3983, 3777, 3322, 3180, 3966, 3231, 3155, 3585, 3420, 3523, 3458, 3573, 3825, 2816, 3942, 3468, 3678, 3226, 3492, 3987, 3788, 3767, 3341, 3907, 3388, 2419, 3841, 3820, 3233, 2637, 3603, 3837, 3428, 3418, 3455, 3221, 3545, 3156, 3801, 3609, 3546, 3743, 3316, 3728, 3772, 3750]. Analyst [1010]. Analytical [806, 718, 3510]. Anastasios [3889]. Anastasiou [3921]. Ancestral [3010]. Anchor [3626, 3665]. Ancillarity [1221, 2334]. Ancillary [1031, 2947]. and/or [3959]. Andi [3918]. Andreas [3921, 3710]. Andrej [3971]. Angelopoulos [3889]. angle [3234]. Angular [1305]. Angular-Linear [1305]. Animal [349, 2664, 79]. Animals [2258]. Anisotropic [2109]. Anna [3711]. Annealing [2712]. ANOVA [1401]. Anru [3825]. Anthony [3890]. Anthropometry [249]. Antoniano [3857]. Antonio [3862, 3972]. any [264, 3644, 55, 62]. Apple [74]. Apples [74]. Applicable [373].

Application [22, 1402, 2278, 724, 61, 964, 3085, 2029, 1362, 114, 2048, 99, 936, 124, 2756, 2569, 92, 33, 2968, 984, 1997, 1304, 2850, 1838, 1324, 881, 1493, 3010, 1708, 5, 2927, 202, 2890, 266, 1752, 281, 2081, 3034, 3282, 3170, 3770, 3180, 3962, 3389, 3359, 3987, 3608, 3193, 1060]. Applications [2475, 1412, 3044, 1747, 2755, 42, 656, 2884, 2232, 863, 2520, 1478, 1147, 2061, 2379, 1583, 1062, 2689, 628, 1632, 1953, 328, 497, 1357, 1743, 2522, 561, 1322, 1088, 707, 2004, 508, 19, 880, 2143, 1722, 1826, 2904, 3245, 3394, 3575, 3983, 3393, 3416, 2816, 3623, 3621, 3767, 3046, 3372, 3940, 3165, 3146]. Applied [59, 158, 410, 411, 501, 44, 131, 2409, 51, 1042, 1066, 55, 62, 1162, 1488, 2261]. Appointment [250]. Approach [2426, 2844, 2213, 1999, 2662, 2997, 2532, 2599, 2485, 2096, 946, 794, 888, 2670, 2592, 771, 772, 2234, 2509, 2325, 2782, 2245, 1891, 782, 2926, 3036, 2893, 2554, 2958, 891, 399, 2794, 548, 1813, 886, 1117, 2966, 2984, 2638, 3050, 1684, 1920, 2292, 769, 2652, 2759, 2825, 2944, 949, 1913, 2729, 3051, 2304, 781, 2129, 3009, 2916, 3234, 3473, 3139, 3339, 3377, 3639, 3578, 3477, 3237, 3161, 3606, 3286, 3630, 3682, 3140, 3121, 3330, 2824, 3608, 3413, 2499, 3303, 3910, 3193]. Approaches [664, 2787, 2788, 3883, 3489]. Appropriateness [1579, 1692]. Approximate [1245, 1533, 3491, 1934, 638, 3526, 1428, 1787, 2232, 503, 1857, 2663, 2228, 584, 1092, 2247, 2127, 2967, 2350, 302, 1997, 1544, 1962, 2136, 1297, 2043, 168, 3653, 3060, 269, 2261, 1495, 3845, 3415, 3194, 3576, 3623, 3744, 3249, 3597, 616]. approximately [3637]. Approximating [1601, 1263, 2802, 2766, 2087]. Approximation [401, 699, 724, 850, 804, 1971, 1352, 2584, 1481, 2770, 2460, 1638, 396, 289, 2246, 301, 698, 390, 3764, 3741, 3182, 3964]. Approximations [1412, 307, 1202, 2069, 2624, 2750, 2617, 94, 1945, 1966, 1100, 2474, 1800, 1885, 3073, 2391, 1537, 2161, 3060, 1159, 2459, 3653]. Arbitrarily [1281]. Arbitrary [2533, 411, 1059]. Arc [360]. Archimedean [3406]. Area [2882, 2993, 2401, 2779, 3255, 3787, 3620]. Areal [2274, 2968, 3168]. Areas [609, 1263]. Argument [599, 936]. Arguments [1403, 597]. Arising [165, 54, 2609, 362, 1476, 227]. Arithmetic [516, 566]. ARL [1235]. Arm [2915]. ARMA [1867, 1505, 1629, 1776]. Armed [1438, 3849]. Armitage [2001]. arms [3606]. Army [123]. Arrangements [47, 1206, 1205, 1148, 2664, 1279, 139]. Array [2884, 2691]. Arrays [2924, 139]. Arrival [411, 491, 477]. Arrivals [474, 866, 1111, 476, 784]. Art [3897]. Arthur [3980]. Article [365]. Artificial [190]. Ascent [2830]. Ascent-based [2830]. Ascolani [3862, 3972]. Aspects [599, 573, 792, 1043, 249, 2615, 789, 1684, 3921, 3938, 3922, 3923, 3924, 3920, 3925, 3926, 3927, 3928, 3929, 3937, 3916, 3919, 3930, 3914, 3931, 3932, 3933, 3934, 3935, 3918, 3913, 3936, 3917, 3915]. Assay [134]. Assays [51]. assess [3171]. Assessing [2275, 2907, 3083, 1612, 2665, 1592, 1170, 3636]. Assessment [1760, 2199, 2571, 2389, 2439, 1169, 1256, 3734, 3731]. Assessments [2272]. Assisted [2844, 3525, 3454, 3578, 3662]. Associate [1459, 776, 1696]. Associated [481, 1659, 2605, 355, 1519, 862, 2495, 1180, 1050, 1103, 372, 545, 1629]. Association [703, 704, 1005, 968, 676, 705, 3266, 2114, 960, 776, 1568, 2696, 820, 3185, 3538, 739]. Associations [136, 100, 3509]. Assumed [3116]. Assumption [3708, 3706, 3709, 309, 3710, 3711, 3702, 3703, 3704, 3712, 3707, 3713, 3714, 3715, 3716, 3717, 3718, 3719, 3720, 3705, 3721, 3722, 3701, 3725, 3723, 3724].

Assumption-lean [3708, 3706, 3709, 3710, 3711, 3702, 3703, 3704, 3712, 3707, 3713, 3714, 3715, 3716, 3717, 3718, 3719, 3720, 3705, 3721, 3722, 3701, 3725, 3723, 3724]. Assumptions [1765, 2091, 3396, 3329]. astronomy [3165]. Asymmetrical [1091, 152, 1206]. Asymptopia [2216]. Asymptotic [748, 995, 153, 2104, 1670, 1011, 2052, 1158, 1291, 754, 1202, 1742, 2264, 1749, 1430, 513, 1394, 2024, 986, 2538, 812, 957, 1055, 357, 407, 809, 1395, 1369, 1692, 1356, 3065, 643, 1113, 1953, 2308, 975, 882, 1723, 3313, 1743, 1645, 1458, 1954, 564, 3169, 2957, 565, 2822, 1805, 544, 1288, 376, 1806, 1629, 1738, 245, 610, 815, 906, 916, 2511, 1664, 3500, 2289, 3539, 1023, 1487, 849, 1772]. Asymptotically [1082, 1477, 2218, 2030, 420, 878, 778, 610, 815, 1545, 1094, 3178]. Asymptotics [3245, 2425, 2178, 2691, 2037, 3785, 3678, 3760]. Asynchronous [1220, 3874, 3322]. Atkinson [2256]. Atom [635]. Attainment [210]. Attention [212, 251]. attractive [3355]. Attributes [56, 2799]. Augmented [3171]. Augusto [3864, 3974]. Aurelien [3925]. Author [2369, 3938, 3981, 3677, 3957, 3870, 3937, 3725, 3902]. Authors [1915, 2561, 2677, 2809, 2432, 2498, 2618, 2742, 2863, 2918, 2974, 3037, 3097, 3141, 3174, 3210, 3250, 3291, 3333, 3382, 3459, 3513, 3556, 3613]. Auto [1056, 1757, 3424, 3064, 3320, 3288, 3610, 3407, 3732, 3446, 3906, 3356]. auto-grouping [3732]. Auto-Logistic [1056]. auto-regression [3064]. auto-regressions [3320, 3610]. auto-regressive [3424, 3288, 3407, 3446, 3906, 3356]. Auto-Repeat [1757]. Autocorrelated [113, 125, 1102, 568, 439, 2397, 359]. Autocorrelation [1398, 1473]. Autocorrelations [147, 1586, 1505, 1364, 1020]. Autocovariance [1909, 3170].

Autocovariances [2089, 3249]. Autologistic [2694]. Automated [2453, 2180, 3347]. Automatic [2398, 466, 3911, 3937, 281, 3194]. Autoregression [1932, 1988, 1396, 1443, 1455]. Autoregressions [2533, 1879, 1948, 2908]. Autoregressive [1595, 2608, 189, 267, 2145, 2240, 158, 588, 295, 254, 1369, 2778, 2491, 2528, 1439, 1702, 2752, 2840, 2150, 1909, 2089, 150, 1705, 1020, 987, 1458, 475, 1998, 2429, 188, 245, 2925, 2507, 3817, 3842, 3819]. Autoregressive-Moving [1439, 1020]. Auxiliary [2458, 832, 929, 1883, 3497, 3464]. Available [3020]. Average [2240, 1598, 1307, 1363, 1506, 1562, 640, 2540, 426, 515, 822, 823, 1439, 2150, 1909, 2089, 1297, 1923, 1020, 1232, 137, 126, 3491, 3424, 3364, 3309, 3407, 2637, 3487, 3356]. Averages [1251, 540, 839, 181]. Averaging [2654, 3547]. Avoiding [2741]. Aware [2322]. Axes [2064]. Axial [1141]. Axiom [1389]. Axioms [1666]. axis [3573]. axis-aligned [3573]. Ayla [3928].

B [3899, 3654, 3220]. Back

[23, 30, 39, 48, 57, 65, 87, 335, 506, 593, 619, 740, 767, 780, 800, 819, 835, 852, 871, 890, 911, 933, 952, 979, 982, 999, 1024, 1027, 1041, 1052, 1077, 1079, 1098, 1137, 1139, 1153, 1168, 1195, 1197, 1214, 1242, 1257, 1259, 1284, 1298, 1313, 1330, 1332, 1345, 1374, 1377, 1409, 1411, 1435, 1453, 1468, 1489, 1511, 1525, 1539, 1555, 1569, 1589, 1608, 1621, 1641, 1661, 1681, 1700, 1724, 1744, 1758, 1773, 1784, 1797, 1807, 1820, 1836, 1846, 1860, 1875, 1889, 1899, 1917, 1927, 1937, 1955, 1973]. Back [1991, 2012, 2034, 2056, 2079, 2101, 2116, 2133, 2157, 2171, 2182, 2195, 2214, 2225, 2237, 2250, 2270, 2284, 2296, 2310, 2329, 2343, 2359, 2370, 2393, 2407, 2420, 2434, 2454, 2469, 2482, 2500, 2514, 2529, 2543, 2562, 2575, 2589, 2601, 2619, 2631, 2643, 2659, 2678, 2700, 2717, 2730, 2744, 2764, 2780, 2797, 2810, 2826, 2835, 2845, 2854, 2866, 2877, 2887, 2900, 2909, 2919, 2928, 2939, 2952, 2964, 2976, 2994, 3005, 3015, 3026, 3039, 3058, 3068, 3080, 3088, 3099, 3110, 3118, 3124, 3133, 3142, 2388]. Back-Fitting [2388]. Backfitting [2815]. Bacterial [239, 605, 240]. Bacteriology [20]. Bag [2253]. Bagged [2839]. Bahadur [1856]. Balance [1732, 1832, 3466]. Balanced [2223, 1459, 1576, 2112, 1386, 1912, 793, 2695, 1829, 877, 1626, 1240, 1696, 392, 1755, 1791, 1531, 69, 2351]. balancing [3491, 3364, 3263]. ball [3876]. Ballot [707]. Bancroft [991]. band [3756]. Banded [3170]. Bandit [1393, 1441, 3849]. Bandits [1833, 2248, 1438]. Bands [2951, 2077, 1029, 2430, 3473, 3300, 3256, 3810]. Bandwidth [2076, 2531, 2217, 408, 1914, 2669, 2204, 447, 2876, 1987, 1968, 2486, 3591, 3910]. bandwidth-free [3910]. Banerjee [3973]. Barker [3694]. Barnard [1061]. Barndorff [2733]. Barndorff-Nielsen [2733]. Barriers [530]. Bartlett [1675, 1781, 2170, 382, 2049, 2149, 670, 1962, 2558]. Bartlett-Type [2149]. Base [2971, 2890]. Base-Line [2971, 2890]. Based [2481, 2577, 3032, 2889, 1459, 1336, 2997, 1625, 2320, 2208, 2582, 2279, 3023, 2170, 2992, 2503, 2038, 1782, 2376, 2686, 279, 2366, 1370, 1307, 1424, 2649, 2656, 2708, 3001, 357, 3011, 2979, 2858, 2293, 3109, 273, 2586, 1497, 1318, 2580, 1815, 2390, 2893, 1379, 1780, 1969, 2880, 2814, 2808, 2761, 2650, 2666, 1550, 1602, 2768, 1274, 2927, 2998, 1696, 2984, 2848, 2187, 2154, 1761, 1987, 1856, 1599, 1248, 2449, 2488, 624, 2351, 2640, 2323, 3434, 3539, 3734, 3449, 2830, 3770, 3475, 3207, 3155, 3183, 3691, 3320, 3806, 3404, 3634, 3181, 3694, 3286, 3397, 3463, 3464, 3220, 3294, 3497]. based [3803, 3197, 3800]. Bases [541]. Basic [3651, 726]. basis [3389]. Basu [3708]. Batches [12]. Bates [2256]. Battey [3706, 3952]. Baye [403]. Bayes [1943, 2392, 1778, 2654, 1479, 1947, 3798, 2548, 2269, 953, 1334, 557, 3601, 2209, 1043, 1387, 2272, 1477, 3055, 2858, 1173, 1997, 1053, 1236, 1415, 2200, 1678, 3731, 53, 1925, 1851, 1769, 1216, 721, 1343, 1449, 2181, 3031, 1563, 2489, 3618, 3209, 3730]. Bayesian [2426, 664, 937, 2721, 3874, 3375, 3032, 1392, 3526, 2364, 2060, 2484, 1564, 3734, 3517, 1336, 2416, 3117, 2227, 2326, 994, 3398, 1402, 1742, 2278, 2450, 3726, 3448, 3564, 3139, 1294, 2787, 3780, 2232, 2458, 1123, 2861, 872, 2398, 2170, 2166, 3400, 794, 888, 2523, 2932, 2503, 988, 3317, 1143, 934, 651, 719, 3094, 3194, 2670, 2362, 492, 3137, 3576, 2209, 2178, 1674, 2049, 2573, 2243, 3469, 3784, 836, 1165, 2127, 3745, 3408, 2349, 3794, 3159, 3090, 812, 849, 957, 1055, 2860, 1134, 1271, 956, 3623, 2442, 2565, 3049, 3761, 1966, 3112, 1355, 2591, 3295, 1483, 3496, 3747]. Bayesian [2664, 2342, 1252, 2490, 1201, 2963, 867, 3512, 3638, 1210, 2829, 3285, 3735, 3744, 3908, 1944, 1638, 2384, 2793, 2086, 1132, 2136, 2417, 2084, 913, 2794, 3331, 693, 901, 1986, 3279, 1804, 713, 2715, 3132, 3426, 2886, 2361, 2418, 2419, 2504, 2772, 2957, 3060, 2728, 2510, 1740, 3778, 1980, 2097, 3050, 2401, 1998, 1107, 1417, 2059, 2537, 3330, 2661, 1520, 3939, 3746, 665, 1181, 672, 2497, 2075, 1588, 2367, 2471, 2613, 2304, 1495, 1671, 1751, 3165, 2785, 3876, 3146, 3196, 3800, 3536, 3964, 3565, 3765]. Be [1042, 355, 921, 822, 1497, 55, 62]. Behaviour [1670, 1762, 434, 490, 2115, 813, 74, 643, 2308, 1566, 2084, 916, 3169]. Behrens [207, 438, 535, 977, 354, 1193, 1157]. Belief [1834, 1558, 1171, 3381]. Belongs [1547]. Ben [3867]. Benjamini [3518]. Bennett [2375, 754]. Berkson [2970, 2881]. Bernoulli [3789, 2030, 1422]. Bernstein [2627]. Bertrand [3859]. Besag [1531]. Best [1450, 1691, 2958, 1907]. Beta [1401, 964, 3610, 1303, 899, 163]. Beta-Binomial [1401]. Beta-negative [3610]. Beta-Optimal [964]. Better [1689, 2258, 1422]. betting [3889, 3967, 3890, 3886, 3891, 3892, 3893, 3894, 3895, 3887, 3896, 3897, 3898, 3899, 3900, 3901, 3888, 3902, 3885]. Between [703, 739, 923, 1228, 1000, 1183, 1429, 1289, 1306, 1031, 2112, 1354, 154, 1666, 1109, 332, 813, 622, 959, 1130, 941, 547, 388, 453, 2002, 535, 1980, 169, 161, 545, 665, 203, 692, 359, 2299, 3434, 2722, 3048, 1766, 3161, 2936, 3686, 2917, 2753, 2813]. between- [2917]. Beveridge [326]. beyond [3873, 3496, 3477, 3504]. Bhansali [3951]. Bias [1028, 2617, 2847, 1982, 2262, 2981, 2896, 2189, 3017, 2822, 2403, 2430, 3561, 3903, 3753, 2637, 3850]. Bias-Calibrated [2403]. Bias-Corrected [2430, 3850]. bias-reducing [3903]. Biased [2611, 2480, 2445, 2390, 3779, 1677, 3136, 3532]. Biased-sample [3779]. Bibliography [9]. BIC [3193]. Bigger [1689]. Bilinear [1507]. Bilodeau [3709]. Bimodality [2695]. Binary [1703, 1637, 1056, 2800, 2377, 1839, 418, 455, 2996, 2242, 632, 1884, 3020, 2085, 2559, 3010, 1592, 3314, 3644, 3418, 3642]. Bingham [1535, 2159, 2912]. Binomial [116, 1401, 1706, 884, 1283, 1997, 244, 1516, 161, 1115, 436, 208, 127, 3290, 3610, 3603]. Bioassay [349]. Bioassays [1012]. Biological [22, 98, 134, 51, 157]. Biology [59]. Biometrika [366, 339]. Biophotonic [3044]. Biorthonormal [881]. Bird [1185, 1184]. Birnbaum [1389]. Birth [265, 2724, 1628, 190, 204, 1668, 274, 1007, 726, 3651]. Birth-and-Death [2724, 190, 204]. Birth-Death [1668, 726]. birth-death-move [3651]. Bivariate [2709, 1960, 686, 754, 861, 1249, 1812, 1872, 1306, 1144, 1573, 1006, 2512, 1109, 1062, 2410, 2816, 2204, 2000, 2757, 2137, 483, 828, 856, 918, 969, 990, 2085, 2233, 1568, 1972, 1121, 3053, 1536, 538, 1115, 1788, 3185, 3159, 3422, 3532, 3612, 3238, 1990]. Black [1995, 3600]. Blair [3709]. Blake [3856]. Blind [2448]. Block [585, 1963, 2201, 2069, 1446, 821, 1576, 1339, 1653, 1603, 1581, 1657, 775, 1036, 1287, 1548, 2540, 1720, 2089, 1125, 1018, 1380, 454, 1145, 216, 1721, 241, 1545, 1935, 1793, 3563, 3762, 3435, 3629, 3479, 3220]. Block-Diagonal [1036]. blocked [3617]. Blocking [2331, 3764]. blockmodels [3904]. Blocks [1963, 448, 2206, 2174, 160, 133, 217]. BLUE [1190]. Blur [2560]. Blur-generated [2560]. Bo [3917]. Bodenham [3922]. Book [3621, 102]. Bootstrap [3852, 2446, 2533, 1849, 1898, 2029, 1996, 2295, 1912, 1792, 1824, 2077, 2480, 2445, 2534, 2509, 2882, 1848, 2506, 2265, 2492, 2460, 2136, 2365, 3742, 2351, 3942, 3283, 3495, 3220, 3546]. Bootstrapping [2624, 3085, 2945, 3305, 1978, 3356]. Bootstraps [2384, 1933]. Borne [1276]. Both [2148, 1887, 2503, 3055, 2230]. Botond [3868]. Bound [635, 2656, 3329, 3353]. Boundaries [985, 2596]. Boundary [2168, 2322, 2291, 2189, 2465, 2415]. Bounded [1194, 2151, 3487, 3889, 3967, 3873, 3890, 3886, 3891, 3892, 3893, 3894, 3895, 3887, 3896, 3897, 3898, 3899, 3900, 3901, 3888, 3902, 3885]. Bounded-Sum [1194]. Bounding [1948]. Bounds [805, 1251, 2915, 697, 1598, 1653, 500, 974, 989, 1652, 1264, 1522, 1696, 2031, 808, 1722, 3985]. box [3600, 1364, 1132, 1037]. Brain [2776, 3962, 3843]. Branch [983]. Branching [983, 863, 282, 708, 760, 1059, 685, 829, 864]. Breakdown [2684, 2125]. Breaking [279]. breaks [3486]. Breeding [179]. Bretz [3671]. Brewer [1213]. Bricks [1720]. bridge [3279]. bridges [3351, 3757]. Brooks [1112, 2682]. Brown [3304]. Brownian [3379, 1185]. budget [3963]. Build [1720]. Building [1230, 2687, 3421]. Buja [3717]. Bulgren [1039]. Bulk [283, 449, 329, 362, 514, 574, 690, 866, 461]. Bulk-Service [514]. Busy [474, 727].

C [3890]. C. [744, 3915]. Cai [3829]. Calculating [124, 170, 263, 426]. Calculation [1781, 2092, 170, 2467, 2694, 2121, 1234]. Calculations [2178, 2155, 124]. Calculator [299]. Calibrated [2844, 2403]. Calibrating [2413, 3760]. Calibration [1557, 1789, 888, 2937, 2591, 2616, 3364, 3537, 3220, 3413, 3984]. Can [1042, 822]. Cancer [166, 2666, 2651]. cannot [355]. Canonical [2027, 1633, 2113, 1365, 1499, 3009]. Capacity [514, 758]. Cape [3823]. Capital [912]. capture [3508]. capture-recapture [3508]. Card [107]. Cards [119, 83]. Care [193, 304, 3034]. Carlo [3120, 2062, 3874, 3807, 3503, 2453, 2681, 2830, 2724, 2227, 3398, 2828, 3680, 2891, 3183, 1644, 2029, 2281, 3315, 2036, 2155, 2573, 3150, 3806, 3745, 2584, 2786, 1896, 280, 908, 3582, 3578, 3904, 2719, 2460, 2417, 3417, 3605, 2504, 2957, 2637, 2059, 2289]. Carolina [183]. Carrier [1276]. Carrier-Borne [1276]. Case [676, 705, 2582, 688, 856, 1723, 693, 1640, 1761, 864, 543, 2352, 301, 692, 2612, 2983, 3422, 3341, 2636, 3221, 3316]. Case-Control [1761, 3341]. Cases [94, 1072]. Casting [3014]. Casualties [143]. Catalano [3864, 3974]. Catalogue [63]. Catching [3193]. Categorical [2042, 1624, 1456, 2832, 1382, 2376, 1733, 1994, 2336, 1740, 3580, 3643]. Causal [2915, 3078, 3321, 3585, 2735, 3815, 2610, 3109, 2692, 2645, 3793, 3771, 3376, 2091, 3108, 2734, 3092, 3418, 3587, 3316, 3880, 3952, 3951, 3687, 3786, 3953, 3943, 3945, 3957, 3947, 3458, 3535, 3567, 3419, 3499, 3950, 3697, 3370, 3751, 3521, 3954, 3844, 3682, 3955, 3948, 3580, 3949, 3956, 3642, 3804, 3310, 3728]. causality [3665, 3242, 3626]. causally [3728]. Causation [103]. cause [3187]. Cells [2448]. Cement [11]. Censored [2709, 2440, 1617, 1273, 2139, 1843, 2623, 2999, 2711, 3054, 1854, 1311, 1949, 1708, 2307, 2290, 454, 463, 2327, 494, 1281, 2954, 3786, 3136, 3642]. Censoring [2208, 1618, 2851, 2941, 2762, 2307, 2942, 2451, 2519, 2696, 3298, 3349]. censoring-by-death [3349]. Central [2368, 2431, 909, 1099, 2025, 3416, 3287]. Centre [1762]. Centred [2083, 2693, 1502]. Centroid [620]. Certain [153, 1670, 560, 86, 106, 841, 749, 2146, 357, 407, 124, 1481, 862, 677, 975, 737, 2233, 1518, 702, 564, 1322, 2181, 227, 1474, 143, 1217]. Chai [3672, 3830]. Chain [1830, 2062, 1239, 2681, 2724, 2227, 632, 2584, 2786, 2645, 2417, 1985, 2504, 2957, 2059, 1903, 327, 2289, 3120, 3874, 3807, 3183, 3745, 3582, 2637]. Chains [2736, 434, 1741, 2479, 3278, 3458]. Chair [4, 111]. Chance [303, 2019]. Change [373, 848, 831, 1459, 232, 1696, 3438, 3542, 3314, 3273, 3288, 3387, 3739, 3344, 3911, 3937, 3596, 3439, 3465, 3373, 3627]. Change-Over [848, 831, 1459, 232, 1696]. change-point [3911, 3937]. change-point-like [3542]. Changeover [1694]. Changepoint [2956, 2239, 3684, 3736, 3739, 3814]. Changepoints [2319]. Changes [3091, 3352, 3286]. Changing [378, 379]. Channel [2212, 2442, 412]. Chaohua [3712]. Chaos [2022, 2015, 2160, 2021, 2019, 2020, 2017]. Chaotic [520, 2221, 2615, 2016]. Characteristic [443, 1764, 1472, 2524, 520, 85, 426, 822, 1497, 1800, 1885, 499]. Characteristics [2653]. Characterization [1440, 1532, 1646, 1148, 2009, 753, 673, 1986, 1979, 3520, 1728, 3152, 3318]. Characterizations [2146, 1559]. Characterizing [2142, 2070, 2468]. Characters [361]. Chart [515]. Charts [459, 595, 1042, 288]. Chavez [3858]. Chebyshev [1800, 1885]. Checking [2099, 3377]. checks [3908]. Chemical [1811, 266]. Chen [3831, 3923]. Chi [755, 816, 427, 1382, 1893, 1733, 1735, 400, 1651]. Chi-Square [755, 816, 1893, 400]. Chi-Squared [427, 1382, 1733, 1735, 1651]. Choi [1039]. Choice [2227, 2938, 1031, 841, 2178, 338, 894, 2276, 1914, 240, 854, 453, 833, 1761, 1449, 1169, 1256, 1288, 1968, 2427, 3591, 3285]. Choleski [214]. Cholesky [2193, 2004]. Choosing [1803, 622, 388, 3551]. Choquet [3358]. Chris [3932]. Christian [3707]. Christine [3672, 3830]. Chromatic [942]. Cialfi [3924]. Circle [1762, 483, 1058]. Circuits [1793]. Circular [1492, 2156, 696, 2162, 1141, 1536, 3138]. Clarification [3036]. Clarke [3859]. Class [1697, 752, 1440, 2879, 1459, 1663, 2425, 2512, 1335, 986, 2243, 2649, 2492, 412, 840, 2559, 1696, 886, 1117, 3053, 3066, 499, 1476, 721, 2327, 1142, 3520, 3778, 3743]. Classes [776, 1238, 3268, 3306, 3392]. Classical [2712, 2970, 2503, 1143, 809, 947, 1400, 479, 3148]. Classification [533, 2278, 794, 2829, 891, 2969, 2666, 157, 2173, 3031, 2688, 3386, 3072, 3431, 3189, 3204, 3695, 3909, 3172, 3545]. Classifications [441, 1057, 652, 750, 2779]. Classifier [3083]. Classifiers [2839, 2987, 3798]. Classifying [751]. Clayton [2259, 2378]. Cliff [1434, 1510]. climate [3286]. Clinical [1734, 3792, 3606]. Clipped [814]. Closed [1241, 536, 117, 3985]. Closely [2017]. Closure [2559]. Cloth [52]. Cluster [2821, 3843, 2893, 2086, 2917, 1220, 718, 1035, 3662]. cluster-randomized [3662]. Clustered [2849, 2523, 2945, 3020, 2763]. Clustering [2985, 2961, 3023, 1905, 3527, 2799, 3057, 2893, 2715, 2886, 1613, 3762, 3435, 3629, 3452]. Clustering-Function-Based [2893]. Clusters [2714, 2588, 3406]. Co [43, 80, 22, 2968, 3070]. Co-Operation [43]. co-ordinate [3070]. Co-Regionalized [2968]. Co-Variance [22]. Coal [18].

Cochran [2555]. Coding [107, 1531, 1810]. Coefficient [1533, 1401, 2789, 2683, 874, 1004, 2118, 272, 469, 1017, 3086, 2949, 62, 790, 723, 825, 2925, 2943, 3850]. Coefficients [752, 93, 2986, 236, 513, 2211, 254, 954, 1423, 179, 2650, 1640, 128, 1226, 927, 203, 2497, 3131]. Coherence [1460]. Coherent [2922, 2280]. Cohort [2625, 2612]. Coin [1677]. Collapsed [1893]. Collapsibility [1763, 2028, 2114, 2211, 2875, 1819, 1368]. collections [3242]. Coloured [1842, 2417]. Column [1983, 1465, 1767, 2112, 1211, 1597, 2540, 1399, 3203, 3529]. Columns [968, 2112, 84]. Combination [331, 877, 2537, 2983, 3650, 3963]. Combinations [841, 1906, 3907]. Combinatorial [1071, 139, 3876]. Combined [1698]. Combining [611, 1140, 1038, 1666, 3694, 3104, 2793, 3106, 3302, 3577]. Comment [366, 382, 1039, 1693, 1112, 1606]. Comments [760, 3223, 518, 1213, 671, 1408]. Commercial [124]. Common [1803, 235, 3456, 3498]. Communication [1726, 1727]. Communications [2615]. Communities [2606, 3579]. community [3972, 3973, 3347, 3974, 3981, 3968, 3975, 3976, 3969, 3977, 3970, 3479, 3971, 3979, 3980, 3978]. Compact [1810]. Compare [2792]. Comparing [3085, 2147, 3235, 1477, 1657, 1950, 1654, 2989, 1360, 1607, 2848, 3816]. Comparison [1093, 664, 373, 1388, 2041, 583, 1012, 1653, 1090, 2626, 325, 2266, 1186, 1155, 2002, 2200, 1030, 1518, 409, 467, 597, 1266, 3726]. Comparisons [1109, 255, 2010, 692]. compartmental [3845]. Compatibility [865]. Compatible [2751]. Competing [2725, 3116, 2487, 2942, 3185, 3410]. Competition [1817, 1596, 857]. Competitive [645]. Competitor [1058]. Complementary [235]. complements [3240]. Complete [1658, 2201, 1404, 1517, 1545, 3530]. Completely [2147]. Completeness [1614]. completion [3323, 3619]. Complex [2805, 2184, 807, 2159, 2912, 1154, 230, 2493, 2794, 26, 3350, 3781, 3540, 3132, 3742, 3804, 3565]. Complexity [2061, 1804, 1809, 2661, 3632]. Compliance [2915, 2955, 2950, 3310]. complicated [3631]. Component [1703, 2758, 2756, 602, 2491, 2528, 860, 2154, 2477, 2869, 3573, 3476, 3388, 3372, 3156, 3772]. Components [559, 2754, 1882, 449, 95, 1075, 1219, 3018, 2874, 3083, 3107, 2032, 2126, 2690, 3323, 1679, 2174, 1543, 659, 2361, 2418, 2241, 1919, 2054, 2837, 436, 3523, 3551, 3307, 3257, 2419, 1327]. Composite [2763, 1283, 3105, 2071, 1217, 3206, 3157, 3942]. Compositional [1542, 3158]. Compound [953, 3408, 715, 747]. Compounds [954]. Comprehensive [2684]. Computation [2104, 140, 2060, 2450, 2770, 1909, 2059, 2933, 215, 3526, 3398, 3194, 3576, 3623]. Computational [202, 3762]. Computationally [2889, 3814]. Computations [1838]. Computed [2533, 2988, 2288]. Computer [493, 582, 1726, 2591, 685, 3537, 387, 3289, 3413]. Computers [281]. Computing [342, 1965, 2089]. Concave [1945, 3245, 3135, 3415, 3683]. Concept [803, 853]. Concepts [1221, 324]. Concerning [838, 930, 943, 2008, 734, 825]. Concise [430]. Concomitants [2029]. Concordance [3454, 1354]. Concordance-assisted [3454]. Concurrence [1572]. concurrent [3478]. Condition [2741, 3232, 3798]. Condition-number-regularized [3232]. Conditional [995, 1023, 1054, 2608, 1670, 3180, 1787, 1857, 1378, 1403, 1394, 2281, 650, 2049, 2149, 2763, 2782, 3253, 2009, 3767, 3302, 2002, 2917, 1907, 1911, 1635, 2219, 2031, 2165, 2443, 2292, 2698, 2261, 1903, 2547, 3566, 3726, 3288, 3191, 3378, 3446, 3269, 2633, 3507, 3846, 3842]. Conditionally [2930, 3018, 2452, 3229]. Conditionals [1960, 1990, 2070, 2468]. Conditioned [1175]. Conditioning [1371, 3509]. Conditions [210, 1059, 3517]. Cones [697]. Confidence [687, 716, 473, 511, 2106, 1118, 2951, 1273, 1789, 1947, 2446, 2533, 341, 2605, 295, 3300, 1849, 1898, 1967, 2905, 1124, 584, 616, 1092, 655, 3499, 1824, 2077, 2534, 625, 745, 313, 3055, 1891, 1173, 2265, 2492, 2460, 3411, 2739, 1264, 2495, 693, 901, 1743, 2047, 2365, 375, 1537, 2838, 244, 1359, 2097, 1072, 808, 572, 665, 1234, 1545, 624, 692, 2551, 1384, 1029, 2430, 2351, 3262, 2899, 3350, 3473, 3290, 3447, 3468, 3637, 3958, 3181, 3422, 3810, 3650, 3907, 3329, 3249, 3376, 3537, 3140, 3121, 3220, 2489, 3649, 3209, 1588]. Confidence-Prior [745]. Confidence-Region [687, 716]. Confirming [374]. Conflict [1789, 1666]. Conformal [3659, 2439, 3961]. Conformalized [3777]. Confounded [47]. Confounders [2622, 3566]. Confounding [1176, 1204, 152, 1189, 299, 3943, 3945, 3219, 3574, 3580]. Congestion [524, 555, 682]. conic [3429]. Conjectures [1548, 1793]. Conjugacy [2649]. Conjugate [1601, 2458, 2571, 2449]. Connected [320, 451, 684, 1320]. Connections [1245, 1870, 3048]. connectivity [3962]. Consecutive [2396]. consequences [3576]. Conservative [1325, 2759, 2825, 2824]. Consideration [1004]. Considerations [52, 1280, 1329, 1958]. Considered [11]. Considering [839]. Consistency [1227, 581, 2127, 2328, 1815, 2627, 507, 2759, 2825, 2729, 2613, 3537, 2824, 3664, 764]. Consistent [1825, 2021, 2076, 3007, 1845, 1709, 420, 3051, 3845, 3178, 3160, 3358, 3407]. Constancy [1215]. Constant [347, 466, 381, 2694]. Constants [380, 3278]. Constituent [14]. Constrained [923, 2036, 921, 3336, 3540]. Constraints [2319, 1191, 3366, 3986, 3963, 3308]. Constructed [1528, 1919, 2054]. Constructing [1625, 3194, 1603, 2534, 886, 1117, 1496, 2600, 3290]. Construction [533, 1658, 1983, 2681, 2191, 2027, 3529, 3617, 510, 1414, 1186, 1236, 392, 162, 1935, 1737, 1065, 3140, 3121, 2489]. Contact [1315]. Contagion [950]. Contagious [442]. Containing [1711, 2403]. Contaminated [2970, 1839, 2674]. Contamination [2955, 2480, 2120]. Contaminations [1944]. Contents [2433, 2745, 2975, 3038, 3098, 3211, 3251, 3292, 3334, 3383, 3460, 3514, 3557, 3614, 3775, 3884, 3666]. context [3747]. Contingency [1514, 1580, 937, 1005, 968, 27, 1924, 615, 571, 1763, 756, 2028, 1893, 346, 608, 655, 960, 230, 1201, 990, 1365, 1463, 563, 1769, 2187, 229, 1585, 1819, 1368, 13]. Contingency-Type [990]. Continually [104]. Continuity [1792, 2319]. Continuous [1440, 145, 305, 2145, 2301, 3424, 2724, 1289, 1619, 2357, 2358, 2318, 1194, 2795, 2063, 884, 85, 2463, 412, 2465, 2962, 2263, 2353, 1159, 2505, 1693, 2859, 2852, 1844, 400, 2546, 3130, 3797, 3440, 3521, 3508, 2459, 3261]. Continuum [2301, 2417, 1919, 2107, 2054]. contour [3297]. Contrast [1814, 3389]. contrast-enhanced [3389]. Contrasts [1118, 1124]. Contribution [607, 461, 323, 28, 3921, 3889, 3967, 3857, 3862, 3972, 3670, 3669, 3973, 3708, 3706, 3952, 3951, 3828, 3709, 3922, 3710, 3829, 3823, 3864, 3974, 3672, 3830, 3831, 3923, 3711, 3924, 3859, 3702, 3890, 3854, 3953, 3703, 3673, 3704, 3712, 3860, 3975, 3920, 3925, 3674, 3886, 3861, 3950, 3825, 3827, 3707, 3713, 3821, 3926, 3891, 3863, 3714, 3927, 3892, 3976, 3969, 3928, 3929, 3671, 3715, 3832, 3893, 3855, 3716, 3894, 3977, 3833, 3916, 3919, 3895, 3954, 3675, 3930, 3858, 3856, 3914, 3887, 3896, 3834, 3717, 3931, 3897, 3932, 3835, 3822, 3718, 3836, 3933, 3868, 3719, 3955, 3865, 3866, 3970, 3934]. contribution [3935, 3948, 3720, 3898, 3949, 3837, 3971, 3899, 3956, 3705, 3867, 3721, 3722, 3900, 3826, 3838, 3901, 3824, 3839, 3918, 3979, 3888, 3676, 3869, 3913, 3936, 3980, 3978, 3917, 3723, 3724, 3840, 3915]. contributions [3592]. Control [585, 459, 595, 2767, 2244, 573, 724, 1576, 2485, 3183, 887, 3003, 131, 85, 2692, 2239, 1493, 81, 3417, 288, 1761, 151, 2759, 2825, 102, 1826, 3441, 3264, 3354, 2983, 3622, 3769, 3641, 3184, 3341, 3905, 2636, 3215, 3580, 2824, 3296, 3598, 3655, 3729, 3221, 3509, 3906]. Controlled [99, 3035, 204, 3488, 3466]. Controllers [709]. Controlling [129, 2213, 3941, 3096, 2963, 3566, 3209]. controls [3689, 3945]. Controversy [1301]. Convergence [2323, 1665, 1578, 2072, 2155, 1756, 2203, 2167, 2479, 2957, 3734, 3324]. Convergent [989]. Convex [322, 921, 777, 1066, 3452, 462, 3324, 3958]. Convolution [928, 3683]. Copula [3116, 3079, 2696, 3406, 3481]. Core [3881, 3781]. Cornfield [1348, 671]. Cornish [1384]. Corpuscle [2109]. Correct [1689]. Corrected [1794, 2430, 3904, 3850]. Correcting [1646]. Correction [1228, 3819, 3818, 3852, 3883, 3990, 3967, 1982, 257, 2170, 641, 3570, 1792, 1562, 1111, 2419, 764, 798, 815, 2558, 658, 3593, 3455]. Corrections [1487, 2156, 147, 427, 1579, 2049, 760, 1771, 1488, 1772]. Correlated [1070, 441, 2320, 1706, 2670, 2722, 669, 2476, 2332, 2466, 634, 425, 2045, 3067, 3502, 3152, 3791]. Correlation [2098, 1118, 1434, 1401, 199, 1194, 236, 513, 874, 1004, 272, 774, 1305, 1510, 2593, 959, 469, 1017, 2113, 1630, 62, 168, 2331, 1729, 865, 790, 723, 825, 798, 757, 590, 654, 3208, 3434, 3416, 3324]. Correlational [1342]. Correlations [2201, 2686, 357, 263, 1984, 1365, 409, 3203]. Correlogram [1307, 269]. Correlograms [2908]. Correspondence [38]. Corresponding [264]. Corrigenda [739, 1605, 848, 455, 1255, 682, 1327, 1604, 1116, 616, 797, 849, 929, 1328, 1329, 690, 384, 646, 3992, 1606, 1117, 1607, 456, 457, 1256]. Corrigendum [1023, 1990, 930, 3757, 2743, 1916, 3175, 3176, 3549, 1112, 1898, 1045, 2468, 3784, 868, 1165, 2055, 1888, 2528, 1060, 1835, 2658, 2224, 2808, 2236, 1926, 1193, 2418, 2419, 3140, 1103, 2054, 3239, 3654, 2499]. Corroboration [496, 868]. Cosmology [399]. Cost [2294, 1975]. Costs [595, 1038, 3034]. Cotton [52, 19]. Count [3020, 2519]. Counter [607, 1878]. Counter-Example [1878]. Counterfactual [2934]. counterfactuals [3659]. Counterpart [3000]. Counting [1910, 2305]. Counts [35, 677, 3153, 3498]. Coupling [3734]. Coupling-based [3734]. couplings [3582]. Covariance [2568, 688, 2387, 3093, 1868, 1036, 1046, 1691, 2203, 1795, 814, 2650, 2687, 1951, 1136, 3766, 2728, 2004, 2472, 2497, 2268, 37, 3072, 3268, 3240, 3881, 3476, 3871, 3182, 3574, 3232, 3484, 3442]. Covariance-Regularized [3072]. Covariances [2069]. Covariate [3525, 3004, 2722, 2792, 3648, 3263, 2692, 2814, 2917, 2726, 1592, 2519, 2536, 2710, 2616, 3776, 3792, 3882, 3984, 3609]. covariate-adaptive [3609]. covariate-adjusted [3792]. Covariate-assisted [3525]. Covariate-Dependent [2519]. Covariates [2447, 2971, 2926, 2762, 2969, 2219, 2263, 2947, 2546, 3434, 3180, 3227, 3663, 3509]. Covariogram [3021]. Cover [746]. Coverage [2606, 2607, 2460, 3445, 3353, 3209]. COVID [3759]. COVID-19 [3759]. CovNet [3766]. Cox [2614, 2743, 2625, 2609, 1573, 3469, 3784, 919, 1037, 3239, 2710]. CP [3782]. CP-decomposition [3782]. Cramér [2264, 1075, 1219, 1327, 1188, 1151, 1323, 1381, 973]. Crane [3972, 3973, 3974, 3975, 3976, 3969, 3977, 3970, 3971, 3979, 3980, 3978]. Credibility [1547]. Cribben [3921]. Criteria [2323, 309, 2973, 2315, 2362, 2622, 2626, 3109, 1449, 1408, 3622, 3697, 3655, 3500]. Criterion [1207, 850, 804, 988, 2395, 1100, 2936, 1654, 1655, 1584, 2411, 1288, 2472, 1737, 3423, 3400, 3567, 3272]. Critical [635, 1433, 864, 978]. criticism [3908]. Criticisms [367]. Crop [285, 67]. Cross [1005, 861, 1802, 2083, 1433, 655, 1952, 2055, 2180, 2283, 1169, 1288, 1919, 2054, 2837, 1588, 2779, 3179, 3367, 3648, 3753, 3597, 3522, 1256]. Cross- [2180]. Cross-Classifications [2779]. Cross-Product [655]. Cross-Ratios [1005]. Cross-Section [2837]. Cross-Sectional [1433]. Cross-Spectrum [861]. Cross-Validated [1919, 2054, 1588]. Cross-Validation [1802, 2083, 1952, 2055, 2283, 1288, 3179, 3367, 3753, 3597, 3522]. Cross-Validatory [1169, 1256]. cross-weighted [3648]. Crosses [2662]. Crossing [3013]. Crossings [701, 840]. Crossover [2128]. CRT [3872]. Crystal [684]. Cumulants [591, 300, 1664, 1772]. Cumulative [303, 3185, 515, 823, 1632, 578, 642, 3260, 3410]. Cure [2941, 3553]. Cured [166]. Curnow [1096]. Current [2290, 2451, 3637]. Currents [2448]. Curvature [1421, 2366, 1692, 1633, 2439, 3788]. Curvatures [415]. Curve [2398, 1427, 1707, 2077, 2293, 1379, 497, 578, 1333, 2399, 1684, 436, 371, 3207, 3374, 3521, 3309]. Curves [406, 3013, 1131, 2423, 2524, 94, 2594, 2113, 1951, 2578, 594]. Curvilinear [206]. Customers [330, 477, 727]. Cusum [1235, 1042]. Cut [1253, 987, 3368]. Cycles [1455]. Cyclic [821, 1206, 1340, 445, 775, 1205, 1432, 1480, 1597, 1414, 648]. Cyclical [1455]. Cylinder [2694]. Cylindrical [1357].

D [3621, 644]. D/E [644]. Dai [3894]. Dam [397, 554, 773, 783, 678, 1071, 604, 532, 519]. Dams [386, 706, 1149]. Dan [3829]. Daniela [3924]. Daniels [513]. Dantzig [3048]. DASSO [3048]. Data [1542, 1425, 2709, 2042, 2721, 2440, 310, 3025, 1456, 1637, 373, 1617, 2145, 1056, 2697, 1010, 1273, 1444, 493, 1479, 1475, 2986, 2970, 3004, 2068, 2402, 3024, 2849, 2961, 2377, 63, 3166, 1503, 1839, 2847, 2990, 1706, 3336, 2437, 2895, 1823, 2832, 1286, 1433, 2217, 2512, 2520, 2275, 2670, 2945, 1382, 2487, 2162, 2376, 2906, 2036, 1733, 1651, 2040, 855, 1699, 2303, 2630, 2707, 2843, 2907, 3020, 2987, 3107, 1843, 2763, 2180, 2979, 2820, 2317, 2447, 2870, 3047, 2968, 2999, 2332, 1766, 1355, 2711, 1984, 2690, 717, 2926, 1544, 2850, 2464, 2342, 2113, 2390, 1994, 2090, 2796, 2466, 2829, 1231]. Data [3052, 1854, 1437, 2085, 7, 882, 2336, 1949, 3010, 2665, 1568, 1630, 2521, 1018, 3082, 1708, 2307, 2892, 2290, 1976, 2287, 2962, 2726, 2305, 1951, 1536, 3073, 2263, 2942, 1277, 2822, 2154, 1740, 1987, 2327, 2288, 2505, 1080, 1711, 1754, 2651, 2766, 2857, 2604, 2451, 2519, 590, 654, 2588, 1599, 1649, 1455, 2859, 1281, 256, 2253, 3051, 3062, 2511, 1776, 2299, 2954, 3034, 2572, 3203, 3665, 3486, 3390, 3163, 3205, 3544, 3322, 3071, 3175, 3176, 3770, 3180, 3652, 3962, 3185, 3275, 3389, 3786, 3471, 3155, 3189, 3527, 3355, 3402, 3168, 3794, 3362, 3528, 2983, 3306, 2816, 3881, 3425, 3591, 3138, 3226]. data [3637, 3283, 3477, 3323, 3768, 3181, 3164, 3792, 3813, 3504, 3987, 3508, 3596, 3779, 3767, 3562, 3173, 3136, 3771, 3331, 3532, 3406, 3663, 3626, 3620, 3182, 3766, 3158, 2636, 3294, 3603, 3643, 3386, 3478, 3814, 3221, 3642, 3577, 3801, 3218, 3565, 3743, 3261, 3123, 3805, 2667]. Data- [1010]. data-adaptive [3596, 3814]. Data-Based [1987]. Data-Dependent [2511]. Data-Driven [2217, 3051, 3166]. databases [3282]. Dated [3010]. dating [3486]. David [3860, 3863, 3866, 3898]. Davison [3890]. Dawid [3953]. Dealing [2556]. Dean [3922]. Death [2724, 190, 204, 1668, 274, 726, 3651, 3418, 3349]. Debiased [3877, 3791, 3491, 3755]. Deblurring [2103].

DeCAMFounder [3880]. December

[111, 80]. Decision

[841, 957, 1055, 1783, 1986, 715, 747, 507, 764, 2181, 949, 781, 3448, 3494, 3939]. Decision-Theoretic [1783]. Decisions [292, 953, 243, 957, 1055]. decomposable [3496]. Decomposition [1534, 2193, 670, 1735, 1880, 2004, 2268, 3782, 3186, 3966, 3427, 3808, 1605]. Decompositions [2686]. Deconvolution [2951, 2750, 2096, 3077, 2784, 2448, 3389, 3346]. Deconvolving [2737]. Decreasing [1553, 921]. deep [3911, 3937, 3878, 3846]. Default [3137]. Defective [129, 145, 85]. Deferred [145]. Deficiency [2190, 1966]. defined [3158]. Defining [1279]. definite [3756, 3202]. Definition [873, 919]. Definitions [2684, 1572, 3282]. Deformable [2922]. Deformation [661, 711]. Deformations [2728]. Degree [894, 3904]. Degrees [1534, 1605, 1711]. Delay [1853, 759, 2914]. Delayed [488, 489, 1833, 3448, 3213]. Delays [431, 508]. Delete [2756]. Deletion [2476, 2756, 2205]. Deletions [2924]. Demand [1038, 604]. Demographic [1034]. denoising [3938, 3912]. Dense [2922, 3478]. Densities [567, 3085, 1945, 3112, 1814, 1076, 3564, 3415, 3607]. Density [2440, 2951, 1742, 2322, 2291, 2674, 2096, 2932, 2806, 2247, 2229, 1639, 1699, 1756, 2630, 2655, 421, 1914, 1346, 370, 2189, 701, 814, 2095, 420, 659, 2219, 1987, 1484, 1322, 1474, 1599, 422, 3552, 3160, 3166, 3135, 3355, 3631, 3340, 3591, 3495]. Deoxyribonucleic [2207]. Departments [250]. Departure [1222, 774, 1046]. Departures [309, 2735, 845, 866]. Dependence [2039, 2512, 1733, 2722, 2539, 2615, 2341, 2713, 2875, 1069, 2753, 3061, 1585, 3267, 3348, 3436, 3318, 3392]. Dependency [1130]. Dependent [1425, 1011, 1156, 2339, 2306, 2716, 3116, 1247, 2684, 2036, 2303, 2180, 553, 2926, 2941, 2796, 1383, 444, 534, 2762, 997, 543, 2537, 2519, 2696, 3092, 2511, 519, 3277, 3332, 3235, 3298, 3959, 3228, 3261]. Depends [313]. Depth [2351, 3404, 3181]. depth-based [3404]. Derandomised [3905]. Derivable [139]. Derivation [991, 309, 430, 795, 2476, 1369, 721, 1164]. Derivative [703, 739, 704, 2322, 2379]. Derivatives [2674, 1914, 1322]. Derived [401, 1433, 1764, 161, 217]. Deriving [1617]. Describing [706]. Design [748, 731, 1732, 2255, 848, 831, 1048, 2727, 1887, 2258, 2362, 2363, 1461, 1912, 2234, 1280, 1329, 1101, 859, 935, 939, 1066, 2342, 2936, 1466, 1333, 762, 2188, 620, 1105, 2510, 1226, 1753, 1422, 241, 1531, 1545, 1496, 728, 1114, 3785, 3678, 3466, 3634, 3764, 3151, 3329, 3451, 3554]. design-based [3634]. Designing [2316, 2315, 3790]. Designs [533, 1084, 585, 2254, 1817, 1983, 47, 1963, 2201, 1459, 876, 2532, 1446, 1625, 2441, 821, 1576, 1135, 1339, 252, 415, 1206, 1340, 732, 730, 1127, 1465, 1767, 2112, 2186, 1694, 2027, 378, 379, 1057, 1211, 1404, 1386, 1565, 2147, 974, 1603, 2073, 1447, 1552, 1581, 1657, 1572, 775, 938, 1087, 1205, 1287, 1432, 1480, 1548, 1597, 2540, 1399, 1452, 1720, 460, 793, 2128, 1148, 1414, 1594, 1709, 2174, 963, 1091, 2074, 1829, 1203, 877, 1125, 1923, 1831, 1267, 242, 1018, 1600, 1240, 1696, 549, 750, 1979, 1145, 216, 1677, 1721, 1755, 217, 2813, 3092, 1935, 2347, 3084, 1793, 1282]. Designs [1266, 1737, 1777, 2124, 1065, 69, 3350, 3449, 2600, 3590, 3321, 3264, 3529, 3617, 3152, 3453, 3425, 3633, 3599, 2206, 3171, 3313, 3343, 3588, 3520, 3742]. Desings [1653]. Desk [170]. Destructive [542]. Detect [2300, 3790, 3286]. Detecting [3486, 3091, 1882, 3506, 3352, 2067, 1363, 960, 2396, 2753, 3242, 3549]. Detection [3433, 676, 705, 3387, 2169, 2239, 2603, 2100, 2236, 2202, 2415, 271, 3542, 3347, 3167, 3205, 3684, 3314, 3288, 3695, 3739, 3911, 3937, 3596, 3192, 3808, 3531]. Determinantal [153, 3326]. Determinants [499].

Determination [1825, 2021, 2832, 557, 811, 1396, 822, 1804, 1443, 700]. Determine [2610]. determined [3412]. Determining [94, 823]. Deterministic [2015, 1663, 3523]. Detre [1348]. Development [204]. Developments [197, 1722]. Deviance [1815, 3272]. Deviance-Based [1815]. Deviation [171, 8, 273, 2840, 21, 169]. Deviations [303, 850, 804, 273, 547]. device [3770]. Devices [202]. Diagnostic [1010, 2581, 2476, 2008]. Diagnostics [2042, 1528, 1892, 1916, 1964, 1718, 2657, 2756, 2409, 2713, 2094, 3576, 3197]. Diagonal [1036]. Diagrams [2692]. Diaz [3713]. Diaz-Ordaz [3713]. Dichotomies [1061]. Dichotomous [1819, 2950]. Dichotomy [1615]. Diciccio [1850]. dictionary [3266]. Didelez [3952, 3951, 3953, 3950, 3954, 3955, 3948, 3949, 3956]. Dietz [3673]. Difference [1429, 410, 411, 2708, 332, 813, 2474, 2814, 2998, 535, 484, 927, 3612]. Difference-Based [2708, 2814]. Difference-in-Differences [2998]. Differences [228, 1581, 1766, 263, 273, 2998, 1545, 3409]. Differencing [2240]. Different [441, 994, 1211, 453, 1678, 2508]. Differentiable [1686]. Differential [3234, 3670, 3669, 3672, 3673, 3677, 3674, 3671, 3675, 2287, 2966, 700, 1268, 927, 3676, 3568, 3668, 3280, 3161, 3294]. Differentiation [2665]. Difficult [2687, 2642]. Diffusion [2889, 1663, 983, 636, 1175, 3351, 3757, 3806, 3745]. Diffusions [3022, 3045, 2391, 3938, 3912, 3530]. dilemma [3193]. Dilution [416, 20, 457]. Dimension [345, 2587, 3103, 2142, 2843, 3219, 2354, 2873, 2016, 1959, 2647, 2633, 2943, 3765, 3486, 3122, 3377, 3147, 3686, 3732, 3814, 3484, 3327]. Dimension-free [3765]. Dimensional [1515, 1397, 3028, 1386, 2894, 2786, 2987, 2586, 1521, 1880, 736, 2814, 2246, 3072, 3783, 3429, 3734, 3157, 3205, 3267, 3544, 3575, 3647, 3983, 3488, 3684, 3199, 3314, 3155, 3204, 3179, 3236, 3396, 3402, 3447, 3736, 3416, 3224, 3380, 3589, 3432, 3324, 3660, 3787, 3344, 3476, 3681, 3504, 3596, 3227, 3329, 3388, 3357, 3663, 3438, 3467, 3574, 3330, 3643, 3683, 3360, 3465, 3509, 3577, 3627, 3875, 3262, 3531, 3276, 3765, 3799, 3772, 3654]. Dimensionality [342, 2907]. Dimensions [2984, 3491, 3586, 3395, 3611, 3338, 3730]. Ding [3704]. Dip [2413]. Direct [598, 3458, 2934, 3035, 2467, 2652, 597, 3008, 3173, 3171]. Directed [2144, 2751, 3108, 3122, 3306, 3247, 3965]. Direction [415]. Directional [2853, 1231, 1895, 1841, 3960]. Directionally [380, 381]. Dirichlet [3094, 1210, 3082]. Dirty [1775]. Disaggregation [1931]. Disarray [1310]. discarding [3188]. Discipline [1500]. Disciplines [645]. Disclosure [2673]. discontinuities [3733]. Discontinuous [1560, 1604]. discordance [3636]. Discordancy [1587, 2299]. Discordant [902]. Discounting [830]. Discoveries [3003, 2842]. Discovering [3127]. Discovery [2213, 2767, 2653, 3090, 2652, 2759, 2825, 2944, 3880, 3441, 3127, 3394, 3941, 3266, 3436, 3354, 3622, 3468, 3769, 3641, 3985, 3184, 3905, 2824, 3296, 3598, 3803, 3655, 3729, 3509]. Discrete [487, 2585, 3004, 2895, 356, 1342, 1356, 1200, 603, 2459, 604, 1180, 2391, 2263, 2353, 1159, 590, 654, 1321, 2859, 256, 519, 3245, 3794, 3747]. Discrete-Valued [2585]. Discretely [2889, 2841]. Discriminant [192, 93, 2229, 2260, 2594, 1016, 179, 1738, 3544, 3204, 3325, 3987, 3172]. Discriminating [1228, 1000, 1183, 2936]. Discrimination [923, 1207, 2209, 1871, 2813, 3084, 11, 3470]. Discriminatory [451]. Discussing [786]. Discussion [3921, 3889, 1850, 1904, 2022, 2062, 2256, 2259, 2375, 2378, 2646, 2733, 2720, 2703, 2788, 2783, 3967, 3857, 3862, 3972, 3670, 2667, 3669, 3973, 3952, 3938, 3951, 3828, 3922, 3829, 3823, 3974, 3672, 3830, 3831, 3923, 3924, 3859, 3981, 3890, 3854, 3953, 3703, 3673, 3677, 3957, 3194, 3975, 3870, 3920, 3925, 3674, 3886, 3861, 3950, 3825, 3827, 3926, 3891, 3863, 3927, 3892, 3976, 3969, 3214, 3928, 3929, 3671, 3832, 3893, 3855, 3937, 3894, 3977, 3161, 3833, 3916, 3919, 3895, 3954, 3675, 3930, 3858, 3856, 3914, 3887, 3896, 3834, 3931, 3897, 3932, 3835, 3822, 3836, 3933, 1606, 3868, 3955]. Discussion [3865, 2682, 3841, 3866, 3970, 3934, 3935, 3948, 3898, 3949, 3837, 3971, 3899, 3956, 3867, 3900, 3826, 3838, 3901, 3824, 3839, 3918, 3979, 3888, 3902, 3676, 3869, 3913, 3936, 3980, 3978, 3917, 3840, 3193, 3915, 2856, 2889, 3864, 2847, 3860, 3213, 2361, 3708, 3706, 3709, 3710, 3711, 3702, 3704, 3712, 3707, 3713, 3821, 3714, 3715, 3716, 3717, 3718, 3719, 3720, 3705, 3721, 3722, 3725, 3723, 3724]. Disease [2456, 2968, 2305, 3226, 3987]. Diseased [35]. Disguised [1021]. Disjunction [2115]. Dispersion [703, 739, 704, 450, 1402, 2853, 1799, 2293, 547, 1180, 361, 2188, 1866]. Displaced [547]. Dissociated [1665]. Dissociatedness [1390]. Distance [2567, 2191, 1782, 628, 1969, 1543, 2327, 1577, 3371, 3526, 3470, 3484]. distance-weighted [3470]. Distances [1969, 3472]. Distinguishing [1748, 2305, 949]. Distributed [1818, 470, 1867, 108, 2962, 3780, 3851]. Distribution [1830, 752, 1440, 1319, 3857, 1532, 3862, 626, 2475, 2704, 404, 1675, 846, 284, 441, 560, 1089, 791, 253, 3859, 694, 411, 1870, 427, 3854, 1430, 1160, 2029, 1646, 843, 1560, 1604, 651, 719, 636, 1628, 108, 207, 1362, 3870, 279, 1038, 130, 2139, 1370, 2571, 2423, 45, 2070, 2468, 674, 1275, 1611, 3861, 874, 1004, 884, 1583, 1852, 2801, 2860, 1062, 633, 2404, 1692, 313, 264, 845, 438, 909, 2689, 903, 758, 786, 862, 101, 204, 1535, 2159, 2912, 1295, 1304, 1188, 1100, 1318, 3855, 1017, 2335, 3256, 1505, 2137, 1210, 2177, 842, 483, 735, 2875, 918, 1631, 2493]. Distribution [1880, 1071, 1364, 3858, 604, 3856, 1463, 464, 1925, 1911, 1020, 1193, 1157, 535, 3868, 3865, 538, 2848, 3866, 484, 790, 2838, 864, 961, 169, 161, 180, 1714, 572, 671, 1115, 70, 376, 301, 3867, 525, 590, 2004, 1281, 256, 1738, 899, 906, 2933, 389, 1545, 327, 3869, 127, 3067, 3539, 3864, 3166, 3346, 3860, 3863, 3813, 3961, 3793, 3519, 3169, 3269, 772]. Distribution-Free [279, 903, 2848, 389, 1545, 3256, 3539, 772]. Distribution-Weighted [3067]. Distributional [3087, 1731, 3054, 1365, 3661, 3802]. Distributions [1254, 1697, 1425, 2709, 153, 1245, 1533, 1161, 1960, 1990, 2704, 401, 1392, 1202, 2681, 3117, 2668, 1619, 1908, 2357, 2358, 2318, 892, 1503, 2264, 1144, 2232, 539, 1309, 2166, 362, 1006, 958, 1433, 1560, 1604, 591, 1045, 1014, 1116, 1478, 2146, 397, 630, 1666, 1147, 836, 1165, 1326, 355, 2649, 2221, 2792, 1369, 1906, 1519, 777, 3109, 895, 353, 384, 1200, 1544, 1238, 1113, 941, 2009, 403, 1155, 412, 1039, 990, 1854, 312, 753, 2389, 1220, 1546, 673, 1180, 1360, 1607, 700, 2848, 2751, 1476, 721, 1105, 768, 1050, 1103, 2324, 372, 495, 545, 227, 1693, 1321, 916]. Distributions [2471, 2740, 400, 1671, 2026, 2087, 433, 1664, 1772, 3381, 3207, 3064, 3304, 3853, 3789, 3235, 2459, 3650, 3158]. Divergence [752, 3032]. Divergence-Based [3032]. Diverging [3075, 3067]. Diversification [3010]. Divisible [2232, 1317]. Division [135, 220, 262, 3992]. DNA [2850]. Does [1832, 1611]. Domain [3085, 2325, 3104, 3389, 3328]. domains [3691, 3631, 3987, 3540, 3766]. dominance [3207]. Dong [3670, 3669, 3672, 3673, 3677, 3712, 3674, 3671, 3675, 3676]. door [3567].

Dosage [94]. Dosage-Response [94]. Dose [2338, 695, 94, 3450, 3264, 3151]. dot [3752]. Double [2532, 1867, 1061, 3367, 2534, 2460, 2752, 2185, 2224, 2448, 479, 3945, 3580]. Double-Autoregressive [2752]. Double-Blind [2448]. Double-Bootstrap [2534, 2460]. Double-Exponential [2185, 2224]. Double-Exponentially [1867]. double-negative [3580]. Double-Sampling [2532]. Doubly [3227, 1538, 3309, 2290, 3577, 3984, 3440]. Downdating [2083]. Dr. [111]. Draper [3860]. Drawing [83]. Drawn [130, 180]. Drift [3359]. Driven [2217, 3051, 3166, 3469, 3784, 3495]. Drop [2242]. Drop- [2242]. Dual [1207, 2147]. Duality [2110, 3365]. Duals [1432]. Due [991, 3753]. Dukes [3703, 3708, 3706, 3709, 3710, 3711, 3702, 3704, 3712, 3707, 3713, 3714, 3715, 3716, 3717, 3718, 3719, 3720, 3705, 3721, 3722, 3725, 3723, 3724]. Dummy [1016]. Duration [444]. Duration-Dependent [444]. Durbin [996, 237]. Dutta [3863]. Duty [131]. Dynamic [3168, 2105, 1393, 1461, 3307, 3548, 2702, 2415, 2123, 1712, 2604, 2813, 3906, 3277, 3795, 3389, 2573, 3359, 3191, 3254, 3184, 3435, 3800, 3875, 3481]. Dynamical [3078, 2239, 2962]. Dynamics [2072, 2019, 3393].

EARMA [1439]. Earthquake [967]. easy [3280]. Ecological [1315, 2257, 2983]. Econometric [2639]. Econometrics [1245]. Economic [841, 154, 338, 150]. Economical [468]. Economics [2577]. Edge [3033]. Edgeworth [1412, 2077]. Editor [1939, 2252]. Editorial [2564, 2312, 2372, 3212, 3252]. Editors [3335, 1862, 1901, 1993, 2058, 2135, 2197, 3293, 3042, 3102, 3144, 3177, 2564, 2621, 2680, 2747, 3616, 3385, 3462, 3516, 3559, 2436, 2502, 2812, 2868, 2312, 2372, 2921, 2978, 3212, 3252]. Editors-2015 [3335].

Edwards [1904]. Effect [559, 303, 2338, 2741, 2955, 2609, 413, 1038, 1733, 1896, 1863, 2461, 2245, 2692, 2460, 2542, 3054, 1091, 1466, 886, 1117, 2443, 3363, 3697, 3797, 3370, 3230, 3692, 3309, 3609, 3690, 3276]. Effective [125, 2142, 94]. Effects [2915, 3087, 2706, 2320, 2291, 2773, 2609, 2934, 1674, 3035, 2480, 1237, 2971, 2610, 1633, 2218, 2695, 3029, 2917, 1831, 3014, 2188, 2822, 1216, 2241, 1753, 2383, 1819, 2710, 2950, 3203, 3600, 3491, 3879, 3449, 3647, 3873, 3364, 3786, 3945, 3458, 3567, 3877, 3617, 3499, 3751, 3440, 3659, 3173, 3171, 2412, 3844, 3771, 3682, 3663, 3281, 3451, 3494, 3487, 3482, 3642, 3906]. Efficiencies [986, 2655, 975, 1634, 1179]. Efficiency [1732, 754, 1444, 2800, 1576, 875, 1598, 1653, 513, 1472, 897, 332, 466, 2540, 1652, 1709, 2691, 380, 381, 1203, 1923, 1831, 1240, 1696, 377, 1856, 610, 815, 69, 2289, 3792, 3694, 3287]. Efficiency-Balanced [1240]. Efficiency-Consistent [1709]. Efficient [2889, 743, 2681, 1082, 2727, 3064, 1044, 3258, 3749, 1673, 2023, 3001, 1480, 3105, 2770, 2030, 3741, 3410, 3153, 3104, 3692, 420, 2694, 551, 2219, 1213, 878, 2428, 778, 1755, 1094, 3883, 3364, 3231, 3489, 3697, 3492, 3686, 3227, 3407, 3487, 3814]. Effort [235]. Eigenmodel [3093]. EigenPrism [3432]. eigenvalue [3511]. Eigenvalues [2876]. Either [921, 3191]. Elastic [2864, 3801, 2834, 2865]. Electric [80, 34]. Electricity [92]. Electronic [493, 582, 299, 387]. Elementary [347, 92]. Elements [1796, 1835]. Elicitation [2450]. Eliminating [1265]. Elimination [1552, 1648, 1337]. Elizabeth [3717]. Ellipsoidal [2247]. Elliptically [1818, 1503]. EM-type [2412]. Embedded [1908, 2050]. embedding [3752, 3371]. Embracing [1919, 2054]. Emission [1920]. Emphasis [2704]. Empirical [3350, 1479, 1947, 3770, 2997, 2723, 2548, 953, 1764, 1472, 2023, 2404, 2245, 3055, 3903, 1497, 1318, 2466, 1236, 1678, 2927, 2998, 2161, 1281, 2075, 2444, 3209, 3730, 2899, 3364, 3398, 3506, 3549, 3779, 3132, 3472, 3746, 3478, 3850, 3353, 3565]. Empirical-Likelihood-Based [2997, 2927, 2998]. Emptiness [554, 603, 1071, 604, 1149, 532]. Encountered [92]. encouragement [3451]. End [2973, 2860, 3109, 3370]. endogenous [3644]. enhanced [3389]. enrichment [3588]. Ensemble [3944, 3431]. ensembles [3512]. Entropy [1207, 1995, 1274, 2510, 1248, 3434, 3286]. entropy-based [3286]. Enumeration [47]. enumerator [3940]. envelope [3336, 3480, 3401]. Envelopes [3248, 3365]. environment [3046]. Environmental [1187, 3198]. Epidemic [310, 1241, 1397, 2190, 1596, 1521, 497, 1315, 1954, 1276]. Epidemics [2337, 2422, 2606, 787, 2861, 1596, 2457, 3845]. Epidemiologic [2930]. Epidemiological [2622]. Epochs [603]. Equal [2069, 3000]. Equalising [330]. Equality [674, 1036, 1997, 3519]. Equally [848, 831, 425]. Equation [638, 410, 411, 946, 636, 1175, 2760, 1268, 2302, 3568, 3361, 3698, 3161, 3294]. Equations [731, 1054, 153, 1245, 1515, 2800, 2849, 191, 556, 1953, 3043, 678, 412, 2131, 2687, 2966, 700, 499, 905, 927, 2536, 3051, 2347, 3280, 3129, 3612]. Equi [1279]. Equi- [1279]. Equilibrium [395, 363, 642]. Equipment [61, 946]. Equivalence [2244, 2362, 1218, 1288, 3306]. Equivalences [3007]. Equivalent [3114, 1389, 1509, 2220, 436, 1094, 1888]. Equivariant [1216]. Erdong [3860]. Ergodic [758, 898, 1668]. Eric [3722]. Erik [3900]. Erlangian [474]. Errata [194, 3656]. Erratum [3665, 3785, 3655]. Error [1977, 1964, 1579, 1272, 3096, 2722, 228, 2967, 2948, 1015, 741, 1868, 2473, 1691, 2460, 2307, 1654, 1232, 2401, 2883, 1922, 1582, 1162, 1488, 2536, 2546, 3423, 3346, 3797, 3578, 3787, 3148, 3597, 3215, 3455, 3547]. Errors [1239, 93, 1102, 1292, 441, 1818, 2750, 2970, 2148, 2386, 2549, 2881, 2321, 1996, 1840, 582, 2708, 2350, 1946, 2476, 2325, 2517, 568, 1383, 179, 290, 1020, 947, 1458, 2957, 1277, 2397, 2045, 1372, 2947, 3429, 3277, 3300, 3278, 3453, 3560, 3791, 3228]. Errors-In-Variables [1292, 2881, 2350, 3429, 3300]. Estiamtes [1790]. Estimability [1501]. Estimable [734]. Estimate [1011, 876, 694, 1430, 2096, 623, 2933, 2547, 2382, 3617, 3597]. Estimated [1235, 86, 1429, 2849, 813, 1497, 3495, 939, 1458]. Estimates [1227, 1457, 611, 166, 1965, 875, 3013, 3278, 1529, 1843, 2461, 273, 1845, 1053, 975, 1949, 2650, 420, 901, 551, 549, 231, 1277, 2164, 1980, 2097, 2405, 1754, 1484, 1599, 2779, 3178, 3317, 3230]. Estimating [2639, 1869, 2587, 2800, 3795, 1619, 2148, 2849, 695, 3786, 2955, 8, 2427, 2663, 3361, 356, 1307, 1863, 2860, 200, 2777, 1704, 2473, 1581, 1586, 2354, 2690, 717, 2692, 2850, 1953, 3792, 3043, 370, 1828, 578, 2085, 814, 3607, 3771, 2760, 3630, 2687, 2400, 1951, 3233, 928, 1002, 2016, 2179, 948, 654, 1959, 2588, 1693, 3965, 2536, 2696, 2947, 3051, 3885, 2347, 3206, 3454, 3453, 3903, 3129, 3281, 3628, 3889, 3967, 3890, 3886, 3891, 3892, 3893, 3894, 3895, 3887, 3896, 3897, 3898, 3899, 3900, 3901, 3888, 3902]. Estimation [1595, 481, 1687, 2709, 1177, 1580, 2922, 1636, 1054, 1245, 260, 2440, 559, 1553, 473, 2337, 1308, 171, 2240, 1292, 2889, 743, 2951, 3056, 1207, 1028, 2339, 861, 1249, 2422, 2606, 493, 1475, 3311, 1977, 2970, 2168, 1405, 3024, 1908, 3087, 892, 296, 1296, 2771, 2548, 2269, 2142, 2291, 725, 1247, 1706, 3023, 2076, 95, 567, 3475, 2674, 2166, 1349, 2029, 1086, 517, 480, 612, 983, 1140, 843, 3304, 2789, 651, 719, 2414, 2494, 2520, 3436, 3396, 293, 2806, 649, 582, 3012, 2048, 2139, 2903, 1707, 2423, 1853, 688, 1275, 830, 771, 1307, 3035, 2981, 3332, 346, 2127, 2722]. Estimation [2229, 1643, 1930, 1884, 408, 2584, 2988, 228, 1910, 1494, 1583, 1593, 1673, 1863, 2077, 2630, 625, 669, 421, 1879, 1946, 742, 2756, 3404, 2368, 2431, 1948, 1034, 3093, 3217, 1906, 2803, 2851, 1032, 67, 2517, 2971, 568, 895, 613, 2711, 1413, 1883, 2218, 2580, 2128, 2335, 1201, 1253, 1346, 2872, 2163, 867, 2752, 2840, 2177, 789, 1039, 1800, 1885, 2176, 2033, 2131, 1668, 2189, 3173, 2150, 1398, 1383, 2138, 3104, 219, 274, 2814, 1209, 2518, 1132, 3629, 2095, 2761, 2993, 3113, 1543, 1522, 3372, 1705, 1180, 1925, 1708, 2307, 2257, 3045, 2579, 1865, 1926, 2768, 659, 2687, 2966, 833, 546]. Estimation [2219, 1268, 2578, 3456, 1647, 1476, 439, 463, 1276, 2154, 1740, 1801, 1987, 2288, 3050, 1920, 965, 494, 2401, 2883, 1714, 1770, 2292, 1712, 962, 2823, 2759, 2825, 542, 3115, 1115, 590, 1474, 1814, 1685, 672, 479, 1577, 1649, 2120, 2859, 256, 2911, 1844, 3076, 1810, 2003, 2348, 2740, 3062, 3642, 2403, 1495, 594, 2302, 558, 597, 1108, 1864, 1776, 2527, 1722, 1698, 2525, 2647, 2444, 2628, 2871, 2943, 2954, 2044, 2950, 3879, 3552, 3160, 3170, 3345, 3391, 3255, 3798, 3726, 3379, 3166, 3135, 3268, 3183, 3064, 3945, 3553, 3179, 3240, 3748, 3631, 3736, 3258, 3280, 3528, 3359]. estimation [3306, 3847, 3538, 3468, 3697, 3881, 3797, 3437, 3591, 3578, 3751, 3257, 3440, 3295, 3324, 3560, 3260, 3477, 3787, 3651, 3147, 3681, 3184, 3492, 3989, 3562, 3287, 3341, 3410, 3632, 3946, 3309, 3388, 3397, 3602, 3583, 3357, 3342, 3438, 3451, 3620, 3403, 3574, 3407, 2824, 3598, 3269, 3487, 3465, 3221, 2785, 3232, 3145, 3850, 3801, 3218, 3800, 3371, 3507, 3728, 3505, 3442, 3982, 3193, 2808]. Estimations [993]. Estimator [1533, 1675, 1650, 2817, 996, 2549, 1316, 2699, 3096, 516, 1387, 1699, 2506, 2896, 2634, 1616, 2553, 610, 815, 2931, 3245, 3812, 3355, 3348, 3644, 3455]. Estimators [2323, 995, 1023, 3852, 1670, 1265, 1762, 2041, 1074, 2322, 1451, 2821, 1335, 2024, 1223, 566, 1178, 1639, 1756, 2655, 2948, 3007, 2332, 1450, 2126, 2328, 1815, 2466, 1236, 2388, 1678, 1297, 1654, 2998, 1954, 2161, 1216, 1467, 1805, 1874, 1516, 1322, 1088, 1629, 1791, 1698, 3429, 3227, 3546]. Euclidean [3371]. evaluating [3906]. Evaluation [964, 777, 2693, 499, 186, 2904, 3883, 3489, 3749, 3370, 3959]. Evans [3952, 3951, 3953, 3950, 3954, 3955, 3948, 3949, 3956]. Evening [80]. Event [2971, 1059, 3299, 3310, 3743]. Events [358, 1289, 1306, 1961, 320, 915, 1130, 761, 941, 547, 677, 2550, 2762, 372, 3270]. Every [3000]. Evidence [2280, 496, 868, 2705]. Evolutionary [177, 2903, 1122, 2518, 714]. Evolutive [219]. Evolving [2959]. Ex [811]. Exact [937, 1533, 404, 2889, 1150, 2264, 2795, 2913, 2281, 1565, 246, 2178, 3469, 3806, 1884, 3762, 2506, 1735, 1880, 1297, 2043, 535, 2848, 2443, 1711, 2004, 1615, 904, 558, 1776, 3299, 3747, 3412, 3876, 3784]. Exactly [297]. Example [355, 1878, 744]. Examples [59, 557, 60]. Exceedances [1929]. Excess [2413, 3455]. Excesses [2981]. Exchange [1465, 1452]. exchangeability [3485]. Exchangeable [1431, 2384, 1406, 3579, 1788, 3444]. Exciting [1049, 1538]. Excursion [3297, 3907]. Exhibition [140]. Existence [1843, 1949, 1516, 1754, 1035, 379]. Exogenous [1656]. expanded [3696]. Expansion [2077, 565, 1629, 3628, 3284]. Expansions [2104, 881, 1365, 1743, 1841, 1384, 1664, 1772]. Expectation [3471, 1180, 2830, 3071, 3175, 3176, 3191]. Expectations [1233]. Expected [746, 2605, 3003, 500, 360, 2051, 720, 2536, 3311, 3847]. Expected-Cover [746]. expectiles [3475, 3358]. Experiment [826, 841, 316, 437, 152, 299]. Experimental [2254, 2255, 210, 2662, 352, 252, 2258, 234, 749, 557, 378, 379, 1057, 1404, 460, 2936, 2510, 1226, 68, 1114, 1065, 3634, 3764].

Experimentation [338, 25, 3448].

Experiments

[748, 731, 43, 1228, 1183, 1817, 1534, 1605, 1941, 1347, 1485, 1513, 2484, 876, 1444, 791, 54, 91, 1135, 1176, 1204, 1339, 1964, 583, 964, 1206, 2315, 2727, 1887, 582, 1461, 1389, 496, 868, 1715, 233, 1280, 1329, 1087, 1205, 1597, 119, 717, 860, 859, 935, 939, 2914, 1594, 2603, 1091, 286, 179, 1104, 648, 452, 1189, 139, 454, 423, 620, 241, 281, 1496, 1591, 79, 26, 387, 3818, 3879, 3289, 3510, 3466, 3959, 3230, 3554, 3662, 3811, 3776, 456]. Expert [1858, 2272, 1838, 2537]. explained [3575]. Explanatory [496, 868]. Explication [795]. Explicit [3129, 1157, 978, 3161, 1193]. Explicit-Limit [978]. Exploration [2006, 2243]. explosion [3419]. Exponent [1529]. Exponential [1960, 1990, 1670, 1246, 1371, 1293, 656, 1006, 1646, 843, 2146, 600, 1971, 1370, 397, 2438, 2130, 1062, 845, 763, 1799, 2293, 1439, 1702, 2009, 1350, 2185, 2224, 702, 1925, 1851, 2307, 2037, 3073, 1647, 961, 1445, 2449, 2261, 2044]. Exponential-Type [702]. Exponentiality [2481, 2038, 1764]. Exponentially [879, 470, 540, 1867, 3746]. Exponents [2221, 2017]. Exposure [2257, 2305, 3455]. exposures [3428]. Expressible [2524]. Expression [235, 2829, 2666]. Expression-Based [2666]. Expressions [631, 797, 822, 1042, 1076]. Extended [2144, 1965, 1827, 3237]. Extending [2384, 3406, 3561]. Extensible [3374]. Extension [1237, 776, 2689, 1111, 865, 216]. Extensions [1005, 189, 492, 2653, 2496, 1640, 2302, 3636]. Extensive [973]. extent [3843]. external [3792]. Extinction [2457]. Extrapolation [1127, 2749]. Extremal [915, 2508, 2179, 3774, 3392]. Extreme [1353, 1961, 2119, 2279, 2714, 2581, 2782, 3311, 3475, 3332, 3610, 3404, 3270, 3197]. Extremes [2935, 2736, 2274, 2513, 2000, 2713, 2406, 3337, 3595, 3553, 3159, 3286, 3965].

F.R.S [4]. Facilities [1757]. Facility [727]. Factor [2392, 1456, 1637, 848, 831, 3828, 1818, 3829, 3823, 3830, 3831, 1135, 394, 571, 1598, 2832, 3827, 3821, 1087, 2540, 185, 3832, 1594, 3833, 3834, 3835, 3822, 3836, 138, 186, 3031, 1755, 3826, 3838, 2003, 3824, 3839, 3840, 3848, 3966, 3798, 3617, 3825, 3731, 3841, 3820, 3837, 3773, 3875]. Factorial [1534, 1605, 1732, 826, 285, 47, 2441, 1176, 1204, 1339, 1206, 2186, 1694, 974, 1603, 1205, 1597, 717, 793, 1652, 963, 1091, 2516, 152, 1831, 1267, 1189, 1279, 139, 299, 217, 3818, 3449, 3321, 3529, 3617, 3313, 3811, 3776]. Factorials [886, 1117, 2188]. Factorization [1246, 915]. Factors [1943, 1778, 1732, 307, 1598, 1653, 1334, 1887, 732, 730, 2186, 2209, 2858, 1997, 1652, 2603, 2200, 1923, 1831, 1769, 1449, 1563, 3523, 3498]. Failure [470, 1731, 946, 2512, 2722, 685, 2090, 2942, 2616, 2546]. fair [3810]. False [2213, 2767, 3003, 2653, 3468, 3641, 2842, 2652, 2759, 2825, 3296, 3598, 3729, 3509, 3441, 3127, 3941, 3436, 3354, 3622, 3769, 3985, 3184, 3905, 2824, 3655, 3209]. Familes [1584]. Familial [1984]. Families [1960, 1990, 1246, 1371, 580, 1033, 1907, 702, 2037, 1037, 1445, 3265, 3222]. Family [2357, 2358, 2318, 1503, 600, 2130, 2282, 2384, 2031, 1321, 2449, 2261]. Familywise [3096, 2626]. Fan [3979]. Fasano [3864, 3974]. Fast [1316, 2228, 3810, 1873, 2345, 2412, 3744, 3192, 3397, 2583, 2933, 3008, 3145, 3238, 3874, 3574, 3845]. faster [3193]. Fault [2596]. Fay [2883]. Fearnhead [3932]. Feature [3028, 3732]. features [3542, 3224]. Feedback [445]. Feng [3975]. fertility [3393]. Festing [2259]. Fibrous [1013]. fidelity [3744]. Fiducial [599, 581, 936, 1043, 355, 602, 403, 312, 744, 558, 597]. Field [46, 1347, 1485, 2484, 35, 1964, 1715, 1591, 3438, 3964]. Fields [2808, 2761, 1816, 2583, 2802, 3568, 3424, 3122, 3161]. Fieller [435, 807]. Filippo [3862, 3972]. Fill [2496]. filling [3374]. Film [3030]. Filter [1424, 3441, 3641]. Filtered [1187]. Filtering [925, 2671, 1842, 2795, 3002, 3530, 3130]. Filters [2018, 2535, 2738, 3022]. Financial [3074, 2577, 2837]. Finding [1341, 224, 257, 1549, 2020, 2207, 3063, 3450, 3264]. fine [3608]. finely [3510]. Finite [1687, 330, 1564, 2566, 2754, 2269, 2166, 934, 2295, 333, 771, 772, 228, 2907, 1278, 903, 895, 758, 1046, 2580, 2282, 360, 2086, 1546, 747, 720, 1359, 1740, 180, 1181, 126, 3760, 532, 3627, 3345, 3687, 3258, 3472, 3742, 3982]. Finitely [1460]. First [32, 111, 554, 788, 50, 1071, 737, 1493, 1149, 532]. First-Passage [788, 737]. Firth [2375]. Fisher [937, 366, 207, 3325, 438, 1535, 1295, 3599, 469, 3650, 1631, 1203, 3607, 365, 1193, 1157, 535, 977, 367, 1384, 3172]. Fisherian [1942]. Fisz [2903]. Fit [2481, 596, 267, 2039, 3044, 1128, 1426, 2156, 588, 2723, 2804, 1672, 1144, 1749, 2275, 2686, 1370, 803, 1695, 407, 1369, 1497, 1318, 1016, 1780, 2177, 1962, 1631, 3000, 1950, 2389, 1559, 1543, 2521, 1323, 2187, 2661, 1338, 188, 245, 400, 2071, 2640, 3225, 3541, 3589, 3909, 3904, 3148, 3467]. Fitted [801]. Fitting [1636, 2093, 406, 2398, 2217, 2923, 1627, 2325, 921, 2672, 990, 2336, 2388, 1333, 1085, 997, 987, 214, 1761, 1684, 2235, 3008, 3139, 3336, 2636]. Fixed [660, 2990, 1237, 3220, 808, 2355]. Fixed-Time [660]. Fixed-Width [808]. Flexible [2548, 3299, 3586]. Flexicast [2914]. flipping [3592]. Florian [3835]. Flow [1312, 509, 3623, 3511]. Fluctuations [639, 681, 680, 204, 287]. fluorescence [3359]. Folk [2345]. Folk-Song [2345]. Follow [2796, 3553]. Follow-Up [2796, 3553]. Following [2573, 2551, 3538]. Fong [3857, 3862, 3864, 3859, 3854, 3860, 3861, 3863, 3855, 3858, 3856, 3868, 3865, 3866, 3867, 3869]. Footrule [1310]. Fop [3980]. Forces [1034]. Forecast [2026, 3358].

Forecasting [1269, 1271, 1483, 1417, 2862, 3170]. Forecasts [2937, 3106, 3153]. Foreman [1213]. Forensic [2280]. Forest [60]. forests [3786, 3638, 3640]. Form [2972, 2027, 874, 1193, 1157]. Formal [822, 1042, 187]. Formation [3, 3503]. Former [355]. Forms [401, 560, 1509, 1888, 1304, 300, 796, 3619]. Formula [237]. Formulae [116, 2540, 426, 1142, 698, 208, 624, 327]. Forster [2378]. forward [3208]. Foundations [917, 2081]. Four [846, 848, 831, 732]. Four-Letter [846]. Fourier [1316, 2096, 424, 502, 710, 1148, 2490, 3052]. FP [1227]. FP- [1227]. Fractal [2191, 2142, 2437, 2354, 1959, 2348]. Fraction [129, 145]. Fractional [826, 2441, 1339, 2186, 963, 2200, 2188]. Fractions [938, 2206, 1414]. Frailty [2699, 2461, 2760, 2367, 3201]. Frames [1942]. Framework [2922, 2934, 2666, 3014, 3539, 3381, 3528, 3586, 3324, 3596, 3808]. frameworks [3243]. Frank [3671]. Frederic [3935]. Free [3056, 279, 1370, 2906, 772, 2968, 903, 2833, 2739, 2848, 389, 1545, 3539, 3680, 3748, 3256, 3302, 3692, 3910, 3765]. Free-Knot [2906, 2739]. Freedom [1534, 1605, 1711]. Frequencies [846, 443, 346, 632, 2207, 2508]. Frequency [253, 3085, 1043, 2325, 1173, 721, 1913, 2785]. Frequentist [664, 994, 2170, 3317, 2049, 2247, 2084, 2401, 3137, 3413]. Frobenius [3693, 3796]. Front [1, 17, 24, 31, 40, 49, 58, 66, 72, 78, 88, 97, 110, 121, 141, 155, 173, 195, 221, 247, 259, 270, 291, 319, 336, 350, 383, 417, 617, 647, 662, 683, 691, 765, 779, 799, 817, 834, 851, 869, 889, 910, 931, 951, 966, 980, 998, 1009, 1025, 1040, 1051, 1063, 1078, 1097, 1119, 1138, 1152, 1166, 1182, 1196, 1199, 1229, 1243, 1244, 1270, 1285, 1299, 1314, 1331, 1344, 1361, 1375, 1391, 1410, 1419, 1436, 1454, 1469, 1490, 1512, 1526, 1540, 1556, 1570, 1590, 1609, 1622, 1642]. Front [1662, 1682, 1701, 1725, 1745, 1759, 1774, 1785, 1798, 1808, 1821, 1837, 1847, 1861, 1876, 1890, 1900, 1918, 1928, 1938, 1956, 1974, 1992, 2013, 2035, 2057, 2080, 2102, 2117, 2134, 2158, 2172, 2183, 2196, 2215, 2226, 2238, 2251, 2271, 2285, 2297, 2311, 2330, 2344, 2360, 2371, 2394, 2408, 2421, 2435, 2455, 2470, 2483, 2501, 2515, 2530, 2544, 2563, 2576, 2590, 2602, 2620, 2632, 2644, 2660, 2679, 2701, 2718, 2731, 2746, 2765, 2781, 2798, 2811, 2827, 2836, 2846, 2855, 2867, 2878, 2888, 2901, 2910, 2920, 2929, 2940, 2953, 2965]. Front [2977, 2995, 3006, 3016, 3027, 3040, 3059, 3069, 3081, 3089, 3100, 3111, 3119, 3125, 3134, 3143, 3567]. Full [1246, 773, 783, 1163, 3508, 1208, 1164, 3982, 3182]. Full-model [3982]. Fully [3073]. Function [1457, 192, 93, 322, 1429, 1296, 725, 2582, 2972, 1764, 1472, 2139, 1583, 1863, 537, 1627, 813, 2404, 1032, 1497, 1318, 1016, 939, 2335, 2893, 370, 1800, 1885, 453, 179, 1132, 420, 1473, 1865, 1926, 2768, 659, 1085, 1002, 3050, 526, 525, 1281, 1303, 1738, 2488, 2348, 163, 2547, 462, 3393, 3387, 3761, 3301, 3946, 3727, 3269, 3218]. Function-Noise [1132]. function-valued [3393]. Functional [1245, 3091, 696, 1739, 2706, 2961, 1073, 341, 1086, 517, 3572, 2520, 2906, 2379, 2907, 2874, 2948, 3107, 2594, 2648, 3698, 146, 867, 2880, 3082, 1976, 2886, 2241, 948, 769, 1008, 3156, 2869, 3750, 2916, 3486, 3163, 3770, 3962, 3275, 3473, 3189, 3527, 3584, 3362, 3307, 3217, 3425, 3477, 3323, 3733, 3810, 3987, 3767, 3602, 3766, 3478, 3799, 3274, 3805, 3656]. Functionals [2023, 1686, 880, 3813, 3417]. Functions [1353, 305, 1246, 1308, 2985, 1202, 1742, 2093, 183, 1996, 1560, 1604, 2959, 2807, 674, 520, 1952, 2055, 2656, 2737, 2777, 2473, 2550, 412, 814, 2233, 2989, 2400, 700, 2705, 734, 1801, 1558, 1406, 245, 2471, 2947, 422, 76, 3206, 3180, 3358, 3748, 3332, 3903, 3651, 3793, 3519, 3871, 3182, 3149, 3386].

Further [704, 395, 1485, 251, 580, 1272, 1517, 1004, 783, 301, 654]. Fused [3552, 2818]. fusion [3780, 3650, 3452, 3643]. Future [803, 1766, 3792].

G [1222, 1500, 465, 1853, 1046, 1500]. G/G/1 [1500]. Gain [69, 3792]. Galton [1805, 978]. Gamma [1628, 1840, 163]. Gao [3712]. Gap [2588]. Gaps [1363, 1506, 1562]. Garrotte [2931]. Gauge [287]. Gauging [151]. Gauss [1640]. Gaussian [2671, 3670, 3669, 3025, 2577, 1292, 3297, 861, 2614, 2743, 3163, 3672, 1882, 3687, 1706, 3673, 3677, 3668, 3691, 2503, 3168, 1362, 2209, 3469, 3784, 2127, 3674, 3122, 3019, 2260, 640, 3033, 2354, 3496, 2770, 3477, 3671, 1735, 3641, 3161, 3675, 1398, 840, 737, 360, 3744, 3607, 3540, 2772, 2583, 2802, 3060, 3778, 1752, 3664, 3239, 3676, 3505, 3442]. Gavioli [3920]. Gavioli-Akilagun [3920]. GEID [1227]. GEID- [1227]. Geisser [671]. Gene [2829, 1940, 3046]. General [1697, 752, 80, 705, 2724, 3087, 3078, 474, 1176, 430, 873, 2741, 1964, 1663, 970, 233, 2387, 2476, 1237, 3761, 1163, 574, 690, 1596, 2872, 1614, 919, 412, 891, 1383, 2693, 1954, 1107, 301, 1321, 1164, 728, 1114, 3539, 3381, 2600, 3528, 3313, 3682, 3808]. generalised [3708, 3706, 3709, 3710, 3711, 3702, 3703, 3704, 3712, 3707, 3713, 3714, 3715, 3716, 3735, 3717, 3718, 3719, 3752, 3720, 3705, 3721, 3722, 3701, 3725, 3723, 3724]. Generalization [3032, 536, 872, 807, 809, 859, 1854, 1417, 904, 188]. Generalizations [2349, 351, 1645, 1162, 1488]. Generalized [3282, 2608, 2453, 2800, 2997, 2612, 3366, 2320, 1620, 1982, 2108, 2884, 1857, 1206, 1340, 1086, 2083, 513, 2275, 629, 3019, 1448, 1486, 2350, 2539, 1894, 2803, 2447, 2648, 976, 1205, 1432, 1597, 2574, 3065, 353, 2586, 2593, 1224, 1771, 2218, 2926, 1713, 2298, 1414, 2462, 1210, 3029, 2958, 1616, 2138, 1441, 2559, 1723, 2960, 2521, 1654, 1279, 2966, 561, 1190, 734, 784, 1866, 2292, 2823, 769, 1922, 590, 2004, 2729, 2734, 2367, 1191, 2302, 3008, 2628, 2286, 3785, 3234, 3567, 3380, 3592, 3219, 3589, 3678, 3562, 3129, 3511, 3145, 3500, 384]. Generalizing [2273]. Generally [877]. Generated [2704, 1342, 1356, 2429, 2560, 3464]. Generating [305, 92, 2206, 412, 700, 525, 880]. Generation [2232, 1267]. generative [3679, 3846]. Genetic [2930, 2667, 3608]. Genetics [165, 487, 636, 863, 1175, 768, 271, 2545]. GENIUS [3988]. GENIUS-MAWII [3988]. Genome [2664]. genomics [3146]. Geoelectrical [2721]. Geography [1230]. Geometric [539, 2146, 1491, 2234, 2843, 2167, 433, 3234, 3473, 3152]. Geometry [2409, 1224, 1771, 1464, 3031, 3788]. George [3710]. geostatistical [3284]. GGM [3641]. Gibbs [2062, 3734, 3225, 2458, 2913, 2072, 2061, 3623, 2203, 3342, 2167, 2331, 2059]. Gilbert [3916]. Gill [2733]. Gini [1370]. Giordano [3974]. Giovanni [3864, 3974]. Gittins [2006, 1833, 1438]. Giudici [2682]. Given [443, 2971, 609]. Giving [826]. Glass [80, 89]. Glasses [44]. Global [3401, 2808, 2761, 3944]. Globally [3364]. Go [1611]. Goal [2405]. gold [3488]. Gongjun [3831]. Good [2084, 2862]. Goodman [1615]. Goodness [2481, 596, 267, 2039, 1128, 1426, 2156, 588, 2723, 1672, 1749, 2275, 2686, 1370, 803, 1695, 407, 1369, 3589, 1497, 1318, 1016, 1962, 1631, 1950, 2389, 1559, 1543, 2521, 1323, 3467, 1338, 188, 245, 400, 2640, 3225, 3909, 3904, 3148]. Goodness-of-Fit [2481, 2039, 1128, 2156, 588, 2723, 1672, 1749, 2686, 1370, 803, 1369, 1497, 1631, 1950, 1543, 1338, 245, 400, 2640, 3589, 3467, 3225, 3909, 3904, 3148]. Goseling [3674]. Gradient [3962, 341, 2220, 3694, 3583, 3497, 3628]. gradient-based [3694, 3497]. gradient-sparse [3628]. Gradients [1940, 3533]. Grading [287]. Graduation [556]. Graph [3691, 942, 1447, 2645, 2751, 3178, 3539, 3682, 3752, 3318]. graph-based [3539]. graph-theoretic [3682]. Graphical [2992, 3595, 2300, 3697, 3033, 263, 1838, 1950, 2537, 1903, 3268, 3244, 3402, 3555, 3496, 3485, 3641, 3357]. Graphs [2861, 3047, 3108, 1903, 2791, 3503, 3444, 3563, 3306, 3496, 3574, 3579, 3505]. GRASP [3909]. Green [503]. Greenford [89]. Greenwood [4, 246, 1645]. Gregor [3956]. grids [3374]. Group [791, 491, 304, 3213, 1690, 2982, 3329, 3622, 3453, 3655]. Grouped [1970, 1475, 882, 1708, 1754, 1281, 2871, 3441, 3406]. Grouping [2156, 147, 3732]. Groups [1211, 3527, 3329]. Grover [3593]. growing [3646]. Growth [194, 605, 2293, 178, 240, 736, 1998, 3972, 3973, 3974, 3180, 3981, 3968, 3975, 3976, 3969, 3977, 3970, 3971, 3979, 3980, 3978]. Gu [3861, 3827]. guarantee [3803]. guarantees [3415, 3324]. Guarding [1372]. Guo [3860, 3950, 3724]. Guosheng [3861].

Haar [2903]. Half [2047, 1338, 1791]. Half-Sample [1338, 1791]. Half-Space [2047]. Hamiltonian [3398, 3150]. Hampson [3214]. Han [3825]. Handling [7]. Hanurav [904]. hard [3499]. Harmonic [659]. Harmonics [1148, 1913]. Hastings [2831, 2775]. Having [1211, 1603, 887, 2047]. Hazard [1457, 2440, 2986, 1764, 2768, 2989, 2031, 1801, 3224]. Hazard-Based [2768]. Hazards [2725, 2955, 2609, 2676, 2569, 2711, 2090, 1407, 2760, 2547, 3447, 3367, 3173]. Head [553]. Head-of-the-Line [553]. Health [2552, 3034]. Hearing [787]. Heather [3706, 3952]. Heavy [2981, 577, 972]. Heavy-Tailed [2981]. Heffernan [2783]. Held [4, 80, 89, 111]. Hellinger [2191]. Hennig [3707]. Heppes [3789]. Hernando [3931]. Herriot [2883]. Heterogeneity [755, 816, 120, 284, 2001, 1688, 640, 1552, 2137, 1337, 2094, 3233, 3801].

Heterogeneous [3665, 2606, 356, 2245, 2873, 3439, 3626, 3167, 3786, 3736, 3877, 3663, 3456, 3494, 3772]. Heteroscedastic [2608, 1967, 2569, 2033, 2124, 3034, 3167, 3277, 3337, 3229, 3288, 3227, 3692]. Heteroscedasticity [2424, 2067, 3506, 3549, 3790, 3446, 3507, 3842]. heteroskedastic [3791]. Hidden [2938, 2738, 2504, 2802, 3880, 3535, 3146]. Hieber [3934]. Hien [3896]. Hierarchical [1514, 2458, 1902, 2526, 2105, 2409, 3093, 1894, 2298, 2282, 1216, 2405, 2367, 3564, 3397, 3365, 3343]. High [1830, 3544, 3684, 3199, 147, 1929, 3402, 3736, 3416, 2894, 2786, 2630, 2843, 2987, 131, 2586, 2814, 1741, 2984, 1216, 2246, 3683, 2125, 3360, 3465, 3072, 3276, 3772, 3491, 3783, 3429, 3734, 3267, 3394, 3575, 3647, 3983, 3488, 3314, 3155, 3204, 3236, 3396, 3447, 3380, 3589, 3432, 3586, 3324, 3660, 3787, 3344, 3476, 3681, 3504, 3596, 3227, 3329, 3388, 3357, 3663, 3438, 3467, 3574, 3395, 3643, 3611, 3814, 3509, 3577, 3484, 3627, 3875, 3262, 3338, 3531, 3730, 3765, 3799, 3654]. High-dimensional [3684, 3736, 3683, 3772, 3783, 3734, 3647, 3983, 3660, 3787, 3681, 3663, 3643, 3627, 3875, 3765, 3799, 3654]. High-Order [1741]. Higher [637, 1463, 2037, 623]. Higher-Order [637, 1463]. Highest [2247]. Highly [2726, 3067, 3688]. Highway [1187]. Hilbert [574, 690, 3767, 3274]. Hines [3713]. Hinkley [1850]. Historical [1373]. History [3004, 3646]. Histosplines [1262]. Hochberg [3518]. Hodges [3716]. Holding [1038]. Holdout [3908]. Hollerith [61]. Holmes [3857, 3862, 3864, 3859, 3854, 3860, 3861, 3863, 3855, 3858, 3856, 3868, 3865, 3866, 3867, 3869]. Homogeneity [528, 86, 2566, 2320, 501, 1068, 522, 3805]. Homogeneous [1778, 265, 441, 2190, 829, 1998, 3149]. homoscedastic [3221]. Hong [3926]. horizon [3727]. Horse [1482]. Hospital [250]. Hot [827]. Hotelling [629]. Howard [3891]. Huk [3863]. HulC [3958]. hulls [3958]. Human [1230, 68, 271]. Hunt [3714]. Hunted [1254]. Hunting [1940]. Hüsler [3304]. Hybrid [3181, 3082, 3507]. Hyperboloid [1880]. Hypercube [2025]. Hypergeometric [1628, 353, 384, 1321]. hypersphere [3158]. Hyperspherical [697, 2853]. Hypoelliptic [3530, 3045]. Hypotheses [192, 2879, 838, 896, 930, 374, 580, 1823, 3096, 2115, 302, 1283, 1163, 2850, 2282, 1907, 2892, 1584, 2356, 1903, 504, 1217, 3441, 3265, 3449, 3873, 3584, 3942, 3222]. Hypothesis [1687, 148, 802, 3032, 3347, 940, 1129, 1710, 3112, 627, 1813, 1669, 668, 1164, 2071, 2748, 3586].

Ian [3714]. ideas [3403]. Idempotent [668]. Identifiability [2725, 3664]. identifiable [3794]. Identification [1319, 2662, 3205, 2280, 3945, 2778, 1796, 1865, 1926, 2890, 1844, 3370, 3257, 3693, 3796, 3781, 3376, 3630, 3627, 1835]. Identifying [2961, 2934, 2040, 3788, 3813]. Identity [2001, 689]. Idleness [645]. Ignorability [2334, 3751]. Ignorable [2068, 2771, 2376, 2447, 2822, 3813]. Ignorance [992]. Igor [3862, 3972, 3890]. II [489, 1291, 705, 1206, 1219, 1327, 379, 772, 1055, 1356, 1596, 220, 262, 62, 589, 186]. III [1414, 1208]. Illustrated [74, 11]. Ilya [3720]. Image [2592, 1779, 2417, 2522, 1985, 2103, 3878, 3907, 3808]. Images [2228, 1884, 3165]. Imaging [3044, 2776, 3389, 3359]. Immigrant [1254]. Immigration [708, 760, 864, 978]. Immune [2190, 1966]. impact [3602, 3636]. Impacts [1187]. Imperfect [2705]. implementations [2412]. Implications [1742, 1628]. Implicit [1644]. Importance [2534, 2137, 3536]. Impossibility [2143]. Imprecise [965]. Improper [1425, 1326, 1372, 3761]. Improve [1699, 2365, 698, 3618]. Improved [1620, 2108, 2895, 2821, 2395, 2126, 3260, 2141, 2558, 3523, 3368, 3220]. Improvement [1873, 289]. Improving [1636, 3024, 106, 2607, 2380, 3795, 3872]. Imputation [2844, 2898, 3011, 2896, 2381, 3504]. imputation-regularized [3504]. Inaccuracy [521]. Inadmissibility [2181]. Inaugural [4]. Inbreeding [284]. Incidence [2456, 3185, 3410]. Incidental [2222]. Including [552]. Inclusion [1680]. Incompatibility [1173]. Incomplete [585, 2337, 821, 1576, 2847, 524, 1286, 555, 682, 775, 1287, 448, 1544, 2336, 1125, 361, 216, 1599, 1649, 163, 2572, 3528, 3986]. Incomplete-Data [2847, 2572]. Incorporating [2212, 2997, 1740, 2241]. Incorrectly [233]. increased [3744]. Increasing [3639]. incremental [3562]. Increments [2354]. Independence [2052, 2069, 1194, 1378, 1403, 2992, 2996, 3028, 631, 797, 2404, 14, 2009, 796, 1022, 1903, 3566, 3378, 3767, 3463, 3484]. Independent [125, 75, 474, 1222, 811, 2684, 3224, 325, 547, 1360, 1607, 484]. Index [2369, 2432, 2498, 2561, 2618, 2677, 2742, 2809, 2863, 2918, 2974, 3037, 3097, 3141, 3174, 3210, 3250, 3291, 3333, 3382, 3459, 3513, 3556, 3613, 1430, 2981, 326, 2553, 2187, 2179, 1438, 2899, 3366, 3224, 3377, 3301, 3227, 3200, 3850, 3208]. Indexed [2444, 3331]. Indices [2006, 1393, 1461, 2595]. Indifference [913]. Indirect [2934, 2801, 2777, 1920, 3458, 3567, 3844]. Individual [415, 3873, 3659, 3343]. individualized [3647, 3454, 3692, 3494, 3963]. Induce [2755]. Induced [2791]. Induction [311]. Industrial [112, 33, 2396, 127, 4, 3, 80, 50]. Industry [18, 10, 5]. Inequalities [322, 554, 971, 351, 1630]. inequality [3411, 3308]. Infection [2305, 1954, 3759].

Infections [726]. Infectious [2456, 2305, 3226]. Inference [3203, 2212, 175, 1260, 2732, 2879, 1392, 3828, 1842, 2841, 1747, 3829, 3823, 2326, 3830, 2930, 3590, 1402, 3831, 3087, 2849, 2938, 2314, 1787, 3646, 1401, 2232, 1857, 2895, 1123, 1942, 2861, 872, 1644, 2526, 2414, 3094, 2738, 2956, 2714, 492, 2376, 667, 2959, 1043, 943, 2675, 2656, 2708, 2801, 3020, 3077, 3001, 3827, 2777, 2557, 2597, 1966, 521, 1154, 2586, 3832, 2664, 2490, 2390, 3147, 2462, 261, 1385, 3833, 3029, 3650, 2793, 2086, 2136, 3834, 1415, 3835, 3822, 1550, 1602, 3836, 713, 2927, 3663, 551, 2504, 2772, 1325, 498, 3060, 2728, 2165, 507, 764, 3472, 2545, 2429, 1406, 2698, 3826]. Inference [3108, 3838, 2734, 2471, 1810, 3092, 3824, 3839, 2511, 2261, 1301, 2946, 3373, 3228, 3308, 3840, 2558, 3972, 3874, 3491, 3973, 3299, 3708, 3706, 3783, 3265, 3709, 3710, 3163, 3525, 3575, 3647, 3974, 3873, 3364, 3770, 3711, 3284, 3687, 3981, 3968, 3702, 3321, 3207, 3246, 3703, 3363, 3704, 3530, 3712, 3699, 3975, 3137, 3273, 3567, 3340, 3877, 2573, 3843, 3469, 3784, 3806, 3745, 3159, 3815, 3825, 3847, 3960, 3368, 3707, 3713, 3714, 3432, 3976, 3969, 3634, 3521, 3496, 3560, 3660, 3747, 3715, 3326, 3716, 3768, 3659, 3872, 3813, 3977, 3533, 3793, 3508, 3562, 3302, 3735, 3717, 3771, 3439, 3376, 3718, 3426, 3719]. inference [3841, 3820, 3970, 3464, 3318, 3580, 3727, 3720, 3837, 3971, 3705, 3746, 3721, 3722, 3446, 3701, 3725, 3742, 3814, 3979, 3845, 3980, 3876, 3577, 3978, 3196, 3804, 3310, 3627, 3910, 3565, 3723, 3690, 3369, 3131, 3261, 3724, 3799, 3821]. Inferences [1984, 825, 2091, 2253, 3619]. Inferential [3679, 3302]. Infinite [1825, 397, 2840, 603, 1071, 840, 526, 519, 3727]. infinite-horizon [3727]. Infinitely [2232, 1317]. Inflation [2472]. Influence [1235, 1760, 2582, 1996, 1427, 2366, 2657, 3083, 1783, 2205, 2151, 2233, 2236, 2202, 2439, 2129, 2192, 2572, 3339, 2499]. Influential [2236, 2202, 3531]. Information [1778, 1580, 2, 16, 29, 41, 64, 77, 96, 109, 122, 142, 149, 156, 164, 174, 182, 196, 209, 222, 238, 258, 278, 308, 334, 368, 398, 432, 458, 471, 485, 505, 527, 618, 663, 712, 766, 818, 870, 932, 981, 1026, 1064, 1120, 1167, 1198, 1258, 1300, 1376, 1420, 1470, 1527, 1571, 1623, 1683, 1746, 1786, 1822, 1859, 1897, 1936, 1989, 2053, 2132, 2194, 2249, 2309, 3041, 3101, 3384, 3399, 3414, 3430, 3443, 3461, 3474, 3483, 3490, 3501, 3550, 3515, 3524, 3534, 3543, 3558, 3571, 3581, 3594, 3604, 3615]. Information [3625, 3635, 3645, 3657, 3667, 3685, 3700, 3737, 3758, 211, 2402, 2997, 482, 2262, 1710, 2038, 1767, 2806, 1044, 667, 496, 795, 868, 3020, 832, 929, 2395, 2610, 1883, 1346, 2163, 1549, 437, 3104, 877, 2467, 2051, 1769, 1279, 2031, 1584, 1213, 2411, 2537, 1563, 3400, 3599, 3493, 3769, 3302, 3606, 3272, 3349, 3500]. Informative [3117, 2242, 3781]. Informativeness [1074]. Infra [2379]. Infra-Red [2379]. Inhomogeneous [1905, 2988, 3076, 2776]. Initial [1059, 2192]. innovation [3064]. Input [887, 773, 783, 880]. Input-Output [880]. Inputs [554, 1071, 1149, 519]. Insecticide [46]. Insight [1433]. Inspection [129, 292, 405, 132, 468, 781, 318]. Instability [1660]. Instantaneous [2525]. Institute [183]. Instrument [3644]. Instrumental [2967, 439, 3795, 3458, 3521, 3989, 3630, 3487, 3642, 3649]. Instruments [114, 3499, 3815, 3521, 3989, 3988]. insufficient [3553]. Insulin [98]. insurgent [3771]. integer [3064, 3610]. integer-valued [3064, 3610]. Integers [443, 1223]. Integral [638, 946, 483, 565, 425]. Integrals [1869, 2524, 2246, 624, 1975]. Integrated [2240, 2322, 2674, 3060]. Integration [2719, 2388, 3417]. Integrative [3961, 3545, 3801]. integro [3612]. integro-difference [3612]. Intensities [1249, 1910]. Intensity [1869, 725, 1538, 1801, 2776, 3651, 3149].

Intentionally [2445]. Inter [1222]. Inter-Departure [1222]. Interaction [1514, 1172, 2093, 1902, 1124, 424, 502, 2140, 229, 3686, 3046, 3247, 3187]. Interactions [27, 571, 1340, 1598, 2112, 608, 2603, 1463, 563, 3617, 3046]. interactive [3493]. Interchange [1465, 1452]. Interchanged [160]. Interclass [1984, 549]. Interdependence [133, 3242]. Interest [73, 2352, 3147]. Interference [536, 364, 396, 440, 3682]. Intergenotypic [1817]. Interim [1891]. Interlaboratory [3115]. intermediate [3943]. Interpolated [2106]. Interpolation [42, 2534, 2342, 128]. Interpretable [3987]. Interpretation [1425, 3078, 1491, 1252, 159, 2388, 1678, 713, 229, 215, 3844, 3752, 3773]. Interpretations [2645]. Interpreting [2705]. Interrupted [552, 1566]. Intersections [759]. Interval [1177, 2212, 1934, 2771, 411, 293, 2139, 625, 669, 1843, 2851, 1032, 1891, 2711, 360, 414, 135, 220, 262, 3992, 1949, 901, 1547, 244, 1980, 2097, 993, 3140, 3121, 3649]. Interval-Censored [2139, 1843, 2711, 1949]. Intervals [746, 2635, 344, 343, 212, 251, 2106, 1118, 1273, 1947, 2446, 1289, 1306, 295, 1849, 1898, 1967, 988, 1124, 246, 1824, 2534, 2607, 313, 3055, 1173, 1130, 2265, 2492, 2460, 547, 2739, 576, 646, 2495, 2084, 2365, 375, 1537, 1359, 2097, 1234, 1588, 2551, 3350, 3290, 3447, 3499, 3637, 3411, 3329, 3249, 3376, 3132, 3262, 3209]. intervention [3451, 3316]. interventional [3306, 3844]. interventions [3585, 3943]. Interviewer [1703]. Intraclass [1118, 1401, 1984]. intractable [3735]. Intrinsic [3540, 798, 3320, 757]. Introduction [2014]. invalid [3499, 3815, 3989, 3988]. Invariance [1254, 2735, 1366, 844, 2516, 3289]. Invariant [1812, 1872, 2221, 2152, 2153, 1519, 1691, 1080, 3070, 3376]. Inventory [512, 641, 1038, 550]. Inverse [3852, 3044, 3056, 2587, 2191, 42, 2787, 2083, 1362, 976, 1735, 950, 2687, 561, 734, 2004, 2103, 2788, 3268, 3779, 3546]. Inverses [1224, 1771, 1190]. Inversion [2721, 2446, 191, 213, 2786, 628, 1350]. Inversions [236]. Invertibility [1448, 1486, 1486]. Invertible [2240]. Investigate [1484]. Investigation [801, 103, 61, 1253, 315, 143]. Investing [3003, 3282]. Investment [912]. Investments [2103]. Invitation [2819]. Involving [1403, 830, 984, 1060, 315, 13]. Ion [2212, 2442]. Iraq [3771]. Iron [131]. Irregular [826, 3047, 2796, 3986]. Irregularly [2145, 2594, 3052, 3390]. Irreproducible [266]. Isadora [3857]. Isolation [2665]. Isotonic [3661, 3587]. Isotropic [3021]. Issue [3384, 3399, 3414, 3430, 3443, 3461, 3474, 3483, 3490, 3501, 3550, 3515, 3524, 3534, 3543, 3558, 3571, 3581, 3594, 3604, 3615, 3625, 3635, 3645, 3657, 3667, 3685, 3700, 3737, 3758]. Items [14]. Iterated [2319, 2265, 2710]. Iteration [2509]. Iterative [1978, 3628]. Iteratively [1643, 3966]. IV [1084, 963]. Ivor [3921].

J [111, 3828, 1112, 3718, 3705, 3722, 3838]. J. [3674]. Jackknife [2817, 2898, 1498, 1996, 1462]. Jackknife-After-Bootstrap [1996]. Jackknifing [1953, 1686, 1874]. Jackson [3927]. James [3927, 3955, 3980]. Jason [3980]. Jeffreys [3032]. Jenkins [1364, 1132]. Jennison [3214]. Jiajin [3975]. Jiajing [3926]. Jiang [3892, 3976]. Jiaqi [3861]. Jiayi [3894]. Jiti [3712]. Jiwei [3723]. Jobs [548]. Johannes [3928, 3893, 3934]. John [821, 3929, 3991]. Joint [3206, 3299, 3652, 1478, 2070, 2468, 1238, 2341, 3357, 3053, 3073, 376, 3455, 2791, 3268, 3813, 3463, 3303]. Jointly [3306]. Jordan [628]. Jorge [3675, 3930]. Joshua [3823]. Journal [3654, 382]. Judged [85]. Judges [1354]. Judgment [234]. Jump [2841, 2681, 2724, 537, 2504, 3806]. jump-diffusion [3806]. Jungbluth [3928]. Junhui [3829]. Jupp [2733].

Kaizheng [3824]. Kalman [2535, 1424]. Kanti [3919]. Kantorovich [3198]. Kaplan [1958]. Karhunen [3284]. Karl [3841]. Karla [3713]. Kaur [3593]. Keeping [74]. Keming [3892]. Kendall [3412]. Kent [3929]. Kernel [1999, 2208, 2291, 2438, 1639, 2655, 1946, 2180, 2204, 2189, 3463, 2365, 1987, 1484, 1855, 1599, 2916, 3816, 3696, 3371, 3274]. Kernel-Based [1599, 3463]. kernel-expanded [3696]. Kernels [1707, 2465, 3607, 3612]. Kind [824]. Kinetics [1811]. Kingman [1606]. knockoff [3641]. Knockoffs [3941, 3488, 3985, 3905]. Knot [2906, 2739]. Knots [1134, 1262]. Knowledge [2184]. Known [2390, 1944]. Kolmogorov [1158, 1291, 1487, 1263, 961, 973]. Kolmogorov-Type [961]. Kolyan [3868]. Kong [2720]. Konstantin [3837]. Korbinian [3837]. Kriging [2990]. Kronecker [3808]. Krusche [3671]. Kruskal [1615]. Kuldeep [3715, 3832]. Kullback [2038]. Kumar [3715, 3832]. Kurtosis [1818, 657].

L [3717, 3705]. Laan [3718]. Label [2556]. labelled [3961]. Labelling [2206]. Labels [1074]. Laboratories [80]. Laboratory [46, 2258]. Lack [3044, 2804, 3541, 1819]. Lack-of-Fit [2804, 3541]. Lagged [1126]. Lagrange [1442, 1504, 1582]. Lagrange-Multiplier [1442, 1504]. Lagrangian [128]. Lags [1126]. Lamp [34]. landscape [3533]. Lane [431, 508]. Langevin [3150, 2391]. Languages [3010]. Laplace [3389, 3073, 3653, 3060, 2246]. Large [570, 43, 3025, 2255, 2990, 3240, 1307, 1363, 71, 984, 1060, 101, 547, 2603, 1958, 1626, 551, 1716, 2766, 3061, 188, 2071, 181, 3525, 3471, 3699, 3368, 3237, 3182, 3294, 3296, 3851, 3545, 3327]. Large- [1307]. Large-Sample [570, 1363, 71, 181]. Large-Scale [43, 3061, 3525, 3368, 3237, 3296, 3851, 3545]. Largest [472, 993]. Larsson [3893]. Lasso [3048, 2982, 3342, 2267, 2818, 2925, 3268, 3555, 3344, 3411, 3154, 3188, 3690]. lasso-type [3188]. Latent [2879, 2523, 1351, 3019, 2803, 2552, 3060, 2353, 2505, 2449, 3972, 3973, 3848, 3297, 3071, 3175, 3176, 3974, 3981, 3968, 3402, 3975, 3794, 3976, 3969, 3816, 3977, 3788, 3129, 3970, 3574, 3971, 3773, 3979, 3980, 3978, 3743]. Latin [1658, 415, 1418, 1404, 1517, 205, 2025, 242, 1600, 393, 217, 84]. Latter [355]. Lattice [1172, 1292, 136, 859, 935, 939, 290, 2694, 2962, 2288, 100, 1531, 1777, 3312, 3123]. Lattices [684, 2175]. Lauritzen [1904, 2646]. Law [2357, 2358, 2318, 3391]. Laws [1519]. Layouts [1823]. leaders [3195]. lean [3708, 3706, 3709, 3710, 3711, 3702, 3703, 3704, 3712, 3707, 3713, 3714, 3715, 3716, 3717, 3718, 3719, 3720, 3705, 3721, 3722, 3701, 3725, 3723, 3724]. Learned [3900]. Learned-Miller [3900]. Learning [3195, 3002, 2837, 2253, 3921, 3600, 3938, 3922, 3923, 3924, 3774, 3454, 3601, 3535, 3920, 3925, 3926, 3927, 3909, 3928, 3929, 3681, 3911, 3937, 3650, 3916, 3919, 3930, 3692, 3640, 3914, 3931, 3932, 3933, 3630, 3934, 3935, 3778, 3732, 3494, 3727, 3918, 3913, 3851, 3936, 3917, 3846, 3915]. Least [801, 601, 1762, 2541, 3103, 1451, 875, 850, 804, 1086, 2083, 1352, 1643, 1715, 2923, 897, 917, 921, 556, 314, 2593, 2672, 2840, 2380, 1616, 1638, 3946, 2465, 2553, 1708, 202, 947, 231, 1277, 214, 315, 1874, 769, 1919, 2054, 3067, 3234, 3812, 3445, 3248, 3360, 3654]. Least-Square [231]. Least-Squares [875, 897, 921, 556, 1616, 1638, 947, 769, 3654]. Leave [1892, 3597, 1916]. Leave- [1892, 1916]. leave-one-out [3597]. Lederer [3928]. Lee [3925]. Leeds [90]. Left [2851].

Lehmann [1293, 1068]. Leibler [2038]. Leng [3654]. Length [758, 515, 823, 360, 2003, 3375]. Lengths [576, 646]. Leser [556]. Lethal [695]. Letter [846]. Level [2924, 2997, 2441, 2832, 1896, 3077, 2143]. Levels [831, 2186, 609, 128, 848]. Leveraging [3650, 3905]. Levinson [2089]. Levy [1553, 3190]. Li [3894, 3977]. Lieshout [3674, 3915]. Life [1842, 3004, 1294, 470, 1067, 2122, 1210, 2176, 1311, 2154]. Life-Tables [1067]. Life-Times [470]. Lifetime [2154, 526, 2816]. Lifting [3011, 3807]. Lifting-Scheme-Based [3011]. Light [272, 3396]. Lijoi [3862, 3972]. like [3542]. Likelihood [481, 716, 995, 1023, 1533, 1970, 1670, 1011, 1246, 1675, 2144, 2879, 2240, 2889, 1156, 615, 166, 2339, 1475, 1747, 2402, 1405, 1932, 1988, 2930, 2997, 2566, 2599, 2723, 1794, 2198, 2706, 2804, 1965, 1073, 2208, 3114, 1620, 2092, 2758, 1247, 1827, 2513, 2895, 1942, 1286, 2078, 1086, 2425, 1320, 2414, 2494, 2663, 2295, 3012, 2036, 2155, 1223, 2049, 2149, 1275, 2592, 1981, 2229, 1643, 1930, 2584, 2596, 2656, 2655, 2860, 1843, 2763, 844, 2032, 2273, 1894, 2774, 2851, 1032, 94, 1878, 984, 1060, 613, 2711, 1100, 3036, 2390, 1385, 2002, 1949, 2010, 2136, 1676, 1297, 1522, 1550]. Likelihood [1602, 2927, 2998, 2726, 546, 2578, 2705, 1476, 2164, 1801, 623, 2638, 1516, 1754, 543, 1714, 1770, 2292, 948, 1072, 1008, 2352, 1629, 2121, 1685, 2859, 2103, 2094, 3197, 2488, 2933, 2355, 1975, 1864, 1776, 2954, 2640, 2899, 2558, 3245, 3299, 3812, 3350, 3157, 3770, 3398, 3135, 3320, 3236, 3806, 3319, 3755, 3237, 3508, 3779, 3132, 3636, 3611, 3746, 3478, 3145, 3850, 3373, 3800, 3353, 3565, 3442]. Likelihood-Based [2889, 2656, 2390, 3197, 3770, 3806, 3800]. Likelihood-Ratio [716]. Likelihoods [2453, 2605, 2170, 2262, 2350, 2065, 1921, 2495, 545, 2766, 1940, 624, 3735, 3845].

Limit [1830, 1478, 1099, 2025, 978, 1455, 3471, 3416, 3762, 3342]. Limitation [2076]. limitations [3510]. Limited [89, 2402, 3165]. Limiting [401, 443, 1742, 362]. Limits [1425, 511, 2831, 294, 341, 584, 616, 1092, 655, 602, 823, 1743, 665, 558, 1384]. Linda [3829]. Lindisfarne [402]. Lindley [492]. Line [523, 2738, 2956, 552, 553, 2971, 464, 2890, 227, 3071, 3175, 3176]. Linear [1595, 746, 731, 1245, 129, 801, 2099, 2098, 2879, 914, 322, 473, 3429, 896, 1102, 2453, 876, 2313, 1202, 2018, 1587, 2986, 2015, 2068, 1405, 588, 2997, 2050, 2148, 3275, 1451, 2771, 1965, 2269, 75, 1073, 1620, 1982, 2108, 2758, 341, 1964, 2884, 850, 804, 1857, 1082, 1086, 517, 1967, 887, 841, 2520, 2683, 2275, 2362, 2363, 1478, 2027, 2281, 191, 1971, 2209, 2769, 3057, 1858, 2114, 771, 1083, 1233, 1387, 2272, 2276, 3107, 625, 669, 1843, 2350, 956, 2387, 2476, 2756, 1768, 2669, 2565, 1868, 1906, 1551, 2778, 2803, 2447, 1704, 2594, 2648, 1163, 1305, 2999]. Linear [1450, 2293, 627, 954, 2574, 1691, 2586, 2218, 2266, 3768, 2298, 1953, 146, 867, 1053, 1969, 2958, 2065, 2033, 2131, 2066, 1383, 1464, 1174, 2559, 1407, 1538, 2960, 1522, 2521, 2236, 2202, 950, 1550, 2051, 1769, 947, 1069, 1635, 3066, 2140, 231, 2822, 1002, 668, 1107, 1449, 1866, 2292, 1712, 948, 1855, 1563, 2823, 769, 1008, 1922, 1501, 203, 2859, 321, 2103, 1738, 188, 245, 2367, 1164, 1191, 1671, 558, 2486, 2192, 2628, 2943, 2899, 2558, 3785, 3234, 3708, 3706, 3709, 3391, 3710, 3544, 3575, 3711, 3702, 3703, 3704, 3712, 3320, 3244, 3367, 3258, 3380, 2234, 3362, 3325, 3592, 3697, 3707, 3713]. linear [3219, 3714, 3589, 3759, 3678, 3715, 3716, 3681, 3301, 3562, 3129, 3227, 3607, 3717, 3718, 3719, 3588, 3520, 3403, 3467, 3720, 3705, 3721, 3722, 3701, 3725, 3482, 3478, 3172, 3145, 3619, 3628, 3850, 3500, 3262, 3547, 3791, 3723, 3131, 3228, 3724, 3123, 3442]. Linear-Utility [746]. Linearization [1717].

Lines [2596, 1046, 146]. link [3260, 3161]. Linkage [271]. linked [3768, 3620]. Linton [3712, 3926]. Liquids [1811]. Literature [225]. Liu [3925, 3833]. LMP [1617]. Local [2831, 1760, 2208, 2617, 2847, 2513, 2992, 2531, 2425, 2217, 2414, 2494, 2663, 650, 667, 649, 2366, 3419, 1952, 2596, 2655, 2669, 3112, 3105, 1838, 3148, 3113, 3190, 2439, 2876, 3318, 2749, 2775, 1968, 2017, 2129, 2486, 3202, 2572, 3284, 3551, 3367, 3521, 3309, 2499, 3218, 2055]. localisation [3843]. localized [3249]. Locally [2027, 2354, 1691, 770, 2220, 2911, 2947, 3457, 3305, 3249]. Location [1093, 373, 1265, 1812, 2340, 975, 828, 856, 2043, 1972, 463, 2165, 2181, 1738, 2382, 2816]. Locations [2753]. Loci [2662]. Loève [3284]. Log [2614, 2743, 2068, 2771, 2853, 1840, 2281, 1083, 2981, 1945, 2266, 1407, 1769, 1563, 3245, 3135, 3415, 3244, 3239]. Log-Concave [1945, 3245, 3135, 3415]. Log-Excesses [2981]. Log-Gaussian [2614, 2743, 3239]. Log-Linear [2068, 2771, 2281, 1083, 2266, 1407, 1769, 1563, 3244]. Log-Normal [1840]. Log-Splines [2853]. Logarithm [108]. Logarithmic [120, 414, 128]. Logic [601, 2081]. Logical [599]. Logistic [1056, 2339, 1977, 2111, 2450, 2281, 2231, 2211, 2221, 1871, 497, 3017, 2085, 2982, 1761]. Lognormal [1275, 633]. Long [2039, 2240, 91, 1853, 2859, 3560]. long-run [3560]. Long-Term [91]. Longitudinal [2697, 2849, 2961, 2377, 2520, 2242, 3019, 2820, 2926, 2796, 2892, 3073, 3051, 3322, 3425, 3386, 3218, 3804, 3303]. Look [1521, 3215, 3470, 3910]. Looking [2017]. Loop [1069]. Lorenzo [3863]. Loss [1730, 167, 1066, 453, 1826, 3761]. Lot [1431, 1181]. Lote [126]. Lovell [2259]. Low [2750, 2112, 2843, 3598, 3262]. Lower [635, 2656, 823, 1263, 1714]. Ltd. [80]. Lu [3915]. Lyapunov [2221, 2017].

M [4, 3674, 363, 3955, 3915, 2492, 1222, 465, 1853, 774, 1046]. M. [3674, 3915]. M/E [774]. M/G/1 [465, 1853, 1046]. M/G/1/N [1222]. Machine [536, 170, 364, 396, 440, 2429, 3921, 3938, 3922, 3923, 3924, 3920, 3925, 3926, 3927, 3928, 3929, 3937, 3916, 3919, 3930, 3914, 3931, 3932, 3933, 3934, 3935, 3918, 3913, 3936, 3917, 3915, 42, 3991]. Machine-Generated [2429]. Machines [193, 212, 251, 304, 124, 466, 380, 381, 548, 1716, 3338]. MacKenzie’s [3916]. Macromolecular [2890]. Made [1695, 367]. Magnetic [2890]. Main [1091, 1117, 653, 3617]. Main-Effect [1091]. Majorants [942]. Majorization [2201]. Making [3561, 3355, 3368]. MALMEM [3547]. Man [280, 886]. Manifold [3807, 3745, 3275, 3150, 3741, 3632]. manifolds [3621]. Manimax [622]. Manimax-Regret [622]. Mann [3216, 1554]. MANOVA [976]. Manufacture [44]. Many [1057, 396, 750, 820, 3847, 3989, 3986, 3988]. Mapping [2968, 3608]. Maps [2221, 2615, 159]. Marbles [2253]. March [89]. Mardia [3919]. Marginal [2709, 3116, 2996, 1320, 3244, 3012, 2461, 2926, 2090, 2760, 1864, 3311, 3482, 3145]. Marginalization [2741, 1123, 1460]. Marginally [2930, 3295]. Mark [3863, 3718, 3836, 3956]. Marked [2992, 2579, 2753, 3225]. Markers [2552]. Markoff [406, 315]. Markov [641, 1830, 3120, 2062, 3874, 3807, 2212, 1239, 2879, 3938, 3912, 2190, 2841, 2736, 405, 2681, 434, 2724, 2227, 2938, 295, 1663, 3183, 3278, 612, 3244, 512, 2738, 2913, 2048, 632, 3745, 2584, 2786, 3306, 3582, 3138, 3660, 782, 3161, 1187, 1099, 2417, 1640, 1985, 1741, 2504, 2479, 2957, 3438, 2583, 2802, 2637, 2059, 327, 3146, 2289, 3846]. Markov-Chain [632]. Markov-Modulated [2913]. Markovian [2656, 1071, 1149]. Marks [2753]. Marshall [1925]. Marta [3864, 3974]. Martin [3893, 3895]. Martingale [3864, 3860, 3853, 3863, 3857, 3862, 3859, 3854, 3870, 3861, 3855, 3858, 3856, 3868, 3865, 3866, 3867, 3869]. Martinussen [3954]. Mask [515]. Masked [2154]. Masking [2205]. Mass [2413]. massive [3563, 3283]. Matched [838, 885, 930, 1602, 1881, 1422, 2947]. Matching [401, 478, 1113]. Materials [18]. Matérn [3568, 3426]. Mateu [3675, 3930]. Mathematical [240, 561, 1935]. Mathematically [1389]. Matrices [191, 213, 1352, 976, 1036, 1796, 1835, 1909, 1180, 1279, 499, 734, 668, 3170, 3416, 3476, 3202]. Matrix [1251, 1402, 1309, 3480, 1868, 1519, 628, 670, 1295, 1691, 1549, 1350, 2467, 2650, 2236, 2202, 561, 138, 2004, 2268, 3782, 3881, 3574, 3619, 3875, 3271, 3505]. matrix-valued [3875]. matrix-variate [3881]. Mats [3705]. Matteo [3974]. Matter [1341, 1, 23, 30, 17, 24, 39, 48, 31, 40, 57, 65, 49, 58, 66, 72, 87, 78, 88, 97, 110, 121, 141, 155, 173, 195, 221, 247, 259, 270, 291, 335, 319, 336, 350, 383, 417, 506, 593, 619, 617, 647, 662, 683, 691, 740, 767, 780, 765, 779, 799, 800, 819, 835, 817, 834, 851, 852, 871, 890, 869, 889, 910, 911, 933, 952, 931, 951, 966, 979, 982, 999, 980, 998, 1009, 1024, 1027, 1041, 1025, 1040, 1051, 1052, 1077, 1079, 1063, 1078, 1097, 1098]. Matter [1137, 1139, 1119, 1138, 1152, 1153, 1168, 1195, 1166, 1182, 1196, 1197, 1214, 1242, 1199, 1229, 1243, 1257, 1259, 1284, 1244, 1270, 1285, 1298, 1313, 1330, 1299, 1314, 1331, 1332, 1345, 1374, 1344, 1361, 1375, 1377, 1409, 1411, 1391, 1410, 1419, 1435, 1453, 1468, 1436, 1454, 1469, 1489, 1511, 1525, 1490, 1512, 1526, 1539, 1555, 1569, 1540, 1556, 1570, 1589, 1608, 1621, 1590, 1609, 1622, 1641, 1661, 1681, 1642, 1662, 1682, 1700, 1724, 1744, 1701, 1725, 1745, 1758, 1773, 1784, 1759, 1774, 1785, 1797, 1807, 1820, 1798, 1808]. Matter [1821, 1836, 1846, 1860, 1837, 1847, 1861, 1875, 1889, 1899, 1876, 1890, 1900, 1917, 1927, 1937, 1918, 1928, 1938, 1955, 1973, 1991, 1956, 1974, 1992, 2012, 2034, 2056, 2013, 2035, 2057, 2079, 2101, 2116, 2133, 2080, 2102, 2117, 2134, 2157, 2171, 2182, 2195, 2158, 2172, 2183, 2196, 2214, 2225, 2237, 2250, 2215, 2226, 2238, 2251, 2270, 2284, 2296, 2310, 2271, 2285, 2297, 2311, 2329, 2343, 2359, 2370, 2330, 2344, 2360, 2371, 2393, 2407, 2420, 2434, 2394, 2408, 2421, 2435, 2454, 2469, 2482, 2500, 2455, 2470, 2483, 2501, 2514]. Matter [2529, 2543, 2562, 2515, 2530, 2544, 2563, 2575, 2589, 2601, 2619, 2576, 2590, 2602, 2620, 2631, 2643, 2659, 2678, 2632, 2644, 2660, 2679, 2700, 2717, 2730, 2744, 2701, 2718, 2731, 2746, 2764, 2780, 2797, 2810, 2765, 2781, 2798, 2811, 2826, 2835, 2845, 2854, 2866, 2827, 2836, 2846, 2855, 2867, 2877, 2887, 2900, 2909, 2919, 2878, 2888, 2901, 2910, 2920, 2928, 2939, 2952, 2964, 2976, 2929, 2940, 2953, 2965, 2977, 2994, 3005, 3015, 3026, 3039, 2995, 3006, 3016, 3027, 3040, 3058, 3068, 3080, 3088, 3099, 3059, 3069, 3081, 3089]. Matter [3100, 3110, 3118, 3124, 3133, 3142, 3111, 3119, 3125, 3134, 3143]. MAWII [3988]. Max [2119, 3246, 3481]. max-copula [3481]. Max-Stable [2119, 3246]. Maximin [3990, 3494, 3740]. maximization [2830, 3071, 3175, 3176, 3191]. Maximizing [2453, 1710, 462, 3437]. Maximum [1070, 481, 995, 1023, 1319, 1533, 1670, 1011, 1241, 1675, 2240, 1156, 615, 166, 2339, 1475, 1405, 2930, 1794, 2198, 1965, 1247, 3135, 1286, 1086, 1995, 2414, 2228, 2686, 2036, 2155, 1223, 1275, 2229, 1643, 1884, 2584, 1843, 2032, 2851, 94, 613, 2711, 2163, 1385, 1949, 1297, 1522, 546, 2578, 1476, 2510, 2164, 623, 1516, 1754, 543, 1714, 1770, 2292, 1629, 1685, 2859, 2103, 2094, 2933, 2355, 1776, 2954, 3442, 3245, 3812, 3145, 948]. Maximum-Likelihood [481, 995, 1023, 1011, 1156, 94, 613, 546, 623, 543]. Maxway [3872]. May [55, 62]. McCullagh [2720]. McIntosh [74]. MCMC [3694, 3738, 3778, 3688]. Mean [330, 171, 3091, 1273, 1160, 2905, 945, 1272, 516, 1863, 391, 2506, 273, 2550, 2176, 578, 840, 2185, 2224, 288, 1654, 375, 1951, 1647, 2838, 893, 2327, 2883, 526, 808, 572, 1142, 1162, 1488, 993, 2734, 2289, 2044, 3163, 3311, 3396, 3387, 3217, 3739, 3578, 3422, 3813, 3287, 3456, 3742, 3627, 3964]. mean-field [3964]. Mean-Squared [2883, 3578]. Mean-Value [1142]. Means [528, 541, 611, 1118, 294, 472, 582, 566, 1477, 3055, 438, 3007, 325, 1828, 535, 642, 3889, 3967, 3267, 3770, 3890, 3886, 3891, 3892, 263, 3893, 3894, 3895, 3887, 3896, 3897, 3898, 3899, 3900, 3901, 3888, 3902, 3885]. Measles [2019]. Measure [1310, 2008, 2324, 2673, 3848, 3235]. Measured [2971, 2546]. Measurement [1005, 968, 1977, 2722, 2967, 2948, 2473, 1922, 1171, 1582, 2536, 3797, 3148, 3455, 3349, 3547]. Measurements [2338, 1731, 1964, 106, 2801, 167, 157, 3299, 3345, 3628]. Measures [1514, 1421, 2568, 1047, 2512, 2114, 2987, 2557, 1305, 1783, 1984, 2872, 2263, 2401, 3031, 2661, 3444, 3408, 3579]. Measuring [704, 2570, 2299]. Mechanical [140]. Mechanism [2447]. Mechanisms [2942, 2822]. Median [1575, 695, 2064, 3001, 94, 1883, 2542, 1723, 376]. Median-Effective [94]. Median-Restricted [3001]. Medians [1719, 2233]. Mediation [3428, 3585, 3943, 3458, 3942, 3316, 3728]. mediator [3316]. mediators [3428, 3728]. Medical [1966]. Medicine [2061]. meet [3665, 3626]. Meeting [3921, 4, 32, 80, 89, 90, 111, 2022, 2062, 3938, 3922, 3923, 3924, 3920, 3925, 50, 3926, 3927, 3928, 3929, 3937, 3916, 3919, 3930, 3914, 3931, 3932, 3933, 3934, 3935, 3918, 3913, 3936, 3917, 3915, 2788, 2667]. Meier [1958]. Meiting [3926]. Meixner [1238]. Members [104]. membership [3699]. Memory [2240, 2859, 3375]. Memoryless [2481]. Mendelian [3988]. Meng [2720]. merge [3330]. Merging [2832, 2806]. Message [2003]. meta [3233, 2637]. meta-analysis [3233, 2637]. Meteorology [67]. Method [801, 1000, 1156, 1426, 493, 1947, 1428, 3004, 347, 2191, 1625, 695, 1176, 106, 557, 581, 297, 2169, 3011, 1237, 2641, 51, 94, 613, 2926, 2206, 510, 2893, 2381, 412, 20, 1921, 7, 2089, 2793, 2808, 2761, 1865, 1926, 1380, 833, 546, 2504, 2188, 214, 878, 1987, 1338, 28, 2045, 904, 2075, 102, 2261, 1496, 198, 2616, 661, 711, 3809, 3160, 2600, 3502, 3523, 3320, 3624, 3649, 3906]. Method-of-Moments [2926]. Methodology [2567, 3346, 3406]. Methods [1005, 2062, 145, 2440, 59, 401, 86, 1812, 2227, 2278, 2828, 2291, 320, 540, 8, 2881, 1644, 2425, 1090, 2663, 311, 2478, 191, 213, 356, 830, 1957, 44, 18, 1603, 2786, 2276, 2303, 2480, 2445, 2737, 2708, 2860, 2882, 3077, 3083, 956, 1848, 92, 2346, 3047, 2000, 984, 1060, 2464, 1201, 2691, 1567, 1969, 2150, 219, 2917, 1680, 2949, 1030, 1851, 5, 2726, 2886, 2173, 1536, 2187, 392, 1080, 68, 2059, 2857, 2125, 19, 1886, 123, 2302, 60, 3120, 3552, 3503, 3450, 3290, 3155, 3150, 3745, 3582, 3440, 3944, 2636, 3353]. Metric [2649, 3198]. Metric-Based [2649]. Metrics [1005]. Metropolis [2831, 2072, 2775]. Michael [3925, 3716, 3980]. Microarray [2850, 3062]. microarrays [3468]. Microbial [726]. Microdata [2673]. microscopy [3359]. Mid [544]. Mid-Ranges [544]. Mill [827, 565]. Miller [3900]. Minima [1952, 2055]. Minimal [1371]. Minimally [3117]. Minimax [1734, 1148, 388, 3391, 3634, 3681]. Minimaxity [1223]. Minimization [595, 3746]. Minimizing [322, 2325]. Minimum [2441, 2905, 1581, 1969, 1208, 1543, 1088, 1814, 1577, 2003, 3434]. Misclassification [1828]. Mises [1295, 1631, 2264, 1075, 1219, 1327, 1188, 1151, 1323, 1381, 973]. Misleading [1403].

Mismeasured [2219]. Missing [1212, 2697, 1003, 493, 3004, 1710, 855, 2656, 2447, 2464, 2342, 1385, 2969, 1630, 1018, 2927, 2726, 2219, 1729, 1277, 2822, 84, 3544, 3216, 3504, 3779]. Missingness [3000, 3736, 3772]. Misspecification [2609, 2570, 1656, 3576]. Misspecified [2568, 2593, 3259, 3692, 3500]. mitigating [3510]. Mitochondrial [2664]. Mixed [2725, 2453, 2326, 2930, 2804, 2758, 1334, 2523, 2657, 884, 3077, 2756, 2317, 1796, 1835, 2462, 3029, 2958, 2138, 2880, 1538, 569, 589, 2822, 2353, 2505, 436, 2383, 2628, 2558, 3785, 3320, 3402, 3219, 3678, 2412, 3331, 3500]. Mixing [2137, 2367, 2933, 1035, 3765]. Mixture [1687, 883, 2697, 2935, 2970, 2535, 2566, 2754, 3023, 2166, 1887, 732, 730, 2295, 1494, 1673, 1699, 2557, 895, 2174, 2941, 2963, 1104, 3082, 1813, 2187, 2556, 1577, 2511, 2087, 2507, 2748, 3178, 3601, 3551, 3476, 3372, 3169, 3982]. Mixtures [1161, 2668, 892, 1601, 539, 2670, 2260, 2088, 1342, 1356, 860, 859, 935, 939, 761, 1039, 248, 3365, 452, 2361, 2418, 1145, 423, 456, 518, 620, 1445, 1343, 3345, 3167, 3235, 3257, 2419]. Modal [2127]. Modality [2413]. Mode [2423]. Model [2725, 1636, 1492, 2212, 1670, 1637, 2844, 536, 1102, 1056, 1118, 1106, 2935, 3056, 2453, 1207, 1750, 2662, 2712, 1812, 2654, 1293, 1587, 2227, 2111, 1074, 2930, 2754, 2441, 2938, 1451, 1230, 3078, 1561, 2847, 1731, 2279, 3023, 1964, 2955, 1309, 2832, 2727, 2609, 1967, 2199, 2676, 1354, 2699, 2376, 944, 3576, 2209, 2769, 2178, 1674, 1312, 2897, 679, 810, 3377, 2570, 894, 2387, 2476, 2756, 1768, 1948, 1868, 1692, 3049, 2623, 240, 1483, 2711, 1691, 2580, 2210, 2552, 3036, 685, 2833, 1053, 2695, 3259, 2176, 1357, 1648, 3693, 3796, 1383, 509, 3000, 2949, 2230, 2200, 1568, 150, 1550, 1020]. Model [2051, 1804, 1001, 1741, 2984, 1458, 1635, 1954, 1136, 3281, 2822, 1716, 1655, 1276, 2638, 2890, 1107, 1417, 1508, 1770, 1752, 2661, 2429, 1288, 1408, 3662, 1582, 1474, 431, 2294, 1685, 1877, 2472, 2120, 2911, 1844, 1372, 2740, 3051, 2355, 1751, 726, 1864, 2235, 2507, 2629, 2871, 2044, 2899, 2904, 2558, 3883, 3785, 3299, 3708, 3706, 3952, 3938, 3812, 3951, 3434, 3709, 3710, 3563, 3658, 3448, 3711, 3564, 3702, 3953, 3703, 3704, 3712, 3957, 3402, 3489, 3950, 3762, 3622, 3707, 3713, 3714, 3759, 3678, 3254, 3715, 3787, 3790, 3716, 3698, 3872, 3285, 3173, 3954, 3435, 3632, 3908, 3421, 3629, 3717]. model [3372, 3718, 3532, 3537, 3406, 3719, 3955, 3653, 3948, 3403, 3479, 3407, 3720, 3949, 3593, 3956, 3705, 3618, 3721, 3722, 3773, 3701, 3725, 3655, 3413, 3850, 3547, 3964, 3481, 3723, 3724, 3982, 3488, 3570]. model-adaptive [3377]. Model-assisted [3662]. Model-Based [2376]. Model-Free [3056, 2833]. model-X [3872, 3488]. Modelling [2015, 3782, 3393, 1961, 2119, 2274, 1503, 3018, 2670, 2853, 2807, 2061, 2967, 3019, 3759, 3747, 1702, 2341, 2090, 2962, 1302, 3073, 3149, 1711, 3643, 2094, 3392, 2776, 3498, 2026, 2011, 2527, 3203, 2667, 3299, 3277, 3275, 3610, 3419, 3270, 3247, 3612, 3294, 3197, 3303]. Models [1595, 1687, 1703, 2671, 1228, 1000, 1183, 2608, 3025, 2577, 2639, 1239, 1347, 2879, 883, 1924, 2145, 2240, 2190, 1564, 2697, 876, 2338, 1475, 1977, 2168, 2306, 2068, 2997, 2093, 2450, 2566, 2611, 2599, 1908, 2050, 2706, 2938, 2804, 2771, 1965, 1811, 2320, 1839, 3114, 1620, 1982, 2108, 1067, 2758, 1523, 2884, 2458, 1867, 199, 1663, 1857, 1929, 1823, 2992, 1573, 1644, 2735, 2996, 2523, 480, 2503, 2321, 1902, 934, 2789, 2526, 512, 641, 2520, 2683, 2275, 2738, 2795, 2295, 2487, 2362, 2363, 2373, 824, 2242, 1782, 2281, 1038, 1971, 2209, 2657, 1352, 2105, 605, 2114]. Models [2243, 2231, 1307, 1363, 1424, 1083, 2981, 2130, 2349, 2584, 2649, 2234, 2656, 2948, 3077, 1448, 1486, 3001, 1843, 2350, 1506, 1562, 2118, 2325, 1482, 2557, 2088, 1237, 2409, 1034, 3033, 1369, 1442, 1504, 1894, 1551, 2245, 2569, 2778, 2803, 2491, 2528, 2317, 2447, 2648, 2968, 1305, 1799, 2293, 984, 1060, 1845, 2574, 1726, 2591, 2586, 2719, 2218, 2266, 2645, 2298, 1505, 2941, 2963, 2462, 2463, 3043, 2137, 1186, 2752, 2840, 3029, 2936, 3086, 1962, 1969, 2958, 2033, 2222, 1264, 2066, 1437, 1660, 2516, 3017, 2138, 1464, 1364, 2089, 2452, 1315, 2880, 1638, 2793, 2553, 1949, 2559, 1132, 2417]. Models [2388, 2794, 1538, 2960, 2521, 3082, 2694, 2374, 3014, 2760, 486, 912, 1656, 1865, 1926, 3021, 1985, 2123, 1813, 2715, 1769, 3053, 997, 1507, 3095, 3066, 2305, 2504, 2751, 2705, 3060, 1216, 2353, 1159, 2031, 1761, 1740, 1558, 2405, 2505, 1516, 1754, 2651, 2883, 1449, 2537, 1866, 2292, 1712, 1855, 1563, 2823, 1922, 2556, 1501, 2604, 654, 1629, 2729, 2813, 2497, 2734, 2045, 967, 2449, 2367, 2696, 2511, 2261, 1191, 1931, 1903, 1671, 1108, 1776, 3008, 2444, 2869, 2628, 2286, 2954, 3034, 2572, 2748, 3178, 3819, 3600, 3234, 3206, 3429, 3912, 3848, 3503, 3297, 2560, 3647, 3071, 3175, 3176, 3366]. models [3185, 3726, 3231, 3284, 3405, 3266, 3352, 3480, 3541, 3064, 3400, 3572, 3656, 3320, 3595, 3774, 3244, 3947, 3447, 3201, 3168, 3601, 3361, 2573, 3224, 3745, 3794, 3377, 3380, 3362, 2816, 3592, 3697, 3152, 3253, 3219, 3589, 3761, 3816, 3496, 3260, 3485, 3326, 3164, 3641, 3301, 3788, 3562, 3259, 3129, 3148, 3227, 3046, 3410, 3302, 2412, 3692, 3331, 3397, 3343, 3357, 3169, 3438, 3241, 3520, 3318, 2636, 3467, 3986, 3664, 3200, 3446, 3386, 3478, 3845, 3145, 3146, 3310, 3500, 3875, 3262, 3547, 3791, 3369, 3507, 3208, 3131, 3228, 3123, 3356, 3817, 3442, 3982]. moderately [3338]. Moderating [1819]. modes [3340]. Modification [1061, 2169, 3690]. Modifications [2078, 1151]. Modified [118, 879, 2566, 1387, 810, 393, 3611]. modular [3579]. Modulated [2913, 3043]. Molecular [2666, 2545]. Moment [3517, 701, 2579, 700, 180, 525, 327, 3726, 3533, 3397, 2633]. moment-adjusted [3533]. moment-based [3397]. Moments [1070, 1451, 317, 2024, 2368, 2431, 1906, 2926, 2185, 2224, 2043, 538, 623, 723, 1050, 1103, 1303, 1722, 3809]. Mondrian [3640]. Monitoring [2857, 1751]. Monod [2727]. Monotone [2697, 3798, 1764, 717, 2400, 1816, 3050]. Monotone-Hazard [1764]. Monotonicity [1689, 1147, 653]. Monro [695, 3624]. Monte [3120, 2062, 3874, 3807, 3503, 2453, 2681, 2830, 2724, 2227, 3398, 2828, 3680, 2891, 3183, 1644, 2029, 2281, 3315, 2036, 2155, 2573, 3150, 3806, 3745, 2584, 2786, 1896, 280, 908, 3582, 3578, 3904, 2719, 2460, 2417, 3417, 3605, 2504, 2957, 2637, 2059, 2289]. Monte-Carlo-assisted [3578]. Moore [2083]. Most [468]. Motion [1185, 3379]. move [3651]. Moving [2240, 540, 1307, 1363, 1506, 1562, 640, 1439, 2150, 1909, 2089, 1297, 1020, 137, 181, 3424, 2573, 3407, 3356]. Moya [3856]. Mr [492, 518]. Mr. [53]. Multi [606, 2377, 1135, 571, 1250, 1058, 709, 893, 525, 993, 1367, 1438, 3798, 3409]. Multi-Armed [1438]. Multi-Factor [1135, 571, 3798]. Multi-normal [525]. Multi-Phase [2377]. multi-resolution [3409]. Multi-Sample [1058]. Multi-Server [1250, 1367]. Multi-Stage [606, 893, 993]. Multi-Term [709]. multiarmed [3776]. Multicollinearity [965]. Multicolour [2228]. multicomponent [3316]. Multidimensional [1514, 2884, 2605, 3047, 2095, 1923, 1985, 1358, 1400, 2352, 1368, 3135, 3766]. Multifactor [1104]. multilayer [3441, 3794]. Multilevel [2374, 3316]. Multimodal [1141]. Multimodality [1484]. Multinomial [743, 1672, 836, 1165, 2266, 1962, 1476, 949, 590, 2004, 2253, 2268]. Multinomials [1749, 3983]. Multinormal [591, 425]. Multiparameter [1717, 913, 2037, 1508]. Multipath [2740]. Multiperiodic [2737]. Multiple [3990, 2213, 2190, 2313, 2902, 1402, 3314, 75, 863, 2956, 3288, 874, 1004, 2040, 957, 1055, 357, 2626, 2968, 628, 2896, 939, 1017, 3518, 854, 2381, 1385, 2282, 1350, 3104, 1030, 157, 1669, 3061, 268, 2527, 215, 3531, 3505, 3542, 3265, 3268, 3278, 3648, 3370, 3477, 3237, 3493, 3733, 3372, 3357, 3740, 3520, 3296, 3545, 3373, 3728]. Multiple- [939]. Multiple-change-point [3314, 3288]. Multiple-Comparison [2626]. Multiple-Disease [2968]. Multiple-Imputation [2896]. multiple-regression [3520]. Multiple-Stage [2190]. Multiple-Type [863]. Multiplex [785]. Multiplicative [2321, 1840, 1352, 2473]. Multiplier [1442, 1504, 1582]. Multipliers [1351]. Multiply [3751, 3580, 3487]. Multiplying [307]. Multiregression [2123]. multiresolution [3139]. Multiresponse [2813]. Multisample [2153, 1141]. Multiscale [3273, 2870, 3047, 3560, 3164, 3423, 3684, 3168]. Multistage [996, 2347, 3394]. Multistate [2623]. Multistep [2790, 2325]. Multistep-Ahead [2790, 2325]. Multitype [2422, 2861]. Multivariate [2475, 2704, 404, 144, 1212, 435, 511, 562, 699, 342, 2935, 3568, 2313, 1557, 1789, 2416, 493, 2986, 2093, 429, 1961, 1503, 2991, 1870, 1309, 3085, 2905, 719, 2526, 3018, 944, 2571, 811, 630, 652, 629, 2231, 2229, 2040, 2169, 3107, 1261, 1328, 2782, 1481, 2152, 2153, 2565, 1509, 1504, 1888, 1704, 777, 909, 763, 628, 627, 806, 1984, 1100, 1612, 2342, 1505, 1994, 2090, 1324, 882, 1909, 2089, 2465, 2406, 1723, 1522, 1443, 1458, 1635, 3063, 1547, 2161, 2848, 494, 927, 653, 572, 672, 1877, 187, 899, 1022, 1108, 2486, 2943, 2044, 3351, 3757, 2382, 3962, 3275, 3789, 3361]. multivariate [3469, 3784, 3257, 3693, 3796, 3406, 3732, 3498, 3481, 3982]. Murphy [2703]. must [1497]. Mutation [239]. mutual [3484]. Mutually [1049]. Muzhe [3841]. my [518, 3900].

N [3889, 1096, 3674, 3915, 130, 2492, 1222]. nanoscale [3359]. Narrowest [3542]. Narrowest-over-threshold [3542]. National [42, 3991]. Natural [1601, 809]. Nature [150]. Navigation [1185, 1184]. Navy [123]. Near [2425, 2379, 3189]. Near-Parametric [2425]. Nearest [1347, 1056, 1386, 2839, 2641, 1225, 1591, 1613]. Nearest-Neighbour [1056]. Nearly [1995, 2073]. Necessary [495]. Negative [317, 397, 2349, 921, 2615, 1647, 1467, 1115, 2852, 127, 2931, 3945, 3610, 3580]. Negativity [2126]. Neighbour [1347, 1056, 2839, 2641, 1591, 1737, 1613]. Neighbouring [1485]. Neighbours [1386]. Nerve [2448]. Nested [2958, 3060, 2031, 2356, 3787]. Net [2864, 2834, 2865]. Network [1757, 2914, 3972, 3973, 3503, 3974, 3981, 3968, 3945, 3975, 3976, 3969, 3768, 3977, 3788, 3970, 3971, 3696, 3979, 3980, 3978]. network-linked [3768]. Networks [1013, 1727, 2173, 1728, 1826, 3347, 3658, 3679, 3699, 3660, 3254, 3485, 3435, 3781, 3247, 3871, 3766, 3778, 3479, 3509, 3878]. Neumann [3991]. Neural [2173, 3778, 3696, 3878]. neuroimaging [3164]. Neutral [2363]. Newton [2346, 3851]. Neyman [382]. Ng [3980]. Nguyen [3896]. Niall [3922]. Nicolae [2720]. Nielsen [2733]. Nikodym [703, 739, 704]. Niu [3834]. Niwen [3724]. NN [1591]. No [366, 2609, 339, 382]. node [3479]. Noise [1842, 1867, 2332, 1132, 2140, 3807, 3166, 3432, 3388]. Noisy [2160, 2020, 2016, 3768, 3619]. Non [1595, 847, 925, 1093, 1778, 731, 559, 395, 2098, 2577, 1553, 473, 441, 2190, 1564, 2951, 3345, 876, 1273, 2018, 3117, 2970, 2015, 2068, 3277, 1794, 2050, 2198, 2915, 1296, 1451, 2771, 1503, 2314, 1135, 2955, 2458, 1663, 2972, 3013, 3943, 1086, 2503, 3553, 3094, 1045, 1014, 1116, 2362, 2363, 2027, 2376, 582, 2769, 2571, 45, 2934, 3340, 614, 2300, 2127, 2349, 3159, 2234, 1494, 2276, 3020, 3077, 625, 669, 1627, 1280, 1329, 1551, 2778, 921, 2447, 3007, 938, 1087, 1163, 3112, 1355, 3257, 3226, 3440, 1304, 3054, 829, 3813, 2554, 2963, 2282, 2936, 853, 2396, 2065, 2131, 828, 856]. Non [918, 969, 1091, 2066, 2693, 2559, 2908, 3136, 2993, 2051, 714, 762, 924, 1085, 1121, 886, 1117, 997, 1360, 1607, 709, 750, 2772, 642, 2822, 1467, 2728, 1980, 3050, 169, 1684, 1998, 1752, 3688, 1577, 2852, 2356, 2933, 1164, 504, 738, 558, 2192, 2931, 3034, 2950, 3517, 2560, 3449, 3364, 3506, 3549, 3300, 3346, 3355, 3551, 3258, 3419, 3408, 3328, 3295, 3560, 3256, 3519, 3607, 3286, 3331, 3602, 3520, 3149, 3269, 3386, 3455, 3498, 3577, 3218, 3146, 3308, 3982]. Non-Additivity [614]. Non-agency [3943]. Non-Bayesian [1980]. non-causal [3419]. Non-Centred [2693]. Non-Compliance [2915, 2955, 2950]. Non-Conjugate [2458, 2571]. Non-Consecutive [2396]. Non-Counterfactual [2934]. Non-Crossing [3013]. Non-Decreasing [1553, 921]. Non-Diffusion [1663]. Non-Elliptically [1503]. Non-Equilibrium [395, 642]. Non-Full [1163, 1164]. Non-Gaussian [2577, 2503, 2127, 2772, 1752]. Non-Hierarchical [2282]. Non-Homogeneous [1778, 441, 2190, 829, 1998, 3149]. Non-Ignorable [2068, 2771, 2376, 2447, 2822, 3813]. Non-Independent [1360, 1607]. Non-Linear [1595, 731, 2098, 473, 876, 2018, 2015, 2050, 1451, 1086, 2362, 2363, 2027, 2769, 2276, 625, 669, 1551, 2778, 2065, 2131, 2066, 2051, 558, 2192, 2234, 3258, 3607, 3520]. Non-Local [3112, 3551]. Non-Negative [2349, 921, 1467, 2852, 2931]. Non-Nested [2356]. Non-Normal [1093, 1045, 1014, 1116, 45, 1304, 2936, 169]. Non-Normality [559, 2300]. Non-Null [504]. Non-Optimal [1280, 1329]. Non-Orthogonal [1135, 582, 938, 1087, 1091, 886, 1117, 750]. Non-Parametric [1553, 1564, 2951, 1273, 3117, 2970, 1296, 2972, 3013, 3094, 1494, 3020, 3077, 1627, 1355, 3054, 2963, 828, 856, 918, 969, 2993, 1085, 1360, 1607, 3050, 1684, 1577, 2933, 3034, 3345, 3277, 3553, 3340, 3159, 3257, 3226, 3440, 3813, 3136, 3449, 3364, 3506, 3549, 3300, 3346, 3355, 3328, 3295, 3560, 3256, 3519, 3286, 3331, 3602, 3269, 3386, 3455, 3218, 3146, 3982]. non-parametrics [3517, 3258, 3408]. non-probability [3577]. Non-Random [2314]. Non-Regular [1794, 2198]. Non-Response [2068, 2771, 2376, 3813]. Non-reversible [3688]. non-separable [2560]. Non-Standard [2554, 2559]. Non-Stationarity [924]. Non-Stationary [847, 925, 853, 2908, 714, 762, 1121, 997, 709, 2728, 738, 3498, 3308]. Non-Zero [3007]. Noncentrality [1234]. Nonconformity [3115]. Nonconsecutive [1776]. nonconvex [3643]. Noniterative [1277]. Nonlinear [1028, 1692, 1702, 3880, 3987]. Nonlinearity [1421]. Nonparametric [2099, 2805, 1308, 2106, 2041, 2790, 2021, 2881, 2424, 2096, 2067, 3748, 3631, 2048, 2139, 1707, 2423, 1853, 1910, 1673, 1699, 2077, 2276, 2607, 2737, 2708, 2801, 2792, 2557, 3797, 2851, 2395, 2597, 1795, 1854, 2793, 2389, 2814, 2388, 2230, 2768, 2578, 2154, 2165, 2397, 2451, 1968, 2367, 2471, 2740, 2629, 2430, 2427]. Nonparametrically [1951]. nonseparable [3871]. Nonzero [1796, 1835]. norm [3693, 3796]. Normal [1093, 1070, 1440, 1161, 1532, 2475, 560, 1089, 342, 1750, 1587, 1402, 2340, 2991, 3023, 887, 472, 2905, 651, 719, 2526, 1045, 1014, 1116, 1840, 130, 2366, 2571, 45, 630, 652, 1477, 1481, 3007, 777, 909, 954, 240, 1304, 300, 609, 2936, 483, 1264, 856, 578, 2185, 2224, 396, 288, 2043, 1911, 2439, 375, 535, 1547, 538, 565, 893, 169, 494, 572, 377, 722, 993, 610, 815, 390, 3178, 3391, 3422, 3456, 525]. normalisation [3763]. Normalised [3848]. Normality [559, 1146, 1428, 1742, 2755, 1150, 1749, 958, 1394, 2300, 1395, 1612, 1723, 1274, 1805, 1806, 1248, 899]. normalization [3584]. Normalized [1780, 3235, 3140, 3121]. Normalizing [2694, 3278]. Normally [108]. North [183]. Note [3, 265, 330, 1983, 1241, 1781, 307, 1485, 699, 885, 171, 754, 86, 2767, 879, 1475, 787, 2160, 2716, 1561, 501, 875, 2092, 237, 1574, 1827, 523, 1351, 945, 1373, 657, 1736, 1478, 2139, 1061, 2366, 326, 1651, 592, 669, 1843, 745, 902, 1506, 1562, 1173, 670, 101, 1101, 1524, 469, 1709, 678, 2065, 1390, 1350, 2185, 2224, 1638, 365, 1518, 1626, 563, 1813, 1769, 561, 1635, 475, 1467, 1037, 922, 1317, 2411, 543, 526, 962, 1721, 1685, 100, 107, 2852, 188, 208, 610, 815, 906, 367]. Notes [1915, 46, 343, 251, 382, 53, 498]. nothing [3310]. November [4, 32, 50]. Nuclear [2890]. Nucleation [266]. Nucleotide [2665]. nugget [3664]. Nuisance [1426, 1221, 1648, 2084, 1743, 1072]. Null [940, 1194, 3096, 2850, 918, 504, 3873, 3942, 3944]. Number [450, 125, 1096, 787, 1095, 1112, 1255, 2186, 1134, 942, 426, 822, 2842, 3063, 2361, 2418, 1716, 1232, 2588, 3075, 3067, 3257, 2419, 3232]. Numbers [2001, 984, 1060, 82, 2603, 286, 126, 13]. Numerical [731, 1054, 1092, 644, 315]. Nutrition [68].

Object [1995]. Objective [2985]. Objects [2799, 3572, 3656]. Observation [2255, 103, 803, 1547, 2519]. Observational [2927, 1592, 1979, 3216, 3306, 3639, 3663]. Observations [585, 1011, 125, 2337, 3091, 1156, 1003, 694, 1247, 2503, 2139, 2684, 3019, 902, 325, 2808, 2761, 361, 1729, 454, 1788, 3530, 3610, 3152, 3456]. Observed [2671, 2889, 2841, 2795, 3022, 2048, 1549, 3045, 2962, 2263, 3323, 3599]. Obtain [128]. Obtained [1011, 2164]. Occasions [201]. Occupancy [442]. Occurrence [876, 679, 810, 967]. Occurring [1001]. Occurs [2971]. Odds [1092, 652, 1648]. Oertzen [3835]. Offset [483]. Oil [2006]. Old [2345, 3222]. Oliver [3712, 3713, 3926]. Olkin [1925]. Ombao [3931]. Omission [2212, 75]. omitted [3561]. omnibus [3519, 3148]. On-Line [2738, 2956, 3071, 3176, 3175]. One [533, 752, 193, 2223, 345, 635, 2771, 304, 2758, 2494, 378, 379, 2234, 2152, 1552, 1173, 2695, 328, 380, 381, 1263, 361, 1547, 1978, 825, 1856, 1582, 415, 3367, 3280, 3597]. One-Parameter [2234]. One-Quadrant [635]. One-Sample [2152, 1263, 1856]. One-Sided [1173, 1582, 3367]. One-Step [2494, 1978, 3280]. One-Way [533, 1552, 2695]. online [3684, 3640]. Only [3020, 2488]. Ontario [74]. Opaque [1290]. Open [1069]. Open-Loop [1069]. Operating [104, 2653, 2524, 426, 822]. Operation [43, 99, 92]. Operations [1593]. Operative [304, 380, 381]. Operator [193]. Operators [2287]. Opinion [1858, 2011]. Optimal [533, 2364, 876, 2924, 3167, 3394, 3647, 529, 2485, 964, 858, 3264, 1465, 2027, 3229, 2243, 2234, 1639, 3622, 1280, 1329, 3453, 1552, 3425, 1287, 1548, 1399, 1452, 3466, 2936, 1668, 578, 2702, 1333, 3343, 3738, 2391, 1979, 3588, 3740, 2510, 1226, 3603, 2944, 1886, 2103, 1968, 1367, 1496, 1114, 1737, 2124, 3963, 3819, 2600, 3204, 3454, 3152, 3437, 3633, 3599, 3692, 3630, 3520, 3494, 3511, 3800, 3817, 3990, 3655]. Optimality [2323, 1963, 2201, 1446, 1730, 2315, 1868, 1447, 1657, 1958, 1793, 3282, 3544, 3762, 3634, 3681]. Optimization [573, 827, 926, 2388, 3504, 3411]. Optimizations [2775]. Optimum [2254, 210, 1096, 1095, 1112, 1255, 2362, 2363, 1565, 832, 460, 806, 551, 298, 2813, 1065, 929].

Oracally [3407]. Oracle [3007]. Orchard [74]. Ord [1434, 1510]. Ordaz [3713]. Order [1830, 2110, 2052, 896, 1869, 1825, 1249, 1336, 2750, 2021, 2533, 317, 2029, 2112, 1038, 632, 1896, 2630, 1396, 254, 1948, 1906, 2968, 2474, 637, 1390, 1209, 1463, 1705, 2037, 2892, 1443, 1741, 564, 1216, 1740, 1080, 2352, 707, 1076, 1262, 2925, 2129, 2904, 3578, 3249, 2499]. Order-Free [2968]. Order-Restricted [2892]. Ordered [1624, 343, 2938, 656, 337, 1082, 554, 1795, 1854, 2406, 2688, 522, 3207, 3554, 3728]. Ordering [449, 3107, 3811, 3818]. Ordinal [3207, 1437, 3418]. Ordinary [1919, 2054, 3360, 3654]. Ordinate [1911, 3070]. Organization [183]. Orientation [1295, 1765, 1881]. Oriented [635]. Oriented- [635]. Origin [1619, 1949]. Ornstein [2577, 2772]. Orthant [2991, 2693]. Orthodox [2958]. Orthogonal [2924, 1135, 430, 2686, 582, 811, 1603, 1583, 1756, 2073, 938, 1087, 1997, 2174, 255, 1091, 1831, 886, 1117, 750, 217, 2947, 76, 3240]. Orthogonality [1732, 1787, 1767, 2317, 1923]. Orthonormal [214]. Oscillation [306]. Oscillations [1278]. Oscillatory [137]. Other [2062, 2072, 18, 128]. out-of-distribution [3961]. out-of-sample [3597]. Out-Patient [250]. Outcome [2339, 2716, 2971, 2796, 1592, 3092, 2950, 3349, 3261]. Outcome-Dependent [2339, 2716, 2796, 3092]. Outcomes [2523, 2353, 3299, 3321, 3418]. Outlier [1587, 3255, 1561, 2657, 1611, 2100, 1748, 1415, 1685, 1671, 3695]. Outliers [1341, 1081, 2040, 2169, 2396, 1851, 3063, 2403, 3961]. Outline [321]. Output [145, 465, 287, 880, 3738]. Outputs [235]. Outs [2242]. Overdispersion [2443]. overfitted [3169]. overlapping [3579]. Overtaking [431]. Owen [3897]. Own [1873].

P [3672, 3830, 3932, 1439, 1582]. PA [1125]. PA-Type [1125]. PAB [1125]. PAB-Type [1125]. Pacchiardi [3863]. Pack [478]. PageRank [3563]. Pair [2947]. Paired [1444, 2722, 3754, 1266]. Paired-Comparison [1266]. Paired-Data [1444]. Pairs [1133, 1602, 1881]. Pairwise [2404, 2554]. Pallavi [3708]. Palm [1716]. Panel [2519, 3799]. Panning [3488]. Paper [366, 592, 760, 1348, 82, 53, 1606, 518, 2646, 2733, 2720, 2703, 2783, 3222, 3126, 3214, 2682]. Papers [664, 1850, 1904, 2256, 2259, 508, 2375, 2378]. Papp [3932]. Paradigm [2120]. Paradox [2392, 1981, 312, 1460]. Paradoxes [1123]. Parallel [955, 3523, 3688]. Parameter [1177, 2316, 2240, 1207, 2306, 1908, 2340, 1965, 1787, 2885, 1349, 843, 3094, 2363, 1223, 1275, 2234, 2023, 2276, 2395, 2317, 313, 1633, 2177, 975, 2066, 1398, 1949, 2043, 1790, 3045, 3002, 2966, 3066, 1276, 2165, 2352, 1649, 3446, 2696, 3075, 692, 597, 2527, 2044, 3178, 3708, 3706, 3170, 3709, 3710, 3711, 3702, 3704, 3712, 3236, 3280, 3707, 3713, 3714, 3243, 3715, 3787, 3716, 3147, 3717, 3718, 3719, 3720, 3705, 3721, 3722, 3725, 3723, 3724]. Parameter-Effects [1633]. Parameterization [2145, 2597, 2331]. Parameterizations [1633]. Parameterizing [3947, 3952, 3951, 3953, 3957, 3950, 3954, 3955, 3948, 3949, 3956]. Parameters [1319, 876, 1426, 2339, 347, 1619, 1150, 295, 1221, 517, 480, 612, 719, 356, 630, 1803, 807, 200, 742, 1586, 895, 984, 1060, 1997, 989, 1497, 2033, 2222, 1648, 2085, 274, 1209, 2084, 693, 1743, 997, 1458, 463, 928, 1002, 1740, 494, 1072, 808, 1115, 1520, 1474, 2294, 1693, 256, 1234, 3075, 2551, 1108, 3473, 3290, 3064, 3244, 3258, 3810, 3343, 3701, 3196, 3262, 3703]. Parametric [570, 1553, 1564, 2951, 1273, 2587, 3117, 1947, 2970, 2611, 2599, 1296, 2279, 2972, 3013, 2425, 3094, 2570, 1494, 2882, 3020, 3077, 1627, 2088, 1355, 3054, 2963, 2463, 828, 856, 918, 969, 1058, 2993, 2760, 1765, 1085, 1360, 1607, 1558, 3050, 1684, 2241, 1577, 2933, 1819, 3034, 3879, 3345, 3449, 3364, 3277, 3966, 3506, 3549, 3300, 3346, 3355, 3553, 3340, 3159, 3377, 3328, 3257, 3226, 3440, 3295, 3560, 3256, 3813, 3519, 3286, 3136, 3331, 3602, 3269, 3386, 3455, 3218, 3146, 3982]. parametrics [3517, 3258, 3408]. Parametrizations [1551]. Part [185, 378, 379, 20, 220, 262, 186]. Partial [676, 705, 2541, 3103, 1204, 960, 2319, 1871, 2553, 201, 1855, 1919, 2054, 3067, 3568, 3248, 3530, 3324, 3161, 3630, 3294]. Partially [2671, 1459, 2986, 2795, 3022, 1944, 3045, 1696, 2688, 2899, 3323, 3301, 3227, 3754, 3850]. Particle [2671, 3120, 2738, 3022, 3130]. Particular [1500, 1634, 1831, 1920, 34]. Partition [230, 2715]. Partitional [443]. Partitioning [960, 2917]. Partitions [345]. Partly [1768, 2711, 2044]. pass [1971]. Passage [788, 786, 737, 1493]. Past [1766]. Path [2960, 1975]. pathway [3942]. Patient [250]. Patients [166]. Patrolled [380, 381]. Patrolling [466]. Pattern [846, 2697, 1230, 2641, 3266, 3192, 3940]. Patterned [499]. Patterns [2697, 1573, 685, 1676, 1240, 1302, 1416, 2776, 3242]. Paul [3932]. PBIB [1459, 1625, 1696]. PCA [3730]. peaks [3750]. peaks-over-threshold [3750]. Pearson [366, 339]. Peculiar [2541]. peer [3945]. Peers [745]. pen [3162]. penalization [3452]. Penalized [3157, 2592, 2229, 1930, 3065, 2757, 2993, 2164, 1801, 3172, 2869, 3236, 3243]. Penalties [2527, 3643]. Peng [3704]. Penrose [2083]. People [787]. Percentage [1988]. Percentile [3852, 2735, 3546]. Percolation [635, 788]. Perfect [2668, 2598, 2496, 2452, 3189, 2658]. Performance [1699, 2987, 1757, 1110, 3883, 3489, 3760].

Performs

[3828, 3829, 3823, 3830, 3831, 3827, 3821, 3832, 3833, 3834, 3835, 3822, 3836, 3826, 3838, 3824, 3839, 3840, 3825, 3841, 3820, 3837]. Period [474, 727]. Periodic [1882, 2959, 2784, 927, 3149]. Periodically [2048]. periodicity [3277]. Periodogram [326, 639, 269, 371]. Perishable [104]. Permanent [1076]. Permutation [309, 1574, 1482, 1537, 3803, 3566, 3313]. Permutation-based [3803]. Permutations [846, 900, 3468]. Personal [1129]. personalized [3563]. Personnel [123]. Perspective [2497]. Perspectives [2503]. Perturbation [2704, 1713, 3355]. Perturbational [1400]. Perturbed [2247]. Peter [3828, 3671]. Pfeffermann [2375]. Phan [3900]. Phase [2339, 2930, 2831, 2377, 348, 3092, 3478, 3261]. Phase-Type [348]. Phenomena [1415]. Phenomenon [1001]. Philip [3899, 3900]. Phillips [3718]. Philological [184]. photon [3165]. photon-limited [3165]. Phylogenetic [2664, 3198]. phylogenies [3433]. Physics [176, 707]. Pictures [1775]. Piecewise [1663, 3759]. Piecewise-Deterministic [1663]. Pierre [3925]. Pierre-Aurelien [3925]. Pietro [3865]. Pilling [3836]. Pistimetric [475, 498]. Pitman [592]. Pivotal [312]. Place [21]. Plackett [608]. Plan [318]. Planar [2333]. Plane [722]. Planning [1228, 1183]. Plans [812, 849, 2692, 1208, 886, 1117, 1181]. Plant [857, 179]. Plant-Breeding [179]. Plants [35]. Plate [2082, 2685]. Plato [451]. Plausibility [1260]. Plot [1485]. Plots [2524, 233, 160]. Plotting [970]. Plug [3096]. Plug-in [3096]. Plummer [2259]. Point [2856, 1412, 621, 1869, 1249, 1289, 1306, 907, 491, 725, 1047, 2992, 1349, 1573, 2333, 136, 2923, 2988, 2860, 1049, 1034, 677, 576, 646, 637, 1538, 1676, 3247, 2579, 1416, 2753, 1317, 2428, 186, 1008, 2759, 2825, 100, 3076, 2776, 692, 1035, 2904, 3542, 3314, 3225, 3273, 3288, 3122, 3319, 3370, 3326, 3344, 3911, 3937, 3596, 3439, 3438, 2824, 3465, 3627, 3531]. Points [1988, 2973, 983, 2684, 3109, 1567, 693, 227, 624, 1114, 3542, 3602, 3373]. Poisons [248]. Poisson [1254, 1778, 1665, 401, 3290, 1090, 1143, 3094, 2913, 773, 783, 992, 2870, 2266, 1059, 685, 829, 1379, 1566, 736, 1220, 1538, 1986, 476, 784, 3149, 2537, 70, 1303, 1007, 3845]. Pole [1185]. Pole-Seeking [1185]. Policy [3959]. Pólya [3564]. Polycross [1829]. Polygons [2415]. Polyhedra [777]. Polymorphism [2665]. Polynomial [1739, 2386, 2549, 2531, 2112, 2217, 1218, 1424, 894, 921, 2473, 3105, 1796, 1835, 1800, 1885, 2739, 2627, 2876, 709, 3336, 2749, 3202]. Polynomial-Projecting [1218, 1424]. Polynomials [430, 214]. Polytomous [2505]. Pooled [3093]. Pooling [2024]. Poor [280]. popularity [3629, 3479]. Population [1254, 194, 487, 116, 2997, 199, 8, 472, 945, 863, 130, 1471, 178, 204, 1099, 2086, 1175, 2665, 288, 2257, 375, 1647, 244, 463, 1359, 1740, 494, 180, 825, 2545, 208, 3687, 3567, 3258, 3422, 3533, 3742]. Populations [923, 239, 1564, 2269, 1905, 2861, 934, 356, 1688, 333, 771, 772, 1283, 622, 903, 2580, 2030, 1141, 388, 453, 55, 62, 535, 1547, 169, 949, 1406, 722, 993]. Portland [11]. Portmanteau [1509, 1888, 3356]. Position [224, 257, 1131]. Positive [1194, 1733, 2103, 3202]. Possible [2862, 3344]. Possibly [1140, 2396, 825, 1776, 3986, 3776]. Post [811, 3538, 2971, 2448, 2263, 1722, 3815, 3230, 3349]. post-outcome [3349]. Post-Randomization [2971]. Post-selection [3538, 3815]. Post-Stratification [1722, 3230]. Post-Synaptic [2448]. Posterior [1943, 3857, 3862, 2635, 1392, 1251, 3859, 3854, 1433, 945, 1014, 3870, 630, 652, 2247, 1233, 2272, 3861, 884, 1395, 1544, 3855, 2695, 3858, 2185, 2224, 3856, 2043, 3868, 3865, 3866, 1105, 1805, 671, 3867, 1806, 916, 2087, 3869, 3864, 3139, 3860, 3853, 3863, 3169, 1045, 1116]. Posteriori [2228, 1884, 495, 545]. Posteriors [3012, 2627]. Pot [857]. Potential [2736, 3321]. Power [626, 1970, 2755, 3012, 1109, 614, 2149, 674, 496, 868, 1896, 1413, 856, 1037, 275, 245, 1664, 1772, 3639, 3933]. powered [3648]. Powerful [2213, 2640]. Powers [316, 516, 357, 407, 882, 905, 3229]. Practical [2213, 191, 213, 424, 502, 2037, 3002]. Practice [2815]. Preadjusted [3269]. precise [3785, 3678]. Precision [1636, 1045, 1014, 1116, 1804]. Predicting [2313, 304]. Prediction [847, 664, 2323, 1239, 2106, 2614, 2743, 2416, 2790, 1610, 540, 1523, 2083, 944, 2130, 3695, 2607, 2882, 3107, 3681, 286, 2949, 1333, 2007, 3882, 1458, 1635, 475, 2428, 1752, 1919, 2054, 1615, 1791, 2045, 2192, 658, 3883, 2916, 3489, 3749, 3578, 3256, 3756, 3376, 3343, 3597, 3239, 3984]. Prediction-based [2323]. Predictions [1169, 1256, 665, 3229]. Predictive [3025, 1747, 2278, 2210, 2086, 1911, 1655, 3908, 3939, 3193]. Predictor [920, 1158, 1291, 1487, 879, 1424, 2958, 3600]. Predictors [1802, 1218, 332, 2648, 1766, 738, 3067, 3602, 3188, 709]. Predictors/Controllers [709]. Preface [3126]. Preliminary [825]. prepivoting [3687]. preprocessing [3753]. Prequential [1942, 2143]. Prescribed [1960, 1990]. Presence [1426, 1150, 2955, 2076, 1221, 2722, 680, 2473, 1883, 2762, 2084, 1743, 659, 3880, 3166, 3943]. Present [1030]. Presentation [298, 3223]. Preservation [1730]. President [4]. Pretesting [1252]. pretreatment [3216]. prevalence [3453]. Prevention [2651]. Previsions [1491]. Price [326]. primary [3147]. Prime [2186]. Principal [3420, 2874, 3370, 1679, 3632, 2287, 2241, 1919, 2054, 2477, 2869, 3523, 3240, 3573, 3307, 3751, 3171, 3388, 3772]. Principle [2110, 1893, 1878, 2465]. Principles [134, 844, 33, 3259]. Prior [1425, 1778, 1580, 3117, 2450, 694, 482, 1433, 3094, 2571, 355, 2649, 884, 745, 3112, 1346, 2185, 2224, 1769, 3636, 2751, 2324, 1563, 1671, 2299, 3139, 3302, 3876]. prior-free [3302]. prior-likelihood [3636]. Priori [557, 3466]. Priorities [552]. Priority [2357, 2358, 645, 2431, 553, 972, 2825, 798, 2824]. Priors [3032, 1601, 1431, 2491, 2181, 1460, 1372, 2511, 3734, 3137, 3551, 3295, 3778, 2528]. privacy [3668, 3670, 3669, 3672, 3673, 3677, 3674, 3671, 3675, 3676]. Priyantha [3676, 3869]. Probabilistic [3921, 3938, 3922, 3923, 3924, 3920, 3925, 2937, 795, 2641, 3926, 3927, 3928, 3929, 3937, 3916, 3919, 3930, 3914, 2794, 3931, 3932, 3933, 3242, 3934, 3935, 3409, 3200, 2477, 3918, 2011, 3913, 3936, 3917, 3915, 3302, 3153, 3388]. Probabilities [1934, 118, 2879, 1074, 2991, 1574, 598, 1129, 1334, 1033, 277, 557, 554, 645, 943, 1481, 1482, 264, 777, 167, 1838, 2693, 2374, 642, 1232, 1841, 1321, 2488]. Probability [1530, 3852, 2635, 2817, 2898, 635, 1689, 697, 1523, 1033, 2373, 944, 385, 99, 1853, 2070, 2468, 500, 970, 1611, 1326, 297, 1639, 2126, 1238, 1155, 753, 1680, 673, 693, 3106, 587, 128, 1547, 184, 475, 565, 1232, 161, 1714, 428, 227, 880, 2081, 2862, 422, 276, 3437, 3779, 3577, 3546, 3261]. probability-dependent [3261]. Probit [118, 2957]. Problem [704, 541, 488, 489, 1207, 1999, 1336, 1293, 994, 347, 1073, 2787, 451, 449, 2881, 1082, 2109, 1261, 1328, 514, 33, 478, 574, 690, 839, 828, 856, 918, 477, 323, 1493, 1441, 1972, 1268, 715, 747, 216, 922, 402, 948, 977, 1110, 1372, 1788, 34, 127, 2385, 3217, 3511, 3353, 246]. Problems [664, 165, 3990, 310, 920, 352, 2191, 1794, 2198, 296, 54, 115, 73, 858, 684, 1140, 1320, 2956, 293, 1382, 2027, 385, 1858, 803, 1803, 92, 2597, 223, 386, 1544, 2464, 1066, 2872, 990, 891, 1727, 440, 25, 399, 10, 53, 2927, 157, 561, 579, 184, 1920, 1322, 2181, 778, 508, 550, 2103, 1728, 3072, 3695, 3815, 3504, 3779, 3740, 3188, 3876, 3349, 3327, 2788]. Procedure [940, 892, 2773, 3003, 1971, 2048, 2653, 970, 1307, 3090, 908, 622, 628, 823, 1253, 1039, 388, 1950, 2989, 587, 586, 1343, 1110, 2944, 1474, 1776, 1613, 3601, 3493, 3218]. Procedures [481, 3990, 1734, 1465, 1833, 1472, 1044, 2626, 1452, 1632, 1324, 1464, 1174, 913, 551, 1978, 1213, 778, 692, 3449, 3354, 3769, 3184, 3740, 3197]. Proceedings [111]. Process [265, 3025, 606, 1869, 1397, 635, 861, 1249, 405, 2274, 724, 725, 983, 1090, 3094, 2913, 1628, 465, 2048, 605, 391, 708, 760, 2245, 363, 774, 190, 204, 2354, 1046, 1439, 609, 1521, 685, 677, 1379, 3043, 370, 497, 1398, 576, 646, 637, 360, 1566, 274, 736, 2140, 2305, 475, 1805, 864, 978, 436, 2449, 1007, 3972, 3973, 3163, 3974, 3849, 3981, 3968, 3975, 3419, 3976, 3969, 3138, 3326, 3977, 3744, 3247, 3970, 3971, 3664, 3979, 3980, 3978]. Processes [847, 1778, 194, 2856, 283, 459, 177, 621, 1553, 2889, 1825, 2841, 487, 2614, 2743, 1289, 1306, 907, 2119, 115, 1047, 503, 1663, 2992, 1349, 1573, 863, 512, 641, 749, 1393, 2923, 2988, 1910, 3019, 282, 992, 1049, 1395, 2088, 348, 862, 178, 782, 898, 1796, 1835, 1059, 1130, 1953, 1346, 829, 328, 1567, 853, 1668, 1187, 840, 1099, 737, 881, 444, 534, 219, 176, 2518, 1220, 1538, 1297, 2579, 1986, 569, 589, 714, 762, 1121, 218, 2772, 2508, 530, 2753, 1317, 266, 1998, 1806, 1814, 508, 2859, 2852, 2911, 1844, 3076, 1035, 390, 738, 2904, 3433]. processes [3423, 3393, 3506, 3225, 3246, 3691, 3304, 3130, 3229, 3288, 3469, 3784, 3806, 3122, 3319, 3548, 3305, 3651, 3733, 3401, 3540, 3190, 3426, 3149, 3239, 3906, 3549]. Procrustes [2333, 1957, 2328, 1713, 1358]. Product [1749, 1124, 566, 655, 2082, 2163, 2715, 34, 3752, 3808]. Product-Type [1124]. Production [304, 131, 235, 5, 19]. Productivity [193, 285, 212, 251]. Products [2322, 170, 235, 842, 60]. Professor [4, 1606]. Profile [2144, 2092, 2170, 2262, 2065, 1921, 2352]. profiles [3699]. Programmes [529]. Programming [321, 1935, 661, 711, 3429, 3756, 3588]. Progressive [658]. Projected [2691, 3966]. Projecting [1218, 1424, 709]. Projection [2595, 3431, 3574, 3494, 3360, 3465, 3654]. projections [3527, 3573]. Projectors [1190]. Proneness [757, 798]. Proof [1646, 1071]. Propagation [2199, 3471]. propensity [3263]. Proper [1425]. Properties [995, 1023, 1011, 113, 1158, 1291, 1487, 1089, 1293, 2541, 588, 2149, 675, 1366, 2349, 812, 849, 957, 1055, 2221, 2839, 2874, 1342, 1356, 1572, 273, 677, 1423, 637, 737, 1723, 289, 702, 901, 1626, 709, 1458, 1954, 269, 544, 137, 117, 698, 1791, 245, 3601, 3651, 3318, 3500]. Property [1254, 2481, 1639, 633, 448, 1510, 706, 1546, 653, 3846]. Proportion [166, 3096, 85, 3007, 2850, 2086, 244, 3436, 3468, 3985]. Proportional [2725, 2955, 2609, 1033, 2676, 2001, 297, 2711, 1648, 1680, 428, 2547, 276, 3447]. Proportions [755, 816, 920, 1401, 2001, 2063, 1494, 1673, 1997, 1577]. Proposal [2681, 3694]. Proposed [492, 675, 613]. Proposer [3669, 3702, 3854, 3886, 3821, 3948, 3913]. Proposers [3969]. Prospective [2716, 2610]. Provable [3427]. Provide [1832]. Provincial [90]. Proximal [3804, 3624]. Prünster [3972]. Prünster’s [3862]. Pseudo [636, 3132]. Pseudo-Distribution [636]. pseudo-empirical-likelihood [3132].

Pseudodata [2291]. Pseudolikelihood [2308, 2554, 2010, 3324]. pseudomarginal [3502]. Pseudomartingale [3043]. Pseudosplines [2277]. Psi [675]. Psychologist [186]. Psychology [352, 198]. public [3282]. Publication [2637]. Punched [119, 83]. pursuit [3208]. Pyramids [3794].

q [1439, 2174]. Qing [3978]. Quadrant [635]. Quadratic [560, 1565, 631, 797, 1304, 300, 1969, 2984, 549, 1467, 796, 2527, 661, 711]. Quadratic-Risk-Based [2984]. Quadratics [632]. Qualitative [1887, 1326]. Quality [74, 85, 2239, 1493, 81, 2188, 19, 102]. Quantal [349, 429, 94, 578, 248, 594]. Quantification [2890, 2698, 3843]. Quantifying [1858, 2192]. Quantile [3457, 3013, 1852, 2410, 3105, 3298, 3842, 3180, 3652, 3541, 3877, 3528, 3404, 3759, 3492, 3683, 3482, 3642, 3509, 3507]. Quantiles [171, 1128, 943, 101, 789, 2474, 1359, 544, 542, 906, 1664, 1772, 3873, 3358, 3332]. Quantitative [1534, 1605, 2662, 2186, 3789]. Quantities [312]. Quantum [2819, 2732]. Quartic [2912]. Quasi [1658, 2844, 2706, 2269, 1827, 2494, 3601, 1517, 205, 3319, 2273, 2346, 2010, 3605, 768, 1343, 3851, 3157, 3315]. Quasi-Bayes [3601, 1343]. Quasi-Complete [1658, 1517]. Quasi-Latin [205]. Quasi-Likelihood [2706, 1827, 2494, 2273, 2010, 3319, 3157]. Quasi-Linear [2269]. Quasi-Model-Assisted [2844]. Quasi-Newton [2346, 3851]. Quasi-Stationary [768, 3605]. quasiconvex [3946]. Quasisymmetric [1924]. Quenouille [189, 865, 518]. Questions [20]. Queue [305, 330, 395, 410, 411, 1500, 413, 490, 1853, 553, 774, 574, 690, 758, 412, 866, 476].

Queueing [283, 536, 347, 474, 491, 512, 641, 1500, 465, 645, 1312, 348, 363, 514, 758, 477, 509, 644, 784]. Queues [250, 1222, 329, 362, 445, 575, 955, 643, 223, 1046, 577, 971, 706, 735, 972, 1111, 461, 642, 433]. Quick [3280].

R [3825, 3891, 1213, 2606]. R. [1096, 1112]. Races [1482]. Rachael [3718]. Radius [1762]. Radon [703, 704]. Rainfall [2274, 679, 810]. Raised [53]. Rajendra [3951]. Ramdas [3889, 3967, 3890, 3886, 3891, 3892, 3893, 3894, 3895, 3887, 3896, 3897, 3898, 3899, 3900, 3901, 3888]. Ramifications [419]. Ramses [3858]. Random [748, 703, 739, 1319, 802, 343, 212, 251, 3503, 2338, 1202, 3431, 2706, 2320, 2314, 1832, 115, 1047, 2232, 2861, 1013, 3130, 985, 2423, 1674, 1122, 681, 942, 1278, 348, 82, 204, 1200, 689, 2218, 3485, 547, 2695, 842, 706, 287, 866, 3000, 135, 220, 262, 3992, 1493, 3014, 900, 1816, 2583, 2802, 1216, 1002, 1359, 634, 543, 372, 1502, 1142, 2555, 100, 83, 2471, 3178, 3889, 3967, 3568, 3424, 3444, 3890, 3572, 3656, 3573, 3332, 3235, 3408, 3886, 3122, 3891, 3892, 3893, 3894, 3987, 3161, 3895, 3887, 3896, 3897, 3438, 3752, 3318, 3898, 3899, 3900, 3579, 3901, 3888]. random [3902, 3885]. Random-Effects [2695, 3014]. random-field [3438]. random-graph [3178]. random-measure [3235]. Random-projection [3431]. Random-walk [3503]. Random-weight [3130]. Randomisation [3873]. randomised [3879, 3844, 3801]. Randomization [626, 1658, 2223, 541, 2364, 792, 3363, 205, 2971, 2074, 1136, 2263, 586, 1496, 3634, 3650, 3682, 3609, 3988]. Randomizations [2902]. randomize [3634]. Randomized [1941, 1513, 2980, 2609, 2147, 2897, 1461, 2073, 2542, 2141, 160, 1380, 454, 2862, 1545, 2950, 3570, 3230, 3588, 3593, 3662]. Randomly [1003, 3054, 3646, 3464]. Randomnes [1416]. Randomness [344, 358, 1379]. Randon [739]. Range [2039, 172, 531, 313, 21, 169]. Ranges [544]. Rank [1617, 686, 896, 1336, 1812, 3449, 2990, 199, 1082, 236, 1427, 1560, 1554, 1604, 1478, 986, 674, 675, 2524, 1163, 2151, 1630, 1019, 1972, 1381, 1499, 1164, 3186, 3877, 3522, 3598]. Rank-based [3449]. Rank-Order [896]. Ranked [1177, 2634]. Ranking [1782, 2649, 742, 206, 197, 778, 198, 3525, 3339, 3368]. Rankings [878, 3358]. Ranks [3001, 1550, 1602, 1790]. Rate [2440, 2213, 3096, 2653, 1756, 2550, 2203, 840, 2989, 1954, 3441, 3127, 3941, 3553, 3354, 3224, 3769, 3641, 3184, 3905, 3729, 3509, 3209]. Rates [2323, 2767, 2663, 2072, 741, 2957, 2652, 2759, 2825, 2824, 3511]. Ratio [716, 626, 1675, 611, 1932, 1988, 2566, 2804, 1620, 2758, 294, 317, 2078, 1092, 45, 500, 2049, 807, 1894, 2774, 1304, 689, 1100, 2308, 2002, 833, 565, 2031, 1232, 1072, 2355, 1698, 2640, 3478, 3373]. Ratio-Statistics [317]. Rational [243]. Ratios [1005, 511, 1202, 2295, 655, 3278, 3432]. Ray [592, 3868]. Rayleigh [3511]. Re [1038, 1708]. Re-Order [1038]. Re-Weighted [1708]. Reactions [266]. read [3126]. Real [2805, 520, 3801]. real-world [3801]. Realibility [1476]. Reality [324]. Realization [190]. Realized [2639, 2610]. reasonable [2427]. Rebaudo [3864, 3974]. Recapture [2599, 2463, 3508]. Receiver [2524]. Reciprocal [1722]. Recognition [2641]. Recombination [2663]. Reconstruction [2991]. Record [253]. Records [279, 316, 1752, 377]. Recovery [2168, 2306, 1767]. Rectangles [751]. Rectangular [2691, 1125, 463]. Rectifying [129, 318]. Recurrence [1515]. Recurrent [915, 2550, 2762, 2457, 372, 3743]. Recursive [1595, 1373, 1879, 1948, 2540, 2150, 1383, 1712, 1649, 738]. Red [2379, 74]. Reduce [2189]. Reduced [3186, 314, 315, 1499, 3522]. Reduced-Rank [1499]. Reducing [2258, 2480, 2823, 3903]. Reduction [3103, 1893, 2873, 2287, 3115, 2647, 2943, 3486, 3122, 3377, 3219, 3147, 3686, 3732, 2633, 3327]. Redundancy [2306]. Reference [1235, 250, 1392, 1942, 18, 1634, 1920, 34]. refitted [3179]. reflected [3379]. Regarding [841, 161]. regeneration [3278]. Regenerative [1001]. regime [3807]. Regimes [2702, 3795, 3454, 3437, 3800]. Region [687, 716, 2485, 909, 1545, 462, 3412, 3808]. Regional [2119]. Regionalization [3423]. Regionalized [2968]. Regions [664, 473, 2533, 2905, 2247, 2341, 2642, 389, 2351, 2899, 3297, 3473, 3404, 3958, 3181, 3411, 3907, 2489]. Registration [2399]. Regression [1624, 1528, 2099, 2805, 920, 1650, 1308, 1102, 3044, 2301, 3056, 2313, 2339, 1215, 2587, 1475, 2541, 2986, 3322, 2111, 2750, 2970, 2041, 2930, 1402, 2093, 2450, 2532, 2625, 3116, 2050, 2386, 2549, 2706, 3103, 1296, 1451, 75, 1610, 1802, 1839, 3405, 418, 455, 482, 1067, 2046, 1579, 2424, 2972, 2531, 2932, 480, 2321, 1143, 1320, 1351, 2067, 1272, 2699, 1131, 824, 2242, 1565, 1858, 811, 2438, 1043, 1178, 2379, 2211, 855, 1803, 1863, 2276, 2708, 625, 669, 1134, 1843, 1126, 1946, 2032, 2461, 2669, 2204, 2565, 1868, 1692, 2569, 2979, 2623, 2395, 2317, 2473, 2319, 2597, 2999, 1766, 314]. Regression [628, 954, 3105, 1783, 2770, 1691, 2926, 939, 2464, 2205, 1066, 3036, 2490, 1994, 2554, 2463, 2550, 146, 854, 1385, 1962, 2466, 1969, 2065, 2380, 2176, 2222, 2066, 1437, 2085, 1324, 2982, 3104, 1638, 2814, 2151, 1466, 2230, 2650, 2993, 1522, 2236, 2202, 1790, 1640, 1020, 1881, 2876, 2290, 947, 1729, 2957, 2031, 3158, 315, 2638, 1754, 1684, 2397, 808, 1919, 2054, 377, 2519, 2107, 1162, 1488, 672, 2267, 3154, 479, 203, 1499, 2045, 1968, 1372, 2925, 1671, 467, 1776, 2710, 359, 3072, 2629, 2685, 1698, 1029, 2430, 2616, 2486, 2871, 2943, 2954, 2546, 3034, 2748, 3312, 3665, 3234, 3783, 3734]. regression [3575, 3186, 3652, 3506, 3549, 3248, 3300, 2427, 3541, 3064, 3179, 3396, 3877, 2816, 3847, 3661, 3152, 3560, 3787, 3790, 3768, 3344, 3256, 3164, 3298, 3569, 3681, 3492, 3512, 3329, 3744, 3946, 3421, 3190, 3602, 3342, 3626, 3241, 3520, 3403, 3522, 3330, 2749, 3683, 3386, 3455, 3482, 3608, 3221, 3587, 3808, 3218, 2633, 3800, 3202, 3791, 3878, 3531, 3276, 3507, 3271, 3308, 3274]. Regressions [2587, 1148, 2873, 2007, 215, 3480, 3320, 3610]. regressive [3424, 3288, 3407, 3446, 3906, 3356]. Regressors [2654, 3847]. Regret [622]. Regular [1794, 2198, 2115, 1529, 1447, 2206, 547, 2857, 3312]. Regularization [2864, 2757, 2960, 2865, 2834, 3375, 3170, 3683, 2785]. Regularized [2890, 3072, 3271, 3204, 3504, 3190, 3232, 3371]. Regulation [926, 839]. Reinforced [1576, 1657, 2175]. Reinforcement [2963, 3727]. Reiss [3304]. reject [3773]. Rejecting [34]. Rejection [1415]. Related [540, 1140, 2478, 1404, 701, 312, 2793, 1323, 2694, 1600, 2173, 2059, 973, 1021, 2471, 3245, 3311, 3564, 3904]. Relation [192, 474, 517, 1692, 67, 340, 1174, 324, 2140]. Relations [801, 914, 154, 1233, 762, 169, 187, 203, 738]. Relationship [1405, 1739, 1073, 341, 1704, 146, 1980, 948, 1008]. Relationships [1245, 2099, 2338, 1215, 696, 2148, 1086, 1433, 867, 1522, 1279, 439, 769, 665, 359]. Relative [1421, 1444, 1131, 584, 616, 986, 2028, 2655, 2987, 975, 1179, 3434, 3201, 3455]. Relatively [874]. Relevance [704, 2698]. Relevant [1964, 2032, 3285, 3352, 3584]. Releveraging [1967]. Reliability [785, 1006, 1062, 1998, 2616]. Reliable [1987]. Remark [1418, 1618, 905, 864].

Remarks [1348, 613]. Remnants [446]. Remote [2786]. Removal [2599, 2005, 421, 2463]. Removals [605]. Removing [2291]. Renewable [3562]. Renewal [638, 606, 449, 809, 3043, 1187, 534, 419, 526]. Repair [785, 380, 381]. Repairmen [396, 1716]. reparametrization [3618]. Repeat [1757]. Repeated [1106, 2338, 1731, 2568, 1964, 2557, 1891, 1450, 2872, 2263, 3299, 3345]. Replace [2756]. Replaced [1015]. Replacement [1530, 2898, 104, 201, 587, 180, 428, 276]. Replicate [826, 1206, 1340, 1205, 1594]. Replicated [2776, 387]. Replicates [1983]. Replication [1405, 1339, 1704, 3731]. Replications [876, 1211, 1282]. Reply [518, 3593, 3938, 3981, 3677, 3957, 3870, 3937, 3841, 3725, 3902]. Report [1862, 1901, 1939, 1993, 2058, 2135, 2197, 2252, 3293, 3335, 3042, 3102, 3144, 3177, 2564, 2621, 2680, 2747, 3616, 3385, 3462, 3516, 3559, 2436, 2502, 2812, 2868, 2312, 2372, 2921, 2978, 3212, 3252]. Reporting [1129]. Represent [240]. Representation [2357, 2358, 2318, 1628, 1424, 2843, 992, 2490, 3495]. Representations [2184, 3358]. Representative [28]. reproducing [3274]. repulsive [3426]. Requiring [212, 251]. Rerandomization [3569, 3633]. Resampling [2828]. Research [4, 3, 80, 111, 59, 73, 2258, 114, 5, 105, 1903, 6, 60]. Reservoir [603]. Reservoirs [1180]. Residual [2856, 1366, 1702, 3036, 1505, 2176, 1364, 1705, 1020, 2638, 2292, 2094, 2355, 2904, 3491, 3644]. Residuals [3225, 873, 1373, 2387, 640, 2319, 919, 439, 2710]. Resistant [1611, 1643, 2120]. Resistivity [2721]. Resnick [3304]. Resolution [1084, 1414, 963, 1208, 3409]. Resolvability [447]. Resolvable [1721, 1282, 1266]. Resolving [1301]. Resonance [2890]. Respect [2244, 2315, 3686]. Response [1703, 2316, 883, 2338, 2068, 2611, 2771, 1618, 2980, 2376, 2028, 2248, 2897, 94, 1101, 2464, 2141, 578, 1727, 2927, 1380, 1516, 1819, 2791, 594, 3434, 3798, 3448, 3570, 3639, 3644, 3151, 3813, 3686, 3593, 3878]. Response-Biased [2611]. Response-Selective [2464]. Responses [1429, 2313, 2800, 349, 1833, 2242, 669, 813, 3054, 248, 3213]. Rest [2091]. Restoration [2228, 1779, 2442, 2522]. Restraints [231]. Restricted [570, 528, 1192, 2804, 1223, 205, 3001, 200, 1264, 2074, 2892, 1685, 1496, 3691, 3403, 3145]. Restrictions [627, 579, 1740, 495, 3391]. Result [704, 991]. Resulting [69]. Results [395, 1048, 660, 580, 277, 1272, 1517, 1092, 2476, 783, 809, 363, 3065, 2008, 1524, 1113, 1017, 644, 476, 1645, 790, 734, 298, 550, 1035, 728, 1065, 3282]. retransformation [2382]. Retrieval [2787]. Retrospective [3126, 3154]. return [3222]. Reuse [1947, 2303]. Reversal [2741]. Reversible [2681, 2724, 2504, 3375, 3183, 3688]. Reversion [768]. Review [225, 1849, 1898, 1362, 3621, 102]. Revised [306]. Revising [1491]. Revisited [1806, 3843]. Reweighted [1643]. Richard [3950]. Richards [613]. Richardson [2646, 3719, 3955]. Ridge [1272, 1178, 1423, 1678, 2107, 1162, 1488, 3791]. Ridge-Type [1178]. Riemann [3150]. Riemannian [3621, 3405]. Right [2440, 3574, 3786, 3136]. Right-Censored [2440, 3786, 3136]. Rigo [3865]. Rinaldo [3979]. Ring [846, 2168]. Risk [1734, 837, 584, 2028, 2984, 2673, 2823, 1088, 2616, 3475, 3746, 3455]. Risk-Reducing [2823]. Risks [2725, 3116, 2487, 616, 132, 2979, 3185, 3298, 3410]. Ritabrata [3863]. Road [509, 431, 508, 3204]. Robbins [695, 3624]. Roberts [2682]. Robins [3955]. Robust [309, 3205, 1428, 2340, 3023, 2373, 3535, 3567, 3057, 1643, 1627, 3847, 3592, 3521, 1413, 2580, 3735, 2150, 1464, 1944, 2595, 3583, 2705, 3451, 2121, 2356, 3221, 1495, 3084, 2628, 3609, 3369, 3255, 3751, 3440, 3227, 3309, 3580, 3487, 3577, 3984, 3988]. Robustifying [2235]. Robustness [1963, 2111, 1074, 2441, 3155, 2626, 1766, 3151, 1358, 1400, 3872, 3694, 2636]. Rock [1660]. Rockware [89]. Rodionov [3890]. Rohe [3828, 3829, 3823, 3830, 3831, 3825, 3832, 3833, 3834, 3835, 3836, 3826, 3838, 3824, 3839, 3840, 3827, 3821, 3822, 3841, 3837]. Role [1036]. Rolling [827]. Romano [1850]. Ronald [366, 365, 367]. Rong [3892]. Room [413]. Root [36, 3968, 1351, 3611, 3972, 3973, 3974, 3981, 3975, 3976, 3969, 3977, 3970, 3971, 3979, 3980, 3978]. Roots [153, 2078, 499]. Rossell [3866]. Rotatability [1667, 2008]. Rotatable [1148]. Rotation [648, 3325]. Roughness [2437]. Rounding [1579]. Routine [342, 106, 387]. Row [1983, 1465, 1767, 2112, 1211, 1597, 2540, 1399, 3203, 3529]. Row- [2540]. Row-and-Column [1983, 1465, 2112, 1211, 1399]. Row-Column [1767, 1597, 3529]. Rows [968, 2112, 84]. Royal [3654, 3, 80, 111]. Rubin [1813]. Rubinstein [3198]. Ruf [3893]. Rule [2555, 3963]. Rules [2130, 957, 1055, 741, 2030, 715, 2181, 34, 3749, 3692, 3190, 3630, 3188]. Ruls [747]. Rumour [787]. Run [515, 823, 3560]. Rungang [3825]. Runs [1342]. Ruodu [3967, 3888]. Ryan [3895].

S [3719, 3955, 3900]. Saddle [1412, 1008]. Saddle-Point [1412]. Saddlepoint [2069, 2624, 1971, 1945, 2474, 1537, 2161, 1841]. Safe [3105]. Saleman [323]. Sam [3933]. Sample [1093, 570, 802, 1128, 1336, 1293, 1947, 2549, 296, 2513, 2674, 1082, 1427, 1560, 1604, 130, 500, 674, 675, 1307, 1363, 2303, 2843, 2987, 71, 2152, 1133, 2245, 1586, 426, 822, 823, 1510, 1188, 2390, 2474, 2634, 828, 856, 918, 969, 1058, 1880, 1263, 1958, 1972, 1381, 1626, 1473, 1274, 790, 544, 1232, 1338, 275, 1856, 1791, 1248, 83, 188, 906, 1788, 2071, 2403, 181, 2616, 2289, 2558, 3539, 3267, 3525, 3849, 3290, 3217, 3660, 3779, 3597, 3636, 3409, 3395, 3760, 3754, 3627, 3910, 3369]. Sampled [2145, 1910, 2594, 2508, 256]. Sampler [2062, 2913, 2072, 2203, 2167, 2331, 2479, 2059, 3874]. Samplers [2724, 2668, 2891, 2598, 2658, 3734, 3183, 3744]. Samples [838, 885, 930, 1564, 1689, 2716, 2314, 8, 1012, 45, 2169, 1015, 101, 1154, 21, 1851, 1626, 55, 62, 551, 463, 494, 180, 825, 70, 1422, 3198, 3754, 3577]. Sampling [1530, 292, 2567, 56, 113, 2817, 2898, 2339, 225, 116, 1074, 158, 588, 2930, 2532, 2611, 2625, 2377, 501, 1523, 1832, 2458, 1290, 2166, 1033, 1013, 2821, 277, 1090, 934, 2373, 2478, 1044, 316, 985, 2571, 2061, 226, 333, 771, 772, 936, 1043, 132, 18, 297, 812, 849, 957, 1055, 1593, 832, 929, 1690, 622, 903, 82, 2896, 1017, 1614, 388, 3017, 12, 2086, 1680, 2025, 1518, 201, 290, 10, 950, 1179, 587, 700, 2583, 1213, 1761, 2510, 718, 1359, 1050, 1753, 428, 298, 2555, 1181, 479, 2120, 281, 126, 1791, 904, 208, 3092, 468, 781, 2347, 127]. sampling [3350, 3398, 3590, 2612, 3563, 3415, 3749, 3815, 3374, 3136, 3532, 3342, 3497, 3742, 3261, 1103]. sandwich [3238]. Sarkar [3828]. Sarvet [3705]. Saving [1593]. Savings [1311]. Sayan [3973]. Scalable [2770, 3536, 3780, 3283, 3597, 3653, 3688]. Scalar [2885]. Scale [1161, 43, 1370, 85, 463, 3061, 1520, 3525, 2816, 3368, 3237, 3182, 3296, 3851, 3545, 3605]. Scale-Free [1370]. Scaling [1637, 2831, 2391, 1358, 1400, 3807]. scan [3192, 3760, 3373]. scanning [3409]. Scans [2203]. Scattered [2808, 2761]. Scatterometer [2787]. sceptical [3731]. Scheduled [477, 866, 1111]. Scheduling [1250, 548, 1367, 1728]. Scheffé [805]. Scheme [3011, 776, 823, 1531, 904, 468, 3688, 3261]. Scheme-Generalization [904]. Schemes [189, 1292, 588, 1625, 295, 254, 578, 2384, 1179, 2331, 117, 188, 245, 1496, 2289]. Schlaflian [1667]. Schmidt [3934]. Schmidt-Hieber [3934]. Schoenberg [3935]. Schwarz [1408]. Scientific [42, 1129, 311]. Score [2320, 2108, 2980, 2691, 1922, 3592, 3263]. Scores [1293, 656, 1058, 3420, 3877]. Scoring [1634, 3358]. Screening [1750, 1999, 3028, 749, 1966, 891, 1886, 3394, 3525, 3224, 3360, 3654]. Scribe [402]. Search [2985, 1465, 1399, 1452]. Searching [1096, 1095, 1112, 1255, 3815]. Seasonal [1307, 1506, 1562]. Sebastian [3673]. Second [2052, 1869, 1249, 824, 2379, 254, 82, 2352, 2129, 2904, 3578, 3249, 2499]. Second-Derivative [2379]. Second-Order [2052, 1869, 1249, 2352, 2129, 2904, 3578, 3249, 2499]. Secondary [2551, 3492, 3341, 3221]. Seconder [3670, 3703, 3855, 3914, 3887, 3822, 3970, 3949]. Section [2837, 4, 3, 80, 111, 50]. Sectional [1433]. Sections [325]. Secure [2606]. Seeking [1185]. Seemingly [2007]. segmentation [3314]. Segmenting [3763]. segments [3205]. Selected [277, 1647, 3617, 3196]. Selecting [472, 2581, 3107, 1422, 3523, 3606, 3649]. Selection [2323, 2864, 2697, 1689, 3056, 2662, 2712, 2416, 2654, 2450, 3103, 2773, 529, 1033, 2531, 2217, 2769, 2023, 2669, 3049, 2395, 2999, 2692, 2210, 3036, 2833, 1324, 2949, 2374, 2876, 579, 2984, 3066, 1655, 1987, 2638, 3031, 1408, 2267, 2472, 2911, 123, 2925, 3051, 3075, 2629, 3008, 2486, 276, 2871, 2834, 3157, 3488, 3199, 3236, 3354, 3815, 3538, 3368, 3761, 3243, 3378, 3638, 3259, 3128, 3632, 3281, 3653, 3215, 3407, 3330, 3154, 3070, 3608, 2785, 3536, 3338, 3964, 3369, 3208, 3765, 3123, 2865]. Selective [3265, 2464, 3690]. Selectivity [2617]. Selector [2204, 3048]. Self [3160, 2807, 2354, 2840, 1538, 2729, 3584, 3140, 3121, 3763]. Self- [1538]. Self-Consistency [2729]. Self-consistent [3160]. Self-Modelling [2807]. self-normalisation [3763]. self-normalization [3584]. self-normalized [3140, 3121]. Self-Similar [2354]. Self-Weighted [2840]. Semi [3883, 3879, 434, 3966, 512, 641, 3489, 263, 603, 1600, 3194, 3749, 3756]. semi-automatic [3194]. semi-definite [3756]. Semi-Infinite [603]. Semi-Latin [1600]. Semi-Markov [641, 434, 512]. Semi-parametric [3879, 3966]. Semi-Squared [263]. Semi-supervised [3883, 3489, 3749]. Semicompeting [2979, 3298]. Semiparametric [1977, 2326, 2599, 3024, 3116, 3087, 2849, 3450, 2046, 2806, 2657, 2769, 2082, 2948, 2245, 2971, 2999, 2574, 2464, 2490, 2390, 2550, 2872, 2381, 3029, 3341, 3532, 3066, 2219, 2890, 2729, 2911, 3187, 3009, 3310, 2954, 3481, 3743, 3261, 3728, 3284, 3064, 3402, 3129, 3046, 3410, 3446, 3145]. Semiparametrically [3001, 3492]. Semisupervised [3575]. sense [3561]. Sensing [2786]. Sensitivities [583]. Sensitivity [2617, 2582, 2538, 2570, 1679, 2794, 1592, 2304, 3546, 3289, 3561, 3852]. Sentencing [145]. Separability [2538]. Separable [2388, 2560, 3254]. Separate [580, 1907, 1584, 3222]. Separating [2917]. Separation [1871, 3031, 2791]. Sequence [1278, 610, 3198, 815]. Sequences [434, 158, 418, 455, 612, 2319, 556, 989, 1208, 2207, 392, 1502, 3375, 3477, 3545]. Sequential [748, 129, 260, 148, 192, 112, 344, 1513, 940, 1089, 1048, 1388, 1192, 116, 2828, 172, 501, 1012, 2891, 3003, 1061, 3354, 3315, 500, 531, 1461, 1593, 1477, 1690, 1283, 2774, 622, 426, 822, 989, 2239, 2265, 1668, 578, 388, 1634, 289, 3002, 375, 747, 1232, 1343, 1677, 962, 778, 977, 1110, 117, 698, 2688, 2862, 208, 1094, 1495, 468, 594, 1217, 2551, 728, 3327, 3503, 3448, 3231, 3680, 3601, 3213, 3359, 3603]. Sequentially [2938, 1919, 2054, 256]. Serial [406, 2404, 2539, 263, 2593, 1729, 723, 1585]. Serially [392]. Series [1595, 1084, 358, 113, 267, 125, 2145, 2039, 1892, 1916, 2585, 2326, 2723, 2790, 1882, 54, 320, 3085, 480, 2503, 2162, 114, 279, 316, 1081, 2769, 154, 575, 1269, 2063, 2349, 1583, 1756, 2707, 2708, 1448, 1486, 407, 897, 1946, 2180, 2325, 681, 1369, 1442, 1504, 2491, 2528, 1342, 1356, 369, 763, 325, 314, 1586, 1845, 119, 1148, 1702, 2713, 1505, 2100, 735, 2396, 814, 1909, 1748, 150, 420, 1705, 1656, 924, 168, 409, 997, 1507, 709, 987, 865, 315, 2016, 105, 416, 457, 70, 2837, 2121, 2125, 1877, 1022, 3654, 1931, 268, 1108, 1864, 1776, 359, 1664, 1772, 2946, 2525]. series [2916, 3170, 3457, 3782, 3314, 3352, 3584, 3162, 3610, 3217, 3328, 3560, 3747, 3495, 3911, 3937, 3249, 3357, 3140, 3121, 3407, 3498, 3373, 3875, 3353, 3481, 3763, 3507, 3308, 3799, 3846]. Server [347, 474, 1250, 1367]. Serves [727]. Service [283, 488, 489, 474, 524, 329, 362, 555, 682, 552, 2801, 348, 514, 412, 81]. Service-Time [474, 412]. Serving [491]. Session [111, 32, 50]. Set [1054, 86, 3379, 378, 379, 1948, 255, 2634, 2588, 83, 2616, 3422]. Sets [3025, 2990, 2605, 325, 2047, 2838, 161, 1072, 572, 2766, 1142, 2688, 3562, 3907, 3882, 3182, 3220, 3294, 3984]. Setting [2784, 2615]. Settings [2559, 3749, 3727]. Several [193, 914, 1567, 693, 727]. Shakeel [3920]. Shape [2333, 2109, 1957, 2159, 2912, 1895, 2493, 2047, 3366, 3621, 3403]. shape-restricted [3403]. shaped [3748]. Shared [2699, 2760, 3201]. sharp [3873]. Sharpening [805, 2630]. Sharpness [2937]. Sheppard [1579]. Sherlock [3932]. Shewhart [102]. shift [3882, 3984]. Shifted [1970, 1619, 1949, 2578, 3433]. shifts [3433, 3802]. Short [2240, 612, 1253, 987]. Short-Cut [1253]. shortfall [3311, 3847]. Shpitser [3720]. Shrinkage [3056, 2541, 1610, 1802, 2216, 1335, 1766, 2283, 2823, 2267, 2348, 2925, 3075, 3734, 3881, 3548, 3190, 3342, 3154, 3371, 3276]. Shrinking [3055]. side [3493, 3769]. Sided [940, 1089, 1388, 1173, 1187, 1582, 3367]. Sides [2230]. Siegmund [3898]. Sign [562, 699, 754, 1872, 2152, 2153, 2949, 2650, 1856, 3592]. Signal [2442, 2890, 3545, 3432, 3808]. signal-to-noise [3432]. Signals [925, 660, 99, 3394, 3628]. Signed [686, 2078, 1478, 1972, 1602, 3108]. Significance [481, 937, 366, 342, 337, 136, 354, 339, 811, 836, 1165, 331, 332, 908, 340, 55, 62, 825, 2944, 37, 76, 3468]. Significant [2773]. Signs [3001]. Similar [54, 2354, 2389, 389]. Similarities [545]. Simple [305, 395, 1397, 1426, 994, 410, 411, 524, 2773, 1832, 1430, 2078, 413, 490, 2699, 874, 1529, 2476, 263, 822, 823, 989, 190, 204, 1521, 510, 706, 1969, 1236, 1921, 444, 534, 2336, 218, 587, 642, 715, 586, 1359, 298, 2555, 2261, 3578, 3608, 3699]. Simplex [1697, 859, 935, 1145, 620]. Simplex-Centroid [620]. Simplex-Lattice [859, 935]. Simplification [1111]. Simplified [908, 431]. Simpson [2392]. Simulated [2712]. simulating [3952, 3951, 3953, 3957, 3947, 3950, 3954, 3955, 3948, 3949, 3956]. Simulation [3351, 2029, 1896, 1385, 2496, 2452, 3391, 2459, 2749, 3757]. Simulations [1912]. Simultaneous [1245, 267, 3103, 1124, 655, 3960, 3985, 1264, 1324, 2759, 2825, 2944, 3115, 1545, 1029, 3131, 3810, 3732, 2824]. Single [2212, 347, 474, 491, 1206, 1340, 2072, 2147, 812, 849, 85, 1205, 1594, 412, 2553, 2665, 1181, 1791, 2003, 692, 2899, 3224, 3377, 3301, 3227, 3850]. Single-Factor [2003]. Single-Index [2553, 2899, 3224, 3377, 3301, 3227, 3850]. Single-Nucleotide [2665]. Single-Server [474]. Single-Site [2072]. Single-Stage [1791]. singular [3186, 3966, 3400, 3574, 3156]. Singularities [1265]. Sir [366, 365, 367]. Siroki [3837].

Site [2072]. sites [3451]. Situations [1280, 1329, 3047, 2554]. Size [1241, 802, 787, 1033, 413, 500, 297, 2843, 2390, 1680, 428, 276, 3406, 3636]. Sizes [841, 2987, 1690, 1188, 1181]. Skew [2475, 2704, 2689]. Skewed [3034]. Skewness [2005]. Slice [2598, 2658, 2479]. Slippage [778, 1110]. Slope [1533, 568]. Slowly [1353]. Small [118, 2549, 1012, 346, 2882, 2987, 1510, 1188, 856, 3104, 2993, 21, 3620, 2838, 2401, 217, 13, 2779, 2558, 3255, 3787, 3327]. Smaller [2126]. Smallest [244]. SMC [3231]. Smirnov [1560, 1604, 1263, 973]. Smirnov-Type [1604]. Smith [3889, 2378, 3967, 3890, 3886, 3891, 3892, 3893, 3894, 3895, 3887, 3896, 3897, 3898, 3899, 3900, 3901, 3888]. Smooth [1457, 2706, 1296, 2321, 2574, 2815, 2400, 1978, 3051, 3415]. Smoothed [1920, 1933]. smoother [3238]. Smoothers [2669]. Smoothing [2585, 2884, 2513, 1718, 2789, 2906, 2438, 1715, 2675, 1779, 1803, 2023, 2276, 2180, 2410, 1768, 2395, 2870, 763, 3105, 3065, 2770, 2757, 3243, 2462, 1350, 2365, 2522, 2642, 2966, 3066, 1684, 1855, 1372, 1588, 2383, 371, 422, 2527, 3030, 3683, 3165, 3621]. Smoothness [2142, 2818, 3008, 3512]. Snell [919]. Soap [3030]. Society [4, 3654, 3, 80, 111]. Soft [2586]. Software [1998]. Sojourn [2623, 2740]. Solution [731, 1054, 638, 347, 1320, 512, 641, 191, 246, 553, 806, 678, 412, 948, 2385, 519, 3815]. Solutions [994, 503, 2103]. Solving [2881, 556, 2131, 700, 905, 3876]. Some [1093, 731, 165, 148, 310, 395, 599, 2577, 46, 59, 177, 343, 914, 52, 1762, 1089, 1048, 573, 792, 660, 352, 1446, 2787, 252, 320, 656, 2279, 1150, 73, 337, 1403, 858, 1573, 1033, 684, 1127, 277, 1320, 1433, 863, 1628, 293, 492, 557, 114, 378, 379, 1386, 356, 1109, 1147, 1223, 531, 2349, 1603, 1529, 1779, 1049, 249, 1280, 2409, 1868, 363, 1552, 1581, 938, 613, 273, 3065, 426, 989, 223, 386, 1100, 1524, 1017, 677, 1379, 328, 789, 1885, 1264, 990, 578, 891, 1184, 1423, 753, 444, 534, 286, 1174, 1949, 440, 2010, 2559, 1680, 289, 673, 242]. Some [53, 644, 476, 218, 1976, 1645, 1954, 1536, 498, 790, 530, 1467, 1159, 544, 1684, 778, 117, 698, 550, 187, 19, 1791, 880, 2143, 1022, 245, 367, 327, 728, 1114, 3734, 3535, 1643, 124, 1329]. Song [2345]. sooner [3193]. Sorting [107]. Sorts [373]. Sound [1967]. Space [2671, 2503, 3028, 2209, 1223, 2130, 3270, 1016, 2047, 1473, 2324, 1752, 2449, 2647, 2560, 3231, 3246, 3204, 3374, 3788, 3871, 3294, 3621]. Space-Time [1473, 3871]. Spaced [848, 831, 3052, 3390]. spaces [3405, 3767, 3472, 3531, 3274]. Spacings [1318, 2163, 1780, 729, 2071]. Sparse [3312, 3444, 3103, 1749, 2275, 3573, 3019, 2763, 2521, 3095, 3511, 3808, 3874, 3234, 3205, 3394, 3322, 3186, 3658, 3359, 3329, 3588, 3522, 3427, 3579, 3465, 3478, 3628, 3805]. Sparsely [2962]. sparsified [3314]. Sparsifying [3325]. Sparsity [2818, 3941, 3512]. Spatial [2856, 3025, 1869, 1172, 2060, 1575, 1434, 1230, 2990, 1905, 2437, 1290, 1349, 1573, 1013, 2217, 2584, 2923, 2988, 3755, 263, 1510, 3651, 2342, 2672, 1567, 3907, 1315, 1676, 2962, 2140, 1302, 1416, 3241, 2728, 2288, 2766, 1844, 2045, 3076, 2776, 3390, 3423, 3652, 3320, 3348, 3122, 3319, 3219, 3591, 3959, 3401, 3421, 3192, 3182, 3296, 3664, 3123]. Spatial-Temporal [2140]. Spatially [2670, 2808, 2761, 3331]. spatio [3771]. spatio-temporal [3771]. Spatiotemporal [2614, 3086, 3612, 2604, 2904, 3206, 3168, 3469, 3784, 3387, 2743, 3239]. Spearman [1310, 3412]. Special [250, 826, 18]. Species [450, 3094, 857]. Specific [1048]. Specification [113, 947, 1877, 34, 11, 3299]. Specified [2930, 2452, 1191, 3295]. Specimens [1290]. Spectacle [44]. Spectra [2326, 2903, 1122, 1049, 1413, 569, 589, 714, 3693, 3796]. Spectral [621, 1003, 1289, 1306, 567, 1316, 3085, 421, 537, 680, 2539, 369, 370, 447, 814, 420, 3021, 659, 1474, 1814, 2268, 3457, 3495, 3752]. spectral-density-driven [3495]. Spectroscopy [2379]. Spectrum [861, 408, 853, 2518]. spending [3603]. Sphere [1127, 1535]. spherelets [3741]. Spherical [1440, 1532, 1309, 1736, 1719, 2064, 1148, 2808, 2761, 1881, 1502]. Spherical-Wavelet-Based [2808, 2761]. Spherically [2838]. Sphericity [2069]. Spline [1010, 2093, 2804, 2789, 2906, 2675, 1134, 1768, 3065, 2770, 2993, 2642, 1684, 1372, 1588, 2383, 2629, 2869, 2385, 3241]. Splines [3074, 1718, 2853, 1736, 2082, 2410, 2565, 2462, 2739, 2685, 3336, 3621, 3243, 3238]. Split [3941, 3290, 3330]. split-and-merge [3330]. Splitting [2088, 3910]. Sports [1634]. Spread [1315]. Sprent [1616]. SPRT [1632]. Square [755, 816, 2223, 36, 1924, 415, 1272, 136, 1893, 242, 1654, 393, 231, 1162, 1488, 100, 1531, 400, 1777]. Squared [427, 2674, 1382, 1733, 263, 1735, 2883, 3578, 1651]. Squares [801, 1658, 601, 1762, 441, 2541, 3103, 1451, 875, 1418, 1086, 2083, 1404, 1517, 1352, 1643, 1715, 205, 2923, 897, 917, 921, 1163, 556, 170, 314, 2593, 2672, 2380, 1616, 1638, 2465, 2553, 1708, 1600, 202, 947, 1277, 214, 315, 1874, 162, 769, 1919, 2054, 217, 1164, 84, 3067, 3812, 3445, 3248, 3946, 3360, 3654]. Srakar [3971]. Srivastava [991]. Stability [3128, 3215, 3446]. Stabilization [2885]. Stable [2357, 2358, 2318, 2119, 1430, 520, 786, 3008, 3246, 3612, 3145]. Stage [606, 1734, 2190, 226, 2086, 2989, 893, 2405, 1753, 993, 1791, 3590, 3499, 3588, 3850]. staggered [3689]. Standard [93, 171, 8, 1996, 1044, 582, 228, 1657, 2517, 273, 2554, 179, 2559, 21, 713, 128, 2957, 1277, 1213, 169, 11, 3278]. Standardization [98]. Standardized [2315, 2335]. Stark [3899]. State [2671, 2212, 2503, 2209, 2048, 2130, 1030, 1752, 2449, 3231]. Statements [713]. States [150]. Stationarity [2752, 924, 3390, 3328, 3249]. Stationary [847, 925, 861, 1249, 1289, 1306, 725, 1047, 887, 397, 2923, 391, 681, 609, 941, 853, 576, 646, 637, 814, 2908, 420, 3871, 569, 589, 714, 762, 1121, 997, 709, 2728, 768, 864, 1814, 2911, 1022, 268, 738, 3457, 3225, 3305, 3249, 3605, 3498, 3308]. Stations [92]. Statistic [1675, 1336, 1370, 1509, 1888, 313, 1188, 2335, 1151, 2308, 1263, 1019, 1323, 1193, 1157, 1645, 535, 376, 301, 2588, 1076, 3155, 3760]. Statistical [446, 664, 1542, 2922, 3, 80, 111, 140, 165, 239, 2475, 310, 2212, 175, 292, 1412, 2732, 59, 120, 2456, 1762, 1172, 1775, 3828, 2841, 1010, 573, 1575, 374, 3829, 352, 3823, 3830, 3831, 1811, 35, 3078, 320, 73, 3246, 1378, 1644, 684, 3699, 234, 311, 1362, 2959, 1043, 1957, 44, 18, 3762, 131, 124, 3827, 249, 710, 302, 3821, 2641, 1894, 92, 51, 33, 185, 2719, 839, 3832, 2615, 3533, 261, 1186, 3833, 447, 3435, 1184, 20, 7, 2969, 159, 176, 81, 25, 399, 3834, 3835, 2666, 1686, 53, 324, 3822, 5, 3836, 128, 1536, 1325, 2705]. Statistical [3464, 3727, 1322, 507, 764, 2857, 416, 457, 2429, 1169, 1256, 2698, 3802, 19, 143, 3826, 3838, 123, 3654, 3824, 3839, 1301, 79, 3619, 3799, 3840, 60, 3921, 3938, 3922, 3647, 3923, 3924, 3780, 3920, 3925, 3825, 3926, 3927, 3928, 3929, 3326, 3937, 3916, 3919, 3930, 3914, 3931, 3932, 3933, 3841, 3820, 3752, 3934, 3935, 3318, 3837, 3511, 3773, 3918, 3913, 3936, 3917, 3915, 2788, 2667, 102]. Statistically [3114, 11]. Statistics [2110, 1665, 112, 1371, 2060, 791, 1620, 183, 1031, 2264, 1749, 317, 2078, 2170, 2029, 1075, 1219, 1327, 3337, 1382, 1478, 2524, 2115, 2988, 2368, 2431, 1369, 1906, 3049, 2774, 2858, 1295, 2474, 1231, 1390, 1209, 1543, 1958, 1381, 2579, 1765, 1881, 561, 564, 961, 2327, 2288, 1358, 180, 973, 1856, 707, 2081, 1262, 389, 2071, 9, 6, 3194, 3285, 3986, 3373].

Status [2552, 2290, 2451, 3637]. Steady [1483, 604, 1417, 1508]. Steel [3863, 3956]. Stein [1265, 3816, 3477, 1678, 744, 1722]. Stein-Type [1265]. Steiner [3956]. Step [2494, 2520, 1978, 3076, 3280]. Steps [2306]. Stepwise [2244, 2774, 3208]. Stereology [1920]. Steven [3891]. Stochastic [459, 2639, 177, 1106, 487, 2985, 2624, 2015, 1402, 724, 2485, 503, 1663, 2861, 2735, 887, 1250, 2584, 1395, 325, 1726, 862, 178, 898, 1953, 1346, 370, 497, 1660, 881, 444, 176, 1538, 1640, 912, 218, 1954, 1809, 2772, 530, 1159, 3294, 266, 927, 1806, 508, 967, 1367, 390, 1728, 2444, 1826, 3433, 3568, 3186, 3393, 3585, 3904, 3533, 3161, 3411, 3435, 3696, 194]. Stochastically [1854, 3207]. Stockout [1038]. Stopping [2828, 858, 720]. Stopping-Time [2828]. Storage [385, 1001]. Store [74]. Storms [2119]. Strata [1941, 1179, 276, 3171]. Strategic [3849]. Strategies [1518]. Strategy [2141, 833, 1367, 3139]. Stratification [755, 816, 3940, 1722, 3420, 3230]. Stratified [2269, 1179, 2726, 1359, 3510, 3749]. Stratum [985, 832, 929]. Stream [941]. streaming [3562]. Streams [761, 880]. Strength [52]. Strikingly [994]. Strimmer [3837]. String [1541, 1606]. Strong [2114, 2759, 2824, 2825, 3188]. Structural [1492, 2587, 1405, 1123, 944, 1424, 1704, 1519, 1546, 2837, 2734, 3486, 3698, 3773]. Structure [2568, 3774, 1694, 1603, 1803, 2387, 774, 1046, 3079, 2203, 1831, 1951, 2331, 2728, 1501, 2497, 3535, 3794, 3476, 3518, 3479]. structure-adaptive [3518]. Structured [2608, 1823, 2861, 3274]. Structures [1834, 2491, 2528, 2593, 1838, 865]. Stuart [237]. Student [484, 148, 3155, 1129]. Studentization [71]. Studentized [1852, 2474]. Student’s [45]. Studies [2456, 2930, 2599, 2625, 154, 2622, 2774, 2463, 20, 2257, 2927, 2726, 1979, 1358, 1400, 2651, 3115, 2947, 3394, 3264, 3639, 3151, 3341, 3804, 3310, 3303]. Study [250, 309, 234, 749, 74, 782, 1385, 150, 1592, 266, 19, 880, 3216]. Studying [1817, 1074]. Sub [56]. Sub-Sampling [56]. Subadditive [898]. Subclass [2001]. Subgroup [1828]. Subject [239, 322, 2148, 2771, 2376, 287, 2307, 3532]. Subjective [934, 234, 1666]. Submodels [1501]. Subpopulations [2873]. Subsampling [2672, 3220]. Subsequent [825]. Subset [826, 1261, 1328, 3192]. Subsets [2333, 2799, 2236, 2202]. Subsidiary [556]. subspace [3287]. Substantive [1903]. Substituted [160]. Substructures [2961]. Success [161, 3731]. Successive [330, 228, 201]. Suffered [143]. Sufficiency [1221, 600, 650, 772, 2272, 2334, 903, 564, 733, 771]. Sufficient [1371, 3122, 313, 2873, 389, 3187, 3327]. Suitable [493]. Sum [2980, 1194, 130, 2319, 1163, 515, 823, 1632, 1164, 3803]. Sumca [3578]. Summary [964, 3020, 3194]. Sums [441, 170, 905]. Sun [3975, 3926]. Sung [2345]. Sunspot [286]. Super [1353]. Super-Slowly [1353]. Superconsistent [3602]. Superimposition [2223]. Superposition [907, 1130]. supersaturated [2600]. Supervised [3732, 3883, 3600, 3489, 3749]. Supplement [711]. Supplementary [482, 1044, 1213]. Supplements [1885]. support [3607, 3338]. supported [3987]. Sure [3028]. Surface [2142, 1073, 2437, 2248, 1101, 1008, 3798, 2785]. Surrogate [2973, 3024, 1618, 3109, 3370]. Survey [2068, 2402, 296, 2376, 936, 1043, 1651, 2896, 2690, 2380, 1466, 1105, 3565]. Survey-Sampling [1043]. Survey-Weighted [2380]. Surveys [2844, 1106, 1450, 806, 3104, 1380, 718, 2403, 3132]. Survival [1479, 2624, 2986, 2279, 1433, 2792, 2461, 2317, 1355, 2941, 2137, 1854, 1568, 3014, 3073, 2327, 2651, 2547, 3777, 3786, 3437, 3226, 3173, 3532, 3406, 3455, 3609].

Survivor [2768]. Susceptibility [2305]. Swallow [3867]. Swendsen [2478]. Switching [2736, 2556, 3193]. Symbolic [2104]. Symmetric [1440, 1397, 1503, 1824, 1521, 1634, 2838, 1145, 1406, 3405, 3202]. Symmetric-Simplex [1145]. Symmetrical [1176, 1204, 974, 1205]. Symmetries [3033]. Symmetry [2704, 1823, 1519, 1105, 1502, 3396]. Synaptic [2448]. synchronization [3962]. Synchronous [1220]. Syndicate [89]. Syndrome [2190, 1966]. synthesis [3939]. Synthetic [3689, 3906]. System [1514, 1515, 104, 474, 524, 887, 1500, 465, 150, 644, 2154, 784]. Systematic [225, 252, 1418, 2112, 306, 107]. Systems [250, 2255, 536, 1056, 1172, 342, 2018, 2624, 2485, 555, 682, 785, 385, 2272, 2184, 758, 786, 1726, 167, 2239, 1838, 1634, 2020, 1069, 2428, 107, 2791]. Szabo [3868].

t [777]. Table [937, 1005, 968, 27, 441, 1294, 304, 608, 483, 1769, 128]. Tables [1514, 1580, 366, 1924, 615, 1067, 571, 1763, 354, 339, 756, 2024, 2028, 1893, 346, 655, 960, 230, 1201, 497, 1365, 1463, 563, 128, 2187, 229, 138, 973, 1585, 1819, 1368, 13, 2637]. Taguchi [2188]. Tail [2736, 2981, 2341, 1263, 1714, 3475, 3348, 3396]. Tail-Areas [1263]. Tailed [2981]. Tailor [1695]. Tailor-Made [1695]. Tails [3053]. Tamás [3932]. Tamhane [1524]. Tan [2720]. Tang [3721]. Tao [3839]. target [2573]. Targeted [3563]. Tasting [1203]. Tawn [2783]. Tchebycheff [351]. Tchetgen [3722]. Tea [1203]. Tea-Tasting [1203]. Technique [118, 410, 411, 233, 786, 185, 416, 457]. Techniques [1215, 1779, 1803, 263, 1895, 3411]. Telephone [1757, 167, 81]. Temperature [2808, 2761]. tempering [3688, 3536]. Template [2922]. Temporal [2140, 3959, 3435, 3771]. Tennis [1634]. Tensor [3555, 2082, 3782, 3966, 3762, 3427].

TeraLasso [3555]. Term [91, 709]. Terminal [957, 1055]. Terms [235, 1873, 1191, 2355]. Test [2481, 937, 755, 816, 626, 404, 344, 528, 189, 366, 373, 686, 699, 2039, 1102, 754, 1128, 1192, 1872, 1293, 1434, 1587, 2446, 2566, 2723, 2320, 1144, 2972, 2170, 1764, 354, 339, 1061, 1370, 1109, 500, 614, 1068, 675, 1363, 836, 1165, 608, 1477, 2792, 640, 908, 1894, 3049, 1657, 1036, 2858, 306, 822, 989, 1510, 1497, 770, 1253, 2308, 828, 856, 918, 969, 1058, 1631, 1854, 1324, 206, 1972, 62, 1274, 1654, 924, 1069, 1360, 1607, 2848, 865, 2154, 893, 2327, 1232, 1841, 825, 416, 457, 377, 798, 757, 2451, 275, 1582, 203, 1615, 1248, 2356, 522]. Test [899, 188, 1094, 504, 1217, 2551, 181, 13, 3390, 3566, 3267, 3216, 3849, 3909, 3328, 3816, 3519, 3650, 3249, 3773]. Testing [1687, 3990, 145, 1146, 802, 3032, 2213, 848, 831, 940, 86, 3044, 1426, 1215, 3983, 1428, 2754, 2424, 3584, 1763, 2038, 3447, 2258, 2894, 537, 1506, 1562, 302, 2404, 3739, 1997, 2574, 2122, 2752, 1210, 1962, 3046, 2066, 1311, 1174, 950, 2892, 1656, 1813, 535, 1729, 1584, 1669, 3403, 2944, 542, 3986, 3061, 3482, 3484, 3846, 2748, 3805, 3347, 3538, 3592, 3453, 3648, 3589, 3586, 3237, 3493, 3476, 3518, 3985, 3961, 3944, 3767, 3372, 3740, 3395, 3296, 3802]. Tests [1093, 481, 570, 687, 716, 596, 148, 1659, 1235, 2098, 112, 358, 46, 267, 914, 1617, 562, 838, 885, 896, 1089, 342, 1842, 2244, 1388, 2156, 2069, 1932, 1988, 588, 2413, 2804, 2108, 501, 580, 2758, 1672, 2279, 1150, 337, 964, 1082, 3085, 2609, 1427, 1560, 1554, 1604, 2281, 2686, 279, 316, 2657, 2001, 986, 1109, 2049, 674, 531, 1733, 331, 1651, 3380, 1695, 1896, 357, 407, 2152, 2153, 1133, 1283, 1868, 2539, 1442, 1504, 2319, 1966, 3112, 426, 1691, 1318, 3054, 1379, 2282, 1780, 2177, 1141, 882, 340, 1950, 2002, 1559, 1907, 289, 55, 62, 168, 947]. Tests [1416, 1537, 1277, 878, 961, 2443, 2327, 653, 1922, 1856, 698, 1022, 245, 3051, 389, 400, 2355, 2071, 37, 76, 2640, 3539, 3225, 3541, 3213, 3942, 3538, 3904, 3148, 3401, 3136, 3313, 3463, 3682, 3467, 3187, 3803, 3478, 3754, 3545, 3609, 3356, 930, 3222]. text [3983]. Textile [114, 105]. Textures [1985]. th [1613, 2633]. thanks [3670, 3669, 3702, 3854, 3703, 3886, 3821, 3969, 3855, 3914, 3887, 3822, 3970, 3948, 3949, 3913]. Their [2577, 2755, 1908, 958, 2686, 114, 1838, 2645, 737, 3010, 1998, 3408, 741, 3651, 499]. theme [3266]. Theorem [435, 2362, 807, 1735, 1614, 1880, 2025, 1640, 721, 978, 707, 1021, 3789, 3416, 403]. Theorems [1717, 1478, 1147, 1218, 753, 673]. Theoretic [1783, 1986, 949, 3448, 3606, 3682]. Theoretical [113, 3415, 2987, 1885, 1909, 2089]. Theory [446, 1353, 239, 395, 211, 638, 607, 342, 837, 309, 103, 158, 224, 257, 1290, 1378, 1013, 1006, 277, 385, 99, 333, 974, 71, 809, 917, 2569, 223, 386, 898, 971, 2719, 1713, 1953, 2163, 1614, 312, 881, 461, 197, 648, 1507, 642, 216, 266, 419, 798, 757, 28, 479, 2497, 2071, 781, 1496, 1114, 433, 1065, 3552, 3445]. Therapy [166]. there [482]. Thick [3162]. Thin [2082, 2685]. thinning [3738]. Thomas [3719, 3955, 3900]. those [1634]. Three [1005, 615, 1010, 1908, 2915, 730, 2048, 1275, 960, 776, 2177, 1880, 857, 1949, 653]. Three-Arm [2915]. Three-Dimensional [1880]. Three-Parameter [1908, 1275, 2177, 1949]. Three-State [2048]. Three-Way [615, 960]. Threshold [2338, 1932, 1988, 2581, 2276, 2000, 2332, 1455, 3542, 3197, 3750]. thresholded [3791, 3276]. Thresholding [2426, 2635, 2586, 3240, 3499]. Thresholds [1929]. throughput [3394]. Thrown [1541, 1606]. Thursday [4]. Tianxi [3977]. Tightness [1816]. tilted [3746]. Tilting [2707, 3083, 3391, 3199, 3339]. Time [1595, 265, 305, 2212, 113, 125, 2145, 2039, 1215, 1892, 1916, 660, 1479, 2585, 2724, 2326, 1289, 2723, 2790, 2828, 1882, 474, 1731, 3085, 329, 480, 2503, 2512, 2795, 2162, 114, 279, 316, 1081, 2769, 154, 2063, 2349, 2707, 2708, 2801, 1448, 1486, 407, 897, 1946, 2180, 254, 2325, 1369, 1442, 1504, 466, 2820, 2491, 2528, 1342, 1356, 553, 369, 763, 314, 1586, 1355, 1845, 786, 839, 2926, 1702, 2713, 547, 1505, 2090, 1379, 2463, 2100, 735, 2396, 412, 380, 381, 1071, 814, 2762, 1727, 1909, 604, 1311, 1748, 2457, 150, 420, 1705, 1473, 1656, 924, 168, 409, 997, 1069, 1507]. Time [2962, 709, 865, 2942, 1159, 720, 768, 2016, 105, 2837, 2121, 2125, 1877, 2859, 2852, 1022, 2740, 1931, 268, 1108, 1864, 1776, 2710, 359, 2946, 2525, 2616, 519, 2546, 2916, 3299, 3170, 3457, 2560, 3782, 3795, 3314, 3139, 3389, 3246, 3352, 3584, 3572, 3656, 3130, 3162, 3610, 3217, 3270, 3328, 3560, 3747, 3495, 3911, 3937, 3508, 3249, 3871, 3357, 3140, 3121, 3407, 3294, 3446, 3428, 3498, 2785, 3373, 3310, 3875, 3353, 3481, 3763, 3507, 3131, 3308, 3799, 3846]. Time-Dependence [1069]. Time-Dependent [553, 2926, 2762, 997, 519]. Time-Homogeneous [265]. Time-Reversion [768]. Time-Series [113, 480, 279, 1369, 1442, 1504, 369, 924, 168, 409, 997]. Time-to-Emptiness [604]. time-to-event [3299, 3310]. Time-Trending [254]. Time-Varying [2837, 2710, 3795, 3572, 3656]. Times [250, 330, 2624, 470, 1222, 2722, 2623, 941, 380, 381, 737, 477, 2307, 2962, 2519]. Tinbergen [150]. Tolerance [664, 746, 988, 909]. Tomography [2819, 2914, 2095, 1920]. Tong [3978]. Too [34]. Tool [2300]. Tools [2104, 1976]. Topics [801]. Torben [3954]. Total [1207, 850, 804, 2757, 2390, 1380, 3697, 3693, 3796]. Toxicological [2258]. Trace [1576, 3778]. Trace-class [3778]. Traces [300]. Tracking [2596, 1998, 2837]. Traffic [660, 99, 1312, 759, 577, 1187, 972, 509, 508]. Trait [2662, 3010]. trajectory [3759].

Transactions [133]. Transfer [3681, 1132, 1865, 1926, 3139]. Transform [1316, 2444]. Transformation [1970, 2316, 36, 120, 958, 2487, 2005, 469, 414, 1037, 1303, 1007, 2525, 3034, 2382, 3162, 3253, 3410, 3310]. transformational [3941]. Transformations [1146, 1528, 1010, 3289, 666, 1428, 2755, 1498, 2321, 1366, 717, 2189, 2230, 733]. Transformed [1269, 2883]. Transforms [272, 1969]. Transient [2831, 490, 1566]. Transition [2879, 1034, 3478, 264]. transport [3819, 3623, 3817]. transported [3451]. transposable [3203]. Transposed [282]. Travelling [323]. Travelling-Saleman [323]. Treated [802, 233, 230]. Treatment [305, 1534, 1605, 848, 831, 3087, 2955, 2609, 2245, 1581, 1657, 2971, 2542, 3054, 2702, 642, 269, 1677, 1520, 1422, 1545, 390, 79, 2950, 3491, 3879, 3449, 3647, 3873, 3364, 3795, 3786, 3363, 3454, 3458, 3877, 3797, 3437, 3440, 3644, 3659, 3230, 3692, 3309, 3630, 3663, 3281, 3494, 3487, 3963, 3801, 3609, 3800, 3906]. treatment-free [3692]. Treatments [1513, 378, 379, 1211, 2147, 1657, 84, 3554, 3418, 3963]. Tree [2608, 3433, 3646, 3774, 3512, 3965]. Tree-Structured [2608]. Trees [3010, 3564, 3747]. Trend [358, 316, 421, 314, 510, 315, 377, 3277, 3759, 3407]. Trend- [315]. Trend-Reduced [314]. Trending [254]. Trends [2946]. Trial [63, 3801]. Trials [1734, 2915, 2542, 232, 161, 2950, 3370, 3792, 3606, 3588]. Triangles [1895]. Triangular [1625, 776, 2791]. Triangulation [2417]. Trimmed [1462, 3812]. Triograms [2757]. Triple [2405]. Triple-Goal [2405]. trouble [3805]. True [3096, 2850, 3803]. Truncated [373, 591, 130, 2139, 3362, 2390, 538, 494, 525, 1281, 3511]. Truncation [2276, 2851, 722, 3418]. Tuesday [111]. Tukey [1118, 614, 2351]. Tumours [2829].

Tune [2345]. Tuning [3236, 3075, 3748]. tuning-free [3748]. Tunnels [1660]. Two [748, 1093, 1530, 664, 192, 1659, 1983, 2212, 373, 1734, 1515, 940, 1089, 1388, 1429, 2339, 1336, 1444, 2924, 3267, 2930, 2754, 2441, 470, 1598, 1082, 2676, 1427, 1560, 1554, 1604, 2520, 3094, 207, 1131, 1971, 1211, 1386, 2001, 1666, 674, 675, 1893, 2115, 226, 957, 1055, 3077, 832, 1477, 332, 813, 1133, 1283, 2245, 955, 1552, 1087, 622, 478, 235, 1997, 3660, 689, 959, 1151, 1201, 1594, 2030, 1187, 828, 856, 918, 388, 453, 736, 1209, 2086, 1972, 1381, 2989, 1360, 1607, 535, 1547, 2848, 484, 2405, 1337, 1753, 825, 3395, 1755, 1856, 431, 508, 2813, 436, 1531, 3076, 3092]. Two [1788, 1282, 3850, 84, 3875, 3539, 3525, 3590, 3849, 3617, 3499, 929, 3622, 3217, 3243, 3686, 3372, 3588, 3409, 3655, 3754, 3261]. Two- [2339, 1082, 918]. two-armed [3849]. two-component [3372]. Two-Dimensional [1515, 1386, 736]. Two-Factor [1598, 1087, 1594, 1755, 3617]. two-group [3622, 3655]. Two-Lane [431, 508]. Two-Level [2924, 2441, 3077]. Two-Order [1209]. Two-Pack [478]. Two-Parameter [3094]. Two-pass [1971]. Two-Phase [2930, 3092, 3261]. Two-Sample [1093, 1336, 1427, 1560, 1604, 674, 675, 1133, 2245, 828, 856, 1972, 1381, 1856, 1788, 3267, 3660, 3395, 3539, 3525, 3849, 3217, 3409, 3754]. Two-Sided [940, 1089, 1388, 1187]. Two-Stage [1734, 226, 2086, 2989, 2405, 1753, 3850, 3590, 3499, 3588]. Two-State [2212]. Two-Step [2520, 3076]. Two-Variety [1531]. Two-Way [2676, 1893, 1552, 1201, 1337, 3875]. Twofold [1109]. Tyler [3838]. Type [1595, 2052, 1265, 791, 503, 1560, 1554, 1604, 863, 1124, 397, 2149, 1178, 348, 770, 990, 702, 242, 1323, 1381, 961, 3568, 3331, 3188, 1125, 2412]. Types [470, 677, 1567, 1748, 143, 727].

Uhlenbeck [2577, 2772].

Uhlenbeck-Based [2577]. Ultimate [132]. Ultrahigh [3028, 3157, 3205, 3179, 3224, 3330]. Unavailability [1963]. Unbalanced [1237, 2820, 2074, 549]. Unbiased [2821, 649, 516, 2368, 2431, 3582, 1450, 770, 2218, 2958, 3578]. Unbiasedness [1072]. Uncategorized [1699]. uncertain [3453]. Uncertainty [2847, 2199, 1170, 3297, 3197]. Uncorrelated [3018, 1366]. Understanding [2438]. Undertones [1184]. Unequal [1530, 2817, 2898, 1523, 277, 2374, 587, 2451]. Unexpected [2207]. Ungrouped [1754]. Uni [380, 381]. Uni-Directionally [380, 381]. Unidentifiable [340]. Unification [1036, 2411, 3780]. Unified [1617, 2532, 1625, 333, 1369, 886, 1117, 2759, 2825, 3478, 3578, 3596, 2824]. Uniform [688, 1945, 1318, 1058, 3374]. Uniformity [63]. Uniformly [3007]. Union [697]. Uniqueness [1233]. Unit [235, 1225]. United [150]. Units [1530, 277, 832, 929, 201]. Univariate [1619, 503, 1109, 753, 673, 1050, 1103, 1693, 3422, 2459]. Universal [3007, 3756]. University [183]. Unknown [1150, 2948, 1134, 1126, 375, 535, 3063, 2361, 2418, 2890, 3166, 3346, 3436, 3422, 3519, 3882, 2419, 3456]. unmeasured [3945, 3580]. Unobserved [1592]. Unordered [1133, 3754]. Unrelated [2007]. Unrestricted [634]. Unstratified [1179]. unsupervised [3753]. Updates [2775]. Updating [2278, 2072, 2331, 3381, 3851]. Upon [284, 1379]. Upper [1598, 1653, 823, 1652, 2488]. Urn [2384]. Usable [3678, 3785]. Usage [1657]. Use [585, 2099, 2639, 1485, 172, 2258, 1382, 1394, 2027, 1044, 378, 379, 1930, 803, 1699, 976, 3112, 119, 1175, 21, 2037, 439, 1213, 973, 590, 3618, 2775, 390, 1864, 127, 69, 493, 3408, 3761]. Used [1779, 822, 150, 515]. Useful [731, 974, 1966, 255, 886, 1117]. Usefulness [2254]. Uses [2577, 1575, 207, 198]. Using [1580, 2440, 305, 2805, 1637, 1842, 1128, 2985, 1273, 1475, 2624, 3024, 2804, 694, 2458, 850, 804, 2437, 2029, 2295, 531, 2708, 1843, 2180, 2368, 2431, 1906, 3049, 2395, 2346, 2610, 622, 1720, 1586, 2354, 2690, 2692, 2914, 2490, 2462, 1549, 2829, 3104, 2089, 2283, 2917, 2993, 2236, 1550, 1602, 1790, 2876, 1458, 475, 2705, 3060, 214, 961, 1711, 1484, 1677, 1422, 1649, 2511, 2355, 1935, 1496, 2616, 3349, 2044, 3433, 3423, 3444, 3506, 3549, 3564, 3321, 3527, 3420, 3945, 3553, 3179, 3258, 3235, 3499, 3815, 3521, 3792, 3411, 3650, 3173, 388, 3607, 2202, 3376, 3663, 3612, 3588, 3158]. using [3403, 3643, 3986, 3487, 3845, 3172]. Usual [2605, 2002, 2495, 3422]. Utility [746, 496, 868]. Utilization [832, 929, 157].

V [3718]. Vacancies [523, 464]. Vaccination [2606]. Vague [1778, 1769, 2181, 1563]. Valid [3637, 2074, 1496, 3649]. Validated [1919, 2054, 1588]. Validation [2625, 1802, 2083, 1952, 2055, 2180, 2389, 2283, 1288, 3179, 3367, 3753, 3597, 3522]. Validatory [1169, 1256]. Validity [2247, 302, 2075]. Value [1353, 2581, 1142, 3186, 3966, 3769, 3220, 3727, 3197]. Valued [2585, 3393, 3064, 3610, 3875]. Values [1235, 1485, 1212, 493, 3004, 253, 2279, 2714, 130, 2656, 2782, 1385, 2192, 3216, 3961, 3905, 3729]. VanderWeele’s [3838]. vanishing [3807]. Vansteelandt [3708, 3706, 3709, 3710, 3711, 3702, 3704, 3712, 3707, 3713, 3714, 3715, 3716, 3717, 3718, 3719, 3720, 3705, 3721, 3722, 3723, 3724, 3703]. Variability [1703, 52, 1013]. Variable [1319, 2864, 802, 3056, 2416, 2450, 3103, 2771, 2531, 2523, 2217, 2028, 2967, 1690, 2803, 514, 2999, 3378, 2833, 3638, 381, 1324, 2353, 1002, 3215, 2505, 161, 1021, 34, 1819, 3338, 2834, 3157, 3311, 3488, 3795, 3199, 3561, 3816, 3521, 3644, 3989, 3129, 3406, 3630, 3330, 3608, 3642, 2785, 3536, 3208, 3765, 2865].

Variables [748, 703, 739, 1624, 1528, 920, 1292, 1202, 1096, 2750, 2148, 2386, 2549, 482, 1618, 2980, 1095, 1112, 1255, 2458, 1194, 337, 2881, 811, 2350, 1304, 689, 959, 1016, 300, 1702, 854, 842, 1357, 1656, 484, 439, 543, 468, 467, 2871, 3880, 3889, 3967, 3600, 3429, 3216, 3890, 3300, 3458, 3535, 3332, 3886, 3891, 3892, 3893, 3894, 3686, 3895, 3329, 3887, 3896, 3897, 3464, 3898, 3899, 3900, 3901, 3360, 3487, 3654, 3888, 3902, 3885, 3649]. Variance [533, 1703, 559, 626, 1659, 1235, 15, 36, 120, 2817, 1825, 441, 1429, 22, 996, 2041, 2758, 394, 95, 1160, 2972, 2885, 651, 3179, 629, 970, 1387, 408, 2082, 2988, 2675, 1863, 391, 2756, 1261, 1328, 602, 639, 640, 680, 813, 2506, 14, 1581, 2126, 1304, 2896, 2690, 2128, 3301, 1379, 2840, 2466, 2634, 2814, 2842, 1705, 1019, 1473, 486, 375, 535, 549, 750, 231, 1277, 544, 2288, 653, 2004, 1791, 2094, 2383, 3062, 1698, 76, 3575, 2427, 3201, 3634, 3560, 3422, 3388, 3233]. Variance-Covariance [2004]. Variance-Heterogeneity [120]. Variance-Time [1379]. Variances [86, 2069, 1068, 710, 3055, 689, 1174, 2164, 522, 3456]. Variate [75, 2232, 108, 273, 1303, 3480, 3881]. Variates [1070, 560, 1735, 425, 820, 3183]. Variation [406, 1529, 74, 2757, 2066, 19, 3363]. Variational [2656, 3809, 3411, 3618, 3964]. Varietal [529]. Varieties [1730]. Variety [3094, 74, 1531]. Varimax [3828, 3829, 3823, 3830, 3831, 3827, 3821, 3832, 3833, 3834, 3835, 3822, 3836, 3826, 3838, 3824, 3839, 3840, 3825, 3841, 3837, 3820]. Variogram [2672]. Various [307, 441, 2203]. VARMA [1796, 1835]. Varying [1353, 2986, 1074, 2789, 2683, 2118, 85, 3086, 1866, 2837, 2710, 3795, 3572, 3656, 3446, 3428, 3850, 3131]. Varying-Coefficient [2789, 2683, 2118, 3086, 3850]. Vector [2905, 897, 959, 1016, 1398, 1297, 1656, 808, 2286, 2385, 3574, 3338]. Vectors [1265, 3627]. Vehicular [99]. Vehicular-Controlled [99]. Version [435, 1872]. Versions [2144]. Versus [2401, 2015, 2149, 3760]. Vertex [3033]. Very [2990]. Vessels [143]. Via [2059, 2925, 2426, 2864, 3852, 3391, 2712, 2587, 2227, 3782, 3658, 3849, 3199, 3314, 3284, 2119, 2142, 3786, 2861, 1286, 3266, 3584, 3278, 3018, 2738, 2670, 3012, 3551, 3573, 2438, 3697, 2404, 2770, 3378, 3911, 3937, 3985, 3533, 3756, 3693, 3796, 3632, 2467, 3583, 3597, 3452, 3342, 721, 2728, 2164, 3511, 3746, 2267, 2588, 2818, 3154, 2348, 3465, 3498, 3484, 3373, 3371, 3878, 3964, 3546, 3690, 3763, 3846, 2865, 2834]. Vibrations [1122]. Vice [4]. Vice-President [4]. View [186]. Viewpoint [102]. Villalobos [3857]. Vintage [3821, 3841, 3820, 3822, 3837, 3828, 3829, 3823, 3830, 3831, 3825, 3827, 3832, 3833, 3834, 3835, 3836, 3826, 3838, 3824, 3839, 3840]. violence [3771]. Visualizing [3600]. Vladimir [3901]. Vol [382]. Volatilities [3018]. Volatility [3074, 2639, 2772, 3361]. Volume [2, 16, 29, 41, 64, 77, 96, 109, 122, 142, 149, 156, 164, 174, 182, 196, 209, 222, 238, 258, 278, 308, 334, 368, 398, 432, 458, 471, 485, 505, 527, 618, 663, 712, 766, 818, 870, 932, 981, 1026, 1064, 1120, 1167, 1198, 1258, 1300, 1376, 1420, 1470, 1527, 1571, 1623, 1683, 1746, 1786, 1822, 1859, 1897, 1936, 1989, 2053, 2132, 2194, 2249, 2309, 2369, 3041, 3101, 3884, 2905, 2433, 2432, 2498, 2618, 2677, 2745, 2742, 2809, 2863, 2918, 2975, 2974, 3038, 3037, 3098, 3097, 3141, 3174, 3211, 3210]. volume [3251, 3292, 3291, 3334, 3333, 3383, 3382, 3460, 3459, 3514, 3513, 3557, 3556, 3614, 3613, 3775, 3666]. Volumes [2605]. vote [3670, 3669, 3702, 3854, 3703, 3886, 3821, 3969, 3855, 3914, 3887, 3822, 3970, 3948, 3949, 3913]. voting [3499]. Vovk [3901]. vs [1764, 1582].

W [1213]. Waiting [250, 330, 329, 413, 552, 1046, 735]. Waiting-Time [735]. Waiting-Times [250]. Wald [1061, 2115, 289, 698]. Walk [1493, 634, 372, 3503]. Walker [3857, 3862, 3864, 3859, 3854, 3860, 3861, 3863, 3855, 3858, 3856, 3868, 3865, 3866, 3867, 3869]. Walking [380, 381]. Walks [706]. Walter [102]. Wang [3967, 2478, 3654, 3824, 3839, 3918, 3979, 3888]. Warfare [115]. Warping [2807, 2592, 3139]. Wasserstein [3526, 3679, 3472]. Waste [3680]. Waste-free [3680]. Watson [1160, 2493, 1805, 978]. Waudby [3889, 3967, 3890, 3891, 3892, 3893, 3894, 3895, 3887, 3896, 3897, 3898, 3899, 3900, 3901, 3888]. Waudby-Smith [3889, 3967, 3890, 3891, 3892, 3893, 3894, 3895, 3887, 3896, 3897, 3898, 3899, 3900, 3901, 3888]. Waudy [3886]. Waudy-Smith [3886]. Wavelet [2426, 2440, 2635, 2216, 2903, 2276, 3011, 2332, 2784, 2880, 2283, 2518, 2808, 2761, 2886, 2497, 2348, 2916]. Wavelet- [3011]. Wavelet-Based [2880]. Wavelets [2805, 2548]. Way [533, 1617, 615, 2676, 1893, 960, 1552, 1201, 2695, 1337, 3875]. weak [3989, 3988, 3356]. Weakly [1047]. Wear [534]. Wear-Dependent [534]. wearable [3770]. Weibull [1870, 2177, 1209]. Weight [496, 3130, 868]. Weighted [541, 1869, 879, 1251, 540, 2512, 2981, 331, 2509, 438, 2840, 2466, 2380, 2136, 1958, 1708, 484, 1874, 2327, 203, 624, 3067, 3506, 3549, 3648, 2636, 3470]. Weighting [3852, 2374, 3364, 3779, 3546]. Weights [2522]. Welch [382, 745, 535]. Wembley [80]. Weng [3935, 3936]. Wermuth [1904]. Wheat [326]. Where [1611, 3812]. Which [1371, 840, 55, 62, 2091, 1082, 355, 822, 1042, 477, 2097, 1262, 727]. while [3566]. White [1882, 3980]. Whitehouse [3925]. Whitney [3216, 1554]. Whittle [3755]. Whole [1225]. Whose [108, 204]. Width [808]. Wiener [1158, 1291, 1487]. Wijayatunga [3676, 3869]. Wilcoxon [275, 301, 504]. Wind [2787]. Winding [466]. Wishart [111, 670, 312, 1021]. Within [349, 2341, 276, 2713, 2917, 2120, 2775]. Within-Animal [349]. within-Cluster [2917]. Without [1530, 1948, 1765, 180, 1072, 973, 428, 276, 3486, 2898, 3813, 3329, 587, 3308]. Wold [3495]. Wong [3711, 3935, 3936]. Wool [73]. Words [846, 2207]. Work [15, 252]. Working [2580]. Works [80, 89, 451]. world [3801]. Wu [3829, 3936]. Wyse [3980].

X’ [3488, 3872]. Xiaowu [3894]. Xiaoyue [3834]. Xin [3978]. Xinyi [3840]. Xu [3972, 3973, 3974, 3831, 3975, 3976, 3969, 3977, 3970, 3971, 3979, 3980, 3978].

Yanbo [3721]. Yang

[3829, 3975, 3833, 3978]. Yarn [19]. year [3437]. years [3272]. Yi [3979]. Yicong [3976]. Yin [3861]. Ying [3827, 3840]. Yingnian [3936]. Yining [3923]. Yinqiu [3827]. Yongmiao [3926]. Youden [2223, 751, 162]. Yu [3892, 3979]. Yuantong [3894]. Yudong [3923]. Yunxiao [3831]. Yuqi [3827].

ZAP [3769]. Zeng [3828, 3829, 3823, 3830, 3831, 3825, 3821, 3832, 3833, 3834, 3835, 3822, 3836, 3841, 3837, 3826, 3838, 3824, 3839, 3840, 3827]. Zero [3007, 1633, 328, 1711]. Zero-One [328]. Zeroes [1711]. Zeros [861, 3876]. Zhang [3825, 3917, 3840]. Zhao [3829, 3723]. Zheng [3976]. Zhilian [3827]. Zhou [3840]. Zhu [3829, 3926].

References

Anonymous:1934:FM

[1] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 1(1):i–ii, ???? 1934. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983592.

Anonymous:1934:VIa

[2] Anonymous. Volume information. Supplement to the Journal of the Royal Statistical Society, 1(1):iii–iv, ???? 1934. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983591.

Anonymous:1934:NFI

[3] Anonymous. Note on the formation of the Industrial and Agricultural Research Section of the Royal Statistical Society. Supplement to the Journal of the Royal Statistical Society, 1(1):1–3, ???? 1934. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983593.

Anonymous:1934:IMI

[4] Anonymous. Inaugural meeting of the Industrial and Agricultural Research Section, held on Thursday, 23rd November, 1933, Professor M. Greenwood, F.R.S., Vice-President of the Society, in the Chair. Supplement to the Journal of the Royal Statistical Society, 1(1):4, ???? 1934. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983594.

Pickard:1934:ASM

[5] R. H. Pickard. The application of statistical methods to production and research in industry. Supplement to the Journal of the Royal Statistical Society, 1(1):5–25, ???? 1934. CODEN ???? ISSN 1466-6162. URL http:/

/www.jstor.org/stable/2983595.

Wishart:1934:SAR

[6] John Wishart. Statistics in agricultural research. Supplement to the Journal of the Royal Statistical Society, 1(1):26– 61, ???? 1934. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2983596.

McKay:1934:NMH

[7] A. T. McKay. A new method of handling statistical data. Supplement to the Journal of the Royal Statistical Society, 1(1):62–75, ???? 1934. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983597.

Davies:1934:MES

[8] O. L. Davies and E. S. Pearson. Methods of estimating from samples the population standard deviation. Supplement to the Journal of the Royal Statistical Society, 1(1):76–93, ???? 1934. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983598.

Wishart:1934:BAS

[9] John Wishart. Bibliography of agricultural statistics, 1931–33. Supplement to the Journal of the Royal Statistical Society, 1(1):94–106, ???? 1934. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983599.

Pearson:1934:SPI

[10] E. S. Pearson. Sampling problems in industry. Supplement to the Journal of the Royal Statistical Society, 1(2):107– 151, ???? 1934. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2983601.

Wilsdon:1934:DSS

[11] B. H. Wilsdon. Discrimination by specification statistically considered and illustrated by the standard specification for Portland cement. Supplement to the Journal of the Royal Statistical Society, 1(2):152–206, ???? 1934. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983602.

McKay:1934:SB

[12] A. T. McKay. Sampling from batches. Supplement to the Journal of the Royal Statistical Society, 1(2):207–216, ???? 1934. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983603.

Yates:1934:CTI

[13] F. Yates. Contingency tables involving small numbers and the χ^{2} test. Supplement to the Journal of the Royal Statistical Society, 1(2):217–235, ???? 1934. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983604.

Irwin:1934:ICI

[14] J. O. Irwin. On the independence of the constituent items in the analysis of variance. Supplement to the Journal of the Royal Statistical Society, 1(2):236– 251, ???? 1934. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983605.

Bartlett:1934:RWA

[15] M. S. Bartlett. Recent work on the analysis of variance. Supplement to the Journal of the Royal Statistical Society, 1(2):252–255, ???? 1934. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983606.

Anonymous:1934:VIb

[16] Anonymous. Volume information. Supplement to the Journal of the Royal Statistical Society, 1(2):256–257, ???? 1934. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983600.

Anonymous:1935:FMa

[17] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 2(1):i–iv, ???? 1935. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983584.

Grumell:1935:SMI

[18] E. S. Grumell. Statistical methods in industry, with special reference to the sampling of coal and other materials. Supplement to the Journal of the Royal Statistical Society, 2(1):1–26, ???? 1935. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983585.

Tippett:1935:SAS

[19] L. H. C. Tippett. Some applications of statistical methods to the study of variation of quality in the production of cotton yarn. Supplement to the Journal of the Royal Statistical Society, 2(1): 27–62, ???? 1935. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2983586.

Matuszewski:1935:SSQ

[20] T. Matuszewski, J. Neyman, and J. Supinska. Statistical studies in questions of bacteriology. Part I. the accuracy of the “Dilution method”. Supplement to the Journal of the Royal Statistical Society, 2(1):63–82, ????

1935. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983587.

Pearson:1935:URP

[21] E. S. Pearson and Joan Haines. The use of range in place of standard deviation in small samples. Supplement to the Journal of the Royal Statistical Society, 2(1):83–98, ???? 1935. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983588.

Brady:1935:BAA

[22] J. Brady. A biological application of the analysis of co-variance. Supplement to the Journal of the Royal Statistical Society, 2(1):99–106, ???? 1935. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983589.

Anonymous:1935:BMa

[23] Anonymous. Back matter. Supplement to the Journal of the Royal Statistical Society, 2(1):??, ???? 1935. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983590.

Anonymous:1935:FMb

[24] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 2(2):i–iv, ???? 1935. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983636.

Neyman:1935:SPA

[25] J. Neyman, K. Iwaszkiewicz, and St. Kolodziejczyk. Statistical problems in agricultural experimentation. Supplement to the Journal of the Royal Statistical Society, 2(2):107–180, ???? 1935. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983637.

Yates:1935:CE

[26] F. Yates. Complex experiments. Supplement to the Journal of the Royal Statistical Society, 2(2):181–247, ???? 1935. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983638.

Bartlett:1935:CTI

[27] M. S. Bartlett. Contingency table interactions. Supplement to the Journal of the Royal Statistical Society, 2(2):248– 252, ???? 1935. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2983639.

Sukhatme:1935:CTR

[28] P. V. Sukhatme. Contribution to the theory of the representative method. Supplement to the Journal of the Royal Statistical Society, 2(2):253–268, ???? 1935. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983640.

Anonymous:1935:VI

[29] Anonymous. Volume information. Supplement to the Journal of the Royal Statistical Society, 2(2):269–iv, ???? 1935. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983635.

Anonymous:1935:BMb

[30] Anonymous. Back matter. Supplement to the Journal of the Royal Statistical Society, 2(2):??, ???? 1935. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983641.

Anonymous:1936:FMa

[31] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 3(1):i–iv, ???? 1936. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983673.

Anonymous:1936:FMS

[32] Anonymous. First meeting of the Session 1935–36, 28th November, 1935. Supplement to the Journal of the Royal Statistical Society, 3(1):1, ???? 1936. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983674.

Jennett:1936:ASP

[33] W. J. Jennett and B. P. Dudding. The application of statistical principles to an industrial problem. Supplement to the Journal of the Royal Statistical Society, 3(1):1–28, ???? 1936. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983675.

Welch:1936:SRR

[34] B. L. Welch. The specification of rules for rejecting too variable a product, with particular reference to an electric lamp problem. Supplement to the Journal of the Royal Statistical Society, 3 (1):29–48, ???? 1936. CODEN ???? ISSN 1466-6162. URL http://www.

jstor.org/stable/2983676.

Cochran:1936:SAF

[35] W. G. Cochran. The statistical analysis of field counts of diseased plants. Supplement to the Journal of the Royal Statistical Society, 3(1):49–67, ???? 1936. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983677.

Bartlett:1936:SRT

[36] M. S. Bartlett. The square root transformation in analysis of variance. Supplement to the Journal of the Royal Statistical Society, 3(1):68–78, ???? 1936. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983678.

Wishart:1936:TSA

[37] John Wishart. Tests of significance in analysis of covariance. Supplement to the Journal of the Royal Statistical Society, 3(1):79–82, ???? 1936. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983679.

Neyman:1936:C

[38] J. Neyman. Correspondence. Supplement to the Journal of the Royal Statistical Society, 3(1):83–85, ???? 1936. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983680.

Anonymous:1936:BMa

[39] Anonymous. Back matter. Supplement to the Journal of the Royal Statistical Society, 3(1):??, ???? 1936. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983681.

Anonymous:1936:FMb

[40] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 3(2):i–iv, ???? 1936. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983665.

Anonymous:1936:VI

[41] Anonymous. Volume information. Supplement to the Journal of the Royal Statistical Society, 3(2):i–204, ???? 1936. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983664.

Comrie:1936:IIS

[42] L. J. Comrie. Inverse interpolation and scientific applications of the National Accounting Machine. Supplement to the Journal of the Royal Statistical Society, 3(2):87–114, ???? 1936. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983666.

Anonymous:1936:COL

[43] Anonymous. Co-operation in largescale experiments. Supplement to the Journal of the Royal Statistical Society, 3(2):115–136, ???? 1936. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983667.

Gould:1936:SMA

[44] C. E. Gould and W. M. Hampton. Statistical methods applied to the manufacture of spectacle glasses. Supplement to the Journal of the Royal Statistical Society, 3(2):137–177, ???? 1936. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983668.

Geary:1936:DSR

[45] R. C. Geary. The distribution of “Student’s” ratio for non-normal samples. Supplement to the Journal of the Royal Statistical Society, 3(2):178–184, ???? 1936. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983669.

Bartlett:1936:SNI

[46] M. S. Bartlett. Some notes on insecticide tests in the laboratory and in the field. Supplement to the Journal of the Royal Statistical Society, 3(2):185–194, ???? 1936. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983670.

Barnard:1936:ECA

[47] M. M. Barnard. An enumeration of the confounded arrangements in the 2 × 2 × 2 ⋯ factorial designs. Supplement to the Journal of the Royal Statistical Society, 3(2):195–202, ???? 1936. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983671.

Anonymous:1936:BMb

[48] Anonymous. Back matter. Supplement to the Journal of the Royal Statistical Society, 3(2):??, ???? 1936. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983672.

Anonymous:1937:FMa

[49] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 4(1):i–iv, ???? 1937. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2984118.

Hartley:1937:FMI

[50] Percival Hartley. First meeting of the Industrial and Agricultural Section, Session 1936–37 November 26th, 1936. Supplement to the Journal of the Royal Statistical Society, 4(1):1, ???? 1937. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2984119.

Irwin:1937:SMA

[51] J. O. Irwin. Statistical method applied to biological assays. Supplement to the Journal of the Royal Statistical Society, 4(1):1–60, ???? 1937. CODEN ???? ISSN 1466-6162. URL http:/

/www.jstor.org/stable/2984120.

Bayes:1937:SCV

[52] A. W. Bayes. Some considerations of the variability of cotton cloth strength. Supplement to the Journal of the Royal Statistical Society, 4(1):61–93, ???? 1937. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2984121. See notes [53].

Pearson:1937:NSS

[53] E. S. Pearson and B. L. Welch. Notes on some statistical problems raised in Mr. Bayes’s paper. Supplement to the Journal of the Royal Statistical Society, 4(1):94–101, ???? 1937. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2984122. See

[52].

Cochran:1937:PAA

[54] W. G. Cochran. Problems arising in the analysis of a series of similar experiments. Supplement to the Journal of the Royal Statistical Society, 4(1):102– 118, ???? 1937. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2984123.

Pitman:1937:STWa

[55] E. J. G. Pitman. Significance tests which may be applied to samples from any populations. Supplement to the Journal of the Royal Statistical Society, 4(1):119–130, ???? 1937. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2984124.

Bartlett:1937:SSA

[56] M. S. Bartlett. Sub-sampling for attributes. Supplement to the Journal of the Royal Statistical Society, 4(1):131– 135, ???? 1937. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2984125.

Anonymous:1937:BMa

[57] Anonymous. Back matter. Supplement to the Journal of the Royal Statistical Society, 4(1):??, ???? 1937. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2984126.

Anonymous:1937:FMb

[58] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 4(2):i–iv, ???? 1937. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983643.

Bartlett:1937:SES

[59] M. S. Bartlett. Some examples of statistical methods of research in agriculture and applied biology. Supplement to the Journal of the Royal Statistical Society, 4(2):137–183, ???? 1937. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983644.

vanRest:1937:ESM

[60] E. D. van Rest. Examples of statistical methods in forest products research. Supplement to the Journal of the Royal Statistical Society, 4(2):184– 209, ???? 1937. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2983645.

Comrie:1937:AHE

[61] L. J. Comrie, G. B. Hey, and H. G. Hudson. The application of Hollerith equipment to an agricultural investigation. Supplement to the Journal of the Royal Statistical Society, 4(2):210–224, ???? 1937. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983646.

Pitman:1937:STWb

[62] E. J. G. Pitman. Significance tests which May be applied to samples from any populations. II. The correlation coefficient test. Supplement to the Journal of the Royal Statistical Society, 4 (2):225–232, ???? 1937. CODEN ???? ISSN 1466-6162. URL http://www.

jstor.org/stable/2983647.

Cochran:1937:CUT

[63] W. G. Cochran. A catalogue of uniformity trial data. Supplement to the Journal of the Royal Statistical Society, 4(2):233–253, ???? 1937. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983648.

Anonymous:1937:VI

[64] Anonymous. Volume information. Supplement to the Journal of the Royal Statistical Society, 4(2):255–iv, ???? 1937. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983642.

Anonymous:1937:BMb

[65] Anonymous. Back matter. Supplement to the Journal of the Royal Statistical Society, 4(2):??, ???? 1937. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983649.

Anonymous:1938:FMa

[66] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 5(1):i–viii, ???? 1938. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983578.

Irwin:1938:CER

[67] J. O. Irwin. Crop estimation and its relation to agricultural meteorology. Supplement to the Journal of the Royal Statistical Society, 5(1):1–45, ???? 1938. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983579.

Simpson:1938:EMH

[68] T. Wake Simpson. Experimental methods and human nutrition. Supplement to the Journal of the Royal Statistical Society, 5(1):46–69, ???? 1938. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983580.

Yates:1938:GER

[69] F. Yates. The gain in efficiency resulting from the use of balanced designs. Supplement to the Journal of the Royal Statistical Society, 5(1):70– 74, ???? 1938. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2983581.

Sukhatme:1938:DSP

[70] P. V. Sukhatme. On the distribution of χ^{2} in samples of the Poisson series. Supplement to the Journal of the Royal Statistical Society, 5(1):75– 79, ???? 1938. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2983582.

Hartley:1938:SLS

[71] H. O. Hartley. Studentization and large-sample theory. Supplement to the Journal of the Royal Statistical Society, 5(1):80–88, ???? 1938. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983583.

Anonymous:1938:FMb

[72] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 5(2):i–viii, ???? 1938. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983651.

Daniels:1938:SPS

[73] H. E. Daniels. Some problems of statistical interest in wool research. Supplement to the Journal of the Royal Statistical Society, 5(2):89–128, ???? 1938. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983652.

Hoblyn:1938:SVK

[74] T. N. Hoblyn. A study of the variation in keeping quality of apples in store: As illustrated by the behaviour of the variety McIntosh Red from an Ontario apple orchard. Supplement to the Journal of the Royal Statistical Society, 5 (2):129–170, ???? 1938. CODEN ???? ISSN 1466-6162. URL http://www.

jstor.org/stable/2983653.

Cochran:1938:OAI

[75] W. G. Cochran. The omission or addition of an independent variate in multiple linear regression. Supplement to the Journal of the Royal Statistical Society, 5(2):171–176, ???? 1938. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983654.

Yates:1938:OFT

[76] F. Yates. Orthogonal functions and tests of significance in the analysis of variance. Supplement to the Journal of the Royal Statistical Society, 5(2):177– 180, ???? 1938. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2983655.

Anonymous:1938:VI

[77] Anonymous. Volume information. Supplement to the Journal of the Royal Statistical Society, 5(2):181–iv, ???? 1938. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983650.

Anonymous:1939:FMa

[78] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 6(1):i–viii, ???? 1939. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983619.

Wishart:1939:STA

[79] John Wishart. Statistical treatment of animal experiments. Supplement to the Journal of the Royal Statistical Society, 6(1):1–22, ???? 1939. CODEN ???? ISSN 1466-6162. URL http:/

                                        /www.jstor.org/stable/2983620.

Anonymous:1939:EMI

[80] Anonymous. Evening meeting of the Industrial and Agricultural Research Section of the Royal Statistical Society held at the Wembley Glass Works and Research Laboratories of the General Electric Co., Ltd., on December 15th, 1938. Supplement to the Journal of the Royal Statistical Society, 6(1):23– 24, ???? 1939. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983621.

Newland:1939:SCQ

[81] W. F. Newland and E. E. Neal. Statistical control of the quality of telephone service. Supplement to the Journal of the Royal Statistical Society, 6(1):25– 50, ???? 1939. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983622.

Kendall:1939:SPR

[82] M. G. Kendall and B. Babington-Smith. Second paper on random sampling numbers. Supplement to the Journal of the Royal Statistical Society, 6 (1):51–61, ???? 1939. CODEN ???? ISSN 1466-6162. URL http://www.

jstor.org/stable/2983623.

Vickery:1939:DRS

[83] C. W. Vickery. On drawing a random sample from a set of punched cards. Supplement to the Journal of the Royal Statistical Society, 6(1):62– 66, ???? 1939. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983624.

Yates:1939:ALS

[84] F. Yates and R. W. Hale. The analysis of Latin squares when two or more rows, columns, or treatments are missing. Supplement to the Journal of the Royal Statistical Society, 6(1):67– 79, ???? 1939. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983625.

Jennett:1939:CPD

[85] W. J. Jennett and B. L. Welch. The control of proportion defective as judged by a single quality characteristic varying on a continuous scale. Supplement to the Journal of the Royal Statistical Society, 6(1):80–88, ???? 1939. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983626.

Bishop:1939:NCM

[86] D. J. Bishop and U. S. Nair. A note on certain methods of testing for the homogeneity of a set of estimated variances. Supplement to the Journal of the Royal Statistical Society, 6(1):89– 99, ???? 1939. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983627.

Anonymous:1939:BM

[87] Anonymous. Back matter. Supplement to the Journal of the Royal Statistical Society, 6(1):??, ???? 1939. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983628.

Anonymous:1939:FMb

[88] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 6(2):i–viii, ???? 1939. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983683.

Anonymous:1939:MHW

[89] Anonymous. Meeting held at the works of the rockware glass syndicate, limited, greenford, March 30th, 1939. Supplement to the Journal of the Royal Statistical Society, 6(2):101, ???? 1939. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983684.

Anonymous:1939:PML

[90] Anonymous. Provincial meeting at Leeds. Supplement to the Journal of the Royal Statistical Society, 6(2):102– 103, ???? 1939. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983685.

Cochran:1939:LTA

[91] W. G. Cochran. Long-term agricultural experiments. Supplement to the Journal of the Royal Statistical Society, 6 (2):104–148, ???? 1939. CODEN ???? ISSN 1466-6162. URL http://www.

jstor.org/stable/2983686.

Ineson:1939:AES

[92] J. L. Ineson. The application of elementary statistical methods to problems encountered in the operation of electricity generating stations. Supplement to the Journal of the Royal Statistical Society, 6(2):149–168, ???? 1939. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983687.

Bartlett:1939:SED

[93] M. S. Bartlett. The standard errors of discriminant function coefficients. Supplement to the Journal of the Royal Statistical Society, 6(2):169–173, ???? 1939. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983688.

Irwin:1939:MLM

[94] J. O. Irwin and E. A. Cheeseman. On the maximum-likelihood method of determining dosage-response curves and approximations to the median-effective dose, in cases of a quantal response.

Supplement to the Journal of the Royal Statistical Society, 6(2):174–185, ???? 1939. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983689.

Daniels:1939:ECV

[95] H. E. Daniels. The estimation of components of variance. Supplement to the Journal of the Royal Statistical Society, 6(2):186–197, ???? 1939. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983690.

Anonymous:1939:VI

[96] Anonymous. Volume information. Supplement to the Journal of the Royal Statistical Society, 6(2):199–iv, ???? 1939. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983682.

Anonymous:1941:FMa

[97] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 7(1):iii–vi, ???? 1941. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983629.

Fieller:1941:BSI

[98] E. C. Fieller. The biological standardization of insulin. Supplement to the Journal of the Royal Statistical Society, 7(1):1–64, ???? 1941. CODEN ???? ISSN 1466-6162. URL http:/

                                        /www.jstor.org/stable/2983630.

Garwood:1941:ATP

[99] F. Garwood. An application of the theory of probability to the operation of vehicular-controlled traffic signals. Supplement to the Journal of the Royal Statistical Society, 7(1):65– 77, ???? 1941. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983631.

Todd:1941:NRA

[100] H. Todd. A note on random associations in a square point lattice. Supplement to the Journal of the Royal Statistical Society, 7(1):78–82, ???? 1941. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983632.

Kendall:1941:NDQ

[101] M. G. Kendall. Note on the distribution of quantiles for large samples. Supplement to the Journal of the Royal Statistical Society, 7(1):83– 85, ???? 1941. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983633.

W:1941:BRB

[102] B. L. W. Book review: Statistical Method from the Viewpoint of Quality Control. by Walter A. Shewhart. Supplement to the Journal of the Royal Statistical Society, 7(1):86–87, ???? 1941. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983634.

Chambers:1941:TOI

[103] E. G. Chambers and G. Udny Yule. Theory and observation in the investigation of accident causation. Supplement to the Journal of the Royal Statistical Society, 7(2):89–109, ???? 1941. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983658.

Campbell:1941:RPM

[104] N. R. Campbell. The replacement of perishable members of a continually operating system. Supplement to the Journal of the Royal Statistical Society, 7(2):110–130, ???? 1941. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983659.

Spencer-Smith:1941:TSM

[105] J. L. Spencer-Smith and H. A. C. Todd. A time series Met with in textile research. Supplement to the Journal of the Royal Statistical Society, 7(2):131– 145, ???? 1941. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2983660.

Daniels:1941:MIC

[106] H. E. Daniels. A method of improving certain routine measurements. Supplement to the Journal of the Royal Statistical Society, 7(2):146–150, ???? 1941. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983661.

Todd:1941:NSC

[107] H. A. C. Todd. A note on systematic coding for card sorting systems. Supplement to the Journal of the Royal Statistical Society, 7(2):151–154, ???? 1941. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983662.

Finney:1941:DVW

[108] D. J. Finney. On the distribution of a variate whose logarithm is normally distributed. Supplement to the Journal of the Royal Statistical Society, 7 (2):155–161, ???? 1941. CODEN ????

ISSN 1466-6162. URL http://www.

jstor.org/stable/2983663.

Anonymous:1941:VI

[109] Anonymous. Volume information. Supplement to the Journal of the Royal Statistical Society, 7(2):163–164, ???? 1941. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983656.

Anonymous:1941:FMb

[110] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 7(2):??, ???? 1941. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983657.

Anonymous:1946:PFM

[111] Anonymous. Proceedings of the first meeting of The Research Section of the Royal Statistical Society, session 1945–1946. held on Tuesday, December 4th, 1945, Dr. J. Wishart in the Chair. Supplement to the Journal of the Royal Statistical Society, 8(1):1, ???? 1946. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983609.

Barnard:1946:STI

[112] G. A. Barnard. Sequential tests in industrial statistics. Supplement to the Journal of the Royal Statistical Society, 8(1):1–26, ???? 1946. CODEN ???? ISSN 1466-6162. URL http:/

/www.jstor.org/stable/2983610.

Bartlett:1946:TSS

[113] M. S. Bartlett. On the theoretical specification and sampling properties of autocorrelated time-series. Supplement to the Journal of the Royal Statistical Society, 8(1):27–41, ???? 1946. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983611.

Foster:1946:SIA

[114] G. A. R. Foster. Some instruments for the analysis of time series and their application to textile research. Supplement to the Journal of the Royal Statistical Society, 8(1):42–61, ???? 1946. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983612.

Cunningham:1946:RPP

[115] L. B. C. Cunningham and W. R. B. Hynd. Random processes in problems of air warfare. Supplement to the Journal of the Royal Statistical Society, 8 (1):62–97, ???? 1946. CODEN ???? ISSN 1466-6162. URL http://www.

jstor.org/stable/2983613.

Burman:1946:SSF

[116] J. P. Burman. Sequential sampling formulae for a binomial population. Supplement to the Journal of the Royal Statistical Society, 8(1):98–103, ???? 1946. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983614.

Stockman:1946:SPC

[117] C. M. Stockman and P. Armitage. Some properties of closed sequential schemes. Supplement to the Journal of the Royal Statistical Society, 8(1):104– 112, ???? 1946. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983615.

Bartlett:1946:MPT

[118] M. S. Bartlett. A modified probit technique for small probabilities. Supplement to the Journal of the Royal Statistical Society, 8(1):113–117, ???? 1946. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983616.

Kempthorne:1946:ASE

[119] O. Kempthorne. The analysis of a series of experiments by the use of punched cards. Supplement to the Journal of the Royal Statistical Society, 8(1):118–127, ???? 1946. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983617.

Bartlett:1946:SAV

[120] M. S. Bartlett and D. G. Kendall. The statistical analysis of varianceheterogeneity and the logarithmic transformation. Supplement to the Journal of the Royal Statistical Society, 8(1):128–138, ???? 1946. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983618.

Anonymous:1946:FM

[121] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 8(1):??, ???? 1946. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983608.

Anonymous:1946:VI

[122] Anonymous. Volume information. Supplement to the Journal of the Royal Statistical Society, 8(1):??, ???? 1946. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983607.

Vernon:1946:SMS

[123] P. E. Vernon. Statistical methods in the selection of Navy and Army personnel. Supplement to the Journal of the Royal Statistical Society, 8(2):139– 153, ???? 1946. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983558.

Hartley:1946:ASC

[124] H. O. Hartley. The application of some commercial calculating machines to certain statistical calculations. Supplement to the Journal of the Royal Statistical Society, 8(2):154–183, ???? 1946. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983559.

Bayley:1946:ENI

[125] G. V. Bayley and J. M. Hammersley. The “Effective” number of independent observations in an autocorrelated time series. Supplement to the Journal of the Royal Statistical Society, 8(2):184– 197, ???? 1946. CODEN ???? ISSN

1466-6162. URL http://www.jstor.
org/stable/2983560.

Vajda:1946:ASN

[126] S. Vajda. Average sampling numbers from finite lote. Supplement to the Journal of the Royal Statistical Society, 8(2):198–201, ???? 1946. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983561.

Wise:1946:UNB

[127] M. E. Wise. The use of the negative binomial distribution in an industrial sampling problem. Supplement to the Journal of the Royal Statistical Society, 8(2):202–211, ???? 1946. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983562.

Richardson:1946:TLC

[128] J. T. Richardson. A table of Lagrangian coefficients for logarithmic interpolation of standard statistical tables to obtain other probability levels. Supplement to the Journal of the Royal Statistical Society, 8(2):212–215, ???? 1946. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983563.

Anscombe:1946:LSR

[129] F. J. Anscombe. Linear sequential rectifying inspection for controlling fraction defective. Supplement to the Journal of the Royal Statistical Society, 8 (2):216–222, ???? 1946. CODEN ???? ISSN 1466-6162. URL http://www.

jstor.org/stable/2983564.

Francis:1946:DSS

[130] V. J. Francis. On the distribution of the sum of n sample values drawn from a truncated normal population. Supplement to the Journal of the Royal Statistical Society, 8(2):223–232, ???? 1946. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983565.

Harding:1946:SCA

[131] E. W. Harding. Statistical control applied to high duty iron production. Supplement to the Journal of the Royal Statistical Society, 8(2):233–243, ???? 1946. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983566.

Grimsey:1946:URS

[132] A. H. R. Grimsey. Ultimate risks in sampling inspection. Supplement to the Journal of the Royal Statistical Society, 8(2):244–250, ???? 1946. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983567.

Stone:1947:IBT

[133] Richard Stone. On the interdependence of blocks of transactions. Supplement to the Journal of the Royal Statistical Society, 9(1):1–45, ???? 1947. CODEN ???? ISSN 1466-6162. URL http://

                                        www.jstor.org/stable/2983570.

Finney:1947:PBA

[134] D. J. Finney. The principles of biological assay. Supplement to the Journal of the Royal Statistical Society, 9(1):46– 91, ???? 1947. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2983571.

Moran:1947:RDI

[135] P. A. P. Moran. The random division of an interval. Supplement to the Journal of the Royal Statistical Society, 9 (1):92–98, ???? 1947. CODEN ???? ISSN 1466-6162. URL http://www.

                                        jstor.org/stable/2983572. See cor-

rigendum [3992].

Finney:1947:SAS

[136] D. J. Finney. The significance of associations in a square point lattice. Supplement to the Journal of the Royal Statistical Society, 9(1):99–103, ???? 1947. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983573.

Spencer-Smith:1947:OPM

[137] J. L. Spencer-Smith. The oscillatory properties of the moving average. Supplement to the Journal of the Royal Statistical Society, 9(1):104–113, ???? 1947. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983574.

Slater:1947:FAM

[138] Patrick Slater. The factor analysis of a matrix of 2×2 tables. Supplement to the Journal of the Royal Statistical Society, 9(1):114–127, ???? 1947. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983575.

Rao:1947:FED

[139] C. Radhakrishna Rao. Factorial experiments derivable from combinatorial arrangements of arrays. Supplement to the Journal of the Royal Statistical Society, 9(1):128–139, ???? 1947. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983576.

Anonymous:1947:EMA

[140] Anonymous. Exhibition of mechanical aids to statistical computation. Supplement to the Journal of the Royal Statistical Society, 9(1):140, ???? 1947. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2983577.

Anonymous:1947:FM

[141] Anonymous. Front matter. Supplement to the Journal of the Royal Statistical Society, 9(1):??, ???? 1947. CODEN ???? ISSN 1466-6162. URL http://

www.jstor.org/stable/2983569.

Anonymous:1947:VIa

[142] Anonymous. Volume information. Supplement to the Journal of the Royal Statistical Society, 9(1):??, ???? 1947. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2983568.

Vajda:1947:SIC

[143] S. Vajda. Statistical investigation of casualties suffered by certain types of vessels. Supplement to the Journal of the Royal Statistical Society, 9(2):141– 175, ???? 1947. CODEN ???? ISSN

                                        1466-6162. URL http://www.jstor.
                                        org/stable/2984112.

Bartlett:1947:MA

[144] M. S. Bartlett. Multivariate analysis. Supplement to the Journal of the Royal Statistical Society, 9(2):176–197, ???? 1947. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2984113.

Anscombe:1947:MDS

[145] F. J. Anscombe, H. J. Godwin, and R. L. Plackett. Methods of deferred sentencing in testing the fraction defective of a continuous output. Supplement to the Journal of the Royal Statistical Society, 9(2):198–217, ???? 1947. CODEN ???? ISSN 1466-

                                        6162. URL http://www.jstor.org/
                                        stable/2984114.

Lindley:1947:RLL

[146] D. V. Lindley. Regression lines and the linear functional relationship. Supplement to the Journal of the Royal Statistical Society, 9(2):218–244, ???? 1947. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2984115.

Daniels:1947:GCH

[147] H. E. Daniels. Grouping corrections for high autocorrelations. Supplement to the Journal of the Royal Statistical Society, 9(2):245–249, ???? 1947. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2984116.

Armitage:1947:SST

[148] P. Armitage. Some sequential tests of Student’s hypothesis. Supplement to the Journal of the Royal Statistical Society, 9(2):250–263, ???? 1947. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2984117.

Anonymous:1947:VIb

[149] Anonymous. Volume information. Supplement to the Journal of the Royal Statistical Society, 9(2):265–266, ???? 1947. CODEN ???? ISSN 1466-

6162. URL http://www.jstor.org/
stable/2984111.

Orcutt:1948:SAN

[150] G. H. Orcutt and J. O. Irwin. A study of the autoregressive nature of the time series used for Tinbergen’s model of the economic system of the United States, 1919–1932. Journal of the Royal Statistical Society. Series B (Methodological), 10(1):1–53, ???? 1948. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983795.

Stevens:1948:CG

[151] W. L. Stevens. Control by gauging. Journal of the Royal Statistical Society. Series B (Methodological), 10(1):

54–108, ???? 1948. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

                                        jstor.org/stable/2983796.

Nair:1948:CAF

[152] K. R. Nair and C. R. Rao. Confounding in asymmetrical factorial experiment. Journal of the Royal Statistical Society. Series B (Methodological), 10(1): 109–131, ???? 1948. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2983797.

Anderson:1948:ADR

[153] T. W. Anderson. The asymptotic distributions of the roots of certain determinantal equations. Journal of the Royal Statistical Society. Series B (Methodological), 10(1):132–139, ???? 1948. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983798.

Geary:1948:SRB

[154] R. C. Geary. Studies in relations between economic time series. Journal of the Royal Statistical Society. Series B (Methodological), 10(1):140–158, ???? 1948. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983799.

Anonymous:1948:FM

[155] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 10(1):??, ???? 1948. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983794.

Anonymous:1948:VIa

[156] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 10(1):

??, ???? 1948. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983793.

Rao:1948:UMM

[157] C. Radhakrishna Rao. The utilization of multiple measurements in problems of biological classification. Journal of the Royal Statistical Society. Series B (Methodological), 10(2):159–203, ???? 1948. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983775.

Champernowne:1948:STA

[158] D. G. Champernowne. Sampling theory applied to autoregressive sequences. Journal of the Royal Statistical Society. Series B (Methodological), 10(2):204–242, ???? 1948. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983776.

Moran:1948:ISM

[159] P. A. P. Moran. The interpretation of statistical maps. Journal of the Royal Statistical Society. Series B (Methodological), 10(2):243–251, ???? 1948. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983777.

Pearce:1948:RBI

[160] S. C. Pearce. Randomized blocks with interchanged and substituted plots. Journal of the Royal Statistical Society. Series B (Methodological), 10(2): 252–256, ???? 1948. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983778.

Skellam:1948:PDD

[161] J. G. Skellam. A probability distribution derived from the binomial distribution by regarding the probability of success as variable between the sets of trials. Journal of the Royal Statistical Society. Series B (Methodological), 10 (2):257–261, ???? 1948. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983779.

Smith:1948:CYS

[162] C. A. B. Smith and H. O. Hartley. The construction of Youden squares. Journal of the Royal Statistical Society. Series B (Methodological), 10(2):262–263, ???? 1948. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983780.

Wise:1948:IBF

[163] M. E. Wise. The incomplete beta function and the incomplete gamma function: An acknowledgment. Journal of the Royal Statistical Society. Series B (Methodological), 10(2):264, ???? 1948. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983781.

Anonymous:1948:VIb

[164] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 10(2): 265, ???? 1948. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

jstor.org/stable/2983774.

Anonymous:1949:SSP

[165] Anonymous. Some statistical problems arising in genetics. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):1–14, ???? 1949. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983693.

Boag:1949:MLE

[166] John W. Boag. Maximum likelihood estimates of the proportion of patients cured by cancer therapy. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):15–53, ???? 1949. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983694.

Kosten:1949:AMP

[167] L. Kosten, J. R. Manning, and F. Garwood. On the accuracy of measurements of probabilities of loss in telephone systems. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):54–67, ???? 1949. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983695.

Quenouille:1949:ATC

[168] M. H. Quenouille. Approximate tests of correlation in time-series. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):68–84, ???? 1949. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983696.

Shone:1949:RBS

[169] K. J. Shone. Relations between the standard deviation and the distribution of range in non-normal populations. Journal of the Royal Statistical Society. Series B (Methodological), 11 (1):85–88, ???? 1949. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983697.

Jowett:1949:CSS

[170] G. H. Jowett. The calculation of sums of squares and products on a desk calculating machine. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):89–90, ???? 1949. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983698.

Benson:1949:NEM

[171] F. Benson. A note on the estimation of mean and standard deviation from quantiles. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):91–100, ???? 1949. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2983699.

Cox:1949:URS

[172] D. R. Cox. The use of the range in sequential analysis. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):101–114, ???? 1949. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983700.

Anonymous:1949:FM

[173] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):??, ???? 1949. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983692.

Anonymous:1949:VIa

[174] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 11(1): ??, ???? 1949. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983691.

Barnard:1949:SI

[175] G. A. Barnard. Statistical inference. Journal of the Royal Statistical Society. Series B (Methodological), 11(2): 115–149, ???? 1949. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984075.

Moyal:1949:SPS

[176] J. E. Moyal. Stochastic processes and statistical physics. Journal of the Royal Statistical Society. Series B (Methodological), 11(2):150–210, ???? 1949. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984076.

Bartlett:1949:SES

[177] M. S. Bartlett. Some evolutionary stochastic processes. Journal of the Royal Statistical Society. Series B (Methodological), 11(2):211–229, ???? 1949. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984077.

Kendall:1949:SPP

[178] David G. Kendall. Stochastic processes and population growth. Journal of the Royal Statistical Society. Series B (Methodological), 11(2):230–282, ???? 1949. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2984078. See errata [194].

Nanda:1949:SED

[179] D. N. Nanda. The standard errors of discriminant function coefficients in plant-breeding experiments. Journal of the Royal Statistical Society. Series B (Methodological), 11(2):283–290, ???? 1949. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984079.

Skellam:1949:DMS

[180] J. G. Skellam. The distribution of the moment statistics of samples drawn without replacement from a finite population. Journal of the Royal Statistical Society. Series B (Methodological), 11 (2):291–296, ???? 1949. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984080.

Wold:1949:LST

[181] H. Wold. A large-sample test for moving averages. Journal of the Royal Statistical Society. Series B (Methodological), 11(2):297–305, ???? 1949. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984081.

Anonymous:1949:VIb

[182] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 11(2): 307, ???? 1949. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

                                        jstor.org/stable/2984074.

Cox:1950:OFI

[183] Gertrude M. Cox. The organization and functions of the Institute of Statistics of the University of North Carolina. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):1–18, ???? 1950. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983830.

Ross:1950:PPP

[184] Alan S. C. Ross. Philological probability problems. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):19–59, ???? 1950. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983831.

Kendall:1950:PFA

[185] M. G. Kendall. Part i: Factor analysis as a statistical technique. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):60–73, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983832.

Smith:1950:PIE

[186] B. Babington Smith. Part II: An evaluation of factor analysis from the point of view of a psychologist. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):73–94, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983833.

Tintner:1950:SFR

[187] Gerhard Tintner. Some formal relations in multivariate analysis. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):95–101, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983834.

Walker:1950:NGL

[188] A. M. Walker. Note on a generalization of the large sample goodness of fit test for linear autoregressive schemes. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):102–107, ???? 1950. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983835.

Bartlett:1950:EQT

[189] M. S. Bartlett and P. H. Diananda. Extensions of Quenouille’s test for autoregressive schemes. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):108–115, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983836.

Kendall:1950:ARS

[190] David G. Kendall. An artificial realization of a simple “Birth-and-Death” process. Journal of the Royal Statistical Society. Series B (Methodological), 12 (1):116–119, ???? 1950. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983837.

Fox:1950:PMS

[191] L. Fox. Practical methods for the solution of linear equations and the inversion of matrices. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):120–136, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983838.

Armitage:1950:SAM

[192] P. Armitage. Sequential analysis with more than two alternative hypotheses, and its relation to discriminant function analysis. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):137–144, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983839.

Ashcroft:1950:PSM

[193] H. Ashcroft. The productivity of several machines under the care of one operator. Journal of the Royal Statistical Society. Series B (Methodological), 12 (1):145–151, ???? 1950. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983840.

Anonymous:1950:ESP

[194] Anonymous. Errata: Stochastic processes and population growth. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):??, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983829. See [178].

Anonymous:1950:FM

[195] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 12(1):??, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983828.

Anonymous:1950:VIa

[196] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 12(1): ??, ???? 1950. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983827.

Moran:1950:RDR

[197] P. A. P. Moran. Recent developments in ranking theory. Journal of the Royal Statistical Society. Series B (Methodological), 12(2):153–162, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983978.

Whitfield:1950:URM

[198] J. W. Whitfield. Uses of the ranking method in psychology. Journal of the Royal Statistical Society. Series B (Methodological), 12(2):163–170, ????

1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983979.

Daniels:1950:RCP

[199] H. E. Daniels. Rank correlation and population models. Journal of the Royal Statistical Society. Series B (Methodological), 12(2):171–191, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983980.

Hammersley:1950:ERP

[200] J. M. Hammersley. On estimating restricted parameters. Journal of the Royal Statistical Society. Series B (Methodological), 12(2):192–240, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983981.

Patterson:1950:SSO

[201] H. D. Patterson. Sampling on successive occasions with partial replacement of units. Journal of the Royal Statistical Society. Series B (Methodological), 12 (2):241–255, ???? 1950. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983982.

Quenouille:1950:CDA

[202] M. H. Quenouille. Computational devices in the application of least squares. Journal of the Royal Statistical Society. Series B (Methodological), 12(2): 256–272, ???? 1950. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983983.

Tintner:1950:TLR

[203] Gerhard Tintner. A test for linear relations between weighted regression coefficients. Journal of the Royal Statistical Society. Series B (Methodological), 12 (2):273–277, ???? 1950. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983984.

Kendall:1950:RFA

[204] David G. Kendall. Random fluctuations in the age-distribution of a population whose development is controlled by the simple “Birth-and-Death” process. Journal of the Royal Statistical Society. Series B (Methodological), 12 (2):278–285, ???? 1950. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983985.

Grundy:1950:RRQ

[205] P. M. Grundy and M. J. R. Healy. Restricted randomization and quasi-Latin squares. Journal of the Royal Statistical Society. Series B (Methodological), 12 (2):286–291, ???? 1950. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983986.

Moran:1950:CRT

[206] P. A. P. Moran. A curvilinear ranking test. Journal of the Royal Statistical Society. Series B (Methodological), 12 (2):292–295, ???? 1950. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983987.

Finney:1950:TNU

[207] D. J. Finney. Two new uses of the behrens–Fisher distribution. Journal of the Royal Statistical Society. Series B (Methodological), 12(2):296–300, ???? 1950. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983988.

Walker:1950:NSS

[208] A. M. Walker. Note on sequential sampling formulae for a binomial population. Journal of the Royal Statistical Society. Series B (Methodological), 12 (2):301–307, ???? 1950. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983989.

Anonymous:1950:VIb

[209] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 12(2): 309–310, ???? 1950. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2983977.

Box:1951:EAO

[210] G. E. P. Box and K. B. Wilson. On the experimental attainment of optimum conditions. Journal of the Royal Statistical Society. Series B (Methodological), 13(1):1–45, ???? 1951. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2983966.

Barnard:1951:TI

[211] G. A. Barnard. The theory of information. Journal of the Royal Statistical Society. Series B (Methodological), 13 (1):46–64, ???? 1951. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983967.

Benson:1951:PMR

[212] F. Benson and D. R. Cox. The productivity of machines requiring attention at random intervals. Journal of the Royal Statistical Society. Series B (Methodological), 13(1):65–82, ???? 1951. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983968.

Fox:1951:MPM

[213] L. Fox and J. G. Hayes. More practical methods for the inversion of matrices. Journal of the Royal Statistical Society. Series B (Methodological), 13(1):83–91, ???? 1951. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983969.

Rushton:1951:LSF

[214] S. Rushton. On least squares fitting by orthonormal polynomials using the Choleski method. Journal of the Royal Statistical Society. Series B (Methodological), 13(1):92–99, ???? 1951. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983970.

Woolf:1951:CIM

[215] Barnet Woolf. Computation and interpretation of multiple regressions. Journal of the Royal Statistical Society. Series B (Methodological), 13(1):100–119, ???? 1951. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983971.

Schutzenberger:1951:EPT

[216] M. P. Schutzenberger. An extension problem in the theory of incomplete block designs. Journal of the Royal Statistical Society. Series B (Methodological), 13(1):120–125, ???? 1951. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983972.

Thompson:1951:FDS

[217] H. R. Thompson and I. D. Dick. Factorial designs in small blocks derived from orthogonal Latin squares. Journal of the Royal Statistical Society. Series B (Methodological), 13(1):126–130, ???? 1951. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983973.

Ramakrishnan:1951:SSS

[218] Alladi Ramakrishnan. Some simple stochastic processes. Journal of the Royal Statistical Society. Series B (Methodological), 13(1):131–140, ???? 1951. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983974.

Moran:1951:EME

[219] P. A. P. Moran. Estimation methods for evolutive processes. Journal of the Royal Statistical Society. Series B (Methodological), 13(1):141–146, ???? 1951. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983975.

Moran:1951:RDI

[220] P. A. P. Moran. The random division of an interval — Part II. Journal of the Royal Statistical Society. Series B (Methodological), 13(1):147–150, ???? 1951. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983976.

Anonymous:1951:FM

[221] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 13(1):??, ???? 1951. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983965.

Anonymous:1951:VIa

[222] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 13(1): ??, ???? 1951. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983964.

Kendall:1951:SPT

[223] David G. Kendall. Some problems in the theory of queues. Journal of the Royal Statistical Society. Series B (Methodological), 13(2):151–185, ???? 1951. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984059.

Daniels:1951:TPF

[224] H. E. Daniels and J. Wishart. The theory of position finding. Journal of the Royal Statistical Society. Series B (Methodological), 13(2):186–207, ???? 1951. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2984060. See correction [257].

Buckland:1951:RLS

[225] William R. Buckland. A review of the literature of systematic sampling. Journal of the Royal Statistical Society. Series B (Methodological), 13(2):208–215, ???? 1951. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984061.

Godambe:1951:TSS

[226] V. P. Godambe. On two-stage sampling. Journal of the Royal Statistical Society. Series B (Methodological), 13 (2):216–218, ???? 1951. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984062.

Sukhatme:1951:CPD

[227] B. V. Sukhatme. On certain probability distributions arising from points on a line. Journal of the Royal Statistical Society. Series B (Methodological), 13 (2):219–232, ???? 1951. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984063.

Guest:1951:ESE

[228] P. G. Guest. The estimation of standard error from successive finite differences. Journal of the Royal Statistical Society. Series B (Methodological), 13 (2):233–237, ???? 1951. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984064.

Simpson:1951:IIC

[229] E. H. Simpson. The interpretation of interaction in contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 13(2):238–241, ???? 1951. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984065.

Lancaster:1951:CCT

[230] H. O. Lancaster. Complex contingency tables treated by the partition of χ^{2}. Journal of the Royal Statistical Society. Series B (Methodological), 13(2): 242–249, ???? 1951. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984066.

Rosenbaum:1951:VLS

[231] S. Rosenbaum. The variance of least-square estimates under linear restraints. Journal of the Royal Statistical Society. Series B (Methodological), 13(2):250–255, ???? 1951. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984067.

Patterson:1951:CT

[232] H. D. Patterson. Change-over trials. Journal of the Royal Statistical Society. Series B (Methodological), 13(2): 256–271, ???? 1951. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984068.

Grundy:1951:GTA

[233] P. M. Grundy. A general technique for the analysis of experiments with incorrectly treated plots. Journal of the Royal Statistical Society. Series B (Methodological), 13(2):272–283, ???? 1951. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984069.

Finney:1951:SJS

[234] D. J. Finney. Subjective judgment in statistical analysis: An experimental study. Journal of the Royal Statistical Society. Series B (Methodological), 13 (2):284–297, ???? 1951. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984070.

Jowett:1951:ECO

[235] G. H. Jowett. The expression of the complementary outputs of two products in terms of a common unit of production effort. Journal of the Royal Statistical Society. Series B (Methodological), 13(2):298–302, ???? 1951. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984071.

Durbin:1951:IRC

[236] J. Durbin and A. Stuart. Inversions and rank correlation coefficients. Journal of the Royal Statistical Society. Series B (Methodological), 13(2):303–309, ???? 1951. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984072. See note [237].

Daniels:1951:NDS

[237] H. E. Daniels. Note on Durbin and Stuart’s formula for E(r_{s}). Journal of the Royal Statistical Society. Series B (Methodological), 13(2):310, ???? 1951. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984073. See [236].

Anonymous:1951:VIb

[238] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 13(2): 311–312, ???? 1951. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984058.

Armitage:1952:STB

[239] P. Armitage. The statistical theory of bacterial populations subject to mutation. Journal of the Royal Statistical Society. Series B (Methodological), 14 (1):1–40, ???? 1952. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984083.

Kendall:1952:CMM

[240] David G. Kendall. On the choice of a mathematical model to represent normal bacterial growth. Journal of the Royal Statistical Society. Series B (Methodological), 14(1):41–44, ???? 1952. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984084.

Tocher:1952:DAB

[241] K. D. Tocher. The design and analysis of block experiments. Journal of the Royal Statistical Society. Series B (Methodological), 14(1):45–100, ???? 1952. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984085.

Pearce:1952:SND

[242] S. C. Pearce. Some new designs of Latin square type. Journal of the Royal Statistical Society. Series B (Methodological), 14(1):101–106, ???? 1952. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984086.

Good:1952:RD

[243] I. J. Good. Rational decisions. Journal of the Royal Statistical Society. Series B (Methodological), 14(1):107–114, ???? 1952. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984087.

Sandelius:1952:CIS

[244] Martin Sandelius. A confidence interval for the smallest proportion of a binomial population. Journal of the Royal Statistical Society. Series B (Methodological), 14(1):115–116, ???? 1952. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984088.

Walker:1952:SPA

[245] A. M. Walker. Some properties of the asymptotic power functions of goodness-of-fit tests for linear autoregressive schemes. Journal of the Royal Statistical Society. Series B (Methodological), 14(1):117–134, ???? 1952. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984089.

Gardner:1952:GPI

[246] A. Gardner. Greenwood’s “Problem of intervals”: An exact solution for N = 3. Journal of the Royal Statistical Society. Series B (Methodological), 14(1): 135–139, ???? 1952. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984090.

Anonymous:1952:FMa

[247] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 14(1):??, ???? 1952. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984082.

Plackett:1952:QRM

[248] R. L. Plackett and P. S. Hewlett. Quantal responses to mixtures of poisons. Journal of the Royal Statistical Society. Series B (Methodological), 14(2): 141–163, ???? 1952. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2983865.

Healy:1952:SSA

[249] M. J. R. Healy. Some statistical aspects of anthropometry. Journal of the Royal Statistical Society. Series B (Methodological), 14(2):164–184, ???? 1952. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983866.

Bailey:1952:SQA

[250] Norman T. J. Bailey. A study of queues and appointment systems in hospital out-patient departments, with special reference to waiting-times. Journal of the Royal Statistical Society. Series B (Methodological), 14(2):185–199, ????

1952. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983867.

Benson:1952:FNP

[251] F. Benson. Further notes on the productivity of machines requiring attention at random intervals. Journal of the Royal Statistical Society. Series B (Methodological), 14(2):200–210, ???? 1952. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983868.

Cox:1952:SRW

[252] D. R. Cox. Some recent work on systematic experimental designs. Journal of the Royal Statistical Society. Series B (Methodological), 14(2):211–219, ???? 1952. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983869.

Chandler:1952:DFR

[253] K. N. Chandler. The distribution and frequency of record values. Journal of the Royal Statistical Society. Series B (Methodological), 14(2):220–228, ???? 1952. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983870.

Hartley:1952:SOA

[254] H. O. Hartley. Second order autoregressive schemes with time-trending coefficients. Journal of the Royal Statistical Society. Series B (Methodological), 14 (2):229–233, ???? 1952. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983871.

Loraine:1952:USO

[255] Phyllis K. Loraine. On a useful set of orthogonal comparisons. Journal of the Royal Statistical Society. Series B (Methodological), 14(2):234–237, ???? 1952. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983872.

Tweedie:1952:EPS

[256] M. C. K. Tweedie. The estimation of parameters from sequentially sampled data on a discrete distribution. Journal of the Royal Statistical Society. Series B (Methodological), 14(2):238–245, ???? 1952. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983873.

Daniels:1952:TPF

[257] H. E. Daniels. Theory of position finding — a correction. Journal of the Royal Statistical Society. Series B (Methodological), 14(2):246, ???? 1952. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983874. See [224].

Anonymous:1952:VI

[258] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 14(2): 247–248, ???? 1952. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2983863.

Anonymous:1952:FMb

[259] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 14(2):??, ???? 1952. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983864.

Anscombe:1953:SE

[260] F. J. Anscombe. Sequential estimation. Journal of the Royal Statistical Society. Series B (Methodological), 15(1):1–29, ???? 1953. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983720.

Lindley:1953:SI

[261] D. V. Lindley. Statistical inference. Journal of the Royal Statistical Society. Series B (Methodological), 15(1):30–76, ???? 1953. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983721.

Moran:1953:RDI

[262] P. A. P. Moran. The random division of an interval — Part II. Journal of the Royal Statistical Society. Series B (Methodological), 15(1):77–80, ???? 1953. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983722.

Jowett:1953:SGT

[263] G. H. Jowett and J. F. Scott. Simple graphical techniques for calculating serial and spatial correlations and means semi-squared differences. Journal of the Royal Statistical Society. Series B (Methodological), 15(1):81–86, ???? 1953. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983723.

Irwin:1953:TPC

[264] J. O. Irwin. On the “Transition probabilities” corresponding to any accident distribution. Journal of the Royal Statistical Society. Series B (Methodological), 15(1):87–89, ???? 1953. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983724.

Armitage:1953:NTH

[265] P. Armitage. A note on the timehomogeneous birth process. Journal of the Royal Statistical Society. Series B (Methodological), 15(1):90–91, ???? 1953. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983725.

Singer:1953:ATS

[266] K. Singer. Application of the theory of stochastic processes to the study of irreproducible chemical reactions and nucleation processes. Journal of the Royal Statistical Society. Series B (Methodological), 15(1):92–106, ???? 1953. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983726.

Bartlett:1953:GFT

[267] M. S. Bartlett and D. V. Rajalakshman. Goodness of fit tests for simultaneous autoregressive series. Journal of the Royal Statistical Society. Series B (Methodological), 15(1):107–124, ???? 1953. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983727.

Whittle:1953:AMS

[268] P. Whittle. The analysis of multiple stationary time series. Journal of the Royal Statistical Society. Series B (Methodological), 15(1):125–139, ???? 1953. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983728.

Sargan:1953:ATP

[269] J. D. Sargan. An approximate treatment of the properties of the correlogram and periodogram. Journal of the Royal Statistical Society. Series B (Methodological), 15(1):140–152, ???? 1953. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983729.

Anonymous:1953:FM

[270] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 15(1):??, ???? 1953. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983719.

Smith:1953:DLH

[271] Cedric A. B. Smith. The detection of linkage in human genetics. Journal of the Royal Statistical Society. Series B (Methodological), 15(2):153–192, ???? 1953. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983767.

Hotelling:1953:NLC

[272] Harold Hotelling. New light on the correlation coefficient and its transforms. Journal of the Royal Statistical Society. Series B (Methodological), 15(2): 193–232, ???? 1953. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983768.

Kamat:1953:SPE

[273] A. R. Kamat. Some properties of estimates for the standard deviation based on deviations from the mean and variate differences. Journal of the Royal Statistical Society. Series B (Methodological), 15(2):233–240, ????

1953. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983769.

Moran:1953:EPB

[274] P. A. P. Moran. The estimation of the parameters of a birth and death process. Journal of the Royal Statistical Society. Series B (Methodological), 15 (2):241–245, ???? 1953. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983770.

Sundrum:1953:PWS

[275] R. M. Sundrum. The power of Wilcoxon’s 2-sample test. Journal of the Royal Statistical Society. Series B (Methodological), 15(2):246–252, ???? 1953. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983771.

Yates:1953:SRW

[276] F. Yates and P. M. Grundy. Selection without replacement from within strata with probability proportional to size. Journal of the Royal Statistical Society. Series B (Methodological), 15 (2):253–261, ???? 1953. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983772.

Durbin:1953:SRS

[277] J. Durbin. Some results in sampling theory when the units are selected with unequal probabilities. Journal of the Royal Statistical Society. Series B (Methodological), 15(2):262–269, ???? 1953. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983773.

Anonymous:1953:VI

[278] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 15(2): 271–272, ???? 1953. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983766.

Foster:1954:DFT

[279] F. G. Foster and A. Stuart. Distributionfree tests in time-series based on the breaking of records. Journal of the Royal Statistical Society. Series B (Methodological), 16(1):1–22, ???? 1954. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984007.

Hammersley:1954:PMM

[280] J. M. Hammersley and K. W. Morton. Poor man’s Monte Carlo. Journal of the Royal Statistical Society. Series B (Methodological), 16(1):23–38, ???? 1954. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984008.

Tocher:1954:AAC

[281] K. D. Tocher. The application of automatic computers to sampling experiments. Journal of the Royal Statistical Society. Series B (Methodological), 16 (1):39–75, ???? 1954. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984009.

Hammersley:1954:TBP

[282] J. M. Hammersley and K. W. Morton. Transposed branching processes. Journal of the Royal Statistical Society. Series B (Methodological), 16(1):76–79, ???? 1954. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984010.

Bailey:1954:QPB

[283] Norman T. J. Bailey. On queueing processes with bulk service. Journal of the Royal Statistical Society. Series B (Methodological), 16(1):80–87, ???? 1954. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984011.

Bennett:1954:DHU

[284] J. H. Bennett. The distribution of heterogeneity upon inbreeding. Journal of the Royal Statistical Society. Series B (Methodological), 16(1):88–99, ???? 1954. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984012.

Banks:1954:FAC

[285] Charlotte Banks. The factorial analysis of crop productivity. Journal of the Royal Statistical Society. Series B (Methodological), 16(1):100–111, ???? 1954. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984013.

Moran:1954:SEP

[286] P. A. P. Moran. Some experiments on the prediction of sunspot numbers. Journal of the Royal Statistical Society. Series B (Methodological), 16(1): 112–117, ???? 1954. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984014.

Mandel:1954:GGS

[287] L. Mandel. Grading with a gauge subject to random output fluctuations.

Journal of the Royal Statistical Society. Series B (Methodological), 16(1): 118–130, ???? 1954. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984015.

Page:1954:CCM

[288] E. S. Page. Control charts for the mean of a normal population. Journal of the Royal Statistical Society. Series B (Methodological), 16(1):131–135, ???? 1954. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984016.

Page:1954:IWA

[289] E. S. Page. An improvement to Wald’s approximation for some properties of sequential tests. Journal of the Royal Statistical Society. Series B (Methodological), 16(1):136–139, ???? 1954. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984017.

Patterson:1954:ELS

[290] H. D. Patterson. The errors of lattice sampling. Journal of the Royal Statistical Society. Series B (Methodological), 16(1):140–149, ???? 1954. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984018.

Anonymous:1954:FM

[291] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 16(1):??, ???? 1954. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984006.

Barnard:1954:SIS

[292] G. A. Barnard. Sampling inspection and statistical decisions. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):151–174, ???? 1954. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984042.

Fieller:1954:SPI

[293] E. C. Fieller. Some problems in interval estimation. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):175–185, ???? 1954. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984043.

Creasy:1954:LRM

[294] Monica A. Creasy. Limits for the ratio of means. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):186–194, ???? 1954. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984044.

David:1954:CIP

[295] S. T. David. Confidence intervals for parameters in Markov autoregressive schemes. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):195–222, ???? 1954. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984045.

Church:1954:PSA

[296] B. M. Church. Problems of sample allocation and estimation in an agricultural survey. Journal of the Royal Statistical Society. Series B (Methodological), 16 (2):223–235, ???? 1954. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984046.

Grundy:1954:MSP

[297] P. M. Grundy. A method of sampling with probability exactly proportional to size. Journal of the Royal Statistical Society. Series B (Methodological), 16 (2):236–238, ???? 1954. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984047.

Stuart:1954:SPO

[298] Alan Stuart. A simple presentation of optimum sampling results. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):239–241, ???? 1954. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984048.

Rowell:1954:AFE

[299] J. G. Rowell. The analysis of a factorial experiment (with confounding) on an electronic calculator. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):242–246, ???? 1954. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984049.

Lancaster:1954:TCQ

[300] H. O. Lancaster. Traces and cumulants of quadratic forms in normal variables. Journal of the Royal Statistical Society. Series B (Methodological), 16 (2):247–254, ???? 1954. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984050.

Sundrum:1954:FAD

[301] R. M. Sundrum. A further approximation to the distribution of Wilcoxon’s statistic in the general case. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):255–260, ???? 1954. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984051.

Hodges:1954:TAV

[302] J. L. Hodges, Jr. and E. L. Lehmann. Testing the approximate validity of statistical hypotheses. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):261–268, ???? 1954. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984052.

Berz:1954:CEC

[303] F. Berz. On the cumulative effect of chance deviations. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):269–284, ???? 1954. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984053.

Cox:1954:TPP

[304] D. R. Cox. A table for predicting the production from a group of machines under the care of one operative. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):285–287, ???? 1954. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984054.

Bailey:1954:CTT

[305] Norman T. J. Bailey. A continuous time treatment of a simple queue using generating functions. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):288–291, ???? 1954. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984055.

Johnston:1954:RTS

[306] J. Johnston. A revised test for systematic oscillation. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):292–295, ???? 1954. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984056.

Bartlett:1954:NMF

[307] M. S. Bartlett. A note on the multiplying factors for various χ^{2} approximations. Journal of the Royal Statistical Society. Series B (Methodological), 16 (2):296–298, ???? 1954. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984057.

Anonymous:1954:VI

[308] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 16(2): 299–301, ???? 1954. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984041.

Box:1955:PTD

[309] G. E. P. Box and S. L. Andersen. Permutation theory in the derivation of robust criteria and the study of departures from assumption. Journal of the Royal Statistical Society. Series B (Methodological), 17(1):1–34, ???? 1955. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983783.

Bailey:1955:SPS

[310] Normal T. J. Bailey. Some problems in the statistical analysis of epidemic data. Journal of the Royal Statistical Society. Series B (Methodological), 17 (1):35–68, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983784.

Fisher:1955:SMS

[311] Ronald Fisher. Statistical methods and scientific induction. Journal of the Royal Statistical Society. Series B (Methodological), 17(1):69–78, ???? 1955. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983785.

Mauldon:1955:PQW

[312] J. G. Mauldon. Pivotal quantities for Wishart’s and related distributions, and a paradox in fiducial theory. Journal of the Royal Statistical Society. Series B (Methodological), 17(1):79–85, ???? 1955. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983786.

Huzurbazar:1955:CIP

[313] V. S. Huzurbazar. Confidence intervals for the parameter of a distribution admitting a sufficient statistic when the range depends on the parameter. Journal of the Royal Statistical Society. Series B (Methodological), 17(1):86–90, ???? 1955. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983787.

Jowett:1955:LSR

[314] G. H. Jowett. Least squares regression analysis for trend-reduced time series. Journal of the Royal Statistical Society. Series B (Methodological), 17 (1):91–104, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983788.

Scott:1955:NIL

[315] J. F. Scott and V. J. Small. A numerical investigation of least squares regression involving trend- reduced Markoff series. Journal of the Royal Statistical Society. Series B (Methodological), 17 (1):105–114, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983789.

Foster:1955:SEP

[316] F. G. Foster and D. Teichroew. A sampling experiment on the powers of the records tests for trend in a time series. Journal of the Royal Statistical Society. Series B (Methodological), 17 (1):115–121, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983790.

David:1955:MNO

[317] H. A. David. Moments of negative order and ratio-statistics. Journal of the Royal Statistical Society. Series B (Methodological), 17(1):122–123, ???? 1955. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983791.

Wurtele:1955:RIP

[318] Zivia S. Wurtele. A rectifying inspection plan. Journal of the Royal Statistical Society. Series B (Methodological), 17(1):124–127, ???? 1955. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983792.

Anonymous:1955:FMa

[319] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 17(1):??, ???? 1955. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983782.

Cox:1955:SSM

[320] D. R. Cox. Some statistical methods connected with series of events. Journal of the Royal Statistical Society. Series B (Methodological), 17(2):129–164, ???? 1955. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983950.

Vajda:1955:OLP

[321] S. Vajda. An outline of linear programming. Journal of the Royal Statistical Society. Series B (Methodological), 17 (2):165–172, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983951.

Beale:1955:MCF

[322] E. M. L. Beale. On minimizing a convex function subject to linear inequalities. Journal of the Royal Statistical Society. Series B (Methodological), 17 (2):173–184, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983952.

Morton:1955:CTS

[323] G. Morton and A. H. Land. A contribution to the “Travelling-Saleman” problem. Journal of the Royal Statistical Society. Series B (Methodological), 17 (2):185–203, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983953.

Pearson:1955:SCR

[324] E. S. Pearson. Statistical concepts in the relation to reality. Journal of the Royal Statistical Society. Series B (Methodological), 17(2):204–207, ???? 1955. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983954.

Jowett:1955:CMS

[325] G. H. Jowett. The comparison of means of sets of observations from sections of independent stochastic series. Journal of the Royal Statistical Society. Series B (Methodological), 17(2):208–227, ???? 1955. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983955.

Gower:1955:NPB

[326] J. C. Gower. A note on the periodogram of the Beveridge wheat price index. Journal of the Royal Statistical Society. Series B (Methodological), 17 (2):228–234, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983956.

Whittle:1955:SDM

[327] P. Whittle. Some distribution and moment formulae for the Markov chain. Journal of the Royal Statistical Society. Series B (Methodological), 17(2): 235–242, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983957.

Lomnicki:1955:SAZ

[328] Z. A. Lomnicki and S. K. Zaremba. Some applications of zero-one processes. Journal of the Royal Statistical Society. Series B (Methodological), 17 (2):243–255, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983958.

Downton:1955:WTB

[329] F. Downton. Waiting time in bulk service queues. Journal of the Royal Statistical Society. Series B (Methodological), 17(2):256–261, ???? 1955.

CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983959.

Bailey:1955:NEM

[330] Normam T. J. Bailey. A note on equalising the mean waiting times of successive customers in a finite queue. Journal of the Royal Statistical Society. Series B (Methodological), 17(2):262–263, ???? 1955. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983960.

Goods:1955:WCS

[331] I. J. Goods. On the weighted combination of significance tests. Journal of the Royal Statistical Society. Series B (Methodological), 17(2):264–265, ???? 1955. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983961.

Healy:1955:SDE

[332] M. J. R. Healy. A significance for the difference in efficiency between two predictors. Journal of the Royal Statistical Society. Series B (Methodological), 17 (2):266–268, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983962.

Godambe:1955:UTS

[333] V. P. Godambe. A unified theory of sampling from finite populations. Journal of the Royal Statistical Society. Series B (Methodological), 17(2):269–278, ???? 1955. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983963.

Anonymous:1955:VI

[334] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 17(2): 279–280, ???? 1955. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983949.

Anonymous:1955:BM

[335] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 17(2):??, ???? 1955. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/4291000.

Anonymous:1955:FMb

[336] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 17(2):??, ???? 1955. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/4290999.

David:1956:STS

[337] F. N. David and N. L. Johnson. Some tests of significance with ordered variables. Journal of the Royal Statistical Society. Series B (Methodological), 18 (1):1–31, ???? 1956. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983731.

Grundy:1956:ECA

[338] P. M. Grundy, M. J. R. Healy, and D. H. Rees. Economic choice of the amount of experimentation. Journal of the Royal Statistical Society. Series B (Methodological), 18(1):32–55, ???? 1956. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983732.

Fisher:1956:TSP

[339] Ronald Fisher. On a test of significance in Pearson’s Biometrika tables (no. 11). Journal of the Royal Statistical Society. Series B (Methodological), 18(1):56–60, ???? 1956. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983733. See

comment [366].

Moran:1956:TSU

[340] P. A. P. Moran. A tests of significance for an unidentifiable relation. Journal of the Royal Statistical Society. Series B (Methodological), 18(1):61–64, ???? 1956. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983734.

Creasy:1956:CLG

[341] Monica A. Creasy. Confidence limits for the gradient in the linear functional relationship. Journal of the Royal Statistical Society. Series B (Methodological), 18(1):65–69, ???? 1956. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983735.

Binet:1956:ATC

[342] F. E. Binet and G. S. Watson. Algebraic theory of the computing routine for tests of significance on the dimensionality of normal multivariate systems. Journal of the Royal Statistical Society. Series B (Methodological), 18 (1):70–78, ???? 1956. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983736.

Barton:1956:SNO

[343] D. E. Barton and F. N. David. Some notes on ordered random intervals. Journal of the Royal Statistical Society. Series B (Methodological), 18(1):79–94, ???? 1956. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983737.

Bartholomew:1956:STR

[344] D. J. Bartholomew. A sequential test for randomness of intervals. Journal of the Royal Statistical Society. Series B (Methodological), 18(1):95–103, ???? 1956. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983738.

Bennett:1956:PMT

[345] J. H. Bennett. Partitions in more than one dimension. Journal of the Royal Statistical Society. Series B (Methodological), 18(1):104–112, ???? 1956. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983739.

Good:1956:ESF

[346] I. J. Good. On the estimation of small frequencies in contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 18(1):113–124, ???? 1956. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983740.

Champernowne:1956:EMS

[347] D. G. Champernowne. An elementary method of solution of the queueing problem with a single server and constant parameters. Journal of the Royal Statistical Society. Series B (Methodological), 18(1):125–128, ???? 1956. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983741.

Jackson:1956:RQP

[348] R. R. P. Jackson. Random queueing processes with phase-type service.

Journal of the Royal Statistical Society. Series B (Methodological), 18(1): 129–132, ???? 1956. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2983742.

Claringbold:1956:WAB

[349] P. J. Claringbold. The within-animal bioassay with quantal responses. Journal of the Royal Statistical Society. Series B (Methodological), 18(1):133–137, ???? 1956. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983743.

Anonymous:1956:FM

[350] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 18(1):??, ???? 1956. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983730.

Mallows:1956:GTI

[351] C. L. Mallows. Generalizations of Tchebycheff’s inequalities. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):139–176, ???? 1956. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983702.

Cane:1956:SSP

[352] Violet R. Cane. Some statistical problems in experimental psychology. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):177–201, ???? 1956. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983703.

Kemp:1956:GHD

[353] C. D. Kemp and A. W. Kemp. Generalized hypergeometric distributions.

Journal of the Royal Statistical Society. Series B (Methodological), 18(2): 202–211, ???? 1956. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983704. See

corrigenda [384].

Fisher:1956:NTB

[354] R. A. Fisher and M. J. R. Healy. New tables of Behrens’ test of significance. Journal of the Royal Statistical Society. Series B (Methodological), 18(2): 212–216, ???? 1956. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983705.

Grundy:1956:FDP

[355] P. M. Grundy. Fiducial distributions and prior distributions: An example in which the former cannot be associated with the latter. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):217–221, ???? 1956. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983706.

Freund:1956:SME

[356] John E. Freund. Some methods of estimating the parameters of discrete heterogeneous populations. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):222–226, ???? 1956. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983707.

Hannan:1956:APC

[357] E. J. Hannan. The asymptotic powers of certain tests based on multiple correlations. Journal of the Royal Statistical Society. Series B (Methodological), 18 (2):227–233, ???? 1956. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983708.

Bartholomew:1956:TRS

[358] D. J. Bartholomew. Tests for randomness in a series of events when the alternative is a trend. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):234–239, ???? 1956. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983709.

Wise:1956:RAR

[359] J. Wise. Regression analysis of relationships between autocorrelated time series. Journal of the Royal Statistical Society. Series B (Methodological), 18 (2):240–256, ???? 1956. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983710.

Miller:1956:EAL

[360] Irwin Miller and John E. Freund. Expected arc length of a Gaussian process on a finite interval. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):257–258, ???? 1956. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983711.

Rao:1956:ADI

[361] C. Radhakrishna Rao. Analysis of dispersion with incomplete observations on one of the characters. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):259–264, ???? 1956. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983712.

Downton:1956:LDA

[362] F. Downton. On limiting distributions arising in bulk service queues. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):265–274, ???? 1956. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983713.

Jackson:1956:SER

[363] R. R. P. Jackson and D. G. Nickols. Some equilibrium results for the queueing process E_{k} /M/1. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):275–279, ???? 1956. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983714.

Naor:1956:MI

[364] P. Naor. On machine interference. Journal of the Royal Statistical Society. Series B (Methodological), 18(2): 280–287, ???? 1956. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983715.

Neyman:1956:NAS

[365] Jerzy Neyman. Note on an article by Sir Ronald Fisher. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):288–294, ???? 1956. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2983716. See comment [382].

Bartlett:1956:CSR

[366] M. S. Bartlett. Comment on Sir Ronald Fisher’s paper: “On a Test of Significance in Pearson’s Biometrika Tables (No. 11)”. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):295–296, ???? 1956. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2983717. See [339] and comment [382].

Welch:1956:NSC

[367] B. L. Welch. Note on some criticisms made by Sir Ronald Fisher. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):297–302, ???? 1956. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983718. See [382].

Anonymous:1956:VI

[368] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 18(2): 303–304, ???? 1956. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983701.

Jenkins:1957:SAT

[369] G. M. Jenkins and M. B. Priestley. The spectral analysis of time-series. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):1–12, ???? 1957. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983992.

Lomnicki:1957:ESD

[370] Z. A. Lomnicki and S. K. Zaremba. On estimating the spectral density function of a stochastic process. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):13–37, ???? 1957. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983993.

Whittle:1957:CPS

[371] P. Whittle. Curve and periodogram smoothing. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):38–63, ???? 1957. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983994.

Skellam:1957:DAR

[372] J. G. Skellam and L. R. Shenton. Distributions associated with random walk and recurrent events. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):64–118, ???? 1957. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983995.

Barton:1957:CTS

[373] D. E. Barton. A comparison of two sorts of test for a change of location applicable to truncated data. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):119–124, ???? 1957. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983996.

Bulmer:1957:CSH

[374] M. G. Bulmer. Confirming statistical hypotheses. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):125–132, ???? 1957. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983997.

Ray:1957:SCI

[375] W. D. Ray. Sequential confidence intervals for the mean of a normal population with unknown variance. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):133–143, ???? 1957. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983998.

Sukhatme:1957:JAD

[376] Balkrishna V. Sukhatme. Joint asymptotic distribution of the median and a U statistic. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):144–148, ???? 1957. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983999.

Stuart:1957:ERT

[377] Alan Stuart. The efficiency of the records test for trend in normal regression. Journal of the Royal Statistical Society. Series B (Methodological), 19 (1):149–153, ???? 1957. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984000.

Freeman:1957:SEDa

[378] G. H. Freeman. Some experimental designs of use in changing from one set of treatments to another, Part I. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):154–162, ???? 1957. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984001.

Freeman:1957:SEDb

[379] G. H. Freeman. Some experimental designs of use in changing from one set of treatments to another: Part II. Existence of the designs. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):163–164, ???? 1957. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984002.

Mack:1957:EMUa

[380] C. Mack, T. Murphy, and N. L. Webb. The efficiency of N machines uni-directionally patrolled by one operative when walking time and repair times are constants. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):166–172, ???? 1957. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984003.

Mack:1957:EMUb

[381] C. Mack. The efficiency of N machines uni-directionally patrolled by one operative when walking time is constant and repair times are variable. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):173–178, ???? 1957. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984004.

Fisher:1957:CNN

[382] R. A. Fisher. Comment on the notes by Neyman, Bartlett, and Welch in this Journal (vol. 18, no. 2, 1956). Journal of the Royal Statistical Society. Series B (Methodological), 19(1): 179, ???? 1957. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

jstor.org/stable/2984005. See [365,

366, 367].

Anonymous:1957:FM

[383] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):??, ???? 1957. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983990.

Kemp:1957:CGH

[384] C. D. Kemp and A. W. Kemp. Corrigenda: Generalized hypergeometric distributions. Journal of the Royal Statistical Society. Series B (Methodological), 19(1):??, ???? 1957. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983991. See [353].

Gani:1957:PPT

[385] J. Gani. Problems in the probability theory of storage systems. Journal of the Royal Statistical Society. Series B (Methodological), 19(2):181–206, ???? 1957. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983814.

Kendall:1957:SPT

[386] David G. Kendall. Some problems in the theory of dams. Journal of the Royal Statistical Society. Series B (Methodological), 19(2):207–233, ???? 1957. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983815.

Yates:1957:RAR

[387] F. Yates, M. J. R. Healy, and S. Lipton. Routine analysis of replicated experiments on an electronic computer. Journal of the Royal Statistical Society. Series B (Methodological), 19(2):234–254, ???? 1957. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983816.

Maurice:1957:MPC

[388] Rita J. Maurice. A minimax procedure for choosing between two populations using sequential sampling. Journal of the Royal Statistical Society. Series B (Methodological), 19(2):255–261, ???? 1957. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983817.

Watson:1957:SSS

[389] G. S. Watson. Sufficient statistics, similar regions and distribution-free tests. Journal of the Royal Statistical Society. Series B (Methodological), 19(2): 262–267, ???? 1957. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983818.

Whittle:1957:UNA

[390] P. Whittle. On the use of the normal approximation in the treatment of stochastic processes. Journal of the Royal Statistical Society. Series B (Methodological), 19(2):268–281, ???? 1957. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983819.

Hannan:1957:VMS

[391] E. J. Hannan. The variance of the mean of a stationary process. Journal of the Royal Statistical Society. Series B (Methodological), 19(2):282–285, ???? 1957. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983820.

Sampford:1957:MCA

[392] M. R. Sampford. Methods of construction and analysis of serially balanced sequences. Journal of the Royal Statistical Society. Series B (Methodological), 19(2):286–304, ???? 1957. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983821.

Rojas:1957:MLS

[393] B. Rojas and R. F. White. The modified Latin square. Journal of the Royal Statistical Society. Series B (Methodological), 19(2):305–317, ???? 1957. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983822.

Creasy:1957:AVA

[394] Monica A. Creasy. Analysis of variance as an alternative to factor analysis. Journal of the Royal Statistical Society. Series B (Methodological), 19 (2):318–325, ???? 1957. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983823.

Bailey:1957:SFR

[395] Norman T. J. Bailey. Some further results in the non-equilibrium theory of a simple queue. Journal of the Royal Statistical Society. Series B (Methodological), 19(2):326–333, ???? 1957. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983824.

Naor:1957:NAM

[396] P. Naor. Normal approximation to machine interference with many repairmen. Journal of the Royal Statistical Society. Series B (Methodological), 19 (2):334–341, ???? 1957. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983825.

Gani:1957:SDN

[397] J. Gani and N. U. Prabhu. Stationary distributions of the negative exponential type for the infinite dam. Journal of the Royal Statistical Society. Series B (Methodological), 19(2):342–351, ???? 1957. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983826.

Anonymous:1957:VI

[398] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 19(2): 353–354, ???? 1957. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983813.

Neyman:1958:SAP

[399] Jerzy Neyman and Elizabeth L. Scott. Statistical approach to problems of cosmology. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):1–43, ???? 1958. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983905.

Watson:1958:CSG

[400] G. S. Watson. On chi-square goodnessof-fit tests for continuous distributions. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):44–72, ???? 1958. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983906.

Barton:1958:MDP

[401] D. E. Barton. The matching distributions: Poisson limiting forms and derived methods of approximation. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):73–92, ???? 1958. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983907.

Silvey:1958:LSP

[402] S. D. Silvey. The lindisfarne scribes’ problem. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):93–101, ???? 1958. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983908.

Lindley:1958:FDB

[403] D. V. Lindley. Fiducial distributions and Bayes’ Theorem. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):102–107, ???? 1958. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983909.

Banerjee:1958:EDT

[404] D. P. Banerjee. On the exact distribution of a test in multivariate analysis. Journal of the Royal Statistical Society. Series B (Methodological), 20 (1):108–110, ???? 1958. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983910.

Broadbent:1958:IMP

[405] S. R. Broadbent. The inspection of a Markov process. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):111–119, ???? 1958. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983911.

Davies:1958:FMS

[406] Hilda M. Davies and G. H. Jowett. The fitting of Markoff serial variation curves. Journal of the Royal Statistical Society. Series B (Methodological), 20 (1):120–142, ???? 1958. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983912.

Hannan:1958:APC

[407] E. J. Hannan. The asymptotic powers of certain tests of goodness of fit for time series. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):143–151, ???? 1958. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983913.

Grenander:1958:BVE

[408] Ulf Grenander. Bandwidth and variance in estimation of the spectrum. Journal of the Royal Statistical Society. Series B (Methodological), 20(1): 152–157, ???? 1958. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983914.

Quenouille:1958:CCT

[409] M. H. Quenouille. The comparison of correlations in time-series. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):158–164, ???? 1958. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983915.

Conolly:1958:DETa

[410] B. W. Conolly. A difference equation technique applied to the simple queue. Journal of the Royal Statistical Society. Series B (Methodological), 20(1): 165–167, ???? 1958. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983916.

Conolly:1958:DETb

[411] B. W. Conolly. A difference equation technique applied to the simple queue with arbitrary arrival interval distribution. Journal of the Royal Statistical Society. Series B (Methodological), 20 (1):168–175, ???? 1958. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983917.

Luchak:1958:CTS

[412] George Luchak. The continuous time solution of the equations of the single channel queue with a general class of service-time distributions by the method of generating functions. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):176–181, ???? 1958. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983918.

Finch:1958:ESW

[413] P. D. Finch. The effect of the size of the waiting room on a simple queue. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):182–186, ???? 1958. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983919.

Moore:1958:IAL

[414] P. G. Moore. Interval analysis and the logarithmic transformation. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):187–192, ???? 1958. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983920.

Cox:1958:ALS

[415] C. P. Cox. The analysis of Latin square designs with individual curvatures in one direction. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):193–204, ???? 1958. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983921.

Stevens:1958:DSS

[416] W. L. Stevens. Dilution series: A statistical test of technique. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):205–214, ???? 1958. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2983922. See corrigenda [457].

Anonymous:1958:FM

[417] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 20(1):??, ???? 1958. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983904.

Cox:1958:RAB

[418] D. R. Cox. The regression analysis of binary sequences. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):215–242, ???? 1958. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2983890. See corrigenda [455].

Smith:1958:RTR

[419] Walter L. Smith. Renewal theory and its ramifications. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):243–302, ???? 1958. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983891.

Parzen:1958:AEC

[420] Emanuel Parzen. On asymptotically efficient consistent estimates of the spectral density function of a stationary time series. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):303–322, ???? 1958. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983892.

Hannan:1958:ESD

[421] E. J. Hannan. The estimation of the spectral density after trend removal. Journal of the Royal Statistical Society. Series B (Methodological), 20(2): 323–333, ???? 1958. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2983893.

Whittle:1958:SPD

[422] P. Whittle. On the smoothing of probability density functions. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):334–343, ???? 1958. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983894.

Scheffe:1958:EM

[423] Henry Scheffé. Experiments with mixtures. Journal of the Royal Statistical Society. Series B (Methodological), 20 (2):344–360, ???? 1958. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983895. See

corrigenda [456], comments [452], and reply [518].

Good:1958:IAP

[424] I. J. Good. The interaction algorithm and practical Fourier analysis. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):361–372, ???? 1958. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983896.

Stuart:1958:ECV

[425] Alan Stuart. Equally correlated variates and the multinormal integral. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):

373–378, ???? 1958. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983897.

Kemp:1958:FCO

[426] K. W. Kemp. Formulae for calculating the operating characteristic and the average sample number of some sequential tests. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):379–386, ???? 1958. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983898.

Darwin:1958:CCS

[427] J. H. Darwin. On corrections to the chi-squared distribution. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):387–392, ???? 1958. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983899.

Stevens:1958:SRP

[428] W. L. Stevens. Sampling without replacement with probability proportional to size. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):393–397, ???? 1958. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983900.

Claringbold:1958:MQA

[429] P. J. Claringbold. Multivariate quantal analysis. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):398–405, ???? 1958. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983901.

Cox:1958:CDG

[430] C. P. Cox. A concise derivation of general orthogonal polynomials. Journal of the Royal Statistical Society. Series B (Methodological), 20(2):406–407, ???? 1958. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983902.

Tanner:1958:SMD

[431] J. C. Tanner. A simplified model for delays in overtaking on a two-lane road. Journal of the Royal Statistical Society. Series B (Methodological), 20(2): 408–414, ???? 1958. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2983903.

Anonymous:1958:VI

[432] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 20(2): 415–416, ???? 1958. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2983889.

Winsten:1959:GDT

[433] C. B. Winsten. Geometric distributions in the theory of queues. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):1–35, ???? 1959. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983924.

Cane:1959:BSS

[434] Violet R. Cane. Behaviour sequences as semi-Markov chains. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):36–58, ???? 1959. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983925.

Bennett:1959:MVF

[435] B. M. Bennett. On a multivariate version of Fieller’s theorem. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):59–62, ???? 1959. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983926.

Vagholkar:1959:PCE

[436] M. K. Vagholkar. The process curve and the equivalent mixed binomial with two components. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):63–66, ???? 1959. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983927.

Mallows:1959:IE

[437] C. L. Mallows. The information in an experiment. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):67–72, ???? 1959. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983928.

James:1959:BFD

[438] G. S. James. The behrens–Fisher distribution and weighted means. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):73–90, ???? 1959. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983929.

Sargan:1959:ERA

[439] J. D. Sargan. The estimation of relationships with autocorrelated residuals by the use of instrumental variables. Journal of the Royal Statistical Society. Series B (Methodological), 21 (1):91–105, ???? 1959. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983930.

Nasr:1959:SPM

[440] Saad K. Nasr. On some problems of machine interference. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):106–113, ???? 1959. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983931.

Bhat:1959:DVS

[441] B. R. Bhat. On the distribution of various sums of squares in an analysis of variance table for different classifications with correlated and non-homogeneous errors. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):114–119, ???? 1959. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983932.

Barton:1959:CO

[442] D. E. Barton and F. N. David. Contagious occupancy. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):120–133, ???? 1959. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983933.

Bodmer:1959:LFI

[443] W. F. Bodmer. The limiting frequencies of integers with a given partitional characteristic. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):134–143, ???? 1959. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983934.

Mercer:1959:SSD

[444] A. Mercer. Some simple durationdependent stochastic processes. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):144–152, ???? 1959. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983935.

Finch:1959:CQF

[445] P. D. Finch. Cyclic queues with feedback. Journal of the Royal Statistical Society. Series B (Methodological), 21 (1):153–157, ???? 1959. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983936.

Aitchison:1959:STR

[446] J. Aitchison. A statistical theory of remnants. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):158–168, ???? 1959. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983937.

Lomnicki:1959:BRS

[447] Z. A. Lomnicki and S. K. Zaremba. Bandwidth and resolvability in statistical spectral analysis. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):169–171, ???? 1959. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983938.

Jones:1959:PIB

[448] R. Morley Jones. On a property of incomplete blocks. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):172–179, ???? 1959. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983939.

Cox:1959:RPB

[449] D. R. Cox. A renewal problem with bulk ordering of components. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):180–189, ???? 1959. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983940.

Barton:1959:DNS

[450] D. E. Barton and F. N. David. The dispersion of a number of species. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):190–194, ???? 1959. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983941.

Cox:1959:DPC

[451] D. R. Cox and L. Brandwood. On a discriminatory problem connected with the works of Plato. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):195–200, ???? 1959. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983942.

Quenouille:1959:EM

[452] M. H. Quenouille. Experiments with mixtures. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):201–202, ???? 1959. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983943.

See reply [518].

Maurice:1959:DLF

[453] Rita Maurice. A different loss function for the choice between two populations. Journal of the Royal Statistical Society. Series B (Methodological), 21 (1):203–213, ???? 1959. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983944.

Sampford:1959:COR

[454] M. R. Sampford and J. Taylor. Censored observations in randomized block experiments. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):214–237, ???? 1959. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983945.

Cox:1959:CRA

[455] D. R. Cox. Corrigenda: The regression analysis of binary sequences. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):238, ???? 1959. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983947. See [418].

Scheffe:1959:CEM

[456] H. Scheffé. Corrigenda: Experiments with mixtures. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):238, ???? 1959. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983948. See [423].

Stevens:1959:CDS

[457] W. L. Stevens. Corrigenda: Dilution series: A statistical test of technique. Journal of the Royal Statistical Society. Series B (Methodological), 21(1):238, ???? 1959. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983946. See [416].

Anonymous:1959:VIa

[458] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 21(1): ??, ???? 1959. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983923.

Barnard:1959:CCS

[459] G. A. Barnard. Control charts and stochastic processes. Journal of the Royal Statistical Society. Series B (Methodological), 21(2):239–271, ???? 1959. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983801.

Kiefer:1959:OED

[460] J. Kiefer. Optimum experimental designs. Journal of the Royal Statistical Society. Series B (Methodological), 21 (2):272–319, ???? 1959. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983802.

Miller:1959:CTB

[461] Rupert G. Miller, Jr. A contribution to the theory of bulk queues. Journal of the Royal Statistical Society. Series B (Methodological), 21(2):320–337, ???? 1959. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983803.

Zoutendijk:1959:MFC

[462] G. Zoutendijk. Maximizing a function in a convex region. Journal of the Royal Statistical Society. Series B (Methodological), 21(2):338–355, ???? 1959. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983804.

Sarhan:1959:ELS

[463] Ahmed E. Sarhan and B. G. Greenberg. Estimation of location and scale parameters for the rectangular population from censored samples. Journal of the Royal Statistical Society. Series B (Methodological), 21(2):356–363, ???? 1959. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983805.

Page:1959:DVL

[464] E. S. Page. The distribution of vacancies on a line. Journal of the Royal Statistical Society. Series B (Methodological), 21(2):364–374, ???? 1959. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983806.

Finch:1959:OPQ

[465] P. D. Finch. The output process of the queueing system M/G/1. Journal of the Royal Statistical Society. Series B (Methodological), 21(2):375–380, ???? 1959. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983807.

Howie:1959:EAW

[466] A. J. Howie and L. R. Shenton. The efficiency of automatic winding machines with constant patrolling time. Journal of the Royal Statistical Society. Series B (Methodological), 21(2):381–395, ???? 1959. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983808.

Williams:1959:CRV

[467] E. J. Williams. The comparison of regression variables. Journal of the Royal Statistical Society. Series B (Methodological), 21(2):396–399, ???? 1959. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983809.

Wetherill:1959:MES

[468] G. B. Wetherill. The most economical sequential sampling scheme for inspection by variables. Journal of the Royal Statistical Society. Series B (Methodological), 21(2):400–408, ???? 1959. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983810.

Laubscher:1959:NFT

[469] N. F. Laubscher. Note on Fisher’s transformation of the correlation coefficient. Journal of the Royal Statistical Society. Series B (Methodological), 21 (2):409–410, ???? 1959. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983811.

Cox:1959:AED

[470] D. R. Cox. The analysis of exponentially distributed life-times with two types of failure. Journal of the Royal Statistical Society. Series B (Methodological), 21(2):411–421, ???? 1959. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983812.

Anonymous:1959:VIb

[471] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 21(2): 423–424, ???? 1959. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2983800.

Dunnett:1960:SLK

[472] C. W. Dunnett. On selecting the largest of k normal population means. Journal of the Royal Statistical Society. Series B (Methodological), 22(1):1–40, ???? 1960. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983876.

Beale:1960:CRN

[473] E. M. L. Beale. Confidence regions in non-linear estimation. Journal of the Royal Statistical Society. Series B (Methodological), 22(1):41–88, ???? 1960. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983877.

Conolly:1960:BPR

[474] B. W. Conolly. The busy period in relation to the single-server queueing system with general independent arrivals and Erlangian service-time. Journal of the Royal Statistical Society. Series B (Methodological), 22(1):89–96, ???? 1960. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983878.

Roy:1960:NPA

[475] A. D. Roy. A note on prediction from an autoregressive process using pistimetric probability. Journal of the Royal Statistical Society. Series B (Methodological), 22(1):97–103, ???? 1960. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983879.

Prabhu:1960:SRQ

[476] N. U. Prabhu. Some results for the queue with Poisson arrivals. Journal of the Royal Statistical Society. Series B (Methodological), 22(1):104–107, ???? 1960. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983880.

Mercer:1960:QPW

[477] A. Mercer. A queueing problem in which the arrival times of the customers are scheduled. Journal of the Royal Statistical Society. Series B (Methodological), 22(1):108–113, ???? 1960. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983881. See correction

[1111].

Joseph:1960:TPM

[478] A. W. Joseph and M. T. L. Bizley. The two-pack matching problem. Journal of the Royal Statistical Society. Series B (Methodological), 22(1):114–130, ???? 1960. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983882.

Tikkiwal:1960:TCR

[479] B. D. Tikkiwal. On the theory of classical regression and double sampling estimation. Journal of the Royal Statistical Society. Series B (Methodological), 22 (1):131–138, ???? 1960. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983883.

Durbin:1960:EPT

[480] J. Durbin. Estimation of parameters in time-series regression models. Journal of the Royal Statistical Society. Series B (Methodological), 22(1):139–153, ???? 1960. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983884.

Aitchison:1960:MLE

[481] J. Aitchison and S. D. Silvey. Maximum-likelihood estimation procedures and associated tests of significance. Journal of the Royal Statistical Society. Series B (Methodological), 22 (1):154–171, ???? 1960. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983885.

Cox:1960:RAW

[482] D. R. Cox. Regression analysis when there is prior information about supplementary variables. Journal of the Royal Statistical Society. Series B (Methodological), 22(1):172–176, ???? 1960. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983886.

Lowe:1960:TIB

[483] J. R. Lowe. A table of the integral of the bivariate normal distribution over an offset circle. Journal of the Royal Statistical Society. Series B (Methodological), 22(1):177–187, ???? 1960. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983887.

Ruben:1960:DWD

[484] Harold Ruben. On the distribution of the weighted difference of two independent student variables. Journal of the Royal Statistical Society. Series B (Methodological), 22(1):188–194, ???? 1960. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983888.

Anonymous:1960:VIa

[485] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 22(1): ??, ???? 1960. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983875.

Plackett:1960:MAV

[486] R. L. Plackett. Models in the analysis of variance. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):195–217, ???? 1960. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984092.

Bodmer:1960:DSP

[487] W. F. Bodmer. Discrete stochastic processes in population genetics. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):218–244, ???? 1960. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984093.

Ben-Israel:1960:PDSa

[488] A. Ben-Israel and P. Naor. A problem of delayed service — I. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):245–269, ???? 1960. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984094.

Ben-Israel:1960:PDSb

[489] A. Ben-Israel and P. Naor. A problem of delayed service — II. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):270–276, ???? 1960. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984095.

Finch:1960:TBS

[490] P. D. Finch. On the transient behaviour of a simple queue. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):277–284, ???? 1960. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984096.

Conolly:1960:QSS

[491] B. W. Conolly. Queueing at a single serving point with group arrival. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):285–298, ???? 1960. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984097.

Fisher:1960:SEB

[492] R. A. Fisher. On some extensions of Bayesian inference proposed by Mr Lindley. Journal of the Royal Statistical Society. Series B (Methodological), 22 (2):299–301, ???? 1960. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984098.

Buck:1960:MEM

[493] S. F. Buck. A method of estimation of missing values in multivariate data suitable for use with an electronic computer. Journal of the Royal Statistical Society. Series B (Methodological), 22 (2):302–306, ???? 1960. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984099.

Singh:1960:EPM

[494] Naunihal Singh. Estimation of parameters of a multivariate normal population from truncated and censored samples. Journal of the Royal Statistical Society. Series B (Methodological), 22 (2):307–311, ???? 1960. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984100.

Sprott:1960:NRD

[495] D. A. Sprott. Necessary restrictions for distributions a posteriori. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):312–318, ???? 1960. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984101.

Good:1960:WEC

[496] I. J. Good. Weight of evidence, corroboration, explanatory power, information and the utility of experiments. Journal of the Royal Statistical Society. Series B (Methodological), 22(2): 319–331, ???? 1960. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984102. See

corrigendum [868].

Mansfield:1960:LPT

[497] Edwin Mansfield and Carlton Hensley. The logistic process: Tables of the stochastic epidemic curve and applications. Journal of the Royal Statistical Society. Series B (Methodological), 22 (2):332–337, ???? 1960. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984103.

Roy:1960:SNP

[498] A. D. Roy. Some notes on pistimetric inference. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):338–347, ???? 1960. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984104.

Roy:1960:EDC

[499] S. N. Roy, B. G. Greenberg, and A. E. Sarhan. Evaluation of determinants, characteristic equations and their roots for a class of patterned matrices. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):348–359, ???? 1960. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984105.

Ghosh:1960:BES

[500] M. N. Ghosh. Bounds for the expected sample size in a sequential probability ratio test. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):360–367, ???? 1960. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984106.

Cox:1960:NTH

[501] D. R. Cox. A note on tests of homogeneity applied after sequential sampling. Journal of the Royal Statistical Society. Series B (Methodological), 22 (2):368–371, ???? 1960. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984107.

Good:1960:IAP

[502] I. J. Good. The interaction algorithm and practical Fourier analysis: An addendum. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):372–375, ???? 1960. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984108.

Daniels:1960:ASG

[503] H. E. Daniels. Approximate solutions of Green’s type for univariate stochastic processes. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):376–401, ???? 1960. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984109.

Wetherill:1960:WTN

[504] G. B. Wetherill. The Wilcoxon test and non-null hypotheses. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):402–418, ???? 1960. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984110.

Anonymous:1960:VIb

[505] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 22(2): 419–420, ???? 1960. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984091.

Anonymous:1961:BM

[506] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):i, ???? 1961. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983862.

Smith:1961:CSI

[507] Cedric A. B. Smith. Consistency in statistical inference and decision. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):1–37, ???? 1961. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/

stable/2983842. See correction [764].

Tanner:1961:TPA

[508] J. C. Tanner. Two papers on applications of stochastic processes to road traffic problems: Delays on a two-lane road. Journal of the Royal Statistical Society. Series B (Methodological), 23 (1):38–63, ???? 1961. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2983843.

Miller:1961:QMR

[509] Alan J. Miller. A queueing model for road traffic flow. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):64–90, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983844.

Leser:1961:SMT

[510] C. E. V. Leser. A simple method of trend construction. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):91–107, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983845.

Bennett:1961:CLM

[511] B. M. Bennett. Confidence limits for multivariate ratios. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):108–112, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983846.

Fabens:1961:SQI

[512] Augustus J. Fabens. The solution of queueing and inventory models by semi-Markov processes. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):113–127, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2983847. See correction [641].

Farlie:1961:AED

[513] D. J. G. Farlie. The asymptotic efficiency of Daniels’s generalized correlation coefficients. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):128–142, ????

1961. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983848.

Jaiswal:1961:BSQ

[514] N. K. Jaiswal. A bulk-service queueing problem with variable capacity. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):143–148, ???? 1961. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2983849.

Kemp:1961:ARL

[515] K. W. Kemp. The average run length of the cumulative sum chart when a V -mask is used. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):149–153, ???? 1961. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983850.

Glasser:1961:UEP

[516] Gerald J. Glasser. An unbiased estimator for powers of the arithmetic mean. Journal of the Royal Statistical Society. Series B (Methodological), 23(1): 154–159, ???? 1961. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2983851.

Dorff:1961:EPL

[517] Martin Dorff and John Gurland. Estimation of the parameters of a linear functional relation. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):160–170, ???? 1961. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2983852.

Scheffe:1961:RMQ

[518] Henry Scheffé. Reply to Mr Quenouille’s comments about my paper on mixtures. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):171–172, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983853. See [452].

Yeo:1961:TDS

[519] G. F. Yeo. The time-dependent solution for an infinite dam with discrete additive inputs. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):173–179, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983854.

Good:1961:RSC

[520] I. J. Good. The real stable characteristic functions and chaotic acceleration. Journal of the Royal Statistical Society. Series B (Methodological), 23(1): 180–183, ???? 1961. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983855.

Kerridge:1961:II

[521] D. F. Kerridge. Inaccuracy and inference. Journal of the Royal Statistical Society. Series B (Methodological), 23 (1):184–194, ???? 1961. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983856.

Vincent:1961:THO

[522] Shirley E. Vincent. A test of homogeneity for ordered variances. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):195–206, ???? 1961. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983857.

Downton:1961:NVL

[523] F. Downton. A note on vacancies on a line. Journal of the Royal Statistical Society. Series B (Methodological), 23 (1):207–214, ???? 1961. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983858.

Cox:1961:SCS

[524] D. R. Cox. A simple congestion system with incomplete service. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):215–222, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983859.

Tallis:1961:MGF

[525] G. M. Tallis. The moment generating function of the truncated multinormal distribution. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):223–229, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983860.

Smith:1961:NRF

[526] Walter L. Smith. A note on the renewal function when the mean renewal lifetime is infinite. Journal of the Royal Statistical Society. Series B (Methodological), 23(1):230–237, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983861.

Anonymous:1961:VI

[527] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 23(1): ??, ???? 1961. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983841.

Bartholomew:1961:THM

[528] D. J. Bartholomew. A test of homogeneity of means under restricted alternatives. Journal of the Royal Statistical Society. Series B (Methodological), 23 (2):239–281, ???? 1961. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984019.

Curnow:1961:OPV

[529] R. N. Curnow. Optimal programmes for varietal selection. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):282–318, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984020.

Saaty:1961:SSP

[530] Thomas L. Saaty. Some stochastic processes with absorbing barriers. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):319–334, ???? 1961. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984021.

Gilchrist:1961:SST

[531] W. G. Gilchrist. Some sequential tests using range. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):335–342, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984022.

Weesakul:1961:FEF

[532] B. Weesakul. First emptiness in a finite dam. Journal of the Royal Statistical Society. Series B (Methodological), 23 (2):343–351, ???? 1961. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984023.

Aitchison:1961:COD

[533] John Aitchison. The construction of optimal designs for the one-way classification analysis of variance. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):352–367, ???? 1961. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984024.

Mercer:1961:SSW

[534] A. Mercer. Some simple weardependent renewal processes. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):368–376, ???? 1961. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984025.

Ray:1961:EDF

[535] W. D. Ray and A. E. N. T. Pitman. An exact distribution of the Fisher– Behrens–Welch statistic for testing the difference between the means of two normal populations with unknown variance. Journal of the Royal Statistical Society. Series B (Methodological), 23 (2):377–384, ???? 1961. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984026.

Benson:1961:CQS

[536] F. Benson and G. Gregory. Closed queueing systems: A generalization of the machine interference model. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):385–393, ???? 1961. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984027.

Hannan:1961:TJS

[537] E. J. Hannan. Testing for a jump in the spectral function. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):394–404, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984028.

Rosenbaum:1961:MTB

[538] S. Rosenbaum. Moments of a truncated bivariate normal distribution. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):405–408, ???? 1961. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984029.

Daniels:1961:MGD

[539] H. E. Daniels. Mixtures of geometric distributions. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):409–413, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984030.

Cox:1961:PEW

[540] D. R. Cox. Prediction by exponentially weighted moving averages and related methods. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):414–422, ???? 1961. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984031.

Barton:1961:RBP

[541] D. E. Barton and C. L. Mallows. The randomization bases of the problem of the amalgamation of weighted means.

Journal of the Royal Statistical Society. Series B (Methodological), 23(2): 423–433, ???? 1961. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984032.

Stout:1961:EQD

[542] H. P. Stout and F. Stern. Estimation from quantiles in destructive testing. Journal of the Royal Statistical Society. Series B (Methodological), 23(2): 434–443, ???? 1961. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984033.

Silvey:1961:NML

[543] S. D. Silvey. A note on maximumlikelihood in the case of dependent random variables. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):444–452, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984034.

Sen:1961:SPA

[544] P. K. Sen. On some properties of the asymptotic variance of the sample quantiles and mid-ranges. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):453–459, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984035.

Sprott:1961:SBL

[545] D. A. Sprott. Similarities between likelihoods and associated distributions a posteriori. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):460–468, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984036.

Richards:1961:MML

[546] F. S. G. Richards. A method of maximum-likelihood estimation. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):469– 475, ???? 1961. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

jstor.org/stable/2984037. See re-

marks [613].

Lewis:1961:IBR

[547] T. Lewis. The intervals between regular events displaced in time by independent random deviations of large dispersion. Journal of the Royal Statistical Society. Series B (Methodological), 23 (2):476–483, ???? 1961. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984038.

Page:1961:ASJ

[548] E. S. Page. An approach to the scheduling of jobs on machines. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):484–492, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984039.

Read:1961:QEI

[549] Robert R. Read. On quadratic estimates of the interclass variance for unbalanced designs. Journal of the Royal Statistical Society. Series B (Methodological), 23(2):493–497, ???? 1961. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984040.

Thatcher:1962:SRI

[550] A. R. Thatcher. Some results on inventory problems. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):1–45, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983744.

Rao:1962:EEO

[551] C. Radhakrishna Rao. Efficient estimates and optimum inference procedures in large samples. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):46–72, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983745.

Gaver:1962:WLI

[552] D. P. Gaver, Jr. A waiting line with interrupted service, including priorities. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):73–90, ???? 1962. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983746.

Jaiswal:1962:TDS

[553] N. K. Jaiswal. Time-dependent solution of the head-of-the-line priority queue. Journal of the Royal Statistical Society. Series B (Methodological), 24 (1):91–101, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983747.

Gani:1962:IFE

[554] J. Gani and R. Pyke. Inequalities for first emptiness probabilities of a dam with ordered inputs. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):102–106, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983748.

Downton:1962:CSI

[555] F. Downton. Congestion systems with incomplete service. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):107–111, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2983749. See corrigenda [682].

Joseph:1962:SSS

[556] A. W. Joseph. Subsidiary sequences for solving Leser’s least-squares graduation equations. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):112–117, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983750.

Fisher:1962:SEB

[557] Ronald Fisher. Some examples of Bayes’ method of the experimental determination of probabilities a priori. Journal of the Royal Statistical Society. Series B (Methodological), 24(1): 118–124, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2983751.

Williams:1962:EFL

[558] E. J. Williams. Exact fiducial limits in non-linear estimation. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):125–139, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983752.

Atiqullah:1962:ENN

[559] M. Atiqullah. On the effect of nonnormality on the estimation of components of variance. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):140–147, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983753.

Bhat:1962:DCQ

[560] B. R. Bhat. On the distribution of certain quadratic forms in normal variates. Journal of the Royal Statistical Society. Series B (Methodological), 24 (1):148–151, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983754.

Rao:1962:NGI

[561] C. Radhakrishna Rao. A note on a generalized inverse of a matrix with applications to problems in mathematical statistics. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):152–158, ???? 1962. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2983755.

Bennett:1962:MST

[562] B. M. Bennett. On multivariate sign tests. Journal of the Royal Statistical Society. Series B (Methodological), 24 (1):159–161, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983756.

Plackett:1962:NIC

[563] R. L. Plackett. A note on interactions in contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):162–166, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983757.

Rohatgi:1962:ASC

[564] Mandakini S. Rohatgi. On the asymptotic sufficiency of certain order statistics. Journal of the Royal Statistical Society. Series B (Methodological), 24 (1):167–176, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983758.

Ruben:1962:NAE

[565] Harold Ruben. A new asymptotic expansion for the normal probability integral and Mill’s ratio. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):177–179, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983759.

Glasser:1962:EPA

[566] Gerald J. Glasser. Estimators for the product of arithmetic means. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):180–184, ???? 1962. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2983760.

Daniels:1962:ESD

[567] H. E. Daniels. The estimation of spectral densities. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):185–198, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983761.

Jenkins:1962:ESW

[568] G. M. Jenkins and J. Chanmugam. The estimation of slope when the errors are autocorrelated. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):199–214, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983762.

Priestley:1962:ASPa

[569] M. B. Priestley. The analysis of stationary processes with mixed spectra — I. Journal of the Royal Statistical Society. Series B (Methodological), 24 (1):215–233, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2983763.

Aitchison:1962:LSR

[570] John Aitchison. Large-sample restricted parametric tests. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):234–250, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983764.

Darroch:1962:IMF

[571] J. N. Darroch. Interactions in multifactor contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):251–263, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2983765.

Stein:1962:CSM

[572] C. M. Stein. Confidence sets for the mean of a multivariate normal distribution. Journal of the Royal Statistical Society. Series B (Methodological), 24 (2):265–296, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984224.

Box:1962:SSA

[573] G. E. P. Box and G. M. Jenkins. Some statistical aspects of adaptive optimization and control. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):297–343, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984225.

Keilson:1962:GBQ

[574] Julian Keilson. The general bulk queue as a Hilbert problem. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):344–358, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2984226. See corrigenda [690].

Ghosal:1962:QS

[575] A. Ghosal. Queues in series. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):359–363, ???? 1962. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984227.

McFadden:1962:LIS

[576] J. A. McFadden. On the lengths of intervals in a stationary point process. Journal of the Royal Statistical Society. Series B (Methodological), 24(2): 364–382, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984228. See

corrigenda [646].

Kingman:1962:QHT

[577] J. F. C. Kingman. On queues in heavy traffic. Journal of the Royal Statistical Society. Series B (Methodological), 24 (2):383–392, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984229.

Marks:1962:SOS

[578] B. L. Marks. Some optimal sequential schemes for estimating the mean of a cumulative normal quantal response curve. Journal of the Royal Statistical Society. Series B (Methodological), 24 (2):393–400, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984230.

Rao:1962:PSR

[579] C. Radhakrishna Rao. Problems of selection with restrictions. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):401–405, ???? 1962. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984231.

Cox:1962:FRT

[580] D. R. Cox. Further results on tests of separate families of hypotheses. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):406–424, ???? 1962. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984232.

Fraser:1962:CFM

[581] D. A. S. Fraser. On the consistency of the fiducial method. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):425–434, ???? 1962. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984233.

Freeman:1962:EMS

[582] G. H. Freeman and J. N. R. Jeffers. Estimation of means and standard errors in the analysis of non-orthogonal experiments by electronic computer. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):435–446, ???? 1962. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984234.

Dar:1962:CSE

[583] S. N. Dar. On the comparison of the sensitivities of experiments. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):447–453, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984235.

Gart:1962:ACL

[584] John J. Gart. Approximate confidence limits for the relative risk. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):454–463, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2984236. See corrigenda [616].

Atiqullah:1962:UCO

[585] M. Atiqullah and D. R. Cox. The use of control observations as an alternative to incomplete block designs. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):464–471, ???? 1962. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984237.

Sandelius:1962:SRP

[586] Martin Sandelius. A simple randomization procedure. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):472–481, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984238.

Rao:1962:SPU

[587] J. N. K. Rao, H. O. Hartley, and W. G. Cochran. On a simple procedure of unequal probability sampling without replacement. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):482–491, ???? 1962. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984239.

Chanda:1962:SPT

[588] K. C. Chanda. Sampling properties of tests of goodness-of-fit for linear autoregressive schemes. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):492–510, ???? 1962. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984240.

Priestley:1962:ASPb

[589] M. B. Priestley. Analysis of stationary processes with mixed spectra — II. Journal of the Royal Statistical Society. Series B (Methodological), 24(2): 511–529, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984241.

Tallis:1962:UGM

[590] G. M. Tallis. The use of a generalized multinomial distribution in the estimation of correlation in discrete data. Journal of the Royal Statistical Society. Series B (Methodological), 24(2): 530–534, ???? 1962. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984242.

Finney:1962:CTM

[591] D. J. Finney. Cumulants of truncated multinormal distributions. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):535–536, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984243.

Gurland:1962:NPR

[592] John Gurland. Note on a paper by ray and Pitman. Journal of the Royal Statistical Society. Series B (Methodological), 24(2):537–538, ???? 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984244.

Anonymous:1963:BMa

[593] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):i–vi, ???? 1963. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984564.

Wetherill:1963:SEQ

[594] G. B. Wetherill. Sequential estimation of quantal response curves. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):1–48, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984541.

Bather:1963:CCM

[595] J. A. Bather. Control charts and the minimization of costs. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):49–80, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984542.

Anscombe:1963:TGF

[596] F. J. Anscombe. Tests of goodness of fit. Journal of the Royal Statistical Society. Series B (Methodological), 25(1): 81–94, ???? 1963. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

                                        jstor.org/stable/2984543.

Williams:1963:CDF

[597] E. J. Williams. A comparison of the direct and fiducial arguments in the estimation of a parameter. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):95–99, ???? 1963. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984544.

Dempster:1963:DP

[598] A. P. Dempster. On direct probabilities. Journal of the Royal Statistical Society. Series B (Methodological), 25 (1):100–110, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984545.

Barnard:1963:SLA

[599] G. A. Barnard. Some logical aspects of the fiducial argument. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):111–114, ???? 1963. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984546.

Fraser:1963:SEF

[600] D. A. S. Fraser. On sufficiency and the exponential family. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):115–123, ???? 1963. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984547.

Barnard:1963:LLS

[601] G. A. Barnard. The logic of least squares. Journal of the Royal Statistical Society. Series B (Methodological), 25 (1):124–127, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984548.

Healy:1963:FLV

[602] M. J. R. Healy. Fiducial limits for a variance component. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):128–130, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984549.

Lloyd:1963:EES

[603] E. H. Lloyd. The epochs of emptiness of a semi-infinite discrete reservoir. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):131–136, ???? 1963. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984550.

Mott:1963:DTE

[604] J. L. Mott. The distribution of the time-to-emptiness of a discrete dam under steady demand. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):137–139, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984551.

Gani:1963:MBG

[605] J. Gani. Models for a bacterial growth process with removals. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):140–149, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984552.

Bartholomew:1963:MSR

[606] D. J. Bartholomew. A multi-stage renewal process. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):150–168, ???? 1963. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984553.

Bentley:1963:CCT

[607] Donald L. Bentley. A contribution to counter theory. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):169–178, ???? 1963. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984554.

Goodman:1963:PTC

[608] Leo A. Goodman. On Plackett’s test for contingency table interactions. Journal of the Royal Statistical Society. Series B (Methodological), 25(1): 179–188, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984555.

Leadbetter:1963:NSP

[609] M. R. Leadbetter. On the normal stationary process — areas outside given levels. Journal of the Royal Statistical Society. Series B (Methodological), 25 (1):189–194, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984556.

Walker:1963:NAE

[610] A. M. Walker. A note on the asymptotic efficiency of an asymptotically normal estimator sequence. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):195–200, ????

1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2984557. See correction [815].

Bennett:1963:CER

[611] B. M. Bennett. On combining estimates of a ratio of means. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):201–205, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984558.

Edwards:1963:EPS

[612] A. W. F. Edwards. Estimation of the parameters in short Markov sequences. Journal of the Royal Statistical Society. Series B (Methodological), 25(1): 206–208, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984559.

Kale:1963:SRM

[613] B. K. Kale. Some remarks on a method of maximum-likelihood estimation proposed by Richards. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):209–212, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984560. See [546].

Ghosh:1963:PTT

[614] M. N. Ghosh and Divakar Sharma. Power of Tukey’s test for nonadditivity. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):213–219, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984561.

Birch:1963:MLT

[615] M. W. Birch. Maximum likelihood in three-way contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):220–233, ???? 1963. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984562.

Gart:1963:CAC

[616] John J. Gart. Corrigenda: Approximate confidence limits for the relative risks. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):234, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984563. See

[584].

Anonymous:1963:FMa

[617] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 25(1):??, ???? 1963. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984540.

Anonymous:1963:VI

[618] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 25(1): ??, ???? 1963. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984539.

Anonymous:1963:BMb

[619] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):i–iv, ???? 1963. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984321.

Scheffe:1963:SCD

[620] Henry Scheffé. The simplex-centroid design for experiments with mixtures. Journal of the Royal Statistical Society. Series B (Methodological), 25(2): 235–263, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984294.

Bartlett:1963:SAP

[621] M. S. Bartlett. The spectral analysis of point processes. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):264–296, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984295.

Johnson:1963:MRP

[622] N. L. Johnson and Rita J. Maurice. A manimax-regret procedure for choosing between two populations using sequential sampling. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):297–304, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984296.

Shenton:1963:HMM

[623] L. R. Shenton and K. Bowman. Higher moments of a maximum-likelihood estimate. Journal of the Royal Statistical Society. Series B (Methodological), 25 (2):305–317, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984297.

Welch:1963:FCP

[624] B. L. Welch and H. W. Peers. On formulae for confidence points based on integrals of weighted likelihoods. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):318–329, ???? 1963. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984298. See note [745].

Halperin:1963:CIE

[625] Max Halperin. Confidence interval estimation in non-linear regression. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):330–333, ???? 1963. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984299.

Atiqullah:1963:RDP

[626] M. Atiqullah. On the randomization distribution and power of the variance ratio test. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):334–347, ???? 1963. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984300.

Kabe:1963:MLH

[627] D. G. Kabe. Multivariate linear hypothesis with linear restrictions. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):348–351, ???? 1963. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984301.

Jowett:1963:AJP

[628] G. H. Jowett. Applications of Jordan’s procedure for matrix inversion in multiple regression and multivariate distance analysis. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):352–357, ???? 1963. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984302.

Ghosh:1963:HGM

[629] M. N. Ghosh. Hotelling’s generalized T^{2} in the multivariate analysis of variance. Journal of the Royal Statistical Society. Series B (Methodological), 25 (2):358–367, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984303.

Geisser:1963:PDM

[630] Seymour Geisser and Jerome Cornfield. Posterior distributions for multivariate normal parameters. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):368–376, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2984304. See comments [671].

Good:1963:IQE

[631] I. J. Good. On the independence of quadratic expressions. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):377–382, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2984305. See corrigenda [797].

Good:1963:QMC

[632] I. J. Good. Quadratics in Markov-chain frequencies, and the binary chain of order 2. Journal of the Royal Statistical Society. Series B (Methodological), 25 (2):383–391, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984306.

Heyde:1963:PLD

[633] C. C. Heyde. On a property of the lognormal distribution. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):392–393, ????

1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984307.

Seth:1963:CUR

[634] Asha Seth. The correlated unrestricted random walk. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):394–400, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984308.

Bishir:1963:LBC

[635] J. Bishir. A lower bound for the critical probability in the one-quadrant oriented- atom percolation process. Journal of the Royal Statistical Society. Series B (Methodological), 25(2): 401–404, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984309.

Ewens:1963:DEP

[636] W. J. Ewens. The diffusion equation and a pseudo-distribution in genetics. Journal of the Royal Statistical Society. Series B (Methodological), 25(2): 405–412, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984310.

McFadden:1963:HOP

[637] J. A. McFadden and Walter Weissblum. Higher-order properties of a stationary point process. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):413–431, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984311.

Bartholomew:1963:ASI

[638] D. J. Bartholomew. An approximate solution of the integral equation of renewal theory. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):432–441, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984312.

Herbst:1963:PAV

[639] Laurence J. Herbst. Periodogram analysis and variance fluctuations. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):442–450, ???? 1963. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984313.

Herbst:1963:TVH

[640] Laurence N. Herbst. A test for variance heterogeneity in the residuals of a Gaussian moving average. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):451–454, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984314.

Fabens:1963:CSQ

[641] A. J. Fabens and A. G. A. D. Perera. A correction to “The Solution of Queueing and Inventory Models by Semi-Markov Processes”. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):455–456, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984315. See [512].

Sack:1963:TNE

[642] R. A. Sack. Treatment of the nonequilibrium theory of simple queues by means of cumulative probabilities. Journal of the Royal Statistical Society. Series B (Methodological), 25(2): 457–463, ???? 1963. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984316.

Keilson:1963:ABQ

[643] J. Keilson. On the asymptotic behaviour of queues. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):464–476, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984317.

Pike:1963:SNR

[644] M. C. Pike. Some numerical results for the queueing system d/E_{k} /1. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):477–488, ???? 1963. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984318.

Gaver:1963:CQI

[645] Donald P. Gaver, Jr. Competitive queueing: Idleness probabilities under priority disciplines. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):489–499, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984319.

McFadden:1963:CLI

[646] J. A. McFadden. Corrigenda: On the lengths of intervals in a stationary point process. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):500, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984320. See [576].

Anonymous:1963:FMb

[647] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 25(2):??, ???? 1963. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984293.

Patterson:1964:TCR

[648] H. D. Patterson. Theory of cyclic rotation experiments. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):1–45, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984602.

Fraser:1964:LUE

[649] D. A. S. Fraser. On local unbiased estimation. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):46–51, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984603.

Fraser:1964:LCS

[650] D. A. S. Fraser. Local conditional sufficiency. Journal of the Royal Statistical Society. Series B (Methodological), 26 (1):52–62, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984604.

Evans:1964:BEV

[651] I. G. Evans. Bayesian estimation of the variance of a normal distribution. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):63–68, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984605.

Geisser:1964:POM

[652] Seymour Geisser. Posterior odds for multivariate normal classifications. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):69–76, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984606.

Srivastava:1964:MPT

[653] J. N. Srivastava. On the monotonicity property of the three main tests for multivariate analysis of variance. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):77–81, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984607.

Tallis:1964:FME

[654] G. M. Tallis. Further models for estimating correlation in discrete data. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):82–85, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984608.

Goodman:1964:SCL

[655] Leo A. Goodman. Simultaneous confidence limits for cross-product ratios in contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):86–102, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984609.

Cox:1964:SAE

[656] D. R. Cox. Some applications of exponential ordered scores. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):103–110, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984610.

Finucan:1964:NK

[657] H. M. Finucan. A note on kurtosis. Journal of the Royal Statistical Society. Series B (Methodological), 26(1): 111–112, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984611.

deBruyn:1964:PPC

[658] C. S. van Dobben de Bruyn. Prediction by progressive correction. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):113–122, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984612.

Priestley:1964:ESD

[659] M. B. Priestley. Estimation of the spectral density function in the presence of harmonic components. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):123–132, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984613.

Buckley:1964:SRF

[660] D. J. Buckley and R. C. Wheeler. Some results for fixed-time traffic signals. Journal of the Royal Statistical Society. Series B (Methodological), 26 (1):133–140, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984614.

Zahl:1964:DMQ

[661] Samuel Zahl. A deformation method for quadratic programming. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):141–160, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984615. See supplement

[711].

Anonymous:1964:FMa

[662] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 26(1):??, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984601.

Anonymous:1964:VI

[663] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 26(1): ??, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984600.

Aitchison:1964:TPC

[664] J. Aitchison. Two papers on the comparison of Bayesian and frequentist approaches to statistical problems of prediction: Bayesian tolerance regions. Journal of the Royal Statistical Society. Series B (Methodological), 26(2): 161–175, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984416.

Thatcher:1964:RBB

[665] A. R. Thatcher. Relationships between Bayesian and confidence limits for predictions. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):176–210, ???? 1964. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984417.

Box:1964:AT

[666] G. E. P. Box and D. R. Cox. An analysis of transformations. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):211–252, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984418.

Fraser:1964:LII

[667] D. A. S. Fraser. On local inference and information. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):253–260, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984419.

Seber:1964:LHI

[668] G. A. F. Seber. The linear hypothesis and idempotent matrices. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):261–266, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984420.

Halperin:1964:NIE

[669] Max Halperin. Note on interval estimation in non-linear regression when responses are correlated. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):267–269, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984421.

Kabe:1964:NBD

[670] D. G. Kabe. A note on the Bartlett decomposition of a Wishart matrix. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):270–273, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984422.

Stone:1964:CPD

[671] M. Stone. Comments on a posterior distribution of Geisser and Cornfield. Journal of the Royal Statistical Society. Series B (Methodological), 26(2): 274–276, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984423. See

[630].

Tiao:1964:BEM

[672] George C. Tiao and Arnold Zellner. On the Bayesian estimation of multivariate regression. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):277–285, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984424.

Patil:1964:CTS

[673] G. P. Patil and V. Seshadri. Characterization theorems for some univariate probability distributions. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):286–292, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984425.

Gibbons:1964:PTSa

[674] Jean D. Gibbons. On the power of two-sample rank tests on the equality of two distribution functions. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):293–304, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984426.

Gibbons:1964:PTSb

[675] Jean D. Gibbons. A proposed twosample rank test: The psi test and its properties. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):305–312, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984427.

Birch:1964:DPA

[676] M. W. Birch. The detection of partial association, I: The 2 × 2 case. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):313–324, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984428.

Lewis:1964:SPC

[677] T. Lewis and L. J. Govier. Some properties of counts of events for certain types of point process. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):325–337, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984429.

Lloyd:1964:NSD

[678] E. H. Lloyd and S. Odoom. A note on the solution of dam equations. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):338–344, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984430.

Green:1964:MRO

[679] J. R. Green. A model for rainfall occurrence. Journal of the Royal Statistical Society. Series B (Methodological), 26 (2):345–353, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984431.

Herbst:1964:SAP

[680] Laurence J. Herbst. Spectral analysis in the presence of variance fluctuations. Journal of the Royal Statistical Society. Series B (Methodological), 26 (2):354–360, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984432.

Herbst:1964:SAF

[681] Laurence J. Herbst. Stationary amplitude fluctuations in a random series. Journal of the Royal Statistical Society. Series B (Methodological), 26(2): 361–364, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984433.

Downton:1964:CCS

[682] F. Downton. Corrigenda: Congestion systems with incomplete service. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):365, ???? 1964. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984434. See [555].

Anonymous:1964:FMb

[683] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 26(2):??, ???? 1964. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984415.

Domb:1964:SSP

[684] C. Domb. Some statistical problems connected with crystal lattices. Journal of the Royal Statistical Society. Series B (Methodological), 26(3):367–397, ???? 1964. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984496.

Lewis:1964:BPP

[685] Peter A. W. Lewis. A branching Poisson process model for the analysis of computer failure patterns. Journal of the Royal Statistical Society. Series B (Methodological), 26(3):398–456, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984497.

Bennett:1964:BSR

[686] B. M. Bennett. A bivariate signed rank test. Journal of the Royal Statistical Society. Series B (Methodological), 26 (3):457–461, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984498.

Aitchison:1964:CRT

[687] J. Aitchison. Confidence-region tests. Journal of the Royal Statistical Society. Series B (Methodological), 26(3): 462–476, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984499.

Geisser:1964:EUC

[688] Seymour Geisser. Estimation in the uniform covariance case. Journal of the Royal Statistical Society. Series B (Methodological), 26(3):477–483, ???? 1964. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984500.

Koop:1964:IVR

[689] J. C. Koop. On an identity for the variances of a ratio of two random variables. Journal of the Royal Statistical Society. Series B (Methodological), 26 (3):484–486, ???? 1964. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984501.

Keilson:1964:CGB

[690] J. Keilson. Corrigenda: The general bulk queue as a Hilbert problem. Journal of the Royal Statistical Society. Series B (Methodological), 26(3):487, ???? 1964. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984502. See [574].

Anonymous:1964:FMc

[691] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 26(3):??, ???? 1964. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984495.

Welch:1965:CBC

[692] B. L. Welch. On comparisons between confidence point procedures in the case of a single parameter. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):1–8, ???? 1965. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984475.

Peers:1965:CPB

[693] H. W. Peers. On confidence points and Bayesian probability points in the case of several parameters. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):9–16, ???? 1965. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984476.

Clutton-Brock:1965:UOE

[694] M. Clutton-Brock. Using the observations to estimate the prior distribution. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):17–27, ???? 1965. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984477.

Cochran:1965:RMM

[695] W. G. Cochran and M. Davis. The Robbins–Monro method for estimating the median lethal dose. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):28–44, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984478.

Chan:1965:CFR

[696] Nai Ng Chan. On circular functional relationships. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):45–56, ???? 1965. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984479.

Choi:1965:PBU

[697] Keewhan Choi. Probability bounds for a union of hyperspherical cones. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):57–73, ???? 1965. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984480.

Tallis:1965:FIW

[698] G. M. Tallis and M. K. Vagholkar. Formulae to improve Wald’s approximation for some properties of sequential tests. Journal of the Royal Statistical Society. Series B (Methodological), 27 (1):74–81, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984481.

Bennett:1965:NAM

[699] B. M. Bennett. Note on a χ^{2}- approximation for the multivariate sign test. Journal of the Royal Statistical Society. Series B (Methodological), 27 (1):82–85, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984482.

Reitsma:1965:DSD

[700] A. Reitsma. The determination of sampling distributions and moment generating functions by solving differential equations. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):86–90, ???? 1965. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984483.

Marsaglia:1965:MCR

[701] G. Marsaglia, A. W. Marshall, and F. Proschan. Moment crossings as related to density crossings. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):91–93, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984484.

Patil:1965:CPE

[702] G. P. Patil and Richard Shorrock. On certain properties of the exponentialtype families. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):94–99, ???? 1965. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984485.

Ali:1965:ABR

[703] S. M. Ali and S. D. Silvey. Association between random variables and the dispersion of a Radon–Nikodym derivative. Journal of the Royal Statistical Society. Series B (Methodological), 27 (1):100–107, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984486. See

corrigenda [739].

Ali:1965:FRR

[704] S. M. Ali and S. D. Silvey. A further result on the relevance of the dispersion of a Radon–Nikodym derivative to the problem of measuring association. Journal of the Royal Statistical Society. Series B (Methodological), 27(1): 108–110, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984487.

Birch:1965:DPA

[705] M. W. Birch. The detection of partial association, II: The general case. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):111–124, ???? 1965. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984488.

Loynes:1965:PRW

[706] R. M. Loynes. On a property of the random walks describing simple queues and dams. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):125–129, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984489.

Takacs:1965:ABT

[707] Lajos Takacs. Applications of a ballot theorem in physics and in order statistics. Journal of the Royal Statistical Society. Series B (Methodological), 27 (1):130–137, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984490.

Heathcote:1965:BPA

[708] C. R. Heathcote. A branching process allowing immigration. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):138–143, ???? 1965. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/

stable/2984491. See corrections and comments [760, 864].

Ray:1965:PPP

[709] W. D. Ray and C. Wyld. Polynomial projecting properties of multiterm Predictors/Controllers in nonstationary time series. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):144–158, ???? 1965. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984492.

Herbst:1965:SFA

[710] L. J. Herbst. The statistical Fourier analysis of variances. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):159–165, ???? 1965. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984493.

Zahl:1965:SDM

[711] Samuel Zahl. Supplement to “A Deformation Method for Quadratic Programming”. Journal of the Royal Statistical Society. Series B (Methodological), 27(1):166–168, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984494. See [661].

Anonymous:1965:VI

[712] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 27(1): ??, ???? 1965. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984474.

Pratt:1965:BIS

[713] John W. Pratt. Bayesian interpretation of standard inference statements. Journal of the Royal Statistical Society. Series B (Methodological), 27(2):169–203, ???? 1965. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984190.

Priestley:1965:ESN

[714] M. B. Priestley. Evolutionary spectra and non-stationary processes. Journal of the Royal Statistical Society. Series B (Methodological), 27(2):204–237, ???? 1965. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984191.

Samuel:1965:SRC

[715] Ester Samuel. On simple rules for the compound decision problem. Journal of the Royal Statistical Society. Series B (Methodological), 27(2):238–244, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984192.

Aitchison:1965:LRC

[716] J. Aitchison. Likelihood-ratio and confidence-region tests. Journal of the Royal Statistical Society. Series B (Methodological), 27(2):245–250, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984193.

Kruskal:1965:AFE

[717] J. B. Kruskal. Analysis of factorial experiments by estimating monotone transformations of the data. Journal of the Royal Statistical Society. Series B (Methodological), 27(2):251–263, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984194.

Sedransk:1965:ASC

[718] J. Sedransk. Analytical surveys with cluster sampling. Journal of the Royal Statistical Society. Series B (Methodological), 27(2):264–278, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984195.

Evans:1965:BEP

[719] I. G. Evans. Bayesian estimation of parameters of a multivariate normal distribution. Journal of the Royal Statistical Society. Series B (Methodological), 27(2):279–283, ???? 1965. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984196.

Savage:1965:FST

[720] I. R. Savage and L. J. Savage. Finite stopping time and finite expected stopping time. Journal of the Royal Statistical Society. Series B (Methodological), 27(2):284–289, ????

1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984197.

Sarndal:1965:DCF

[721] Carl-Erik Särndal. Derivation of a class of frequency distributions via Bayes’s theorem. Journal of the Royal Statistical Society. Series B (Methodological), 27(2):290–300, ???? 1965. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984198.

Tallis:1965:PTN

[722] G. M. Tallis. Plane truncation in normal populations. Journal of the Royal Statistical Society. Series B (Methodological), 27(2):301–307, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984199.

Shenton:1965:MSC

[723] L. R. Shenton and Whitney L. Johnson. Moments of a serial correlation coefficient. Journal of the Royal Statistical Society. Series B (Methodological), 27 (2):308–320, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984200.

Comer:1965:ASA

[724] J. P. Comer, Jr. Application of stochastic approximation to process control. Journal of the Royal Statistical Society. Series B (Methodological), 27(2): 321–331, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984201.

Cox:1965:EIF

[725] D. R. Cox. On the estimation of the intensity function of a stationary point process. Journal of the Royal Statistical Society. Series B (Methodological), 27 (2):332–337, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984202.

Williams:1965:BBD

[726] T. Williams. The basic birth-death model for microbial infections. Journal of the Royal Statistical Society. Series B (Methodological), 27(2):338–360, ???? 1965. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984203.

Welch:1965:BPF

[727] P. D. Welch. On the busy period of a facility which serves customers of several types. Journal of the Royal Statistical Society. Series B (Methodological), 27 (2):361–370, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984204.

Whittle:1965:SGR

[728] P. Whittle. Some general results in sequential design. Journal of the Royal Statistical Society. Series B (Methodological), 27(3):371–394, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345792.

Pyke:1965:S

[729] R. Pyke. Spacings. Journal of the Royal Statistical Society. Series B (Methodological), 27(3):395–449, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345793.

Draper:1965:MDT

[730] Norman R. Draper and Willard Lawrence. Mixture designs for three factors. Journal of the Royal Statistical Society. Series B (Methodological), 27 (3):450–465, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345794.

Albasiny:1965:NSS

[731] E. L. Albasiny. The numerical solution of some non-linear equations, useful in the design of experiments. Journal of the Royal Statistical Society. Series B (Methodological), 27(3):466–472, ???? 1965. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345795.

Draper:1965:MDF

[732] N. R. Draper and W. E. Lawrence. Mixture designs for four factors. Journal of the Royal Statistical Society. Series B (Methodological), 27(3):473–478, ???? 1965. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345796.

Sprott:1965:TS

[733] D. A. Sprott. Transformations and sufficiency. Journal of the Royal Statistical Society. Series B (Methodological), 27 (3):479–485, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345797.

Searle:1965:ARC

[734] S. R. Searle. Additional results concerning estimable functions and generalized inverse matrices. Journal of the Royal Statistical Society. Series B (Methodological), 27(3):486–490, ???? 1965. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345798.

Loynes:1965:WTD

[735] R. M. Loynes. On the waiting-time distribution for queues in series. Journal of the Royal Statistical Society. Series B (Methodological), 27(3):491–496, ???? 1965. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345799.

Morgan:1965:TDP

[736] R. W. Morgan and D. J. A. Welsh. A two-dimensional Poisson growth process. Journal of the Royal Statistical Society. Series B (Methodological), 27 (3):497–504, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345800.

Mehr:1965:CPG

[737] C. B. Mehr and J. A. McFadden. Certain properties of Gaussian processes and their first-passage times. Journal of the Royal Statistical Society. Series B (Methodological), 27(3):505–522, ???? 1965. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345801.

Whittle:1965:RRP

[738] P. Whittle. Recursive relations for predictors of non-stationary processes. Journal of the Royal Statistical Society. Series B (Methodological), 27(3): 523–532, ???? 1965. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345802.

Ali:1965:CAB

[739] S. M. Ali and S. D. Silvey. Corrigenda: Association between random variables and the dispersion of a Randon–Nikodym derivative. Journal of the Royal Statistical Society.

Series B (Methodological), 27(3):533, ???? 1965. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345803. See [703].

Anonymous:1966:BMa

[740] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):i–iv, ???? 1966. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984292.

Hills:1966:ART

[741] M. Hills. Allocation rules and their error rates. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):1–31, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984268.

Hartigan:1966:ERP

[742] J. A. Hartigan. Estimation by ranking parameters. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):32–44, ???? 1966. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984269.

Bhat:1966:EME

[743] B. R. Bhat and N. V. Kulkarni. On efficient multinomial estimation. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):45–52, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984270.

Pinkham:1966:FEC

[744] R. S. Pinkham. On a fiducial example of C. Stein. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):53–54, ???? 1966. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984271.

Hartigan:1966:NCP

[745] J. A. Hartigan. Note on the confidenceprior of Welch and Peers. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):55–56, ???? 1966. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984272. See [624].

Aitchison:1966:ECL

[746] J. Aitchison. Expected-cover and linear-utility tolerance intervals. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):57–62, ???? 1966. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984273.

Samuel:1966:SCR

[747] Ester Samuel. Sequential compound ruls for the finite decision problem. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):63–72, ???? 1966. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984274.

Abramson:1966:ASD

[748] Lee R. Abramson. Asymptotic sequential design of experiments with two random variables. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):73–87, ???? 1966. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984275.

Finney:1966:ESC

[749] D. J. Finney. An experimental study of certain screening processes. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):88–109, ???? 1966. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984276.

Rees:1966:AVD

[750] David H. Rees. The analysis of variance of designs with many nonorthogonal classifications. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):110–117, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984277.

Preece:1966:CYR

[751] D. A. Preece. Classifying Youden rectangles. Journal of the Royal Statistical Society. Series B (Methodological), 28 (1):118–130, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984278.

Ali:1966:GCC

[752] S. M. Ali and S. D. Silvey. A general class of coefficients of divergence of one distribution from another. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):131–142, ???? 1966. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984279.

Menon:1966:CTS

[753] M. V. Menon. Characterization theorems for some univariate probability distributions. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):143–145, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984280.

Bhattacharyya:1966:NAE

[754] G. K. Bhattacharyya. A note on the asymptotic efficiency of Bennett’s bivariate sign test. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):146–149, ???? 1966. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984281.

Armitage:1966:CST

[755] P. Armitage. The chi-square test for heterogeneity of proportions, after adjustment for stratification. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):150–163, ???? 1966. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984282.

Gart:1966:AAC

[756] John J. Gart. Alternative analyses of contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):164–179, ???? 1966. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984283.

Subrahmaniam:1966:TIC

[757] K. Subrahmaniam. A test for “Intrinsic correlation” in the theory of accident proneness. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):180–189, ???? 1966. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/

stable/2984284. See correction and acknowledgement of priority [798].

Keilson:1966:EQL

[758] J. Keilson. The ergodic queue length distribution for queueing systems with finite capacity. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):190–201, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984285.

Hawkes:1966:DTI

[759] A. G. Hawkes. Delay at traffic intersections. Journal of the Royal Statistical Society. Series B (Methodological), 28 (1):202–212, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984286.

Heathcote:1966:CCP

[760] C. R. Heathcote. Corrections and comments on the paper “A Branching Process Allowing Immigration”. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):213–217, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984287. See [708].

Leadbetter:1966:SEM

[761] M. R. Leadbetter. On streams of events and mixtures of streams. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):218–227, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984288.

Priestley:1966:DRN

[762] M. B. Priestley. Design relations for non-stationary processes. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):228–240, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984289.

Jones:1966:ESM

[763] Richard H. Jones. Exponential smoothing for multivariate time series. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):241–251, ???? 1966. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984290.

Smith:1966:CCS

[764] Cedric A. B. Smith. Correction: Consistency in statistical inference and decision. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):252, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984291. See

[507].

Anonymous:1966:FMa

[765] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 28(1):??, ???? 1966. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984267.

Anonymous:1966:VI

[766] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 28(1): ??, ???? 1966. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984266.

Anonymous:1966:BMb

[767] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 28(2):i–v, ???? 1966. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984382.

Seneta:1966:QSD

[768] E. Seneta. Quasi-stationary distributions and time-reversion in genetics. Journal of the Royal Statistical Society. Series B (Methodological), 28(2): 253–277, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984371.

Sprent:1966:GLS

[769] P. Sprent. A generalized least-squares approach to linear functional relationships. Journal of the Royal Statistical Society. Series B (Methodological), 28 (2):278–297, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984372.

Krishnan:1966:LUT

[770] Marakatha Krishnan. Locally unbiased type M test. Journal of the Royal Statistical Society. Series B (Methodological), 28(2):298–309, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984373.

Godambe:1966:NASa

[771] V. P. Godambe. A new approach to sampling from finite populations. I. Sufficiency and linear estimation. Journal of the Royal Statistical Society. Series B (Methodological), 28(2):310–319, ???? 1966. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984374.

Godambe:1966:NASb

[772] V. P. Godambe. A new approach to sampling from finite populations. II. Distribution-free sufficiency. Journal of the Royal Statistical Society. Series B (Methodological), 28(2):320–328, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984375.

Hasofer:1966:AFDa

[773] A. M. Hasofer. The almost full dam with Poisson input. Journal of the Royal Statistical Society. Series B (Methodological), 28(2):329–335, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984376.

Jenkins:1966:CSD

[774] J. H. Jenkins. On the correlation structure of the departure process of the M/E_{λ} /1 queue. Journal of the Royal Statistical Society. Series B (Methodological), 28(2):336–344, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984377.

John:1966:CIB

[775] J. A. John. Cyclic incomplete block designs. Journal of the Royal Statistical Society. Series B (Methodological), 28 (2):345–360, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984378.

John:1966:ETA

[776] Peter W. M. John. An extension of the triangular association scheme to three associate classes. Journal of the Royal Statistical Society. Series B (Methodological), 28(2):361–365, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984379.

John:1966:EPC

[777] S. John. On the evaluation of probabilities of convex polyhedra under multivariate normal and t distributions. Journal of the Royal Statistical Society. Series B (Methodological), 28(2): 366–369, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984380.

Srivastava:1966:SAE

[778] M. S. Srivastava. Some asymptotically efficient sequential procedures for ranking and slippage problems. Journal of the Royal Statistical Society. Series B (Methodological), 28(2):370–380, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984381.

Anonymous:1966:FMb

[779] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 28(2):??, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984370.

Anonymous:1966:BMc

[780] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):i–v, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984454.

Wetherill:1966:DTA

[781] G. B. Wetherill and G. E. G. Campling. The decision theory approach to sampling inspection. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):381–416, ????

1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984436.

Kingman:1966:ASM

[782] J. F. C. Kingman. An approach to the study of Markov processes. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):417–447, ???? 1966. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984437.

Hasofer:1966:AFDb

[783] A. M. Hasofer. The almost full dam with Poisson input: Further results. Journal of the Royal Statistical Society. Series B (Methodological), 28(3): 448–455, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984438.

Shanbhag:1966:GQS

[784] D. N. Shanbhag. On a generalized queueing system with Poisson arrivals. Journal of the Royal Statistical Society. Series B (Methodological), 28(3): 456–458, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984439.

Downton:1966:RMS

[785] F. Downton. The reliability of multiplex systems with repair. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):459–476, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984440.

Keilson:1966:TDP

[786] Julian Keilson. A technique for discussing the passage time distribution for stable systems. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):477–486, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984441.

Cane:1966:NSE

[787] Violet R. Cane. A note on the size of epidemics and the number of people hearing a rumour. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):487–490, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984442.

Hammersley:1966:FPP

[788] J. M. Hammersley. First-passage percolation. Journal of the Royal Statistical Society. Series B (Methodological), 28 (3):491–496, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984443.

Loynes:1966:SAE

[789] R. M. Loynes. Some aspects of the estimation of quantiles. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):497–512, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984444.

Ruben:1966:SNR

[790] Harold Ruben. Some new results on the distribution of the sample correlation coefficient. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):513–525, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984445.

Calinski:1966:DTS

[791] T. Calinski. On the distribution of the F -type statistics in the analysis of a group of experiments. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):526–542, ???? 1966. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984446.

Box:1966:SAR

[792] G. E. P. Box and Irwin Guttman. Some aspects of randomization. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):543–558, ???? 1966. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984447.

Kshirsagar:1966:BFD

[793] A. M. Kshirsagar. Balanced factorial designs. Journal of the Royal Statistical Society. Series B (Methodological), 28 (3):559–567, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984448.

Dunsmore:1966:BAC

[794] I. R. Dunsmore. A Bayesian approach to classification. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):568–577, ???? 1966. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984449.

Good:1966:DPE

[795] I. J. Good. A derivation of the probabilistic explication of information. Journal of the Royal Statistical Society. Series B (Methodological), 28 (3):578–581, ???? 1966. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984450.

Shanbhag:1966:IQF

[796] D. N. Shanbhag. On the independence of quadratic forms. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):582–583, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984451.

Good:1966:IQE

[797] I. J. Good. On the independence of quadratic expressions: Corrigenda. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):584, ???? 1966. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984452. See [631].

Subrahmaniam:1966:CAP

[798] K. Subrahmaniam. Correction and acknowledgement of priority: A Test for “Intrinsic Correlation” in the Theory of Accident Proneness. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):585, ???? 1966. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984453. See [757].

Anonymous:1966:FMc

[799] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 28(3):??, ???? 1966. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984435.

Anonymous:1967:BMa

[800] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series B (Methodological), 29(1):i–vi, ???? 1967. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984583.

Anscombe:1967:TIL

[801] F. J. Anscombe. Topics in the investigation of linear relations fitted by the method of least squares. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):1–52, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984567.

Bartholomew:1967:HTW

[802] D. J. Bartholomew. Hypothesis testing when the sample size is treated as a random variable. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):53–82, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984568.

Guttman:1967:UCF

[803] Irwin Guttman. The use of the concept of a future observation in goodness-of-fit problems. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):83–100, ???? 1967. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984569.

Davies:1967:LAU

[804] M. Davies. Linear approximation using the criterion of least total deviations. Journal of the Royal Statistical Society. Series B (Methodological), 29 (1):101–109, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984570.

Bohrer:1967:SSB

[805] Robert Bohrer. On sharpening Scheffé bounds. Journal of the Royal Statistical Society. Series B (Methodological), 29 (1):110–114, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984571.

Kokan:1967:OAM

[806] A. R. Kokan and Sanaullah Khan. Optimum allocation in multivariate surveys: An analytical solution. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):115–125, ???? 1967. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984572.

Halperin:1967:GFT

[807] Max Halperin. A generalization of Fieller’s theorem to the ratio of complex parameters. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):126–131, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984573.

Srivastava:1967:FWC

[808] M. S. Srivastava. On fixed-width confidence bounds for regression parameters and mean vector. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):132–140, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984574.

Heyde:1967:ARR

[809] C. C. Heyde. Asymptotic renewal results for a natural generalization of classical renewal theory. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):141–150, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984575.

Green:1967:MMR

[810] J. R. Green. A modified model for rainfall occurrence. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):151–153, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984576.

Geary:1967:EPD

[811] R. C. Geary. Ex post determination of significance in multivariate regression when the independent variables are orthogonal. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):154–161, ???? 1967. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984577.

Hald:1967:APB

[812] A. Hald. Asymptotic properties of Bayesian single sampling plans. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):162–173, ???? 1967. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984578. See corrigenda

[849].

Herzberg:1967:BVF

[813] Agnes M. Herzberg. The behaviour of the variance function of the difference between two estimated responses. Journal of the Royal Statistical Society. Series B (Methodological), 29(1): 174–179, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984579.

McNeil:1967:ECS

[814] D. R. McNeil. Estimating the covariance and spectral density functions from a clipped stationary time series. Journal of the Royal Statistical Society. Series B (Methodological), 29(1): 180–195, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984580.

Walker:1967:CNA

[815] A. M. Walker. Correction: A Note on the Asymptotic Efficiency of an Asymptotically Normal Estimator Sequence. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):196, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984581. See

[610].

Armitage:1967:ACS

[816] P. Armitage. Addendum: The chisquare test for heterogeneity of proportions, after adjustment for stratification. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):197, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984582.

Anonymous:1967:FMa

[817] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):??, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984566.

Anonymous:1967:VI

[818] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 29(1):

??, ???? 1967. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984565.

Anonymous:1967:BMb

[819] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):i–iv, ???? 1967. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984599.

Williams:1967:AAA

[820] E. J. Williams. The analysis of association among many variates. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):199–242, ???? 1967. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984585.

Clatworthy:1967:JCI

[821] Willard H. Clatworthy. On John’s cyclic incomplete block designs. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):243–247, ???? 1967. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984586.

Kemp:1967:FEW

[822] Kenneth W. Kemp. Formal expressions which can be used for the determination of the operating characteristic and average sample number of a simple sequential test. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):248–262, ???? 1967. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984587.

Kemp:1967:SPD

[823] Kenneth W. Kemp. A simple procedure for determining upper and lower limits for the average sample run length of a cumulative sum scheme. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):263–265, ???? 1967. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984588.

Fisk:1967:MSK

[824] P. R. Fisk. Models of the second kind in regression analysis. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):266–281, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984589.

Srivastava:1967:ICP

[825] S. R. Srivastava and T. A. Bancroft. Inferences concerning a population correlation coefficient from one or possibly two samples subsequent to a preliminary test of significance. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):282–291, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984590.

Banerjee:1967:SSG

[826] K. S. Banerjee and W. T. Federer. On a special subset giving an irregular fractional replicate of a 2^{n} factorial experiment. Journal of the Royal Statistical Society. Series B (Methodological), 29 (2):292–299, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984591.

Dowson:1967:OHR

[827] D. C. Dowson. Optimization of a hot rolling mill. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):300–319, ???? 1967. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984592.

Mardia:1967:NPT

[828] K. V. Mardia. A non-parametric test for the bivariate two-sample location problem. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):320–342, ???? 1967. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984593.

Lewis:1967:NHB

[829] Peter A. W. Lewis. Non-homogeneous branching Poisson processes. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):343–354, ???? 1967. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984594.

Gilchrist:1967:MEI

[830] W. G. Gilchrist. Methods of estimation involving discounting. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):355–369, ???? 1967. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984595.

Berenblut:1967:CDT

[831] I. I. Berenblut. A change-over design for testing a treatment factor at four equally spaced levels. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):370–373, ????

1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2984596. See corrigenda [848].

Hanurav:1967:OUA

[832] T. V. Hanurav. Optimum utilization of auxiliary information: πps sampling of two units from a stratum. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):374–391, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984597. See addenda and

corrigenda [929].

Rao:1967:CSR

[833] T. J. Rao. On the choice of a strategy for the ratio method of estimation. Journal of the Royal Statistical Society. Series B (Methodological), 29(2): 392–397, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984598.

Anonymous:1967:FMb

[834] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 29(2):??, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984584.

Anonymous:1967:BMc

[835] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):i–vi, ???? 1967. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984399.

Good:1967:BST

[836] I. J. Good. A Bayesian significance test for multinomial distributions. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):399–431, ???? 1967. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984384. See [1165].

Borch:1967:TR

[837] Karl Borch. The theory of risk. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):432–467, ???? 1967. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984385.

Bennett:1967:THC

[838] B. M. Bennett. Tests of hypotheses concerning matched samples. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):468–474, ???? 1967. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984386. See corrigen-

dum [930].

Kounias:1967:CST

[839] Eustratios Kounias. Considering statistical and time averages in a regulation problem. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):475–488, ???? 1967. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984387.

McFadden:1967:CGP

[840] J. A. McFadden. On a class of Gaussian processes for which the mean rate of crossings is infinite. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):489–502, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984388.

Ericson:1967:ECE

[841] W. A. Ericson. On the economic choice of experiment sizes for decision regarding certain linear combinations. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):503–512, ???? 1967. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984389.

Lomnicki:1967:DPR

[842] Z. A. Lomnicki. On the distribution of products of random variables. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):513–524, ???? 1967. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984390.

El-Sayyad:1967:EPE

[843] G. M. El-Sayyad. Estimation of the parameter of an exponential distribution. Journal of the Royal Statistical Society. Series B (Methodological), 29 (3):525–532, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984391.

Hartigan:1967:LIP

[844] J. A. Hartigan. The likelihood and invariance principles. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):533–539, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984392.

Jackson:1967:ADE

[845] O. A. Y. Jackson. An analysis of departures from the exponential distribution. Journal of the Royal Statistical Society. Series B (Methodological), 29 (3):540–549, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984393.

Barton:1967:FLW

[846] D. E. Barton and F. N. David. Fourletter words: The distribution of pattern frequencies in ring permutations. Journal of the Royal Statistical Society. Series B (Methodological), 29(3): 550–569, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984394.

Abdrabbo:1967:PNS

[847] N. A. Abdrabbo and M. B. Priestley. On the prediction of non-stationary processes. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):570–585, ???? 1967. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984395.

Berenblut:1967:CCD

[848] I. I. Berenblut. Corrigenda: A Change-Over Design for Testing a Treatment Factor at Four Equally Spaced Levels. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):586, ???? 1967. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984397. See [831].

Hald:1967:CAP

[849] A. Hald. Corrigenda: Asymptotic properties of Bayesian single sampling plans. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):586, ???? 1967. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984396. See

[812].

Davies:1967:ALA

[850] M. Davies. Acknowledgement: Linear approximation using the criterion of least total deviations. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):587, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984398.

Anonymous:1967:FMc

[851] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 29(3):??, ???? 1967. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984383.

Anonymous:1968:BMa

[852] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):i–iii, ???? 1968. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984473.

Loynes:1968:CSN

[853] R. M. Loynes. On the concept of the spectrum for non-stationary processes. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):1–30, ???? 1968. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984457.

Lindley:1968:CVM

[854] D. V. Lindley. The choice of variables in multiple regression. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):31–66, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984458.

Haitovsky:1968:MDR

[855] Yoel Haitovsky. Missing data in regression analysis. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):67–82, ???? 1968. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984459.

Mardia:1968:SSP

[856] K. V. Mardia. Small sample power of a non-parametric test for the bivariate two-sample location problem in the normal case. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):83–92, ???? 1968. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984460.

McGilchrist:1968:PCT

[857] C. A. McGilchrist. Plant competition: Three species per pot. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):93–107, ???? 1968. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984461.

DeGroot:1968:SPO

[858] Morris H. DeGroot. Some problems of optimal stopping. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):108–122, ???? 1968. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984462.

Lambrakis:1968:EMG

[859] D. P. Lambrakis. Experiments with mixtures: A generalization of the simplex-lattice design. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):123–136, ????

1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984463.

Lambrakis:1968:ECM

[860] D. P. Lambrakis. Experiments with p-component mixtures. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):137–144, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984464.

Brillinger:1968:ECS

[861] David R. Brillinger. Estimation of the cross-spectrum of a stationary bivariate Gaussian process from its zeros. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):145–159, ???? 1968. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984465.

Kemp:1968:DAC

[862] Adrienne W. Kemp and C. D. Kemp. On a distribution associated with certain stochastic processes. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):160–163, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984466.

Ewens:1968:SAM

[863] W. J. Ewens. Some applications of multiple-type branching processes in population genetics. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):164–175, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984467.

Seneta:1968:SDB

[864] E. Seneta. The stationary distribution of a branching process allowing immigration: A remark on the critical case. Journal of the Royal Statistical Society. Series B (Methodological), 30(1): 176–179, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984468. See

[708].

Rosenhead:1968:EQT

[865] Jonathan Rosenhead. An extension of Quenouille’s test for the compatibility of correlation structures in time series. Journal of the Royal Statistical Society. Series B (Methodological), 30(1): 180–184, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984469.

Mercer:1968:QRA

[866] A. Mercer. A queue with random arrivals and scheduled bulk departures. Journal of the Royal Statistical Society. Series B (Methodological), 30(1): 185–189, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984470.

Lindley:1968:BEL

[867] D. V. Lindley and G. M. El-Sayyad. The Bayesian estimation of a linear functional relationships. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):190–202, ???? 1968. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984471.

Good:1968:CWE

[868] I. J. Good. Corrigendum: Weight of evidence, corroboration, explanatory power, information and the utility of experiments. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):203, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984472. See [496].

Anonymous:1968:FMa

[869] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 30(1):??, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984456.

Anonymous:1968:VI

[870] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 30(1): ??, ???? 1968. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984455.

Anonymous:1968:BMb

[871] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):i–iv, ???? 1968. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984521.

Dempster:1968:GBI

[872] A. P. Dempster. A generalization of Bayesian inference. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):205–247, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984504.

Cox:1968:GDR

[873] D. R. Cox and E. J. Snell. A general definition of residuals. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):248–275, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984505.

Gurland:1968:RSF

[874] John Gurland. A relatively simple form of the distribution of the multiple correlation coefficient. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):276–283, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984506.

Cox:1968:NEL

[875] D. R. Cox and D. V. Hinkley. A note on the efficiency of least-squares estimates. Journal of the Royal Statistical Society. Series B (Methodological), 30 (2):284–289, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984507.

Box:1968:ORO

[876] M. J. Box. The occurrence of replications in optimal designs of experiments to estimate parameters in non-linear models. Journal of the Royal Statistical Society. Series B (Methodological), 30 (2):290–302, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984508.

Nelder:1968:CIG

[877] J. A. Nelder. The combination of information in generally balanced designs. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):

303–311, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984509.

Sen:1968:AET

[878] Pranab Kumar Sen. Asymptotically efficient tests by the method of n rankings. Journal of the Royal Statistical Society. Series B (Methodological), 30 (2):312–317, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984510.

Blight:1968:NME

[879] B. J. N. Blight. A note on a modified exponentially weighted predictor. Journal of the Royal Statistical Society. Series B (Methodological), 30(2): 318–320, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984511.

Vere-Jones:1968:SAP

[880] D. Vere-Jones. Some applications of probability generating functionals to the study of input-output streams. Journal of the Royal Statistical Society. Series B (Methodological), 30(2): 321–333, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984512.

Mehr:1968:ABE

[881] C. B. Mehr. An application of biorthonormal expansions in theory of stochastic processes. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):334–337, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984513.

McNeil:1968:APM

[882] D. R. McNeil. The asymptotic powers of multivariate tests with grouped data. Journal of the Royal Statistical Society. Series B (Methodological), 30 (2):338–348, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984514.

Becker:1968:MRM

[883] N. G. Becker. Models for the response of a mixture. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):349–358, ???? 1968. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984515.

Hald:1968:MBD

[884] A. Hald. The mixed binomial distribution and the posterior distribution of p for a continuous prior distribution. Journal of the Royal Statistical Society. Series B (Methodological), 30(2): 359–367, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984516.

Bennett:1968:NTM

[885] B. M. Bennett. Note on X^{2} tests for matched samples. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):368–370, ???? 1968. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984517.

Raktoe:1968:UAC

[886] B. L. Raktoe and W. T. Federer. A unified approach for constructing a useful class of non-orthogonal man effect plans in k^{n} factorials. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):371–380, ???? 1968. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984518. See corrigenda

[1117].

Dowson:1968:LCL

[887] D. C. Dowson. On the linear control of a linear system having a normal stationary stochastic input. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):381–395, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984519.

Dunsmore:1968:BAC

[888] I. R. Dunsmore. A Bayesian approach to calibration. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):396–405, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984520.

Anonymous:1968:FMb

[889] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 30(2):??, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984503.

Anonymous:1968:BMc

[890] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):i–iv, ???? 1968. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984265.

Marshall:1968:GAS

[891] Albert W. Marshall and Ingram Olkin. A general approach to some screening and classification problems. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):407–443, ???? 1968. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984246.

Choi:1968:EPM

[892] Keewhan Choi and W. G. Bulgren. An estimation procedure for mixtures of distributions. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):444–460, ???? 1968. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/

stable/2984247. See comment [1039].

Scott:1968:MST

[893] A. Scott. A multi-stage test for a normal mean. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):461–468, ???? 1968. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984248.

Hager:1968:CDP

[894] H. Hager and C. Antle. The choice of the degree of a polynomial model. Journal of the Royal Statistical Society. Series B (Methodological), 30(3): 469–471, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984249.

Kabir:1968:EPF

[895] A. B. M. Lutful Kabir. Estimation of parameters of a finite mixture of distributions. Journal of the Royal Statistical Society. Series B (Methodological), 30 (3):472–482, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984250.

Bennett:1968:ROT

[896] B. M. Bennett. Rank-order tests of linear hypotheses. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):483–489, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984251.

Hannan:1968:LSE

[897] E. J. Hannan. Least-squares efficiency for vector time series. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):490–498, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984252.

Kingman:1968:ETS

[898] J. F. C. Kingman. The ergodic theory of subadditive stochastic processes. Journal of the Royal Statistical Society. Series B (Methodological), 30(3): 499–510, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984253.

Wagle:1968:MBD

[899] B. Wagle. Multivariate beta distribution and a test for multivariate normality. Journal of the Royal Statistical Society. Series B (Methodological), 30 (3):511–516, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984254.

Plackett:1968:RP

[900] R. L. Plackett. Random permutations. Journal of the Royal Statistical Society. Series B (Methodological), 30(3): 517–534, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984255.

Peers:1968:CPB

[901] H. W. Peers. Confidence properties of Bayesian interval estimates. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):535–544, ???? 1968. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984256.

Hartigan:1968:NDO

[902] J. A. Hartigan. Note on discordant observations. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):545–550, ???? 1968. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984257.

Joshi:1968:DFS

[903] V. M. Joshi. Distribution-free sufficiency in sampling finite populations. Journal of the Royal Statistical Society. Series B (Methodological), 30(3): 551–555, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984258.

Vijayan:1968:ESS

[904] K. Vijayan. An exact πps sampling scheme-generalization of a method of hanurav. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):556–566, ???? 1968. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984259.

Searle:1968:RSE

[905] S. R. Searle. A remark on solving equations in sums of powers. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):567–569, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984260.

Walker:1968:NAD

[906] A. M. Walker. A note on the asymptotic distribution of sample quantiles. Journal of the Royal Statistical Society. Series B (Methodological), 30(3): 570–575, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984261.

Cinlar:1968:SPP

[907] E. Cinlar and R. A. Agnew. On the superposition of point processes. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):576–581, ???? 1968. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984262.

Hope:1968:SMC

[908] Adery C. A. Hope. A simplified Monte Carlo significance test procedure. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):582–598, ???? 1968. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984263.

John:1968:CTR

[909] S. John. A central tolerance region for the multivariate normal distribution. Journal of the Royal Statistical Society. Series B (Methodological), 30 (3):599–601, ???? 1968. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984264.

Anonymous:1968:FMc

[910] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 30(3):??, ???? 1968. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984245.

Anonymous:1969:BMa

[911] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):i–iii, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984343.

Plackett:1969:SMC

[912] R. L. Plackett. Stochastic models of capital investment. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):1–28, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984324.

Novick:1969:MBI

[913] Melvin R. Novick. Multiparameter Bayesian indifference procedures. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):29–64, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984325.

Basu:1969:STS

[914] A. P. Basu. On some tests for several linear relations. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):65–71, ????

1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984326.

Heyde:1969:EFR

[915] C. C. Heyde. On extremal factorization and recurrent events. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):72–79, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984327.

Walker:1969:ABP

[916] A. M. Walker. On the asymptotic behaviour of posterior distributions. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):80–88, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984328.

Hill:1969:FTL

[917] B. M. Hill. Foundations for the theory of least squares. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):89–97, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984329.

Mardia:1969:NDN

[918] K. V. Mardia. On the null distribution of a non-parametric test for the bivariate two- sample problem. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):98–102, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984330.

Loynes:1969:CSG

[919] R. M. Loynes. On Cox and Snell’s general definition of residuals. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):103–106, ???? 1969. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984331.

Becker:1969:RPW

[920] N. G. Becker. Regression problems when the predictor variables are proportions. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):107–112, ???? 1969. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984332.

Hudson:1969:LSF

[921] Derek J. Hudson. Least-squares fitting of a polynomial constrained to be either non-negative non-decreasing or convex. Journal of the Royal Statistical Society. Series B (Methodological), 31(1): 113–118, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984333.

Scott:1969:NAP

[922] Alastair Scott. A note on an allocation problem. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):119–122, ???? 1969. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984334.

Anderson:1969:CDB

[923] J. A. Anderson. Constrained discrimination between k populations. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):123–139, ???? 1969. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984335.

Priestley:1969:TNS

[924] M. B. Priestley and T. Subba Rao. A test for non-stationarity of time-series. Journal of the Royal Statistical Society. Series B (Methodological), 31(1): 140–149, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984336.

Abdrabbo:1969:FNS

[925] N. A. Abdrabbo and M. B. Priestley. Filtering non-stationary signals. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):150–159, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984337.

Dunsmore:1969:RO

[926] I. R. Dunsmore. Regulation and optimization. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):160–170, ???? 1969. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984338.

Siskind:1969:MSD

[927] V. Siskind. Multivariate stochastic differential and difference equations with periodic coefficients. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):171–180, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984339.

Sclove:1969:EPC

[928] S. L. Sclove and J. van Ryzin. Estimating the parameters of a convolution. Journal of the Royal Statistical Society. Series B (Methodological), 31 (1):181–191, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984340.

Hanurav:1969:ACO

[929] T. V. Hanurav. Addenda and corrigenda: Optimum utilization of auxiliary information: πps sampling of two units from a stratum. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):192–194, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984341. See [832].

Bennett:1969:CTH

[930] B. M. Bennett. Corrigendum: Tests of hypotheses concerning matched samples. Journal of the Royal Statistical Society. Series B (Methodological), 31 (1):194, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984342. See

[838].

Anonymous:1969:FMa

[931] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 31(1):??, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984323.

Anonymous:1969:VI

[932] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 31(1): ??, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984322.

Anonymous:1969:BMb

[933] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):i–iii, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984223.

Ericson:1969:SBM

[934] W. A. Ericson. Subjective Bayesian models in sampling finite populations. Journal of the Royal Statistical Society. Series B (Methodological), 31(2): 195–233, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984206.

Lambrakis:1969:EMA

[935] D. P. Lambrakis. Experiments with mixtures: An alternative to the simplex-lattice design. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):234–245, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984207.

Godambe:1969:FAA

[936] V. P. Godambe. A fiducial argument with application to survey sampling. Journal of the Royal Statistical Society. Series B (Methodological), 31(2): 246–260, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984208.

Altham:1969:EBA

[937] Patricia M. E. Altham. Exact Bayesian analysis of a 2 × 2 contingency table, and Fisher’s “Exact” significance test. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):

261–269, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984209.

John:1969:SNO

[938] Peter W. M. John. Some nonorthogonal fractions of 2^{n} designs. Journal of the Royal Statistical Society. Series B (Methodological), 31(2): 270–275, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984210.

Lambrakis:1969:EME

[939] D. P. Lambrakis. Experiments with mixtures: Estimated regression function of the multiple- lattice design. Journal of the Royal Statistical Society. Series B (Methodological), 31(2): 276–284, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984211.

Billard:1969:SPT

[940] L. Billard and M. K. Vagholkar. A sequential procedure for testing a null hypothesis against a two-sided alternative hypothesis. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):285–294, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984212.

Leadbetter:1969:DTB

[941] M. R. Leadbetter. On the distributions of the times between events in a stationary stream of events. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):295–302, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984213.

Holgate:1969:MCN

[942] P. Holgate. Majorants of the chromatic number of a random graph. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):303–309, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984214.

Green:1969:ICP

[943] J. R. Green. Inference concerning probabilities and quantiles. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):310–316, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984215.

Fraser:1969:SPP

[944] D. A. S. Fraser and M. Safiul Haq. Structural probability and prediction for the multivariate model. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):317–331, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984216.

Ericson:1969:NPM

[945] W. A. Ericson. A note on the posterior mean of a population mean. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):332–334, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984217.

Downton:1969:IEA

[946] F. Downton. An integral equation approach to equipment failure. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):335–349, ????

1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984218.

Ramsey:1969:TSE

[947] J. B. Ramsey. Tests for specification errors in classical linear leastsquares regression analysis. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):350–371, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984219.

Solari:1969:MLS

[948] Mary E. Solari. The “Maximum likelihood solution” of the problem of estimating a linear functional relationship. Journal of the Royal Statistical Society. Series B (Methodological), 31(2): 372–375, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984220.

Stromberg:1969:DAM

[949] John L. Stromberg. Distinguishing among multinomial populations: A decision theoretic approach. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):376–387, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984221.

Peritz:1969:TLC

[950] E. Peritz and E. Shlom. Testing for linear contagion–inverse sampling. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):388–395, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984222.

Anonymous:1969:FMb

[951] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 31(2):??, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984205.

Anonymous:1969:BMc

[952] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):i–ii, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984358.

Copas:1969:CDE

[953] J. B. Copas. Compound decisions and empirical Bayes. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):397–425, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984345.

Kabe:1969:LCN

[954] D. G. Kabe. Linear compounds of normal linear regression coefficients. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):426–431, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984346.

Hunter:1969:TQP

[955] Jeffrey J. Hunter. Two queues in parallel. Journal of the Royal Statistical Society. Series B (Methodological), 31 (3):432–445, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984347.

Hartigan:1969:LBM

[956] J. A. Hartigan. Linear Bayesian methods. Journal of the Royal Statistical Society. Series B (Methodological), 31 (3):446–454, ???? 1969. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984348.

Hald:1969:APB

[957] A. Hald and N. Keiding. Asymptotic properties of Bayesian decision rules for two terminal decisions and multiple sampling. I. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):455–471, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984349.

Draper:1969:DTT

[958] N. R. Draper and D. R. Cox. On distributions and their transformation to normality. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):472–476, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984350.

Kshirsagar:1969:CBT

[959] A. M. Kshirsagar. Correlation between two vector variables. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):477–485, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984351.

Goodman:1969:PDP

[960] Leo A. Goodman. On partitioning χ^{2} and detecting partial association in three-way contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):486–498, ???? 1969. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984352.

Seshadri:1969:TED

[961] V. Seshadri, M. Csorgo, and M. A. Stephens. Tests for the exponential distribution using Kolmogorov-type statistics. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):499–509, ???? 1969. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984353.

Solomon:1969:NSE

[962] Daniel L. Solomon. A note on sequential estimation. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):510–513, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984354.

Margolin:1969:RIF

[963] Barry H. Margolin. Resolution IV fractional factorial designs. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):514–523, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984355.

Davies:1969:BOT

[964] Robert B. Davies. Beta-optimal tests and an application to the summary evaluation of experiments. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):524–538, ???? 1969. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984356.

Silvey:1969:MIE

[965] S. D. Silvey. Multicollinearity and imprecise estimation. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):539–552, ???? 1969. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984357.

Anonymous:1969:FMc

[966] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 31(3):??, ???? 1969. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984344.

Vere-Jones:1970:SME

[967] D. Vere-Jones. Stochastic models for earthquake occurrence. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):1–62, ???? 1970. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984402.

Altham:1970:MAR

[968] Patricia M. E. Altham. The measurement of association of rows and columns for an r × s contingency table. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):63–73, ???? 1970. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984403.

Mardia:1970:BNP

[969] K. V. Mardia. A bivariate nonparametric c-sample test. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):74–87, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984404.

Gnanadesikan:1970:PPP

[970] R. Gnanadesikan and M. B. Wilk. A probability plotting procedure for general analysis of variance. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):88–101, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984405.

Kingman:1970:ITQ

[971] J. F. C. Kingman. Inequalities in the theory of queues. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):102–110, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984406.

Mazumdar:1970:PQH

[972] Sati Mazumdar. On priority queues in heavy traffic. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):111–114, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984407.

Stephens:1970:UKS

[973] M. A. Stephens. Use of the Kolmogorov–Smirnov, Cramér–von Mises and related statistics without extensive tables. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):115–122, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984408.

Gulati:1970:BUT

[974] B. R. Gulati and E. G. Kounias. On bounds useful in the theory of symmetrical factorial designs. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):123–133, ???? 1970. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984409.

Loynes:1970:ARE

[975] R. M. Loynes. On the asymptotic relative efficiencies of certain location parameter estimates. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):134–136, ???? 1970. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984410.

John:1970:UGI

[976] J. A. John. Use of generalized inverse matrices in MANOVA. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):137–143, ???? 1970. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984411.

Srivastava:1970:SAB

[977] M. S. Srivastava. On a sequential analogue of the behrens–Fisher problem. Journal of the Royal Statistical Society. Series B (Methodological), 32(1): 144–148, ???? 1970. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984412.

Seneta:1970:ELT

[978] E. Seneta. An explicit-limit theorem for the critical Galton–Watson process with immigration. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):149–152, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984413.

Anonymous:1970:BMa

[979] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):153–iv, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984414.

Anonymous:1970:FMa

[980] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 32(1):??, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984401.

Anonymous:1970:VI

[981] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 32(1): ??, ???? 1970. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984400.

Anonymous:1970:BMb

[982] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):i–ii, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984538.

Edwards:1970:EBP

[983] A. W. F. Edwards. Estimation of the branch points of a branching diffusion process. Journal of the Royal Statistical Society. Series B (Methodological), 32 (2):155–174, ???? 1970. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984523.

Kalbfleisch:1970:ALM

[984] John D. Kalbfleisch and D. A. Sprott. Application of likelihood methods to models involving large numbers of parameters. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):175–208, ???? 1970. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984524.

See corrigendum [1060].

Fuller:1970:SRS

[985] Wayne A. Fuller. Sampling with random stratum boundaries. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):209–226, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984525.

Gastwirth:1970:ARE

[986] Joseph L. Gastwirth. On asymptotic relative efficiencies of a class of rank tests. Journal of the Royal Statistical Society. Series B (Methodological), 32 (2):227–232, ???? 1970. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984526.

Ray:1970:SCF

[987] W. D. Ray. A short cut in the fitting of autoregressive series. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):233–235, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984527.

Easterling:1970:ACB

[988] Robert G. Easterling and David L. Weeks. An accuracy criterion for Bayesian tolerance intervals. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):236–240, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984528.

Kemp:1970:CSB

[989] Kenneth W. Kemp. Convergent sequences of bounds for some parameters of a simple sequential test. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):241–253, ???? 1970. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984529.

Mardia:1970:SPF

[990] K. V. Mardia. Some problems of fitting for contingency-type bivariate distributions. Journal of the Royal Statistical Society. Series B (Methodological), 32 (2):254–264, ???? 1970. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984530.

Arnold:1970:ADR

[991] Barry C. Arnold. An alternative derivation of a result due to Srivastava and bancroft. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):265–267, ???? 1970. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984531.

Hasofer:1970:RIP

[992] A. M. Hasofer. On the representation of ignorance in Poisson processes. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):268–271, ???? 1970. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984532.

Tong:1970:MSI

[993] Yung Liang Tong. Multi-stage interval estimations of the largest mean of k normal populations. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):272–277, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984533.

Chambers:1970:SPS

[994] M. L. Chambers. A simple problem with strikingly different frequentist and Bayesian solutions. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):278–282, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984534.

Andersen:1970:APC

[995] Erling Bernhard Andersen. Asymptotic properties of conditional maximumlikelihood estimators. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):283–301, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2984535. See corrigendum [1023].

Brewer:1970:DNM

[996] K. R. W. Brewer and M. Hanif. Durbin’s new multistage variance estimator. Journal of the Royal Statistical Society. Series B (Methodological), 32 (2):302–311, ???? 1970. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984536.

Rao:1970:FNS

[997] T. Subba Rao. The fitting of non-stationary time-series models with time-dependent parameters. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):312–322, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984537.

Anonymous:1970:FMb

[998] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 32(2):??, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984522.

Anonymous:1970:BMc

[999] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 32(3):i–iii, ???? 1970. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984369.

Atkinson:1970:MDB

[1000] A. C. Atkinson. A method for discriminating between models. Journal of the Royal Statistical Society. Series B (Methodological), 32(3):323–353, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984360.

Prabhu:1970:RPO

[1001] N. U. Prabhu and Michael Rubinovitch. On a regenerative phenomenon occurring in a storage model. Journal of the Royal Statistical Society. Series B (Methodological), 32(3): 354–361, ???? 1970. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984361.

Sclove:1970:EPL

[1002] Stanley L. Sclove and John Van Ryzin. Estimating the parameters of a linear function of a random variable. Journal of the Royal Statistical Society. Series B (Methodological), 32(3):362–368, ???? 1970. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984362.

Bloomfield:1970:SAR

[1003] P. Bloomfield. Spectral analysis with randomly missing observations. Journal of the Royal Statistical Society. Series B (Methodological), 32(3):369–380, ???? 1970. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984363.

Gurland:1970:FCD

[1004] John Gurland and Roy Milton. Further consideration of the distribution of the multiple correlation coefficient. Journal of the Royal Statistical Society. Series B (Methodological), 32(3):381–394, ???? 1970. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984364.

Altham:1970:MAC

[1005] Patricia M. E. Altham. The measurement of association in a contingency table: Three extensions of the crossratios and metrics methods. Journal of the Royal Statistical Society. Series B (Methodological), 32(3):395–407, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984365.

Downton:1970:BED

[1006] F. Downton. Bivariate exponential distributions in reliability theory. Journal of the Royal Statistical Society. Series B (Methodological), 32(3):408–417, ???? 1970. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984366.

Waugh:1970:TBP

[1007] W. A. O’N. Waugh. Transformation of a birth process into a Poisson process. Journal of the Royal Statistical Society. Series B (Methodological), 32 (3):418–431, ???? 1970. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984367.

Sprent:1970:SPL

[1008] P. Sprent. The saddle point of the likelihood surface for a linear functional relationship. Journal of the Royal Statistical Society. Series B (Methodological), 32(3):432–434, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984368.

Anonymous:1970:FMc

[1009] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 32(3):??, ???? 1970. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984359.

Boneva:1971:STT

[1010] Liliana I. Boneva, David Kendall, and Ivan Stefanov. Spline transformations: Three new diagnostic aids for the statistical data- analyst. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):1–71, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2986005.

Bar-Shalom:1971:APM

[1011] Yaakov Bar-Shalom. On the asymptotic properties of the maximumlikelihood estimate obtained from dependent observations. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):72–77, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2986006.

Davis:1971:CSB

[1012] Miles Davis. Comparison of sequential bioassays in small samples. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):78–87, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2986007.

Dodson:1971:SVT

[1013] C. T. J. Dodson. Spatial variability and the theory of sampling in random fibrous networks. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):88–94, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2986008.

Finucan:1971:PPN

[1014] H. M. Finucan. Posterior precision for non-normal distributions. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):95–97, ???? 1971. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2986009. See corrigenda

[1045, 1116].

Hartigan:1971:EAR

[1015] J. A. Hartigan. Error analysis by replaced samples. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):98–110, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2986010.

Kshirsagar:1971:GFD

[1016] A. M. Kshirsagar. Goodness of fit of a discriminant function from the vector space of dummy variables. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):111–116, ???? 1971. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2986011.

Lee:1971:SRS

[1017] Yoong-Sin Lee. Some results on the sampling distribution of the multiple correlation coefficient. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):117–130, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2986012.

Pearce:1971:BDM

[1018] S. C. Pearce and J. N. R. Jeffers. Block designs and missing data. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):131–136, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2986013.

Peritz:1971:SRA

[1019] E. Peritz. On a statistic for rank analysis of variance. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):137–139, ????

1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2986014.

Pierce:1971:DRA

[1020] David A. Pierce. Distribution of residual autocorrelations in the regression model with autoregressivemoving average errors. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):140–146, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2986015.

Tan:1971:DWV

[1021] W. Y. Tan and Irwin Guttman. A disguised Wishart variable and a related theorem. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):147–152, ???? 1971. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2986016.

Wahba:1971:STI

[1022] Grace Wahba. Some tests of independence for stationary multivariate time series. Journal of the Royal Statistical Society. Series B (Methodological), 33 (1):153–166, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2986017.

Andersen:1971:CAP

[1023] E. B. Andersen. Corrigendum: Asymptotic properties of conditional maximum-likelihood estimators. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):167, ???? 1971. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2986018. See [995].

Anonymous:1971:BMa

[1024] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):169–iv, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2986019.

Anonymous:1971:FMa

[1025] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 33(1):??, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2986004.

Anonymous:1971:VI

[1026] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Methodological), 33(1): ??, ???? 1971. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2986003.

Anonymous:1971:BMb

[1027] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 33(2):i–v, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985014.

Box:1971:BNE

[1028] M. J. Box. Bias in nonlinear estimation. Journal of the Royal Statistical Society. Series B (Methodological), 33 (2):171–201, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985002.

Wynn:1971:SCB

[1029] Henry P. Wynn and P. Bloomfield.

Simultaneous confidence bands in regression analysis. Journal of the Royal Statistical Society. Series B (Methodological), 33(2):202–217, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985003.

ONeill:1971:PSM

[1030] R. O’Neill and G. B. Wetherill. The present state of multiple comparison methods. Journal of the Royal Statistical Society. Series B (Methodological), 33(2):218–250, ???? 1971. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985004.

Cox:1971:CBA

[1031] D. R. Cox. The choice between alternative ancillary statistics. Journal of the Royal Statistical Society. Series B (Methodological), 33(2):251–255, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985005.

Hudson:1971:IEL

[1032] D. J. Hudson. Interval estimation from the likelihood function. Journal of the Royal Statistical Society. Series B (Methodological), 33(2):256–262, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985006.

Dodds:1971:SFS

[1033] D. J. Dodds and J. G. Fryer. Some families of selection probabilities for sampling with probability proportional to size. Journal of the Royal Statistical Society. Series B (Methodological), 33 (2):263–274, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985007.

Hoem:1971:PEF

[1034] Jan M. Hoem. Point estimation of forces of transition in demographic models. Journal of the Royal Statistical Society. Series B (Methodological), 33 (2):275–289, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985008.

Westcott:1971:EMR

[1035] M. Westcott. On existence and mixing results for cluster point processes. Journal of the Royal Statistical Society. Series B (Methodological), 33(2): 290–300, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985009.

John:1971:TEB

[1036] S. John. A test of equality of block-diagonal covariance matrices and its role of unification. Journal of the Royal Statistical Society. Series B (Methodological), 33(2):301–306, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985010.

Schlesselman:1971:PFN

[1037] J. Schlesselman. Power families: A note on the Box and Cox transformation. Journal of the Royal Statistical Society. Series B (Methodological), 33 (2):307–311, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985011.

Foster:1971:EDD

[1038] F. G. Foster, J. V. Rosenhead, and V. Siskind. The effect of the demand distribution in inventory models combining holding stockout and re-order costs. Journal of the Royal Statistical Society. Series B (Methodological), 33 (2):312–325, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985012.

MacDonald:1971:CEP

[1039] P. D. M. MacDonald. Comment on “An Estimation Procedure for Mixtures of Distributions” by Choi and Bulgren. Journal of the Royal Statistical Society. Series B (Methodological), 33(2): 326–329, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985013. See

[892].

Anonymous:1971:FMb

[1040] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 33(2):??, ???? 1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985001.

Anonymous:1971:BMc

[1041] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 33(3):i–iii, ???? 1971. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984688.

Kemp:1971:FEW

[1042] K. W. Kemp. Formal expressions which can be applied to Cusum charts. Journal of the Royal Statistical Society. Series B (Methodological), 33(3):331–360, ???? 1971. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984679.

Godambe:1971:BFF

[1043] V. P. Godambe and Mary E. Thompson. Bayes, fiducial and frequency aspects of statistical inference in regression analysis in survey-sampling. Journal of the Royal Statistical Society. Series B (Methodological), 33(3):361–390, ???? 1971. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984680.

Foreman:1971:EUS

[1044] E. K. Foreman and K. R. W. Brewer.

The efficient use of supplementary information in standard sampling procedures. Journal of the Royal Statistical Society. Series B (Methodological), 33 (3):391–400, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984681. See

comments [1213].

Finucan:1971:CPP

[1045] H. M. Finucan. Corrigendum: Posterior precision for non-normal distributions. Journal of the Royal Statistical Society. Series B (Methodological), 33(3):400, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984682. See

[1014].

King:1971:CSD

[1046] Ralph A. King. The covariance structure of the departure process from M/G/1 queues with finite waiting lines. Journal of the Royal Statistical Society. Series B (Methodological), 33(3): 401–405, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984683.

Daley:1971:WSP

[1047] D. J. Daley. Weakly stationary point processes and random measures. Journal of the Royal Statistical Society. Series B (Methodological), 33(3):406–428, ???? 1971. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984684.

Bohrer:1971:SSR

[1048] Robert Bohrer. Some specific results in sequential design. Journal of the Royal Statistical Society. Series B (Methodological), 33(3):429–437, ???? 1971. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984685.

Hawkes:1971:PSS

[1049] Alan G. Hawkes. Point spectra of some mutually exciting point processes. Journal of the Royal Statistical Society. Series B (Methodological), 33(3): 438–443, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984686.

Shenton:1971:SMM

[1050] L. R. Shenton, K. O. Bowman, and D. Sheehan. Sampling moments of moments associated with univariate distributions. Journal of the Royal Statistical Society. Series B (Methodological), 33 (3):444–457, ???? 1971. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984687. See

[1103].

Anonymous:1971:FMc

[1051] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 33(3):??, ????

1971. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984678.

Anonymous:1972:BMa

[1052] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 34(1):i–iii, ???? 1972. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985058.

Lindley:1972:BEL

[1053] D. V. Lindley and A. F. M. Smith.

Bayes estimates for the linear model. Journal of the Royal Statistical Society. Series B (Methodological), 34(1):1–41, ???? 1972. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985048.

Andersen:1972:NSS

[1054] E. B. Andersen. The numerical solution of a set of conditional estimation equations. Journal of the Royal Statistical Society. Series B (Methodological), 34(1):42–54, ???? 1972. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985049.

Hald:1972:APB

[1055] A. Hald and N. Keiding. Asymptotic properties of Bayesian decision rules for two terminal decisions and multiple sampling. II. Journal of the Royal Statistical Society. Series B (Methodological), 34(1):55–74, ???? 1972. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985050.

Besag:1972:NNS

[1056] J. E. Besag. Nearest-neighbour systems and the auto-logistic model for binary data. Journal of the Royal Statistical Society. Series B (Methodological), 34 (1):75–83, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985051.

Freeman:1972:EDM

[1057] G. H. Freeman. Experimental designs with many classifications. Journal of the Royal Statistical Society. Series B (Methodological), 34(1):84–101, ???? 1972. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985052.

Mardia:1972:MSU

[1058] K. V. Mardia. A multi-sample uniform scores test on a circle and its parametric competitor. Journal of the Royal Statistical Society. Series B (Methodological), 34(1):102–113, ???? 1972. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985053.

Lawrence:1972:AEI

[1059] A. J. Lawrence. Arbitrary event initial conditions for branching Poisson processes. Journal of the Royal Statistical Society. Series B (Methodological), 34 (1):114–123, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985054.

Kalbfleisch:1972:CAL

[1060] J. D. Kalbfleisch and D. A. Sprott. Corrigendum: Application of likelihood methods to models involving large numbers of parameters. Journal of the Royal Statistical Society. Series B (Methodological), 34(1):124–125, ???? 1972. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985055. See [984].

Fryer:1972:NBM

[1061] M. J. Fryer. A note on Barnard’s modification to Wald’s sequential test for double dichotomies. Journal of the Royal Statistical Society. Series B (Methodological), 34(1):126–129, ???? 1972. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985056.

Hawkes:1972:BED

[1062] Alan G. Hawkes. A bivariate exponential distribution with applications to reliability. Journal of the Royal Statistical Society. Series B (Methodological), 34(1):129–131, ???? 1972. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985057.

Anonymous:1972:FMa

[1063] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 34(1):??, ???? 1972. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985047.

Anonymous:1972:VI

[1064] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Methodological), 34(1): ??, ???? 1972. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985046.

Wynn:1972:RTC

[1065] Henry P. Wynn. Results in the theory and construction of D-optimum experimental designs. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):133–147, ???? 1972. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985179.

Laycock:1972:CLA

[1066] P. J. Laycock. Convex loss applied to design in regression problems. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):148–186, ???? 1972. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985180.

Cox:1972:RML

[1067] D. R. Cox. Regression models and lifetables. Journal of the Royal Statistical Society. Series B (Methodological), 34 (2):187–220, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2985181.

Ghosh:1972:LTH

[1068] B. K. Ghosh. On Lehmann’s test for homogeneity of variances. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):221–234, ???? 1972. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985182.

Rao:1972:TTD

[1069] T. Subba Rao and H. Tong. A test for time-dependence of linear open-loop systems. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):235–250, ???? 1972. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2985183.

Afonja:1972:MMC

[1070] Biyi Afonja. The moments of the maximum of correlated normal and tvariates. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):251–262, ???? 1972. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985184.

Matthews:1972:CPD

[1071] Jane P. Matthews. A combinatorial proof of the distribution of the time to first emptiness of an infinite dam with Markovian inputs. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):263–267, ???? 1972. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985185.

Spjotvoll:1972:ULR

[1072] Emil Spjotvoll. Unbiasedness of likelihood ratio confidence sets in cases without nuisance parameters. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):268–273, ???? 1972. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985186.

Copas:1972:LSL

[1073] J. B. Copas. The likelihood surface in the linear functional relationship problem. Journal of the Royal Statistical Society. Series B (Methodological), 34 (2):274–278, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985187.

Cassel:1972:MSR

[1074] Claes-Magnus Cassel and Carl-Erik Sarndal. A model for studying robustness of estimators and informativeness of labels in sampling with varying probabilities. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):279–289, ???? 1972. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2985188.

Durbin:1972:CCM

[1075] J. Durbin and M. Knott. Components of Cramér–von Mises statistics. I. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):290–307, ???? 1972. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985189.

Vaughan:1972:PEO

[1076] R. J. Vaughan and W. N. Venables. Permanent expressions for order statistic densities. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):308–310, ???? 1972. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985190.

Anonymous:1972:BMb

[1077] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):??, ???? 1972. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985191.

Anonymous:1972:FMb

[1078] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 34(2):??, ???? 1972. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985178.

Anonymous:1972:BMc

[1079] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 34(3):i–ii, ????

1972. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985087.

Sibson:1972:OIM

[1080] Robin Sibson. Order invariant methods for data analysis. Journal of the Royal Statistical Society. Series B (Methodological), 34(3):311–349, ???? 1972. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985070.

Fox:1972:OTS

[1081] A. J. Fox. Outliers in time series. Journal of the Royal Statistical Society. Series B (Methodological), 34(3):350–363, ???? 1972. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985071.

Deshpande:1972:LOR

[1082] Jayant V. Deshpande. Linear ordered rank tests which are asymptotically efficient for the two- sample problem. Journal of the Royal Statistical Society. Series B (Methodological), 34(3): 364–370, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985072.

Gokhale:1972:ALL

[1083] D. V. Gokhale. Analysis of log-linear models. Journal of the Royal Statistical Society. Series B (Methodological), 34 (3):371–376, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985073.

Anderson:1972:RID

[1084] D. A. Anderson and J. N. Srivastava.

Resolution IV designs of the 2^{m} × 3 series. Journal of the Royal Statistical Society. Series B (Methodological), 34 (3):377–384, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985074.

Priestley:1972:NPF

[1085] M. B. Priestley and M. T. Chao. Nonparametric function fitting. Journal of the Royal Statistical Society. Series B (Methodological), 34(3):385–392, ???? 1972. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985075.

Dolby:1972:GLS

[1086] G. R. Dolby. Generalized least squares and maximum likelihood estimation of non-linear functional relationships. Journal of the Royal Statistical Society. Series B (Methodological), 34(3): 393–400, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985076.

John:1972:TFE

[1087] J. A. John and T. M. F. Smith. Twofactor experiments in non-orthogonal designs. Journal of the Royal Statistical Society. Series B (Methodological), 34 (3):401–409, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985077.

Swartz:1972:MRE

[1088] G. B. Swartz and V. K. Murthy. Minimum risk estimators with applications. Journal of the Royal Statistical Society. Series B (Methodological), 34(3): 410–416, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985078.

Billard:1972:PST

[1089] L. Billard. Properties of some two-sided sequential tests for the normal distribution. Journal of the Royal Statistical Society. Series B (Methodological), 34 (3):417–426, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985079.

El-Sayyad:1972:CMS

[1090] G. M. El-Sayyad. A comparison of methods of sampling from a Poisson process. Journal of the Royal Statistical Society. Series B (Methodological), 34 (3):427–430, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985080.

Margolin:1972:NOM

[1091] Barry H. Margolin. Non-orthogonal main-effect designs for asymmetrical factorial experiments. Journal of the Royal Statistical Society. Series B (Methodological), 34(3):431–440, ???? 1972. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985081.

Gart:1972:NRA

[1092] John J. Gart and Donald G. Thomas.

Numerical results on approximate confidence limits for the odds ratio. Journal of the Royal Statistical Society. Series B (Methodological), 34(3):441–447, ???? 1972. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985082.

Afifi:1972:CST

[1093] A. A. Afifi and P. J. Kim. Comparison of some two-sample location tests for non-normal alternatives. Journal of the Royal Statistical Society. Series B (Methodological), 34(3):448–455, ???? 1972. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985083.

Weiss:1972:AES

[1094] L. Weiss and J. Wolfowitz. An asymptotically efficient, sequential equivalent of the t-test. Journal of the Royal Statistical Society. Series B (Methodological), 34(3):456–460, ???? 1972. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985084.

Curnow:1972:NVW

[1095] R. N. Curnow. The number of variables when searching for an optimum. Journal of the Royal Statistical Society. Series B (Methodological), 34(3):461–476, ???? 1972. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985085. See comments

[1096] and corrections [1112, 1255].

Brooks:1972:RCN

[1096] R. J. Brooks, D. J. Finney, S. C. Pearce, F. Yates, and R. N. Curnow. On R. N. Curnow’s “The Number of Variables when Searching for an Optimum”. Journal of the Royal Statistical Society. Series B (Methodological), 34 (3):477–481, ???? 1972. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985086. See

[1095, 1112].

Anonymous:1972:FMc

[1097] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 34(3):??, ???? 1972. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985069.

Anonymous:1973:BMa

[1098] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):i–iv, ???? 1973. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985141.

McNeil:1973:CLA

[1099] Donald R. McNeil and Siegfried Schach. Central limit analogues for Markov population processes. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):1–23, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985121.

Korin:1973:SAD

[1100] B. P. Korin and E. H. Stevens. Some approximations for the distribution of a multivariate likelihood ratio criterion. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):24–27, ???? 1973. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985122.

Kupper:1973:NAR

[1101] Lawrence L. Kupper. A note on the admissibility of a response surface design. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):28–32, ???? 1973. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985123.

Berenblut:1973:NTA

[1102] I. I. Berenblut and G. I. Webb. A new test for autocorrelated errors in the linear regression model. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):33–50, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985124.

Shenton:1973:CSM

[1103] L. R. Shenton, K. O. Bowman, and D. Sheehan. Corrigendum: Sampling moments of moments associated with univariate distributions. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):50, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985125. See [1050].

Nigam:1973:MME

[1104] A. K. Nigam. Multifactor mixture experiments. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):51–56, ???? 1973. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985126.

Scott:1973:SDS

[1105] Alastair Scott and T. M. F. Smith. Survey design, symmetry and posterior distributions. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):57–60, ???? 1973. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985127.

Blight:1973:SMR

[1106] B. J. N. Blight and A. J. Scott. A stochastic model for repeated surveys. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):61–66, ???? 1973. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985128.

Smith:1973:GBL

[1107] A. F. M. Smith. A general Bayesian linear model. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):67–75, ???? 1973. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985129.

Wilson:1973:EPM

[1108] G. Tunnicliffe Wilson. The estimation of parameters in multivariate time series models. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):76–85, ???? 1973. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985130.

Genizi:1973:SPC

[1109] Abraham Genizi. Some power comparisons between bivariate tests and a twofold univariate test. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):86–96, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985131.

Srivastava:1973:PSP

[1110] M. S. Srivastava. The performance of a sequential procedure for a slippage problem. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):97–103, ???? 1973. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985132.

Mercer:1973:QSA

[1111] A. Mercer. Queues with scheduled arrivals: A correction, simplification and extension. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):104–116, ????

1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985133. See [477].

Curnow:1973:CCR

[1112] R. N. Curnow. Corrigendum: Comment by R. J. Brooks on “The Number of Variables When Searching for an Optimum”. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):116, ???? 1973. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985134. See

[1095].

Lanke:1973:ARM

[1113] Jan Lanke. Asymptotic results on matching distributions. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):117–122, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985135.

Whittle:1973:SGP

[1114] P. Whittle. Some general points in the theory of optimal experimental design. Journal of the Royal Statistical Society. Series B (Methodological), 35(1): 123–130, ???? 1973. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985136.

Subrahmaniam:1973:EPB

[1115] Kocherlakota Subrahmaniam and Kathleen Subrahmaniam. On the estimation of the parameters in the bivariate negative binomial distribution. Journal of the Royal Statistical Society. Series B (Methodological), 35(1): 131–146, ???? 1973. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985137.

Finucan:1973:CPP

[1116] H. M. Finucan. Corrigenda: Posterior precision for non-normal distributions. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):146, ???? 1973. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985139. See

[1014].

Raktoe:1973:CUA

[1117] B. L. Raktoe and W. T. Federer.

Corrigenda: A unified approach for constructing a useful class of nonorthogonal main effect plans in k^{n} factorials. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):146, ???? 1973. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985138. See

[886].

Bhargava:1973:TCI

[1118] R. P. Bhargava and M. S. Srivastava.

On Tukey’s confidence intervals for the contrasts in the means of the intraclass correlation model. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):147–152, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985140.

Anonymous:1973:FMa

[1119] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 35(1):??, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985120.

Anonymous:1973:VI

[1120] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 35(1): ??, ???? 1973. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985119.

Priestley:1973:ABN

[1121] M. B. Priestley and H. Tong. On the analysis of bivariate non-stationary processes. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):153–166, ???? 1973. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984905.

Hammond:1973:ESR

[1122] J. K. Hammond. Evolutionary spectra in random vibrations. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):167–188, ???? 1973. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984906.

Dawid:1973:MPB

[1123] A. P. Dawid, M. Stone, and J. V. Zidek. Marginalization paradoxes in Bayesian and structural inference. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):189–233, ???? 1973. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984907.

Gabriel:1973:SCI

[1124] K. R. Gabriel, J. Putter, and Y. Wax. Simultaneous confidence intervals for product-type interaction contrasts. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):

234–244, ???? 1973. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984908.

Paik:1973:PTI

[1125] U. B. Paik and W. T. Federer. On PA-Type incomplete block designs and PAB-Type rectangular designs. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):245–251, ???? 1973. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984909.

Hannan:1973:LRU

[1126] E. J. Hannan and P. M. Robinson.

Lagged regression with unknown lags. Journal of the Royal Statistical Society. Series B (Methodological), 35(2): 252–267, ???? 1973. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984910.

Draper:1973:SDE

[1127] Norman R. Draper and Agnes M.

Herzberg. Some designs for extrapolation outside a sphere. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):268–276, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984911.

Bofinger:1973:GFT

[1128] Eve Bofinger. Goodness-of-fit test using sample quantiles. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):277–284, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984912.

Dickey:1973:SRP

[1129] James Dickey. Scientific reporting and personal probabilities: Student’s hypothesis. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):285–305, ???? 1973. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984913.

Lawrence:1973:DIB

[1130] A. J. Lawrence. Dependency of intervals between events in superposition processes. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):306–315, ???? 1973. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984914.

Finucan:1973:RPT

[1131] H. M. Finucan. The relative position of the two regression curves. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):316–322, ???? 1973. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984915.

Newbold:1973:BEB

[1132] Paul Newbold. Bayesian estimation of Box–Jenkins transfer function-noise models. Journal of the Royal Statistical Society. Series B (Methodological), 35 (2):323–336, ???? 1973. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984916.

Hinkley:1973:TST

[1133] D. V. Hinkley. Two-sample tests with unordered pairs. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):337–346, ????

1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984917.

Halpern:1973:BSR

[1134] Elkan F. Halpern. Bayesian spline regression when the number of knots is unknown. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):347–360, ???? 1973. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984918.

Cotter:1973:MFE

[1135] Sarah C. Cotter, J. A. John, and T. M. F. Smith. Multi-factor experiments in non-orthogonal designs. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):361–367, ???? 1973. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984919.

Robinson:1973:ACU

[1136] J. Robinson. The analysis of covariance under a randomization model. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):368–376, ???? 1973. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984920.

Anonymous:1973:BMb

[1137] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):377–ii, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984921.

Anonymous:1973:FMb

[1138] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 35(2):??, ????

1973. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984904.

Anonymous:1973:BMc

[1139] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 35(3):i–ii, ???? 1973. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985118.

Efron:1973:CPR

[1140] B. Efron and C. Morris. Combining possibly related estimation problems. Journal of the Royal Statistical Society. Series B (Methodological), 35(3): 379–421, ???? 1973. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2985106.

Mardia:1973:MTM

[1141] K. V. Mardia and B. D. Spurr. Multisample tests for multimodal and axial circular populations. Journal of the Royal Statistical Society. Series B (Methodological), 35(3):422–436, ???? 1973. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985107.

Streit:1973:MVF

[1142] F. Streit. Mean-value formulae for a class of random sets. Journal of the Royal Statistical Society. Series B (Methodological), 35(3):437–444, ???? 1973. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985108.

El-Sayyad:1973:BCA

[1143] G. M. El-Sayyad. Bayesian and classical analysis of Poisson regression. Journal of the Royal Statistical Society. Series B (Methodological), 35(3):445–451, ???? 1973. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985109.

Dahiya:1973:TFB

[1144] Ram C. Dahiya and John Gurland. A test of fit for bivariate distributions. Journal of the Royal Statistical Society. Series B (Methodological), 35(3): 452–465, ???? 1973. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985110.

Saxena:1973:SSB

[1145] S. K. Saxena and A. K. Nigam.

Symmetric-simplex block designs for mixtures. Journal of the Royal Statistical Society. Series B (Methodological), 35(3):466–472, ???? 1973. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985111.

Atkinson:1973:TTN

[1146] A. C. Atkinson. Testing transformations to normality. Journal of the Royal Statistical Society. Series B (Methodological), 35(3):473–479, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985112.

Ghosh:1973:SMT

[1147] B. K. Ghosh. Some monotonicity theorems for χ^{2}, F and t distributions with applications. Journal of the Royal Statistical Society. Series B (Methodological), 35(3):480–492, ???? 1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985113.

Kupper:1973:MDF

[1148] Lawrence L. Kupper. Minimax designs for Fourier series and spherical harmonics regressions: A characterization of rotatable arrangements. Journal of the Royal Statistical Society. Series B (Methodological), 35(3):493–500, ???? 1973. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985114.

Shanbhag:1973:FED

[1149] D. N. Shanbhag. On the first emptiness of dams with Markovian inputs. Journal of the Royal Statistical Society. Series B (Methodological), 35(3):501–506, ???? 1973. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985115.

Csorgo:1973:SET

[1150] M. Csorgo, V. Seshadri, and M. Yalovsky. Some exact tests for normality in the presence of unknown parameters. Journal of the Royal Statistical Society. Series B (Methodological), 35(3):507–522, ???? 1973. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985116.

Lehman:1973:TMC

[1151] Eugène H. Lehman. On two modifications of the Cramér–von Mises statistic. Journal of the Royal Statistical Society. Series B (Methodological), 35(3): 523, ???? 1973. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

                                        jstor.org/stable/2985117.

Anonymous:1973:FMc

[1152] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 35(3):??, ????

1973. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985105.

Anonymous:1974:BMa

[1153] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 36(1):i, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984779.

Kish:1974:ICS

[1154] Leslie Kish and Martin Richard Frankel. Inference from complex samples. Journal of the Royal Statistical Society. Series B (Methodological), 36 (1):1–37, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984767.

Lindsey:1974:CPD

[1155] J. K. Lindsey. Comparison of probability distributions. Journal of the Royal Statistical Society. Series B (Methodological), 36(1):38–47, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984768.

Bhat:1974:MML

[1156] B. R. Bhat. On the method of maximum-likelihood for dependent observations. Journal of the Royal Statistical Society. Series B (Methodological), 36(1):48–53, ???? 1974. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984769.

Rahman:1974:EFD

[1157] M. Rahman and A. K. Md. Ehsanes Saleh. Explicit form of the distribution of the Behrens–Fisher d-statistic.

Journal of the Royal Statistical Society. Series B (Methodological), 36(1): 54–60, ???? 1974. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

jstor.org/stable/2984770. See cor-

rigendum [1193].

Bhansali:1974:APW

[1158] R. J. Bhansali. Asymptotic properties of the Wiener–Kolmogorov predictor. I. Journal of the Royal Statistical Society. Series B (Methodological), 36(1):61–73, ???? 1974. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984771. See corrections

[1487].

Sargan:1974:SDA

[1159] J. D. Sargan. Some discrete approximations to continuous time stochastic models. Journal of the Royal Statistical Society. Series B (Methodological), 36(1):74–90, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984772.

Dent:1974:MVW

[1160] Warren T. Dent and James D. Broffitt. The mean and variance of Watson’s distribution. Journal of the Royal Statistical Society. Series B (Methodological), 36(1):91–98, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984773.

Andrews:1974:SMN

[1161] D. F. Andrews and C. L. Mallows. Scale mixtures of normal distributions. Journal of the Royal Statistical Society. Series B (Methodological), 36(1): 99–102, ???? 1974. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

jstor.org/stable/2984774.

Theobald:1974:GMS

[1162] C. M. Theobald. Generalizations of mean square error applied to ridge regression. Journal of the Royal Statistical Society. Series B (Methodological), 36(1):103–106, ???? 1974. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984775. See corrections

[1488].

John:1974:SSN

[1163] J. A. John and T. M. F. Smith.

Sum of squares in non-full rank general linear hypotheses. Journal of the Royal Statistical Society. Series B (Methodological), 36(1):107–108, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984776.

Wedderburn:1974:ADS

[1164] R. W. M. Wedderburn. Alternative derivation of the sum of squares in a non-full rank general linear hypothesis. Journal of the Royal Statistical Society. Series B (Methodological), 36(1): 108–109, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984777.

Good:1974:CBS

[1165] I. J. Good. Corrigendum: A Bayesian significance test for multinomial distributions. Journal of the Royal Statistical Society. Series B (Methodological), 36(1):109, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984778. See

Anonymous:1974:FMa

[1166] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 36(1):??, ???? 1974. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984766.

Anonymous:1974:VI

[1167] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 36(1): ??, ???? 1974. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984765.

Anonymous:1974:BMb

[1168] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 36(2):i–ii, ???? 1974. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984822.

Stone:1974:CVC

[1169] M. Stone. Cross-validatory choice and assessment of statistical predictions. Journal of the Royal Statistical Society. Series B (Methodological), 36(2): 111–147, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984809. See

corrigenda [1256].

Tversky:1974:AU

[1170] Amos Tversky. Assessing uncertainty. Journal of the Royal Statistical Society. Series B (Methodological), 36(2): 148–159, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984810.

Suppes:1974:MB

[1171] Patrick Suppes. The measurement of belief. Journal of the Royal Statistical Society. Series B (Methodological), 36 (2):160–191, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984811.

Besag:1974:SIS

[1172] Julian Besag. Spatial interaction and the statistical analysis of lattice systems. Journal of the Royal Statistical Society. Series B (Methodological), 36 (2):192–236, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984812.

Joshi:1974:NIO

[1173] V. M. Joshi. A note on the incompatibility of one-sided Bayes and frequency confidence intervals. Journal of the Royal Statistical Society. Series B (Methodological), 36(2):237–242, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984813.

Naik:1974:SPT

[1174] Umesh D. Naik. On some procedures for testing a linear relation among variances. Journal of the Royal Statistical Society. Series B (Methodological), 36 (2):243–257, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984814.

Narain:1974:CDE

[1175] P. Narain. The conditioned diffusion equation and its use in population genetics. Journal of the Royal Statistical Society. Series B (Methodological), 36 (2):258–266, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984815.

Cotter:1974:GMC

[1176] Sarah C. Cotter. A general method of confounding for symmetrical factorial experiments. Journal of the Royal Statistical Society. Series B (Methodological), 36(2):267–276, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984816.

Alam:1974:IER

[1177] K. Alam and K. M. Lal Saxena. On interval estimation of a ranked parameter. Journal of the Royal Statistical Society. Series B (Methodological), 36 (2):277–283, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984817.

Goldstein:1974:RTE

[1178] M. Goldstein and A. F. M. Smith. Ridge-type estimators for regression analysis. Journal of the Royal Statistical Society. Series B (Methodological), 36(2):284–291, ???? 1974. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984818.

Ramachandran:1974:ASR

[1179] Geetha Ramachandran and T. J. Rao. Allocation to strata and relative efficiencies of stratified and unstratified πP S sampling schemes. Journal of the Royal Statistical Society. Series B (Methodological), 36(2):292–298, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984819.

Pegram:1974:EDM

[1180] G. G. S. Pegram. Expectation and dispersion matrices for the estimation of distributions associated with discrete reservoirs. Journal of the Royal Statistical Society. Series B (Methodological), 36(2):299–304, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984820.

Thyregod:1974:BSS

[1181] P. Thyregod. Bayesian single sampling acceptance plans for finite lot sizes. Journal of the Royal Statistical Society. Series B (Methodological), 36(2): 305–319, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984821.

Anonymous:1974:FMb

[1182] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 36(2):??, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984808.

Atkinson:1974:PED

[1183] A. C. Atkinson and D. R. Cox.

Planning experiments for discriminating between models. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):321–348, ???? 1974. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984923. See correction

[1228].

Matthews:1974:BNS

[1184] G. V. T. Matthews. On bird navigation, with some statistical undertones.

Journal of the Royal Statistical Society. Series B (Methodological), 36(3): 349–364, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984924.

Kendall:1974:PSB

[1185] David G. Kendall. Pole-seeking Brownian motion and bird navigation. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):365–417, ???? 1974. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984925.

Lindsey:1974:CCS

[1186] J. K. Lindsey. Construction and comparison of statistical models. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):418–425, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984926.

Marcus:1974:EIH

[1187] Allan H. Marcus. Environmental impacts of highway traffic as two-sided filtered Markov renewal processes. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):426–429, ???? 1974. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984927.

Knott:1974:DCM

[1188] M. Knott. The distribution of the Cramér–von Mises statistic for small sample sizes. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):430–438, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984928.

Raja:1974:CFE

[1189] J. Sundara Raja. Confounding in factorial experiments. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):439–441, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984929.

Rao:1974:PGI

[1190] C. Radhakrishna Rao. Projectors, generalized inverses and the BLUE’s. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):442–448, ???? 1974. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984930.

Wedderburn:1974:GLM

[1191] R. W. M. Wedderburn. Generalized linear models specified in terms of constraints. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):449–454, ???? 1974. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984931.

Brown:1974:RST

[1192] B. M. Brown. A restricted sequential test. Journal of the Royal Statistical Society. Series B (Methodological), 36 (3):455–465, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984932.

Rahman:1974:CEF

[1193] M. Rahman and A. K. M. E. Saleh.

Corrigendum: Explicit form of the distribution of the Behrens–Fisher d statistic. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):466, ???? 1974. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984933. See

[1157].

Darroch:1974:INC

[1194] J. N. Darroch and I. R. James. F -independence and null correlation of continuous, bounded-sum, positive variables. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):467–483, ???? 1974. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984934.

Anonymous:1974:BMc

[1195] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):??, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984935.

Anonymous:1974:FMc

[1196] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 36(3):??, ???? 1974. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984922.

Anonymous:1975:BMa

[1197] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):i–ii, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985000.

Anonymous:1975:VI

[1198] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):i– iv, ???? 1975. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984984.

Anonymous:1975:FMa

[1199] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):vii–viii, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984985.

Kingman:1975:RDD

[1200] J. F. C. Kingman. Random discrete distributions. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):1–22, ???? 1975. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984986.

Leonard:1975:BEM

[1201] Tom Leonard. Bayesian estimation methods for two-way contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 37 (1):23–37, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984987.

Broffitt:1975:AAD

[1202] J. D. Broffitt and W. T. Dent. Asymptotic approximations for distributions of ratios of linear functions of random variables. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):38–48, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984988.

Morton:1975:EFT

[1203] R. Morton. On the efficiency of Fisher’s tea-tasting designs. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):49–53, ????

1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984989.

Cotter:1975:PCS

[1204] Sarah C. Cotter. Partial confounding in symmetrical factorial experiments. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):54–62, ???? 1975. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984990.

John:1975:SRF

[1205] J. A. John and A. M. Dean. Single replicate factorial experiments in generalized cyclic designs: I. symmetrical arrangements. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):63–71, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984991.

Dean:1975:SRF

[1206] A. M. Dean and J. A. John. Single replicate factorial experiments in generalized cyclic designs: II. Asymmetrical arrangements. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):72–76, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984992.

Borth:1975:TEC

[1207] David M. Borth. A total entropy criterion for the dual problem of model discrimination and parameter estimation. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):77–87, ???? 1975. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984993.

Pajak:1975:MFS

[1208] Thomas F. Pajak and Sidney Addelman. Minimum full sequences of 2^{n−m} resolution III plans. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):88–95, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984994.

Murthy:1975:EWP

[1209] V. K. Murthy and G. B. Swartz. Estimation of Weibull parameters from two-order statistics. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):96–102, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984995.

Lochner:1975:GDD

[1210] Robert H. Lochner. A generalized Dirichlet distribution in Bayesian life testing. Journal of the Royal Statistical Society. Series B (Methodological), 37 (1):103–113, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984996.

Freeman:1975:RCD

[1211] G. H. Freeman. Row-and-column designs with two groups of treatments having different replications. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):114–128, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984997.

Beale:1975:MVM

[1212] E. M. L. Beale and R. J. A. Little.

Missing values in multivariate analysis.

Journal of the Royal Statistical Society. Series B (Methodological), 37(1): 129–145, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984998.

Scott:1975:CEU

[1213] Alastair Scott and T. M. F. Smith. Comments on “The Efficient Use of Supplementary Information in Standard Sampling Procedures” by E. K. Foreman and K. R. W. Brewer. Journal of the Royal Statistical Society. Series B (Methodological), 37(1):146–148, ???? 1975. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984999. See [1044].

Anonymous:1975:BMb

[1214] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):i–ii, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984903.

Brown:1975:TTC

[1215] R. L. Brown, J. Durbin, and J. M. Evans. Techniques for testing the constancy of regression relationships over time. Journal of the Royal Statistical Society. Series B (Methodological), 37 (2):149–192, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984889.

Sahai:1975:BEE

[1216] Hardeo Sahai. Bayes equivariant estimators in high order hierarchical random effects models. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):193–197, ????

1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984890.

Whitehead:1975:STC

[1217] John Whitehead. A sequential test for certain composite hypotheses. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):198–204, ???? 1975. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984891.

Godolphin:1975:ETP

[1218] E. J. Godolphin and P. J. Harrison.

Equivalence theorems for polynomialprojecting predictors. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):205–215, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984892.

Durbin:1975:CCM

[1219] J. Durbin, M. Knott, and C. C.

Taylor. Components of Cramér– von Mises statistics. II. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):216–237, ???? 1975. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984893. See corrigenda

[1327].

Oakes:1975:SAD

[1220] David Oakes. Synchronous and asynchronous distributions for Poisson cluster processes. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):238–247, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984894.

Dawid:1975:CSA

[1221] A. P. Dawid. On the concepts of sufficiency and ancillarity in the presence of nuisance parameters. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):248–258, ???? 1975. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984895.

Daley:1975:IID

[1222] D. J. Daley and D. N. Shanbhag. Independent inter-departure times in M/G/1/N queues. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):259–263, ???? 1975. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984896.

Ghosh:1975:AMS

[1223] Malay Ghosh. Admissibility and minimaxity of some maximum likelihood estimators when the parameter space is restricted to integers. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):264–271, ???? 1975. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984897.

Kruskal:1975:GGI

[1224] William Kruskal. The geometry of generalized inverses. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):272–283, ???? 1975. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984898. See corrections

[1771].

Richardson:1975:ANW

[1225] J. W. Richardson. Addition to the nearest whole unit. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):284–287, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984899.

Sibson:1975:COE

[1226] Robin Sibson and Alan Kenny. Coefficients in D-optimal experimental design. Journal of the Royal Statistical Society. Series B (Methodological), 37 (2):288–292, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984900.

Agren:1975:CGF

[1227] Anders Agren. The consistency of the GEID-(FP-) estimates. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):293–295, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984901.

Anonymous:1975:CPE

[1228] Anonymous. Correction: Planning experiments for discriminating between models. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):296, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984902. See

[1183].

Anonymous:1975:FMb

[1229] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 37(2):??, ???? 1975. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984888.

Cliff:1975:MBA

[1230] A. D. Cliff and J. K. Ord. Model building and the analysis of spatial pattern in human geography. Journal of the Royal Statistical Society. Series B (Methodological), 37(3):297–348, ???? 1975. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984781.

Mardia:1975:SDD

[1231] K. V. Mardia. Statistics of directional data. Journal of the Royal Statistical Society. Series B (Methodological), 37 (3):349–393, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984782.

Siegmund:1975:EPA

[1232] D. Siegmund. Error probabilities and average sample number of the sequential probability ratio test. Journal of the Royal Statistical Society. Series B (Methodological), 37(3):394–401, ???? 1975. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984783.

Goldstein:1975:URL

[1233] M. Goldstein. Uniqueness relations for linear posterior expectations. Journal of the Royal Statistical Society. Series B (Methodological), 37(3):402–405, ???? 1975. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984784.

Venables:1975:CCI

[1234] W. Venables. Calculation of confidence intervals for noncentrality parameters.

Journal of the Royal Statistical Society. Series B (Methodological), 37(3): 406–412, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984785.

Bagshaw:1975:IRV

[1235] Michael Bagshaw and Richard A. Johnson. The influence of reference values and estimated variance on the ARL of Cusum tests. Journal of the Royal Statistical Society. Series B (Methodological), 37(3):413–420, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984786.

Maritz:1975:CSE

[1236] J. S. Maritz and T. Lwin. Construction of simple empirical Bayes estimators. Journal of the Royal Statistical Society. Series B (Methodological), 37 (3):421–425, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984787.

Hochberg:1975:EMG

[1237] Yosef Hochberg. An extension of the T -method to general unbalanced models of fixed effects. Journal of the Royal Statistical Society. Series B (Methodological), 37(3):426–433, ???? 1975. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984788.

Lancaster:1975:JPD

[1238] H. O. Lancaster. Joint probability distributions in the Meixner classes. Journal of the Royal Statistical Society. Series B (Methodological), 37(3):434–443, ???? 1975. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984789.

Bartholomew:1975:EPM

[1239] D. J. Bartholomew. Errors of prediction for Markov chain models. Journal of the Royal Statistical Society. Series B (Methodological), 37(3):444–456, ???? 1975. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984790.

Puri:1975:PEB

[1240] P. D. Puri and A. K. Nigam. On patterns of efficiency-balanced designs. Journal of the Royal Statistical Society. Series B (Methodological), 37(3): 457–458, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984791.

Barbour:1975:NMS

[1241] Andrew D. Barbour. A note on the maximum size of a closed epidemic. Journal of the Royal Statistical Society. Series B (Methodological), 37(3): 459–460, ???? 1975. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984792.

Anonymous:1975:BMc

[1242] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 37(3):??, ???? 1975. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984793.

Anonymous:1975:FMc

[1243] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 37(3):??, ???? 1975. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984780.

Anonymous:1976:FMa

[1244] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):i–ii, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984824.

Anderson:1976:ELF

[1245] T. W. Anderson. Estimation of linear functional relationships: Approximate distributions and connections with simultaneous equations in econometrics. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):1–36, ???? 1976. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984825.

Barndorff-Nielsen:1976:FLF

[1246] O. Barndorff-Nielsen. Factorization of likelihood functions for full exponential families. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):37–44, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984826.

Crowder:1976:MLE

[1247] Martin J. Crowder. Maximum likelihood estimation for dependent observations. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):45–53, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984827.

Vasicek:1976:TNB

[1248] Oldrich Vasicek. A test for normality based on sample entropy. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):54–59, ????

1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984828.

Brillinger:1976:ESO

[1249] David R. Brillinger. Estimation of the second-order intensities of a bivariate stationary point process. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):60–66, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984829.

Glazebrook:1976:MSS

[1250] K. D. Glazebrook and P. Nash. On multi-server stochastic scheduling. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):67–72, ???? 1976. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984830.

Chamberlain:1976:MWA

[1251] Gary Chamberlain and Edward E. Leamer. Matrix weighted averages and posterior bounds. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):73–84, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984831.

Leamer:1976:BIP

[1252] Edward E. Leamer and Gary Chamberlain. A Bayesian interpretation of pretesting. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):85–94, ???? 1976. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984832.

Leonard:1976:ITP

[1253] Tom Leonard and Keith Ord. An investigation of the F -test procedure as an estimation short-cut. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):95–98, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984833.

Abakuks:1976:IPP

[1254] Andris Abakuks. An invariance property of the Poisson distributions for a hunted immigrant population. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):99–101, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984834.

Curnow:1976:CNV

[1255] R. N. Curnow. Corrigenda: The number of variables when searching for an optimum. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):102, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984836. See

[1095].

Stone:1976:CCV

[1256] M. Stone. Corrigenda: Crossvalidatory choice and assessment of statistical predictions. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):102, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984835. See [1169].

Anonymous:1976:BMa

[1257] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 38(1):??, ???? 1976. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984837.

Anonymous:1976:VI

[1258] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 38(1): ??, ???? 1976. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984823.

Anonymous:1976:BMb

[1259] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 38(2):i–ii, ???? 1976. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985026.

Barndorff-Nielsen:1976:PI

[1260] O. Barndorff-Nielsen. Plausibility inference. Journal of the Royal Statistical Society. Series B (Methodological), 38 (2):103–131, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2985016.

Hawkins:1976:SPM

[1261] Douglas M. Hawkins. The subset problem in multivariate analysis of variance. Journal of the Royal Statistical Society. Series B (Methodological), 38(2): 132–139, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2985017. See

corrigenda [1328].

Wahba:1976:HKW

[1262] Grace Wahba. Histosplines with knots which are order statistics. Journal of the Royal Statistical Society. Series B (Methodological), 38(2):140–151, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985018.

Pelz:1976:ALT

[1263] Wolfgang Pelz and I. J. Good. Approximating the lower tail-areas of the Kolmogorov–Smirnov one-sample statistic. Journal of the Royal Statistical Society. Series B (Methodological), 38(2):152–156, ???? 1976. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985019.

Marcus:1976:SSC

[1264] Ruth Marcus and Eric Peritz. Some simultaneous confidence bounds in normal models with restricted alternatives. Journal of the Royal Statistical Society. Series B (Methodological), 38(2): 157–165, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985020.

Berger:1976:ESS

[1265] James O. Berger and M. E. Bock. Eliminating singularities of Stein-type estimators of location vectors. Journal of the Royal Statistical Society. Series B (Methodological), 38(2):166–170, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985021.

Williams:1976:RPC

[1266] E. R. Williams. Resolvable pairedcomparison designs. Journal of the Royal Statistical Society. Series B (Methodological), 38(2):171–174, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985022.

Patterson:1976:GFD

[1267] H. D. Patterson. Generation of factorial designs. Journal of the Royal Statistical Society. Series B (Methodological), 38(2):175–179, ???? 1976. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985023.

Robinson:1976:APD

[1268] P. M. Robinson. The aliasing problem in differential equation estimation. Journal of the Royal Statistical Society. Series B (Methodological), 38(2): 180–188, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2985024.

Granger:1976:FTS

[1269] C. W. J. Granger and P. Newbold. Forecasting transformed series. Journal of the Royal Statistical Society. Series B (Methodological), 38(2):189–203, ???? 1976. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985025.

Anonymous:1976:FMb

[1270] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 38(2):??, ???? 1976. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985015.

Harrison:1976:BF

[1271] P. J. Harrison and C. F. Stevens. Bayesian forecasting. Journal of the Royal Statistical Society. Series B (Methodological), 38(3):205–247, ???? 1976. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984970.

Farebrother:1976:FRM

[1272] R. W. Farebrother. Further results on the mean square error of ridge regression. Journal of the Royal Statistical Society. Series B (Methodological), 38 (3):248–250, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984971.

Breth:1976:NPC

[1273] M. Breth. Non-parametric confidence intervals for a mean using censored data. Journal of the Royal Statistical Society. Series B (Methodological), 38 (3):251–254, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984972.

Prescott:1976:TNB

[1274] P. Prescott. On a test for normality based on sample entropy. Journal of the Royal Statistical Society. Series B (Methodological), 38(3):254–256, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984973.

Giesbrecht:1976:MLE

[1275] F. Giesbrecht and O. Kempthorne.

Maximum likelihood estimation in the three-parameter lognormal distribution. Journal of the Royal Statistical Society. Series B (Methodological), 38 (3):257–264, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984974.

Saunders:1976:PEC

[1276] Roy Saunders and Richard J. Kryscio.

Parameter estimation for the carrierborne epidemic model. Journal of the Royal Statistical Society. Series B (Methodological), 38(3):265–269, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984975.

Rubin:1976:NLS

[1277] Donald B. Rubin. Noniterative least squares estimates, standard errors and F -tests for analyses of variance with missing data. Journal of the Royal Statistical Society. Series B (Methodological), 38(3):270–274, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984976.

Imhof:1976:OFR

[1278] J. P. Imhof. Oscillations in the finite random sequence. Journal of the Royal Statistical Society. Series B (Methodological), 38(3):275–278, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984977.

Raktoe:1976:AMG

[1279] B. L. Raktoe. On alias matrices and generalized defining relationships of equi- information factorial arrangements. Journal of the Royal Statistical Society. Series B (Methodological), 38 (3):279–283, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984978.

Herzberg:1976:SCO

[1280] Agnes M. Herzberg and David F. Andrews. Some considerations in the optimal design of experiments in non-optimal situations. Journal of the Royal Statistical Society. Series B (Methodological), 38(3):284–289, ???? 1976. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984979. See corrigenda

[1329].

Turnbull:1976:EDF

[1281] Bruce W. Turnbull. The empirical distribution function with arbitrarily grouped, censored and truncated data. Journal of the Royal Statistical Society. Series B (Methodological), 38(3): 290–295, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984980.

Williams:1976:RDT

[1282] E. R. Williams, H. D. Patterson, and J. A. John. Resolvable designs with two replications. Journal of the Royal Statistical Society. Series B (Methodological), 38(3):296–301, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984981.

Hoel:1976:STC

[1283] D. G. Hoel, G. H. Weiss, and R. Simon. Sequential tests for composite hypotheses with two binomial populations. Journal of the Royal Statistical Society. Series B (Methodological), 38 (3):302–308, ???? 1976. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984982.

Anonymous:1976:BMc

[1284] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 38(3):??, ???? 1976. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984983.

Anonymous:1976:FMc

[1285] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 38(3):??, ???? 1976. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984969.

Dempster:1977:MLI

[1286] A. P. Dempster, N. M. Laird, and D. B. Rubin. Maximum likelihood from incomplete data via the EM algorithm. Journal of the Royal Statistical Society. Series B (Methodological), 39(1):1–38, ???? 1977. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984875.

John:1977:OIB

[1287] J. A. John and Toby J. Mitchell. Optimal incomplete block designs. Journal of the Royal Statistical Society. Series B (Methodological), 39(1):39–43, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984876.

Stone:1977:AEC

[1288] M. Stone. An asymptotic equivalence of choice of model by cross-validation and Akaike’s criterion. Journal of the Royal Statistical Society. Series B (Methodological), 39(1):44–47, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984877.

Charnock:1977:CTS

[1289] D. M. Charnock. Continuous time spectral analysis of intervals between events in stationary point processes. Journal of the Royal Statistical Society. Series B (Methodological), 39(1):48–55, ???? 1977. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984878.

Davy:1977:STO

[1290] Pamela Davy and R. E. Miles. Sampling theory for opaque spatial specimens. Journal of the Royal Statistical Society. Series B (Methodological), 39 (1):56–65, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984879.

Bhansali:1977:APW

[1291] R. J. Bhansali. Asymptotic properties of the Wiener–Kolmogorov predictor. II. Journal of the Royal Statistical Society. Series B (Methodological), 39 (1):66–72, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984880.

Besag:1977:EVE

[1292] Julian Besag. Errors-in-variables estimation for Gaussian lattice schemes. Journal of the Royal Statistical Society. Series B (Methodological), 39(1):73–78, ???? 1977. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984881.

Burr:1977:PES

[1293] P. Burr and D. H. Young. Properties of the exponential scores test for the k-sample problem under the Lehmann model. Journal of the Royal Statistical Society. Series B (Methodological), 39(1):79–85, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984882.

Cornfield:1977:BLT

[1294] Jerome Cornfield and Katherine Detre. Bayesian life table analysis. Journal of the Royal Statistical Society. Series B (Methodological), 39(1):86–94, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/

stable/2984883. See remarks [1348].

Khatri:1977:MFM

[1295] C. G. Khatri and K. V. Mardia. The von Mises–Fisher matrix distribution in orientation statistics. Journal of the Royal Statistical Society. Series B (Methodological), 39(1):95–106, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984884.

Clark:1977:NPE

[1296] R. M. Clark. Non-parametric estimation of a smooth regression function. Journal of the Royal Statistical Society. Series B (Methodological), 39(1): 107–113, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984885.

Osborn:1977:EAM

[1297] Denise R. Osborn. Exact and approximate maximum likelihood estimators for vector moving average processes. Journal of the Royal Statistical Society. Series B (Methodological), 39(1): 114–118, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984886.

Anonymous:1977:BMa

[1298] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 39(1):??, ????

1977. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984887.

Anonymous:1977:FMa

[1299] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 39(1):??, ???? 1977. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984874.

Anonymous:1977:VI

[1300] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Methodological), 39(1): ??, ???? 1977. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984873.

Wilkinson:1977:RCS

[1301] G. N. Wilkinson. On resolving the controversy in statistical inference. Journal of the Royal Statistical Society. Series B (Methodological), 39(2):119–171, ???? 1977. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984795.

Ripley:1977:MSP

[1302] B. D. Ripley. Modelling spatial patterns. Journal of the Royal Statistical Society. Series B (Methodological), 39 (2):172–212, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984796.

Veevers:1977:MBF

[1303] A. Veevers and M. C. K. Tweedie. On the moments of a beta function transformation of a Poisson variate. Journal of the Royal Statistical Society. Series B (Methodological), 39(2):213–216, ???? 1977. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984797.

Khuri:1977:DQF

[1304] A. I. Khuri and I. J. Good. The distribution of quadratic forms in nonnormal variables and an application to the variance ratio. Journal of the Royal Statistical Society. Series B (Methodological), 39(2):217–221, ???? 1977. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984798.

Johnson:1977:MMA

[1305] Richard A. Johnson and Thomas Wehrly. Measures and models for angular correlation and angular-linear correlation. Journal of the Royal Statistical Society. Series B (Methodological), 39 (2):222–229, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984799.

Charnock:1977:SAI

[1306] D. M. Charnock. Spectral analysis of intervals between events in stationary bivariate point processes. Journal of the Royal Statistical Society. Series B (Methodological), 39(2):230–237, ???? 1977. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984800.

Godolphin:1977:PES

[1307] E. J. Godolphin. A procedure for estimating seasonal moving average models based on large- sample estimation of the correlogram. Journal of the Royal Statistical Society. Series B (Methodological), 39(2):238–247, ???? 1977. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984801.

Benedetti:1977:NER

[1308] Jacqueline K. Benedetti. On the nonparametric estimation of regression functions. Journal of the Royal Statistical Society. Series B (Methodological), 39(2):248–253, ???? 1977. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984802.

Dawid:1977:SMD

[1309] A. P. Dawid. Spherical matrix distributions and a multivariate model. Journal of the Royal Statistical Society. Series B (Methodological), 39(2):254–261, ???? 1977. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984803.

Diaconis:1977:SFM

[1310] Persi Diaconis and R. L. Graham.

Spearman’s footrule as a measure of disarray. Journal of the Royal Statistical Society. Series B (Methodological), 39(2):262–268, ???? 1977. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984804.

Muenz:1977:TSC

[1311] Larry R. Muenz and Sylvan B. Green.

Time savings in censored life testing. Journal of the Royal Statistical Society. Series B (Methodological), 39(2): 269–275, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984805.

Gipps:1977:QMT

[1312] P. G. Gipps. A queueing model for traffic flow. Journal of the Royal Statistical Society. Series B (Methodological), 39 (2):276–282, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984806.

Anonymous:1977:BMb

[1313] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 39(2):??, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984807.

Anonymous:1977:FMb

[1314] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 39(2):??, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984794.

Mollison:1977:SCM

[1315] Denis Mollison. Spatial contact models for ecological and epidemic spread. Journal of the Royal Statistical Society. Series B (Methodological), 39(3): 283–326, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2985089.

DeJong:1977:FFT

[1316] Piet De Jong. The fast Fourier transform spectral estimator. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):327–330, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985090.

Shanbhag:1977:NID

[1317] D. N. Shanbhag and M. Westcott. A note on infinitely divisible point processes. Journal of the Royal Statistical Society. Series B (Methodological), 39 (3):331–332, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985091.

Koziol:1977:GFT

[1318] James A. Koziol. On goodness of fit tests based on the empirical distribution function for uniform spacings. Journal of the Royal Statistical Society. Series B (Methodological), 39(3): 333–336, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985092.

Anderson:1977:IPD

[1319] T. W. Anderson and S. G. Ghurye.

Identification of parameters by the distribution of a maximum random variable. Journal of the Royal Statistical Society. Series B (Methodological), 39 (3):337–342, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985093.

El-Shaarawi:1977:MLS

[1320] A. El-Shaarawi. Marginal likelihood solution to some problems connected with regression analysis. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):343–348, ???? 1977. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985094.

Tripathi:1977:GFD

[1321] Ram C. Tripathi and John Gurland.

A general family of discrete distributions with hypergeometric probabilities. Journal of the Royal Statistical Society. Series B (Methodological), 39 (3):349–356, ???? 1977. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985095.

Singh:1977:AED

[1322] R. S. Singh. Applications of estimators of a density and its derivatives to certain statistical problems. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):357–363, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985096.

Pettitt:1977:CMT

[1323] A. N. Pettitt. A Cramér–von Mises type goodness of fit statistic related to √b_{1} and b_{2}. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):364–370, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985097.

McKay:1977:VSM

[1324] R. J. McKay. Variable selection in multivariate regression: An application of simultaneous test procedures. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):371–380, ???? 1977. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985098.

Robinson:1977:CSI

[1325] G. K. Robinson. Conservative statistical inference. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):381–386, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985099.

Goodman:1977:QPI

[1326] T. N. T. Goodman. Qualitative probability and improper distributions. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):387–393, ???? 1977. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985100.

Durbin:1977:CCC

[1327] J. Durbin, M. Knott, and C. C.

Taylor. Corrigenda: Components of Cramér–von Mises statistics. II. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):394, ???? 1977. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985103. See [1219].

Hawkins:1977:CSP

[1328] Douglas M. Hawkins. Corrigenda:

The subset problem in multivariate analysis of variance. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):394, ???? 1977. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985101. See [1261].

Herzberg:1977:CSC

[1329] A. M. Herzberg and D. F. Andrews.

Corrigenda: Some considerations in the optimal design of experiments in non-optimal situations. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):394, ???? 1977. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985102. See [1280].

Anonymous:1977:BMc

[1330] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):??, ???? 1977. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985104.

Anonymous:1977:FMc

[1331] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 39(3):??, ???? 1977. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985088.

Anonymous:1978:BMa

[1332] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 40(1):i–ii, ???? 1978. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984872.

OHagan:1978:CFO

[1333] A. O’Hagan and J. F. C. Kingman. Curve fitting and optimal design for prediction. Journal of the Royal Statistical Society. Series B (Methodological), 40(1):1–42, ???? 1978. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984861.

Dickey:1978:BFM

[1334] J. M. Dickey and E. Gunel. Bayes factors from mixed probabilities. Journal of the Royal Statistical Society. Series B (Methodological), 40(1):43–46, ???? 1978. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984862.

Farebrother:1978:CSE

[1335] R. W. Farebrother. A class of shrinkage estimators. Journal of the Royal Statistical Society. Series B (Methodological), 40(1):47–49, ???? 1978. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984863.

Brooks:1978:BAT

[1336] R. J. Brooks. Bayesian analysis of a two-sample problem based on the rank order statistic. Journal of the Royal Statistical Society. Series B (Methodological), 40(1):50–57, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984864.

Singh:1978:TWE

[1337] M. Singh and A. Dey. Two-way elimination of heterogeneity. Journal of the Royal Statistical Society. Series B (Methodological), 40(1):58–63, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984865.

Stephens:1978:HSM

[1338] M. A. Stephens. On the half-sample method for goodness-of-fit. Journal of the Royal Statistical Society. Series B (Methodological), 40(1):64–70, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984866.

Cotter:1978:BDF

[1339] Sarah C. Cotter. Block designs for fractional factorial experiments with replication. Journal of the Royal Statistical Society. Series B (Methodological), 40 (1):71–78, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984867.

Dean:1978:AIS

[1340] A. M. Dean. The analysis of interactions in single replicate generalized cyclic designs. Journal of the Royal Statistical Society. Series B (Methodological), 40(1):79–84, ????

1978. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984868.

Andrews:1978:FOM

[1341] David F. Andrews and Daryl Pregibon. Finding the outliers that matter. Journal of the Royal Statistical Society. Series B (Methodological), 40(1):85–93, ???? 1978. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984869.

Jacobs:1978:DTSa

[1342] P. A. Jacobs and P. A. W. Lewis. Discrete time series generated by mixtures. I: Correlational and runs properties. Journal of the Royal Statistical Society. Series B (Methodological), 40(1): 94–105, ???? 1978. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

                                        jstor.org/stable/2984870.

Smith:1978:QBS

[1343] A. F. M. Smith and U. E. Makov. A quasi-Bayes sequential procedure for mixtures. Journal of the Royal Statistical Society. Series B (Methodological), 40(1):106–112, ???? 1978. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984871.

Anonymous:1978:FMa

[1344] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 40(1):??, ???? 1978. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984860.

Anonymous:1978:BMb

[1345] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 40(2):i–ii, ????

1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984764.

Leonard:1978:DES

[1346] Tom Leonard. Density estimation, stochastic processes and prior information. Journal of the Royal Statistical Society. Series B (Methodological), 40 (2):113–146, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984749.

Bartlett:1978:NNM

[1347] M. S. Bartlett. Nearest neighbour models in the analysis of field experiments. Journal of the Royal Statistical Society. Series B (Methodological), 40(2): 147–174, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984750.

Kalbfleisch:1978:RPC

[1348] J. D. Kalbfleisch and R. J. MacKay.

Remarks on a paper by Cornfield and Detre. Journal of the Royal Statistical Society. Series B (Methodological), 40 (2):175–177, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984751. See

[1294].

Diggle:1978:PES

[1349] Peter J. Diggle. On parameter estimation for spatial point processes. Journal of the Royal Statistical Society. Series B (Methodological), 40(2):178–181, ???? 1978. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984752.

McKenzie:1978:NMI

[1350] E. McKenzie and M. M. Peterson. A note on a matrix inversion in multiple exponential smoothing. Journal of the Royal Statistical Society. Series B (Methodological), 40(2):182–183, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984753.

Eplett:1978:NAM

[1351] W. J. R. Eplett. A note about the multipliers in latent root regression. Journal of the Royal Statistical Society. Series B (Methodological), 40(2):184–185, ???? 1978. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984754.

Gabriel:1978:LSA

[1352] K. R. Gabriel. Least squares approximation of matrices by additive and multiplicative models. Journal of the Royal Statistical Society. Series B (Methodological), 40(2):186–196, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984755.

Anderson:1978:SSV

[1353] C. W. Anderson. Super-slowly varying functions in extreme value theory. Journal of the Royal Statistical Society. Series B (Methodological), 40(2): 197–202, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984756.

Feigin:1978:MCB

[1354] Paul D. Feigin and Ayala Cohen. On a model for concordance between judges. Journal of the Royal Statistical Society. Series B (Methodological), 40(2):

203–213, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984757.

Kalbfleisch:1978:NPB

[1355] John D. Kalbfleisch. Non-parametric Bayesian analysis of survival time data. Journal of the Royal Statistical Society. Series B (Methodological), 40(2): 214–221, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984758.

Jacobs:1978:DTSb

[1356] Patricia A. Jacobs and Peter A. W.

Lewis. Discrete time series generated by mixtures II: Asymptotic properties. Journal of the Royal Statistical Society. Series B (Methodological), 40(2): 222–228, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984759.

Mardia:1978:MCV

[1357] K. V. Mardia and T. W. Sutton.

A model for cylindrical variables with applications. Journal of the Royal Statistical Society. Series B (Methodological), 40(2):229–233, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984760.

Sibson:1978:SRM

[1358] Robin Sibson. Studies in the robustness of multidimensional scaling: Procrustes statistics. Journal of the Royal Statistical Society. Series B (Methodological), 40(2):234–238, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984761.

Sedransk:1978:CIQ

[1359] J. Sedransk and J. Meyer. Confidence intervals for the quantiles of a finite population: Simple random and stratified simple random sampling. Journal of the Royal Statistical Society. Series B (Methodological), 40(2):239–252, ???? 1978. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984762.

Raviv:1978:NPT

[1360] Alona Raviv. A non-parametric test for comparing two non-independent distributions. Journal of the Royal Statistical Society. Series B (Methodological), 40 (2):253–261, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984763. See

corrigenda [1607].

Anonymous:1978:FMb

[1361] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 40(2):??, ???? 1978. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984748.

Folks:1978:IGD

[1362] J. L. Folks and R. S. Chhikara. The inverse Gaussian distribution and its statistical application — a review. Journal of the Royal Statistical Society. Series B (Methodological), 40(3):263–289, ???? 1978. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984691.

Godolphin:1978:LST

[1363] E. J. Godolphin. A large-sample test for detecting gaps in moving average models. Journal of the Royal Statistical Society. Series B (Methodological), 40 (3):290–295, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984692.

McLeod:1978:DRA

[1364] A. I. McLeod. On the distribution of residual autocorrelations in Box–Jenkins models. Journal of the Royal Statistical Society. Series B (Methodological), 40(3):296–302, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984693.

ONeill:1978:DEC

[1365] M. E. O’Neill. Distributional expansions for canonical correlations from contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 40(3):303–312, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984694.

Godolphin:1978:IPU

[1366] E. J. Godolphin and M. De Tullio.

Invariance properties of uncorrelated residual transformations. Journal of the Royal Statistical Society. Series B (Methodological), 40(3):313–321, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984695.

Weber:1978:OSM

[1367] Richard R. Weber and Peter Nash.

An optimal strategy in multi-server stochastic scheduling. Journal of the Royal Statistical Society. Series B (Methodological), 40(3):322–327, ???? 1978. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984696.

Whittemore:1978:CMC

[1368] Alice S. Whittemore. Collapsibility of multidimensional contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 40(3): 328–340, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984697.

Hosking:1978:UDA

[1369] J. R. M. Hosking. A unified derivation of the asymptotic distributions of goodness-of-fit statistics for autoregressive time-series models. Journal of the Royal Statistical Society. Series B (Methodological), 40(3):341–349, ???? 1978. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984698.

Gail:1978:SFG

[1370] M. H. Gail and J. L. Gastwirth. A scale-free goodness-of-fit test for the exponential distribution based on the Gini statistic. Journal of the Royal Statistical Society. Series B (Methodological), 40(3):350–357, ???? 1978. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984699.

Berman:1978:EFC

[1371] Mark Berman. Exponential families and conditioning on statistics which are not minimal sufficient. Journal of the Royal Statistical Society. Series B (Methodological), 40(3):358–363, ???? 1978. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984700.

Wahba:1978:IPS

[1372] Grace Wahba. Improper priors, spline smoothing and the problem of guarding against model errors in regression. Journal of the Royal Statistical Society. Series B (Methodological), 40(3): 364–372, ???? 1978. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984701.

Farebrother:1978:HNR

[1373] R. W. Farebrother. An historical note on recursive residuals. Journal of the Royal Statistical Society. Series B (Methodological), 40(3):373–375, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984702.

Anonymous:1978:BMc

[1374] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 40(3):??, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984703.

Anonymous:1978:FMc

[1375] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 40(3):??, ???? 1978. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984690.

Anonymous:1978:VI

[1376] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Methodological), 40(3): ??, ???? 1978. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984689.

Anonymous:1979:BMa

[1377] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 41(1):i–ii, ???? 1979. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984731.

Dawid:1979:CIS

[1378] A. P. Dawid. Conditional independence in statistical theory. Journal of the Royal Statistical Society. Series B (Methodological), 41(1):1–31, ???? 1979. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984718.

Liebetrau:1979:STR

[1379] Albert M. Liebetrau. Some tests of randomness based upon the variance-time curve of the Poisson process. Journal of the Royal Statistical Society. Series B (Methodological), 41(1):32–39, ???? 1979. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984719.

Raghavarao:1979:BTR

[1380] D. Raghavarao and W. T. Federer. Block total response as an alternative to the randomized response method in surveys. Journal of the Royal Statistical Society. Series B (Methodological), 41(1):40–45, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984720.

Pettitt:1979:TSC

[1381] A. N. Pettitt. Two-sample Cramér– von Mises type rank statistics. Journal of the Royal Statistical Society. Series B (Methodological), 41(1):46–53, ????

1979. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984721.

Fienberg:1979:UCS

[1382] Stephen E. Fienberg. The use of chisquared statistics for categorical data problems. Journal of the Royal Statistical Society. Series B (Methodological), 41(1):54–64, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984722.

McGilchrist:1979:REG

[1383] C. A. McGilchrist and R. L. Sandland.

Recursive estimation of the general linear model with dependent errors. Journal of the Royal Statistical Society. Series B (Methodological), 41(1):65–68, ???? 1979. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984723.

Winterbottom:1979:CFE

[1384] Alan Winterbottom. Cornish–Fisher expansions for confidence limits. Journal of the Royal Statistical Society. Series B (Methodological), 41(1):69–75, ???? 1979. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984724.

Little:1979:MLI

[1385] Roderick J. A. Little. Maximum likelihood inference for multiple regression with missing values: A simulation study. Journal of the Royal Statistical Society. Series B (Methodological), 41 (1):76–87, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984725.

Freeman:1979:STD

[1386] G. H. Freeman. Some two-dimensional designs balanced for nearest neighbours. Journal of the Royal Statistical Society. Series B (Methodological), 41 (1):88–95, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984726.

Goldstein:1979:VML

[1387] Michael Goldstein. The variance modified linear Bayes estimator. Journal of the Royal Statistical Society. Series B (Methodological), 41(1):96–100, ???? 1979. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984727.

Borgan:1979:CTS

[1388] Ornulf Borgan. Comparison of two sequential tests for two-sided alternatives. Journal of the Royal Statistical Society. Series B (Methodological), 41 (1):101–106, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984728.

Godambe:1979:BAM

[1389] V. P. Godambe. On Birnbaum’s axiom of mathematically equivalent experiments. Journal of the Royal Statistical Society. Series B (Methodological), 41 (1):107–110, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984729.

McGinley:1979:NDO

[1390] W. G. McGinley. A note on dissociatedness and order statistics. Journal of the Royal Statistical Society. Series B (Methodological), 41(1):111–112, ???? 1979. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984730.

Anonymous:1979:FMa

[1391] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 41(1):??, ???? 1979. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984717.

Bernardo:1979:RPD

[1392] Jose M. Bernardo. Reference posterior distributions for Bayesian inference. Journal of the Royal Statistical Society. Series B (Methodological), 41 (2):113–147, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985028.

Gittins:1979:BPD

[1393] J. C. Gittins. Bandit processes and dynamic allocation indices. Journal of the Royal Statistical Society. Series B (Methodological), 41(2):148–177, ???? 1979. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985029. See [1438].

Fligner:1979:UCA

[1394] Michael A. Fligner and Thomas P.

Hettmansperger. On the use of conditional asymptotic normality. Journal of the Royal Statistical Society. Series B (Methodological), 41(2):178–183, ???? 1979. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985030.

Heyde:1979:APN

[1395] C. C. Heyde and I. M. Johnstone.

On asymptotic posterior normality for stochastic processes. Journal of the Royal Statistical Society. Series B (Methodological), 41(2):184–189, ???? 1979. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985031.

Hannan:1979:DOA

[1396] E. J. Hannan and B. G. Quinn. The determination of the order of an autoregression. Journal of the Royal Statistical Society. Series B (Methodological), 41(2):190–195, ???? 1979. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985032.

Billard:1979:SDS

[1397] L. Billard, H. Lacayo, and Naftali A. Langberg. The symmetric mdimensional simple epidemic process. Journal of the Royal Statistical Society. Series B (Methodological), 41(2): 196–202, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985033.

McDunnough:1979:EAP

[1398] Philip McDunnough. The estimation of an autocorrelation parameter of a Gaussian vector process. Journal of the Royal Statistical Society. Series B (Methodological), 41(2):203–209, ???? 1979. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985034.

Jones:1979:ASO

[1399] Byron Jones. Algorithms to search for optimal row-and-column designs. Journal of the Royal Statistical Society. Series B (Methodological), 41(2):210–216, ???? 1979. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985035.

Sibson:1979:SRM

[1400] Robin Sibson. Studies in the robustness of multidimensional scaling: Perturbational analysis of classical scaling. Journal of the Royal Statistical Society. Series B (Methodological), 41(2): 217–229, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985036.

Crowder:1979:IAI

[1401] Martin J. Crowder. Inference about the intraclass correlation coefficient in the beta-binomial ANOVA for proportions. Journal of the Royal Statistical Society. Series B (Methodological), 41 (2):230–234, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985037.

Chen:1979:BIN

[1402] Chan-Fu Chen. Bayesian inference for a normal dispersion matrix and its application to stochastic multiple regression analysis. Journal of the Royal Statistical Society. Series B (Methodological), 41(2):235–248, ???? 1979. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985038.

Dawid:1979:SMA

[1403] A. P. Dawid. Some misleading arguments involving conditional independence. Journal of the Royal Statistical Society. Series B (Methodological), 41 (2):249–252, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985039.

Freeman:1979:CLS

[1404] G. H. Freeman. Complete Latin squares and related experimental designs. Journal of the Royal Statistical Society. Series B (Methodological), 41 (2):253–262, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2985040.

Chan:1979:MLE

[1405] Lai K. Chan and Tak K. Mak. Maximum likelihood estimation of a linear structural relationship with replication. Journal of the Royal Statistical Society. Series B (Methodological), 41(2): 263–268, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2985041.

Sugden:1979:ISF

[1406] R. A. Sugden. Inference on symmetric functions of exchangeable populations. Journal of the Royal Statistical Society. Series B (Methodological), 41 (2):269–273, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2985042.

Oakes:1979:LLH

[1407] David Oakes. On log-linear hazards. Journal of the Royal Statistical Society. Series B (Methodological), 41(2): 274–275, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2985043.

Stone:1979:CMS

[1408] M. Stone. Comments on model selection criteria of Akaike and Schwarz. Journal of the Royal Statistical Society. Series B (Methodological), 41(2): 276–278, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2985044.

Anonymous:1979:BMb

[1409] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 41(2):??, ???? 1979. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985045.

Anonymous:1979:FMb

[1410] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 41(2):??, ???? 1979. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985027.

Anonymous:1979:BMc

[1411] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 41(3):i–ii, ???? 1979. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985068.

Barndorff-Nielsen:1979:ESP

[1412] O. Barndorff-Nielsen and D. R. Cox.

Edgeworth and saddle-point approximations with statistical applications. Journal of the Royal Statistical Society. Series B (Methodological), 41(3): 279–312, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985061.

Kleiner:1979:REP

[1413] B. Kleiner and R. D. Martin. Robust estimation of power spectra. Journal of the Royal Statistical Society. Series B (Methodological), 41(3):313–351, ???? 1979. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985062.

Lewis:1979:CRI

[1414] S. M. Lewis. The construction of resolution III fractions from generalized cyclic designs. Journal of the Royal Statistical Society. Series B (Methodological), 41(3):352–357, ???? 1979. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985063.

OHagan:1979:ORP

[1415] A. O’Hagan. On outlier rejection phenomena in Bayes inference. Journal of the Royal Statistical Society. Series B (Methodological), 41(3):358–367, ???? 1979. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985064.

Ripley:1979:TRS

[1416] B. D. Ripley. Tests of ‘randomness’ for spatial point patterns. Journal of the Royal Statistical Society. Series B (Methodological), 41(3):368–374, ???? 1979. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985065.

Smith:1979:GBS

[1417] J. Q. Smith. A generalization of the Bayesian steady forecasting model. Journal of the Royal Statistical Society. Series B (Methodological), 41(3): 375–387, ???? 1979. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2985066.

Cox:1979:RSL

[1418] D. R. Cox. A remark on systematic Latin squares. Journal of the Royal Statistical Society. Series B (Methodological), 41(3):388–389, ????

1979. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985067.

Anonymous:1979:FMc

[1419] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 41(3):??, ???? 1979. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985060.

Anonymous:1979:VI

[1420] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Methodological), 41(3): ??, ???? 1979. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985059.

Bates:1980:RCM

[1421] Douglas M. Bates and Donald G.

Watts. Relative curvature measures of nonlinearity. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):1–25, ???? 1980. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2984733.

Tamhane:1980:SBB

[1422] Ajit C. Tamhane. Selecting the better Bernoulli treatment using a matched samples design. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):26–30, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984734.

McDonald:1980:SAP

[1423] Gary C. McDonald. Some algebraic properties of ridge coefficients. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):31–34, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984735.

Godolphin:1980:SRP

[1424] E. J. Godolphin and J. M. Stone. On the structural representation for polynomial-projecting predictor models based on the Kalman filter. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):35–45, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984736.

Akaike:1980:IIP

[1425] Hirotugu Akaike. The interpretation of improper prior distributions as limits of data dependent proper prior distributions. Journal of the Royal Statistical Society. Series B (Methodological), 42 (1):46–52, ???? 1980. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984737.

Braun:1980:SMT

[1426] Henry Braun. A simple method for testing goodness of fit in the presence of nuisance parameters. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):53–63, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984738.

Eplett:1980:ICT

[1427] W. J. R. Eplett. An influence curve for two-sample rank tests. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):64–70, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984739.

Carroll:1980:RMT

[1428] Raymond J. Carroll. A robust method for testing transformations to achieve approximate normality. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):71–78, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984740.

Box:1980:VFD

[1429] G. E. P. Box and N. R. Draper. The variance function of the difference between two estimated responses. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):79–82, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984741.

DeHaan:1980:SAE

[1430] L. De Haan and S. I. Resnick. A simple asymptotic estimate for the index of a stable distribution. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):83–87, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984742.

Eaves:1980:EPL

[1431] D. M. Eaves. On exchangeable priors in lot acceptance. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):88–93, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984743.

John:1980:GCD

[1432] J. A. John. Generalized cyclic designs with b < v as duals of cyclic designs. Journal of the Royal Statistical Society.

Series B (Methodological), 42(1):94–95, ???? 1980. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984744.

Elandt-Johnson:1980:SPP

[1433] Regina C. Elandt-Johnson. Some prior and posterior distributions in survival analysis: A critical insight on relationships derived from cross-sectional data. Journal of the Royal Statistical Society. Series B (Methodological), 42(1): 96–106, ???? 1980. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

                                        jstor.org/stable/2984745.

Burridge:1980:COT

[1434] P. Burridge. On the Cliff–Ord test for spatial correlation. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):107–108, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984746.

Anonymous:1980:BMa

[1435] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):??, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984747.

Anonymous:1980:FMa

[1436] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 42(1):??, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984732.

McCullagh:1980:RMO

[1437] Peter McCullagh. Regression models for ordinal data. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):109–142, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984952.

Whittle:1980:MAB

[1438] P. Whittle. Multi-armed bandits and the Gittins index. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):143–149, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984953. See [1393].

Lawrance:1980:EAM

[1439] A. J. Lawrance and P. A. W. Lewis.

The exponential autoregressive-moving average EARMA (p,q) process. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):150–161, ???? 1980. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984954.

Ali:1980:CND

[1440] Mir M. Ali. Characterization of the normal distribution among the continuous symmetric spherical class. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):162–164, ???? 1980. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984955.

Nash:1980:GBP

[1441] Peter Nash. A generalized bandit problem. Journal of the Royal Statistical Society. Series B (Methodological), 42 (2):165–169, ???? 1980. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984956.

Hosking:1980:LMT

[1442] J. R. M. Hosking. Lagrange-multiplier tests of time-series models. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):170–181, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984957.

Quinn:1980:ODM

[1443] B. G. Quinn. Order determination for a multivariate autoregression. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):182–185, ???? 1980. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984958.

Brooks:1980:RET

[1444] R. J. Brooks. On the relative efficiency of two paired-data experiments. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):186–191, ???? 1980. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984959.

Shaked:1980:MEF

[1445] Moshe Shaked. On mixtures from exponential families. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):192–198, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984960.

Cheng:1980:OSB

[1446] Ching-Shui Cheng. On the Eoptimality of some block designs. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):199–204, ???? 1980. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984961.

Jacroux:1980:ORG

[1447] Mike Jacroux. On the E-optimality of regular graph designs. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):205–209, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984962.

Hallin:1980:IGI

[1448] Marc Hallin. Invertibility and generalized invertibility of time series models. Journal of the Royal Statistical Society. Series B (Methodological), 42(2): 210–212, ???? 1980. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984963.

Smith:1980:BFC

[1449] A. F. M. Smith and D. J. Spiegelhalter. Bayes factors and choice criteria for linear models. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):213–220, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984964.

Jones:1980:BLU

[1450] Roger G. Jones. Best linear unbiased estimators for repeated surveys. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):221–226, ???? 1980. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984965.

Clarke:1980:MLS

[1451] G. P. Y. Clarke. Moments of the least squares estimators in a nonlinear regression model. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):227–237, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984966.

Jones:1980:EIP

[1452] Byron Jones and John A. Eccleston. Exchange and interchange procedures to search for optimal designs. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):238–243, ???? 1980. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984967.

Anonymous:1980:BMb

[1453] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):??, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984968.

Anonymous:1980:FMb

[1454] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 42(2):??, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984951.

Tong:1980:TAL

[1455] H. Tong and K. S. Lim. Threshold autoregression, limit cycles and cyclical data. Journal of the Royal Statistical Society. Series B (Methodological), 42 (3):245–292, ???? 1980. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2985164.

Bartholomew:1980:FAC

[1456] D. J. Bartholomew. Factor analysis for categorical data. Journal of the Royal Statistical Society. Series B (Methodological), 42(3):293–321, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985165.

Anderson:1980:SEH

[1457] J. A. Anderson and A. Senthilselvan.

Smooth estimates for the hazard function. Journal of the Royal Statistical Society. Series B (Methodological), 42 (3):322–327, ???? 1980. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985166.

Reinsel:1980:APP

[1458] Greg Reinsel. Asymptotic properties of prediction errors for the multivariate autoregressive model using estimated parameters. Journal of the Royal Statistical Society. Series B (Methodological), 42(3):328–333, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985167.

Blaisdell:1980:PBC

[1459] E. A. Blaisdell, Jr. and D. Raghavarao.

Partially balanced change-over designs based on m-associate class PBIB designs. Journal of the Royal Statistical Society. Series B (Methodological), 42 (3):334–338, ???? 1980. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985168.

Sudderth:1980:FAP

[1460] William D. Sudderth. Finitely additive priors, coherence and the marginalization paradox. Journal of the Royal Statistical Society. Series B (Methodological), 42(3):339–341, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985169.

Glazebrook:1980:RDA

[1461] K. D. Glazebrook. On randomized dynamic allocation indices for the sequential design of experiments. Journal of the Royal Statistical Society. Series B (Methodological), 42(3):342–346, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985170.

Hinkley:1980:TJ

[1462] David Hinkley and Hai-Li Wang. A trimmed jackknife. Journal of the Royal Statistical Society. Series B (Methodological), 42(3):347–356, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985171.

ONeill:1980:DHO

[1463] M. E. O’Neill. The distribution of higher-order interactions in contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 42(3):357–365, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985172.

McKean:1980:GRP

[1464] Joseph W. McKean and Ronald M. Schrader. The geometry of robust procedures in linear models. Journal of the Royal Statistical Society. Series B (Methodological), 42(3):366–371, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985173.

Eccleston:1980:EIP

[1465] John A. Eccleston and Byron Jones.

Exchange and interchange procedures to search for optimal row-and-column designs. Journal of the Royal Statistical Society. Series B (Methodological), 42 (3):372–376, ???? 1980. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985174.

Nathan:1980:ESD

[1466] G. Nathan and D. Holt. The effect of survey design on regression analysis. Journal of the Royal Statistical Society. Series B (Methodological), 42(3): 377–386, ???? 1980. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985175.

Sankaranarayanan:1980:NAS

[1467] K. Sankaranarayanan. A note on the admissibility of some non-negative quadratic estimators. Journal of the Royal Statistical Society. Series B (Methodological), 42(3):387–389, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985176.

Anonymous:1980:BMc

[1468] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 42(3):??, ???? 1980. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985177.

Anonymous:1980:FMc

[1469] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 42(3):??, ???? 1980. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985163.

Anonymous:1980:VI

[1470] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 42(3): ??, ???? 1980. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985162.

Holgate:1981:PA

[1471] P. Holgate. Population algebras. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):1–19, ???? 1981. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985143.

Feuerverger:1981:EEC

[1472] Andrey Feuerverger and Philip Mc-Dunnough. On the efficiency of empirical characteristic function procedures. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):20–27, ???? 1981. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985144.

Pfeifer:1981:VSS

[1473] Phillip E. Pfeifer and Stuart Jay Deutsch. Variance of the sample spacetime autocorrelation function. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):28–33, ???? 1981. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985145.

Taniguchi:1981:EPP

[1474] Masanobu Taniguchi. An estimation procedure of parameters of a certain spectral density model. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):34–40, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985146.

Burridge:1981:NML

[1475] J. Burridge. A note on maximum likelihood estimation for regression models using grouped data. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):41–45, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985147.

Samaniego:1981:MLE

[1476] F. J. Samaniego and L. E. Jones. Maximum likelihood estimation for a class of multinomial distributions arising in realibility. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):46–52, ???? 1981. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985148.

Hayre:1981:ABS

[1477] L. S. Hayre. An asymptotically Bayes sequential test for comparing two normal means. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):53–60, ???? 1981. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985149.

Fligner:1981:NLT

[1478] Michael A. Fligner. A note on limit theorems for joint distributions with applications to linear signed rank statistics. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):61–64, ???? 1981. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985150.

Burridge:1981:EBA

[1479] J. Burridge. Empirical Bayes analysis of survival time data. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):65–75, ???? 1981. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985151.

John:1981:ECD

[1480] J. A. John. Efficient cyclic designs. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):76–80, ???? 1981. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985152.

Henery:1981:ACM

[1481] R. J. Henery. An approximation to certain multivariate normal probabilities. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):81–85, ???? 1981. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985153.

Henery:1981:PPM

[1482] R. J. Henery. Permutation probabilities as models for horse races. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):86–91, ???? 1981. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985154.

Key:1981:BSF

[1483] P. B. Key and E. J. Godolphin. On the Bayesian steady forecasting model. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):92–96, ???? 1981. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2985155.

Silverman:1981:UKD

[1484] B. W. Silverman. Using kernel density estimates to investigate multimodality. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):97–99, ???? 1981. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985156.

Bartlett:1981:FNU

[1485] M. S. Bartlett. A further note on the use of neighbouring plot values in the analysis of field experiments. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):100–102, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985157.

Hallin:1981:AIG

[1486] M. Hallin. Addendum to “Invertibility and generalized invertibility of time series models”. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):103, ???? 1981. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2985158.

Bhansali:1981:CAP

[1487] R. J. Bhansali. Corrections: Asymptotic properties of the Wiener– Kolmogorov predictor. I. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):104, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985159. See [1158].

Theobald:1981:CGM

[1488] C. M. Theobald. Corrections: Generalizations of mean square error applied to ridge regression. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):104, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985160. See [1162].

Anonymous:1981:BMa

[1489] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):??, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985161.

Anonymous:1981:FMa

[1490] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 43(1):??, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985142.

Goldstein:1981:RPG

[1491] Michael Goldstein. Revising previsions: A geometric interpretation. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):105–130, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984839.

Anderson:1981:CSM

[1492] Dorothy A. Anderson. The circular structural model. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):131–141, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984840.

Munford:1981:FPP

[1493] A. G. Munford. A first passage problem in a random walk with a quality control application. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):142–146, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984841.

Hall:1981:NPE

[1494] Peter Hall. On the non-parametric estimation of mixture proportions. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):147–156, ???? 1981. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984842.

West:1981:RSA

[1495] M. West. Robust sequential approximate Bayesian estimation. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):157–166, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984843.

White:1981:MCV

[1496] L. V. White and W. J. Welch. A method for constructing valid restricted randomization schemes using the theory of D-optimal design of experiments. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):167–172, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984844.

Koutrouvelis:1981:GFT

[1497] I. A. Koutrouvelis and John Kellermeier. A goodness-of-fit test based on the empirical characteristic function when parameters must be estimated. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):

173–176, ???? 1981. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984845.

Cressie:1981:TJ

[1498] Noel Cressie. Transformations and the jackknife. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):177–182, ???? 1981. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984846.

Tso:1981:RRR

[1499] M. K.-S. Tso. Reduced-rank regression and canonical analysis. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):183–189, ???? 1981. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984847.

Fakinos:1981:GGQ

[1500] D. Fakinos. The G/G/1 queueing system with a particular queue discipline. Journal of the Royal Statistical Society. Series B (Methodological), 43(2): 190–196, ???? 1981. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984848.

Stewart:1981:ESL

[1501] I. Stewart and H. P. Wynn. The estimability structure of linear models and submodels. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):197–207, ???? 1981. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984849.

Smith:1981:RSC

[1502] A. F. M. Smith. On random sequences with centred spherical symmetry. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):208–209, ???? 1981. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984850.

Cook:1981:FDM

[1503] R. Dennis Cook and Mark E. Johnson.

A family of distributions for modelling non-elliptically symmetric multivariate data. Journal of the Royal Statistical Society. Series B (Methodological), 43 (2):210–218, ???? 1981. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984851.

Hosking:1981:LMT

[1504] J. R. M. Hosking. Lagrange-multiplier tests of multivariate time-series models. Journal of the Royal Statistical Society. Series B (Methodological), 43(2): 219–230, ???? 1981. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984852.

Li:1981:DRA

[1505] W. K. Li and A. I. McLeod. Distribution of the residual autocorrelations in multivariate ARMA time series models. Journal of the Royal Statistical Society. Series B (Methodological), 43(2): 231–239, ???? 1981. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984853.

Harvey:1981:NTG

[1506] A. C. Harvey and J. Tomenson. A note on testing for gaps in seasonal moving average models. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):240–243, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2984854. See corrigendum [1162].

Rao:1981:TBT

[1507] T. Subba Rao. On the theory of bilinear time series models. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):244–255, ???? 1981. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984855.

Smith:1981:MSM

[1508] J. Q. Smith. The multiparameter steady model. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):256–260, ???? 1981. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984856.

Hosking:1981:EFM

[1509] J. R. M. Hosking. Equivalent forms of the multivariate portmanteau statistic. Journal of the Royal Statistical Society. Series B (Methodological), 43(2): 261–262, ???? 1981. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2984857. See

corrigendum [1888].

King:1981:SSP

[1510] Maxwell L. King. A small sample property of the Cliff–Ord test for spatial correlation. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):263–264, ???? 1981. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2984858.

Anonymous:1981:BMb

[1511] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):??, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984859.

Anonymous:1981:FMb

[1512] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 43(2):??, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984838.

Bather:1981:RAT

[1513] J. A. Bather. Randomized allocation of treatments in sequential experiments. Journal of the Royal Statistical Society. Series B (Methodological), 43(3): 265–292, ???? 1981. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984938.

Agresti:1981:HSI

[1514] Alan Agresti. A hierarchical system of interaction measures for multidimensional contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 43(3):293–301, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984939.

Besag:1981:STD

[1515] Julian Besag. On a system of two-dimensional recurrence equations. Journal of the Royal Statistical Society. Series B (Methodological), 43(3): 302–309, ???? 1981. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984940.

Silvapulle:1981:EML

[1516] Mervyn J. Silvapulle. On the existence of maximum likelihood estimators for the binomial response models. Journal of the Royal Statistical Society. Series B (Methodological), 43(3):310–313, ???? 1981. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984941.

Freeman:1981:FRQ

[1517] G. H. Freeman. Further results on quasi-complete Latin squares. Journal of the Royal Statistical Society. Series B (Methodological), 43(3):314–320, ???? 1981. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984942.

Padmawar:1981:NCC

[1518] V. R. Padmawar. A note on the comparison of certain sampling strategies. Journal of the Royal Statistical Society. Series B (Methodological), 43(3): 321–326, ???? 1981. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984943.

Jensen:1981:IDA

[1519] D. R. Jensen and I. J. Good. Invariant distributions associated with matrix laws under structural symmetry. Journal of the Royal Statistical Society. Series B (Methodological), 43(3): 327–332, ???? 1981. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2984944.

Sweeting:1981:SPB

[1520] Trevor Sweeting. Scale parameters: A Bayesian treatment. Journal of the Royal Statistical Society. Series B (Methodological), 43(3):333–338, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984945.

Lefevre:1981:ALS

[1521] Cl. Lefevre. Another look at the symmetric m-dimensional simple epidemic process. Journal of the Royal Statistical Society. Series B (Methodological), 43 (3):339–341, ???? 1981. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2984946.

Patefield:1981:MLR

[1522] W. M. Patefield. Multivariate linear relationships: Maximum likelihood estimation and regression bounds. Journal of the Royal Statistical Society. Series B (Methodological), 43(3):342–352, ???? 1981. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984947.

Cumberland:1981:PMU

[1523] William G. Cumberland and Richard M.

Royall. Prediction models and unequal probability sampling. Journal of the Royal Statistical Society. Series B (Methodological), 43(3):353–367, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984948.

Lam:1981:NSR

[1524] K. Lam and K. L. Mehra. A note on some results of Tamhane. Journal of the Royal Statistical Society. Series B (Methodological), 43(3):368–369, ???? 1981. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984949.

Anonymous:1981:BMc

[1525] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 43(3):??, ????

1981. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984950.

Anonymous:1981:FMc

[1526] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 43(3):??, ???? 1981. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984937.

Anonymous:1981:VI

[1527] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 43(3): ??, ???? 1981. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984936.

Atkinson:1982:RDT

[1528] A. C. Atkinson. Regression diagnostics, transformations and constructed variables. Journal of the Royal Statistical Society. Series B (Methodological), 44 (1):1–36, ???? 1982. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2984705.

Hall:1982:SSE

[1529] Peter Hall. On some simple estimates of an exponent of regular variation. Journal of the Royal Statistical Society. Series B (Methodological), 44(1):37–42, ???? 1982. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984706.

Agarwal:1982:UPS

[1530] S. K. Agarwal, Pranesh Kumar, and A. Dey. On unequal probability sampling of two units without replacement. Journal of the Royal Statistical Society.

Series B (Methodological), 44(1):43–46, ???? 1982. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984707.

Veevers:1982:TVS

[1531] A. Veevers and M. Zafar-Yab. A twovariety square lattice design balanced under Besag’s coding scheme. Journal of the Royal Statistical Society. Series B (Methodological), 44(1):47–48, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984708.

Arnold:1982:ACS

[1532] Steven F. Arnold and James Lynch.

On Ali’s characterization of the spherical normal distribution. Journal of the Royal Statistical Society. Series B (Methodological), 44(1):49–51, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984709.

Anderson:1982:EAD

[1533] T. W. Anderson and Takamitsu Sawa.

Exact and approximate distributions of the maximum likelihood estimator of a slope coefficient. Journal of the Royal Statistical Society. Series B (Methodological), 44(1):52–62, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984710.

Bailey:1982:DTD

[1534] R. A. Bailey. The decomposition of treatment degrees of freedom in quantitative factorial experiments. Journal of the Royal Statistical Society. Series B (Methodological), 44(1):63–70, ???? 1982. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984711. See corrigenda

[1605].

Kent:1982:FBD

[1535] John T. Kent. The Fisher–Bingham distribution on the sphere. Journal of the Royal Statistical Society. Series B (Methodological), 44(1):71–80, ???? 1982. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984712.

Rivest:1982:SSM

[1536] Louis-Paul Rivest. Some statistical methods for bivariate circular data. Journal of the Royal Statistical Society. Series B (Methodological), 44(1):81–90, ???? 1982. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2984713.

Robinson:1982:SAP

[1537] J. Robinson. Saddlepoint approximations for permutation tests and confidence intervals. Journal of the Royal Statistical Society. Series B (Methodological), 44(1):91–101, ???? 1982. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984714.

Ogata:1982:LIM

[1538] Yosihiko Ogata and Hirotugu Akaike. On linear intensity models for mixed doubly stochastic Poisson and self- exciting point processes. Journal of the Royal Statistical Society. Series B (Methodological), 44(1):102–107, ???? 1982. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2984715.

Anonymous:1982:BMa

[1539] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 44(1):108, ???? 1982. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2984716.

Anonymous:1982:FMa

[1540] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 44(1):??, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2984704.

Kingman:1982:TS

[1541] J. F. C. Kingman. The thrown string.

Journal of the Royal Statistical Society. Series B (Methodological), 44(2): 109–138, ???? 1982. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345820. See

corrigenda [1606].

Aitchison:1982:SAC

[1542] J. Aitchison. The statistical analysis of compositional data. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):139–177, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345821.

Parr:1982:MDE

[1543] William C. Parr and William R. Schucany. Minimum distance estimation and components of goodness-of-fit statistics. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):178–189, ???? 1982. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345822.

Laird:1982:APD

[1544] Nan M. Laird and Thomas A. Louis. Approximate posterior distributions for incomplete data problems. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):190–200, ???? 1982. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345823.

Wei:1982:ADF

[1545] L. J. Wei. Asymptotically distributionfree simultaneous confidence region of treatment differences in a randomized complete block design. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):201–208, ???? 1982. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345824.

Panaretos:1982:SPF

[1546] John Panaretos. On a structural property of finite distributions. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):209–211, ???? 1982. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345825.

Rigby:1982:CIP

[1547] R. A. Rigby. A credibility interval for the probability that a new observation belongs to one of two multivariate normal populations. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):212–220, ???? 1982. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345826.

John:1982:COB

[1548] J. A. John and E. R. Williams. Conjectures for optimal block designs. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):221–225, ???? 1982. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345827.

Louis:1982:FOI

[1549] Thomas A. Louis. Finding the observed information matrix when using the EM algorithm. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):226–233, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345828.

Pettitt:1982:ILM

[1550] A. N. Pettitt. Inference for the linear model using a likelihood based on ranks. Journal of the Royal Statistical Society. Series B (Methodological), 44 (2):234–243, ???? 1982. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345829.

Hougaard:1982:PNL

[1551] Philip Hougaard. Parametrizations of non-linear models. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):244–252, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345830.

Jacroux:1982:SOD

[1552] Mike Jacroux. Some E-optimal designs for the one-way and two-way elimination of heterogeneity. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):253–261, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345831.

Basawa:1982:NPE

[1553] I. V. Basawa and P. J. Brockwell. Non-parametric estimation for nondecreasing Levy processes. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):262–269, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345832.

Eplett:1982:TMW

[1554] W. J. R. Eplett. Two Mann– Whitney type rank tests. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):270–286, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345833.

Anonymous:1982:BMb

[1555] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):??, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345834.

Anonymous:1982:FMb

[1556] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 44(2):??, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345819.

Brown:1982:MC

[1557] P. J. Brown. Multivariate calibration. Journal of the Royal Statistical Society. Series B (Methodological), 44(3):

287–321, ???? 1982. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345489.

Shafer:1982:BFP

[1558] Glenn Shafer. Belief functions and parametric models. Journal of the Royal Statistical Society. Series B (Methodological), 44(3):322–352, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345490.

OReilly:1982:CGF

[1559] Federico J. O’Reilly and Michael A.

Stephens. Characterizations and goodness of fit tests. Journal of the Royal Statistical Society. Series B (Methodological), 44(3):353–360, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345491.

Eplett:1982:DST

[1560] W. J. R. Eplett. The distributions of Smirnov type two-sample rank tests for discontinuous distribution functions. Journal of the Royal Statistical Society. Series B (Methodological), 44(3): 361–369, ???? 1982. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345492. See

corrigenda [1604].

Cook:1982:NAO

[1561] R. D. Cook, N. Holschuh, and S. Weisberg. A note on an alternative outlier model. Journal of the Royal Statistical Society. Series B (Methodological), 44 (3):370–376, ???? 1982. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345493.

Harvey:1982:NTG

[1562] A. C. Harvey and J. Tomenson. A note on testing for gaps in seasonal moving average models: Correction. Journal of the Royal Statistical Society. Series B (Methodological), 44(3):376, ???? 1982. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345494. See [1506].

Spiegelhalter:1982:BFL

[1563] D. J. Spiegelhalter and A. F. M. Smith. Bayes factors for linear and log-linear models with vague prior information. Journal of the Royal Statistical Society. Series B (Methodological), 44(3): 377–387, ???? 1982. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345495.

Binder:1982:NPB

[1564] David A. Binder. Non-parametric Bayesian models for samples from finite populations. Journal of the Royal Statistical Society. Series B (Methodological), 44(3):388–393, ???? 1982. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345496.

Gaffke:1982:EOD

[1565] N. Gaffke and O. Krafft. Exact Doptimum designs for quadratic regression. Journal of the Royal Statistical Society. Series B (Methodological), 44 (3):394–397, ???? 1982. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345497.

Milne:1982:TBI

[1566] Claire Milne. Transient behaviour of the interrupted Poisson process. Journal of the Royal Statistical Society. Series B (Methodological), 44(3):398–405, ???? 1982. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345498.

Lotwick:1982:MAS

[1567] H. W. Lotwick and B. W. Silverman.

Methods for analysing spatial processes of several types of points. Journal of the Royal Statistical Society. Series B (Methodological), 44(3):406–413, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345499.

Oakes:1982:MAB

[1568] David Oakes. A model for association in bivariate survival data. Journal of the Royal Statistical Society. Series B (Methodological), 44(3):414–422, ???? 1982. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345500.

Anonymous:1982:BMc

[1569] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 44(3):??, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345501.

Anonymous:1982:FMc

[1570] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 44(3):??, ???? 1982. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345488.

Anonymous:1982:VI

[1571] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Methodological), 44(3): ??, ???? 1982. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345487.

Jarrett:1983:DPC

[1572] R. G. Jarrett. Definitions and properties for m-concurrence designs. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):1–10, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345616.

Diggle:1983:BCP

[1573] Peter J. Diggle and Robin K. Milne. Bivariate Cox processes: Some models for bivariate spatial point patterns. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):11–21, ???? 1983. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345617.

Dansie:1983:NPP

[1574] B. R. Dansie. A note on permutation probabilities. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):22–24, ???? 1983. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2345618.

Brown:1983:SUS

[1575] B. M. Brown. Statistical uses of the spatial median. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):25–30, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345619.

Constantine:1983:TEC

[1576] Gregory M. Constantine. On the trace efficiency for control of reinforced balanced incomplete block designs. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):31–36, ???? 1983. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345620.

Titterington:1983:MDN

[1577] D. M. Titterington. Minimum distance non-parametric estimation of mixture proportions. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):37–46, ???? 1983. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345621.

Boyles:1983:CEA

[1578] Russell A. Boyles. On the convergence of the EM algorithm. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):47–50, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345622.

Dempster:1983:RER

[1579] Arthur P. Dempster and Donald B. Rubin. Rounding error in regression: The appropriateness of Sheppard’s corrections. Journal of the Royal Statistical Society. Series B (Methodological), 45 (1):51–59, ???? 1983. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345623.

Albert:1983:ECT

[1580] James H. Albert and Arjun K. Gupta.

Estimation in contingency tables using prior information. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):60–69, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345624.

Jacroux:1983:SMV

[1581] Mike Jacroux. Some minimum variance block designs for estimating treatment differences. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):70–76, ???? 1983. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2345625.

Tanaka:1983:OSL

[1582] Katsuto Tanaka. The one-sided Lagrange multiplier test of the AR (p) model vs the AR (p) model with measurement error. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):77–80, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345626.

Hall:1983:OSD

[1583] Peter Hall. Orthogonal series distribution function estimation, with applications. Journal of the Royal Statistical Society. Series B (Methodological), 45 (1):81–88, ???? 1983. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345627.

Sawyer:1983:TSF

[1584] K. R. Sawyer. Testing separate familes of hypotheses: An information criterion. Journal of the Royal Statistical Society. Series B (Methodological), 45 (1):89–99, ???? 1983. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345628.

Tavare:1983:SDC

[1585] Simon Tavare. Serial dependence in contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):100–106, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345629.

Kabaila:1983:ETS

[1586] Paul Kabaila. On estimating time series parameters using sample autocorrelations. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):107–119, ???? 1983. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345630.

Butler:1983:ODT

[1587] R. Butler. Outlier discordancy test in the normal linear model. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):120–132, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345631.

Wahba:1983:BCI

[1588] Grace Wahba. Bayesian “Confidence intervals” for the cross-validated smoothing spline. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):133–150, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345632.

Anonymous:1983:BMa

[1589] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):??, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345633.

Anonymous:1983:FMa

[1590] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 45(1):??, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345615.

Wilkinson:1983:NNN

[1591] G. N. Wilkinson, S. R. Eckert, T. W. Hancock, and O. Mayo. Nearest neighbour (NN) analysis of field experiments. Journal of the Royal Statistical Society. Series B (Methodological), 45 (2):151–211, ???? 1983. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345523.

Rosenbaum:1983:ASU

[1592] P. R. Rosenbaum and D. B. Rubin. Assessing sensitivity to an unobserved binary covariate in an observational study with binary outcome. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):212–218, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345524.

Hall:1983:SES

[1593] Peter Hall. Sequential estimation saving sampling operations. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):219–223, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345525.

Lewis:1983:SRD

[1594] S. M. Lewis, A. M. Dean, and P. H. Lewis. Single replicate designs for two-factor experiments. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):224–227, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345526.

Aase:1983:REN

[1595] Knut K. Aase. Recursive estimation in non-linear time series models of autoregressive type. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):228–237, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345527.

Kendall:1983:ECI

[1596] W. S. Kendall and I. W. Saunders. Epidemics in competition II: The general epidemic. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):238–244, ???? 1983. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345528.

John:1983:FEG

[1597] J. A. John and S. M. Lewis. Factorial experiments in generalized cyclic row-column designs. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):245–251, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345529.

Dean:1983:UBA

[1598] A. M. Dean and S. M. Lewis. Upper bounds for average efficiency factors of two-factor interactions. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):252–257, ???? 1983. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345530.

Titterington:1983:KBD

[1599] D. M. Titterington and G. M. Mill. Kernel-based density estimates from incomplete data. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):258–266, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345531.

Preece:1983:SLS

[1600] D. A. Preece and G. H. Freeman. Semi-Latin squares and related designs. Journal of the Royal Statistical Society. Series B (Methodological), 45(2): 267–277, ???? 1983. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345532.

Dalal:1983:APM

[1601] S. R. Dalal and W. J. Hall. Approximating priors by mixtures of natural conjugate priors. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):278–286, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345533.

Pettitt:1983:LBI

[1602] A. N. Pettitt. Likelihood based inference using signed ranks for matched pairs. Journal of the Royal Statistical Society. Series B (Methodological), 45 (2):287–296, ???? 1983. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345534.

Gupta:1983:SNM

[1603] S. C. Gupta. Some new methods for constructing block designs having orthogonal factorial structure. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):297–307, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345535.

Eplett:1983:CDS

[1604] W. J. R. Eplett. Corrigenda: The distributions of Smirnov-type two-sample rank tests for discontinuous distribution functions. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):308–309, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345536. See [1560].

Bailey:1983:CDT

[1605] R. A. Bailey. Corrigenda: Decomposition of treatment degrees of freedom in quantitative factorial experiments. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):309, ???? 1983. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345537. See [1534].

Raftery:1983:CCD

[1606] A. E. Raftery. Corrigenda: Comment in the discussion of Professor Kingman’s paper “The Thrown String”. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):309, ???? 1983. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345538. See [1541].

Raviv:1983:CNP

[1607] Alona Raviv. Corrigenda: A nonparametric test for comparing two nonindependent distributions. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):309, ????

1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345539. See [1360].

Anonymous:1983:BMb

[1608] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):??, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345540.

Anonymous:1983:FMb

[1609] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 45(2):??, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345522.

Copas:1983:RPS

[1610] J. B. Copas. Regression, prediction and shrinkage. Journal of the Royal Statistical Society. Series B (Methodological), 45(3):311–354, ???? 1983. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345402.

Goldstein:1983:ORD

[1611] Michael Goldstein. Outlier resistant distribution: Where does the probability go? Journal of the Royal Statistical Society. Series B (Methodological), 45 (3):355–357, ???? 1983. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345403.

Koziol:1983:AMN

[1612] James A. Koziol. On assessing multivariate normality. Journal of the Royal Statistical Society. Series B (Methodological), 45(3):358–361, ????

1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345404.

Wong:1983:NNC

[1613] M. Anthony Wong and Tom Lane. A k th nearest neighbour clustering procedure. Journal of the Royal Statistical Society. Series B (Methodological), 45 (3):362–368, ???? 1983. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345405.

Liu:1983:GCT

[1614] Tzen-Ping Liu. A general completeness theorem in sampling theory. Journal of the Royal Statistical Society. Series B (Methodological), 45(3):369–372, ???? 1983. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345406.

Turek:1983:ETG

[1615] Richard J. Turek and Ronald C. Suich.

An exact test on the Goodman–Kruskal λ for prediction on a dichotomy. Journal of the Royal Statistical Society. Series B (Methodological), 45(3):373–379, ???? 1983. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345407.

Mak:1983:SGL

[1616] Tak K. Mak. On Sprent’s generalized least-squares estimator. Journal of the Royal Statistical Society. Series B (Methodological), 45(3):380–383, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345408.

Basu:1983:UWD

[1617] A. P. Basu, J. K. Ghosh, and P. K.

Sen. A unified way of deriving LMP rank tests from censored data. Journal of the Royal Statistical Society. Series B (Methodological), 45(3):384–390, ???? 1983. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345409.

Cox:1983:RCS

[1618] D. R. Cox. A remark on censoring and surrogate response variables. Journal of the Royal Statistical Society. Series B (Methodological), 45(3):391–393, ???? 1983. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345410.

Cheng:1983:EPC

[1619] R. C. H. Cheng and N. A. K. Amin. Estimating parameters in continuous univariate distributions with a shifted origin. Journal of the Royal Statistical Society. Series B (Methodological), 45 (3):394–403, ???? 1983. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345411. See

comment [1693].

Cordeiro:1983:ILR

[1620] Gauss M. Cordeiro. Improved likelihood ratio statistics for generalized linear models. Journal of the Royal Statistical Society. Series B (Methodological), 45(3):404–413, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345412.

Anonymous:1983:BMc

[1621] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 45(3):??, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345413.

Anonymous:1983:FMc

[1622] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 45(3):??, ???? 1983. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345401.

Anonymous:1983:VI

[1623] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Methodological), 45(3): ??, ???? 1983. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345400.

Anderson:1984:ROC

[1624] J. A. Anderson. Regression and ordered categorical variables. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):1–30, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345457.

Cheng:1984:UMC

[1625] Ching-Shui Cheng, G. M. Constantine, and A. S. Hedayat. A unified method for constructing PBIB designs based on triangular and L_{2} schemes. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):31–37, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345458.

Pfeffermann:1984:NLS

[1626] D. Pfeffermann. Note on large sample properties of balanced samples. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):38–41, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345459.

Hardle:1984:RNP

[1627] W. Hardle and T. Gasser. Robust nonparametric function fitting. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):42–51, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345460.

Ferreri:1984:HBP

[1628] Carlo Ferreri. On the hypergeometric birth process and some implications about the gamma distribution representation. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):52–57, ???? 1984. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345461.

Tanaka:1984:AEA

[1629] Katsuto Tanaka. An asymptotic expansion associated with the maximum likelihood estimators in ARMA models. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):58–67, ???? 1984. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345462.

Papaioannou:1984:IRC

[1630] Takis Papaioannou and Sotiris Loukas. Inequalities on rank correlation with missing data. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):68–71, ???? 1984. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345463.

Mardia:1984:GFT

[1631] K. V. Mardia, D. Holmes, and J. T.

Kent. A goodness-of-fit test for the von Mises–Fisher distribution. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):72–78, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345464.

Khan:1984:CSP

[1632] Rasul A. Khan. On cumulative sum procedures and the SPRT with applications. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):79–85, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345465.

Kass:1984:CPZ

[1633] Robert E. Kass. Canonical parameterizations and zero parameter-effects curvature. Journal of the Royal Statistical Society. Series B (Methodological), 46 (1):86–92, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345466.

Miles:1984:SSA

[1634] Roger E. Miles. Symmetric sequential analysis: The efficiencies of sports scoring systems (with particular reference to those of tennis). Journal of the Royal Statistical Society. Series B (Methodological), 46(1):93–108, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345467.

Reinsel:1984:NCP

[1635] Gregory Reinsel. A note on conditional prediction in the multivariate linear model. Journal of the Royal Statistical Society. Series B (Methodological), 46 (1):109–117, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345468.

Altham:1984:IPE

[1636] P. M. E. Altham. Improving the precision of estimation by fitting a model. Journal of the Royal Statistical Society. Series B (Methodological), 46(1): 118–119, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345469.

Bartholomew:1984:SBD

[1637] D. J. Bartholomew. Scaling binary data using a factor model. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):120–123, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345470.

Mouchart:1984:NLS

[1638] M. Mouchart and L. Simar. A note on least-squares approximation in the Bayesian analysis of regression models. Journal of the Royal Statistical Society. Series B (Methodological), 46(1): 124–133, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345471.

Hall:1984:OPK

[1639] Peter Hall. An optimal property of kernel estimators of a probability density. Journal of the Royal Statistical Society. Series B (Methodological), 46(1): 134–138, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345472.

Pfeffermann:1984:EGM

[1640] D. Pfeffermann. On extensions of the Gauss–Markov theorem to the case of stochastic regression coefficients. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):139–148, ???? 1984. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345473.

Anonymous:1984:BMa

[1641] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):??, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345474.

Anonymous:1984:FMa

[1642] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 46(1):??, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345456.

Green:1984:IRL

[1643] P. J. Green. Iteratively reweighted least squares for maximum likelihood estimation, and some robust and resistant alternatives. Journal of the Royal Statistical Society. Series B (Methodological), 46(2):149–192, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345503.

Diggle:1984:MCM

[1644] Peter J. Diggle and Richard J. Gratton. Monte Carlo methods of inference for implicit statistical models. Journal of the Royal Statistical Society. Series B (Methodological), 46(2):193–227, ???? 1984. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345504.

Rao:1984:ARG

[1645] J. S. Rao and Morgan Kuo. Asymptotic results on the Greenwood statistic and some of its generalizations. Journal of the Royal Statistical Society. Series B (Methodological), 46(2):228–237, ???? 1984. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345505.

Dufour:1984:CPC

[1646] Robert Dufour, Urs R. Maag, and Constance van Eeden. Correcting a proof of a characterization of the exponential distribution. Journal of the Royal Statistical Society. Series B (Methodological), 46(2):238–241, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345506.

Sackrowitz:1984:EMS

[1647] Harold Sackrowitz and Ester Samuel-Cahn. Estimation of the mean of a selected negative exponential population. Journal of the Royal Statistical Society. Series B (Methodological), 46 (2):242–249, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345507.

McCullagh:1984:ENP

[1648] Peter McCullagh. On the elimination of nuisance parameters in the proportional odds model. Journal of the Royal Statistical Society. Series B (Methodological), 46(2):250–256, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345508.

Titterington:1984:RPE

[1649] D. M. Titterington. Recursive parameter estimation using incomplete data. Journal of the Royal Statistical Society. Series B (Methodological), 46(2): 257–267, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345509.

Bellhouse:1984:ARE

[1650] D. R. Bellhouse and V. M. Joshi. On the admissibility of the regression estimator. Journal of the Royal Statistical Society. Series B (Methodological), 46 (2):268–269, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345510.

Gross:1984:NCS

[1651] W. F. Gross. A note on “Chi-Squared tests with survey data”. Journal of the Royal Statistical Society. Series B (Methodological), 46(2):270–272, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345511.

Lewis:1984:UBF

[1652] S. M. Lewis and A. M. Dean. Upper bounds for factorial efficiency factors. Journal of the Royal Statistical Society. Series B (Methodological), 46(2): 273–278, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345512.

Dean:1984:CUB

[1653] A. M. Dean and S. M. Lewis. A comparison of upper bounds for efficiency factors of block desings. Journal of the Royal Statistical Society. Series B (Methodological), 46(2):279–283, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345513.

Price:1984:TCE

[1654] J. Michael Price. A test for comparing estimators under the generalized mean square error criterion. Journal of the Royal Statistical Society. Series B (Methodological), 46(2):284–295, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345514.

SanMartini:1984:PMS

[1655] A. San Martini and F. Spezzaferri. A predictive model selection criterion. Journal of the Royal Statistical Society. Series B (Methodological), 46(2): 296–303, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345515.

Poskitt:1984:TMV

[1656] D. S. Poskitt and A. R. Tremayne. Testing misspecification in vector time series models with exogenous variables. Journal of the Royal Statistical Society. Series B (Methodological), 46(2): 304–315, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345516.

Jacroux:1984:OUR

[1657] Mike Jacroux. On the optimality and usage of reinforced block designs for comparing test treatments with a standard treatment. Journal of the Royal Statistical Society. Series B (Methodological), 46(2):316–322, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345517.

Bailey:1984:QCL

[1658] R. A. Bailey. Quasi-complete Latin squares: Construction and randomization. Journal of the Royal Statistical Society. Series B (Methodological), 46 (2):323–334, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345518.

Azzalini:1984:TNT

[1659] A. Azzalini and D. R. Cox. Two new tests associated with analysis of variance. Journal of the Royal Statistical Society. Series B (Methodological), 46 (2):335–343, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345519.

McCullagh:1984:SMR

[1660] Peter McCullagh and Peter Lang.

Stochastic models for rock instability in tunnels. Journal of the Royal Statistical Society. Series B (Methodological), 46 (2):344–352, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345520.

Anonymous:1984:BMb

[1661] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 46(2):??, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345521.

Anonymous:1984:FMb

[1662] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 46(2):??, ????

1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345502.

Davis:1984:PDM

[1663] M. H. A. Davis. Piecewisedeterministic Markov processes: A general class of non-diffusion stochastic models. Journal of the Royal Statistical Society. Series B (Methodological), 46 (3):353–388, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345677.

Withers:1984:AED

[1664] C. S. Withers. Asymptotic expansions for distributions and quantiles with power series cumulants. Journal of the Royal Statistical Society. Series B (Methodological), 46(3):389–396, ???? 1984. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345678. See corrections

[1772].

Barbour:1984:PCD

[1665] A. D. Barbour and G. K. Eagleson. Poisson convergence for dissociated statistics. Journal of the Royal Statistical Society. Series B (Methodological), 46(3):397–402, ???? 1984. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345679.

Genest:1984:CBT

[1666] Christian Genest. A conflict between two axioms for combining subjective distributions. Journal of the Royal Statistical Society. Series B (Methodological), 46(3):403–405, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345680.

Draper:1984:SR

[1667] Norman R. Draper. Schlaflian rotatability. Journal of the Royal Statistical Society. Series B (Methodological), 46 (3):406–411, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345681.

Manjunath:1984:OSE

[1668] S. M. Manjunath. Optimal sequential estimation for ergodic birth-death processes. Journal of the Royal Statistical Society. Series B (Methodological), 46 (3):412–418, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345682.

Sawyer:1984:MHT

[1669] K. R. Sawyer. Multiple hypothesis testing. Journal of the Royal Statistical Society. Series B (Methodological), 46 (3):419–424, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345683.

Bar-Lev:1984:ABC

[1670] Shaul K. Bar-Lev. Asymptotic behaviour of conditional maximum likelihood estimators in a certain exponential model. Journal of the Royal Statistical Society. Series B (Methodological), 46(3):425–430, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345684.

West:1984:OMP

[1671] Mike West. Outlier models and prior distributions in Bayesian linear regression. Journal of the Royal Statistical Society. Series B (Methodological), 46 (3):431–439, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345685.

Cressie:1984:MGF

[1672] Noel Cressie and Timothy R. C. Read. Multinomial goodness-of-fit tests. Journal of the Royal Statistical Society. Series B (Methodological), 46 (3):440–464, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345686.

Hall:1984:ENE

[1673] Peter Hall and D. M. Titterington. Efficient nonparametric estimation of mixture proportions. Journal of the Royal Statistical Society. Series B (Methodological), 46(3):465–473, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345687.

Ghosh:1984:NBA

[1674] Malay Ghosh and Glen Meeden. A new Bayesian analysis of a random effects model. Journal of the Royal Statistical Society. Series B (Methodological), 46 (3):474–482, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345688.

Barndorff-Nielsen:1984:BAL

[1675] O. E. Barndorff-Nielsen and D. R. Cox. Bartlett adjustments to the likelihood ratio statistic and the distribution of the maximum likelihood estimator. Journal of the Royal Statistical Society. Series B (Methodological), 46 (3):483–495, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345689.

Ogata:1984:LAS

[1676] Yosihiko Ogata and Masaharu Tanemura. Likelihood analysis of spatial point patterns. Journal of the Royal Statistical Society. Series B (Methodological), 46(3):496–518, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345690.

Smith:1984:STA

[1677] Richard L. Smith. Sequential treatment allocation using biased coin designs. Journal of the Royal Statistical Society. Series B (Methodological), 46 (3):519–543, ???? 1984. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345691.

Oman:1984:DEB

[1678] Samuel D. Oman. A different empirical Bayes interpretation of ridge and Stein estimators. Journal of the Royal Statistical Society. Series B (Methodological), 46(3):544–557, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345692.

Krzanowski:1984:SPC

[1679] W. J. Krzanowski. Sensitivity of principal components. Journal of the Royal Statistical Society. Series B (Methodological), 46(3):558–563, ???? 1984. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345693.

Nigam:1984:SMI

[1680] A. K. Nigam, Pranesh Kumar, and V. K. Gupta. Some methods of inclusion probability proportional to size sampling. Journal of the Royal Statistical Society. Series B (Methodological), 46(3):564–571, ???? 1984. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2345694.

Anonymous:1984:BMc

[1681] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 46(3):??, ???? 1984. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345695.

Anonymous:1984:FMc

[1682] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 46(3):??, ???? 1984. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345676.

Anonymous:1984:VI

[1683] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 46(3): ??, ???? 1984. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345675.

Silverman:1985:SAS

[1684] B. W. Silverman. Some aspects of the spline smoothing approach to non-parametric regression curve fitting. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):1–52, ???? 1985. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345542.

Thompson:1985:NRM

[1685] R. Thompson. A note on restricted maximum likelihood estimation with an alternative outlier model. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):53–55, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345543.

Parr:1985:JDS

[1686] William C. Parr. Jackknifing differentiable statistical functionals. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):56–66, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345544.

Aitkin:1985:EHT

[1687] Murray Aitkin and Donald B. Rubin. Estimation and hypothesis testing in finite mixture models. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):67–75, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345545.

Gilula:1985:AHA

[1688] Zvi Gilula. On the analysis of heterogeneity among populations. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):76–83, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345546.

Bofinger:1985:MPC

[1689] Eve Bofinger. Monotonicity of the probability of correct selection or are bigger samples better? Journal of the Royal Statistical Society. Series B (Methodological), 47(1):84–89, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345547.

Hayre:1985:GSS

[1690] Lakhbir S. Hayre. Group sequential sampling with variable group sizes. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):90–97, ???? 1985. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345548.

King:1985:LBI

[1691] Maxwell L. King and Grant H. Hillier. Locally best invariant tests of the error covariance matrix of the linear regression model. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):98–102, ???? 1985. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2345549.

Hougaard:1985:AAD

[1692] P. Hougaard. The appropriateness of the asymptotic distribution in a nonlinear regression model in relation to curvature. Journal of the Royal Statistical Society. Series B (Methodological), 47 (1):103–114, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345550.

Titterington:1985:CEP

[1693] D. M. Titterington. Comment on “Estimating Parameters in Continuous Univariate Distributions”. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):115–116, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345551. See [1619].

Fletcher:1985:CDF

[1694] D. J. Fletcher and J. A. John. Changeover designs and factorial structure. Journal of the Royal Statistical Society. Series B (Methodological), 47 (1):117–124, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345552.

Hall:1985:TMT

[1695] Peter Hall. Tailor-made tests of goodness of fit. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):125–131, ???? 1985. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345553.

Raghavarao:1985:EBP

[1696] D. Raghavarao and E. A. Blaisdell, Jr. Efficiency bounds for partially balanced change-over designs based on Massociate class PBIB designs. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):132–135, ???? 1985. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345554.

Aitchison:1985:GCD

[1697] J. Aitchison. A general class of distributions on the simplex. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):136–146, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345555.

Wu:1985:VEC

[1698] C. F. Jeff Wu. Variance estimation for the combined ratio and combined regression estimators. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):147–154, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345556.

Hall:1985:UUD

[1699] Peter Hall and D. M. Titterington. The use of uncategorized data to improve the performance of a nonparametric estimator of a mixture density. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):155–163, ???? 1985. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345557.

Anonymous:1985:BMa

[1700] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):??, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345558.

Anonymous:1985:FMa

[1701] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 47(1):??, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345541.

Lawrance:1985:MRA

[1702] A. J. Lawrance and P. A. W. Lewis. Modelling and residual analysis of nonlinear autoregressive time series in exponential variables. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):165–202, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345560.

Anderson:1985:VCM

[1703] Dorothy A. Anderson and Murray Aitkin. Variance component models with binary response: Interviewer variability. Journal of the Royal Statistical Society. Series B (Methodological), 47 (2):203–210, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345561.

Isogawa:1985:EML

[1704] Yoshiko Isogawa. Estimating a multivariate linear structural relationship with replication. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):211–215, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345562.

Paulsen:1985:ERV

[1705] Jostein Paulsen and Dag Tjostheim.

On the estimation of residual variance and order in autoregressive time series. Journal of the Royal Statistical Society. Series B (Methodological), 47(2): 216–228, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345563.

Crowder:1985:GEC

[1706] Martin Crowder. Gaussian estimation for correlated binomial data. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):229–237, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345564.

Gasser:1985:KNC

[1707] T. Gasser, H-G. Muller, and V. Mammitzsch. Kernels for nonparametric curve estimation. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):238–252, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345565.

Pettitt:1985:RWL

[1708] A. N. Pettitt. Re-weighted least squares estimation with censored and grouped data: An application of the EM algorithm. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):253–260, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345566.

Lewis:1985:NEC

[1709] S. M. Lewis and A. M. Dean. A note on efficiency-consistent designs. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):261–262, ???? 1985. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345567.

Eaves:1985:MMI

[1710] D. M. Eaves. On maximizing missing information about a hypothesis. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):263–266, ???? 1985. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345568.

Siegel:1985:MDC

[1711] Andrew F. Siegel. Modelling data containing exact zeroes using zero degrees of freedom. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):267–271, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345569.

Snyder:1985:RED

[1712] R. D. Snyder. Recursive estimation of dynamic linear models. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):272–276, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345570.

Langron:1985:PTG

[1713] S. P. Langron and A. J. Collins.

Perturbation theory for generalized Procrustes analysis. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):277–284, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345571.

Smith:1985:MLE

[1714] Richard L. Smith and Ishay Weissman.

Maximum likelihood estimation of the lower tail of a probability distribution. Journal of the Royal Statistical Society. Series B (Methodological), 47(2): 285–298, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345572.

Green:1985:AFE

[1715] Peter Green, Christopher Jennison, and Allan Seheult. Analysis of field experiments by least squares smoothing. Journal of the Royal Statistical Society. Series B (Methodological), 47 (2):299–315, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345573.

Salia:1985:PML

[1716] R. N. Salia and M. A. Shashiashvili.

On Palm’s model with a large number of machines and repairmen. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):316–322, ???? 1985. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345574.

Brown:1985:MLT

[1717] B. M. Brown. Multiparameter linearization theorems. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):323–331, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345575.

Eubank:1985:DSS

[1718] R. L. Eubank. Diagnostics for smoothing splines. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):332–341, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345576.

Fisher:1985:SM

[1719] N. I. Fisher. Spherical medians. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):342–348, ???? 1985. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345577.

Jones:1985:UBB

[1720] Byron Jones. Using bricks to build block designs. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):349–356, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345578.

Speed:1985:NAR

[1721] T. P. Speed, E. R. Williams, and H. D. Patterson. A note on the analysis of resolvable block designs. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):357–361, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345579.

Wooff:1985:BRM

[1722] David A. Wooff. Bounds on reciprocal moments with applications and developments in Stein estimation and post-stratification. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):362–371, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345580.

Oja:1985:APG

[1723] H. Oja and A. Niinimaa. Asymptotic properties of the generalized median in the case of multivariate normality. Journal of the Royal Statistical Society. Series B (Methodological), 47(2): 372–377, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345581.

Anonymous:1985:BMb

[1724] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):??, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345582.

Anonymous:1985:FMb

[1725] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 47(2):??, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345559.

Kelly:1985:SMC

[1726] F. P. Kelly. Stochastic models of computer communication systems. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):379–395, ???? 1985. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345773.

Mitrani:1985:RTP

[1727] I. Mitrani. Response time problems in communication networks. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):396–406, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345774.

Whittle:1985:SCP

[1728] P. Whittle. Scheduling and characterization problems for stochastic networks. Journal of the Royal Statistical Society. Series B (Methodological), 47 (3):407–428, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345775.

Robinson:1985:TSC

[1729] P. M. Robinson. Testing for serial correlation in regression with missing observations. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):429–437, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345776.

Constantine:1985:POU

[1730] Gregory M. Constantine. Preservation of optimality under loss of varieties. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):

438–446, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345777.

Crowder:1985:DMR

[1731] Martin Crowder. A distributional model for repeated failure time measurements. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):447–452, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345778.

Bailey:1985:BOE

[1732] R. A. Bailey. Balance, orthogonality and efficiency factors in factorial design. Journal of the Royal Statistical Society. Series B (Methodological), 47 (3):453–458, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345779.

Gleser:1985:EPD

[1733] Leon J. Gleser and David S. Moore.

The effect of positive dependence on chi-squared tests for categorical data. Journal of the Royal Statistical Society. Series B (Methodological), 47(3): 459–465, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345780.

Bather:1985:MRT

[1734] John Bather and Gordon Simons.

The minimax risk for two-stage procedures in clinical trials. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):466–475, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345781.

Letac:1985:ECS

[1735] G. Letac, V. Seshadri, and G. A. Whitmore. An exact chi-squared decomposition theorem for inverse Gaussian variates. Journal of the Royal Statistical Society. Series B (Methodological), 47 (3):476–481, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345782.

Fisher:1985:NSS

[1736] Nicholas I. Fisher and Toby Lewis. A note on spherical splines. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):482–488, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345783.

Williams:1985:CCO

[1737] E. R. Williams. A criterion for the construction of optimal neighbour designs. Journal of the Royal Statistical Society. Series B (Methodological), 47 (3):489–497, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345784.

Vlachonikolis:1985:ADL

[1738] Ioannis G. Vlachonikolis. On the asymptotic distribution of the location linear discriminant function. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):498–509, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345785.

Chan:1985:PFR

[1739] Lai K. Chan and Tak K. Mak. On the polynomial functional relationship. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):

510–518, ???? 1985. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345786.

Sedransk:1985:BEF

[1740] J. Sedransk, J. Monahan, and H. Y.

Chiu. Bayesian estimation of finite population parameters in categorical data models incorporating order restrictions. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):519–527, ???? 1985. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345787.

Raftery:1985:MHO

[1741] Adrian E. Raftery. A model for high-order Markov chains. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):528–539, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345788.

Chen:1985:ANL

[1742] Chan-Fu Chen. On asymptotic normality of limiting density functions with Bayesian implications. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):540–546, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345789.

Peers:1985:AEC

[1743] H. W. Peers and Mohammad Iqbal.

Asymptotic expansions for confidence limits in the presence of nuisance parameters, with applications. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):547–554, ???? 1985. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345790.

Anonymous:1985:BMc

[1744] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):??, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345791.

Anonymous:1985:FMc

[1745] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 47(3):??, ???? 1985. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345772.

Anonymous:1985:VI

[1746] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 47(3): ??, ???? 1985. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345771.

Butler:1986:PLI

[1747] Ronald W. Butler. Predictive likelihood inference with applications. Journal of the Royal Statistical Society. Series B (Methodological), 48(1):1–38, ???? 1986. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345635.

Muirhead:1986:DOT

[1748] C. R. Muirhead. Distinguishing outlier types in time series. Journal of the Royal Statistical Society. Series B (Methodological), 48(1):39–47, ???? 1986. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345636.

Dale:1986:ANG

[1749] Jocelyn R. Dale. Asymptotic normality of goodness-of-fit statistics for sparse product multinomials. Journal of the Royal Statistical Society. Series B (Methodological), 48(1):48–59, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345637.

Boys:1986:SNM

[1750] R. J. Boys and I. R. Dunsmore. Screening in a normal model. Journal of the Royal Statistical Society. Series B (Methodological), 48(1):60–69, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345638.

West:1986:BMM

[1751] Mike West. Bayesian model monitoring. Journal of the Royal Statistical Society. Series B (Methodological), 48 (1):70–78, ???? 1986. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345639.

Smith:1986:NGS

[1752] R. L. Smith and J. E. Miller. A non-Gaussian state space model and application to prediction of records. Journal of the Royal Statistical Society. Series B (Methodological), 48(1):79–88, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345640.

Skinner:1986:DET

[1753] C. J. Skinner. Design effects of two-stage sampling. Journal of the Royal Statistical Society. Series B (Methodological), 48(1):89–99, ????

1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345641.

Silvapulle:1986:EML

[1754] M. J. Silvapulle and J. Burridge. Existence of maximum likelihood estimates in regression models for grouped and ungrouped data. Journal of the Royal Statistical Society. Series B (Methodological), 48(1):100–106, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345642.

Suen:1986:ETF

[1755] Chung-Yi Suen and I. M. Chakravarti. Efficient two-factor balanced designs. Journal of the Royal Statistical Society. Series B (Methodological), 48(1): 107–114, ???? 1986. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345643.

Hall:1986:RCO

[1756] Peter Hall. On the rate of convergence of orthogonal series density estimators. Journal of the Royal Statistical Society. Series B (Methodological), 48(1): 115–122, ???? 1986. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345644.

Kelly:1986:ARF

[1757] F. P. Kelly. On auto-repeat facilities and telephone network performance. Journal of the Royal Statistical Society. Series B (Methodological), 48(1): 123–132, ???? 1986. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345645.

Anonymous:1986:BMa

[1758] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 48(1):??, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345646.

Anonymous:1986:FMa

[1759] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 48(1):??, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345634.

Cook:1986:ALI

[1760] R. Dennis Cook. Assessment of local influence. Journal of the Royal Statistical Society. Series B (Methodological), 48 (2):133–169, ???? 1986. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345711.

Scott:1986:FLM

[1761] A. J. Scott and C. J. Wild. Fitting logistic models under case-control or choice based sampling. Journal of the Royal Statistical Society. Series B (Methodological), 48(2):170–182, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345712.

Berman:1986:SBS

[1762] Mark Berman and David Culpin. The statistical behaviour of some least squares estimators of the centre and radius of a circle. Journal of the Royal Statistical Society. Series B (Methodological), 48(2):183–196, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345713.

Ducharme:1986:TCC

[1763] Gilles R. Ducharme and Yves Lepage. Testing collapsibility in contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 48 (2):197–205, ???? 1986. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345714.

Epps:1986:TEV

[1764] T. W. Epps and L. B. Pulley. A test of exponentiality vs. monotone-hazard alternatives derived from the empirical characteristic function. Journal of the Royal Statistical Society. Series B (Methodological), 48(2):206–213, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345715.

Prentice:1986:OSP

[1765] Michael J. Prentice. Orientation statistics without parametric assumptions. Journal of the Royal Statistical Society. Series B (Methodological), 48(2): 214–222, ???? 1986. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345716.

Jones:1986:RSP

[1766] M. C. Jones and J. B. Copas. On the robustness of shrinkage predictors in regression to differences between past and future data. Journal of the Royal Statistical Society. Series B (Methodological), 48(2):223–237, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345717.

Eccleston:1986:RRC

[1767] J. A. Eccleston and J. A. John.

Recovery of row and column information in row-column designs with adjusted orthogonality. Journal of the Royal Statistical Society. Series B (Methodological), 48(2):238–243, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345718.

Heckman:1986:SSP

[1768] Nancy E. Heckman. Spline smoothing in a partly linear model. Journal of the Royal Statistical Society. Series B (Methodological), 48(2):244–248, ???? 1986. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345719.

Raftery:1986:NBF

[1769] A. E. Raftery. A note on Bayes factors for log-linear contingency table models with vague prior information. Journal of the Royal Statistical Society. Series B (Methodological), 48(2):249–250, ???? 1986. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345720.

Smith:1986:MLE

[1770] Richard L. Smith. Maximum likelihood estimation for the NEAR(2) model. Journal of the Royal Statistical Society. Series B (Methodological), 48(2): 251–257, ???? 1986. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345721.

Kruskal:1986:CGG

[1771] William Kruskal. Corrections: The geometry of generalized inverses. Journal of the Royal Statistical Society.

Series B (Methodological), 48(2):258, ???? 1986. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345722. See [1224].

Withers:1986:CAE

[1772] C. S. Withers. Corrections: Asymptotic expansions for distributions and quantiles with power series cumulants. Journal of the Royal Statistical Society. Series B (Methodological), 48(2):258, ???? 1986. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345723. See [1664].

Anonymous:1986:BMb

[1773] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 48(2):??, ???? 1986. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345724.

Anonymous:1986:FMb

[1774] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 48(2):??, ???? 1986. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345710.

Besag:1986:SAD

[1775] Julian Besag. On the statistical analysis of dirty pictures. Journal of the Royal Statistical Society. Series B (Methodological), 48(3):259–302, ???? 1986. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345426.

Wincek:1986:EML

[1776] Michael A. Wincek and Gregory C. Reinsel. An exact maximum likelihood estimation procedure for regression — ARMA time series models with possibly nonconsecutive data. Journal of the Royal Statistical Society. Series B (Methodological), 48(3):303–313, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345427.

Williams:1986:ALS

[1777] E. R. Williams, D. Ratcliff, and P. H. van Ewijk. The analysis of lattice square designs. Journal of the Royal Statistical Society. Series B (Methodological), 48(3):314–321, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345428.

Akman:1986:BFN

[1778] V. E. Akman and A. E. Raftery. Bayes factors for non-homogeneous Poisson processes with vague prior information. Journal of the Royal Statistical Society. Series B (Methodological), 48(3): 322–329, ???? 1986. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345429.

Hall:1986:SST

[1779] Peter Hall and D. M. Titterington.

On some smoothing techniques used in image restoration. Journal of the Royal Statistical Society. Series B (Methodological), 48(3):330–343, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345430.

Lockhart:1986:TFB

[1780] R. A. Lockhart, F. J. O’Reilly, and M. A. Stephens. Tests of fit based on normalized spacings. Journal of the Royal Statistical Society. Series B (Methodological), 48(3):344–352, ???? 1986. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345431.

Barndorff-Nielsen:1986:NCB

[1781] O. E. Barndorff-Nielsen and P. Blaesild. A note on the calculation of Bartlett adjustments. Journal of the Royal Statistical Society. Series B (Methodological), 48(3):353–358, ???? 1986. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345432.

Fligner:1986:DBR

[1782] M. A. Fligner and J. S. Verducci. Distance based ranking models. Journal of the Royal Statistical Society. Series B (Methodological), 48(3):359–369, ???? 1986. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345433.

Kempthorne:1986:DTM

[1783] Peter J. Kempthorne. Decisiontheoretic measures of influence in regression. Journal of the Royal Statistical Society. Series B (Methodological), 48(3):370–378, ???? 1986. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2345434.

Anonymous:1986:BMc

[1784] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 48(3):??, ???? 1986. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345435.

Anonymous:1986:FMc

[1785] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 48(3):??, ???? 1986. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345425.

Anonymous:1986:VI

[1786] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Methodological), 48(3): ??, ???? 1986. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345424.

Cox:1987:POA

[1787] D. R. Cox and N. Reid. Parameter orthogonality and approximate conditional inference. Journal of the Royal Statistical Society. Series B (Methodological), 49(1):1–39, ???? 1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345476.

Wei:1987:TSP

[1788] L. J. Wei. Two-sample problem with bivariate exchangeable observations. Journal of the Royal Statistical Society. Series B (Methodological), 49 (1):40–45, ???? 1987. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345477.

Brown:1987:CCM

[1789] Philip J. Brown and Rolf Sundberg.

Confidence and conflict in multivariate calibration. Journal of the Royal Statistical Society. Series B (Methodological), 49(1):46–57, ???? 1987. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345478.

Pettitt:1987:ERP

[1790] A. N. Pettitt. Estiamtes for a regression parameter using ranks. Journal of the Royal Statistical Society. Series B (Methodological), 49(1):58–67, ???? 1987. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345479.

Valliant:1987:SPP

[1791] Richard Valliant. Some prediction properties of balanced half-sample variance estimators in single-stage sampling. Journal of the Royal Statistical Society. Series B (Methodological), 49 (1):68–81, ???? 1987. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345480.

Hall:1987:BCC

[1792] Peter Hall. On the bootstrap and continuity correction. Journal of the Royal Statistical Society. Series B (Methodological), 49(1):82–89, ???? 1987. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345481.

Wild:1987:COC

[1793] Peter Wild. On circuits and optimality conjectures for block designs. Journal of the Royal Statistical Society. Series B (Methodological), 49(1):90–94, ???? 1987. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345482.

Cheng:1987:CML

[1794] R. C. H. Cheng and T. C. Iles. Corrected maximum likelihood in nonregular problems. Journal of the Royal Statistical Society. Series B (Methodological), 49(1):95–101, ????

1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345483.

Marcus:1987:NAC

[1795] Ruth Marcus and Abraham Genizi.

Nonparametric analysis of covariance with ordered alternatives. Journal of the Royal Statistical Society. Series B (Methodological), 49(1):102–111, ???? 1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345484.

Koreisha:1987:INE

[1796] S. G. Koreisha and T. M. Pukkila.

Identification of nonzero elements in the polynomial matrices of mixed VARMA processes. Journal of the Royal Statistical Society. Series B (Methodological), 49(1):112–126, ???? 1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2345485. See corrigendum [1835].

Anonymous:1987:BMa

[1797] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 49(1):??, ???? 1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345486.

Anonymous:1987:FMa

[1798] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 49(1):??, ???? 1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345475.

Jorgensen:1987:EDM

[1799] Bent Jorgensen. Exponential dispersion models. Journal of the Royal Statistical Society. Series B (Methodological), 49(2):127–162, ???? 1987. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345415.

Madan:1987:CPA

[1800] D. B. Madan and E. Seneta. Chebyshev polynomial approximations and characteristic function estimation. Journal of the Royal Statistical Society. Series B (Methodological), 49(2):163–169, ???? 1987. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345416.

Senthilselvan:1987:PLE

[1801] A. Senthilselvan. Penalized likelihood estimation of hazard and intensity functions. Journal of the Royal Statistical Society. Series B (Methodological), 49(2):170–174, ???? 1987. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2345417.

Copas:1987:CVS

[1802] J. B. Copas. Cross-validation shrinkage of regression predictors. Journal of the Royal Statistical Society. Series B (Methodological), 49(2):175–183, ???? 1987. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345418.

Hall:1987:CST

[1803] Peter Hall and D. M. Titterington. Common structure of techniques for choosing smoothing parameters in regression problems. Journal of the Royal Statistical Society. Series B (Methodological), 49(2):184–198, ???? 1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345419.

Poskitt:1987:PCB

[1804] D. S. Poskitt. Precision, complexity and Bayesian model determination. Journal of the Royal Statistical Society. Series B (Methodological), 49(2): 199–208, ???? 1987. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345420.

Scott:1987:PAN

[1805] David Scott. On posterior asymptotic normality and asymptotic normality of estimators for the Galton–Watson process. Journal of the Royal Statistical Society. Series B (Methodological), 49 (2):209–214, ???? 1987. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345421.

Sweeting:1987:APN

[1806] T. J. Sweeting and A. O. Adekola.

Asymptotic posterior normality for stochastic processes revisited. Journal of the Royal Statistical Society. Series B (Methodological), 49(2):215–222, ???? 1987. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345422.

Anonymous:1987:BMb

[1807] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 49(2):??, ???? 1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345423.

Anonymous:1987:FMb

[1808] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 49(2):??, ???? 1987. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345414.

Rissanen:1987:SC

[1809] Jorma Rissanen. Stochastic complexity. Journal of the Royal Statistical Society. Series B (Methodological), 49(3): 223–239, ???? 1987. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2985991.

Wallace:1987:EIC

[1810] C. S. Wallace and P. R. Freeman. Estimation and inference by compact coding. Journal of the Royal Statistical Society. Series B (Methodological), 49 (3):240–265, ???? 1987. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2985992.

Clifford:1987:SMC

[1811] Peter Clifford, N. J. B. Green, and M. J. Pilling. Statistical models of chemical kinetics in liquids. Journal of the Royal Statistical Society. Series B (Methodological), 49(3):266–300, ???? 1987. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2985993.

Brown:1987:AIR

[1812] B. M. Brown and T. P. Hettmansperger. Affine invariant rank methods in the bivariate location model. Journal of the Royal Statistical Society. Series B (Methodological), 49(3):301–310, ???? 1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985994.

Quinn:1987:NAR

[1813] B. G. Quinn, G. J. McLachlan, and N. L. Hjort. A note on the Aitkin– Rubin approach to hypothesis testing in mixture models. Journal of the Royal Statistical Society. Series B (Methodological), 49(3):311–314, ???? 1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985995.

Taniguchi:1987:MCE

[1814] Masanobu Taniguchi. Minimum contrast estimation for spectral densities of stationary processes. Journal of the Royal Statistical Society. Series B (Methodological), 49(3):315–325, ???? 1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985996.

Lenth:1987:CDB

[1815] R. V. Lenth and P. J. Green. Consistency of deviance-based M-estimators. Journal of the Royal Statistical Society. Series B (Methodological), 49(3): 326–330, ???? 1987. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985997.

Ritov:1987:TMR

[1816] Y. Ritov. Tightness of monotone random fields. Journal of the Royal Statistical Society. Series B (Methodological), 49(3):331–333, ???? 1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985998.

Azais:1987:DES

[1817] J. M. Azais. Designs of experiments for studying intergenotypic competition. Journal of the Royal Statistical Society. Series B (Methodological), 49 (3):334–345, ???? 1987. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2985999.

Browne:1987:AKF

[1818] M. W. Browne and A. Shapiro. Adjustments for kurtosis in factor analysis with elliptically distributed errors. Journal of the Royal Statistical Society. Series B (Methodological), 49(3): 346–352, ???? 1987. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2986000.

Wermuth:1987:PCL

[1819] Nanny Wermuth. Parametric collapsibility and the lack of moderating effects in contingency tables with a dichotomous response variable. Journal of the Royal Statistical Society. Series B (Methodological), 49(3):353–364, ???? 1987. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2986001.

Anonymous:1987:BMc

[1820] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 49(3):??, ???? 1987. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2986002.

Anonymous:1987:FMc

[1821] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 49(3):??, ????

1987. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985990.

Anonymous:1987:VI

[1822] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Methodological), 49(3): ??, ???? 1987. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985989.

Dawid:1988:SMH

[1823] A. P. Dawid. Symmetry models and hypotheses for structured data layouts. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):1–34, ???? 1988. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345805.

Hall:1988:SBC

[1824] Peter Hall. On symmetric bootstrap confidence intervals. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):35–45, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345806.

Bhansali:1988:COD

[1825] R. J. Bhansali. Consistent order determination for processes with infinite variance. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):46–60, ???? 1988. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345807.

Zachary:1988:CSL

[1826] Stan Zachary. Control of stochastic loss networks, with applications. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):61–73, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345808.

Davidian:1988:NEQ

[1827] M. Davidian and R. J. Carroll. A note on extended quasi-likelihood. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):74–82, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345809.

Mak:1988:ESM

[1828] Tak K. Mak. Estimating subgroup means with misclassification. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):83–92, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345810.

Morgan:1988:BPD

[1829] J. P. Morgan. Balanced polycross designs. Journal of the Royal Statistical Society. Series B (Methodological), 50 (1):93–104, ???? 1988. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345811.

Adke:1988:LDH

[1830] S. R. Adke and S. R. Deshmukh. Limit distribution of a high order Markov chain. Journal of the Royal Statistical Society. Series B (Methodological), 50 (1):105–108, ???? 1988. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345812.

Paterson:1988:EFP

[1831] L. J. Paterson. The efficiency factors of particular effects in designs with orthogonal factorial structure. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):109–117, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345813.

Cumberland:1988:DSR

[1832] W. G. Cumberland and R. M. Royall. Does simple random sampling provide adequate balance? Journal of the Royal Statistical Society. Series B (Methodological), 50(1):118–124, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345814.

Eick:1988:GPB

[1833] Stephen G. Eick. Gittins procedures for bandits with delayed responses. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):125–132, ???? 1988. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345815.

Goldstein:1988:ABS

[1834] Michael Goldstein. Adjusting belief structures. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):133–154, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345816.

Koreisha:1988:CIN

[1835] S. G. Koreisha and T. M. Pukkila.

Corrigendum: Identification of nonzero elements in the polynomial matrices of mixed VARMA processes. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):155, ???? 1988. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345817. See [1796].

Anonymous:1988:BMa

[1836] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):156, ???? 1988. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345818.

Anonymous:1988:FMa

[1837] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 50(1):??, ???? 1988. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345804.

Lauritzen:1988:LCP

[1838] S. L. Lauritzen and D. J. Spiegelhalter. Local computations with probabilities on graphical structures and their application to expert systems. Journal of the Royal Statistical Society. Series B (Methodological), 50(2):157–224, ???? 1988. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345762.

Copas:1988:BRM

[1839] J. B. Copas. Binary regression models for contaminated data. Journal of the Royal Statistical Society. Series B (Methodological), 50(2):225–265, ???? 1988. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345763.

Firth:1988:MEL

[1840] David Firth. Multiplicative errors: Log-normal or gamma? Journal of the Royal Statistical Society. Series B (Methodological), 50(2):266–268, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345764.

Skovgaard:1988:SED

[1841] Ib M. Skovgaard. Saddlepoint expansions for directional test probabilities. Journal of the Royal Statistical Society. Series B (Methodological), 50(2): 269–280, ???? 1988. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345765.

Blackwell:1988:IAL

[1842] Louis M. Blackwell and Nozer D.

Singpurwalla. Inference from accelerated life tests using filtering in coloured noise. Journal of the Royal Statistical Society. Series B (Methodological), 50 (2):281–292, ???? 1988. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345766.

Hamada:1988:NEM

[1843] M. Hamada and S. K. Tse. A note on the existence of maximum likelihood estimates in linear regression models using interval-censored data. Journal of the Royal Statistical Society. Series B (Methodological), 50(2):293–296, ???? 1988. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345767.

Vecchia:1988:EMI

[1844] A. V. Vecchia. Estimation and model identification for continuous spatial processes. Journal of the Royal Statistical Society. Series B (Methodological), 50(2):297–312, ???? 1988. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345768.

Karlsen:1988:CET

[1845] Hans Karlsen and Dag Tjostheim. Consistent estimates for the NEAR(2) and NLAR(2) time series models. Journal of the Royal Statistical Society. Series B (Methodological), 50(2):313–320, ???? 1988. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345769.

Anonymous:1988:BMb

[1846] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 50(2):??, ???? 1988. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345770.

Anonymous:1988:FMb

[1847] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 50(2):??, ???? 1988. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345761.

Hinkley:1988:BM

[1848] David V. Hinkley. Bootstrap methods. Journal of the Royal Statistical Society. Series B (Methodological), 50(3): 321–337, ???? 1988. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345698. See

discussion [1850].

Diciccio:1988:RBC

[1849] Thomas J. Diciccio and Joseph P. Romano. A review of bootstrap confidence intervals. Journal of the Royal Statistical Society. Series B (Methodological), 50(3):338–354, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2345699. See discussion [1850] and corrigenda [1898].

Anonymous:1988:DPH

[1850] Anonymous. Discussion of the papers by Hinkley and Diciccio and Romano. Journal of the Royal Statistical Society. Series B (Methodological), 50(3): 355–370, ???? 1988. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345700. See

[1848, 1849].

Pettit:1988:BMO

[1851] L. I. Pettit. Bayes methods for outliers in exponential samples. Journal of the Royal Statistical Society. Series B (Methodological), 50(3):371–380, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345701.

Hall:1988:DSQ

[1852] Peter Hall and Simon J. Sheather. On the distribution of a studentized quantile. Journal of the Royal Statistical Society. Series B (Methodological), 50 (3):381–391, ???? 1988. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345702.

Gaver:1988:NEP

[1853] D. P. Gaver and P. A. Jacobs. Nonparametric estimation of the probability of a long delay in the M/G/1 queue. Journal of the Royal Statistical Society. Series B (Methodological), 50(3):392–402, ???? 1988. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345703.

Mau:1988:GNT

[1854] Jochen Mau. A generalization of a nonparametric test for stochastically ordered distributions to censored survival data. Journal of the Royal Statistical Society. Series B (Methodological), 50 (3):403–412, ???? 1988. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345704.

Speckman:1988:KSP

[1855] Paul Speckman. Kernel smoothing in partial linear models. Journal of the Royal Statistical Society. Series B (Methodological), 50(3):413–436, ???? 1988. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345705.

Tableman:1988:BET

[1856] Mara Tableman and Thomas P. Hettmansperger. Bahadur efficiency of two-sample tests based on onesample sign statistics. Journal of the Royal Statistical Society. Series B (Methodological), 50(3):437–444, ???? 1988. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345706.

Davison:1988:ACI

[1857] A. C. Davison. Approximate conditional inference in generalized linear models. Journal of the Royal Statistical Society. Series B (Methodological), 50 (3):445–461, ???? 1988. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345707.

Garthwaite:1988:QEO

[1858] Paul H. Garthwaite and James M. Dickey. Quantifying expert opinion in linear regression problems. Journal of the Royal Statistical Society. Series B (Methodological), 50(3):462–474, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345708.

Anonymous:1988:VI

[1859] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Methodological), 50(3): 475–476, ???? 1988. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345696.

Anonymous:1988:BMc

[1860] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 50(3):??, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345709.

Anonymous:1988:FMc

[1861] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 50(3):??, ???? 1988. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345697.

Anonymous:1989:RE

[1862] Anonymous. Report of the Editors.

Journal of the Royal Statistical Society. Series B (Methodological), 51(1):1– 2, ???? 1989. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345836.

Hall:1989:VFE

[1863] Peter Hall and R. J. Carroll. Variance function estimation in regression: The effect of estimating the mean. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):3–14, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345837.

Wilson:1989:UML

[1864] G. Tunnicliffe Wilson. On the use of marginal likelihood in time series model estimation. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):15–27, ???? 1989. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345838.

Poskitt:1989:MEI

[1865] D. S. Poskitt. A method for the estimation and identification of transfer function models. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):29–46, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2345839. See corrigendum [1926].

Smyth:1989:GLM

[1866] Gordon K. Smyth. Generalized linear models with varying dispersion. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):47–60, ???? 1989. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345840.

Damsleth:1989:AMD

[1867] Eivind Damsleth and A. H. El-Shaarawi. ARMA models with doubleexponentially distributed noise. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):61–69, ???? 1989. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345841.

Honda:1989:OST

[1868] Yuzo Honda. On the optimality of some tests of the error covariance matrix in the linear regression model. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):71–79, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345842.

Berman:1989:EWI

[1869] Mark Berman and Peter Diggle. Estimating weighted integrals of the second-order intensity of a spatial point process. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):81–92, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345843.

Crowder:1989:MDW

[1870] Martin Crowder. A multivariate distribution with Weibull connections. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):93–107, ???? 1989. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345844.

Lesaffre:1989:PSL

[1871] E. Lesaffre and A. Albert. Partial separation in logistic discrimination. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):109–116, ???? 1989. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345845.

Brown:1989:AIB

[1872] B. M. Brown and T. P. Hettmansperger. An affine invariant bivariate version of the sign test. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):117–125, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345846.

Meilijson:1989:FIE

[1873] Isaac Meilijson. A fast improvement to the EM algorithm on its own terms. Journal of the Royal Statistical Society. Series B (Methodological), 51(1): 127–138, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345847.

Shao:1989:JWL

[1874] Jun Shao. Jackknifing weighted least squares estimators. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):139–156, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345848.

Anonymous:1989:BMa

[1875] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):??, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345849.

Anonymous:1989:FMa

[1876] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 51(1):??, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345835.

Tiao:1989:MSM

[1877] George C. Tiao and Ruey S. Tsay.

Model specification in multivariate time series. Journal of the Royal Statistical Society. Series B (Methodological), 51(2):157–213, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345602.

Joshi:1989:CEA

[1878] V. M. Joshi. A counter-example against the likelihood principle. Journal of the Royal Statistical Society. Series B (Methodological), 51(2):215–216, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345603.

Hannan:1989:REA

[1879] E. J. Hannan, A. J. McDougall, and D. S. Poskitt. Recursive estimation of autoregressions. Journal of the Royal Statistical Society. Series B (Methodological), 51(2):217–233, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345604.

Massam:1989:EDT

[1880] Hélène Massam. An exact decomposition theorem for a sample from the three-dimensional hyperboloid distribution. Journal of the Royal Statistical Society. Series B (Methodological), 51(2):235–240, ???? 1989. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345605.

Prentice:1989:SRM

[1881] Michael J. Prentice. Spherical regression on matched pairs of orientation statistics. Journal of the Royal Statistical Society. Series B (Methodological), 51(2):241–248, ???? 1989. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345606.

Chiu:1989:DPC

[1882] Shean-Tsong Chiu. Detecting periodic components in a white Gaussian time series. Journal of the Royal Statistical Society. Series B (Methodological), 51 (2):249–259, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345607.

Kuk:1989:MEP

[1883] Anthony Y. C. Kuk and T. K. Mak. Median estimation in the presence of auxiliary information. Journal of the Royal Statistical Society. Series B (Methodological), 51(2):261–269, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345608.

Greig:1989:EMP

[1884] D. M. Greig, B. T. Porteous, and A. H. Seheult. Exact maximum a posteriori estimation for binary images. Journal of the Royal Statistical Society. Series B (Methodological), 51(2):271–279, ???? 1989. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345609.

Madan:1989:CPA

[1885] D. B. Madan and E. Seneta. Chebyshev polynomial approximations for characteristic function estimation: Some theoretical supplements. Journal of the Royal Statistical Society. Series B (Methodological), 51(2):281–285, ????

1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345610.

Turkman:1989:OSM

[1886] K. F. Turkman and M. A. Amaral Turkman. Optimal screening methods. Journal of the Royal Statistical Society. Series B (Methodological), 51(2): 287–295, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345611.

Donev:1989:DEB

[1887] A. N. Donev. Design of experiments with both mixture and qualitative factors. Journal of the Royal Statistical Society. Series B (Methodological), 51 (2):297–302, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345612.

Hosking:1989:CEF

[1888] J. R. M. Hosking. Corrigendum:

Equivalent forms of the multivariate portmanteau statistic. Journal of the Royal Statistical Society. Series B (Methodological), 51(2):303, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345613. See [1509].

Anonymous:1989:BMb

[1889] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 51(2):??, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345614.

Anonymous:1989:FMb

[1890] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 51(2):??, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345601.

Jennison:1989:IAR

[1891] Christopher Jennison and Bruce W. Turnbull. Interim analyses: The repeated confidence interval approach. Journal of the Royal Statistical Society. Series B (Methodological), 51(3): 305–361, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345448.

Bruce:1989:LDT

[1892] Andrew G. Bruce and R. Douglas Martin. Leave-k-out diagnostics for time series. Journal of the Royal Statistical Society. Series B (Methodological), 51 (3):363–424, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345449. See

corrigendum [1916].

Gilula:1989:CTW

[1893] Zvi Gilula and Abba M. Krieger. Collapsed two-way contingency tables and the chi-square reduction principle. Journal of the Royal Statistical Society. Series B (Methodological), 51(3): 425–433, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345450.

Hosoya:1989:HSM

[1894] Yuzo Hosoya. Hierarchical statistical models and a generalized likelihood ratio test. Journal of the Royal Statistical Society. Series B (Methodological), 51 (3):435–447, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345451.

Mardia:1989:SAT

[1895] K. V. Mardia. Shape analysis of triangles through directional techniques. Journal of the Royal Statistical Society. Series B (Methodological), 51(3): 449–458, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345452.

Hall:1989:ESO

[1896] Peter Hall and D. M. Titterington. The effect of simulation order on level accuracy and power of Monte Carlo tests. Journal of the Royal Statistical Society. Series B (Methodological), 51(3): 459–467, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345453.

Anonymous:1989:VI

[1897] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Methodological), 51(3): 469–470, ???? 1989. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345446.

Diciccio:1989:CRB

[1898] Thomas J. Diciccio and Joseph P. Romano. Corrigendum: A review of bootstrap confidence intervals. Journal of the Royal Statistical Society. Series B (Methodological), 51(3):470, ???? 1989. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345454. See [1849].

Anonymous:1989:BMc

[1899] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 51(3):??, ????

1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345455.

Anonymous:1989:FMc

[1900] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 51(3):??, ???? 1989. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345447.

Anonymous:1990:RE

[1901] Anonymous. Report of the Editors. Journal of the Royal Statistical Society. Series B (Methodological), 52(1):1– 2, ???? 1990. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345648.

Edwards:1990:HIM

[1902] David Edwards. Hierarchical interaction models. Journal of the Royal Statistical Society. Series B (Methodological), 52(1):3–20, ???? 1990. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345649.

See [1904].

Wermuth:1990:SRH

[1903] Nanny Wermuth and Steffen Lilholt Lauritzen. On substantive research hypotheses, conditional independence graphs and graphical chain models. Journal of the Royal Statistical Society. Series B (Methodological), 52(1):21–50, ???? 1990. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345650. See [1904].

Anonymous:1990:DPE

[1904] Anonymous. Discussion of the papers by Edwards, and Wermuth and Lauritzen. Journal of the Royal Statistical Society. Series B (Methodological), 52 (1):51–72, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345651. See

[1901, 1903].

Cuzick:1990:SCI

[1905] Jack Cuzick and Robert Edwards. Spatial clustering for inhomogeneous populations. Journal of the Royal Statistical Society. Series B (Methodological), 52 (1):73–104, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345652.

Hosking:1990:MAE

[1906] J. R. M. Hosking. L-moments: Analysis and estimation of distributions using linear combinations of order statistics. Journal of the Royal Statistical Society. Series B (Methodological), 52 (1):105–124, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345653.

Pace:1990:BCT

[1907] L. Pace and A. Salvan. Best conditional tests for separate families of hypotheses. Journal of the Royal Statistical Society. Series B (Methodological), 52 (1):125–134, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345654.

Cheng:1990:EMT

[1908] R. C. H. Cheng and T. C. Iles. Embedded models in three-parameter distributions and their estimation. Journal of the Royal Statistical Society. Series B (Methodological), 52(1):135–149, ???? 1990. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345655.

Mittnik:1990:CTA

[1909] Stefan Mittnik. Computation of theoretical autocovariance matrices of multivariate autoregressive moving average time series. Journal of the Royal Statistical Society. Series B (Methodological), 52(1):151–155, ???? 1990. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345656.

Guttorp:1990:NEI

[1910] Peter Guttorp and Mary Lou Thompson. Nonparametric estimation of intensities for sampled counting processes. Journal of the Royal Statistical Society. Series B (Methodological), 52 (1):157–173, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345657.

Pettit:1990:CPO

[1911] L. I. Pettit. The conditional predictive ordinate for the normal distribution. Journal of the Royal Statistical Society. Series B (Methodological), 52 (1):175–184, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345658.

Graham:1990:BDB

[1912] R. L. Graham, D. V. Hinkley, P. W. M. John, and S. Shi. Balanced design of bootstrap simulations. Journal of the Royal Statistical Society. Series B (Methodological), 52(1):185–202, ???? 1990. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345659.

Truong-Van:1990:NAF

[1913] B. Truong-Van. A new approach to frequency analysis with amplified harmonics. Journal of the Royal Statistical Society. Series B (Methodological), 52 (1):203–221, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345660.

Hardle:1990:BCD

[1914] Wolfgang Hardle, J. S. Marron, and M. P. Wand. Bandwidth choice for density derivatives. Journal of the Royal Statistical Society. Series B (Methodological), 52(1):223–232, ???? 1990. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345661.

Anonymous:1990:NA

[1915] Anonymous. Notes for authors. Journal of the Royal Statistical Society. Series B (Methodological), 52(1):233–236, ???? 1990. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345662.

Bruce:1990:CLD

[1916] Andrew G. Bruce and R. Douglas Martin. Corrigendum: Leave-k-Out diagnostics for time series. Journal of the Royal Statistical Society. Series B (Methodological), 52(1):236, ???? 1990. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345663. See [1892].

Anonymous:1990:BMa

[1917] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 52(1):??, ????

1990. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345664.

Anonymous:1990:FMa

[1918] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 52(1):??, ???? 1990. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345647.

Stone:1990:CRC

[1919] M. Stone and R. J. Brooks. Continuum regression: Cross-validated sequentially constructed prediction embracing ordinary least squares, partial least squares and principal components regression. Journal of the Royal Statistical Society. Series B (Methodological), 52(2):237–269, ???? 1990. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2345437. See corrigendum [2054].

Silverman:1990:SEA

[1920] B. W. Silverman, M. C. Jones, J. D. Wilson, and D. W. Nychka. A smoothed EM approach to indirect estimation problems, with particular, reference to stereology and emission tomography. Journal of the Royal Statistical Society. Series B (Methodological), 52(2):271–324, ???? 1990. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345438.

McCullagh:1990:SMA

[1921] Peter McCullagh and Robert Tibshirani. A simple method for the adjustment of profile likelihoods. Journal of the Royal Statistical Society. Series B (Methodological), 52(2):325–344, ???? 1990. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345439.

Stefanski:1990:STG

[1922] Leonard A. Stefanski and Raymond J.

Carroll. Score tests in generalized linear measurement error models. Journal of the Royal Statistical Society. Series B (Methodological), 52(2):345–359, ???? 1990. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345440.

Park:1990:AEF

[1923] D. K. Park and A. M. Dean. Average efficiency factors and adjusted orthogonality in multidimensional designs. Journal of the Royal Statistical Society. Series B (Methodological), 52 (2):361–368, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345441.

Becker:1990:QMA

[1924] Mark P. Becker. Quasisymmetric models for the analysis of square contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 52(2):369–378, ???? 1990. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345442.

Pena:1990:BEM

[1925] Edsel A. Pena and Arjun K. Gupta.

Bayes estimation for the Marshall– Olkin exponential distribution. Journal of the Royal Statistical Society. Series B (Methodological), 52(2):379–389, ???? 1990. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345443.

Poskitt:1990:CME

[1926] D. S. Poskitt. Corrigendum: A method for the estimation and identification of transfer function models. Journal of the Royal Statistical Society. Series B (Methodological), 52(2):391, ???? 1990. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345444. See [1865].

Anonymous:1990:BMb

[1927] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 52(2):392, ???? 1990. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345445.

Anonymous:1990:FMb

[1928] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 52(2):??, ???? 1990. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345436.

Davison:1990:MEH

[1929] A. C. Davison and R. L. Smith. Models for exceedances over high thresholds. Journal of the Royal Statistical Society. Series B (Methodological), 52 (3):393–442, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345667.

Green:1990:UEP

[1930] Peter J. Green. On use of the EM for penalized likelihood estimation. Journal of the Royal Statistical Society. Series B (Methodological), 52(3):443–452, ???? 1990. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345668.

Wei:1990:DTS

[1931] William W. S. Wei and Daniel O.

Stram. Disaggregation of time series models. Journal of the Royal Statistical Society. Series B (Methodological), 52 (3):453–467, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345669.

Chan:1990:LRT

[1932] K. S. Chan and H. Tong. On likelihood ratio tests for threshold autoregression. Journal of the Royal Statistical Society. Series B (Methodological), 52(3): 469–476, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345670.

Young:1990:ASB

[1933] G. A. Young. Alternative smoothed bootstraps. Journal of the Royal Statistical Society. Series B (Methodological), 52(3):477–484, ???? 1990. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345671.

Barndorff-Nielsen:1990:AIP

[1934] O. E. Barndorff-Nielsen. Approximate interval probabilities. Journal of the Royal Statistical Society. Series B (Methodological), 52(3):485–496, ???? 1990. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345672.

Whitaker:1990:CBD

[1935] D. Whitaker, C. M. Triggs, and J. A.

John. Construction of block designs using mathematical programming. Journal of the Royal Statistical Society. Series B (Methodological), 52(3):497–503, ???? 1990. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345673.

Anonymous:1990:VI

[1936] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 52(3): 505–506, ???? 1990. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345665.

Anonymous:1990:BMc

[1937] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 52(3):??, ???? 1990. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345674.

Anonymous:1990:FMc

[1938] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 52(3):??, ???? 1990. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345666.

Anonymous:1991:RE

[1939] Anonymous. Report of the Editor. Journal of the Royal Statistical Society. Series B (Methodological), 53(1):1– 2, ???? 1991. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345726.

Thomas:1991:GHG

[1940] Alun Thomas. Gene hunting with gradients of likelihoods. Journal of the Royal Statistical Society. Series B (Methodological), 53(1):3–26, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345727.

Bailey:1991:SRE

[1941] R. A. Bailey. Strata for randomized experiments. Journal of the Royal Statistical Society. Series B (Methodological), 53(1):27–78, ???? 1991. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345728.

Dawid:1991:FIL

[1942] A. P. Dawid. Fisherian inference in likelihood and prequential frames of reference. Journal of the Royal Statistical Society. Series B (Methodological), 53 (1):79–109, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345729.

Aitkin:1991:PBF

[1943] Murray Aitkin. Posterior Bayes factors. Journal of the Royal Statistical Society. Series B (Methodological), 53 (1):111–142, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345730.

Moreno:1991:RBA

[1944] E. Moreno and J. A. Cano. Robust Bayesian analysis with ϵcontaminations partially known. Journal of the Royal Statistical Society. Series B (Methodological), 53(1):143–155, ???? 1991. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345731.

Jensen:1991:USA

[1945] Jens Ledet Jensen. Uniform saddlepoint approximations and log-concave densities. Journal of the Royal Statistical Society. Series B (Methodological), 53(1):157–172, ???? 1991. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345732.

Hart:1991:KRE

[1946] Jeffrey D. Hart. Kernel regression estimation with time series errors. Journal of the Royal Statistical Society. Series B (Methodological), 53(1):173–187, ???? 1991. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345733.

Carlin:1991:SRM

[1947] Bradley P. Carlin and Alan E. Gelfand. A sample reuse method for accurate parametric empirical Bayes confidence intervals. Journal of the Royal Statistical Society. Series B (Methodological), 53(1):189–200, ???? 1991. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345734.

Hemerly:1991:ROE

[1948] E. M. Hemerly and M. H. A. Davis. Recursive order estimation of autoregressions without bounding the model set. Journal of the Royal Statistical Society. Series B (Methodological), 53(1): 201–210, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345735.

Nakamura:1991:EML

[1949] T. Nakamura. Existence of maximum likelihood estimates for intervalcensored data from some threeparameter models with a shifted origin. Journal of the Royal Statistical Society. Series B (Methodological), 53 (1):211–220, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345736.

Mudholkar:1991:GPC

[1950] Govind S. Mudholkar, Georgia D. Kollia, C. Thomas Lin, and Kartik R. Patel. A graphical procedure for comparing goodness-of-fit tests. Journal of the Royal Statistical Society. Series B (Methodological), 53(1):221–232, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345737.

Rice:1991:EMC

[1951] John A. Rice and B. W. Silverman. Estimating the mean and covariance structure nonparametrically when the data are curves. Journal of the Royal Statistical Society. Series B (Methodological), 53(1):233–243, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345738.

Hall:1991:LMC

[1952] Peter Hall and J. S. Marron. Local minima in cross-validation functions. Journal of the Royal Statistical Society. Series B (Methodological), 53(1): 245–252, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345739. See

corrigendum [2055].

Lele:1991:JLE

[1953] Subhash Lele. Jackknifing linear estimating equations: Asymptotic theory and applications in stochastic processes. Journal of the Royal Statistical Society. Series B (Methodological), 53 (1):253–267, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345740.

Rida:1991:APS

[1954] Wasima N. Rida. Asymptotic properties of some estimators for the infection rate in the general stochastic epidemic model. Journal of the Royal Statistical Society. Series B (Methodological), 53 (1):269–283, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345741.

Anonymous:1991:BMa

[1955] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 53(1):??, ???? 1991. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345742.

Anonymous:1991:FMa

[1956] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 53(1):??, ???? 1991. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345725.

Goodall:1991:PMS

[1957] Colin Goodall. Procrustes methods in the statistical analysis of shape. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):285–339, ???? 1991. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345744.

Pepe:1991:WKM

[1958] Margaret Sullivan Pepe and Thomas R. Fleming. Weighted Kaplan–Meier statistics: Large sample and optimality considerations. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):341–352, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345745.

Taylor:1991:EDF

[1959] Charles C. Taylor and S. James Taylor.

Estimating the dimension of a fractal. Journal of the Royal Statistical Society. Series B (Methodological), 53(2): 353–364, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345746.

Arnold:1991:BDC

[1960] B. C. Arnold and D. J. Strauss. Bivariate distributions with conditionals in prescribed exponential families. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):365–375, ???? 1991. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345747. See corrigen-

dum [1990].

Coles:1991:MEM

[1961] Stuart G. Coles and Jonathan A. Tawn.

Modelling extreme multivariate events. Journal of the Royal Statistical Society. Series B (Methodological), 53(2): 377–392, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345748.

Lugtenburg:1991:ABA

[1962] D. Lugtenburg. An approximate Bartlett adjustment for testing the goodness of fit of multinomial regression models. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):393–398, ????

1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345749.

Bhaumik:1991:ORU

[1963] Dulal K. Bhaumik and Dexter C. Whittinghill III. Optimality and robustness to the unavailability of blocks in block designs. Journal of the Royal Statistical Society. Series B (Methodological), 53 (2):399–407, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345750.

Cullis:1991:EMD

[1964] B. R. Cullis, C. A. McGilchrist, and A. C. Gleeson. Error model diagnostics in the general linear model relevant to the analysis of repeated measurements and field experiments. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):409–416, ???? 1991. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345751.

Clarkson:1991:CEM

[1965] Douglas B. Clarkson and Robert I. Jennrich. Computing extended maximum likelihood estimates for linear parameter models. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):417–426, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345752.

Johnson:1991:BIM

[1966] Wesley O. Johnson and Joseph L. Gastwirth. Bayesian inference for medical screening tests: Approximations useful for the analysis of acquired immune deficiency syndrome. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):427–439, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345753.

Dorfman:1991:SCI

[1967] Alan H. Dorfman. Sound confidence intervals in the heteroscedastic linear model through releveraging. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):441–452, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345754.

Vieu:1991:NRO

[1968] Philippe Vieu. Nonparametric regression: Optimal local bandwidth choice. Journal of the Royal Statistical Society. Series B (Methodological), 53(2): 453–464, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345755.

Luong:1991:MDM

[1969] Andrew Luong. Minimum distance methods based on quadratic distances for transforms in simple linear regression models. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):465–471, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345756.

Atkinson:1991:GLS

[1970] Anthony C. Atkinson and Luis R.

Pericchi. Grouped likelihood for the shifted power transformation. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):473–482, ???? 1991. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345757.

Fraser:1991:ELM

[1971] D. A. S. Fraser, N. Reid, and A. Wong. Exponential linear models: A two-pass procedure for saddlepoint approximation. Journal of the Royal Statistical Society. Series B (Methodological), 53 (2):483–492, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345758.

Peters:1991:BSR

[1972] Dawn Peters and Ronald H. Randles. A bivariate signed rank test for the twosample location problem. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):493–504, ???? 1991. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345759.

Anonymous:1991:BMb

[1973] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):??, ???? 1991. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345760.

Anonymous:1991:FMb

[1974] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 53(2):??, ???? 1991. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345743.

Whittle:1991:LCP

[1975] P. Whittle. Likelihood and cost as path integrals. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):505–538, ???? 1991. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345585.

Ramsay:1991:STF

[1976] J. O. Ramsay and C. J. Dalzell. Some tools for functional data analysis. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):539–572, ???? 1991. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345586.

Carroll:1991:SEL

[1977] R. J. Carroll and M. P. Wand. Semiparametric estimation in logistic measurement error models. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):573–585, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345587.

Schucany:1991:OSB

[1978] William R. Schucany and Suojin Wang.

One-step bootstrapping for smooth iterative procedures. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):587–596, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345588.

Rosenbaum:1991:COD

[1979] Paul R. Rosenbaum. A characterization of optimal designs for observational studies. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):597–610, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345589.

Severini:1991:RBB

[1980] Thomas A. Severini. On the relationship between Bayesian and non-Bayesian interval estimates. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):611–618, ???? 1991. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345590.

Goldstein:1991:LP

[1981] M. Goldstein and J. V. Howard. A likelihood paradox. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):619–628, ???? 1991. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345591.

Cordeiro:1991:BCG

[1982] Gauss M. Cordeiro and Peter Mc-Cullagh. Bias correction in generalized linear models. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):629–643, ???? 1991. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345592.

Bailey:1991:NCR

[1983] R. A. Bailey and H. D. Patterson. A note on the construction of row-and-column designs with two replicates. Journal of the Royal Statistical Society. Series B (Methodological), 53(3): 645–648, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345593.

Konishi:1991:IMM

[1984] Sadanori Konishi, C. G. Khatri, and C. Radhakrishna Rao. Inferences on multivariate measures of interclass and intraclass correlations in familial data. Journal of the Royal Statistical Society. Series B (Methodological), 53(3): 649–659, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345594.

Qian:1991:MMC

[1985] W. Qian and D. M. Titterington.

Multidimensional Markov chain models for image textures. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):661–674, ???? 1991. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345595.

Polson:1991:BDT

[1986] Nicholas G. Polson and Gareth O.

Roberts. A Bayesian decision theoretic characterization of Poisson processes. Journal of the Royal Statistical Society. Series B (Methodological), 53(3): 675–682, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345596.

Sheather:1991:RDB

[1987] S. J. Sheather and M. C. Jones. A reliable data-based bandwidth selection method for kernel density estimation. Journal of the Royal Statistical Society. Series B (Methodological), 53(3): 683–690, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345597.

Chan:1991:PPL

[1988] K. S. Chan. Percentage points of likelihood ratio tests for threshold autoregression. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):691–696, ???? 1991. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2345598.

Anonymous:1991:VI

[1989] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 53(3): 697–699, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345583.

Arnold:1991:CBD

[1990] B. C. Arnold and D. J. Strauss. Corrigendum: Bivariate distributions with conditionals in prescribed exponential families. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):700, ???? 1991. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345599. See

[1960].

Anonymous:1991:BMc

[1991] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):??, ???? 1991. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345600.

Anonymous:1991:FMc

[1992] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 53(3):??, ???? 1991. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345584.

Anonymous:1992:RE

[1993] Anonymous. Report of the Editors. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):1– 2, ???? 1992. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345946.

Liang:1992:MRA

[1994] Kung-Yee Liang, Scott L. Zeger, and Bahjat Qaqish. Multivariate regression analyses for categorical data. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):3–40, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345947.

Donoho:1992:MEN

[1995] David L. Donoho, Iain M. Johnstone, Jeffrey C. Hoch, and Alan S. Stern. Maximum entropy and the nearly black object. Journal of the Royal Statistical Society. Series B (Methodological), 54 (1):41–81, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345948.

Efron:1992:JAB

[1996] Bradley Efron. Jackknife-after-bootstrap standard errors and influence functions. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):83–127, ???? 1992. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345949.

Kass:1992:ABF

[1997] Robert E. Kass and Suresh K.

Vaidyanathan. Approximate Bayes factors and orthogonal parameters, with application to testing equality of two binomial proportions. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):129–144, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345950.

Singpurwalla:1992:NHA

[1998] Nozer D. Singpurwalla and Refik Soyer. Non-homogeneous autoregressive processes for tracking (software) reliability growth, and their Bayesian analysis. Journal of the Royal Statistical Society. Series B (Methodological), 54(1): 145–156, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345951.

Boys:1992:KAS

[1999] Richard J. Boys. On a kernel approach to a screening problem. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):157–169, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345952.

Joe:1992:BTM

[2000] Harry Joe, Richard L. Smith, and Ishay Weissman. Bivariate threshold methods for extremes. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):171–183, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345953.

Gart:1992:IAT

[2001] John J. Gart. The identity of Armitage’s two tests of heterogeneity of proportions for proportional subclass numbers. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):185–187, ???? 1992. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2345954.

Mukerjee:1992:CBC

[2002] Rahul Mukerjee. Comparison between the conditional likelihood ratio and the usual likelihood ratio tests. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):189–194, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345955.

Wallace:1992:SFA

[2003] C. S. Wallace and P. R. Freeman.

Single-factor analysis by minimum message length estimation. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):195–209, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345956.

Tanabe:1992:ECD

[2004] Kunio Tanabe and Masahiko Sagae.

An exact Cholesky decomposition and the generalized inverse of the variancecovariance matrix of the multinomial distribution, with applications. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):211–219, ???? 1992. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345957.

Hall:1992:RST

[2005] Peter Hall. On the removal of skewness by transformation. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):221–228, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345958.

Benkherouf:1992:GIO

[2006] L. Benkherouf, K. D. Glazebrook, and R. W. Owen. Gittins indices and oil exploration. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):229–241, ???? 1992. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2345959.

Percy:1992:PSU

[2007] David F. Percy. Prediction for seemingly unrelated regressions. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):243–252, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345960.

Khuri:1992:DRC

[2008] Andre I. Khuri. Diagnostic results concerning a measure of rotatability. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):253–267, ???? 1992. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345961.

Liang:1992:CED

[2009] Ta Chen Liang and N. Balakrishnan. A characterization of exponential distributions through conditional independence. Journal of the Royal Statistical Society. Series B (Methodological), 54 (1):269–271, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345962.

Nelder:1992:LQL

[2010] J. A. Nelder and Y. Lee. Likelihood, quasi-likelihood and pseudolikelihood: Some comparisons. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):273–284, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345963.

West:1992:MPA

[2011] Mike West and Jo Crosse. Modelling probabilistic agent opinion. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):285–299, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345964.

Anonymous:1992:BMa

[2012] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):??, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345965.

Anonymous:1992:FMa

[2013] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 54(1):??, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345945.

Anonymous:1992:I

[2014] Anonymous. Introduction. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):301–302, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346129.

Casdagli:1992:CDV

[2015] Martin Casdagli. Chaos and deterministic versus stochastic non-linear modelling. Journal of the Royal Statistical Society. Series B (Methodological), 54 (2):303–328, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346130.

Smith:1992:EDN

[2016] Richard L. Smith. Estimating dimension in noisy chaotic time series. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):329–351, ???? 1992. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346131.

Wolff:1992:LLE

[2017] Rodney C. L. Wolff. Local Lyapunov exponents: Looking closely at chaos. Journal of the Royal Statistical Society. Series B (Methodological), 54(2): 353–371, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346132.

Broomhead:1992:LFN

[2018] D. S. Broomhead, J. P. Huke, and M. R. Muldoon. Linear filters and non-linear systems. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):373–382, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346133.

Grenfell:1992:CCM

[2019] B. T. Grenfell. Chance and chaos in measles dynamics. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):383–398, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346134.

Nychka:1992:FCN

[2020] Douglas Nychka, Stephen Ellner, A. Ronald Gallant, and Daniel McCaffrey. Finding chaos in noisy systems. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):

399–426, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346135.

Cheng:1992:CNO

[2021] B. Cheng and H. Tong. On consistent nonparametric order determination and chaos. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):427–449, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346136.

Anonymous:1992:DMC

[2022] Anonymous. Discussion on the meeting on chaos. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):451–474, ???? 1992. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2346137.

Hall:1992:EFE

[2023] Peter Hall and Iain Johnstone. Empirical functionals and efficient smoothing parameter selection. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):475–530, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346138.

Gart:1992:PTA

[2024] John J. Gart. Pooling 2 × 2 tables: Asymptotic moments of estimators. Journal of the Royal Statistical Society. Series B (Methodological), 54 (2):531–539, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346139.

Owen:1992:CLT

[2025] Art B. Owen. A central limit theorem for Latin hypercube sampling. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):541–551, ???? 1992. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346140.

West:1992:MAF

[2026] Mike West. Modelling agent forecast distributions. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):553–567, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346141.

Ford:1992:UCF

[2027] I. Ford, B. Torsney, and C. F. J. Wu. The use of a canonical form in the construction of locally optimal designs for non-linear problems. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):569–583, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346142.

Geng:1992:CRR

[2028] Zhi Geng. Collapsibility of relative risk in contingency tables with a response variable. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):585–593, ???? 1992. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2346143.

Do:1992:DEU

[2029] Kim-Anh Do and Peter Hall. Distribution estimation using concomitants of order statistics, with application to Monte Carlo simulation for the bootstrap. Journal of the Royal Statistical Society. Series B (Methodological), 54 (2):595–607, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346144.

Li:1992:AEA

[2030] Zhaohai Li and Cun-Hui Zhang.

Asymptotically efficient allocation rules for two Bernoulli populations. Journal of the Royal Statistical Society. Series B (Methodological), 54(2): 609–616, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346145.

Sasieni:1992:IBC

[2031] Peter Sasieni. Information bounds for the conditional hazard ratio in a nested family of regression models. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):617–635, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346146.

Helland:1992:MLR

[2032] Inge S. Helland. Maximum likelihood regression on relevant components. Journal of the Royal Statistical Society. Series B (Methodological), 54 (2):637–647, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346147.

Mak:1992:EPH

[2033] Tak K. Mak. Estimation of parameters in heteroscedastic linear models. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):649–655, ???? 1992. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346148.

Anonymous:1992:BMb

[2034] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):??, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346149.

Anonymous:1992:FMb

[2035] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 54(2):??, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346128.

Geyer:1992:CMC

[2036] Charles J. Geyer and Elizabeth A. Thompson. Constrained Monte Carlo maximum likelihood for dependent data. Journal of the Royal Statistical Society. Series B (Methodological), 54 (3):657–699, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345852.

Pierce:1992:PUH

[2037] Donald A. Pierce and Dawn Peters. Practical use of higher order asymptotics for multiparameter exponential families. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):701–737, ???? 1992. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2345853.

Ebrahimi:1992:TEB

[2038] Nader Ebrahimi, Mohamed Habibullah, and Ehsan S. Soofi. Testing exponentiality based on Kullback– Leibler information. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):739–748, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345854.

Beran:1992:GFT

[2039] Jan Beran. A goodness-of-fit test for time series with long range dependence. Journal of the Royal Statistical Society. Series B (Methodological), 54(3): 749–760, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345855.

Hadi:1992:IMO

[2040] Ali S. Hadi. Identifying multiple outliers in multivariate data. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):761–771, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345856.

Carter:1992:CVE

[2041] C. K. Carter and G. K. Eagleson. A comparison of variance estimators in nonparametric regression. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):773–780, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345857.

Andersen:1992:DCD

[2042] Erling B. Andersen. Diagnostics in categorical data analysis. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):781–791, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345858.

Pericchi:1992:EAP

[2043] L. R. Pericchi and A. F. M. Smith. Exact and approximate posterior moments for a normal location parameter. Journal of the Royal Statistical Society. Series B (Methodological), 54(3): 793–804, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345859.

Zhao:1992:MMP

[2044] Lue Ping Zhao, Ross L. Prentice, and Steven G. Self. Multivariate mean parameter estimation by using a partly exponential model. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):805–811, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345860.

Vecchia:1992:NMP

[2045] A. V. Vecchia. A new method of prediction for spatial regression models with correlated errors. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):813–830, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345861.

Cuzick:1992:SAR

[2046] Jack Cuzick. Semiparametric additive regression. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):831–843, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345862.

Perron:1992:CSH

[2047] François Perron. Confidence sets having the shape of a half-space. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):845–852, ???? 1992. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345863.

Frydman:1992:NEP

[2048] Halina Frydman. A nonparametric estimation procedure for a periodically observed three-state Markov process, with application to aids. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):853–866, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345864.

Ghosh:1992:BFB

[2049] J. K. Ghosh and Rahul Mukerjee.

Bayesian and frequentist Bartlett corrections for likelihood ratio and conditional likelihood ratio tests. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):867–875, ???? 1992. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345865.

Cheng:1992:EMN

[2050] R. C. H. Cheng, B. E. Evans, and T. C. Iles. Embedded models in non-linear regression. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):877–888, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345866.

Polson:1992:EAI

[2051] Nicholas G. Polson. On the expected amount of information from a non-linear model. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):889–895, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345867.

Barndorff-Nielsen:1992:TSO

[2052] O. E. Barndorff-Nielsen and P. Blaesild. A type of second-order asymptotic independence. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):897–901, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345868.

Anonymous:1992:VI

[2053] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 54(3): 903–905, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345850.

Stone:1992:CCR

[2054] M. Stone and R. J. Brooks. Corrigendum: Continuum regression: Crossvalidated sequentially constructed prediction embracing ordinary least squares, partial least squares and principal components regression. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):906–907, ???? 1992. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345869. See [1919].

Hall:1992:CLM

[2055] Peter Hall and J. S. Marron. Corrigendum: Local minima in cross-validation functions. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):907, ???? 1992. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345870. See

[1952].

Anonymous:1992:BMc

[2056] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):??, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345871.

Anonymous:1992:FMc

[2057] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 54(3):??, ???? 1992. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345851.

Anonymous:1993:RE

[2058] Anonymous. Report of the Editors.

Journal of the Royal Statistical Society. Series B (Methodological), 55(1): 1, ???? 1993. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346062.

Smith:1993:BCG

[2059] A. F. M. Smith and G. O. Roberts.

Bayesian computation via the Gibbs sampler and related Markov chain Monte Carlo methods. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):3–23, ???? 1993. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346063. See discussion

[2062].

Besag:1993:SSB

[2060] Julian Besag and Peter J. Green. Spatial statistics and Bayesian computation. Journal of the Royal Statistical Society. Series B (Methodological), 55 (1):25–37, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346064.

Gilks:1993:MCA

[2061] W. R. Gilks, D. G. Clayton, D. J. Spiegelhalter, N. G. Best, and A. J. McNeil. Modelling complexity: Applications of Gibbs sampling in medicine. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):39–52, ???? 1993. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346065.

Anonymous:1993:DMG

[2062] Anonymous. Discussion on the meeting on the Gibbs sampler and other Markov chain Monte Carlo methods. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):53–102, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346066. See [2059].

Grunwald:1993:TSC

[2063] Gary K. Grunwald, Adrian E. Raftery, and Peter Guttorp. Time series of continuous proportions. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):103–116, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346067.

Fisher:1993:SMA

[2064] N. I. Fisher, A. D. Lunn, and S. J. Davies. Spherical median axes. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):117–124, ???? 1993. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346068.

Macaskill:1993:NAP

[2065] G. T. Macaskill. A note on adjusted profile likelihoods in non-linear regression. Journal of the Royal Statistical Society. Series B (Methodological), 55 (1):125–131, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346069.

McCabe:1993:TPV

[2066] B. P. M. McCabe and S. J. Leybourne.

Testing for parameter variation in nonlinear regression models. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):133–144, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346070.

Eubank:1993:DHN

[2067] R. L. Eubank and Will Thomas.

Detecting heteroscedasticity in nonparametric regression. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):145–155, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346071.

Chambers:1993:LLM

[2068] R. L. Chambers and A. H. Welsh.

Log-linear models for survey data with non-ignorable non-response. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):157–170, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346072.

Butler:1993:SAT

[2069] Ronald W. Butler, S. Huzurbazar, and James G. Booth. Saddlepoint approximations for tests of block independence, sphericity and equal variances and covariances. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):171–183, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346073.

Gelman:1993:CJP

[2070] Andrew Gelman and T. P. Speed. Characterizing a joint probability distribution by conditionals. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):185–188, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2346074. See corrigendum [2468].

Wells:1993:LST

[2071] Martin T. Wells, S. Rao Jammalamadaka, and Ram C. Tiwari. Large sample theory of spacings statistics for tests of fit for the composite hypothesis. Journal of the Royal Statistical Society. Series B (Methodological), 55 (1):189–203, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346075.

Frigessi:1993:CRG

[2072] Arnoldo Frigessi, Patrizia Di Stefano, Chii-Ruey Hwang, and Shuenn-Jyi Sheu. Convergence rates of the Gibbs sampler, the Metropolis algorithm and other single-site updating dynamics. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):

205–219, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346076.

Hooper:1993:NOR

[2073] Peter M. Hooper. Nearly orthogonal randomized designs. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):221–236, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346077.

Monod:1993:VRR

[2074] Herve Monod and R. A. Bailey.

Valid restricted randomization for unbalanced designs. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):237–251, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346078.

Vovk:1993:EVB

[2075] V. G. Vovk and V. V. V’Yugin. On the empirical validity of the Bayesian method. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):253–266, ???? 1993. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2346079.

Dabrowski:1993:CEP

[2076] Andre Robert Dabrowski and David McDonald. Consistent estimation in the presence of bandwidth limitation. Journal of the Royal Statistical Society. Series B (Methodological), 55(1): 267–289, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346080.

Hall:1993:EEB

[2077] Peter Hall. On Edgeworth expansion and bootstrap confidence bands in nonparametric curve estimation. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):291–304, ???? 1993. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346081.

DiCiccio:1993:SMS

[2078] Thomas J. DiCiccio and Michael A. Martin. Simple modifications for signed roots of likelihood ratio statistics. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):305–316, ???? 1993. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346082.

Anonymous:1993:BMa

[2079] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):??, ???? 1993. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346083.

Anonymous:1993:FMa

[2080] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 55(1):??, ???? 1993. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346061.

Vovk:1993:LPA

[2081] V. G. Vovk. A logic of probability, with application to the foundations of statistics. Journal of the Royal Statistical Society. Series B (Methodological), 55 (2):317–351, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346196.

Gu:1993:SAV

[2082] Chong Gu and Grace Wahba. Semiparametric analysis of variance with tensor product thin plate splines. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):353–368, ???? 1993. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346197.

Dunne:1993:DMP

[2083] T. T. Dunne and M. Stone. Downdating the Moore–Penrose generalized inverse for cross-validation of centred least squares prediction. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):369–375, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346198.

Nicolaou:1993:BIG

[2084] Anna Nicolaou. Bayesian intervals with good frequentist behaviour in the presence of nuisance parameters. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):377–390, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346199.

McDonald:1993:ELR

[2085] Barry W. McDonald. Estimating logistic regression parameters for bivariate binary data. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):391–397, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346200.

Nandram:1993:BPI

[2086] B. Nandram and J. Sedransk. Bayesian predictive inference for a finite population proportion: Two-stage cluster sampling. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):399–408, ???? 1993. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2346201.

West:1993:APD

[2087] Mike West. Approximating posterior distributions by mixture. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):409–422, ???? 1993. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346202.

Hill:1993:PMS

[2088] Bruce M. Hill. Parametric models for A_{n}: Splitting processes and mixtures. Journal of the Royal Statistical Society. Series B (Methodological), 55(2): 423–433, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2346203.

Mittnik:1993:CTA

[2089] Stefan Mittnik. Computing theoretical autocovariances of multivariate autoregressive moving average models by using a block Levinson method. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):435–440, ???? 1993. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346204.

Liang:1993:MMH

[2090] Kung-Yee Liang, Steven G. Self, and Yue-Cune Chang. Modelling marginal hazards in multivariate failure time data. Journal of the Royal Statistical Society. Series B (Methodological), 55 (2):441–453, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346205.

Stone:1993:AWC

[2091] Richard Stone. The assumptions on which causal inferences rest. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):455–466, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346206.

Cox:1993:NCA

[2092] D. R. Cox and N. Reid. A note on the calculation of adjusted profile likelihood. Journal of the Royal Statistical Society. Series B (Methodological), 55 (2):467–471, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346207.

Chen:1993:FMR

[2093] Zehua Chen. Fitting multivariate regression functions by interaction spline models. Journal of the Royal Statistical Society. Series B (Methodological), 55 (2):473–491, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346208.

Verbyla:1993:MVH

[2094] A. P. Verbyla. Modelling variance heterogeneity: Residual maximum likelihood and diagnostics. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):493–508, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346209.

OSullivan:1993:MDE

[2095] Finbarr O’Sullivan and Yudi Pawitan. Multidimensional density estimation by tomography. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):509–521, ???? 1993. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346210.

Diggle:1993:FAN

[2096] Peter J. Diggle and Peter Hall. A Fourier approach to nonparametric deconvolution of a density estimate. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):523–531, ???? 1993. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346211.

Severini:1993:BIE

[2097] Thomas A. Severini. Bayesian interval estimates which are also confidence intervals. Journal of the Royal Statistical Society. Series B (Methodological), 55 (2):533–540, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2346212.

Barbour:1993:CTN

[2098] A. D. Barbour and M. Costi. Correlation tests for non-linear alternatives. Journal of the Royal Statistical Society. Series B (Methodological), 55(2): 541–548, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2346213.

Azzalini:1993:UNR

[2099] Adelchi Azzalini and Adrian Bowman. On the use of nonparametric regression for checking linear relationships. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):549–557, ???? 1993. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346214.

Ljung:1993:ODT

[2100] Greta M. Ljung. On outlier detection in time series. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):559–567, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346215.

Anonymous:1993:BMb

[2101] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):??, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346216.

Anonymous:1993:FMb

[2102] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 55(2):??, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346195.

Vardi:1993:IDO

[2103] Y. Vardi and D. Lee. From image deblurring to optimal investments: Maximum likelihood solutions for positive linear inverse problems. Journal of the Royal Statistical Society. Series B (Methodological), 55(3):569–612, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345873.

Andrews:1993:TSC

[2104] David F. Andrews and James E. Stafford. Tools for the symbolic computation of asymptotic expansions. Journal of the Royal Statistical Society. Series B (Methodological), 55(3):613–627, ???? 1993. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345874.

Gamerman:1993:DHM

[2105] Dani Gamerman and Helio S. Migon. Dynamic hierarchical models. Journal of the Royal Statistical Society. Series B (Methodological), 55(3):629–642, ???? 1993. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345875.

Beran:1993:INP

[2106] Rudolf Beran and Peter Hall. Interpolated nonparametric prediction intervals and confidence intervals. Journal of the Royal Statistical Society. Series B (Methodological), 55(3):643–652, ???? 1993. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345876.

Sundberg:1993:CRR

[2107] Rolf Sundberg. Continuum regression and ridge regression. Journal of the Royal Statistical Society. Series B (Methodological), 55(3):653–659, ???? 1993. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345877.

Cordeiro:1993:IST

[2108] Gauss M. Cordeiro, Silvia L. De Paula Ferrari, and Gilberto A. Paula. Improved score tests for generalized linear models. Journal of the Royal Statistical Society. Series B (Methodological), 55 (3):661–674, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345878.

Girling:1993:SAA

[2109] A. J. Girling. Shape analysis for the anisotropic corpuscle problem. Journal of the Royal Statistical Society. Series B (Methodological), 55(3):675–686, ???? 1993. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345879.

Balasubramanian:1993:DPO

[2110] K. Balasubramanian and N. Balakrishnan. Duality principle in order statistics. Journal of the Royal Statistical Society. Series B (Methodological), 55 (3):687–691, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345880.

Carroll:1993:RLR

[2111] R. J. Carroll and Shane Pederson. On robustness in the logistic regression model. Journal of the Royal Statistical Society. Series B (Methodological), 55 (3):693–706, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345881.

Edmondson:1993:SRC

[2112] R. N. Edmondson. Systematic rowand-column designs balanced for low order polynomial interactions between rows and columns. Journal of the Royal Statistical Society. Series B (Methodological), 55(3):707–723, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345882.

Leurgans:1993:CCA

[2113] S. E. Leurgans, R. A. Moyeed, and B. W. Silverman. Canonical correlation analysis when the data are curves. Journal of the Royal Statistical Society. Series B (Methodological), 55(3): 725–740, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345883.

Geng:1993:SCA

[2114] Zhi Geng and Chooichiro Asano. Strong collapsibility of association measures in linear models. Journal of the Royal Statistical Society. Series B (Methodological), 55(3):741–747, ???? 1993. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345884.

Glonek:1993:BWS

[2115] G. F. V. Glonek. On the behaviour of Wald statistics for the disjunction of two regular hypotheses. Journal of the Royal Statistical Society. Series B (Methodological), 55(3):749–755, ???? 1993. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345885.

Anonymous:1993:BMc

[2116] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 55(3):??, ???? 1993. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345886.

Anonymous:1993:FMc

[2117] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 55(3):??, ????

1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345872.

Hastie:1993:VCM

[2118] Trevor Hastie and Robert Tibshirani.

Varying-coefficient models. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):757–796, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345993.

Coles:1993:RME

[2119] Stuart G. Coles. Regional modelling of extreme storms via max-stable processes. Journal of the Royal Statistical Society. Series B (Methodological), 55 (4):797–816, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345994.

Titterington:1993:CMR

[2120] D. M. Titterington. A contamination model and resistant estimation within the sampling paradigm. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):817–827, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345995.

Taplin:1993:RLC

[2121] Ross H. Taplin. Robust likelihood calculation for time series. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):829–836, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345996.

Lindley:1993:ALT

[2122] Dennis V. Lindley and Nozer D.

Singpurwalla. Adversarial life testing.

Journal of the Royal Statistical Society. Series B (Methodological), 55(4): 837–847, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345997.

Queen:1993:MDM

[2123] Catriona M. Queen and Jim Q. Smith. Multiregression dynamic models. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):849–870, ???? 1993. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345998.

Wong:1993:HOD

[2124] Weng Kee Wong and R. Dennis Cook. Heteroscedastic G-optimal designs. Journal of the Royal Statistical Society. Series B (Methodological), 55 (4):871–880, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345999.

Tatum:1993:HBM

[2125] Lawrence G. Tatum and Clifford M. Hurvich. High breakdown methods of time series analysis. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):881–896, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346000.

Kelly:1993:IEV

[2126] Robert J. Kelly and Thomas Mathew. Improved estimators of variance components with smaller probability of negativity. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):897–911, ???? 1993. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346001.

Gordon:1993:ANG

[2127] N. J. Gordon and A. F. M. Smith. Approximate non-Gaussian Bayesian estimation and modal consistency. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):913–918, ???? 1993. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346002.

Kunert:1993:EVC

[2128] J. Kunert and B. P. Utzig. Estimation of variance in crossover designs. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):919–927, ???? 1993. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346003.

Wu:1993:SOA

[2129] Xizhi Wu and Zhen Luo. Secondorder approach to local influence. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):929–936, ???? 1993. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346004. See corrigen-

dum [2499].

Grunwald:1993:PRE

[2130] Gary K. Grunwald, Peter Guttorp, and Adrian E. Raftery. Prediction rules for exponential family state space models. Journal of the Royal Statistical Society. Series B (Methodological), 55(4): 937–943, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346005.

Mak:1993:SNL

[2131] Tak K. Mak. Solving non-linear estimation equations. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):945–955, ???? 1993. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346006.

Anonymous:1993:VI

[2132] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 55(4): 957–960, ???? 1993. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345991.

Anonymous:1993:BMd

[2133] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):??, ???? 1993. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346007.

Anonymous:1993:FMd

[2134] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 55(4):??, ???? 1993. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345992.

Anonymous:1994:RE

[2135] Anonymous. Report of the Editors. Journal of the Royal Statistical Society. Series B (Methodological), 56(1): 1, ???? 1994. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346024.

Newton:1994:ABI

[2136] Michael A. Newton and Adrian E.

Raftery. Approximate Bayesian inference with the weighted likelihood bootstrap. Journal of the Royal Statistical Society. Series B (Methodological), 56 (1):3–48, ???? 1994. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346025.

Lindeboom:1994:HMB

[2137] Maarten Lindeboom and Gerard J. Van Den Berg. Heterogeneity in models for bivariate survival: The importance of the mixing distribution. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):49–60, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346026.

McGilchrist:1994:EGM

[2138] C. A. McGilchrist. Estimation in generalized mixed models. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):61–69, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346027.

Frydman:1994:NNE

[2139] Halina Frydman. A note on nonparametric estimation of the distribution function from interval-censored and truncated observations. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):71–74, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346028.

Renshaw:1994:LST

[2140] Eric Renshaw. The linear spatialtemporal interaction process and its relation to 1/ω-noise. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):75–91, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346029.

Mangat:1994:IRR

[2141] N. S. Mangat. An improved randomized response strategy. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):93–95, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346030.

Constantine:1994:CSS

[2142] A. G. Constantine and Peter Hall. Characterizing surface smoothness via estimation of effective fractal dimension. Journal of the Royal Statistical Society. Series B (Methodological), 56 (1):97–113, ???? 1994. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2346031.

Vovk:1994:PLI

[2143] V. G. Vovk and V. V. V’Yugin. Prequential level of impossibility with some applications. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):115–123, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346032.

Barndorff-Nielsen:1994:AVP

[2144] O. E. Barndorff-Nielsen. Adjusted versions of profile likelihood and directed likelihood, and extended likelihood. Journal of the Royal Statistical Society. Series B (Methodological), 56 (1):125–140, ???? 1994. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346033.

Belcher:1994:PCT

[2145] J. Belcher, J. S. Hampton, and G. Tunnicliffe Wilson. Parameterization of continuous time autoregressive models for irregularly sampled time series data. Journal of the Royal Statistical Society. Series B (Methodological), 56(1): 141–155, ???? 1994. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346034.

Fosam:1994:CCE

[2146] E. B. Fosam and D. N. Shanbhag.

Certain characterizations of exponential and geometric distributions. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):157–160, ???? 1994. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346035.

Gerami:1994:CRD

[2147] A. Gerami and S. M. Lewis. Completely randomized designs for comparing dual with single treatments. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):161–165, ???? 1994. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346036.

Cheng:1994:ELR

[2148] Chi-Lun Cheng and John W. Van Ness. On estimating linear relationships when both variables are subject to errors. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):167–183, ???? 1994. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2346037.

Ghosh:1994:AVC

[2149] J. K. Ghosh and Rahul Mukerjee. Adjusted versus conditional likelihood: Power properties and Bartlett-type adjustment. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):185–188, ???? 1994. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2346038.

McDougall:1994:RMR

[2150] A. J. McDougall. Robust methods for recursive autoregressive moving average estimation. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):189–207, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346039.

Naranjo:1994:BIR

[2151] J. D. Naranjo and T. P. Hettmansperger. Bounded influence rank regression. Journal of the Royal Statistical Society. Series B (Methodological), 56(1): 209–220, ???? 1994. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2346040.

Hettmansperger:1994:AIMa

[2152] Thomas P. Hettmansperger, Jukka Nyblom, and Hannu Oja. Affine invariant multivariate one-sample sign tests. Journal of the Royal Statistical Society. Series B (Methodological), 56(1): 221–234, ???? 1994. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2346041.

Hettmansperger:1994:AIMb

[2153] Thomas P. Hettmansperger and Hannu Oja. Affine invariant multivariate multisample sign tests. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):235–249, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346042.

Schabe:1994:NEC

[2154] Hendrik Schabe. Nonparametric estimation of component lifetime based on masked system life test data. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):251–259, ???? 1994. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346043.

Geyer:1994:CMC

[2155] Charles J. Geyer. On the convergence of Monte Carlo maximum likelihood calculations. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):261–274, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346044.

Brown:1994:GCC

[2156] B. M. Brown. Grouping corrections for circular goodness-of-fit tests. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):275–283, ???? 1994. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346045.

Anonymous:1994:BMa

[2157] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):??, ????

1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346046.

Anonymous:1994:FMa

[2158] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 56(1):??, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346023.

Kent:1994:CBD

[2159] John T. Kent. The complex Bingham distribution and shape analysis. Journal of the Royal Statistical Society. Series B (Methodological), 56(2):285–299, ???? 1994. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345900.

Chan:1994:NNC

[2160] K. S. Chan and H. Tong. A note on noisy chaos. Journal of the Royal Statistical Society. Series B (Methodological), 56(2):301–311, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345901.

Ronchetti:1994:ESA

[2161] Elvezio Ronchetti and A. H. Welsh. Empirical saddlepoint approximations for multivariate M-estimators. Journal of the Royal Statistical Society. Series B (Methodological), 56(2):313–326, ???? 1994. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345902.

Fisher:1994:TSA

[2162] N. I. Fisher and A. J. Lee. Time series analysis of circular data. Journal of the Royal Statistical Society. Series B (Methodological), 56(2):327–339, ???? 1994. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345903.

Lind:1994:ITM

[2163] Niels C. Lind. Information theory and maximum product of spacings estimation. Journal of the Royal Statistical Society. Series B (Methodological), 56 (2):341–343, ???? 1994. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345904.

Segal:1994:VMP

[2164] Mark R. Segal, Peter Bacchetti, and Nicholas P. Jewell. Variances for maximum penalized likelihood estimates obtained via the EM algorithm. Journal of the Royal Statistical Society. Series B (Methodological), 56(2):345–352, ???? 1994. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345905.

Severini:1994:NCI

[2165] Thomas A. Severini. Nonparametric conditional inference for a location parameter. Journal of the Royal Statistical Society. Series B (Methodological), 56(2):353–362, ???? 1994. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345906.

Diebolt:1994:EFM

[2166] Jean Diebolt and Christian P. Robert.

Estimation of finite mixture distributions through Bayesian sampling. Journal of the Royal Statistical Society. Series B (Methodological), 56(2):363–375, ???? 1994. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345907.

Roberts:1994:GCG

[2167] Gareth O. Roberts and Nicholas G. Polson. On the geometric convergence of the Gibbs sampler. Journal of the Royal Statistical Society. Series B (Methodological), 56(2):377–384, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345908.

Catchpole:1994:BER

[2168] E. A. Catchpole and B. J. T. Morgan. Boundary estimation in ring recovery models. Journal of the Royal Statistical Society. Series B (Methodological), 56 (2):385–391, ???? 1994. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2345909.

Hadi:1994:MMD

[2169] Ali S. Hadi. A modification of a method for the detection of outliers in multivariate samples. Journal of the Royal Statistical Society. Series B (Methodological), 56(2):393–396, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345910.

DiCiccio:1994:FBB

[2170] Thomas J. DiCiccio and Steven E. Stern. Frequentist and Bayesian Bartlett correction of test statistics based on adjusted profile likelihoods. Journal of the Royal Statistical Society. Series B (Methodological), 56(2): 397–408, ???? 1994. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345911.

Anonymous:1994:BMb

[2171] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 56(2):??, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345912.

Anonymous:1994:FMb

[2172] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 56(2):??, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345899.

Ripley:1994:NNR

[2173] B. D. Ripley. Neural networks and related methods for classification. Journal of the Royal Statistical Society. Series B (Methodological), 56(3):409–456, ???? 1994. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346118.

Lewis:1994:MDQ

[2174] S. M. Lewis, A. M. Dean, N. R. Draper, and P. Prescott. Mixture designs for q components in orthogonal blocks. Journal of the Royal Statistical Society. Series B (Methodological), 56(3):457–467, ???? 1994. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346119.

Pearce:1994:RL

[2175] S. C. Pearce. Reinforced lattices. Journal of the Royal Statistical Society. Series B (Methodological), 56(3):469–476, ???? 1994. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346120.

Maguluri:1994:EMR

[2176] Gangaji Maguluri and Cun-Hui Zhang. Estimation in the mean residual life regression model. Journal of the Royal Statistical Society. Series B (Methodological), 56(3):477–489, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346121.

Lockhart:1994:ETF

[2177] Richard A. Lockhart and Michael A. Stephens. Estimation and tests of fit for the three-parameter Weibull distribution. Journal of the Royal Statistical Society. Series B (Methodological), 56 (3):491–500, ???? 1994. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2346122.

Gelfand:1994:BMC

[2178] A. E. Gelfand and D. K. Dey. Bayesian model choice: Asymptotics and exact calculations. Journal of the Royal Statistical Society. Series B (Methodological), 56(3):501–514, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346123.

Smith:1994:EEI

[2179] Richard L. Smith and Ishay Weissman. Estimating the extremal index. Journal of the Royal Statistical Society. Series B (Methodological), 56(3):515–528, ???? 1994. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346124.

Hart:1994:AKS

[2180] Jeffrey D. Hart. Automated kernel smoothing of dependent data by using time series cross- validation. Journal of the Royal Statistical Society. Series B (Methodological), 56(3):529–542, ???? 1994. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346125.

Smith:1994:ICB

[2181] J. Q. Smith. The inadmissibility of certain Bayes decision rules under vague priors in location problems. Journal of the Royal Statistical Society. Series B (Methodological), 56(3):543–548, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346126.

Anonymous:1994:BMc

[2182] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 56(3):??, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346127.

Anonymous:1994:FMc

[2183] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 56(3):??, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346117.

Grenander:1994:RKC

[2184] Ulf Grenander and Michael I. Miller.

Representations of knowledge in complex systems. Journal of the Royal Statistical Society. Series B (Methodological), 56(4):549–603, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346184.

Mitchell:1994:NPM

[2185] Ann F. S. Mitchell. A note on posterior moments for a normal mean with double-exponential prior. Journal of the Royal Statistical Society. Series B (Methodological), 56(4):605–610, ???? 1994. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346185. See corrigen-

dum [2224].

Edmondson:1994:FFD

[2186] R. N. Edmondson. Fractional factorial designs for factors with a prime number of quantitative levels. Journal of the Royal Statistical Society. Series B (Methodological), 56(4):611–622, ???? 1994. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346186.

Rudas:1994:NIF

[2187] Tamas Rudas, Clifford C. Clogg, and Bruce G. Lindsay. A new index of fit based on mixture methods for the analysis of contingency tables. Journal of the Royal Statistical Society. Series B (Methodological), 56(4):623–639, ???? 1994. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346187.

Rosenbaum:1994:DEF

[2188] Paul R. Rosenbaum. Dispersion effects from fractional factorials in Taguchi’s method of quality design. Journal of the Royal Statistical Society. Series B (Methodological), 56(4):641–652, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346188.

Marron:1994:TRB

[2189] J. S. Marron and D. Ruppert. Transformations to reduce boundary bias in kernel density estimation. Journal of the Royal Statistical Society. Series B (Methodological), 56(4):653–671, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346189.

Billard:1994:MSN

[2190] L. Billard and Zhen Zhao. Multiplestage non-homogeneous Markov models for the acquired immune deficiency syndrome epidemic. Journal of the Royal Statistical Society. Series B (Methodological), 56(4):673–686, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346190.

Chen:1994:IPF

[2191] Jiahua Chen and J. D. Kalbfleisch. Inverse problems in fractal construction: Hellinger distance method. Journal of the Royal Statistical Society. Series B (Methodological), 56(4):687–700, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346191.

Yao:1994:QII

[2192] Qiwei Yao and Howell Tong. Quantifying the influence of initial values on non-linear prediction. Journal of the Royal Statistical Society. Series B (Methodological), 56(4):701–725, ???? 1994. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346192.

Farebrother:1994:CD

[2193] Richard William Farebrother. The Cholesky decomposition of P − pp′. Journal of the Royal Statistical Society. Series B (Methodological), 56(4): 727, ???? 1994. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

                                        jstor.org/stable/2346193.

Anonymous:1994:VI

[2194] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 56(4): 729–731, ???? 1994. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2346182.

Anonymous:1994:BMd

[2195] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 56(4):??, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346194.

Anonymous:1994:FMd

[2196] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 56(4):??, ???? 1994. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346183.

Anonymous:1995:RE

[2197] Anonymous. Report of the Editors: 1994. Journal of the Royal Statistical Society. Series B (Methodological), 57 (1):1, ???? 1995. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

                                        jstor.org/stable/2346085.

Cheng:1995:NRM

[2198] R. C. H. Cheng and L. Traylor. Nonregular maximum likelihood problems. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):3–44, ???? 1995. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346086.

Draper:1995:APM

[2199] David Draper. Assessment and propagation of model uncertainty. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):45–97, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346087.

OHagan:1995:FBF

[2200] Anthony O’Hagan. Fractional Bayes factors for model comparison. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):99–138, ???? 1995. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346088.

Bhaumik:1995:MOC

[2201] Dulal K. Bhaumik. Majorization and D-optimality for complete block designs under correlations. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):139–143, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346089.

Pena:1995:DIS

[2202] Daniel Pena and Victor J. Yohai. The detection of influential subsets in linear regression by using an influence matrix. Journal of the Royal Statistical Society. Series B (Methodological), 57 (1):145–156, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346090. See

corrigendum [2236].

Liu:1995:CSC

[2203] Jun S. Liu, Wing H. Wong, and Augustine Kong. Covariance structure and convergence rate of the Gibbs sampler with various scans. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):157–169, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346091.

Herrmann:1995:BSB

[2204] Eva Herrmann, Joachim Engel, M. P. Wand, and Theo Gasser. A bandwidth selector for bivariate kernel regression. Journal of the Royal Statistical Society. Series B (Methodological), 57(1): 171–180, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346092.

Lawrence:1995:DIM

[2205] A. J. Lawrence. Deletion influence and masking in regression. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):181–189, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346093.

Laycock:1995:MGL

[2206] P. J. Laycock and P. J. Rowley. A method for generating and labelling all regular fractions or blocks for q^{n−m}- designs. Journal of the Royal Statistical Society. Series B (Methodological), 57 (1):191–204, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346094.

Prum:1995:FWU

[2207] Bernard Prum, François Rodolphe, and Elisabeth de Turckheim. Finding words with unexpected frequencies in deoxyribonucleic acid sequences. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):205–220, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346095.

Copas:1995:LLB

[2208] J. B. Copas. Local likelihood based on kernel censoring. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):221–235, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346096.

Fruhwirth-Schnatter:1995:BMD

[2209] Sylvia Fruhwirth-Schnatter. Bayesian model discrimination and Bayes factors for linear Gaussian state space models. Journal of the Royal Statistical Society. Series B (Methodological), 57(1): 237–246, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346097.

Laud:1995:PMS

[2210] Purushottam W. Laud and Joseph G.

Ibrahim. Predictive model selection. Journal of the Royal Statistical Society. Series B (Methodological), 57(1): 247–262, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346098.

Guo:1995:CLR

[2211] Jianhua Guo and Zhi Geng. Collapsibility of logistic regression coefficients. Journal of the Royal Statistical Society. Series B (Methodological), 57(1): 263–267, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2346099.

Ball:1995:SIT

[2212] F. G. Ball and S. S. Davies. Statistical inference for a two-state Markov model of a single ion channel, incorporating time interval omission. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):269–287, ???? 1995. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346100.

Benjamini:1995:CFD

[2213] Yoav Benjamini and Yosef Hochberg. Controlling the false discovery rate: A practical and powerful approach to multiple testing. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):289–300, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346101.

Anonymous:1995:BMa

[2214] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):??, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346102.

Anonymous:1995:FMa

[2215] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 57(1):??, ????

1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346084.

Donoho:1995:WSA

[2216] David L. Donoho, Iain M. Johnstone, Gerard Kerkyacharian, and Dominique Picard. Wavelet shrinkage: Asymptopia? Journal of the Royal Statistical Society. Series B (Methodological), 57 (2):301–369, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345967.

Fan:1995:DDB

[2217] Jianqing Fan and Irene Gijbels. Datadriven bandwidth selection in local polynomial fitting: Variable bandwidth and spatial adaptation. Journal of the Royal Statistical Society. Series B (Methodological), 57(2):371–394, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345968.

Kuk:1995:AUE

[2218] Anthony Y. C. Kuk. Asymptotically unbiased estimation in generalized linear models with random effects. Journal of the Royal Statistical Society. Series B (Methodological), 57(2):395–407, ???? 1995. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345969.

Robins:1995:SEE

[2219] James M. Robins, Fushing Hsieh, and Whitney Newey. Semiparametric efficient estimation of a conditional density with missing or mismeasured covariates. Journal of the Royal Statistical Society. Series B (Methodological), 57(2):409–424, ???? 1995. CODEN JSTBAJ. ISSN 0035-9246. URL http:

                                        //www.jstor.org/stable/2345970.

Lange:1995:GAL

[2220] Kenneth Lange. A gradient algorithm locally equivalent to the EM algorithm. Journal of the Royal Statistical Society. Series B (Methodological), 57(2): 425–437, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345971.

Hall:1995:PID

[2221] Peter Hall and Rodney C. L. Wolff. Properties of invariant distributions and Lyapunov exponents for chaotic logistic maps. Journal of the Royal Statistical Society. Series B (Methodological), 57(2):439–452, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345972.

Mak:1995:RMI

[2222] Tak K. Mak. Regression models with incidental parameters. Journal of the Royal Statistical Society. Series B (Methodological), 57(2):453–457, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345973.

Bailey:1995:RBS

[2223] R. A. Bailey, D. A. Preece, and C. A. Rowley. Randomization for a balanced superimposition of one Youden square on another. Journal of the Royal Statistical Society. Series B (Methodological), 57(2):459–469, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345974.

Mitchell:1995:CNP

[2224] Ann F. S. Mitchell. Corrigendum: A note on posterior moments for a normal mean with double-exponential prior. Journal of the Royal Statistical Society. Series B (Methodological), 57(2):471, ???? 1995. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345975. See [2185].

Anonymous:1995:BMb

[2225] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 57(2):??, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345976.

Anonymous:1995:FMb

[2226] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 57(2):??, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345966.

Carlin:1995:BMC

[2227] Bradley P. Carlin and Siddhartha Chib. Bayesian model choice via Markov chain Monte Carlo methods. Journal of the Royal Statistical Society. Series B (Methodological), 57(3): 473–484, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346151.

Ferrari:1995:FAM

[2228] Pablo A. Ferrari, Arnoldo Frigessi, and Paula Gonzaga de Sa. Fast approximate maximum a posteriori restoration of multicolour images. Journal of the Royal Statistical Society. Series B (Methodological), 57(3):485–500, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346152.

Granville:1995:MDA

[2229] Vincent Granville and Jean Paul Rasson. Multivariate discriminant analysis and maximum penalized likelihood density estimation. Journal of the Royal Statistical Society. Series B (Methodological), 57(3):501–517, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346153.

Nychka:1995:NTB

[2230] Douglas Nychka and David Ruppert. Nonparametric transformations for both sides of a regression model. Journal of the Royal Statistical Society. Series B (Methodological), 57(3): 519–532, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346154.

Glonek:1995:MLM

[2231] G. F. V. Glonek and P. McCullagh. Multivariate logistic models. Journal of the Royal Statistical Society. Series B (Methodological), 57(3):533–546, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346155.

Damien:1995:ARV

[2232] Paul Damien, Purushottam W. Laud, and Adrian F. M. Smith. Approximate random variate generation from infinitely divisible distributions with applications to Bayesian inference. Journal of the Royal Statistical Society. Series B (Methodological), 57(3):547–563, ???? 1995. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346156.

Niinimaa:1995:IFC

[2233] A. Niinimaa and H. Oja. On the influence functions of certain bivariate medians. Journal of the Royal Statistical Society. Series B (Methodological), 57 (3):565–574, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346157.

Haines:1995:GAO

[2234] Linda M. Haines. A geometric approach to optimal design for oneparameter non-linear models. Journal of the Royal Statistical Society. Series B (Methodological), 57(3):575–598, ???? 1995. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346158.

Windham:1995:RMF

[2235] Michael P. Windham. Robustifying model fitting. Journal of the Royal Statistical Society. Series B (Methodological), 57(3):599–609, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346159.

Pena:1995:CDI

[2236] Daniel Pena and Victor J. Yohai. Corrigendum: The detection of influential subsets in linear regression by using an influence matrix. Journal of the Royal Statistical Society. Series B (Methodological), 57(3):611, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346160. See [2202].

Anonymous:1995:BMc

[2237] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 57(3):??, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346161.

Anonymous:1995:FMc

[2238] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 57(3):??, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346150.

Lai:1995:SCD

[2239] Tze Leung Lai. Sequential changepoint detection in quality control and dynamical systems. Journal of the Royal Statistical Society. Series B (Methodological), 57(4):613–658, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345934.

Beran:1995:MLE

[2240] Jan Beran. Maximum likelihood estimation of the differencing parameter for invertible short and long memory autoregressive integrated moving average models. Journal of the Royal Statistical Society. Series B (Methodological), 57(4):659–672, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345935.

Silverman:1995:IPE

[2241] B. W. Silverman. Incorporating parametric effects into functional principal components analysis. Journal of the Royal Statistical Society. Series B (Methodological), 57(4):673–689, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345936.

Fitzmaurice:1995:RML

[2242] Garrett M. Fitzmaurice, Geert Molenberghs, and Stuart R. Lipsitz. Regression models for longitudinal binary responses with informative drop- outs. Journal of the Royal Statistical Society. Series B (Methodological), 57(4): 691–704, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345937.

Glazebrook:1995:CBM

[2243] K. D. Glazebrook and R. J. Boys. A class of Bayesian models for optimal exploration. Journal of the Royal Statistical Society. Series B (Methodological), 57(4):705–720, ???? 1995. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345938.

Bofinger:1995:ERC

[2244] Eve Bofinger and Mark Bofinger.

Equivalence with respect to a control: Stepwise tests. Journal of the Royal Statistical Society. Series B (Methodological), 57(4):721–733, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345939.

Hsieh:1995:EPA

[2245] Fushing Hsieh. The empirical process approach for semiparametric twosample models with heterogeneous treatment effect. Journal of the Royal Statistical Society. Series B (Methodological), 57(4):735–748, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345940.

Shun:1995:LAH

[2246] Zhenming Shun and Peter McCullagh. Laplace approximation of high dimensional integrals. Journal of the Royal Statistical Society. Series B (Methodological), 57(4):749–760, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345941.

Ghosh:1995:PEH

[2247] J. K. Ghosh and Rahul Mukerjee. On perturbed ellipsoidal and highest posterior density regions with approximate frequentist validity. Journal of the Royal Statistical Society. Series B (Methodological), 57(4):761–769, ???? 1995. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345942.

Ginebra:1995:RSB

[2248] Josep Ginebra and Murray K. Clayton. Response surface bandits. Journal of the Royal Statistical Society. Series B (Methodological), 57(4):771–784, ???? 1995. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345943.

Anonymous:1995:VI

[2249] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 57(4): 785–787, ???? 1995. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345932.

Anonymous:1995:BMd

[2250] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 57(4):??, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345944.

Anonymous:1995:FMd

[2251] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 57(4):??, ???? 1995. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345933.

Anonymous:1996:RE

[2252] Anonymous. Report of the Editor:

1995. Journal of the Royal Statistical Society. Series B (Methodological), 58 (1):1, ???? 1996. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

jstor.org/stable/2346163.

Walley:1996:IMD

[2253] Peter Walley. Inferences from multinomial data: Learning about a bag of marbles. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):3–57, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346164.

Atkinson:1996:UOE

[2254] A. C. Atkinson. The usefulness of optimum experimental designs. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):59–76, ???? 1996. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346165. See discussion

[2256].

Bates:1996:EDO

[2255] R. A. Bates, R. J. Buck, E. Riccomagno, and H. P. Wynn. Experimental design and observation for large systems. Journal of the Royal Statistical Society. Series B (Methodological), 58 (1):77–94, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346166. See

discussion [2256].

Anonymous:1996:DPA

[2256] Anonymous. Discussion of the papers by Atkinson, and Bates et al. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):95–111, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2346167. See [2254, 2255].

Plummer:1996:EPE

[2257] Martyn Plummer and David Clayton. Estimation of population exposure in ecological studies. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):113–126, ???? 1996. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346168. See discussion

[2259].

Festing:1996:RUL

[2258] Michael F. W. Festing and David P. Lovell. Reducing the use of laboratory animals in toxicological research and testing by better experimental design. Journal of the Royal Statistical Society. Series B (Methodological), 58 (1):127–140, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346169. See

discussion [2259].

Anonymous:1996:DPP

[2259] Anonymous. Discussion of the papers by Plummer and Clayton, and Festing and Lovell. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):141–153, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2346170. See [2258, 2257].

Hastie:1996:DAG

[2260] Trevor Hastie and Robert Tibshirani.

Discriminant analysis by Gaussian mixtures. Journal of the Royal Statistical Society. Series B (Methodological), 58 (1):155–176, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346171.

Waterman:1996:SAM

[2261] Richard P. Waterman and Bruce G.

Lindsay. A simple and accurate method for approximate conditional inference applied to exponential family models. Journal of the Royal Statistical Society. Series B (Methodological), 58(1): 177–188, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346172.

DiCiccio:1996:IBA

[2262] Thomas J. DiCiccio, Michael A. Martin, Steven E. Stern, and G. Alastair Young. Information bias and adjusted profile likelihoods. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):189–203, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346173.

Rochon:1996:ACO

[2263] James Rochon. Accounting for covariates observed post randomization for discrete and continuous repeated measures data. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):205–219, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346174.

Csorgo:1996:EAD

[2264] Sandor Csorgo and Julian J. Faraway. The exact and asymptotic distributions of Cramér–von Mises statistics. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):221–234, ???? 1996. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346175.

Lee:1996:SIB

[2265] Stephen M. S. Lee and G. Alastair Young. Sequential iterated bootstrap confidence intervals. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):235–251, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346176.

Lang:1996:CMP

[2266] Joseph B. Lang. On the comparison of multinomial and Poisson log-linear models. Journal of the Royal Statistical Society. Series B (Methodological), 58 (1):253–266, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http://

                                        www.jstor.org/stable/2346177.

Tibshirani:1996:RSS

[2267] Robert Tibshirani. Regression shrinkage and selection via the lasso. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):267–288, ???? 1996. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346178.

Watson:1996:SDC

[2268] Geoffrey S. Watson. Spectral decomposition of the covariance matrix of a multinomial. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):289–291, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346179.

Cocchi:1996:QLB

[2269] Daniela Cocchi and Michel Mouchart.

Quasi-linear Bayes estimation in stratified finite populations. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):293–300, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346180.

Anonymous:1996:BMa

[2270] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):??, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346181.

Anonymous:1996:FMa

[2271] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 58(1):??, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346162.

Goldstein:1996:BLS

[2272] M. Goldstein and A. O’Hagan. Bayes linear sufficiency and systems of expert posterior assessments. Journal of the Royal Statistical Society. Series B (Methodological), 58(2):301–316, ????

1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345978.

Heyde:1996:QLG

[2273] C. C. Heyde and R. Morton. Quasilikelihood and generalizing the EM algorithm. Journal of the Royal Statistical Society. Series B (Methodological), 58(2):317–327, ???? 1996. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345979.

Coles:1996:MEA

[2274] Stuart G. Coles and Jonathan A. Tawn. Modelling extremes of the areal rainfall process. Journal of the Royal Statistical Society. Series B (Methodological), 58 (2):329–347, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345980.

Farrington:1996:AGF

[2275] C. P. Farrington. On assessing goodness of fit of generalized linear models to sparse data. Journal of the Royal Statistical Society. Series B (Methodological), 58(2):349–360, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345981.

Hall:1996:CSP

[2276] Peter Hall and Prakash Patil. On the choice of smoothing parameter, threshold and truncation in nonparametric regression by non-linear wavelet methods. Journal of the Royal Statistical Society. Series B (Methodological), 58 (2):361–377, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345982.

Hastie:1996:P

[2277] Trevor Hastie. Pseudosplines. Journal of the Royal Statistical Society. Series B (Methodological), 58(2):379–396, ???? 1996. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345983.

Chen:1996:PUM

[2278] Rong Chen and Jun S. Liu. Predictive updating methods with application to Bayesian classification. Journal of the Royal Statistical Society. Series B (Methodological), 58(2):397–415, ???? 1996. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345984.

Crowder:1996:STB

[2279] Martin Crowder. Some tests based on extreme values for a parametric survival model. Journal of the Royal Statistical Society. Series B (Methodological), 58(2):417–424, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345985.

Dawid:1996:CAF

[2280] A. P. Dawid and J. Mortera. Coherent analysis of forensic identification evidence. Journal of the Royal Statistical Society. Series B (Methodological), 58 (2):425–443, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2345986.

Forster:1996:MCE

[2281] Jonathan J. Forster, John W. McDonald, and Peter W. F. Smith. Monte Carlo exact conditional tests for loglinear and logistic models. Journal of the Royal Statistical Society. Series B (Methodological), 58(2):445–453, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345987.

Liu:1996:MTN

[2282] Wei Liu. Multiple tests of a nonhierarchical finite family of hypotheses. Journal of the Royal Statistical Society. Series B (Methodological), 58(2): 455–461, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345988.

Nason:1996:WSU

[2283] G. P. Nason. Wavelet shrinkage using cross-validation. Journal of the Royal Statistical Society. Series B (Methodological), 58(2):463–479, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345989.

Anonymous:1996:BMb

[2284] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 58(2):??, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345990.

Anonymous:1996:FMb

[2285] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 58(2):??, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345977.

Yee:1996:VGA

[2286] T. W. Yee and C. J. Wild. Vector generalized additive models. Journal of the Royal Statistical Society. Series B (Methodological), 58(3):481–493, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345888.

Ramsay:1996:PDA

[2287] J. O. Ramsay. Principal differential analysis: Data reduction by differential operators. Journal of the Royal Statistical Society. Series B (Methodological), 58(3):495–508, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345889.

Sherman:1996:VES

[2288] Michael Sherman. Variance estimation for statistics computed from spatial lattice data. Journal of the Royal Statistical Society. Series B (Methodological), 58(3):509–523, ???? 1996. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345890.

Yue:1996:AES

[2289] Huibin Yue and K. S. Chan. Asymptotic efficiency of the sample mean in Markov chain Monte Carlo schemes. Journal of the Royal Statistical Society. Series B (Methodological), 58(3): 525–539, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345891.

Rabinowitz:1996:RDC

[2290] Daniel Rabinowitz and Nicholas P.

Jewell. Regression with doubly censored current status data. Journal of the Royal Statistical Society. Series B (Methodological), 58(3):541–550, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345892.

Cowling:1996:PMR

[2291] Ann Cowling and Peter Hall. On pseudodata methods for removing boundary effects in kernel density estimation. Journal of the Royal Statistical Society. Series B (Methodological), 58(3): 551–563, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345893.

Smyth:1996:CLA

[2292] Gordon K. Smyth and Arunas P. Verbyla. A conditional likelihood approach to residual maximum likelihood estimation in generalized linear models. Journal of the Royal Statistical Society. Series B (Methodological), 58(3):565–572, ???? 1996. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345894.

Jorgensen:1996:LGC

[2293] Bent Jorgensen, Rodrigo Labouriau, and Soren Lundbye-Christensen. Linear growth curve analysis based on exponential dispersion models. Journal of the Royal Statistical Society. Series B (Methodological), 58(3):573–592, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2345895.

Taylor:1996:CAP

[2294] Jeremy M. G. Taylor, Arminda L. Siqueira, and Robert T. Weiss. The cost of adding parameters to a model. Journal of the Royal Statistical Society. Series B (Methodological), 58(3): 593–607, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2345896.

Feng:1996:UBL

[2295] Z. D. Feng and C. E. McCulloch.

Using bootstrap likelihood ratios in finite mixture models. Journal of the Royal Statistical Society. Series B (Methodological), 58(3):609–617, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345897.

Anonymous:1996:BMc

[2296] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 58(3):??, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345898.

Anonymous:1996:FMc

[2297] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 58(3):??, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345887.

Lee:1996:HGL

[2298] Y. Lee and J. A. Nelder. Hierarchical generalized linear models. Journal of the Royal Statistical Society. Series B (Methodological), 58(4):619–678, ???? 1996. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346105.

Young:1996:MDB

[2299] K. D. S. Young and L. I. Pettit. Measuring discordancy between prior and data. Journal of the Royal Statistical Society. Series B (Methodological), 58 (4):679–689, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346106.

Ghosh:1996:NGT

[2300] Sucharita Ghosh. A new graphical tool to detect non-normality. Journal of the Royal Statistical Society. Series B (Methodological), 58(4):691–702, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346107.

Bjorkstrom:1996:CRA

[2301] Anders Bjorkstrom and Rolf Sundberg. Continuum regression is not always continuous. Journal of the Royal Statistical Society. Series B (Methodological), 58(4):703–710, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346108.

Wild:1996:AEG

[2302] C. J. Wild and T. W. Yee. Additive extensions to generalized estimation equation methods. Journal of the Royal Statistical Society. Series B (Methodological), 58(4):711–725, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346109.

Hall:1996:SRM

[2303] Peter Hall and Bingyi Jing. On sample reuse methods for dependent data. Journal of the Royal Statistical Society. Series B (Methodological), 58(4): 727–737, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2346110.

Weiss:1996:ABS

[2304] Robert Weiss. An approach to Bayesian sensitivity analysis. Journal of the Royal Statistical Society. Series B (Methodological), 58(4):739–750, ???? 1996. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346111.

Rhodes:1996:CPM

[2305] Philip H. Rhodes, M. Elizabeth Halloran, and Ira M. Longini, Jr. Counting process models for infectious disease data: Distinguishing exposure to infection from susceptibility. Journal of the Royal Statistical Society. Series B (Methodological), 58(4):751–762, ???? 1996. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346112.

Catchpole:1996:SPR

[2306] E. A. Catchpole, S. N. Freeman, and B. J. T. Morgan. Steps to parameter redundancy in age-dependent recovery models. Journal of the Royal Statistical Society. Series B (Methodological), 58 (4):763–774, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346113.

Phillips:1996:ECE

[2307] M. J. Phillips and T. J. Sweeting. Estimation for censored exponential data when the censoring times are subject to error. Journal of the Royal Statistical Society. Series B (Methodological), 58 (4):775–783, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346114.

Liang:1996:ABP

[2308] Kung-Yee Liang and Steven G. Self. On the asymptotic behaviour of the pseudolikelihood ratio test statistic. Journal of the Royal Statistical Society. Series B (Methodological), 58(4):785–796, ???? 1996. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2346115.

Anonymous:1996:VI

[2309] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Methodological), 58(4): 797–799, ???? 1996. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2346103.

Anonymous:1996:BMd

[2310] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 58(4):??, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346116.

Anonymous:1996:FMd

[2311] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 58(4):??, ???? 1996. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346104.

Jones:1997:ERE

[2312] M. C. Jones and G. A. Young. Editorial: Report of the Editors — 1996. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):1– 2, ???? 1997. CODEN JSTBAJ. ISSN

                                        0035-9246. URL http://www.jstor.
                                        org/stable/2345914.

Breiman:1997:PMR

[2313] Leo Breiman and Jerome H. Friedman. Predicting multivariate responses in multiple linear regression. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):3–54, ????

1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345915.

Copas:1997:INR

[2314] J. B. Copas and H. G. Li. Inference for non-random samples. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):55–95, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345916.

Dette:1997:DER

[2315] Holger Dette. Designing experiments with respect to ‘standardized’ optimality criteria. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):97–110, ???? 1997. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345917.

Atkinson:1997:DRT

[2316] Anthony C. Atkinson and R. Dennis Cook. Designing for a response transformation parameter. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):111–124, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345918.

Hutton:1997:POM

[2317] J. L. Hutton and P. J. Solomon. Parameter orthogonality in mixed regression models for survival data. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):125–136, ???? 1997. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345919.

Cheng:1997:CRF

[2318] R. C. H. Cheng and W. B. Liu. A continuous representation of the family of stable law distributions. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):137–145, ???? 1997. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345920.

Jandhyala:1997:IPS

[2319] V. K. Jandhyala and I. B. MacNeill. Iterated partial sum sequences of regression residuals and tests for changepoints with continuity constraints. Journal of the Royal Statistical Society. Series B (Methodological), 59(1): 147–156, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345921.

Commenges:1997:GST

[2320] D. Commenges and H. Jacqmin-Gadda. Generalized score test of homogeneity based on correlated random effects models. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):157–171, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345922.

Eagleson:1997:TSR

[2321] G. K. Eagleson and H. G. Müller. Transformations for smooth regression models with multiplicative errors. Journal of the Royal Statistical Society. Series B (Methodological), 59(1): 173–189, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345923.

Cheng:1997:BAE

[2322] Ming-Yen Cheng. Boundary aware estimators of integrated density derivative products. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):191–203, ???? 1997. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2345924.

Altman:1997:OPB

[2323] Naomi Altman and Christian Léger.

On the optimality of prediction-based selection criteria and the convergence rates of estimators. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):205–216, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345925.

Sibisi:1997:PDM

[2324] Sibusiso Sibisi and John Skilling. Prior distributions on measure space. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):217–235, ???? 1997. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2345926.

Haywood:1997:FTS

[2325] J. Haywood and G. Tunnicliffe Wilson. Fitting time series models by minimizing multistep-ahead errors: A frequency domain approach. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):237–254, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345927.

Carter:1997:SBI

[2326] C. K. Carter and R. Kohn. Semiparametric Bayesian inference for time series with mixed spectra. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):255–268, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345928.

Shen:1997:WMS

[2327] Yu Shen and Thomas R. Fleming. Weighted mean survival test statistics: A class of distance tests for censored survival data. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):269–280, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345929.

Kent:1997:CPE

[2328] John T. Kent and Kanti V. Mardia. Consistency of Procrustes estimators. Journal of the Royal Statistical Society. Series B (Methodological), 59(1): 281–290, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2345930.

Anonymous:1997:BMa

[2329] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):??, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345931.

Anonymous:1997:FMa

[2330] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 59(1):??, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2345913.

Roberts:1997:USC

[2331] G. O. Roberts and S. K. Sahu.

Updating schemes, correlation structure, blocking and parameterization for the Gibbs sampler. Journal of the Royal Statistical Society. Series B (Methodological), 59(2):291–317, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346048.

Johnstone:1997:WTE

[2332] Iain M. Johnstone and Bernard W. Silverman. Wavelet threshold estimators for data with correlated noise. Journal of the Royal Statistical Society. Series B (Methodological), 59(2):319–351, ???? 1997. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346049.

Dryden:1997:PSA

[2333] Ian L. Dryden, Mohammad Reza Faghihi, and Charles C. Taylor. Procrustes shape analysis of planar point subsets. Journal of the Royal Statistical Society. Series B (Methodological), 59 (2):353–374, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346050.

Heitjan:1997:ISA

[2334] Daniel F. Heitjan. Ignorability, sufficiency and ancillarity. Journal of the Royal Statistical Society. Series B (Methodological), 59(2):375–381, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346051.

Lee:1997:EDF

[2335] Stephen M. S. Lee and G. Alastair Young. Estimation of the distribution function of a standardized statistic. Journal of the Royal Statistical Society. Series B (Methodological), 59(2): 383–400, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346052.

Molenberghs:1997:SFA

[2336] Geert Molenberghs and Els Goetghebeur. Simple fitting algorithms for incomplete categorical data. Journal of the Royal Statistical Society. Series B (Methodological), 59(2):401–414, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346053.

Becker:1997:EEI

[2337] Niels G. Becker and Abraham M. Hasofer. Estimation in epidemics with incomplete observations. Journal of the Royal Statistical Society. Series B (Methodological), 59(2):415–429, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346054.

Brockhoff:1997:RET

[2338] Per M. Brockhoff and Hans-Georg Müller. Random effect threshold models for dose–response relationships with repeated measurements. Journal of the Royal Statistical Society. Series B (Methodological), 59(2):431–446, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346055.

Breslow:1997:MLE

[2339] Norman E. Breslow and Richard Holubkov. Maximum likelihood estimation of logistic regression parameters under two- phase, outcome-dependent sampling. Journal of the Royal Statistical Society. Series B (Methodological), 59 (2):447–461, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346056.

Choy:1997:RAN

[2340] S. T. Boris Choy and Adrian F. M.

Smith. On robust analysis of a normal location parameter. Journal of the Royal Statistical Society. Series B (Methodological), 59(2):463–474, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346057.

Ledford:1997:MDW

[2341] Anthony W. Ledford and Jonathan A.

Tawn. Modelling dependence within joint tail regions. Journal of the Royal Statistical Society. Series B (Methodological), 59(2):475–499, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346058.

Le:1997:BMS

[2342] Nhu D. Le, Weimin Sun, and James V.

Zidek. Bayesian multivariate spatial interpolation with data missing by design. Journal of the Royal Statistical Society. Series B (Methodological), 59 (2):501–510, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2346059.

Anonymous:1997:BMb

[2343] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 59(2):??, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346060.

Anonymous:1997:FMb

[2344] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 59(2):??, ???? 1997. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346047.

Meng:1997:EAO

[2345] Xiao-Li Meng and David Van Dyk. The EM algorithm — an old folk-song sung to a fast new tune. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):511–567, ???? 1997. CODEN JSTBAJ. ISSN 0035-

                                        9246. URL http://www.jstor.org/
                                        stable/2346009.

Jamshidian:1997:AEA

[2346] Mortaza Jamshidian and Robert I. Jennrich. Acceleration of the EM algorithm by using quasi-Newton methods. Journal of the Royal Statistical Society. Series B (Methodological), 59(3): 569–587, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2346010.

Whittemore:1997:MSD

[2347] Alice S. Whittemore. Multistage sampling designs and estimating equations. Journal of the Royal Statistical Society. Series B (Methodological), 59(3): 589–602, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

                                        /www.jstor.org/stable/2346011.

Wang:1997:FFE

[2348] Yazhen Wang. Fractal function estimation via wavelet shrinkage. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):603–613, ???? 1997. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346012.

Grunwald:1997:SPG

[2349] Gary K. Grunwald, Kais Hamza, and Rob J. Hyndman. Some properties and generalizations of non-negative Bayesian time series models. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):615–626, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346013.

Hanfelt:1997:ALG

[2350] John J. Hanfelt and Kung-Yee Liang.

Approximate likelihoods for generalized linear errors-in-variables models. Journal of the Royal Statistical Society. Series B (Methodological), 59(3): 627–637, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346014.

Yeh:1997:BCR

[2351] Arthur B. Yeh and Kesar Singh. Balanced confidence regions based on Tukey’s depth and the bootstrap. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):639–652, ???? 1997. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346015.

Stern:1997:SOA

[2352] Steven E. Stern. A second-order adjustment to the profile likelihood in the case of a multidimensional parameter of interest. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):653–665, ???? 1997. CODEN JSTBAJ. ISSN 0035-9246. URL http:

//www.jstor.org/stable/2346016.

Sammel:1997:LVM

[2353] Mary Dupuis Sammel, Louise M. Ryan, and Julie M. Legler. Latent variable models for mixed discrete and continuous outcomes. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):667–678, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346017.

Kent:1997:EFD

[2354] John T. Kent and Andrew T. A. Wood. Estimating the fractal dimension of a locally self-similar Gaussian process by using increments. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):679–699, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346018.

Welham:1997:LRT

[2355] S. J. Welham and R. Thompson. Likelihood ratio tests for fixed model terms using residual maximum likelihood. Journal of the Royal Statistical Society. Series B (Methodological), 59(3): 701–714, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2346019.

Victoria-Feser:1997:RTN

[2356] Maria-Pia Victoria-Feser. A robust test for non-nested hypotheses. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):715–727, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346020.

Cheng:1997:APA

[2357] R. C. H. Cheng and W. B. Liu. Acknowledgement of priority: Acknowledgement of priority: A continuous representation of the family of stable law distributions. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):729, ???? 1997. CODEN JSTBAJ. ISSN 0035-9246.

Cheng:1997:APC

[2358] R. C. H. Cheng and W. B. Liu. Acknowledgement of priority: A continuous representation of the family of stable law distributions. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):729, ???? 1997. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2346021.

Anonymous:1997:BMc

[2359] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):??, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346022.

Anonymous:1997:FMc

[2360] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 59(3):??, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2346008.

Richardson:1997:BAM

[2361] Sylvia Richardson and Peter J. Green.

On Bayesian analysis of mixtures with an unknown number of components (with discussion). Journal of the Royal Statistical Society. Series B (Methodological), 59(4):731–792, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2985194. See corrigenda [2418, 2419].

Firth:1997:BOD

[2362] D. Firth and J. P. Hinde. On Bayesian D-optimum design criteria and the equivalence theorem in non-linear models. Journal of the Royal Statistical Society. Series B (Methodological), 59(4): 793–797, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985195.

Firth:1997:PNO

[2363] D. Firth and J. P. Hinde. Parameter neutral optimum design for non-linear models. Journal of the Royal Statistical Society. Series B (Methodological), 59 (4):799–811, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985196.

Berry:1997:OBR

[2364] Scott M. Berry and Joseph B. Kadane. Optimal Bayesian randomization. Journal of the Royal Statistical Society. Series B (Methodological), 59 (4):813–819, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985197.

Polansky:1997:KSI

[2365] Alan M. Polansky and William. R. Schucany. Kernel smoothing to improve bootstrap confidence intervals. Journal of the Royal Statistical Society. Series B (Methodological), 59(4): 821–838, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http:/

/www.jstor.org/stable/2985198.

Fung:1997:NLI

[2366] Wing K. Fung and C. W. Kwan. A note on local influence based on normal curvature. Journal of the Royal Statistical Society. Series B (Methodological), 59 (4):839–843, ???? 1997. CODEN JST-BAJ. ISSN 0035-9246. URL http://

www.jstor.org/stable/2985199.

Walker:1997:HGL

[2367] Stephen G. Walker and Bani K.

Mallick. Hierarchical generalized linear models and frailty models with Bayesian nonparametric mixing. Journal of the Royal Statistical Society. Series B (Methodological), 59(4):845–860, ???? 1997. CODEN JSTBAJ. ISSN

0035-9246. URL http://www.jstor.
org/stable/2985200.

Heffernan:1997:UEC

[2368] Peter M. Heffernan. Unbiased estimation of central moments by using U-statistics. Journal of the Royal Statistical Society. Series B (Methodological), 59(4):861–863, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/

stable/2985201. See acknowledgement of priority [2431].

Anonymous:1997:IAI

[2369] Anonymous. Index: Author index: Volume 59, 1997. Journal of the Royal Statistical Society. Series B (Methodological), 59(4):865–867, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985192.

Anonymous:1997:BMd

[2370] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Methodological), 59(4):??, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985202.

Anonymous:1997:FMd

[2371] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Methodological), 59(4):??, ???? 1997. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/
stable/2985193.

Jones:1998:ERE

[2372] M. C. Jones and G. A. Young. Editorial: Report of the Editors — 1997. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1):1–2, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985967.

Firth:1998:RMP

[2373] D. Firth and K. E. Bennett. Robust models in probability sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1):3–21, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985968.

Pfeffermann:1998:WUS

[2374] D. Pfeffermann, C. J. Skinner, D. J. Holmes, H. Goldstein, and J. Rasbash. Weighting for unequal selection probabilities in multilevel models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1): 23–40, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985969. See dis-

cussion [2375].

Anonymous:1998:DPFa

[2375] Anonymous. Discussion on the papers by Firth and Bennett and Pfeffermann et al. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1):41–56, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985970. See

[2374].

Forster:1998:MBI

[2376] Jonathan J. Forster and Peter W. F.

Smith. Model-based inference for categorical survey data subject to nonignorable non-response. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1):57– 70, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985971.

Clayton:1998:ALB

[2377] David Clayton, David Spiegelhalter, Graham Dunn, and Andrew Pickles. Analysis of longitudinal binary data from multi-phase sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1): 71–87, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985972. See dis-

cussion [2378].

Anonymous:1998:DPFb

[2378] Anonymous. Discussion on the papers by Forster and Smith and Clayton et al. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1):89–102, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985973. See

[2377].

Goutis:1998:SDF

[2379] Constantinos Goutis. Second-derivative functional regression with applications to near infra-red spectroscopy. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1): 103–114, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985974.

Magee:1998:ISW

[2380] Lonnie Magee. Improving surveyweighted least squares regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1): 115–126, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985975.

Lipsitz:1998:SMM

[2381] Stuart R. Lipsitz, Lue Ping Zhao, and Geert Molenberghs. A semiparametric method of multiple imputation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1): 127–144, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985976.

Chakraborty:1998:ATR

[2382] Biman Chakraborty and Probal Chaudhuri. On an adaptive transformation– retransformation estimate of multivariate location. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1):145–157, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985977.

Wang:1998:MES

[2383] Yuedong Wang. Mixed effects smoothing spline analysis of variance. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1): 159–174, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985978.

Muliere:1998:EFB

[2384] Pietro Muliere and Stephen Walker.

Extending the family of Bayesian bootstraps and exchangeable urn schemes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60 (1):175–182, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985979.

Yee:1998:ASV

[2385] Thomas W. Yee. On an alternative solution to the vector spline problem. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60 (1):183–188, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985980.

Cheng:1998:PRE

[2386] Chi-Lung Cheng and Hans Schneeweiss.

Polynomial regression with errors in the variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1):189–199, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985981.

Haslett:1998:RLM

[2387] John Haslett and Kevin Hayes. Residuals for the linear model with general covariance structure. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1): 201–215, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985982.

Nielsen:1998:OII

[2388] J. P. Nielsen and O. B. Linton. An optimization interpretation of integration and back-fitting estimators for separable nonparametric models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1): 217–222, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985983.

Munk:1998:NVS

[2389] Axel Munk and Claudia Czado. Nonparametric validation of similar distributions and assessment of goodness of fit. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1):223–241, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985984.

Li:1998:SLB

[2390] Gang Li and Jing Qin. Semiparametric likelihood-based inference for biased and truncated data when the total sample size is known. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1): 243–254, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985985.

Roberts:1998:OSD

[2391] Gareth O. Roberts and Jeffrey S.

Rosenthal. Optimal scaling of discrete approximations to Langevin diffusions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60 (1):255–268, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985986.

Aitkin:1998:SPB

[2392] Murray Aitkin. Simpson’s paradox and the Bayes factor. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(1):269– 270, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985987.

Anonymous:1998:BMa

[2393] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60 (1):??, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985988.

Anonymous:1998:FMa

[2394] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60 (1):??, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985966.

Hurvich:1998:SPS

[2395] Clifford M. Hurvich, Jeffrey S. Simonoff, and Chih-Ling Tsai. Smoothing parameter selection in nonparametric regression using an improved Akaike information criterion. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2): 271–293, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985940.

Luceno:1998:DPN

[2396] Alberto Luceño. Detecting possibly non-consecutive outliers in industrial time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2):295–310, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985941.

Smith:1998:ANR

[2397] Michael Smith, Chi-Ming Wong, and Robert Kohn. Additive nonparametric regression with autocorrelated errors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2):311–331, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985942.

Denison:1998:ABC

[2398] D. G. T. Denison, B. K. Mallick, and A. F. M. Smith. Automatic Bayesian curve fitting. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2):333– 350, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985943.

Ramsay:1998:CR

[2399] J. O. Ramsay and Xiaochun Li. Curve registration. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2):351–363, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985944.

Ramsay:1998:ESM

[2400] J. O. Ramsay. Estimating smooth monotone functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2):365–375, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985945.

Singh:1998:BVF

[2401] A. C. Singh, D. M. Stukel, and D. Pfeffermann. Bayesian versus frequentist measures of error in small area estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2):377–396, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985946.

Chambers:1998:LIL

[2402] Raymond L. Chambers, Alan H. Dorfman, and Suojin Wang. Limited information likelihood analysis of survey data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2):397–411, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985947.

Welsh:1998:BCE

[2403] A. H. Welsh and Elvezio Ronchetti. Bias-calibrated estimation from sample surveys containing outliers. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2): 413–428, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985948.

Hong:1998:TPS

[2404] Yongmiao Hong. Testing for pairwise serial independence via the empirical distribution function. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2): 429–453, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985949.

Shen:1998:TGE

[2405] Wei Shen and Thomas A. Louis. Triplegoal estimates in two-stage hierarchical models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2):455–471, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985950.

Nadarajah:1998:OME

[2406] S. Nadarajah, C. W. Anderson, and J. A. Tawn. Ordered multivariate extremes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(2):473–496, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985951.

Anonymous:1998:BMb

[2407] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60 (2):??, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985952.

Anonymous:1998:FMb

[2408] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60 (2):??, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985939.

Hodges:1998:SAG

[2409] James S. Hodges. Some algebra and geometry for hierarchical models, applied to diagnostics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(3): 497–536, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985928.

He:1998:BQS

[2410] Xuming He, Pin Ng, and Stephen Portnoy. Bivariate quantile smoothing splines. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(3):537–550, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985929.

Shi:1998:NUA

[2411] Peide Shi and Chih-Ling Tsai. A note on the unification of the Akaike information criterion. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(3):551–558, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985930.

Meng:1998:FET

[2412] Xiao-Li Meng and David van Dyk. Fast EM-type implementations for mixed effects models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(3):559–578, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985931.

Cheng:1998:CEM

[2413] Ming-Yen Cheng and Peter Hall. Calibrating the excess mass and dip tests of modality. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(3):579–589, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985932.

Fan:1998:LML

[2414] Jianqing Fan, Mark Farmen, and Irene Gijbels. Local maximum likelihood estimation and inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(3): 591–608, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985933.

Pievatolo:1998:BDT

[2415] A. Pievatolo and Peter J. Green.

Boundary detection through dynamic polygons. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(3):609–626, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985934.

Brown:1998:MBV

[2416] P. J. Brown, M. Vannucci, and T. Fearn. Multivariate Bayesian variable selection and prediction. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(3): 627–641, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985935.

Nicholls:1998:BIA

[2417] Geoff K. Nicholls. Bayesian image analysis with Markov chain Monte Carlo and coloured continuum triangulation models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(3):643–659, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985936.

Richardson:1998:CBA

[2418] Sylvia Richardson and Peter J. Green.

Corrigendum: On Bayesian analysis of mixtures with an unknown number of components. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(3):661, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/

stable/2985937. See [2361, 2419].

Richardson:1998:CCB

[2419] S. Richardson and P. J. Green. Correction: Corrigendum: On Bayesian analysis of mixtures with an unknown number of components. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(3):661, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2361, 2418].

Anonymous:1998:BMc

[2420] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60 (3):662, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2985938.

Anonymous:1998:FMc

[2421] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60 (3):??, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2985927.

Britton:1998:EME

[2422] Tom Britton. Estimation in multitype epidemics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(4):663–679, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985955.

Gasser:1998:NEM

[2423] Theo Gasser, Peter Hall, and Brett Presnell. Nonparametric estimation of the mode of a distribution of random curves. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(4):681–691, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985956.

Dette:1998:THN

[2424] Holger Dette and Axel Munk. Testing heteroscedasticity in nonparametric regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(4):693–708, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985957.

Eguchi:1998:CLL

[2425] Shinto Eguchi and John Copas. A class of local likelihood methods and near-parametric asymptotics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(4): 709–724, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985958.

Abramovich:1998:WTB

[2426] F. Abramovich, T. Sapatinas, and B. W. Silverman. Wavelet thresholding via a Bayesian approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(4): 725–749, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2985959.

Dette:1998:EVN

[2427] Holger Dette, Axel Munk, and Thorsten Wagner. Estimating the variance in nonparametric regression — what is a reasonable choice? Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(4): 751–764, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2985960.

Skouras:1998:EPP

[2428] Kostas Skouras and A. Philip Dawid. On efficient point prediction systems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60 (4):765–780, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2985961.

Stockis:1998:SIM

[2429] Jean-Pierre Stockis and Howell Tong. On the statistical inference of a machine-generated autoregressive AR(1) model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(4):781–796, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2985962.

Xia:1998:BCC

[2430] Yingcun Xia. Bias-corrected confidence bands in nonparametric regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(4):797–811, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985963.

Heffernan:1998:APU

[2431] Peter M. Heffernan. Acknowledgement of priority: Unbiased estimation of central moments by using U-statistics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(4):813, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2985964. See

[2368].

Anonymous:1998:IAV

[2432] Anonymous. Index of authors, volume 60, 1998. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(4):814–816, ???? 1998. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2985953.

Anonymous:1998:CV

[2433] Anonymous. Contents of volume 60, 1998. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60(4):817–818, ???? 1998. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:1998:BMd

[2434] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series B (Statistical Methodology), 60 (4):??, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985965.

Anonymous:1998:FMd

[2435] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 60 (4):??, ???? 1998. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2985954.

Firth:1999:RE

[2436] D. Firth and M. C. Jones. Report of the Editors: 1998. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (1):1–2, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680733.

Davies:1999:FAS

[2437] Steve Davies and Peter Hall. Fractal analysis of surface roughness by using spatial data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1):3–37, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680734.

Gijbels:1999:UES

[2438] I. Gijbels, A. Pope, and M. P. Wand. Understanding exponential smoothing via kernel regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1):39– 50, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680735.

Poon:1999:CNC

[2439] Wai-Yin Poon and Yat Sun Poon.

Conformal normal curvature and assessment of local influence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1): 51–61, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680736.

Antoniadis:1999:DHR

[2440] Anestis Antoniadis, Gerard Grégoire, and Guy Nason. Density and hazard rate estimation for right-censored data by using wavelet methods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1): 63–84, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680737.

Cheng:1999:MAM

[2441] Ching-Shui Cheng, David M. Steinberg, and Don X. Sun. Minimum aberration and model robustness for twolevel fractional factorial designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1): 85–93, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680738.

Hodgson:1999:BRI

[2442] M. E. A. Hodgson. A Bayesian restoration of an ion channel signal. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1): 95–114, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680739.

Severini:1999:EOE

[2443] Thomas A. Severini. On the effect of overdispersion on exact conditional tests. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1):115–126, ???? 1999. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680740.

Yao:1999:ETE

[2444] Qiwei Yao and Byron J. T. Morgan. Empirical transform estimation for indexed stochastic models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1): 127–141, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680741.

Hall:1999:IBB

[2445] Peter Hall and Brett Presnell. Intentionally biased bootstrap methods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (1):143–158, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680742.

Carpenter:1999:TIB

[2446] James Carpenter. Test inversion bootstrap confidence intervals. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1):

159–172, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680743.

Ibrahim:1999:MCG

[2447] Joseph G. Ibrahim, Stuart R. Lipsitz, and Ming-Hui Chen. Missing covariates in generalized linear models when the missing data mechanism is non-ignorable. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1):173– 190, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680744.

Poskitt:1999:DBD

[2448] D. S. Poskitt, K. Dogancay, and Shin-Ho Chung. Double-blind deconvolution: The analysis of post-synaptic currents in nerve cells. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1):191– 212, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680745.

Vidoni:1999:EFS

[2449] Paolo Vidoni. Exponential family state space models based on a conjugate latent process. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1):213– 221, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680746.

Chen:1999:PEV

[2450] Ming-Hui Chen, Joseph G. Ibrahim, and Constantin Yiannoutsos. Prior elicitation, variable selection and Bayesian computation for logistic regression models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1):223–242, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680747.

Sun:1999:NTC

[2451] Jianguo Sun. A nonparametric test for current status data with unequal censoring. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1):243–250, ???? 1999. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680748.

Moller:1999:PSC

[2452] Jesper Møller. Perfect simulation of conditionally specified models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1): 251–264, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680749.

Booth:1999:MGL

[2453] James G. Booth and James P. Hobert. Maximizing generalized linear mixed model likelihoods with an automated Monte Carlo EM algorithm. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(1): 265–285, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680750.

Anonymous:1999:BMa

[2454] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (1):??, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680751.

Anonymous:1999:FMa

[2455] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (1):??, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680732.

Becker:1999:SSI

[2456] Niels G. Becker and Tom Britton. Statistical studies of infectious disease incidence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(2):287–307, April 1999. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2680642.

Naasell:1999:TER

[2457] Ingemar Nåsell. On the time to extinction in recurrent epidemics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(2): 309–330, April 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680643.

Damien:1999:GSB

[2458] Paul Damien, Jon Wakefield, and Stephen Walker. Gibbs sampling for Bayesian non-conjugate and hierarchical models by using auxiliary variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (2):331–344, April 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680644.

Luceno:1999:DAC

[2459] Alberto Luceño. Discrete approximations to continuous univariate distributions — an alternative to simulation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (2):345–352, April 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680645.

Lee:1999:EMC

[2460] Stephen M. S. Lee and G. Alastair Young. The effect of Monte Carlo approximation on coverage error of double-bootstrap confidence intervals. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (2):353–366, April 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680646.

Henderson:1999:EFM

[2461] Robin Henderson and Paul Oman. Effect of frailty on marginal regression estimates in survival analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(2): 367–379, April 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680647.

Lin:1999:IGA

[2462] Xihong Lin and Daowen Zhang. Inference in generalized additive mixed models by using smoothing splines. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (2):381–400, April 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680648.

Lin:1999:PRM

[2463] D. Y. Lin and Paul S. F. Yip. Parametric regression models for continuous time removal and recapture studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (2):401–411, April 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680649.

Lawless:1999:SMR

[2464] J. F. Lawless, J. D. Kalbfleisch, and C. J. Wild. Semiparametric methods for response-selective and missing data problems in regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(2): 413–438, April 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680650.

Muller:1999:MBK

[2465] H. G. Müller and U. Stadtmüller. Multivariate boundary kernels and a continuous least squares principle. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(2): 439–458, April 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680651.

Lumley:1999:WEA

[2466] Thomas Lumley and Patrick Heagerty. Weighted empirical adaptive variance estimators for correlated data regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(2):459–477, April 1999. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680652.

Oakes:1999:DCI

[2467] David Oakes. Direct calculation of the information matrix via the EM algorithm. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(2):479–482, April 1999. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680653.

Gelman:1999:CCJ

[2468] A. Gelman and T. P. Speed. Corrigendum: Characterizing a joint probability distribution by conditionals. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(2):483, April 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680654. See

[2070].

Anonymous:1999:BMb

[2469] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (2):??, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680655.

Anonymous:1999:FMb

[2470] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (2):??, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680641.

Walker:1999:BNI

[2471] Stephen G. Walker, Paul Damien, Purushottam W. Laud, and Adrian F. M. Smith. Bayesian nonparametric inference for random distributions and related functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(3): 485–527, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680720.

Tibshirani:1999:CIC

[2472] Robert Tibshirani and Keith Knight.

The covariance inflation criterion for adaptive model selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(3): 529–546, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680721.

Iturria:1999:PRE

[2473] Stephen J. Iturria, Raymond J. Carroll, and David Firth. Polynomial regression and estimating functions in the presence of multiplicative measurement error. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(3):547–561, ???? 1999. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680722.

Ma:1999:SAD

[2474] Chunsheng Ma and John Robinson. Saddlepoint approximations for the difference of order statistics and studentized sample quantiles. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(3): 563–577, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680723.

Azzalini:1999:SAM

[2475] A. Azzalini and A. Capitanio. Statistical applications of the multivariate skew normal distribution. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(3): 579–602, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680724.

Haslett:1999:SDD

[2476] John Haslett. A simple derivation of deletion diagnostic results for the general linear model with correlated errors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (3):603–609, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680725.

Tipping:1999:PPC

[2477] Michael E. Tipping and Christopher M.

Bishop. Probabilistic principal component analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(3):611–622, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680726.

Fishman:1999:ASW

[2478] George S. Fishman. An analysis of Swendsen–Wang and related sampling methods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(3):623–641, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680727.

Roberts:1999:CSS

[2479] Gareth O. Roberts and Jeffrey S.

Rosenthal. Convergence of slice sampler Markov chains. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(3):643–660, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680728.

Hall:1999:BBM

[2480] Peter Hall and Brett Presnell. Biased bootstrap methods for reducing the effects of contamination. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(3): 661–680, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680729.

Ahmad:1999:GFT

[2481] Ibrahim A. Ahmad and Ibrahim A. Alwasel. A goodness-of-fit test for exponentiality based on the memoryless property. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(3):681–689, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680730.

Anonymous:1999:BMc

[2482] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (3):??, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680731.

Anonymous:1999:FMc

[2483] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (3):??, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680719.

Besag:1999:BAA

[2484] Julian Besag and David Higdon. Bayesian analysis of agricultural field experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4):691–746, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680704.

Dacre:1999:ARA

[2485] Marcus Dacre, Kevin Glazebrook, and Jose Niño-Mora. The achievable region approach to the optimal control of stochastic systems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4): 747–791, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680705.

Yang:1999:MBS

[2486] Lijian Yang and Rolf Tschernig. Multivariate bandwidth selection for local linear regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4):793– 815, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680706.

Fine:1999:ACR

[2487] J. P. Fine. Analysing competing risks data with transformation models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (4):817–830, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680707.

Walley:1999:UPB

[2488] Peter Walley and Serafin Moral. Upper probabilities based only on the likelihood function. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4):831–847, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680708.

Sweeting:1999:CBC

[2489] Trevor J. Sweeting. On the construction of Bayes–confidence regions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4): 849–861, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680709.

Lenk:1999:BIS

[2490] Peter J. Lenk. Bayesian inference for semiparametric regression using a Fourier representation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4):863– 879, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680710.

Huerta:1999:PCS

[2491] Gabriel Huerta and Mike West. Priors and component structures in autoregressive time series models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4): 881–899, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680711. See cor-

rigendum [2528].

Lee:1999:CMB

[2492] Stephen M. S. Lee. On a class of m out of n bootstrap confidence intervals. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (4):901–911, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680712.

Mardia:1999:CWD

[2493] K. V. Mardia and I. L. Dryden.

The complex Watson distribution and shape analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4):913–926, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680713.

Fan:1999:OSL

[2494] Jianqing Fan and Jianwei Chen. Onestep local quasi-likelihood estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (4):927–943, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680714.

Mukerjee:1999:CIA

[2495] Rahul Mukerjee and N. Reid. On confidence intervals associated with the usual and adjusted likelihoods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4): 945–953, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680715.

Moller:1999:EFA

[2496] Jesper Moller and Katja Schladitz. Extensions of Fill’s algorithm for perfect simulation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4):955–969, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680716.

Vannucci:1999:CSW

[2497] Marina Vannucci and Fabio Corradi. Covariance structure of wavelet coefficients: Theory and models in a Bayesian perspective. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4):971–986, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680717.

Anonymous:1999:IAV

[2498] Anonymous. Index of authors, volume 61, 1999. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4):987–989, ???? 1999. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680702.

Wu:1999:CSO

[2499] X. Wu and Z. Luo. Corrigendum: Second-order approach to local influence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61(4):990, ???? 1999. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2129].

Anonymous:1999:BMd

[2500] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 61 (4):??, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680718.

Anonymous:1999:FMd

[2501] Anonymous. Front matter. Journal of the Royal Statistical Society.

Series B (Statistical Methodology), 61 (4):??, ???? 1999. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680703.

Firth:2000:RE

[2502] D. Firth and M. C. Jones. Report of the Editors: 1999. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (1):1–2, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680675.

Durbin:2000:TSA

[2503] J. Durbin and S. J. Koopman. Time series analysis of non-Gaussian observations based on state space models from both classical and Bayesian perspectives. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(1):3–56, ???? 2000. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2680676.

Robert:2000:BIH

[2504] Christian P. Robert, Tobias Rydén, and D. M. Titterington. Bayesian inference in hidden Markov models through the reversible jump Markov chain Monte Carlo method. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(1): 57–75, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680677.

Shi:2000:LVM

[2505] Jian-Qing Shi and Sik-Yum Lee. Latent variable models with mixed continuous and polytomous data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(1): 77–87, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680678.

Hutson:2000:EBM

[2506] Alan D. Hutson and Michael D. Ernst. The exact bootstrap mean and variance of an L-estimator. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(1): 89–94, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680679.

Wong:2000:MAM

[2507] Chun Shan Wong and Wai Keung Li. On a mixture autoregressive model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (1):95–115, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680680.

Robinson:2000:EAP

[2508] Michael E. Robinson and Jonathan A. Tawn. Extremal analysis of processes sampled at different frequencies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(1): 117–135, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680681.

Hall:2000:WBA

[2509] Peter Hall and Yoshihiko Maesono. A weighted bootstrap approach to bootstrap iteration. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(1):137–144, ???? 2000. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680682.

Sebastiani:2000:MES

[2510] Paola Sebastiani and Henry P. Wynn.

Maximum entropy sampling and optimal Bayesian experimental design. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (1):145–157, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680683.

Wasserman:2000:AIM

[2511] Larry Wasserman. Asymptotic inference for mixture models using datadependent priors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(1):159–180, ???? 2000. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680684.

Fan:2000:CWD

[2512] Juanjuan Fan, Ross L. Prentice, and Li Hsu. A class of weighted dependence measures for bivariate failure time data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(1):181–190, ???? 2000. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2680685.

Davison:2000:LLS

[2513] A. C. Davison and N. I. Ramesh. Local likelihood smoothing of sample extremes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(1):191–208, ???? 2000. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680686.

Anonymous:2000:BMa

[2514] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (1):??, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680687.

Anonymous:2000:FMa

[2515] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (1):??, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680674.

McCullagh:2000:IFM

[2516] Peter McCullagh. Invariance and factorial models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2):209–256, ???? 2000. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3088857.

Jamshidian:2000:SEE

[2517] Mortaza Jamshidian and Robert I. Jennrich. Standard errors for EM estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2):257–270, ???? 2000. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3088858.

Nason:2000:WPA

[2518] Guy P. Nason, Rainer von Sachs, and Gerald Kroisandt. Wavelet processes and adaptive estimation of the evolutionary wavelet spectrum. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2): 271–292, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088859.

Sun:2000:RAP

[2519] Jianguo Sun and L. J. Wei. Regression analysis of panel count data with covariate-dependent observation and censoring times. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2):293– 302, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088860.

Fan:2000:TSE

[2520] Jianqing Fan and Jin-Ting Zhang.

Two-step estimation of functional linear models with applications to longitudinal data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2):303–322, ???? 2000. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3088861.

Paul:2000:GFG

[2521] Sudhir R. Paul and Dianliang Deng. Goodness of fit of generalized linear models to sparse data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2): 323–333, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088862.

Polzehl:2000:AWS

[2522] Jörg Polzehl and Vladimir G. Spokoiny. Adaptive weights smoothing with applications to image restoration. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2): 335–354, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088863.

Dunson:2000:BLV

[2523] David B. Dunson. Bayesian latent variable models for clustered mixed outcomes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2):355–366, ???? 2000. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3088864.

Girling:2000:RSE

[2524] A. J. Girling. Rank statistics expressible as integrals under P –P -plots and receiver operating characteristic curves. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2):367–382, ???? 2000. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3088865.

Xia:2000:EIT

[2525] Yingcun Xia, Howell Tong, W. K.

Li, and Li-Xing Zhu. On the estimation of an instantaneous transformation for time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2): 383–397, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088866.

Everson:2000:IMN

[2526] Philip J. Everson and Carl N. Morris. Inference for multivariate normal hierarchical models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2):399– 412, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088867.

Wood:2000:MSP

[2527] S. N. Wood. Modelling and smoothing parameter estimation with multiple quadratic penalties. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2): 413–428, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088868.

Huerta:2000:CPC

[2528] G. Huerta and M. West. Corrigendum: Priors and component structures in autoregressive time series models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(2):429, ???? 2000. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3088869. See

[2491].

Anonymous:2000:BMb

[2529] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (2):??, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088870.

Anonymous:2000:FMb

[2530] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (2):??, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088856.

Doksum:2000:VBS

[2531] Kjell Doksum, Derick Peterson, and Alex Samarov. On variable bandwidth selection in local polynomial regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (3):431–448, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680689.

Chen:2000:UAR

[2532] Yi-Hau Chen and Hung Chen. A unified approach to regression analysis under double-sampling designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(3): 449–460, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680690.

Choi:2000:BCR

[2533] Edwin Choi and Peter Hall. Bootstrap confidence regions computed from autoregressions of arbitrary order. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(3): 461–477, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680691.

Hall:2000:IIW

[2534] Peter Hall, Stephen M.-S. Lee, and G. Alastair Young. Importance of interpolation when constructing doublebootstrap confidence intervals. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(3): 479–491, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680692.

Chen:2000:MKF

[2535] Rong Chen and Jun S. Liu. Mixture Kalman filters. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(3):493–508, ???? 2000. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680693.

Wang:2000:EEE

[2536] C.-Y. Wang and Margaret Sullivan Pepe. Expected estimating equations to accommodate covariate measurement error. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(3):509–524, ???? 2000. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680694.

Smith:2000:BPM

[2537] Jim Q. Smith and Álvaro E. Faria, Jr. Bayesian Poisson models for the graphical combination of dependent expert information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(3):525– 544, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680695.

Gastwirth:2000:ASS

[2538] Joseph L. Gastwirth, Abba M. Krieger, and Paul R. Rosenbaum. Asymptotic separability in sensitivity analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (3):545–555, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680696.

Hong:2000:GST

[2539] Yongmiao Hong. Generalized spectral tests for serial dependence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(3): 557–574, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680697.

John:2000:RFA

[2540] J. A. John and D. Whitaker. Recursive formulae for the average efficiency factor in block and row- column designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (3):575–583, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680698.

Butler:2000:PSP

[2541] Neil A. Butler and Michael C. Denham. The peculiar shrinkage properties of partial least squares regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (3):585–593, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680699.

Lee:2000:MTE

[2542] Myoung jae Lee. Median treatment effect in randomized trials. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(3): 595–604, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680700.

Anonymous:2000:BMc

[2543] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (3):??, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680701.

Anonymous:2000:FMc

[2544] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (3):??, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680688.

Stephens:2000:IMP

[2545] Matthew Stephens and Peter Donnelly. Inference in molecular population genetics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4):605–655, ???? 2000. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680611.

Zhou:2000:FTR

[2546] Haibo Zhou and C.-Y. Wang. Failure time regression with continuous covariates measured with error. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4): 657–665, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680612.

Xu:2000:PHE

[2547] Ronghui Xu and John O’Quigley. Proportional hazards estimate of the conditional survival function. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4): 667–680, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680613.

Clyde:2000:FEB

[2548] Merlise Clyde and Edward I. George. Flexible empirical Bayes estimation for wavelets. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4):681–698, ???? 2000. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680614.

Cheng:2000:SSE

[2549] Chi-Lun Cheng, Hans Schneeweiss, and Markus Thamerus. A small sample estimator for a polynomial regression with errors in the variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4): 699–709, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680615.

Lin:2000:SRM

[2550] D. Y. Lin, L. J. Wei, I. Yang, and Z. Ying. Semiparametric regression for the mean and rate functions of recurrent events. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4): 711–730, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680616.

Whitehead:2000:CIS

[2551] John Whitehead, Susan Todd, and W. J. Hall. Confidence intervals for secondary parameters following a sequential test. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4):731–745, ???? 2000. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680617.

Lee:2000:MML

[2552] Mei-Ling Ting Lee, Victor DeGruttola, and David Schoenfeld. A model for markers and latent health status. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4):

747–762, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680618.

Naik:2000:PLS

[2553] Prasad Naik and Chih-Ling Tsai. Partial least squares estimator for singleindex models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4):763–771, ???? 2000. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680619.

Liang:2000:RAU

[2554] Kung-Yee Liang and Jing Qin. Regression analysis under non-standard situations: A pairwise pseudolikelihood approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4):773–786, ???? 2000. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680620.

Sugden:2000:CRS

[2555] R. A. Sugden, T. M. F. Smith, and R. P. Jones. Cochran’s rule for simple random sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4):787– 793, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680621.

Stephens:2000:DLS

[2556] Matthew Stephens. Dealing with label switching in mixture models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4):

795–809, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680622.

Hettmansperger:2000:ANI

[2557] T. P. Hettmansperger and Hoben Thomas. Almost nonparametric inference for repeated measures in mixture models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4):811–825, ???? 2000. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2680623.

Zucker:2000:ISS

[2558] David M. Zucker, Offer Lieberman, and Orly Manor. Improved small sample inference in the mixed linear model: Bartlett correction and adjusted likelihood. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4):827–838, ???? 2000. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2680624.

Neuhaus:2000:CCB

[2559] John M. Neuhaus. Closure of the class of binary generalized linear models in some non-standard settings. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4): 839–846, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680625.

Brown:2000:BGN

[2560] Patrick E. Brown, Kjetil F. Kåresen, Gareth O. Roberts, and Stefano Tonellato. Blur-generated non-separable space–time models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4):847–860, ???? 2000. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680626.

Anonymous:2000:IA

[2561] Anonymous. Index of authors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62(4): 861–863, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680609.

Anonymous:2000:BMd

[2562] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (4):??, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680627.

Anonymous:2000:FMd

[2563] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 62 (4):??, ???? 2000. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680610.

Davison:2001:ERE

[2564] A. C. Davison and D. Firth. Editorial: Report of the Editors — 2000. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(1):1–2, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680629.

Holmes:2001:BRM

[2565] C. C. Holmes and B. K. Mallick.

Bayesian regression with multivariate linear splines. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(1):3–17, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680630.

Chen:2001:MLR

[2566] Hanfeng Chen, Jiahua Chen, and John D. Kalbfleisch. A modified likelihood ratio test for homogeneity in finite mixture models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(1): 19–29, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680631.

Barry:2001:DSM

[2567] Simon C. Barry and A. H. Welsh. Distance sampling methodology. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(1): 31–53, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680632.

Crowder:2001:RMA

[2568] Martin Crowder. On repeated measures analysis with misspecified covariance structure. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(1):55–62, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680633.

Hsieh:2001:HHR

[2569] Fushing Hsieh. On heteroscedastic hazards regression models: Theory and application. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(1):63–79, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680634.

Gustafson:2001:MSP

[2570] Paul Gustafson. On measuring sensitivity to parametric model misspecification. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(1):81–94, ???? 2001. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680635.

Garthwaite:2001:NCP

[2571] Paul H. Garthwaite and Shafeeqah A. Al-Awadhi. Non-conjugate prior distribution assessment for multivariate normal sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(1):95–110, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680636.

Zhu:2001:LII

[2572] Hong-Tu Zhu and Sik-Yum Lee. Local influence for incomplete-data models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (1):111–126, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680637.

Gilks:2001:FMT

[2573] Walter R. Gilks and Carlo Berzuini.

Following a moving target — Monte Carlo inference for dynamic Bayesian models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(1):127–146, ???? 2001. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2680638.

Kauermann:2001:TGL

[2574] Göran Kauermann and Gerhard Tutz.

Testing generalized linear and semiparametric models against smooth alternatives. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(1):147–166, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680639.

Anonymous:2001:BMa

[2575] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (1):??, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680640.

Anonymous:2001:FMa

[2576] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (1):??, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680628.

Barndorff-Nielsen:2001:NGO

[2577] Ole E. Barndorff-Nielsen and Neil Shephard. Non-Gaussian Ornstein– Uhlenbeck-based models and some of their uses in financial economics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2): 167–241, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680596.

Ronn:2001:NML

[2578] Birgitte B. Rønn. Nonparametric maximum likelihood estimation for shifted curves. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2):243–259, ???? 2001. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680597.

Politis:2001:MES

[2579] Dimitris N. Politis and Michael Sherman. Moment estimation for statistics from marked point processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2): 261–275, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680598.

Kuk:2001:REF

[2580] Anthony Y. C. Kuk and A. H. Welsh. Robust estimation for finite populations based on a working model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2):

277–292, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680599.

Guillou:2001:DST

[2581] Armelle Guillou and Peter Hall. A diagnostic for selecting the threshold in extreme value analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2): 293–305, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680600.

Critchley:2001:IAB

[2582] Frank Critchley, Richard A. Atkinson, Guobing Lu, and Elenice Biazi. Influence analysis based on the case sensitivity function. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2):307– 323, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680601.

Rue:2001:FSG

[2583] Håvard Rue. Fast sampling of Gaussian Markov random fields. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2): 325–338, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680602.

Gu:2001:MLE

[2584] Ming Gao Gu and Hong-Tu Zhu. Maximum likelihood estimation for spatial models by Markov chain Monte Carlo stochastic approximation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2): 339–355, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680603.

Cai:2001:SDV

[2585] Zongwu Cai, Qiwei Yao, and Wenyang Zhang. Smoothing for discrete-valued time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2):357–375, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680604.

Klinger:2001:IHD

[2586] Artur Klinger. Inference in high dimensional generalized linear models based on soft thresholding. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2): 377–392, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680605.

Bura:2001:ESD

[2587] Efstathia Bura and R. Dennis Cook. Estimating the structural dimension of regressions via parametric inverse regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2):393–410, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680606.

Tibshirani:2001:ENC

[2588] Robert Tibshirani, Guenther Walther, and Trevor Hastie. Estimating the number of clusters in a data set via the gap statistic. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(2):411– 423, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680607.

Anonymous:2001:BMb

[2589] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (2):??, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680608.

Anonymous:2001:FMb

[2590] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (2):??, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680595.

Kennedy:2001:BCC

[2591] Marc C. Kennedy and Anthony O’Hagan. Bayesian calibration of computer models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(3):425–464, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680584.

Glasbey:2001:PLA

[2592] C. A. Glasbey and K. V. Mardia. A penalized likelihood approach to image warping. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(3):465–514, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680585.

Koreisha:2001:GLS

[2593] Sergio G. Koreisha and Yue Fang. Generalized least squares with misspecified serial correlation structures. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(3): 515–531, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680586.

James:2001:FLD

[2594] Gareth M. James and Trevor J. Hastie. Functional linear discriminant analysis for irregularly sampled curves. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(3): 533–550, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680587.

Nason:2001:RPI

[2595] Guy P. Nason. Robust projection indices. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(3):551–567, ???? 2001. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/2680588.

Hall:2001:LLT

[2596] Peter Hall, Liang Peng, and Christian Rau. Local likelihood tracking of fault lines and boundaries. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(3): 569–582, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680589.

Jiang:2001:PIN

[2597] Wenxin Jiang, Victor Kipnis, Douglas Midthune, and Raymond J. Carroll. Parameterization and inference for nonparametric regression problems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (3):583–591, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680590.

Mira:2001:PSS

[2598] A. Mira, J. Møller, and G. O. Roberts.

Perfect slice samplers. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(3):593– 606, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680591. See cor-

rigendum [2658].

Chen:2001:PSM

[2599] Kani Chen. Parametric and semiparametric models for recapture and removal studies: A likelihood approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (3):607–619, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680592.

Butler:2001:GMC

[2600] Neil A. Butler, Roger Mead, Kent M. Eskridge, and Steven G. Gilmour. A general method of constructing E(s^{2})- optimal supersaturated designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(3): 621–632, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680593.

Anonymous:2001:BMc

[2601] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (3):??, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680594.

Anonymous:2001:FMc

[2602] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (3):??, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680583.

Lewis:2001:DIE

[2603] S. M. Lewis and A. M. Dean. Detection of interactions in experiments on large numbers of factors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4): 633–672, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680658.

Stroud:2001:DMS

[2604] Jonathan R. Stroud, Peter Müller, and Bruno Sansó. Dynamic models for spatiotemporal data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4): 673–689, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680659.

Datta:2001:EVM

[2605] Gauri S. Datta and Thomas J. DiCiccio. On expected volumes of multidimensional confidence sets associated with the usual and adjusted likelihoods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (4):691–703, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680660.

Britton:2001:EHC

[2606] Tom Britton. Epidemics in heterogeneous communities: Estimation of R_{O} and secure vaccination coverage. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4): 705–715, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680661.

Hall:2001:ICA

[2607] Peter Hall and Andrew Rieck. Improving coverage accuracy of nonparametric prediction intervals. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4): 717–725, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680662.

Audrino:2001:TSG

[2608] Francesco Audrino and Peter Bühlmann. Tree-structured generalized autoregressive conditional heteroscedastic models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4):727–744, ???? 2001. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680663.

DiRienzo:2001:EMM

[2609] A. G. DiRienzo and S. W. Lagakos. Effects of model misspecification on tests of no randomized treatment effect arising from Cox’s proportional hazards model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4):745–757, ???? 2001. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680664.

Joffe:2001:UIR

[2610] Marshall M. Joffe. Using information on realized effects to determine prospective causal effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4): 759–774, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680665.

Chen:2001:PMR

[2611] Kani Chen. Parametric models for response-biased sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4):775– 789, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680666.

Chen:2001:GCC

[2612] Kani Chen. Generalized case–cohort sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4):791–809, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/2680667.

Walker:2001:BC

[2613] Stephen Walker and Nils Lid Hjort.

On Bayesian consistency. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4): 811–821, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680668.

Brix:2001:SPL

[2614] Anders Brix and Peter J. Diggle. Spatiotemporal prediction for log-Gaussian Cox processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4):823–841, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/

stable/2680669. See corrigenda [3239, 3239].

Lawrance:2001:SAC

[2615] A. J. Lawrance and N. Balakrishna.

Statistical aspects of chaotic maps with negative dependence in a communications setting. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4):843–853, ????

2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680670.

Xie:2001:RSC

[2616] Sharon X. Xie, C. Y. Wang, and Ross L. Prentice. A risk set calibration method for failure time regression by using a covariate reliability sample. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (4):855–870, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/2680671.

Copas:2001:LSA

[2617] John Copas and Shinto Eguchi. Local sensitivity approximations for selectivity bias. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4):871–895, ???? 2001. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/2680672.

Anonymous:2001:IAV

[2618] Anonymous. Index of authors, volume 63, 2001. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63(4):897–899, ???? 2001. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2680656.

Anonymous:2001:BMd

[2619] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (4):??, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680673.

Anonymous:2001:FMd

[2620] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 63 (4):??, ???? 2001. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/2680657.

Davison:2002:RE

[2621] A. C. Davison and D. Firth. Report of the Editors — 2001. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(1): 1–2, January 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088845.

Geng:2002:CCE

[2622] Zhi Geng, Jianhua Guo, and Wing-Kam Fung. Criteria for confounders in epidemiological studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(1): 3–15, January 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088846.

Huang:2002:CRM

[2623] Yijian Huang. Censored regression with the multistate accelerated sojourn times model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(1):17–29, January 2002. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3088847.

Butler:2002:BST

[2624] Ronald W. Butler and Douglas A. Bronson. Bootstrapping survival times in stochastic systems by using saddlepoint approximations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(1):31– 49, January 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088848.

Chen:2002:CRC

[2625] Yi-Hau Chen. Cox regression in cohort studies with validation sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(1): 51–62, January 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088849.

Holland:2002:FRC

[2626] Burt Holland and Siu Hung Cheung. Familywise robustness criteria for multiple-comparison procedures. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(1): 63–77, January 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088850.

Petrone:2002:CBP

[2627] Sonia Petrone and Larry Wasserman. Consistency of Bernstein polynomial posteriors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(1):79–100, January 2002. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3088851.

Yau:2002:REG

[2628] Kelvin K. W. Yau and Anthony Y. C.

Kuk. Robust estimation in generalized linear mixed models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(1):101– 117, January 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088852.

Wood:2002:MSS

[2629] Sally Wood, Robert Kohn, Tom Shively, and Wenxin Jiang. Model selection in spline nonparametric regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(1):119–139, January 2002. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3088853.

Hall:2002:HOD

[2630] Peter Hall and Michael C. Minnotte.

High order data sharpening for density estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(1):141–157, January 2002. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3088854.

Anonymous:2002:BMa

[2631] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64 (1):??, ???? 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088855.

Anonymous:2002:FMa

[2632] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64 (1):??, ???? 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088844.

Yin:2002:DRC

[2633] Xiangrong Yin and R. Dennis Cook. Dimension reduction for the conditional k th moment in regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(2): 159–175, May 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088793.

MacEachern:2002:NRS

[2634] Steven N. MacEachern, Ömer Öztürk, Douglas A. Wolfe, and Gregory V. Stark. A new ranked set sample estimator of variance. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(2): 177–188, May 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088794.

Barber:2002:PPI

[2635] Stuart Barber, Guy P. Nason, and Bernard W. Silverman. Posterior probability intervals for wavelet thresholding. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(2):189–205, May 2002. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3088795.

Scott:2002:RWM

[2636] Alastair Scott and Chris Wild. On the robustness of weighted methods for fitting models to case–control data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(2): 207–219, May 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088796.

Shi:2002:PBM

[2637] Jian Qing Shi and John Copas. Publication bias and meta-analysis for 2 × 2 tables: an average Markov chain Monte Carlo EM algorithm. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(2): 221–236, May 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088797.

Shi:2002:RMS

[2638] Peide Shi and Chih-Ling Tsai. Regression model selection: A residual likelihood approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(2): 237–252, May 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088798.

Barndorff-Nielsen:2002:EAR

[2639] Ole E. Barndorff-Nielsen and Neil Shephard. Econometric analysis of realized volatility and its use in estimating stochastic volatility models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(2): 253–280, May 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088799.

Zhang:2002:PGF

[2640] Jin Zhang. Powerful goodness-of-fit tests based on the likelihood ratio. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64 (2):281–294, May 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088800.

Holmes:2002:PNN

[2641] C. C. Holmes and N. M. Adams. A probabilistic nearest neighbour method for statistical pattern recognition. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64 (2):295–306, May 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088801.

Ramsay:2002:SSD

[2642] Tim Ramsay. Spline smoothing over difficult regions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(2):307–319, May 2002. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3088802.

Anonymous:2002:BMb

[2643] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64 (2):??, ???? 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088803.

Anonymous:2002:FMb

[2644] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64 (2):??, ???? 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088792.

Lauritzen:2002:CGM

[2645] Steffen L. Lauritzen and Thomas S.

Richardson. Chain graph models and their causal interpretations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3): 321–348, August 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088778. See dis-

cussion [2646].

Anonymous:2002:DPL

[2646] Anonymous. Discussion on the paper by Lauritzen and Richardson. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3): 348–361, August 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2645].

Xia:2002:AED

[2647] Yingcun Xia, Howell Tong, W. K. Li, and Li-Xing Zhu. An adaptive estimation of dimension reduction space. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3):363–410, August 2002. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3088779.

James:2002:GLM

[2648] Gareth M. James. Generalized linear models with functional predictors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3):411–432, August 2002. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3088780.

Gupta:2002:CCP

[2649] Jayanti Gupta and Paul Damien. Conjugacy class prior distributions on metric-based ranking models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3): 433–445, August 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088781.

Ollila:2002:ERC

[2650] Esa Ollila, Hannu Oja, and Thomas P. Hettmansperger. Estimates of regression coefficients based on the sign covariance matrix. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3):447–466, August 2002. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3088782.

Sinha:2002:MSD

[2651] Debajyoti Sinha, Joseph G. Ibrahim, and Ming-Hui Chen. Models for survival data from cancer prevention studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3):467–477, August 2002. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3088783.

Storey:2002:DAF

[2652] John D. Storey. A direct approach to false discovery rates. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3):479– 498, August 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088784.

Genovese:2002:OCE

[2653] Christopher Genovese and Larry Wasserman. Operating characteristics and extensions of the false discovery rate procedure. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3):499–517, August 2002. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3088785.

Brown:2002:BMA

[2654] P. J. Brown, M. Vannucci, and T. Fearn. Bayes model averaging with selection of regressors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3): 519–536, August 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088786.

Hall:2002:REK

[2655] Peter Hall and Terence Tao. Relative efficiencies of kernel and local likelihood density estimators. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3): 537–547, August 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088787.

Hall:2002:AVL

[2656] Peter Hall, K. Humphreys, and D. M. Titterington. On the adequacy of variational lower bound functions for likelihood-based inference in Markovian models with missing values. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3): 549–564, August 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088788.

Fung:2002:IDO

[2657] Wing-Kam Fung, Zhong-Yi Zhu, Bo-Cheng Wei, and Xuming He. Influence diagnostics and outlier tests for semiparametric mixed models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3): 565–579, August 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088789.

Mira:2002:CPS

[2658] A. Mira, J. Møller, and G. O. Roberts. Corrigendum: Perfect slice samplers. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(3):581, August 2002. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3088790. See

[2598].

Anonymous:2002:BMc

[2659] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64 (3):??, ???? 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088791.

Anonymous:2002:FMc

[2660] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64 (3):??, ???? 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088777.

Spiegelhalter:2002:BMM

[2661] David J. Spiegelhalter, Nicola G. Best, Bradley P. Carlin, and Angelika Van Der Linde. Bayesian measures of model complexity and fit. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):583– 639, October 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088806.

Broman:2002:MSA

[2662] Karl W. Broman and Terence P. Speed.

A model selection approach for the identification of quantitative trait loci in experimental crosses. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):641– 656, October 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088807.

Fearnhead:2002:ALM

[2663] Paul Fearnhead and Peter Donnelly.

Approximate likelihood methods for estimating local recombination rates. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):657–680, October 2002. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3088808.

Larget:2002:BPI

[2664] Bret Larget, Donald L. Simon, and Joseph B. Kadane. Bayesian phylogenetic inference from animal mitochondrial genome arrangements. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):681– 693, October 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088809.

Nicholson:2002:APD

[2665] George Nicholson, Albert V. Smith, Frosti Jónsson, Ómar Gústafsson, Kári Stefánsson, and Peter Donnelly. Assessing population differentiation and isolation from single-nucleotide polymorphism data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):695–715, October 2002. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3088810.

Parmigiani:2002:SFE

[2666] Giovanni Parmigiani, Elizabeth S. Garrett, Ramaswamy Anbazhagan, and Edward Gabrielson. A statistical framework for expression-based molecular classification in cancer. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):717– 736, October 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088811. See dis-

cussion [2667].

Balding:2002:DMS

[2667] David J. Balding, Andrew D. Carothers, Jonathan L. Marchini, Lon R. Cardon, Atam Vetta, Bob Griffiths, B. S. Weir, W. G. Hill, Darlene Goldstein, Korbinian Strimmer, Simon Myers, Mark A. Beaumont, C. A. Glasbey, C. D. Mayer, Richard Durrett, Rasmus Nielsen, P. M. Visscher, S. A. Knott, C. S. Haley, Roderick D. Ball, Christine A. Hackett, Susan Holmes, Dirk Husmeier, Ritsert C. Jansen, Cajo J. F. ter Braak, Chris A. Maliepaard, Martin P. Boer, Paul Joyce, Na Li, Matthew Stephens, George A. Marcoulides, Zvi Drezner, Kanti Mardia, Gilean McVean, Xiao-Li Meng, Michael F. Ochs, Mark Pagel, Naijun Sha, Marina Vannucci, Mikko J. Sillanpää, Scott Sisson, Brian S. Yandell, Chunfang Jin, Jaya M. Satagopan, Patrick J. Gaffney, Zhao-Bang Zeng, Karl W. Broman, Terence P. Speed, Paul Fearnhead, Peter Donnelly, Bret Larget, Donald L. Simon, Joseph B. Kadane, George Nicholson, Albert V. Smith, Frosti Jónsson, Ómar Gústafsson, Kárl Stefánsson, Peter Donnelly, Giovanni Parmigiani, Elizabeth S. Garrett, Ramaswamy Anbazhagan, and Edward Gabrielson. Discussion on the meeting on ‘Statistical modelling and analysis of genetic data’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):737–775, October 2002. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3088812. See [2666].

Casella:2002:PSM

[2668] G. Casella, K. L. Mengersen, C. P. Robert, and D. M. Titterington. Perfect samplers for mixtures of distributions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):777–790, October 2002. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3088813.

Hengartner:2002:BSL

[2669] Nicolas W. Hengartner, Marten H. Wegkamp, and Eric Matzner-Løber. Bandwidth selection for local linear regression smoothers. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):791– 804, October 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088814.

Fernandez:2002:MSC

[2670] Carmen Fernández and Peter J. Green. Modelling spatially correlated data via mixtures: A Bayesian approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4): 805–826, October 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088815.

Andrieu:2002:PFP

[2671] Christophe Andrieu and Arnaud Doucet. Particle filtering for partially observed Gaussian state space models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):827–836, October 2002. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3088816.

Lee:2002:LSV

[2672] Yoon Dong Lee and Soumendra N.

Lahiri. Least squares variogram fitting by spatial subsampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4): 837–854, October 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088817.

Skinner:2002:MDR

[2673] C. J. Skinner and M. J. Elliot. A measure of disclosure risk for microdata. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):855–867, October 2002. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3088818.

Delaigle:2002:EIS

[2674] A. Delaigle and I. Gijbels. Estimation of integrated squared density derivatives from a contaminated sample. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):869–886, October 2002. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3088819.

Guo:2002:ISS

[2675] Wensheng Guo. Inference in smoothing spline analysis of variance. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4): 887–898, October 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088820.

Efron:2002:TWP

[2676] Bradley Efron. The two-way proportional hazards model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):899– 909, October 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088821.

Anonymous:2002:IAV

[2677] Anonymous. Index of authors, volume 64, 2002. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64(4):911–916, October 2002. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3088804.

Anonymous:2002:BMd

[2678] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64 (4):??, ???? 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088822.

Anonymous:2002:FMd

[2679] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 64 (4):??, ???? 2002. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088805.

Davison:2003:RE

[2680] A. C. Davison and R. Henderson. Report of the Editors, 2002. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1): 1–2, February 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088824.

Brooks:2003:ECR

[2681] S. P. Brooks, P. Giudici, and G. O.

Roberts. Efficient construction of reversible jump Markov chain Monte Carlo proposal distributions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1): 3–39, February 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088825. See dis-

cussion [2682].

Robert:2003:DPB

[2682] Christian P. Robert, Xiao-Li Meng, Jesper Møller, Jeffrey S. Rosenthal, C. Jennison, M. A. Hurn, F. Al-Awadhi, Peter McCullagh, Christophe Andrieu, Arnaud Doucet, Petros Dellaportas, Ioulia Papageorgiou, Ricardo S. Ehlers, Elena A. Erosheva, Stephen E. Fienberg, Jonathan J. Forster, Roger C. Gill, Nial Friel, Peter Green, David Hastie, R. King, Hans R. Künsch, N. A. Lazar, and C. Osinski. Discussion on the paper by Brooks, Giudici and Roberts. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1): 39–55, February 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2681].

Fan:2003:AVC

[2683] Jianqing Fan, Qiwei Yao, and Zongwu Cai. Adaptive varying-coefficient linear models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):57–80, February 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3088826.

Genton:2003:CDB

[2684] Marc G. Genton and André Lucas. Comprehensive definitions of breakdown points for independent and dependent observations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):81– 94, February 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088827.

Wood:2003:TPR

[2685] Simon N. Wood. Thin plate regression splines. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):95–114, February 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3088828.

Fortiana:2003:GFT

[2686] J. Fortiana and A. Grané. Goodnessof-fit tests based on maximum correlations and their orthogonal decompositions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):115–126, February 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3088829.

Qu:2003:BAE

[2687] Annie Qu and Bruce G. Lindsay. Building adaptive estimating equations when inverse of covariance estimation is difficult. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):127–142, February 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3088830.

Tatsuoka:2003:SCP

[2688] Curtis Tatsuoka and Thomas Ferguson. Sequential classification on partially ordered sets. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):143–157, February 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3088831.

Jones:2003:SED

[2689] M. C. Jones and M. J. Faddy. A skew extension of the t-distribution, with applications. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):159–174, February 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3088832.

Korn:2003:EVC

[2690] Edward L. Korn and Barry I.

Graubard. Estimating variance components by using survey data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1): 175–190, February 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088833.

Li:2003:EPS

[2691] Haihong Li, Bruce G. Lindsay, and Richard P. Waterman. Efficiency of projected score methods in rectangular array asymptotics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):191– 208, February 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088834.

Kuroki:2003:CSE

[2692] Manabu Kuroki and Masami Miyakawa. Covariate selection for estimating the causal effect of control plans by using causal diagrams. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):209–222, February 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3088835.

Miwa:2003:EGN

[2693] Tetsuhisa Miwa, A. J. Hayter, and Satoshi Kuriki. The evaluation of general non-centred orthant probabilities. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):223–234, February 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3088836.

Pettitt:2003:ECN

[2694] A. N. Pettitt, N. Friel, and R. Reeves. Efficient calculation of the normalizing constant of the autologistic and related models on the cylinder and lattice.

Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):235–246, February 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3088837.

Liu:2003:PBB

[2695] Jiannong Liu and James S. Hodges.

Posterior bimodality in the balanced one-way random-effects model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1): 247–255, February 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088838.

Wang:2003:EAP

[2696] Weijing Wang. Estimating the association parameter for copula models under dependent censoring. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):257– 273, February 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088839.

Birmingham:2003:PMS

[2697] Jolene Birmingham, Andrea Rotnitzky, and Garrett M. Fitzmaurice. Pattern– mixture and selection models for analysing longitudinal data with monotone missing patterns. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):275– 297, February 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3088840.

Sundberg:2003:CSI

[2698] Rolf Sundberg. Conditional statistical inference and quantification of relevance. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1):299–315, February 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3088841.

Fine:2003:SES

[2699] Jason P. Fine, David V. Glidden, and Kristine E. Lee. A simple estimator for a shared frailty regression model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(1): 317–329, February 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088842.

Anonymous:2003:BMa

[2700] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65 (1):??, ???? 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088843.

Anonymous:2003:FMa

[2701] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65 (1):??, ???? 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3088823.

Murphy:2003:ODT

[2702] S. A. Murphy. Optimal dynamic treatment regimes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2):331–355, May 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.
org/stable/3647509. See discussion

[2703].

Anonymous:2003:DPM

[2703] Anonymous. Discussion on the paper by Murphy. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2):355–366, May 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2702].

Azzalini:2003:DGP

[2704] Adelchi Azzalini and Antonella Capitanio. Distributions generated by perturbation of symmetry with emphasis on a multivariate skew t-distribution. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65 (2):367–389, May 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647510.

Royall:2003:ISE

[2705] Richard Royall and Tsung-Shan Tsou.

Interpreting statistical evidence by using imperfect models: Robust adjusted likelihood functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2): 391–404, May 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647511.

Chiou:2003:FQL

[2706] Jeng-Min Chiou, Hans-Georg Müller, and Jane-Ling Wang. Functional quasi-likelihood regression models with smooth random effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2):405–423, May 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647512.

Hall:2003:DTT

[2707] Peter Hall and Qiwei Yao. Data tilting for time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2):425–442, May 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647513.

Hall:2003:UDB

[2708] Peter Hall and Ingrid Van Keilegom. Using difference-based methods for inference in nonparametric regression with time series errors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2): 443–456, May 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647514.

Akritas:2003:EBM

[2709] Michael G. Akritas and Ingrid Van Keilegom. Estimation of bivariate and marginal distributions with censored data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2):457–471, May 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647515.

Winnett:2003:IRT

[2710] Angela Winnett and Peter Sasieni. Iterated residuals and time-varying covariate effects in Cox regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2): 473–488, May 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647516.

Kim:2003:MLE

[2711] Jong S. Kim. Maximum likelihood estimation for the proportional hazards model with partly interval-censored data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2):489–502, May 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647517.

Brooks:2003:CMS

[2712] S. P. Brooks, N. Friel, and R. King.

Classical model selection via simulated annealing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2):503–520, May 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647518.

Ledford:2003:DDW

[2713] Anthony W. Ledford and Jonathan A.

Tawn. Diagnostics for dependence within time series extremes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2):

521–543, May 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647519.

Ferro:2003:ICE

[2714] Christopher A. T. Ferro and Johan Segers. Inference for clusters of extreme values. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2):545–556, May 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647520.

Quintana:2003:BCP

[2715] Fernando A. Quintana and Pilar L. Iglesias. Bayesian clustering and product partition models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2): 557–574, May 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647521.

Chen:2003:NPA

[2716] Hua Yun Chen. A note on the prospective analysis of outcome-dependent samples. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(2):575–584, May 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647522.

Anonymous:2003:BMb

[2717] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65 (2):??, ???? 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647523.

Anonymous:2003:FMb

[2718] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65 (2):??, ???? 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647508.

Kong:2003:TSM

[2719] A. Kong, P. McCullagh, X.-L. Meng, D. Nicolae, and Z. Tan. A theory of statistical models for Monte Carlo integration. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(3):585–604, August 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647541. See

discussion [2720].

Anonymous:2003:DPK

[2720] Anonymous. Discussion on the paper by Kong, McCullagh, Meng, Nicolae and Tan. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(3):604–618, August 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2719].

Andersen:2003:BIG

[2721] Kim E. Andersen, Stephen P. Brooks, and Martin B. Hansen. Bayesian inversion of geoelectrical resistivity data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(3):619–642, August 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647542.

Gorfine:2003:EDB

[2722] Malka Gorfine, Li Hsu, and Ross L. Prentice. Estimation of dependence between paired correlated failure times in the presence of covariate measurement error. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(3):643–661, August 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647543.

Chen:2003:ELG

[2723] Song Xi Chen, Wolfgang Härdle, and Ming Li. An empirical likelihood goodness-of-fit test for time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(3):663–678, August 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647544.

Cappe:2003:RJB

[2724] Olivier Cappé, Christian P. Robert, and Tobias Rydén. Reversible jump, birth-and-death and more general continuous time Markov chain Monte Carlo samplers. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(3):679–700, August 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647545.

Abbring:2003:IMP

[2725] Jaap H. Abbring and Gerard J. Van Den Berg. The identifiability of the mixed proportional hazards competing risks model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(3):701–710, August 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647546.

Rathouz:2003:LMM

[2726] Paul J. Rathouz. Likelihood methods for missing covariate data in highly stratified studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(3):711–723, August 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647547.

Dette:2003:EDE

[2727] Holger Dette, Viatcheslav B. Melas, Andrey Pepelyshev, and Nikolai Strigul. Efficient design of experiments in the Monod model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(3): 725–742, August 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647548.

Schmidt:2003:BIN

[2728] Alexandra M. Schmidt and Anthony O’Hagan. Bayesian inference for nonstationary spatial covariance structure via spatial deformations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(3):743– 758, August 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647549.

Tsodikov:2003:SMG

[2729] A. Tsodikov. Semiparametric models: A generalized self-consistency approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(3):759–774, August 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647550.

Anonymous:2003:BMc

[2730] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65 (3):??, ???? 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647551.

Anonymous:2003:FMc

[2731] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65 (3):??, ???? 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647540.

Barndorff-Nielsen:2003:QSI

[2732] Ole E. Barndorff-Nielsen, Richard D. Gill, and Peter E. Jupp. On quantum statistical inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):775– 804, November 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647584. See dis-

cussion [2733].

Anonymous:2003:DPB

[2733] Anonymous. Discussion on the paper by Barndorff-Nielsen, Gill and Jupp.

Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):805–816, November 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2732].

Vansteelandt:2003:CIG

[2734] S. Vansteelandt and E. Goetghebeur. Causal inference with generalized structural mean models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):817– 835, November 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647585.

Dobbin:2003:ASD

[2735] Kevin K. Dobbin and Thomas A.

Louis. Accommodating stochastic departures from percentile invariance in causal models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):837–849, November 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647586.

Bortot:2003:EMC

[2736] Paola Bortot and Stuart Coles. Extremes of Markov chains with tail switching potential. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):851– 867, November 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647587.

Hall:2003:NMD

[2737] Peter Hall and Jiying Yin. Nonparametric methods for deconvolving multiperiodic functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):869– 886, November 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647588.

Fearnhead:2003:LIH

[2738] Paul Fearnhead and Peter Clifford. On-line inference for hidden Markov models via particle filters. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):887– 899, November 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647589.

Mao:2003:FKP

[2739] Wenxin Mao and Linda H. Zhao. Freeknot polynomial splines with confidence intervals. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):901–919, November 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647590.

Wang:2003:NES

[2740] Weijing Wang. Nonparametric estimation of the sojourn time distributions for a multipath model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):921– 935, November 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647591.

Cox:2003:GCA

[2741] D. R. Cox and Nanny Wermuth. A general condition for avoiding effect reversal after marginalization. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):937– 941, November 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647592.

Anonymous:2003:IAV

[2742] Anonymous. Index of authors, volume 65, 2003. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):943–945, November 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647582.

Brix:2003:CSP

[2743] Anders Brix and Peter J. Diggle. Corrigendum: Spatiotemporal prediction for log-Gaussian Cox processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65 (4):946, November 2003. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647593. See

[2614].

Anonymous:2003:BMd

[2744] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65 (4):??, ???? 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647594.

Anonymous:2003:CV

[2745] Anonymous. Contents of volume 65, 2003. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65(4):??, November 2003. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2003:FMd

[2746] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 65 (4):??, ???? 2003. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647583.

Davison:2004:RE

[2747] A. C. Davison and R. Henderson. Report of the Editors, 2003. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1): 1–2, February 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647622.

Zhu:2004:HTM

[2748] Hong-Tu Zhu and Heping Zhang. Hypothesis testing in mixture regression models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):3–16, February 2004. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647623.

Staudenmayer:2004:LPR

[2749] John Staudenmayer and David Ruppert. Local polynomial regression and simulation–extrapolation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):17– 30, February 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647624.

Carroll:2004:LOA

[2750] Raymond J. Carroll and Peter Hall.

Low order approximations in deconvolution and regression with errors in variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):31–46, February 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647625.

Roverato:2004:CPD

[2751] Alberto Roverato and Guido Consonni.

Compatible prior distributions for directed acyclic graph models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1): 47–61, February 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647626.

Ling:2004:ETS

[2752] Shiqing Ling. Estimation and testing stationarity for double-autoregressive models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):63–78, February 2004. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647627.

Schlather:2004:DDB

[2753] Martin Schlather, Paulo J. Ribeiro, Jr., and Peter J. Diggle. Detecting dependence between marks and locations of marked point processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):

79–93, February 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647628.

Chen:2004:TFM

[2754] Hanfeng Chen, Jiahua Chen, and John D. Kalbfleisch. Testing for a finite mixture model with two components. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):95–115, February 2004. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647629.

Chen:2004:PTI

[2755] Willa W. Chen and Rohit S. Deo. Power transformations to induce normality and their applications. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1): 117–130, February 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647630.

Haslett:2004:ADR

[2756] John Haslett and Dominic Dillane. Application of ‘delete = replace’ to deletion diagnostics for variance component estimation in the linear mixed model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):131–143, February 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647631.

Koenker:2004:PTT

[2757] Roger Koenker and Ivan Mizera. Penalized triograms: Total variation regularization for bivariate smoothing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1): 145–163, February 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647632.

Crainiceanu:2004:LRT

[2758] Ciprian M. Crainiceanu and David Ruppert. Likelihood ratio tests in linear mixed models with one variance component. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):165–185, February 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647633.

Storey:2004:SCC

[2759] John D. Storey, Jonathan E. Taylor, and David Siegmund. Strong control, conservative point estimation and simultaneous conservative consistency of false discovery rates: A unified approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):187–205, February 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647634.

Pipper:2004:EEP

[2760] Christian Bressen Pipper and Torben Martinussen. An estimating equation for parametric shared frailty models with marginal additive hazards. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1): 207–220, February 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647635.

Oh:2004:EGT

[2761] Hee-Seok Oh and Ta-Hsin Li. Estimation of global temperature fields from scattered observations by a spherical-wavelet-based spatially adaptive method. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):221–238, February 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/

stable/3647636. See corrigendum [2808].

Miloslavsky:2004:REA

[2762] Maja Miloslavsky, Sündüz Keleş, Mark J. van der Laan, and Steve Butler. Recurrent events analysis in the presence of time-dependent covariates and dependent censoring. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):239– 257, February 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647637.

Hanfelt:2004:CCL

[2763] John J. Hanfelt. Composite conditional likelihood for sparse clustered data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(1):259–273, February 2004. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647638.

Anonymous:2004:BMa

[2764] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series B (Statistical Methodology), 66 (1):??, ???? 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647639.

Anonymous:2004:FMa

[2765] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66 (1):??, ???? 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647621.

Stein:2004:ALL

[2766] Michael L. Stein, Zhiyi Chi, and Leah J. Welty. Approximating likelihoods for large spatial data sets. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2): 275–296, May 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647525.

Black:2004:NAC

[2767] M. A. Black. A note on the adaptive control of false discovery rates. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2): 297–304, May 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647526.

Prentice:2004:HBN

[2768] Ross L. Prentice, F. Zoe Moodie, and Jianrong Wu. Hazard-based nonparametric survivor function estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66 (2):305–319, May 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647527.

Gao:2004:SNL

[2769] Jiti Gao and Howell Tong. Semiparametric non-linear time series model selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2):321–336, May 2004. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647528.

Kim:2004:SSG

[2770] Young-Ju Kim and Chong Gu. Smoothing spline Gaussian regression: More scalable computation via efficient approximation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2):337–356, May 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647529.

Clarke:2004:IEL

[2771] Paul S. Clarke and Peter W. F. Smith. Interval estimation for loglinear models with one variable subject to non-ignorable non-response. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2): 357–368, May 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647530.

Roberts:2004:BIN

[2772] Gareth O. Roberts, Omiros Papaspiliopoulos, and Petros Dellaportas. Bayesian inference for non-Gaussian Ornstein–Uhlenbeck stochastic volatility processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2):369–393, May 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647531.

Cox:2004:SPS

[2773] D. R. Cox and Man Yu Wong. A simple procedure for the selection of significant effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2):395–400, May 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647532.

Huang:2004:SLR

[2774] Wenzheng Huang. Stepwise likelihood ratio statistics in sequential studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66 (2):401–409, May 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647533.

Tjelmeland:2004:ULO

[2775] Håkon Tjelmeland and Jo Eidsvik. On the use of local optimizations within Metropolis–Hastings updates. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2): 411–427, May 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647534.

Wager:2004:MSI

[2776] C. G. Wager, B. A. Coull, and N. Lange. Modelling spatial intensity for replicated inhomogeneous point patterns in brain imaging. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2): 429–446, May 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647535.

Heggland:2004:EFI

[2777] Knut Heggland and Arnoldo Frigessi. Estimating functions in indirect inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2):447–462, May 2004. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647536.

Huang:2004:INL

[2778] Jianhua Z. Huang and Lijian Yang. Identification of non-linear additive autoregressive models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2):463–477, May 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647537.

Zhang:2004:SAE

[2779] Li-Chun Zhang and Raymond L. Chambers. Small area estimates for cross-classifications. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(2):479–496, May 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647538.

Anonymous:2004:BMb

[2780] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series B (Statistical Methodology), 66 (2):??, ???? 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647539.

Anonymous:2004:FMb

[2781] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66 (2):??, ???? 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647524.

Heffernan:2004:CAM

[2782] Janet E. Heffernan and Jonathan A.

Tawn. A conditional approach for multivariate extreme values. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3): 497–529, August 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647494. See dis-

cussion [2783].

Anonymous:2004:DPH

[2783] Anonymous. Discussion on the paper by Heffernan and Tawn. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3): 530–546, August 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2782].

Johnstone:2004:WDP

[2784] Iain M. Johnstone, Gérard Kerkyacharian, Dominique Picard, and Marc Raimondo. Wavelet deconvolution in a periodic setting. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3):547–573, August 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647495.

Wolfe:2004:BVS

[2785] Patrick J. Wolfe, Simon J. Godsill, and Wee-Jing Ng. Bayesian variable selection and regularization for time– frequency surface estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3): 575–589, August 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647496.

Haario:2004:MCM

[2786] H. Haario, M. Laine, M. Lehtinen, E. Saksman, and J. Tamminen. Markov chain Monte Carlo methods for high dimensional inversion in remote sensing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3):591–607, August 2004. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647497.

Cornford:2004:BAS

[2787] Dan Cornford, Lehel Csató, David J. Evans, and Manfred Opper. Bayesian analysis of the scatterometer wind retrieval inverse problem: Some new approaches. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3):609–626, August 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.
org/stable/3647498. See discussion

[2788].

Anonymous:2004:DMS

[2788] Anonymous. Discussion on the meeting on ‘Statistical approaches to inverse problems’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3):627–652, August 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2787].

Eubank:2004:SSE

[2789] R. L. Eubank, Chunfeng Huang, Y. Muñoz Maldonado, Naisyin Wang, Suojin Wang, and R. J. Buchanan. Smoothing spline estimation in varyingcoefficient models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3):653–667, August 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647499.

Chen:2004:NMA

[2790] Rong Chen, Lijian Yang, and Christian Hafner. Nonparametric multistepahead prediction in time series analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3):669–686, August 2004. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647500.

Wermuth:2004:JRG

[2791] Nanny Wermuth and D. R. Cox. Joint response graphs and separation induced by triangular systems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3): 687–717, August 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647501.

Heller:2004:NTC

[2792] Glenn Heller and E. S. Venkatraman. A nonparametric test to compare survival distributions with covariate adjustment. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3):719–733, August 2004. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647502.

Muller:2004:MCI

[2793] Peter Müller, Fernando Quintana, and Gary Rosner. A method for combining inference across related nonparametric Bayesian models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3): 735–749, August 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647503.

Oakley:2004:PSA

[2794] Jeremy E. Oakley and Anthony O’Hagan. Probabilistic sensitivity analysis of complex models: A Bayesian approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3):751–769, August 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647504.

Fearnhead:2004:EFP

[2795] Paul Fearnhead and Loukia Meligkotsidou. Exact filtering for partially observed continuous time models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3): 771–789, August 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647505.

Lin:2004:ALD

[2796] Haiqun Lin, Daniel O. Scharfstein, and Robert A. Rosenheck. Analysis of longitudinal data with irregular, outcome-dependent follow-up. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(3): 791–813, August 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647506.

Anonymous:2004:BMc

[2797] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66 (3):??, ???? 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647507.

Anonymous:2004:FMc

[2798] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66 (3):??, ???? 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647493.

Friedman:2004:COS

[2799] Jerome H. Friedman and Jacqueline J.

Meulman. Clustering objects on subsets of attributes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(4):815–849, November 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647651.

Chaganty:2004:EGE

[2800] N. Rao Chaganty and Harry Joe. Efficiency of generalized estimating equations for binary responses. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(4):851– 860, November 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647652.

Hall:2004:NIA

[2801] Peter Hall and Juhyun Park. Nonparametric inference about service time distribution from indirect measurements. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(4):861–875, November 2004. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647653.

Rue:2004:AHG

[2802] Håvard Rue, Ingelin Steinsland, and Sveinung Erland. Approximating hidden Gaussian Markov random fields. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(4):877–892, November 2004. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647654.

Huber:2004:EGL

[2803] Philippe Huber, Elvezio Ronchetti, and Maria-Pia Victoria-Feser. Estimation of generalized linear latent variable models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(4):893–908, November 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647655.

Claeskens:2004:RLR

[2804] Gerda Claeskens. Restricted likelihood ratio lack-of-fit tests using mixed spline models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(4):909–926, November 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647656.

Barber:2004:RNR

[2805] Stuart Barber and Guy P. Nason. Real nonparametric regression using complex wavelets. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(4):927–939, November 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647657.

Fokianos:2004:MIS

[2806] Konstantinos Fokianos. Merging information for semiparametric density estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(4):941–958, November 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647658.

Gervini:2004:SMW

[2807] Daniel Gervini and Theo Gasser. Selfmodelling warping functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(4):959– 971, November 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647659.

Oh:2004:CEG

[2808] Hee-Seok Oh and Ta-Hsin Li. Corrigendum: Estimation of global temperature fields from scattered observations by a spherical-wavelet-based spatially adaptive method. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66 (4):973, November 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647660. See

[2761].

Anonymous:2004:IAV

[2809] Anonymous. Index of authors, volume 66, 2004. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66(4):974–977, November 2004. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647649.

Anonymous:2004:BMd

[2810] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66 (4):??, ???? 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647661.

Anonymous:2004:FMd

[2811] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 66 (4):??, ???? 2004. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647650.

Henderson:2005:RE

[2812] R. Henderson and A. Wood. Report of the Editors, 2004. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1): 1–2, February 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647596.

Ucinski:2005:ODD

[2813] Dariusz Uciński and Barbara Bogacka.

T -optimum designs for discrimination between two multiresponse dynamic models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1):3–18, February 2005. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647597.

Munk:2005:DBV

[2814] Axel Munk, Nicolai Bissantz, Thorsten Wagner, and Gudrun Freitag. On difference-based variance estimation in nonparametric regression when the covariate is high dimensional. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1): 19–41, February 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647598.

Nielsen:2005:SBP

[2815] Jens Perch Nielsen and Stefan Sperlich.

Smooth backfitting in practice. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1): 43–61, February 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647599.

He:2005:BLS

[2816] Wenqing He and Jerald F. Lawless. Bivariate location–scale models for regression analysis, with applications to lifetime data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1):63–78, February 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647600.

Berger:2005:JVE

[2817] Yves G. Berger and Chris J. Skinner. A jackknife variance estimator for unequal probability sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1): 79–89, February 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647601.

Tibshirani:2005:SSF

[2818] Robert Tibshirani, Michael Saunders, Saharon Rosset, Ji Zhu, and Keith Knight. Sparsity and smoothness via the fused lasso. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1):91–108, February 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647602.

Artiles:2005:IQT

[2819] L. M. Artiles, R. D. Gill, and M. I. Guţă. An invitation to quantum tomography. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1):109–134, February 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647603.

Huang:2005:ALD

[2820] Wenzheng Huang and Garrett M. Fitzmaurice. Analysis of longitudinal data unbalanced over time. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1):135– 155, February 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647604.

Dryver:2005:IUE

[2821] Arthur L. Dryver and Steven K.

Thompson. Improved unbiased estimators in adaptive cluster sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1): 157–166, February 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647605.

Saha:2005:ABL

[2822] Chandan Saha and Michael P. Jones.

Asymptotic bias in the linear mixed effects model under non-ignorable missing data mechanisms. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1):167– 182, February 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647606.

Spitzner:2005:RRS

[2823] Dan J. Spitzner. Risk-reducing shrinkage estimation for generalized linear models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1):183–196, February 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647607.

Storey:2005:APA

[2824] John D. Storey, Jonathan E. Taylor, and David Siegmund. Acknowledgment of priority: Acknowledgement of priority: Strong control, conservative point estimation and simultaneous conservative consistency of false discovery rates: a unified approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1): 197, February 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Storey:2005:APS

[2825] John D. Storey, Jonathan E. Taylor, and David Siegmund. Acknowledgement of priority: Strong control, conservative point estimation and simultaneous conservative consistency of false discovery rates: A unified approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(1):197, ???? 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647608.

Anonymous:2005:BMa

[2826] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (1):??, ???? 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647609.

Anonymous:2005:FMa

[2827] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (1):??, ???? 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647595.

Chen:2005:STR

[2828] Yuguo Chen, Junyi Xie, and Jun S.

Liu. Stopping-time resampling for sequential Monte Carlo methods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(2): 199–217, April 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647574.

Mallick:2005:BCT

[2829] Bani K. Mallick, Debashis Ghosh, and Malay Ghosh. Bayesian classification of tumours by using gene expression data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (2):219–234, April 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647575.

Caffo:2005:ABM

[2830] Brian S. Caffo, Wolfgang Jank, and Galin L. Jones. Ascent-based Monte Carlo expectation– maximization. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(2):235–251, April 2005. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647576.

Christensen:2005:SLT

[2831] Ole F. Christensen, Gareth O. Roberts, and Jeffrey S. Rosenthal. Scaling limits for the transient phase of local Metropolis–Hastings algorithms. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(2): 253–268, April 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647577.

Dellaportas:2005:MDC

[2832] Petros Dellaportas and Claudia Tarantola. Model determination for categorical data with factor level merging. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (2):269–283, April 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647578.

Li:2005:MFV

[2833] Lexin Li, R. Dennis Cook, and Christopher J. Nachtsheim. Model-free variable selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(2):285–299, April 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647579.

Zou:2005:RVS

[2834] Hui Zou and Trevor Hastie. Regularization and variable selection via the elastic net. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(2):301–320, April 2005. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647580.

Anonymous:2005:BMb

[2835] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (2):??, ???? 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647581.

Anonymous:2005:FMb

[2836] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (2):??, ???? 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647573.

Talih:2005:SLT

[2837] Makram Talih and Nicolas Hengartner. Structural learning with timevarying components: Tracking the cross-section of financial time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (3):321–341, June 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647663.

Samworth:2005:SCS

[2838] Richard Samworth. Small confidence sets for the mean of a spherically symmetric distribution. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(3):343–361, June 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647664.

Hall:2005:PBN

[2839] Peter Hall and Richard J. Samworth. Properties of bagged nearest neighbour classifiers. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(3):363–379, June 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647665.

Ling:2005:SWL

[2840] Shiqing Ling. Self-weighted least absolute deviation estimation for infinite variance autoregressive models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(3): 381–393, June 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647666.

Bladt:2005:SID

[2841] Mogens Bladt and Michael Sørensen. Statistical inference for discretely observed Markov jump processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(3): 395–410, June 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647667.

Owen:2005:VNF

[2842] Art B. Owen. Variance of the number of false discoveries. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(3): 411–426, June 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647668.

Hall:2005:GRH

[2843] Peter Hall, J. S. Marron, and Amnon Neeman. Geometric representation of high dimension, low sample size data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (3):427–444, June 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647669.

Beaumont:2005:CIS

[2844] Jean-François Beaumont. Calibrated imputation in surveys under a quasimodel-assisted approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(3): 445–458, June 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647670.

Anonymous:2005:BMc

[2845] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (3):??, ???? 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647671.

Anonymous:2005:FMc

[2846] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (3):??, ???? 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647662.

Copas:2005:LMU

[2847] John Copas and Shinto Eguchi. Local model uncertainty and incompletedata bias (with discussion). Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(4):459– 513, September 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647641.

Rosenbaum:2005:EDF

[2848] Paul R. Rosenbaum. An exact distribution-free test comparing two multivariate distributions based on adjacency. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(4):515–530, September 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647642.

Chiou:2005:EEE

[2849] Jeng-Min Chiou and Hans-Georg Müller. Estimated estimating equations: Semiparametric inference for clustered and longitudinal data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (4):531–553, September 2005. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647643.

Langaas:2005:EPT

[2850] Mette Langaas, Bo Henry Lindqvist, and Egil Ferkingstad. Estimating the proportion of true null hypotheses, with application to DNA microarray data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(4):555–572, September 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647644.

Hudgens:2005:NML

[2851] Michael G. Hudgens. On nonparametric maximum likelihood estimation with interval censoring and left truncation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(4):573–587, September 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647645.

Tsai:2005:NNN

[2852] Henghsiu Tsai and K. S. Chan. A note on non-negative continuous time processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(4):589–597, September 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647646.

Ferreira:2005:MDD

[2853] José T. A. S. Ferreira and Mark F. J. Steel. Modelling directional dispersion through hyperspherical logsplines. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(4):599–616, September 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647647.

Anonymous:2005:BMd

[2854] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series B (Statistical Methodology), 67 (4):??, ???? 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647648.

Anonymous:2005:FMd

[2855] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (4):??, ???? 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647640.

Baddeley:2005:RAS

[2856] A. Baddeley, R. Turner, J. Møller, and M. Hazelton. Residual analysis for spatial point processes (with discussion). Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(5):617–666, November 2005. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647612.

Stein:2005:SMR

[2857] Michael L. Stein. Statistical methods for regular monitoring data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(5):667– 687, November 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647613.

Johnson:2005:BFB

[2858] Valen E. Johnson. Bayes factors based on test statistics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(5):689– 701, November 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647614.

Tsai:2005:MLE

[2859] Henghsiu Tsai and K. S. Chan. Maximum likelihood estimation of linear continuous time long memory processes with discrete time data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(5):703– 716, November 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647615.

Hall:2005:BLM

[2860] Peter Hall and Julian Z. Wang.

Bayesian likelihood methods for estimating the end point of a distribution. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(5):717–729, November 2005. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647616.

Demiris:2005:BIS

[2861] Nikolaos Demiris and Philip D. O’Neill.

Bayesian inference for stochastic multitype epidemics in structured populations via random graphs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(5):731– 745, November 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647617.

Vovk:2005:GRS

[2862] Vladimir Vovk and Glenn Shafer. Good randomized sequential probability forecasting is always possible. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(5):747– 763, November 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647618.

Anonymous:2005:IAV

[2863] Anonymous. Index of authors, volume 67, 2005. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(5):765–767, November 2005. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647610.

Anonymous:2005:ARV

[2864] Anonymous. Addendum: Regularization and variable selection via the elastic net. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67(5):768, ???? 2005. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3647619.

Zou:2005:AAR

[2865] Hui Zou and Trevor Hastie. Addendum: Addendum: Regularization and variable selection via the elastic net. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (5):768, November 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2005:BMe

[2866] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (5):??, ???? 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647620.

Anonymous:2005:FMe

[2867] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 67 (5):??, ???? 2005. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647611.

Henderson:2006:RE

[2868] R. Henderson and A. Wood. Report of the Editors, 2005. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(1): 1–2, February 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647553.

Yao:2006:PSM

[2869] Fang Yao and Thomas C. M. Lee.

Penalized spline models for functional principal component analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(1): 3–25, February 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647554.

Jansen:2006:MPD

[2870] Maarten Jansen. Multiscale Poisson data smoothing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(1):27–48, February 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647555.

Yuan:2006:MSE

[2871] Ming Yuan and Yi Lin. Model selection and estimation in regression with grouped variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(1):49– 67, February 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647556.

Lin:2006:SEG

[2872] Xihong Lin and Raymond J. Carroll. Semiparametric estimation in general repeated measures problems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(1): 69–88, February 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647557.

Ni:2006:SDR

[2873] Liqiang Ni and R. Dennis Cook. Sufficient dimension reduction in regressions across heterogeneous subpopulations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(1):89–107, February 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647558.

Hall:2006:PFP

[2874] Peter Hall and Mohammad Hosseini-Nasab. On properties of functional principal components analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(1): 109–126, February 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647559.

Ma:2006:CDD

[2875] Zongming Ma, Xianchao Xie, and Zhi Geng. Collapsibility of distribution dependence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(1):127–133, February 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647560.

Prewitt:2006:BSL

[2876] Kathryn Prewitt and Sharon Lohr.

Bandwidth selection in local polynomial regression using eigenvalues. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(1): 135–154, February 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647561.

Anonymous:2006:BMa

[2877] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68 (1):??, ???? 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647562.

Anonymous:2006:FMa

[2878] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68 (1):??, ???? 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647552.

Bartolucci:2006:LIC

[2879] Francesco Bartolucci. Likelihood inference for a class of latent Markov models under linear hypotheses on the transition probabilities. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(2): 155–178, April 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647564.

Morris:2006:WBF

[2880] Jeffrey S. Morris and Raymond J. Carroll. Wavelet-based functional mixed models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(2):179–199, April 2006. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647565.

Delaigle:2006:NMS

[2881] Aurore Delaigle, Peter Hall, and Peihua Qiu. Nonparametric methods for solving the Berkson errors-in-variables problem. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(2):201–220, April 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647566.

Hall:2006:PBM

[2882] Peter Hall and Tapabrata Maiti. On parametric bootstrap methods for small area prediction. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(2):221–238, April 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647567.

Slud:2006:MSE

[2883] Eric V. Slud and Tapabrata Maiti.

Mean-squared error estimation in transformed Fay–Herriot models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(2): 239–257, April 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3647568.

Currie:2006:GLA

[2884] I. D. Currie, M. Durban, and P. H. C.

Eilers. Generalized linear array models with applications to multidimensional smoothing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(2):259–280, April 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3647569.

DiCiccio:2006:VSS

[2885] Thomas J. DiCiccio, Anna Clara Monti, and G. Alastair Young. Variance stabilization for a scalar parameter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(2):281–303, April 2006. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3647570.

Ray:2006:FCB

[2886] Shubhankar Ray and Bani Mallick.

Functional clustering by Bayesian wavelet methods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(2):305–332, April 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3647571.

Anonymous:2006:BMb

[2887] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68 (2):??, ???? 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647572.

Anonymous:2006:FMb

[2888] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68 (2):??, ???? 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3647563.

Beskos:2006:ECE

[2889] Alexandros Beskos, Omiros Papaspiliopoulos, Gareth O. Roberts, and Paul Fearnhead. Exact and computationally efficient likelihood-based estimation for discretely observed diffusion processes (with discussion). Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(3): 333–382, June 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3879281.

Sima:2006:RSM

[2890] Diana M. Sima and Sabine Van Huffel. Regularized semiparametric model identification with application to nuclear magnetic resonance signal quantification with unknown macromolecular base-line. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(3):383–409, June 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3879282.

DelMoral:2006:SMC

[2891] Pierre Del Moral, Arnaud Doucet, and Ajay Jasra. Sequential Monte Carlo samplers. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(3):411–436, June 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3879283.

Pilla:2006:TOR

[2892] Ramani S. Pilla, Annie Qu, and Catherine Loader. Testing for orderrestricted hypotheses in longitudinal data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(3):437–455, June 2006. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3879284.

Li:2006:NAC

[2893] Baibing Li. A new approach to cluster analysis: The clustering-function-based method. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(3):457–476, June 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3879285.

Goeman:2006:TAH

[2894] Jelle J. Goeman, Sara A. Van De Geer, and Hans C. Van Houwelingen. Testing against a high dimensional alternative. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(3):477–493, June 2006. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3879286.

Davison:2006:ILI

[2895] A. C. Davison, D. A. S. Fraser, and N. Reid. Improved likelihood inference for discrete data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(3): 495–508, June 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3879287.

Kim:2006:BMI

[2896] Jae Kwang Kim, J. Michael Brick, Wayne A. Fuller, and Graham Kalton. On the bias of the multiple-imputation variance estimator in survey sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68 (3):509–521, June 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3879288.

Gjestvang:2006:NRR

[2897] Christopher R. Gjestvang and Sarjinder Singh. A new randomized response model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(3):523–530, June 2006. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/3879289. See

correction [3570] and reply [3593].

Berger:2006:AJI

[2898] Yves G. Berger and J. N. K. Rao.

Adjusted jackknife for imputation under unequal probability sampling without replacement. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(3):531–547, June 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3879290.

Zhu:2006:ELC

[2899] Lixing Zhu and Liugen Xue. Empirical likelihood confidence regions in a partially linear single-index model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(3): 549–570, June 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3879291.

Anonymous:2006:BMc

[2900] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68 (3):??, ???? 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3879292.

Anonymous:2006:FMc

[2901] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68 (3):??, ???? 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3879280.

Brien:2006:MR

[2902] C. J. Brien and R. A. Bailey. Multiple randomizations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(4):571–609, September 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3879294.

Fryzlewicz:2006:HFE

[2903] Piotr Fryzlewicz and Guy P. Nason. Haar–Fisz estimation of evolutionary wavelet spectra. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(4):611– 634, September 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3879295.

Zhuang:2006:SOR

[2904] Jiancang Zhuang. Second-order residual analysis of spatiotemporal point processes and applications in model evaluation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(4):635–653, September 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3879296.

Efron:2006:MVC

[2905] Bradley Efron. Minimum volume confidence regions for a multivariate normal mean vector. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(4):655–670, September 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3879297.

Gervini:2006:FKS

[2906] Daniel Gervini. Free-knot spline smoothing for functional data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(4):671– 687, September 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3879298.

Hall:2006:AFD

[2907] Peter Hall and Céline Vial. Assessing the finite dimensionality of functional data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(4):689–705, September 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3879299.

Nielsen:2006:CNS

[2908] Bent Nielsen. Correlograms for nonstationary autoregressions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(4):707– 720, September 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3879300.

Anonymous:2006:BMd

[2909] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68 (4):??, ???? 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3879301.

Anonymous:2006:FMd

[2910] Anonymous. Front matter. Journal of the Royal Statistical Society.

Series B (Statistical Methodology), 68 (4):??, ???? 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3879293.

VanBellegem:2006:SEM

[2911] Sébastien Van Bellegem and Rainer Dahlhaus. Semiparametric estimation by model selection for locally stationary processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(5):721–746, November 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/3879272.

Kent:2006:CBQ

[2912] J. T. Kent, K. V. Mardia, and P. Mc-Donnell. The complex Bingham quartic distribution and shape analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(5):747–765, November 2006. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3879273.

Fearnhead:2006:EGS

[2913] Paul Fearnhead and Chris Sherlock. An exact Gibbs sampler for the Markovmodulated Poisson process. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(5):767– 784, November 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3879274.

Lawrence:2006:NDT

[2914] Earl Lawrence, George Michailidis, and Vijayan N. Nair. Network delay tomography using flexicast experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(5):785–813, November 2006. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3879275.

Cheng:2006:BCE

[2915] Jing Cheng and Dylan S. Small.

Bounds on causal effects in three-arm trials with non-compliance. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(5):815– 836, November 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3879276.

Antoniadis:2006:FWK

[2916] Anestis Antoniadis, Efstathios Paparoditis, and Theofanis Sapatinas. A functional wavelet–kernel approach for time series prediction. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(5):837– 857, November 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/3879277.

Neuhaus:2006:SBW

[2917] John M. Neuhaus and Charles E. Mc-Culloch. Separating between- and within-cluster covariate effects by using conditional and partitioning methods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(5):859–872, November 2006. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/3879278.

Anonymous:2006:IAV

[2918] Anonymous. Index of authors, volume 68, 2006. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68(5):873–875, November 2006. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/3879270.

Anonymous:2006:BMe

[2919] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68 (5):??, ???? 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3879279.

Anonymous:2006:FMe

[2920] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 68 (5):??, ???? 2006. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/3879271.

Robert:2007:RE

[2921] C. P. Robert and A. T. A. Wood. Report of the Editors, 2006. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(1): 1–2, February 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/4623250.

Allassonniere:2007:TCS

[2922] S. Allassonnière, Y. Amit, and A. Trouvé. Towards a coherent statistical framework for dense deformable template estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(1):3–29, February 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://www.jstor.
org/stable/4623251.

Guan:2007:LSF

[2923] Yongtao Guan and Michael Sherman.

On least squares fitting for stationary spatial point processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(1): 31–49, February 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623252.

Butler:2007:OAD

[2924] Neil A. Butler and Victorino M.

Ramos. Optimal additions to and deletions from two-level orthogonal arrays. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(1):51–61, February 2007. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/4623253.

Wang:2007:RCA

[2925] Hansheng Wang, Guodong Li, and Chih-Ling Tsai. Regression coefficient and autoregressive order shrinkage and selection via the lasso. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(1): 63–78, February 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623254.

Lai:2007:MRA

[2926] Tze Leung Lai and Dylan Small. Marginal regression analysis of longitudinal data with time-dependent covariates: A generalized methodof-moments approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(1):79– 99, February 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/4623255.

Qin:2007:ELB

[2927] Jing Qin and Biao Zhang. Empiricallikelihood-based inference in missing response problems and its application in observational studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(1):101– 122, February 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/4623256.

Anonymous:2007:BMa

[2928] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (1):??, ???? 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/4623257.

Anonymous:2007:FMa

[2929] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (1):??, ???? 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/4623249.

Chatterjee:2007:MLI

[2930] Nilanjan Chatterjee and Yi-Hau Chen.

Maximum likelihood inference on a mixed conditionally and marginally specified regression model for genetic epidemiologic studies with two-phase sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(2):123–142, April 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/4623259.

Yuan:2007:NNG

[2931] Ming Yuan and Yi Lin. On the nonnegative garrotte estimator. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(2): 143–161, April 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623260.

Dunson:2007:BDR

[2932] David B. Dunson, Natesh Pillai, and Ju-Hyun Park. Bayesian density regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(2):163–183, April 2007. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/4623261.

Wang:2007:FCN

[2933] Yong Wang. On fast computation of the non-parametric maximum likelihood estimate of a mixing distribution. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(2):185–198, April 2007. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/4623262.

Geneletti:2007:IDI

[2934] Sara Geneletti. Identifying direct and indirect effects in a non-counterfactual framework. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(2):199–215, April 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4623263.

Boldi:2007:MMM

[2935] M.-O. Boldi and A. C. Davison. A mixture model for multivariate extremes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (2):217–229, April 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/4623264.

Lopez-Fidalgo:2007:OED

[2936] J. López-Fidalgo, C. Tommasi, and P. C. Trandafir. An optimal experimental design criterion for discriminating between non-normal models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(2): 231–242, April 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/4623265.

Gneiting:2007:PFC

[2937] Tilmann Gneiting, Fadoua Balabdaoui, and Adrian E. Raftery. Probabilistic forecasts, calibration and sharpness. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (2):243–268, April 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623266.

Chopin:2007:IMC

[2938] Nicolas Chopin. Inference and model choice for sequentially ordered hidden Markov models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(2):269–284, April 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/4623267.

Anonymous:2007:BMb

[2939] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (2):??, ???? 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623268.

Anonymous:2007:FMb

[2940] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (2):??, ???? 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623258.

Li:2007:MCS

[2941] Yi Li, Ram C. Tiwari, and Subharup Guha. Mixture cure survival models with dependent censoring. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(3): 285–306, June 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623270.

Rotnitzky:2007:AFT

[2942] Andrea Rotnitzky, Andres Farall, Andrea Bergesio, and Daniel Scharfstein. Analysis of failure time data under competing censoring mechanisms. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (3):307–327, June 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623271.

Yuan:2007:DRC

[2943] Ming Yuan, Ali Ekici, Zhaosong Lu, and Renato Monteiro. Dimension reduction and coefficient estimation in multivariate linear regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(3): 329–346, June 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623272.

Storey:2007:ODP

[2944] John D. Storey. The optimal discovery procedure: A new approach to simultaneous significance testing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(3): 347–368, June 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623273.

Field:2007:BCD

[2945] C. A. Field and A. H. Welsh. Bootstrapping clustered data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(3):

369–390, June 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623274.

Wu:2007:ITT

[2946] Wei Biao Wu and Zhibiao Zhao. Inference of trends in time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(3): 391–410, June 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623275.

Wang:2007:OLA

[2947] Molin Wang and John J. Hanfelt.

Orthogonal locally ancillary estimating functions for matched pair studies and errors in covariates. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(3): 411–428, June 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623276.

Hall:2007:SEF

[2948] Peter Hall and Yanyuan Ma. Semiparametric estimators of functional measurement error models with unknown error. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(3):429–446, June 2007. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/4623277.

Nott:2007:CSP

[2949] David J. Nott and Anthony Y. C.

Kuk. Coefficient sign prediction methods for model selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(3): 447–461, June 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623278.

vanderLaan:2007:ETE

[2950] Mark J. van der Laan, Alan Hubbard, and Nicholas P. Jewell. Estimation of treatment effects in randomized trials with non-compliance and a dichotomous outcome. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(3): 463–482, June 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623279.

Bissantz:2007:NPC

[2951] Nicolai Bissantz, Lutz Dümbgen, Hajo Holzmann, and Axel Munk. Nonparametric confidence bands in deconvolution density estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(3): 483–506, June 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623280.

Anonymous:2007:BMc

[2952] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (3):??, ???? 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623281.

Anonymous:2007:FMc

[2953] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (3):??, ???? 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623269.

Zeng:2007:MLE

[2954] D. Zeng and D. Y. Lin. Maximum likelihood estimation in semiparametric regression models with censored data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(4):507–564, September 2007. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/4623283.

Cuzick:2007:EET

[2955] Jack Cuzick, Peter Sasieni, Jonathan Myles, and Jonathan Tyrer. Estimating the effect of treatment in a proportional hazards model in the presence of non-compliance and contamination. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(4):565–588, September 2007. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/4623284.

Fearnhead:2007:LIM

[2956] Paul Fearnhead and Zhen Liu. Online inference for multiple changepoint problems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(4):589–605, September 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/4623285.

Roy:2007:CRA

[2957] Vivekananda Roy and James P. Hobert. Convergence rates and asymptotic standard errors for Markov chain Monte Carlo algorithms for Bayesian probit regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(4):607–623, September 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4623286.

Ma:2007:NGL

[2958] Renjun Ma and Bent Jørgensen. Nested generalized linear mixed models: An orthodox best linear unbiased predictor approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(4):625– 641, September 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/4623287.

Genton:2007:SIE

[2959] Marc G. Genton and Peter Hall. Statistical inference for evolving periodic functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(4):643–657, September 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4623288.

Park:2007:RPA

[2960] Mee Young Park and Trevor Hastie. L_{1}-regularization path algorithm for generalized linear models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(4):659– 677, September 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623289.

Chiou:2007:FCI

[2961] Jeng-Min Chiou and Pai-Ling Li. Functional clustering and identifying substructures of longitudinal data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(4):679– 699, September 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623290.

Rasmussen:2007:CTM

[2962] Jakob G. Rasmussen, Jesper Møller, Brian H. Aukema, Kenneth F. Raffa, and Jun Zhu. Continuous time modelling of dynamical spatial lattice data observed at sparsely distributed times. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(4):701–713, September 2007. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/4623291.

Lijoi:2007:CRB

[2963] Antonio Lijoi, Ramsés H. Mena, and Igor Prünster. Controlling the reinforcement in Bayesian non-parametric mixture models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(4):715–740, September 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/4623292.

Anonymous:2007:BMd

[2964] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series B (Statistical Methodology), 69 (4):??, ???? 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/4623293.

Anonymous:2007:FMd

[2965] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (4):??, ???? 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/4623282.

Ramsay:2007:PED

[2966] J. O. Ramsay, G. Hooker, D. Campbell, and J. Cao. Parameter estimation for differential equations: A generalized smoothing approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(5):741– 796, November 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/4623296.

Gustafson:2007:MEM

[2967] Paul Gustafson. Measurement error modelling with an approximate instrumental variable. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(5):797–815, November 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4623297.

Jin:2007:OFC

[2968] Xiaoping Jin, Sudipto Banerjee, and Bradley P. Carlin. Order-free coregionalized areal data models with application to multiple-disease mapping.

Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(5):817–838, November 2007. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/4623298.

Mojirsheibani:2007:SCM

[2969] Majid Mojirsheibani and Zahra Montazeri. Statistical classification with missing covariates. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(5):839–857, November 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/4623299.

Carroll:2007:NPR

[2970] Raymond J. Carroll, Aurore Delaigle, and Peter Hall. Non-parametric regression estimation from data contaminated by a mixture of Berkson and classical errors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(5):859–878, November 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/4623300.

Jemiai:2007:SET

[2971] Yannis Jemiai, Andrea Rotnitzky, Bryan E. Shepherd, and Peter B. Gilbert. Semiparametric estimation of treatment effects given base-line covariates on an outcome measured after a post-randomization event occurs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(5):879–901, November 2007. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/4623301.

Dette:2007:NTP

[2972] Holger Dette, Natalie Neumeyer, and Ingrid Van Keilegom. A new test for the parametric form of the variance function in non-parametric regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(5):903–917, November 2007. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/4623302.

Chen:2007:CSE

[2973] Hua Chen, Zhi Geng, and Jinzhu Jia. Criteria for surrogate end points. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (5):919–932, November 2007. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/4623303.

Anonymous:2007:IAV

[2974] Anonymous. Index of authors, volume 69, 2007. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(5):933–935, November 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/4623294.

Anonymous:2007:CV

[2975] Anonymous. Contents of volume 69, 2007. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69(5):937–938, November 2007. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2007:BMe

[2976] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series B (Statistical Methodology), 69 (5):??, ???? 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623304.

Anonymous:2007:FMe

[2977] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 69 (5):??, ???? 2007. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/4623295.

Robert:2008:RE

[2978] C. P. Robert and A. T. A. Wood.

Report of the Editors, 2007. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1): 1–2, February 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203807.

Hsieh:2008:RAB

[2979] Jin-Jian Hsieh, Weijing Wang, and A. Adam Ding. Regression analysis based on semicompeting risks data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70 (1):3–20, February 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203808.

Cruyff:2008:ARR

[2980] Maarten J. L. F. Cruyff, Ardo Van Den Hout, and Peter G. M. Van Der Heijden. The analysis of randomized response sum score variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):

21–30, February 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203809.

Gomes:2008:TIE

[2981] M. Ivette Gomes, Laurens De Haan, and Lígia Henriques Rodrigues. Tail index estimation for heavy-tailed models: Accommodation of bias in weighted log-excesses. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):31–52, February 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20203810.

Meier:2008:GLL

[2982] Lukas Meier, Sara Van De Geer, and Peter Bühlmann. The group lasso for logistic regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):53– 71, February 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203811.

Haneuse:2008:CEC

[2983] Sebastien J.-P. A. Haneuse and Jonathan C. Wakefield. The combination of ecological and case–control data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):73–93, February 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/20203812.

Ray:2008:MSH

[2984] Surajit Ray and Bruce G. Lindsay. Model selection in high dimensions: A quadratic-risk-based approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1): 95–118, February 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203813.

Booth:2008:CUO

[2985] James G. Booth, George Casella, and James P. Hobert. Clustering using objective functions and stochastic search. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):119–139, February 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/20203814.

Cai:2008:PLH

[2986] Jianwen Cai, Jianqing Fan, Jiancheng Jiang, and Haibo Zhou. Partially linear hazard regression with varying coefficients for multivariate survival data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):141–158, February 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/20203815.

Hall:2008:TMR

[2987] Peter Hall, Yvonne Pittelkow, and Malay Ghosh. Theoretical measures of relative performance of classifiers for high dimensional data with small sample sizes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):159–173, February 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/20203816.

Guan:2008:VES

[2988] Yongtao Guan. Variance estimation for statistics computed from inhomogeneous spatial point processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1): 175–190, February 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203817.

Qiu:2008:TSP

[2989] Peihua Qiu and Jun Sheng. A twostage procedure for comparing hazard rate functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):191–208, February 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20203818.

Cressie:2008:FRK

[2990] Noel Cressie and Gardar Johannesson. Fixed rank kriging for very large spatial data sets. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):209– 226, February 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203819.

Craig:2008:NRM

[2991] Peter Craig. A new reconstruction of multivariate normal orthant probabilities. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):227–243, February 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20203820.

Didelez:2008:GMM

[2992] Vanessa Didelez. Graphical models for marked point processes based on local independence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):245– 264, February 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203821.

Opsomer:2008:NPS

[2993] J. D. Opsomer, G. Claeskens, M. G.

Ranalli, G. Kauermann, and F. J. Breidt. Non-parametric small area estimation using penalized spline regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(1):265–286, February 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/20203822.

Anonymous:2008:BMa

[2994] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70 (1):??, ???? 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203823.

Anonymous:2008:FMa

[2995] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70 (1):??, ???? 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203806.

Drton:2008:BMM

[2996] Mathias Drton and Thomas S. Richardson. Binary models for marginal independence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(2):287–309, April 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/20203826.

Chaudhuri:2008:GLM

[2997] Sanjay Chaudhuri, Mark S. Handcock, and Michael S. Rendall. Generalized linear models incorporating population level information: An empiricallikelihood-based approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(2): 311–328, April 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203827.

Qin:2008:ELB

[2998] Jing Qin and Biao Zhang. Empiricallikelihood-based difference-in-differences estimators. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(2):329–349, April 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/20203828.

Johnson:2008:VSS

[2999] Brent A. Johnson. Variable selection in semiparametric linear regression with censored data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(2): 351–370, April 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203829.

Molenberghs:2008:EMR

[3000] Geert Molenberghs, Caroline Beunckens, Cristina Sotto, and Michael G. Kenward. Every missingness not at random model has a missingness at random counterpart with equal fit. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70 (2):371–388, April 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203830.

Hallin:2008:SEI

[3001] Marc Hallin, Catherine Vermandele, and Bas J. M. Werker. Semiparametrically efficient inference based on signs and ranks for median-restricted models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(2):389–412, April 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/20203831.

Polson:2008:PFS

[3002] Nicholas G. Polson, Jonathan R.

Stroud, and Peter Müller. Practical filtering with sequential parameter learning. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(2):413–428, April 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/20203832.

Foster:2008:IPS

[3003] Dean P. Foster and Robert A. Stine.

α-investing: A procedure for sequential control of expected false discoveries. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(2):429–444, April 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/20203833.

Catchpole:2008:NMA

[3004] E. A. Catchpole, B. J. T. Morgan, and G. Tavecchia. A new method for analysing discrete life history data with missing covariate values. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(2): 445–460, April 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203834.

Anonymous:2008:BMb

[3005] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70 (2):??, ???? 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203835.

Anonymous:2008:FMb

[3006] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70 (2):??, ???? 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203825.

Jin:2008:PNZ

[3007] Jiashun Jin. Proportion of non-zero normal means: Universal oracle equivalences and uniformly consistent estimators. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(3):461–493, July 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/20203838.

Wood:2008:FSD

[3008] Simon N. Wood. Fast stable direct fitting and smoothness selection for generalized additive models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(3): 495–518, July 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203839.

Xia:2008:SAC

[3009] Yingcun Xia. A semiparametric approach to canonical analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(3): 519–543, July 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203840.

Nicholls:2008:DAT

[3010] Geoff K. Nicholls and Russell D. Gray.

Dated ancestral trees from binary trait data and their application to the diversification of languages. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(3): 545–566, July 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203841.

Heaton:2008:WLS

[3011] T. J. Heaton and B. W. Silverman. A wavelet- or lifting-scheme-based imputation method. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(3): 567–587, July 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203842.

Friel:2008:MLE

[3012] N. Friel and A. N. Pettitt. Marginal likelihood estimation via power posteriors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(3):589–607, July 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/20203843.

Dette:2008:NCN

[3013] Holger Dette and Stanislav Volgushev. Non-crossing non-parametric estimates of quantile curves. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(3): 609–627, July 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203844.

Pilla:2008:CRE

[3014] Ramani S. Pilla, Yongdai Kim, and Hakbae Lee. On casting random-effects models in a survival framework. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(3): 629–642, July 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203845.

Anonymous:2008:BMc

[3015] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70 (3):??, ???? 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203846.

Anonymous:2008:FMc

[3016] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70 (3):??, ???? 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203837.

McCullagh:2008:SBL

[3017] Peter McCullagh. Sampling bias and logistic models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(4):643–677, September 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/20203849.

Fan:2008:MMV

[3018] Jianqing Fan, Mingjin Wang, and Qiwei Yao. Modelling multivariate volatilities via conditionally uncorrelated components. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(4):679–702, September 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/20203850.

Hall:2008:MSG

[3019] Peter Hall, Hans-Georg Müller, and Fang Yao. Modelling sparse generalized longitudinal observations with latent Gaussian processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(4):703– 723, September 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203851.

Hall:2008:NPI

[3020] Peter Hall and Tapabrata Maiti. Nonparametric inference for clustered binary and count data when only summary information is available. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(4):725– 738, September 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203852.

Powojowski:2008:ISA

[3021] Miro R. Powojowski. Isotropic spectral additive models of the covariogram. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(4):739–753, September 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/20203853.

Fearnhead:2008:PFP

[3022] Paul Fearnhead, Omiros Papaspiliopoulos, and Gareth O. Roberts. Particle filters for partially observed diffusions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(4):755–777, September 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/20203854.

Cuesta-Albertos:2008:REN

[3023] J. A. Cuesta-Albertos, C. Matrán, and A. Mayo-Iscar. Robust estimation in the normal mixture model based on robust clustering. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(4):779– 802, September 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203855.

Chen:2008:ISE

[3024] Song Xi Chen, Denis H. Y. Leung, and Jing Qin. Improving semiparametric estimation by using surrogate data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(4):803–823, September 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/20203856.

Banerjee:2008:GPP

[3025] Sudipto Banerjee, Alan E. Gelfand, Andrew O. Finley, and Huiyan Sang. Gaussian predictive process models for large spatial data sets. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(4):825– 848, September 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203857.

Anonymous:2008:BMd

[3026] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(4): ??, September 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203858.

Anonymous:2008:FMd

[3027] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(4):

??, September 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203848.

Fan:2008:SIS

[3028] Jianqing Fan and Jinchi Lv. Sure independence screening for ultrahigh dimensional feature space. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):849– 911, November 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203862.

Lombardia:2008:SIG

[3029] María José Lombardía and Stefan Sperlich. Semiparametric inference in generalized mixed effects models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):913– 930, November 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203863.

Wood:2008:SFS

[3030] Simon N. Wood, Mark V. Bravington, and Sharon L. Hedley. Soap film smoothing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):931–955, November 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20203864.

Smith:2008:SMG

[3031] Jim Q. Smith, Paul E. Anderson, and Silvia Liverani. Separation measures and the geometry of Bayes factor selection for classification. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):957– 980, November 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203865.

Bayarri:2008:GJD

[3032] M. J. Bayarri and G. García-Donato.

Generalization of Jeffreys divergencebased priors for Bayesian hypothesis testing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):981–1003, November 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/20203866.

Hojsgaard:2008:GGM

[3033] Søren Højsgaard and Steffen L. Lauritzen. Graphical Gaussian models with edge and vertex symmetries. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):1005– 1027, November 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203867.

Zhou:2008:NPH

[3034] Xiao-Hua Zhou, Huazhen Lin, and Eric Johnson. Non-parametric heteroscedastic transformation regression models for skewed data with an application to health care costs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):1029– 1047, November 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203868.

Goetgeluk:2008:ECD

[3035] Sylvie Goetgeluk, Stijn Vansteelandt, and Els Goetghebeur. Estimation of controlled direct effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):1049– 1066, November 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203869.

Leng:2008:CRM

[3036] Chenlei Leng, Peide Shi, and Chih-Ling Tsai. Clarification: Regression model selection: A residual likelihood approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):1067, November 2008. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/20203870.

Anonymous:2008:IAV

[3037] Anonymous. Index of authors, volume 70, 2008. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):1069–1071, November 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2008:CV

[3038] Anonymous. Contents of volume 70, 2008. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):1073–1074, November 2008. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2008:BMe

[3039] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5):

??, November 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203871.

Anonymous:2008:FMe

[3040] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70(5): ??, November 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203861.

Anonymous:2008:VI

[3041] Anonymous. Volume information.

Journal of the Royal Statistical Society. Series B (Statistical Methodology), 70 (5):??, November 2008. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203860.

Casella:2009:RE

[3042] G. Casella and C. P. Robert. Report of the Editors, 2008. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1): 1–2, January 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203874.

Lin:2009:PEE

[3043] Fengchang Lin and Jason P. Fine.

Pseudomartingale estimating equations for modulated renewal process models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1):3–23, January 2009. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/20203875.

Bissantz:2009:TLF

[3044] Nicolai Bissantz, Gerda Claeskens, Hajo Holzmann, and Axel Munk. Testing for lack of fit in inverse regression — with applications to biophotonic imaging. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1):25–48, January 2009. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/20203876.

Pokern:2009:PEP

[3045] Yvo Pokern, Andrew M. Stuart, and Petter Wiberg. Parameter estimation for partially observed hypoelliptic diffusions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1):49–73, January 2009. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/20203877.

Maity:2009:TSM

[3046] Arnab Maity, Raymond J. Carroll, Enno Mammen, and Nilanjan Chatterjee. Testing in semiparametric models with interaction, with applications to gene–environment interactions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1): 75–96, January 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203878.

Jansen:2009:MMD

[3047] Maarten Jansen, Guy P. Nason, and B. W. Silverman. Multiscale methods for data on graphs and irregular multidimensional situations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1):

97–125, January 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203879.

James:2009:DCB

[3048] Gareth M. James, Peter Radchenko, and Jinchi Lv. DASSO: Connections between the Dantzig selector and lasso. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1):127–142, January 2009. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/20203880.

Hu:2009:BMS

[3049] Jianhua Hu and Valen E. Johnson.

Bayesian model selection using test statistics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1):143–158, January 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/20203881.

Shively:2009:BAN

[3050] Thomas S. Shively, Thomas W. Sager, and Stephen G. Walker. A Bayesian approach to non-parametric monotone function estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1):159–175, January 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/20203882.

Wang:2009:CMS

[3051] Lan Wang and Annie Qu. Consistent model selection and data-driven smooth tests for longitudinal data in the estimating equations approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1):177–190, January 2009. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/20203883.

Matsuda:2009:FAI

[3052] Yasumasa Matsuda and Yoshihiro Yajima. Fourier analysis of irregularly spaced data on R^{d}. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1):191– 217, January 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203884.

Ramos:2009:NCM

[3053] Alexandra Ramos and Anthony Ledford. A new class of models for bivariate joint tails. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1):219–241, January 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/20203885.

Lee:2009:NPT

[3054] Myoung-Jae Lee. Non-parametric tests for distributional treatment effect for randomly censored responses. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1): 243–264, January 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/20203886.

Hwang:2009:EBC

[3055] J. T. Gene Hwang, Jing Qiu, and Zhigen Zhao. Empirical Bayes confidence intervals shrinking both means and variances. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1):265–285, January 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/20203887.

Bondell:2009:SIR

[3056] Howard D. Bondell and Lexin Li.

Shrinkage inverse regression estimation for model-free variable selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1): 287–299, January 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203888.

Garcia-Escudero:2009:RLC

[3057] L. A. García-Escudero, A. Gordaliza, R. San Martín, S. Van Aelst, and R. Zamar. Robust linear clustering. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(1): 301–318, January 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203889.

Anonymous:2009:BMa

[3058] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71 (1):??, January 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203890.

Anonymous:2009:FMa

[3059] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71 (1):??, January 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/20203873.

Rue:2009:ABI

[3060] Håvard Rue, Sara Martino, and Nicolas Chopin. Approximate Bayesian inference for latent Gaussian models by using integrated nested Laplace approximations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(2):319–392, April 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40247579.

Sun:2009:LSM

[3061] Wenguang Sun and T. Tony Cai. Large-scale multiple testing under dependence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(2):393–424, April 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40247580.

Wang:2009:VEA

[3062] Yuedong Wang, Yanyuan Ma, and Raymond J. Carroll. Variance estimation in the analysis of microarray data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(2): 425–445, April 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40247581.

Riani:2009:FUN

[3063] Marco Riani, Anthony C. Atkinson, and Andrea Cerioli. Finding an unknown number of multivariate outliers. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71 (2):447–466, April 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40247582.

Drost:2009:EEA

[3064] Feike C. Drost, Ramon van den Akker, and Bas J. M. Werker. Efficient estimation of auto-regression parameters and innovation distributions for semiparametric integer-valued AR(p) models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(2):467–485, April 2009. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/40247583.

Kauermann:2009:SAR

[3065] Göran Kauermann, Tatyana Krivobokova, and Ludwig Fahrmeir. Some asymptotic results on generalized penalized spline smoothing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(2):487–503, April 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40247584.

Reiss:2009:SPS

[3066] Philip T. Reiss and R. Todd Ogden.

Smoothing parameter selection for a class of semiparametric linear models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71 (2):505–523, April 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40247585.

Zhu:2009:DWP

[3067] Li-Ping Zhu and Li-Xing Zhu. On distribution-weighted partial least squares with diverging number of highly correlated predictors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(2): 525–548, April 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40247586.

Anonymous:2009:BMb

[3068] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71 (2):??, April 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40247587.

Anonymous:2009:FMb

[3069] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71 (2):??, April 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40247578.

Tyler:2009:ICO

[3070] David E. Tyler, Frank Critchley, Lutz Dümbgen, and Hannu Oja. Invariant co-ordinate selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(3): 549–592, June 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40247589.

Cappe:2009:LEM

[3071] Olivier Cappé and Eric Moulines.

On-line expectation–maximization algorithm for latent data models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(3): 593–613, June 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40247590.  See

corrigenda [3175, 3176].

Witten:2009:CRR

[3072] Daniela M. Witten and Robert Tibshirani. Covariance-regularized regression and classification for high dimensional problems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(3):615–636, June 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40247591.

Rizopoulos:2009:FEL

[3073] Dimitris Rizopoulos, Geert Verbeke, and Emmanuel Lesaffre. Fully exponential Laplace approximations for the joint modelling of survival and longitudinal data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(3):637–654, June 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40247592.

Audrino:2009:SFV

[3074] Francesco Audrino and Peter Bühlmann.

Splines for financial volatility. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(3): 655–670, June 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40247593.

Wang:2009:STP

[3075] Hansheng Wang, Bo Li, and Chenlei Leng. Shrinkage tuning parameter selection with a diverging number of parameters. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(3):671–683, June 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40247594.

Waagepetersen:2009:TSE

[3076] Rasmus Waagepetersen and Yongtao Guan. Two-step estimation for inhomogeneous spatial point processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71 (3):685–702, June 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40247595.

Hall:2009:DMN

[3077] Peter Hall and Tapabrata Maiti. Deconvolution methods for nonparametric inference in two-level mixed models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(3):703–718, June 2009. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/40247596.

Commenges:2009:GDS

[3078] Daniel Commenges and Anne Gégout-Petit. A general dynamical statistical model with causal interpretation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(3): 719–736, June 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40247597.

Kluppelberg:2009:CSA

[3079] Claudia Klüppelberg and Gabriel Kuhn. Copula structure analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(3): 737–753, June 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40247598.

Anonymous:2009:BMc

[3080] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71 (3):??, June 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40247599.

Anonymous:2009:FMc

[3081] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71 (3):??, June 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40247588.

Petrone:2009:HDM

[3082] Sonia Petrone, Michele Guindani, and Alan E. Gelfand. Hybrid Dirichlet mixture models for functional data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(4):755–782, September 2009. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/40247601.

Hall:2009:TMA

[3083] Peter Hall, D. M. Titterington, and Jing-Hao Xue. Tilting methods for assessing the influence of components in a classifier. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(4):783–803, September 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40247602.

Wiens:2009:RDD

[3084] Douglas P. Wiens. Robust discrimination designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(4):805–829, September 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40247603.

Dette:2009:BFD

[3085] Holger Dette and Efstathios Paparoditis. Bootstrapping frequency domain tests in multivariate time series with an application to comparing spectral densities. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(4):831– 857, September 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40247604.

Lu:2009:AVC

[3086] Zudi Lu, Dag Johan Steinskog, Dag Tjøstheim, and Qiwei Yao. Adaptively varying-coefficient spatiotemporal models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(4):859–880, September 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40247605.

Cheng:2009:SEI

[3087] Jing Cheng, Jing Qin, and Biao Zhang.

Semiparametric estimation and inference for distributional and general treatment effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(4):881–904, September 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40247606.

Anonymous:2009:BMd

[3088] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(4): ??, September 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40247607.

Anonymous:2009:FMd

[3089] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(4): ??, September 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40247600.

Guindani:2009:BDP

[3090] Michele Guindani, Peter Müller, and Song Zhang. A Bayesian discovery procedure. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(5):905–925, November 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40541562.

Berkes:2009:DCM

[3091] István Berkes, Robertas Gabrys, Lajos Horváth, and Piotr Kokoszka. Detecting changes in the mean of functional observations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(5):927–946, November 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40541563.

Wang:2009:CIO

[3092] Weiwei Wang, Daniel Scharfstein, Zhiqiang Tan, and Ellen J. MacKenzie. Causal inference in outcome-dependent two-phase sampling designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(5):947– 969, November 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40541564.

Hoff:2009:HEP

[3093] Peter D. Hoff. A hierarchical eigenmodel for pooled covariance estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(5):971–992, November 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40541565.

Favaro:2009:BNP

[3094] Stefano Favaro, Antonio Lijoi, Ramsés H. Mena, and Igor Prünster. Bayesian non-parametric inference for species variety with a two-parameter Poisson– Dirichlet process prior. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(5):993– 1008, November 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40541566.

Ravikumar:2009:SAM

[3095] Pradeep Ravikumar, John Lafferty, Han Liu, and Larry Wasserman. Sparse additive models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(5):1009–1030, November 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://www.jstor.
org/stable/40541567.

Finner:2009:CFE

[3096] Helmut Finner and Veronika Gontscharuk.

Controlling the familywise error rate with plug-in estimator for the proportion of true null hypotheses. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(5):1031– 1048, November 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40541568.

Anonymous:2009:IAV

[3097] Anonymous. Index of authors, volume 71, 2009. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(5):1049–1052, November 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2009:CV

[3098] Anonymous. Contents of volume 71, 2009. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(5):1053–1054, November 2009. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2009:BMe

[3099] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(5): ??, November 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40541569.

Anonymous:2009:FMe

[3100] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71(5): ??, November 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40541561.

Anonymous:2009:VI

[3101] Anonymous. Volume information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 71 (5):??, November 2009. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40541560.

Casella:2010:RE

[3102] G. Casella and C. P. Robert. Report of the Editors — 2009. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(1): 1–2, January 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40541571.

Chun:2010:SPL

[3103] Hyonho Chun and Sündüz Keleş.

Sparse partial least squares regression for simultaneous dimension reduction and variable selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(1): 3–25, January 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40541572.

Merkouris:2010:CIM

[3104] Takis Merkouris. Combining information from multiple surveys by using regression for efficient small domain estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(1):27–48, January 2010. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/40541573.

Kai:2010:LCQ

[3105] Bo Kai, Runze Li, and Hui Zou. Local composite quantile regression smoothing: An efficient and safe alternative to local polynomial regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(1): 49–69, January 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40541574.

Ranjan:2010:CPF

[3106] Roopesh Ranjan and Tilmann Gneiting. Combining probability forecasts. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72 (1):71–91, January 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40541575.

Hall:2010:OSC

[3107] Peter Hall and You-Jun Yang. Ordering and selecting components in multivariate or functional data linear prediction. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(1):93–110, January 2010. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/40541576.

VanderWeele:2010:SDA

[3108] Tyler J. VanderWeele and James M. Robins. Signed directed acyclic graphs for causal inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(1):111–127, January 2010. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40541577.

Ju:2010:CSE

[3109] Chuan Ju and Zhi Geng. Criteria for surrogate end points based on causal distributions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(1):129– 142, January 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40541578.

Anonymous:2010:BMa

[3110] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72 (1):??, January 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40541579.

Anonymous:2010:FMa

[3111] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72 (1):??, January 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40541570.

Johnson:2010:UNL

[3112] Valen E. Johnson and David Rossell.

On the use of non-local prior densities in Bayesian hypothesis tests. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(2): 143–170, March 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40541581.

Park:2010:LAE

[3113] Juhyun Park and Burkhardt Seifert.

Local additive estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(2): 171–191, March 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40541582.

Copas:2010:LSE

[3114] John Copas and Shinto Eguchi. Likelihood for statistically equivalent models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(2):193–217, March 2010. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/40541583.

Strawderman:2010:SER

[3115] William E. Strawderman and Andrew L. Rukhin. Simultaneous estimation and reduction of nonconformity in interlaboratory studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(2): 219–234, March 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40541584.

Chen:2010:SMR

[3116] Yi-Hau Chen. Semiparametric marginal regression analysis for dependent competing risks under an assumed copula. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(2):235–251, March 2010. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/40541585.

Bush:2010:MIP

[3117] Christopher A. Bush, Juhee Lee, and Steven N. MacEachern. Minimally informative prior distributions for nonparametric Bayesian analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(2): 253–268, March 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40541586.

Anonymous:2010:BMb

[3118] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72 (2):??, March 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40541587.

Anonymous:2010:FMb

[3119] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72 (2):??, March 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40541580.

Andrieu:2010:PMC

[3120] Christophe Andrieu, Arnaud Doucet, and Roman Holenstein. Particle Markov chain Monte Carlo methods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72 (3):269–342, June 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40802151.

Shao:2010:SNA

[3121] Xiaofeng Shao. A self-normalized approach to confidence interval construction in time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(3):343–366, June 2010. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40802152. See corrigendum

[3140].

Guan:2010:SDR

[3122] Yongtao Guan and Hansheng Wang.

Sufficient dimension reduction for spatial point processes directed by Gaussian random fields. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(3):367–387, June 2010. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40802153.

Zhu:2010:SSL

[3123] Jun Zhu, Hsin-Cheng Huang, and Perla E. Reyes. On selection of spatial linear models for lattice data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(3): 389–402, June 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40802154.

Anonymous:2010:BMc

[3124] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72 (3):??, June 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40802155.

Anonymous:2010:FMc

[3125] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72 (3):??, June 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40802150.

Firth:2010:PRR

[3126] David Firth. Preface: ‘retrospective read paper’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(4):403, September 2010. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40802218.

Benjamini:2010:DFD

[3127] Yoav Benjamini. Discovering the false discovery rate. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(4):405–416, September 2010. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40802219.

Meinshausen:2010:SS

[3128] Nicolai Meinshausen and Peter Bühlmann.

Stability selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(4):417–473, September 2010. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40802220.

Ma:2010:EEE

[3129] Yanyuan Ma and Marc G. Genton. Explicit estimating equations for semiparametric generalized linear latent variable models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(4):475–495, September 2010. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/40802221.

Fearnhead:2010:RWP

[3130] Paul Fearnhead, Omiros Papaspiliopoulos, Gareth O. Roberts, and Andrew Stuart. Random-weight particle filtering of continuous time processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(4):497–512, September 2010. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/40802222.

Zhou:2010:SIL

[3131] Zhou Zhou and Wei Biao Wu. Simultaneous inference of linear models with time varying coefficients. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(4):513– 531, September 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40802223.

Rao:2010:BPE

[3132] J. N. K. Rao and Changbao Wu. Bayesian pseudo-empirical-likelihood intervals for complex surveys. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(4):533– 544, September 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40802224.

Anonymous:2010:BMd

[3133] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(4): ??, September 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40802225.

Anonymous:2010:FMd

[3134] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(4): ??, September 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40802217.

Cule:2010:MLE

[3135] Madeleine Cule, Richard Samworth, and Michael Stewart. Maximum likelihood estimation of a multidimensional log-concave density. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(5):545– 607, November 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40925417.

Ning:2010:NPT

[3136] Jing Ning, Jing Qin, and Yu Shen.

Non-parametric tests for right-censored data with biased sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(5):609– 630, November 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40925418.

Fraser:2010:DPB

[3137] D. A. S. Fraser, N. Reid, E. Marras, and G. Y. Yi. Default priors for Bayesian and frequentist inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(5):631–654, November 2010. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/40925419.

Kato:2010:MPC

[3138] Shogo Kato. A Markov process for circular data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(5):655–672, November 2010. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/40925420.

Claeskens:2010:MAT

[3139] Gerda Claeskens, Bernard W. Silverman, and Leen Slaets. A multiresolution approach to time warping achieved by a Bayesian prior–posterior transfer fitting strategy. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(5):673– 694, November 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/40925421.

Shao:2010:CSN

[3140] Xiaofeng Shao. Corrigendum: A selfnormalized approach to confidence interval construction in time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72 (5):695–696, November 2010. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http:

                                        //www.jstor.org/stable/40925422.

See [3121].

Anonymous:2010:IAV

[3141] Anonymous. Index of authors, volume 72, 2010. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(5):697–699, November 2010. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2010:BMe

[3142] Anonymous. Back matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(5):

??, November 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40925423.

Anonymous:2010:FMe

[3143] Anonymous. Front matter. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 72(5): ??, November 2010. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/40925416.

Casella:2011:RE

[3144] G. Casella and G. Roberts. Report of the Editors — 2010. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(1): 1–2, January 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41057422.

Wood:2011:FSR

[3145] Simon N. Wood. Fast stable restricted maximum likelihood and marginal likelihood estimation of semiparametric generalized linear models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(1): 3–36, January 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41057423.

Yau:2011:BNP

[3146] C. Yau, O. Papaspiliopoulos, G. O.

Roberts, and C. Holmes. Bayesian non-parametric hidden Markov models with applications in genomics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(1):

37–57, January 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41057424.

Li:2011:IPP

[3147] Lexin Li, Liping Zhu, and Lixing Zhu. Inference on the primary parameter of interest with the aid of dimension reduction estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(1): 59–80, January 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41057425.

Ma:2011:LOG

[3148] Yanyuan Ma, Jeffrey D. Hart, Ryan Janicki, and Raymond J. Carroll. Local and omnibus goodness-of-fit tests in classical measurement error models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73 (1):81–98, January 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41057426.

Shao:2011:MNH

[3149] Nan Shao and Keh-Shin Lii. Modelling non-homogeneous Poisson processes with almost periodic intensity functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(1):99–122, January 2011. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/41057427.

Girolami:2011:RML

[3150] Mark Girolami and Ben Calderhead. Riemann manifold Langevin and Hamiltonian Monte Carlo methods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(2):123–214, March 2011. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/41057430.

Li:2011:RDD

[3151] Pengfei Li and Douglas P. Wiens. Robustness of design in dose–response studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(2):215–238, March 2011. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/41057431.

Holland-Letz:2011:GCO

[3152] Tim Holland-Letz, Holger Dette, and Andrey Pepelyshev. A geometric characterization of optimal designs for regression models with correlated observations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(2):239–252, March 2011. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/41057432.

McCabe:2011:EPF

[3153] Brendan P. M. McCabe, Gael M. Martin, and David Harris. Efficient probabilistic forecasts for counts. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(2): 253–272, March 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41057433.

Tibshirani:2011:RSS

[3154] Robert Tibshirani. Regression shrinkage and selection via the lasso: a retrospective. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(3):273–282, June 2011. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/41262671.

Delaigle:2011:RAM

[3155] Aurore Delaigle, Peter Hall, and Jiashun Jin. Robustness and accuracy of methods for high dimensional data analysis based on Student’s t-statistic. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73 (3):283–301, June 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41262672.

Yang:2011:FSC

[3156] Wenjing Yang, Hans-Georg Müller, and Ulrich Stadtmüller. Functional singular component analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(3): 303–324, June 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41262673.

Bradic:2011:PCQ

[3157] Jelena Bradic, Jianqing Fan, and Weiwei Wang. Penalized composite quasi-likelihood for ultrahigh dimensional variable selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(3): 325–349, June 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41262674.

Scealy:2011:RCD

[3158] J. L. Scealy and A. H. Welsh. Regression for compositional data by using distributions defined on the hypersphere. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(3):351–375, June 2011. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/41262675.

Guillotte:2011:NPB

[3159] Simon Guillotte, François Perron, and Johan Segers. Non-parametric Bayesian inference on bivariate extremes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(3):377–406, June 2011. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/41262676.

Bernacchia:2011:SCM

[3160] Alberto Bernacchia and Simone Pigolotti. Self-consistent method for density estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(3):407–422, June 2011. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/41262677.

Lindgren:2011:ELB

[3161] Finn Lindgren, Håvard Rue, Johan Lindström, John T. Kent, Peter J. Diggle, J. B. Illian, D. P. Simpson, Tilmann Gneiting, Michael Scheuerer, R. Furrer, E. Furrer, D. Nychka, Paul Fearnhead, Peter Challenor, Yiannis Andrianakis, Gemma Stephenson, Jesper Møller, Xiangping Hu, Daniel Simpson, David Bolin, Patrick E. Brown, Michela Cameletti, Sara Martino, Daniel Cooley, Jennifer A. Hoeting, Rosa M. Crujeiras, Andrés Prieto, Marco A. R. Ferreira, Geir-Arne Fuglstad, Andrew Gelman, Peter Guttorp, Barnali Das, Ben Haaland, David J. Nott, John Haslett, Chaitanya Joshi, Vincent Garreta, Michael Höhle, L. Ippoliti, R. J. Martin, R. J. Bhansali, Venkata K. Jandhyala, Stergios B. Fotopoulos, Mikyoung Jun, Håvard Wahl Kongsgård, Giovanna Jona Lasinio, Alessio Pollice, Chihoon Lee, Wayne T. Lee, Cari G. Kaufman, Bo Li, Marc G. Genton, Georg Lindgren, K. V. Mardia, Jorge Mateu, Debashis Mondal, Werner G. Müller, Helmut Waldl, Alessandro Ottavi, Omiros Papaspiliopoulos, Emilio Porcu, Marc Saez, Alexandra M. Schmidt, Alfred Stein, Paul Switzer, Kamil Turkman, Christopher K. Wikle, and Mevin B. Hooten. An explicit link between Gaussian fields and Gaussian Markov random fields: the stochastic partial differential equation approach [with discussion]. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(4):423– 498, September 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41262260.

Fryzlewicz:2011:TPT

[3162] P. Fryzlewicz and H.-S. Oh. Thick pen transformation for time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(4):499– 529, September 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41262261.

Bunea:2011:AIM

[3163] Florentina Bunea, Andrada E. Ivanescu, and Marten H. Wegkamp. Adaptive inference for the mean of a Gaussian process in functional data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(4):531– 558, September 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41262262.

Li:2011:MAR

[3164] Yimei Li, Hongtu Zhu, Dinggang Shen, Weili Lin, John H. Gilmore, and Joseph G. Ibrahim. Multiscale adaptive regression models for neuroimaging data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(4):559–578, September 2011. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/41262263.

White:2011:BSP

[3165] John Thomas White and Subhashis Ghosal. Bayesian smoothing of photonlimited images with applications in astronomy. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(4):579–599, September 2011. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/41262264.

Comte:2011:DDD

[3166] F. Comte and C. Lacour. Data-driven density estimation in the presence of additive noise with unknown distribution. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(4):601–627, September 2011. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/41262265.

Cai:2011:ODH

[3167] T. Tony Cai, X. Jessie Jeng, and Jiashun Jin. Optimal detection of heterogeneous and heteroscedastic mixtures. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(5):629–662, November 2011. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/41262268.

Ferreira:2011:DMS

[3168] Marco A. R. Ferreira, Scott H. Holan, and Adelmo I. Bertolde. Dynamic multiscale spatiotemporal models for Gaussian areal data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(5):663– 688, November 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41262269.

Rousseau:2011:ABP

[3169] Judith Rousseau and Kerrie Mengersen. Asymptotic behaviour of the posterior distribution in overfitted mixture models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(5):689–710, November 2011. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/41262270.

Bickel:2011:BRA

[3170] Peter J. Bickel and Yulia R. Gel.

Banded regularization of autocovariance matrices in application to parameter estimation and forecasting of time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(5):711–728, November 2011. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/41262271.

Mattei:2011:ADA

[3171] Alessandra Mattei and Fabrizia Mealli.

Augmented designs to assess principal strata direct effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(5):729– 752, November 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41262272.

Witten:2011:PCU

[3172] Daniela M. Witten and Robert Tibshirani. Penalized classification using Fisher’s linear discriminant. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(5):753– 772, November 2011. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41262273.

Martinussen:2011:EDE

[3173] Torben Martinussen, Stijn Vansteelandt, Mette Gerster, and Jacob von Bornemann Hjelmborg. Estimation of direct effects for survival data by using the Aalen additive hazards model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(5):773–788, November 2011. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/41262274.

Anonymous:2011:IAV

[3174] Anonymous. Index of authors, volume 73, 2011. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(5):789–791, November 2011. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cappe:2011:CLEa

[3175] Olivier Cappé and Eric Moulines. Corrigendum: On-line expectation– maximization algorithm for latent data models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(5):792, November 2011. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3071].

Cappe:2011:CLEb

[3176] Olivier Cappé and Eric Moulines. Corrigendum: On-line expectation– maximization algorithm for latent data models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 73(5):792, November 2011. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/41262275. See [3071].

Casella:2012:RE

[3177] G. Casella and G. Roberts. Report of the Editors — 2011. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(1): 1–2, January 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41430926.

Ambroise:2012:NCA

[3178] Christophe Ambroise and Catherine Matias. New consistent and asymptotically normal parameter estimates for random-graph mixture models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(1): 3–35, January 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41430927.

Fan:2012:VEU

[3179] Jianqing Fan, Shaojun Guo, and Ning Hao. Variance estimation using refitted cross-validation in ultrahigh dimensional regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(1):37–65, January 2012. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/41430928.

Chen:2012:CQA

[3180] Kehui Chen and Hans-Georg Müller.

Conditional quantile analysis when covariates are functions, with application to growth data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(1):67– 89, January 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41430929.

Lee:2012:HCR

[3181] Stephen M. S. Lee. Hybrid confidence regions based on data depth. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(1): 91–109, January 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41430930.

Sang:2012:FSA

[3182] Huiyan Sang and Jianhua Z. Huang. A full scale approximation of covariance functions for large spatial data sets. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(1):111–132, January 2012. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/41430931.

Dellaportas:2012:CVE

[3183] Petros Dellaportas and Ioannis Kontoyiannis. Control variates for estimation based on reversible Markov chain Monte Carlo samplers. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(1): 133–161, January 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41430932.

Liang:2012:ADA

[3184] Kun Liang and Dan Nettleton. Adaptive and dynamic adaptive procedures for false discovery rate control and estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(1):163–182, January 2012. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/41430933.

Cheng:2012:CIA

[3185] Yu Cheng and Jason P. Fine. Cumulative incidence association models for bivariate competing risks data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(2): 183–202, March 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41430936.

Chen:2012:RRS

[3186] Kun Chen, Kung-Sik Chan, and Nils Chr. Stenseth. Reduced rank stochastic regression with a sparse singular value decomposition. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(2): 203–221, March 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41430937.

Vansteelandt:2012:STS

[3187] Stijn Vansteelandt, Tyler J. Vander-Weele, and James M. Robins. Semiparametric tests for sufficient cause interaction. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(2):223–244, March 2012. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/41430938.

Tibshirani:2012:SRD

[3188] Robert Tibshirani, Jacob Bien, Jerome Friedman, Trevor Hastie, Noah Simon, Jonathan Taylor, and Ryan J. Tibshirani. Strong rules for discarding predictors in lasso-type problems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(2):

245–266, March 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41430939.

Delaigle:2012:ANP

[3189] Aurore Delaigle and Peter Hall. Achieving near perfect classification for functional data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(2):267–286, March 2012. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/41430940.

Polson:2012:LSR

[3190] Nicholas G. Polson and James G. Scott. Local shrinkage rules, Lévy processes and regularized regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(2): 287–311, March 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41430941.

He:2012:DEC

[3191] Yunxiao He and Chuanhai Liu. The dynamic ‘expectation — conditional maximization either’ algorithm. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(2): 313–336, March 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41430942.

Neill:2012:FSS

[3192] Daniel B. Neill. Fast subset scan for spatial pattern detection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(2):

337–360, March 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41430943.

vanErven:2012:CFS

[3193] Tim van Erven, Peter Grünwald, and Steven de Rooij. Catching up faster by switching sooner: a predictive approach to adaptive estimation with an application to the AIC–BIC dilemma [with discussion]. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(3):361–417, June 2012. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/41674638.

Fearnhead:2012:CSS

[3194] Paul Fearnhead and Dennis Prangle. Constructing summary statistics for approximate Bayesian computation: semi-automatic approximate Bayesian computation [with discussion]. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(3): 419–474, June 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41674639.

Mougeot:2012:LL

[3195] Mathilde Mougeot, Dominique Picard, and Karine Tribouley. Learning out of leaders. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(3):475–513, June 2012. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/41674640.

Yekutieli:2012:ABI

[3196] Daniel Yekutieli. Adjusted Bayesian inference for selected parameters. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(3): 515–541, June 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41674641.

Wadsworth:2012:LBP

[3197] J. L. Wadsworth and J. A. Tawn. Likelihood-based procedures for threshold diagnostics and uncertainty in extreme value modelling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(3):543–567, June 2012. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/41674642.

Evans:2012:PKR

[3198] Steven N. Evans and Frederick A. Matsen. The phylogenetic Kantorovich– Rubinstein metric for environmental sequence samples. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(3):569–592, June 2012. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/41674643.

Cho:2012:HDV

[3199] Haeran Cho and Piotr Fryzlewicz. High dimensional variable selection via tilting. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(3):593–622, June 2012. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/41674644.

Thas:2012:PIM

[3200] Olivier Thas, Jan De Neve, Lieven Clement, and Jean-Pierre Ottoy. Probabilistic index models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(4):623– 671, September 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41674606.

Farrington:2012:RFV

[3201] C. Paddy Farrington, Steffen Unkel, and Karim Anaya-Izquierdo. The relative frailty variance and shared frailty models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(4):673–696, September 2012. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/41674607.

Yuan:2012:LPR

[3202] Ying Yuan, Hongtu Zhu, Weili Lin, and J. S. Marron. Local polynomial regression for symmetric positive definite matrices. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(4):697– 719, September 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41674608.

Allen:2012:ITD

[3203] Genevera I. Allen and Robert Tibshirani. Inference with transposable data: modelling the effects of row and column correlations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(4):721– 743, September 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41674609.

Fan:2012:RCH

[3204] Jianqing Fan, Yang Feng, and Xin Tong. A road to classification in high dimensional space: the regularized optimal affine discriminant. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(4):745– 771, September 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/41674610.

Cai:2012:RDI

[3205] T. Tony Cai, X. Jessie Jeng, and Hongzhe Li. Robust detection and identification of sparse segments in ultrahigh dimensional data analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(5):773–797, November 2012. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/41674631.

Bai:2012:JCE

[3206] Yun Bai, Peter X.-K. Song, and T. E. Raghunathan. Joint composite estimating functions in spatiotemporal models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(5):799–824, November 2012. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/41674632.

Davidov:2012:ODC

[3207] Ori Davidov and Amir Herman. Ordinal dominance curve based inference for stochastically ordered distributions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(5):825–847, November 2012. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/41674633.

Zhong:2012:CPF

[3208] Wenxuan Zhong, Tingting Zhang, Yu Zhu, and Jun S. Liu. Correlation pursuit: forward stepwise variable selection for index models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(5):849– 870, November 2012. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/41674634.

Zhao:2012:EBF

[3209] Zhigen Zhao and J. T. Gene Hwang.

Empirical Bayes false coverage rate controlling confidence intervals. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74 (5):871–891, November 2012. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/41674635.

Anonymous:2012:IAV

[3210] Anonymous. Index of authors, volume 74, 2012. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(5):893–895, November 2012. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2012:CV

[3211] Anonymous. Contents of volume 74, 2012. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 74(5):896–897, November 2012. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Roberts:2013:ERE

[3212] G. O. Roberts and I. Van Keilegom. Editorial: Report of the Editors — 2012. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(1):1–2, January 2013. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/23361012.

Hampson:2013:GST

[3213] Lisa V. Hampson and Christopher Jennison. Group sequential tests for delayed responses (with discussion). Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75 (1):3–39, January 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/23361013. See dis-

cussion [3214].

Julious:2013:DPH

[3214] Steven A. Julious, Munyaradzi Dimairo, Nigel Stallard, Ekkehard Glimm, Frank Bretz, Paul Gallo, Giorgos Minas, John A. D. Aston, Marc Vandemeulebroecke, Martin Posch, Steven Gilmour, Peter Bauer, Bertrand Clarke, Jay Bartroff, Carl-Fredrik Burman, D. S. Coad, David Draper, Jorge Mateu, Francisco J. Rodríguez-Cortés, Kit C. B. Roes, Weng Kee Wong, Oleksandr Sverdlov, Roger Lewis, Alessandro Zini, Emilio Porcu, and Roberto D’Ercole. Discussion on the paper by Hampson and Jennison. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(1):40–54, January 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/23361274. See [3213].

Shah:2013:VSE

[3215] Rajen D. Shah, Richard J. Samworth, and R. J. Samsworth. Variable selection with error control: another look at stability selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(1): 55–80, January 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/23361014.

Chen:2013:MWT

[3216] Song Xi Chen, Jing Qin, and Cheng Yong Tang. Mann–Whitney test with adjustments to pretreatment variables for missing values and observational study. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(1):81–102, January 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/23361015.

Horvath:2013:EMF

[3217] Lajos Horváth, Piotr Kokoszka, and Ron Reeder. Estimation of the mean of functional time series and a twosample problem. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(1):103–122, January 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.org/
stable/23361016.

Yao:2013:NLE

[3218] Weixin Yao and Runze Li. New local estimation procedure for a nonparametric regression function for longitudinal data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(1):123–138, January 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/23361017.

Hughes:2013:DRA

[3219] John Hughes and Murali Haran. Dimension reduction and alleviation of confounding for spatial generalized linear mixed models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(1):139–159, January 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/23361018.

Shao:2013:FBS

[3220] Xiaofeng Shao and Dimitris N. Politis. Fixed b subsampling and the block bootstrap: improved confidence sets based on p-value calibration. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(1): 161–184, January 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

                                        jstor.org/stable/23361019.

Wei:2013:REH

[3221] Jiawei Wei, Raymond J. Carroll, Ursula U. Müller, Ingrid Van Keilegom, and Nilanjan Chatterjee. Robust estimation for homoscedastic regression in the secondary analysis of case — control data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(1):185– 206, January 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/23361020.

Cox:2013:ROP

[3222] D. R. Cox. A return to an old paper: ‘Tests of separate families of hypotheses’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(2):207–211, March 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL http://www.jstor.
org/stable/23360923. See comments

[3223].

Kent:2013:CP

[3223] J. T. Kent and Grayham E. Mizon.

Comments on the presentation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(2): 212–215, March 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/23361330.  See

[3222].

Gorst-Rasmussen:2013:ISS

[3224] Anders Gorst-Rasmussen and Thomas Scheike. Independent screening for single-index hazard rate models with ultrahigh dimensional features. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(2): 217–245, March 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/23360924.

Coeurjolly:2013:RGF

[3225] Jean-François Coeurjolly and Frédéric Lavancier. Residuals and goodness-of-fit tests for stationary marked Gibbs point processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(2):247–276, March 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/23360925.

Kenah:2013:NPS

[3226] Eben Kenah. Non-parametric survival analysis of infectious disease data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(2):277–303, March 2013. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        www.jstor.org/stable/23360926.

Ma:2013:DRE

[3227] Yanyuan Ma and Liping Zhu. Doubly robust and efficient estimators for heteroscedastic partially linear singleindex models allowing high dimensional covariates. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(2):305–322, March 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/23360927.

Zhou:2013:ILM

[3228] Zhou Zhou and Xiaofeng Shao. Inference for linear models with dependent errors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(2):323–343, March 2013. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/23360928.

Francq:2013:OPP

[3229] Christian Francq and Jean-Michel Zakoïan. Optimal predictions of powers of conditionally heteroscedastic processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(2):345–367, March 2013. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

www.jstor.org/stable/23360929.

Miratrix:2013:ATE

[3230] Luke W. Miratrix, Jasjeet S. Sekhon, and Bin Yu. Adjusting treatment effect estimates by post-stratification in randomized experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(2): 369–396, March 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://www.

jstor.org/stable/23360930.

Chopin:2013:SEA

[3231] N. Chopin, P. E. Jacob, and O. Papaspiliopoulos. SMC^{2}: an efficient algorithm for sequential analysis of state space models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(3):397–426, June 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Won:2013:CNR

[3232] Joong-Ho Won, Johan Lim, Seung-Jean Kim, and Bala Rajaratnam. Condition-number-regularized covariance estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(3):427–450, June 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rukhin:2013:EHV

[3233] Andrew L. Rukhin. Estimating heterogeneity variance in meta-analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75 (3):451–469, June 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Augugliaro:2013:DGL

[3234] Luigi Augugliaro, Angelo M. Mineo, and Ernst C. Wit. Differential geometric least angle regression: a differential geometric approach to sparse generalized linear models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(3):471–498, June 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Griffin:2013:CDU

[3235] J. E. Griffin, M. Kolossiatis, and M. F. J. Steel. Comparing distributions by using dependent normalized random-measure mixtures. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(3): 499–529, June 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Fan:2013:TPS

[3236] Yingying Fan and Cheng Yong Tang. Tuning parameter selection in high dimensional penalized likelihood. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(3): 531–552, June 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lee:2013:ELA

[3237] Youngjo Lee and Jan F. Bjørnstad. Extended likelihood approach to largescale multiple testing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(3):553–575, June 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Xiao:2013:FBS

[3238] Luo Xiao, Yingxing Li, and David Ruppert. Fast bivariate P -splines: the sandwich smoother. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(3):577–599, June 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Taylor:2013:CSP

[3239] Benjamin M. Taylor and Peter J. Diggle. Corrigendum: Spatiotemporal prediction for log-Gaussian Cox processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75 (3):601–602, June 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2614].

Fan:2013:LCE

[3240] Jianqing Fan, Yuan Liao, and Martina Mincheva. Large covariance estimation by thresholding principal orthogonal complements. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(4):603– 680, September 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Sangalli:2013:SSR

[3241] Laura M. Sangalli, James O. Ramsay, and Timothy O. Ramsay. Spatial spline regression models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(4):681– 703, September 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ramsahai:2013:PCD

[3242] Roland R. Ramsahai. Probabilistic causality and detecting collections of interdependence patterns. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(4):705– 723, September 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Krivobokova:2013:SPS

[3243] Tatyana Krivobokova. Smoothing parameter selection in two frameworks for penalized splines. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(4):725– 741, September 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Evans:2013:MLL

[3244] Robin J. Evans and Thomas S. Richardson. Marginal log-linear parameters for graphical Markov models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(4):743–768, September 2013. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Balabdaoui:2013:ADL

[3245] Fadoua Balabdaoui, Hanna Jankowski, Kaspar Rufibach, and Marios Pavlides.

Asymptotics of the discrete log-concave maximum likelihood estimator and related applications. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(4):769–790, September 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Davis:2013:SIM

[3246] Richard A. Davis, Claudia Klüppelberg, and Christina Steinkohl. Statistical inference for max-stable processes in space and time. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(5):791–819, November 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Perry:2013:PPM

[3247] Patrick O. Perry and Patrick J. Wolfe.

Point process modelling for directed interaction networks. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(5):821– 849, November 2013. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cook:2013:EPL

[3248] R. D. Cook, I. S. Helland, and Z. Su.

Envelopes and partial least squares regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(5):851–877, November 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Nason:2013:TSO

[3249] Guy Nason. A test for second-order stationarity and approximate confidence intervals for localized autocovariances for locally stationary time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75 (5):879–904, November 2013. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2013:IAI

[3250] Anonymous. Index of authors: Index of authors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(5):905–907, November 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2013:CV

[3251] Anonymous. Contents of volume 75, 2013. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 75(5):909–910, November 2013. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Roberts:2014:ERE

[3252] G. O. Roberts and I. Van Keilegom. Editorial: Report of the Editors — 2013. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1):1–2, January 2014. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Hothorn:2014:CTM

[3253] Torsten Hothorn, Thomas Kneib, and Peter Bühlmann. Conditional transformation models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1):3–27, January 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Krivitsky:2014:SMD

[3254] Pavel N. Krivitsky and Mark S. Handcock. A separable model for dynamic networks. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1):29–46, January 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chambers:2014:ORS

[3255] Ray Chambers, Hukum Chandra, Nicola Salvati, and Nikos Tzavidis. Outlier robust small area estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76 (1):47–69, January 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lei:2014:DFP

[3256] Jing Lei and Larry Wasserman.

Distribution-free prediction bands for non-parametric regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1): 71–96, January 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kasahara:2014:NPI

[3257] Hiroyuki Kasahara and Katsumi Shimotsu. Non-parametric identification and estimation of the number of components in multivariate mixtures. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1): 97–111, January 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Goga:2014:EEN

[3258] Camelia Goga and Anne Ruiz-Gazen.

Efficient estimation of non-linear finite population parameters by using nonparametrics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1):113–140, January 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lv:2014:MSP

[3259] Jinchi Lv and Jun S. Liu. Model selection principles in misspecified models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1):141–167, January 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kosmidis:2014:IEC

[3260] Ioannis Kosmidis. Improved estimation in cumulative link models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1): 169–196, January 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhou:2014:SID

[3261] Haibo Zhou, Wangli Xu, Donglin Zeng, and Jianwen Cai. Semiparametric inference for data with a continuous outcome from a two-phase probabilitydependent sampling scheme. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1):197– 215, January 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhang:2014:CIL

[3262] Cun-Hui Zhang and Stephanie S. Zhang. Confidence intervals for low dimensional parameters in high dimensional linear models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1):217– 242, January 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Imai:2014:CBP

[3263] Kosuke Imai and Marc Ratkovic. Covariate balancing propensity score. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1):243–263, January 2014. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dette:2014:ODD

[3264] Holger Dette, Christine Kiss, Norbert Benda, and Frank Bretz. Optimal designs for dose finding studies with an active control. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1):265–295, January 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Benjamini:2014:SIM

[3265] Yoav Benjamini and Marina Bogomolov. Selective inference on multiple families of hypotheses. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(1):297– 318, January 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Deng:2014:APD

[3266] Ke Deng, Zhi Geng, and Jun S. Liu. Association pattern discovery via theme dictionary models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(2):319–347, March 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cai:2014:TST

[3267] T. Tony Cai, Weidong Liu, and Yin Xia. Two-sample test of high dimensional means under dependence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(2): 349–372, March 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Danaher:2014:JGL

[3268] Patrick Danaher, Pei Wang, and Daniela M. Witten. The joint graphical lasso for inverse covariance estimation across multiple classes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(2): 373–397, March 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Veraverbeke:2014:PNP

[3269] Noël Veraverbeke, Irène Gijbels, and Marek Omelka. Preadjusted nonparametric estimation of a conditional distribution function. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(2):399–438, March 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Huser:2014:STM

[3270] R. Huser and A. C. Davison. Space– time modelling of extreme events. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(2):439–461, March 2014. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhou:2014:RMR

[3271] Hua Zhou and Lexin Li. Regularized matrix regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(2):463–483, March 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Spiegelhalter:2014:DIC

[3272] David J. Spiegelhalter, Nicola G.

Best, Bradley P. Carlin, and Angelika van der Linde. The deviance information criterion: 12 years on. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(3): 485–493, June 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Frick:2014:MCP

[3273] Klaus Frick, Axel Munk, and Hannes Sieling. Multiscale change point inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(3):495–580, June 2014. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhu:2014:SFA

[3274] Hongxiao Zhu, Fang Yao, and Hao Helen Zhang. Structured functional additive regression in reproducing kernel Hilbert spaces. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(3):581–603, June 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chiou:2014:LMM

[3275] Jeng-Min Chiou and Hans-Georg Müller. Linear manifold modelling of multivariate functional data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(3): 605–626, June 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zheng:2014:HDT

[3276] Zemin Zheng, Yingying Fan, and Jinchi Lv. High dimensional thresholded regression and shrinkage effect. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(3): 627–649, June 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chen:2014:NPA

[3277] Yu-Chun Chen, Ming-Yen Cheng, and Hau-Tieng Wu. Non-parametric and adaptive modelling of dynamic periodicity and trend with heteroscedastic and dependent errors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(3): 651–682, June 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Doss:2014:ESE

[3278] Hani Doss and Aixin Tan. Estimates and standard errors for ratios of normalizing constants from multiple Markov chains via regeneration. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(4):683–712, September 2014. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Polson:2014:BB

[3279] Nicholas G. Polson, James G. Scott, and Jesse Windle. The Bayesian bridge. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(4):713–733, September 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Hall:2014:QEO

[3280] Peter Hall and Yanyuan Ma. Quick and easy one-step parameter estimation in differential equations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(4):735– 748, September 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rolling:2014:MSE

[3281] Craig A. Rolling and Yuhong Yang.

Model selection for estimating treatment effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(4):749–769, September 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Aharoni:2014:GID

[3282] Ehud Aharoni and Saharon Rosset.

Generalized α-investing: definitions, optimality results and application to public databases. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(4):771–794, September 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kleiner:2014:SBM

[3283] Ariel Kleiner, Ameet Talwalkar, Purnamrita Sarkar, and Michael I. Jordan. A scalable bootstrap for massive data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(4):795–816, September 2014. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chu:2014:SIG

[3284] Tingjin Chu, Haonan Wang, and Jun Zhu. On semiparametric inference of geostatistical models via local Karhunen–Loève expansion. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(4):817– 832, September 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Marin:2014:RSB

[3285] Jean-Michel Marin, Natesh S. Pillai, Christian P. Robert, and Judith Rousseau. Relevant statistics for Bayesian model choice. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(5):833– 859, November 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Naveau:2014:NPE

[3286] Philippe Naveau, Armelle Guillou, and Théo Rietsch. A non-parametric entropy-based approach to detect changes in climate extremes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(5):861– 884, November 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ma:2014:EEC

[3287] Yanyuan Ma and Liping Zhu. On estimation efficiency of the central mean subspace. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(5):885–901, November 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Fryzlewicz:2014:MCP

[3288] P. Fryzlewicz and S. Subba Rao.

Multiple-change-point detection for auto-regressive conditional heteroscedastic processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(5):903–924, November 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Borgonovo:2014:TIS

[3289] E. Borgonovo, S. Tarantola, E. Plischke, and M. D. Morris. Transformations and invariance in the sensitivity analysis of computer experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(5):925–947, November 2014. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Decrouez:2014:SSM

[3290] Geoffrey Decrouez and Peter Hall.

Split sample methods for constructing confidence intervals for binomial and Poisson parameters. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(5):949– 975, November 2014. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2014:IAV

[3291] Anonymous. Index of authors, volume 76, 2014. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(5):977–979, November 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2014:CV

[3292] Anonymous. Contents of volume 76, 2014. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 76(5):981–982, November 2014. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2015:RER

[3293] Anonymous. Report of the Editors — 2014: Report of the Editors — 2014. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1):1, January 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Sigrist:2015:SPD

[3294] Fabio Sigrist, Hans R. Künsch, and Werner A. Stahel. Stochastic partial differential equation based modelling of large space–time data sets. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1): 3–33, January 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kessler:2015:MSP

[3295] David C. Kessler, Peter D. Hoff, and David B. Dunson. Marginally specified priors for non-parametric Bayesian estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1):35–58, January 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Sun:2015:FDC

[3296] Wenguang Sun, Brian J. Reich, T. Tony Cai, Michele Guindani, and Armin Schwartzman. False discovery control in large-scale spatial multiple testing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1):59–83, January 2015. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bolin:2015:ECU

[3297] David Bolin and Finn Lindgren. Excursion and contour uncertainty regions for latent Gaussian models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1):85– 106, January 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Li:2015:QRA

[3298] Ruosha Li and Limin Peng. Quantile regression adjusting for dependent censoring from semicompeting risks. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1): 107–130, January 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Barrett:2015:JMR

[3299] Jessica Barrett, Peter Diggle, Robin Henderson, and David Taylor-Robinson. Joint modelling of repeated measurements and time-to-event outcomes: flexible model specification and exact likelihood inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1):131– 148, January 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Delaigle:2015:CBN

[3300] Aurore Delaigle, Peter Hall, and Farshid Jamshidi. Confidence bands in non-parametric errors-in-variables regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1):149–169, January 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lian:2015:VFP

[3301] Heng Lian, Hua Liang, and Raymond J. Carroll. Variance function partially linear single-index models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1): 171–194, January 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Martin:2015:CIM

[3302] Ryan Martin and Chuanhai Liu. Conditional inferential models: combining information for prior-free probabilistic inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1):195–217, January 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhang:2015:JMA

[3303] Weiping Zhang, Chenlei Leng, and Cheng Yong Tang. A joint modelling approach for longitudinal studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1):219–238, January 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Engelke:2015:EHR

[3304] Sebastian Engelke, Alexander Malinowski, Zakhar Kabluchko, and Martin Schlather. Estimation of Hüsler–Reiss distributions and Brown–Resnick processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1):239–265, January 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kreiss:2015:BLS

[3305] Jens-Peter Kreiss and Efstathios Paparoditis. Bootstrapping locally stationary processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1):267–290, January 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Hauser:2015:JIO

[3306] Alain Hauser and Peter Bühlmann.

Jointly interventional and observational data: estimation of interventional Markov equivalence classes of directed acyclic graphs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(1):291– 318, January 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Hormann:2015:DFP

[3307] Siegfried Hörmann, Lukasz Kidziński, and Marc Hallin. Dynamic functional principal components. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(2): 319–348, March 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhou:2015:INS

[3308] Zhou Zhou. Inference for nonstationary time series regression with or without inequality constraints. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(2): 349–371, March 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ogburn:2015:DRE

[3309] Elizabeth L. Ogburn, Andrea Rotnitzky, and James M. Robins. Doubly robust estimation of the local average treatment effect curve. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(2): 373–396, March 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Yu:2015:STM

[3310] Wen Yu, Kani Chen, Michael E. Sobel, and Zhiliang Ying. Semiparametric transformation models for causal inference in time-to-event studies with all-or-nothing compliance. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(2): 397–415, March 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cai:2015:EME

[3311] Juan-Juan Cai, John H. J. Einmahl, Laurens de Haan, and Chen Zhou. Estimation of the marginal expected shortfall: the mean when a related variable is extreme. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(2):417–442, March 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Abramovich:2015:SAR

[3312] Felix Abramovich and Tal Lahav.

Sparse additive regression on a regular lattice. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(2):443–459, March 2015. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Pauly:2015:APT

[3313] Markus Pauly, Edgar Brunner, and Frank Konietschke. Asymptotic permutation tests in general factorial designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(2):461–473, March 2015. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cho:2015:MCP

[3314] Haeran Cho and Piotr Fryzlewicz.

Multiple-change-point detection for high dimensional time series via sparsified binary segmentation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(2): 475–507, March 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Gerber:2015:SQM

[3315] Mathieu Gerber and Nicolas Chopin.

Sequential quasi Monte Carlo. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(3): 509–579, June 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zheng:2015:CMA

[3316] Cheng Zheng and Xiao-Hua Zhou.

Causal mediation analysis in the multilevel intervention and multicomponent mediator case. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(3):581–615, June 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Efron:2015:FAB

[3317] Bradley Efron. Frequentist accuracy of Bayesian estimates. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(3): 617–646, June 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Schweinberger:2015:LDR

[3318] Michael Schweinberger and Mark S. Handcock. Local dependence in random graph models: characterization, properties and statistical inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77 (3):647–676, June 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Guan:2015:QLS

[3319] Yongtao Guan, Abdollah Jalilian, and Rasmus Waagepetersen. Quasilikelihood for spatial point processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77 (3):677–697, June 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dutta:2015:HLM

[3320] Somak Dutta and Debashis Mondal. An h-likelihood method for spatial mixed linear models based on intrinsic auto-regressions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(3):699–726, June 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dasgupta:2015:CIF

[3321] Tirthankar Dasgupta, Natesh S. Pillai, and Donald B. Rubin. Causal inference from 2^{K} factorial designs by using potential outcomes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(4):727– 753, September 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cao:2015:RAS

[3322] Hongyuan Cao, Donglin Zeng, and Jason P. Fine. Regression analysis of sparse asynchronous longitudinal data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(4):755–776, September 2015. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kraus:2015:CCP

[3323] David Kraus. Components and completion of partially observed functional data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(4):777–801, September 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Khare:2015:CPF

[3324] Kshitij Khare, Sang-Yun Oh, and Bala Rajaratnam. A convex pseudolikelihood framework for high dimensional partial correlation estimation with convergence guarantees. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(4):803– 825, September 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Hao:2015:SFL

[3325] Ning Hao, Bin Dong, and Jianqing Fan. Sparsifying the Fisher linear discriminant by rotation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(4):827– 851, September 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lavancier:2015:DPP

[3326] Frédéric Lavancier, Jesper Møller, and Ege Rubak. Determinantal point process models and statistical inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(4):853–877, September 2015. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Yin:2015:SSD

[3327] Xiangrong Yin and Haileab Hilafu. Sequential sufficient dimension reduction for large p, small n problems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(4):879– 892, September 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Jin:2015:NNP

[3328] Lei Jin, Suojin Wang, and Haiyan Wang. A new non-parametric stationarity test of time series in the time domain. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(5):893–922, November 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Meinshausen:2015:GBC

[3329] Nicolai Meinshausen. Group bound:

confidence intervals for groups of variables in sparse high dimensional regression without assumptions on the design. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(5):923–945, November 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Song:2015:SMB

[3330] Qifan Song and Faming Liang. A split-and-merge Bayesian variable selection approach for ultrahigh dimensional regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(5):947–972, November 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Papageorgiou:2015:BNP

[3331] Georgios Papageorgiou, Sylvia Richardson, and Nicky Best. Bayesian nonparametric models for spatially indexed data of mixed type. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(5):973– 999, November 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Gong:2015:EEQ

[3332] Jinguo Gong, Yadong Li, Liang Peng, and Qiwei Yao. Estimation of extreme quantiles for functions of dependent random variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(5):1001– 1024, November 2015. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2015:IAV

[3333] Anonymous. Index of authors, volume 77, 2015. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(5):1025–1027, November 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2015:CV

[3334] Anonymous. Contents of volume 77, 2015. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 77(5):1029–1030, November 2015. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2016:RER

[3335] Anonymous. Report of the editors-2015. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1):1–2, January 2016. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Daouia:2016:DEF

[3336] Abdelaati Daouia, Hohsuk Noh, and Byeong U. Park. Data envelope fitting with constrained polynomial splines. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78 (1):3–30, January 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Einmahl:2016:SHE

[3337] John H. J. Einmahl, Laurens de Haan, and Chen Zhou. Statistics of heteroscedastic extremes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1):31– 51, January 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhang:2016:VSS

[3338] Xiang Zhang, Yichao Wu, Lan Wang, and Runze Li. Variable selection for support vector machines in moderately high dimensions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1):53–76, January 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Genton:2016:TAR

[3339] Marc G. Genton and Peter Hall. A tilting approach to ranking influence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78 (1):77–97, January 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Genovese:2016:NPI

[3340] Christopher R. Genovese, Marco Perone-Pacifico, Isabella Verdinelli, and Larry Wasserman. Non-parametric inference for density modes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1): 99–126, January 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ma:2016:SES

[3341] Yanyuan Ma and Raymond J. Carroll.

Semiparametric estimation in the secondary analysis of case-control studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1):127–151, January 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rajaratnam:2016:LRE

[3342] Bala Rajaratnam, Steven Roberts, Doug Sparks, and Onkar Dalal. Lasso regression: estimation and shrinkage via the limit of Gibbs sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1): 153–174, January 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Prus:2016:ODP

[3343] Maryna Prus and Rainer Schwabe. Optimal designs for the prediction of individual parameters in hierarchical models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1):175–191, January 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lee:2016:LHD

[3344] Sokbae Lee, Myung Hwan Seo, and Youngki Shin. The lasso for high dimensional regression with a possible change point. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1):193–210, January 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bonhomme:2016:NPE

[3345] Stéphane Bonhomme, Koen Jochmans, and Jean-Marc Robin. Non-parametric estimation of finite mixtures from repeated measurements. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1):211– 229, January 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Delaigle:2016:MNP

[3346] Aurore Delaigle and Peter Hall.

Methodology for non-parametric deconvolution when the error distribution is unknown. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1):231–252, January 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bickel:2016:HTA

[3347] Peter J. Bickel and Purnamrita Sarkar.

Hypothesis testing for automated community detection in networks. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1): 253–273, January 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Einmahl:2016:EST

[3348] John H. J. Einmahl, Anna Kiriliouk, Andrea Krajina, and Johan Segers. An M-estimator of spatial tail dependence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1):275–298, January 2016. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Yang:2016:UPO

[3349] Fan Yang and Dylan S. Small. Using post-outcome measurement information in censoring-by-death problems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(1):299–318, January 2016. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Berger:2016:ELC

[3350] Y. G. Berger and O. De La Riva Torres. Empirical likelihood confidence intervals for complex sampling designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(2):319–341, March 2016. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bladt:2016:SMD

[3351] Mogens Bladt, Samuel Finch, and Michael Sørensen. Simulation of multivariate diffusion bridges. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(2): 343–369, March 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See corrigendum [3757].

Dette:2016:DRC

[3352] Holger Dette and Dominik Wied. Detecting relevant changes in time series models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(2):371–394, March 2016. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhang:2016:CBP

[3353] Xianyang Zhang and Xiaofeng Shao. On the coverage bound problem of empirical likelihood methods for time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(2):395–421, March 2016. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

GSell:2016:SSP

[3354] Max Grazier G’Sell, Stefan Wager, Alexandra Chouldechova, and Robert Tibshirani. Sequential selection procedures and false discovery rate control. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(2):423–444, March 2016. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Doosti:2016:MNP

[3355] Hassan Doosti and Peter Hall. Making a non-parametric density estimator more attractive, and more accurate, by data perturbation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(2): 445–462, March 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhu:2016:BPT

[3356] Ke Zhu. Bootstrapping the portmanteau tests in weak auto-regressive moving average models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(2):463– 485, March 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Qiu:2016:JEM

[3357] Huitong Qiu, Fang Han, Han Liu, and Brian Caffo. Joint estimation of multiple graphical models from high dimensional time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(2):487–504, March 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ehm:2016:QEC

[3358] Werner Ehm, Tilmann Gneiting, Alexander Jordan, and Fabian Krüger. Of quantiles and expectiles: consistent scoring functions, Choquet representations and forecast rankings. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(3): 505–562, June 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Hartmann:2016:DES

[3359] Alexander Hartmann, Stephan Huckemann, Jörn Dannemann, Oskar Laitenberger, Claudia Geisler, Alexander Egner, and Axel Munk. Drift estimation in sparse sequential dynamic imaging, with application to nanoscale fluorescence microscopy. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(3):563–587, June 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2016:HDO

[3360] Xiangyu Wang and Chenlei Leng. High dimensional ordinary least squares projection for screening variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(3): 589–611, June 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See corrigendum [3654].

Francq:2016:EMV

[3361] Christian Francq and Jean-Michel Zakoïan. Estimating multivariate volatility models equation by equation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78 (3):613–635, June 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Hall:2016:TLM

[3362] Peter Hall and Giles Hooker. Truncated linear models for functional data. Journal of the Royal Statistical Society.

Series B (Statistical Methodology), 78 (3):637–653, June 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ding:2016:RIT

[3363] Peng Ding, Avi Feller, and Luke Miratrix. Randomization inference for treatment effect variation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(3): 655–671, June 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chan:2016:GEN

[3364] Kwun Chuen Gary Chan, Sheung Chi Phillip Yam, and Zheng Zhang. Globally efficient non-parametric inference of average treatment effects by empirical balancing calibration weighting. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(3):673–700, June 2016. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Polson:2016:MEH

[3365] Nicholas G. Polson and James G. Scott.

Mixtures, envelopes and hierarchical duality. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(4):701–727, September 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chen:2016:GAI

[3366] Yining Chen and Richard J. Samworth.

Generalized additive and index models with shape constraints. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(4):729– 754, September 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Gamiz:2016:DOS

[3367] María Luz Gámiz, Enno Mammen, María Dolores Martínez Miranda, and Jens Perch Nielsen. Double one-sided cross-validation of local linear hazards. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(4):755–779, September 2016. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Henderson:2016:MCI

[3368] Nicholas C. Henderson and Michael A. Newton. Making the cut: improved ranking and selection for large-scale inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(4):781–804, September 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhelonkin:2016:RIS

[3369] Mikhail Zhelonkin, Marc G. Genton, and Elvezio Ronchetti. Robust inference in sample selection models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78 (4):805–827, September 2016. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Jiang:2016:PCE

[3370] Zhichao Jiang, Peng Ding, and Zhi Geng. Principal causal effect identification and surrogate end point evaluation by multiple trials. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(4):829– 848, September 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhang:2016:DSE

[3371] Luwan Zhang, Grace Wahba, and Ming Yuan. Distance shrinkage and Euclidean embedding via regularized kernel estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(4):849–867, September 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Patra:2016:ETC

[3372] Rohit Kumar Patra and Bodhisattva Sen. Estimation of a two-component mixture model with applications to multiple testing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(4):869–893, September 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Yau:2016:IMC

[3373] Chun Yip Yau and Zifeng Zhao. Inference for multiple change points in time series via likelihood ratio scan statistics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(4):895–916, September 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

He:2016:EGU

[3374] Zhijian He and Art B. Owen. Extensible grids: uniform sampling on a space filling curve. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(4):917– 931, September 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bacallado:2016:BRL

[3375] Sergio Bacallado, Vijay Pande, Stefano Favaro, and Lorenzo Trippa. Bayesian regularization of the length of memory in reversible sequences. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(4):933– 946, September 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Peters:2016:CIU

[3376] Jonas Peters, Peter Bühlmann, and Nicolai Meinshausen. Causal inference by using invariant prediction: identification and confidence intervals. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78 (5):947–1012, November 2016. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Guo:2016:MCP

[3377] Xu Guo, Tao Wang, and Lixing Zhu. Model checking for parametric singleindex models: a dimension reduction model-adaptive approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(5):1013– 1035, November 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lee:2016:VSA

[3378] Kuang-Yao Lee, Bing Li, and Hongyu Zhao. Variable selection via additive conditional independence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(5):1037– 1055, November 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cholaquidis:2016:SER

[3379] Alejandro Cholaquidis, Ricardo Fraiman, Gábor Lugosi, and Beatriz Pateiro-López. Set estimation from reflected Brownian motion. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(5):1057–1078, November 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Guo:2016:THD

[3380] Bin Guo and Song Xi Chen. Tests for high dimensional generalized linear models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(5):1079–1102, November 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bissiri:2016:GFU

[3381] P. G. Bissiri, C. C. Holmes, and S. G.

Walker. A general framework for updating belief distributions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(5):1103– 1130, November 2016. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2016:IAV

[3382] Anonymous. Index of authors, volume 78, 2016. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(5):1131–1134, November 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2016:CV

[3383] Anonymous. Contents of volume 78, 2016. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(5):1135–1136, November 2016. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2017:IIa

[3384] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1): 1–2, January 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dunson:2017:RE

[3385] David Dunson and Piotr Fryzlewicz. Report of the Editors — 2016. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1): 3–4, January 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Vogt:2017:CNP

[3386] Michael Vogt and Oliver Linton. Classification of non-parametric regression functions in longitudinal data models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79 (1):5–27, January 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Gromenko:2017:DCS

[3387] Oleksandr Gromenko, Piotr Kokoszka, and Matthew Reimherr. Detection of change in the spatiotemporal mean function. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1):29–50, January 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Passemier:2017:ENV

[3388] Damien Passemier, Zhaoyuan Li, and Jianfeng Yao. On estimation of the noise variance in high dimensional probabilistic principal component analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1):51–67, January 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Comte:2017:LDB

[3389] Fabienne Comte, Charles-A. Cuenod, Marianna Pensky, and Yves Rozenholc. Laplace deconvolution on the basis of time domain data and its application to dynamic contrast-enhanced imaging. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79 (1):69–94, January 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bandyopadhyay:2017:TSI

[3390] Soutir Bandyopadhyay and Suhasini Subba Rao. A test for stationarity for irregularly spaced spatial data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1): 95–123, January 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Botev:2017:NLU

[3391] Z. I. Botev. The normal law under linear restrictions: simulation and estimation via minimax tilting. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1):

125–148, January 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wadsworth:2017:MAE

[3392] J. L. Wadsworth, J. A. Tawn, A. C. Davison, and D. M. Elton. Modelling across extremal dependence classes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1):149–175, January 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chen:2017:MFV

[3393] Kehui Chen, Pedro Delicado, and Hans-Georg Müller. Modelling function-valued stochastic processes, with applications to fertility dynamics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1):177–196, January 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cai:2017:OSD

[3394] T. Tony Cai and Wenguang Sun. Optimal screening and discovery of sparse signals with applications to multistage high throughput studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1):197– 223, January 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Stadler:2017:TST

[3395] Nicolas Städler and Sach Mukherjee. Two-sample testing in high dimensions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1):225–246, January 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Fan:2017:EHD

[3396] Jianqing Fan, Quefeng Li, and Yuyan Wang. Estimation of high dimensional mean regression in the absence of symmetry and light tail assumptions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1): 247–265, January 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Perry:2017:FMB

[3397] Patrick O. Perry. Fast momentbased estimation for hierarchical models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1):267–291, January 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chaudhuri:2017:HMC

[3398] Sanjay Chaudhuri, Debashis Mondal, and Teng Yin. Hamiltonian Monte Carlo sampling in Bayesian empirical likelihood computation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(1):293– 320, January 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2017:IIb

[3399] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2): 321–322, March 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Drton:2017:BIC

[3400] Mathias Drton and Martyn Plummer. A Bayesian information criterion for singular models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2):323–380, March 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Myllymaki:2017:GET

[3401] Mari Myllymäki, Tomáš Mrkvička, Pavel Grabarnik, Henri Seijo, and Ute Hahn. Global envelope tests for spatial processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2):381–404, March 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Fan:2017:HDS

[3402] Jianqing Fan, Han Liu, Yang Ning, and Hui Zou. High dimensional semiparametric latent graphical model for mixed data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2):405–421, March 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Sen:2017:TAL

[3403] Bodhisattva Sen and Mary Meyer. Testing against a linear regression model using ideas from shaperestricted estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2):423–448, March 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

He:2017:EED

[3404] Yi He and John H. J. Einmahl. Estimation of extreme depth-based quantile regions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2):449–461, March 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cornea:2017:RMR

[3405] Emil Cornea, Hongtu Zhu, Peter Kim, Joseph G. Ibrahim, and for the Alzheimer’s Disease Neuroimaging Initiative. Regression models on Riemannian symmetric spaces. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2): 463–482, March 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Prenen:2017:EAC

[3406] Leen Prenen, Roel Braekers, and Luc Duchateau. Extending the Archimedean copula methodology to model multivariate survival data grouped in clusters of variable size. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2):483–505, March 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Shao:2017:OEE

[3407] Qin Shao and Lijian Yang. Oracally efficient estimation and consistent model selection for auto-regressive moving average time series with trend. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2): 507–524, March 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Griffin:2017:CRM

[3408] Jim E. Griffin and Fabrizio Leisen. Compound random measures and their use in Bayesian non-parametrics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2): 525–545, March 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Soriano:2017:PMR

[3409] Jacopo Soriano and Li Ma. Probabilistic multi-resolution scanning for two-sample differences. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2):547– 572, March 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Mao:2017:EES

[3410] Lu Mao and D. Y. Lin. Efficient estimation of semiparametric transformation models for the cumulative incidence of competing risks. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2): 573–587, March 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lu:2017:CIR

[3411] Shu Lu, Yufeng Liu, Liang Yin, and Kai Zhang. Confidence intervals and regions for the lasso by using stochastic variational inequality techniques in optimization. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2):589–611, March 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Schreyer:2017:ERD

[3412] Manuela Schreyer, Roland Paulin, and Wolfgang Trutschnig. On the exact region determined by Kendall’s τ and Spearman’s ρ. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2):613–633, March 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wong:2017:FAC

[3413] Raymond K. W. Wong, Curtis B. Storlie, and Thomas C. M. Lee. A frequentist approach to computer model calibration. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(2):635–648, March 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2017:IIc

[3414] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3): 649–650, June 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dalalyan:2017:TGA

[3415] Arnak S. Dalalyan. Theoretical guarantees for approximate sampling from smooth and log-concave densities. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79 (3):651–676, June 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Gao:2017:HDC

[3416] Jiti Gao, Xiao Han, Guangming Pan, and Yanrong Yang. High dimensional correlation matrices: the central limit theorem and its applications. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3): 677–693, June 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Oates:2017:CFM

[3417] Chris J. Oates, Mark Girolami, and Nicolas Chopin. Control functionals for Monte Carlo integration. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3): 695–718, June 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2017:CAO

[3418] Linbo Wang, Thomas S. Richardson, and Xiao-Hua Zhou. Causal analysis of ordinal treatments and binary outcomes under truncation by death. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79 (3):719–735, June 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Gourieroux:2017:LEM

[3419] Christian Gouriéroux and Jean-Michel Zakoïan. Local explosion modelling by non-causal process. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3):737–756, June 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ding:2017:PSA

[3420] Peng Ding and Jiannan Lu. Principal stratification analysis using principal scores. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3):757–777, June 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Nandy:2017:AMB

[3421] Siddhartha Nandy, Chae Young Lim, and Tapabrata Maiti. Additive model building for spatial regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3): 779–800, June 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Leeb:2017:AUC

[3422] Hannes Leeb and Paul Kabaila. Admissibility of the usual confidence set for the mean of a univariate or bivariate normal population: the unknown variance case. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3):801–813, June 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bradley:2017:RMS

[3423] Jonathan R. Bradley, Christopher K.

Wikle, and Scott H. Holan. Regionalization of multiscale spatial processes by using a criterion for spatial aggregation error. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3):815–832, June 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Brockwell:2017:CAR

[3424] Peter J. Brockwell and Yasumasa Matsuda. Continuous auto-regressive moving average random fields on R^{n}. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3):

833–857, June 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ji:2017:ODL

[3425] Hao Ji and Hans-Georg Müller. Optimal designs for longitudinal and functional data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3):859–876, June 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rao:2017:BIM

[3426] Vinayak Rao, Ryan P. Adams, and David D. Dunson. Bayesian inference for Matérn repulsive processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3): 877–897, June 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Sun:2017:PST

[3427] Will Wei Sun, Junwei Lu, Han Liu, and Guang Cheng. Provable sparse tensor decomposition. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3):899–916, June 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

VanderWeele:2017:MAT

[3428] Tyler J. VanderWeele and Eric J. Tchetgen Tchetgen. Mediation analysis with time varying exposures and mediators. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3):917–938, June 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Belloni:2017:LCP

[3429] Alexandre Belloni, Mathieu Rosenbaum, and Alexandre B. Tsybakov. Linear and conic programming estimators in high dimensional errors-in-variables models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(3):939–956, June 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2017:IId

[3430] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79 (4):957–958, September 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.
1111/rssb.12210/full.

Cannings:2017:RPE

[3431] Timothy I. Cannings and Richard J.

Samworth. Random-projection ensemble classification. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4):959– 1035, September 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.
1111/rssb.12228/full.

Janson:2017:EIH

[3432] Lucas Janson, Rina Foygel Barber, and Emmanuel Candès. EigenPrism: inference for high dimensional signalto-noise ratios. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4):1037–1065, September 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        onlinelibrary.wiley.com/doi/10.
                                        1111/rssb.12203/full.

Bastide:2017:DAS

[3433] Paul Bastide, Mahendra Mariadassou, and Stéphane Robin. Detection of adaptive shifts on phylogenies by using shifted stochastic processes on a tree. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4):1067–1093, September 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

                                        URL http://onlinelibrary.wiley.
                                        com/doi/10.1111/rssb.12206/full.

Bhattacharya:2017:MRE

[3434] Bhaskar Bhattacharya and Mohammad Al-talib. A minimum relative entropy based correlation model between the response and covariates. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4):1095–1118, September 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

                                        URL http://onlinelibrary.wiley.
                                        com/doi/10.1111/rssb.12199/full.

Matias:2017:SCT

[3435] Catherine Matias and Vincent Miele. Statistical clustering of temporal networks through a dynamic stochastic block model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4):1119– 1141, September 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http:/

                                        /onlinelibrary.wiley.com/doi/10.
                                        1111/rssb.12200/full.

Fan:2017:EFD

[3436] Jianqing Fan and Xu Han. Estimation of the false discovery proportion with unknown dependence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4):1143–1164, September 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

URL http://onlinelibrary.wiley.
com/doi/10.1111/rssb.12204/full.

Jiang:2017:EOT

[3437] Runchao Jiang, Wenbin Lu, Rui Song, and Marie Davidian. On estimation of optimal treatment regimes for maximizing t-year survival probability. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4):1165–1185, September 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

URL http://onlinelibrary.wiley.
com/doi/10.1111/rssb.12201/full.

Roy:2017:CPE

[3438] Sandipan Roy, Yves Atchadé, and George Michailidis. Change point estimation in high dimensional Markov random-field models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4):1187– 1206, September 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.
1111/rssb.12205/full.

Pein:2017:HCP

[3439] Florian Pein, Hannes Sieling, and Axel Munk. Heterogeneous change point inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4):1207–1227, September 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

URL http://onlinelibrary.wiley.
com/doi/10.1111/rssb.12202/full.

Kennedy:2017:NPM

[3440] Edward H. Kennedy, Zongming Ma, Matthew D. McHugh, and Dylan S. Small. Non-parametric methods for doubly robust estimation of continuous treatment effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4): 1229–1245, September 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.
1111/rssb.12212/full.

Barber:2017:PFM

[3441] Rina Foygel Barber and Aaditya Ramdas. The p-filter: multilayer false discovery rate control for grouped hypotheses. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4):1247–1268, September 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.
1111/rssb.12218/full.

Zwiernik:2017:MLE

[3442] Piotr Zwiernik, Caroline Uhler, and Donald Richards. Maximum likelihood estimation for linear Gaussian covariance models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(4): 1269–1292, September 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.
1111/rssb.12217/full.

Anonymous:2017:IIe

[3443] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79 (5):1293–1294, November 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Caron:2017:SGU

[3444] François Caron and Emily B. Fox.

Sparse graphs using exchangeable random measures. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1295–1366, November 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Care:2017:CTL

[3445] Algo Carè, Simone Garatti, and Marco C. Campi. A coverage theory for least squares. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1367– 1389, November 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Truquet:2017:PSS

[3446] Lionel Truquet. Parameter stability and semiparametric inference in time varying auto-regressive conditional heteroscedasticity models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1391– 1414, November 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Fang:2017:TCI

[3447] Ethan X. Fang, Yang Ning, and Han Liu. Testing and confidence intervals for high dimensional proportional hazards models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1415– 1437, November 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chick:2017:BDT

[3448] Stephen Chick, Martin Forster, and Paolo Pertile. A Bayesian decision theoretic model of sequential experimentation with delayed response. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1439– 1462, November 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Brunner:2017:RBP

[3449] Edgar Brunner, Frank Konietschke, Markus Pauly, and Madan L. Puri. Rank-based procedures in factorial designs: hypotheses about nonparametric treatment effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79 (5):1463–1485, November 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Clertant:2017:SDF

[3450] M. Clertant and J. O’Quigley. Semiparametric dose finding methods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79 (5):1487–1508, November 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rudolph:2017:REE

[3451] Kara E. Rudolph and Mark J. van der Laan. Robust estimation of encouragement design intervention effects transported across sites. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1509– 1525, November 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Radchenko:2017:CCF

[3452] Peter Radchenko and Gourab Mukherjee. Convex clustering via l_{1} fusion penalization. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1527– 1546, November 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Huang:2017:OGT

[3453] Shih-Hao Huang, Mong-Na Lo Huang, Kerby Shedden, and Weng Kee Wong. Optimal group testing designs for estimating prevalence with uncertain testing errors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1547–1563, November 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Fan:2017:CAL

[3454] Caiyun Fan, Wenbin Lu, Rui Song, and Yong Zhou. Concordance-assisted learning for estimating optimal individualized treatment regimes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1565– 1582, November 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2017:JNP

[3455] Ching-Yun Wang, Harry Cullings, Xiao Song, and Kenneth J. Kopecky. Joint non-parametric correction estimator for excess relative risk regression in survival analysis with exposure measurement error. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1583– 1599, November 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rukhin:2017:ECM

[3456] Andrew L. Rukhin. Estimation of the common mean from heterogeneous normal observations with unknown variances. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1601–1618, November 2017. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Birr:2017:QSA

[3457] Stefan Birr, Stanislav Volgushev, Tobias Kley, Holger Dette, and Marc Hallin. Quantile spectral analysis for locally stationary time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79 (5):1619–1643, November 2017. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Frolich:2017:DIT

[3458] Markus Frölich and Martin Huber. Direct and indirect treatment effects– causal chains and mediation analysis with instrumental variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1645– 1666, November 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2017:IAI

[3459] Anonymous. Index of authors: Index of authors, volume 79, 2017. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1667– 1671, November 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2017:CCV

[3460] Anonymous. Contents: Contents of volume 79, 2017. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 79(5):1673– 1675, November 2017. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2018:IIa

[3461] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1): 1–2, January 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dunson:2018:RE

[3462] David Dunson and Piotr Fryzlewicz.

Report of the Editors — 2017. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1): 3–4, January 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Pfister:2018:KBT

[3463] Niklas Pfister, Peter Bühlmann, Bernhard Schölkopf, and Jonas Peters. Kernel-based tests for joint independence. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1):5–31, January 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Schouten:2018:SIB

[3464] Barry Schouten. Statistical inference based on randomly generated auxiliary variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1):33–56, January 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2018:HDC

[3465] Tengyao Wang and Richard J. Samworth. High dimensional change point estimation via sparse projection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1): 57–83, January 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kallus:2018:OPB

[3466] Nathan Kallus. Optimal a priori balance in the design of controlled experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1):85–112, January 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Shah:2018:GFT

[3467] Rajen D. Shah and Peter Bühlmann. Goodness-of-fit tests for high dimensional linear models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1):113– 135, January 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Hemerik:2018:FDP

[3468] Jesse Hemerik and Jelle J. Goeman.

False discovery proportion estimation by permutations: confidence for significance analysis of microarrays. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1): 137–155, January 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Goncalves:2018:EBI

[3469] Flávio B. Gonçalves and Dani Gamerman. Exact Bayesian inference in spatiotemporal Cox processes driven by multivariate Gaussian processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1): 157–175, January 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See corrigendum [3784].

Wang:2018:ALD

[3470] Boxiang Wang and Hui Zou. Another look at distance-weighted discrimination. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1):177–198, January 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dehaene:2018:EPL

[3471] Guillaume Dehaene and Simon Barthelmé.

Expectation propagation in the large data limit. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1):199–217, January 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Sommerfeld:2018:IEW

[3472] Max Sommerfeld and Axel Munk. Inference for empirical Wasserstein distances on finite spaces. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1):219– 238, January 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Choi:2018:GAC

[3473] Hyunphil Choi and Matthew Reimherr. A geometric approach to confidence regions and bands for functional parameters. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(1):239–260, January 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2018:IIb

[3474] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(2): 261–262, March 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Daouia:2018:ETR

[3475] Abdelaati Daouia, Stéphane Girard, and Gilles Stupfler. Estimation of tail risk based on extreme expectiles. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(2): 263–292, March 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Li:2018:STC

[3476] Weiming Li and Jianfeng Yao. On structure testing for component covariance matrices of a high dimensional mixture. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(2):293–318, March 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Koudstaal:2018:MGS

[3477] Mark Koudstaal and Fang Yao. From multiple Gaussian sequences to functional data and beyond: a Stein estimation approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(2):319–342, March 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2018:UEL

[3478] Honglang Wang, Ping-Shou Zhong, Yuehua Cui, and Yehua Li. Unified empirical likelihood ratio tests for functional concurrent linear models and the phase transition from sparse to dense functional data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(2):343–364, March 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Sengupta:2018:BMN

[3479] Srijan Sengupta and Yuguo Chen. A block model for node popularity in networks with community structure. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(2): 365–386, March 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ding:2018:MVR

[3480] Shanshan Ding and R. Dennis Cook.

Matrix variate regressions and envelope models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(2):387–408, March 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhao:2018:SDM

[3481] Zifeng Zhao and Zhengjun Zhang. Semiparametric dynamic max-copula model for multivariate time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(2):409–432, March 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2018:TML

[3482] Huixia Judy Wang, Ian W. McKeague, and Min Qian. Testing for marginal linear effects in quantile regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(2): 433–452, March 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2018:IIc

[3483] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(3): 453–454, June 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Yao:2018:TMI

[3484] Shun Yao, Xianyang Zhang, and Xiaofeng Shao. Testing mutual independence in high dimension via distance covariance. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(3):455–480, June 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lauritzen:2018:RNG

[3485] Steffen Lauritzen, Alessandro Rinaldo, and Kayvan Sadeghi. Random networks, graphical models and exchangeability. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(3):481–508, June 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Aue:2018:DDS

[3486] Alexander Aue, Gregory Rice, and Ozan Sönmez. Detecting and dating structural breaks in functional data without dimension reduction. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(3): 509–529, June 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2018:BEM

[3487] Linbo Wang and Eric Tchetgen Tchetgen. Bounded, efficient and multiply robust estimation of average treatment effects using instrumental variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(3):531–550, June 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Candes:2018:PGM

[3488] Emmanuel Candès, Yingying Fan, Lucas Janson, and Jinchi Lv. Panning for gold: ‘model-X’ knockoffs for high dimensional controlled variable selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80 (3):551–577, June 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Gronsbell:2018:SSA

[3489] Jessica L. Gronsbell and Tianxi Cai. Semi-supervised approaches to efficient evaluation of model prediction performance. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(3):579–594, June 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2018:IId

[3490] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80 (4):595–596, September 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Athey:2018:ARB

[3491] Susan Athey, Guido W. Imbens, and Stefan Wager. Approximate residual balancing: debiased inference of average treatment effects in high dimensions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(4):597–623, September 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Liang:2018:SEE

[3492] Liang Liang, Yanyuan Ma, Ying Wei, and Raymond J. Carroll. Semiparametrically efficient estimation in quantile regression of secondary analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(4):625–648, September 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lei:2018:AIP

[3493] Lihua Lei and William Fithian. AdaPT: an interactive procedure for multiple testing with side information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(4):649–679, September 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Shi:2018:MPL

[3494] Chengchun Shi, Rui Song, Wenbin Lu, and Bo Fu. Maximin projection learning for optimal treatment decision with heterogeneous individualized treatment effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(4):681– 702, September 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Krampe:2018:EWR

[3495] Jonas Krampe, Jens-Peter Kreiss, and Efstathios Paparoditis. Estimated Wold representation and spectraldensity-driven bootstrap for time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(4):703–726, September 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Khare:2018:BIG

[3496] Kshitij Khare, Bala Rajaratnam, and Abhishek Saha. Bayesian inference for Gaussian graphical models beyond decomposable graphs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(4):727– 747, September 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Titsias:2018:AGB

[3497] Michalis K. Titsias and Omiros Papaspiliopoulos. Auxiliary gradientbased sampling algorithms. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(4):749– 767, September 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2018:MNS

[3498] Fangfang Wang and Haonan Wang. Modelling non-stationary multivariate time series of counts via common factors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(4):769–791, September 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Guo:2018:CIC

[3499] Zijian Guo, Hyunseung Kang, T. Tony Cai, and Dylan S. Small. Confidence intervals for causal effects with invalid instruments by using two-stage hard thresholding with voting. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(4):793– 815, September 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Yu:2018:API

[3500] Dalei Yu, Xinyu Zhang, and Kelvin K. W. Yau. Asymptotic properties and information criteria for misspecified generalized linear mixed models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(4):817–836, September 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2018:IIe

[3501] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80 (5):837–838, November 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Deligiannidis:2018:CPM

[3502] George Deligiannidis, Arnaud Doucet, and Michael K. Pitt. The correlated pseudomarginal method. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):839– 870, November 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bloem-Reddy:2018:RWM

[3503] Benjamin Bloem-Reddy and Peter Orbanz. Random-walk models of network formation and sequential Monte Carlo methods for graphs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):871– 898, November 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Liang:2018:IRO

[3504] Faming Liang, Bochao Jia, Jingnan Xue, Qizhai Li, and Ye Luo. An imputation-regularized optimization algorithm for high dimensional missing data problems and beyond. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):899–926, November 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhu:2018:MMG

[3505] Yunzhang Zhu and Lexin Li. Multiple matrix Gaussian graphs estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):927–950, November 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chown:2018:DHN

[3506] Justin Chown and Ursula U. Müller. Detecting heteroscedasticity in nonparametric regression using weighted empirical processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):951– 974, November 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See corrigendum [3549].

Zheng:2018:HQR

[3507] Yao Zheng, Qianqian Zhu, Guodong Li, and Zhijie Xiao. Hybrid quantile regression estimation for time series models with conditional heteroscedasticity. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):975–993, November 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Liu:2018:FLI

[3508] Yang Liu, Yukun Liu, Pengfei Li, and Jing Qin. Full likelihood inference for abundance from continuous time capture-recapture data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):995– 1014, November 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Xie:2018:FDR

[3509] Jichun Xie and Ruosha Li. False discovery rate control for high dimensional networks of quantile associations conditioning on covariates. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):1015– 1034, November 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Fogarty:2018:MAL

[3510] Colin B. Fogarty. On mitigating the analytical limitations of finely stratified experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):1035– 1056, November 2018. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Tan:2018:SGE

[3511] Kean Ming Tan, Zhaoran Wang, Han Liu, and Tong Zhang. Sparse generalized eigenvalue problem: optimal statistical rates via truncated Rayleigh flow. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):1057–1086, November 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Linero:2018:BRT

[3512] Antonio R. Linero and Yun Yang.

Bayesian regression tree ensembles that adapt to smoothness and sparsity. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):1087–1110, November 2018. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2018:IAV

[3513] Anonymous. Index of authors, volume 80, 2018. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):1111–1113, November 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2018:CV

[3514] Anonymous. Contents of volume 80, 2018. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 80(5):1115–1116, November 2018. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2019:IIa

[3515] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(1): 1–2, February 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dunson:2019:RE

[3516] David Dunson and Simon Wood. Report of the Editors — 2018. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(1): 3–4, February 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bornn:2019:MCB

[3517] Luke Bornn, Neil Shephard, and Reza Solgi. Moment conditions and Bayesian non-parametrics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(1):5–43, February 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Li:2019:MTS

[3518] Ang Li and Rina Foygel Barber. Multiple testing with the structure-adaptive Benjamini–Hochberg algorithm. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(1): 45–74, February 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Luedtke:2019:ONP

[3519] Alex Luedtke, Marco Carone, and Mark J. van der Laan. An omnibus non-parametric test of equality in distribution for unknown functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(1): 75–99, February 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Schmidt:2019:CCK

[3520] Dennis Schmidt. Characterization of c-, L- and φ_{k}-optimal designs for a class of non-linear multiple-regression models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(1):101–120, February 2019. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kennedy:2019:RCI

[3521] Edward H. Kennedy, Scott Lorch, and Dylan S. Small. Robust causal inference with continuous instruments using the local instrumental variable curve. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(1):121–143, February 2019. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

She:2019:CVS

[3522] Yiyuan She and Hoang Tran. On crossvalidation for sparse reduced rank regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(1):145–161, February 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dobriban:2019:DPA

[3523] Edgar Dobriban and Art B. Owen. Deterministic parallel analysis: an improved method for selecting factors and principal components. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(1):163– 183, February 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2019:IIb

[3524] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(2): 185–186, April 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cai:2019:CAR

[3525] T. Tony Cai, Wenguang Sun, and Weinan Wang. Covariate-assisted ranking and screening for large-scale twosample inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(2):187–234, April 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bernton:2019:ABC

[3526] Espen Bernton, Pierre E. Jacob, Mathieu Gerber, and Christian P. Robert. Approximate Bayesian computation with the Wasserstein distance. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(2): 235–269, April 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Delaigle:2019:CFD

[3527] Aurore Delaigle, Peter Hall, and Tung Pham. Clustering functional data into groups by using projections. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(2): 271–304, April 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Han:2019:GFQ

[3528] Peisong Han, Linglong Kong, Jiwei Zhao, and Xingcai Zhou. A general framework for quantile estimation with incomplete data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(2):305–333, April 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Godolphin:2019:CRC

[3529] J. D. Godolphin. Construction of row-column factorial designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(2): 335–360, April 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ditlevsen:2019:HDF

[3530] Susanne Ditlevsen and Adeline Samson. Hypoelliptic diffusions: filtering and inference from complete and partial observations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(2):361–384, April 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhao:2019:MIP

[3531] Junlong Zhao, Chao Liu, Lu Niu, and Chenlei Leng. Multiple influential point detection in high dimensional regression spaces. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(2):385–408, April 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Piao:2019:SMB

[3532] Jin Piao, Jing Ning, and Yu Shen. Semiparametric model for bivariate survival data subject to biased sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(2):409–429, April 2019. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Liang:2019:SIP

[3533] Tengyuan Liang and Weijie J. Su. Statistical inference for the population landscape via moment-adjusted stochastic gradients. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(2):431–456, April 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2019:IIc

[3534] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(3): 457–458, July 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Frot:2019:RCS

[3535] Benjamin Frot, Preetam Nandy, and Marloes H. Maathuis. Robust causal structure learning with some hidden variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(3):459–487, July 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zanella:2019:SIT

[3536] Giacomo Zanella and Gareth Roberts.

Scalable importance tempering and Bayesian variable selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(3): 489–517, July 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Plumlee:2019:CMC

[3537] Matthew Plumlee. Computer model calibration with confidence and consistency. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(3):519–545, July 2019. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Heller:2019:PSE

[3538] Ruth Heller, Amit Meir, and Nilanjan Chatterjee. Post-selection estimation and testing following aggregate association tests. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(3):547–573, July 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bhattacharya:2019:GAF

[3539] Bhaswar B. Bhattacharya. A general asymptotic framework for distributionfree graph-based two-sample tests.

Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81 (3):575–602, July 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Niu:2019:IGP

[3540] Mu Niu, Pokman Cheung, Lizhen Lin, Zhenwen Dai, Neil Lawrence, and David Dunson. Intrinsic Gaussian processes on complex constrained domains. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(3):603–627, July 2019. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dong:2019:LFT

[3541] Chen Dong, Guodong Li, and Xingdong Feng. Lack-of-fit tests for quantile regression models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(3):629–648, July 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Baranowski:2019:NTD

[3542] Rafal Baranowski, Yining Chen, and Piotr Fryzlewicz. Narrowest-over-threshold detection of multiple change points and change-point-like features. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81 (3):649–672, July 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2019:IId

[3543] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81 (4):673–674, September 2019. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cai:2019:HDL

[3544] T. Tony Cai and Linjun Zhang. High dimensional linear discriminant analysis: optimality, adaptive algorithm and missing data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(4):675–705, September 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Xiang:2019:SCI

[3545] Dongdong Xiang, Sihai Dave Zhao, and T. Tony Cai. Signal classification for the integrative analysis of multiple sequences of large-scale multiple tests. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(4):707–734, September 2019. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhao:2019:SAI

[3546] Qingyuan Zhao, Dylan S. Small, and Bhaswar B. Bhattacharya. Sensitivity analysis for inverse probability weighting estimators via the percentile bootstrap. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(4):735–761, September 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See correction [3852].

Zhang:2019:MMA

[3547] Xinyu Zhang, Yanyuan Ma, and Raymond J. Carroll. MALMEM: model averaging in linear measurement error models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(4):763–779, September 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kowal:2019:DSP

[3548] Daniel R. Kowal, David S. Matteson, and David Ruppert. Dynamic shrinkage processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(4):781–804, September 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chown:2019:CDH

[3549] Justin Chown and Ursula U. Müller. Corrigendum: Detecting heteroscedasticity in non-parametric regression using weighted empirical processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81 (4):805–806, September 2019. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3506].

Anonymous:2019:II

[3550] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81 (5):807–808, November 2019. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Fuquene:2019:CMC

[3551] Jairo Fúquene, Mark Steel, and David Rossell. On choosing mixture components via non-local priors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(5):809– 837, November 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bassett:2019:FDE

[3552] Robert Bassett and James Sharpnack.

Fused density estimation: theory and methods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(5):839–860, November 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Escobar-Bach:2019:NPC

[3553] Mikael Escobar-Bach and Ingrid Van Keilegom. Non-parametric cure rate estimation under insufficient followup by using extremes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(5):861– 880, November 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Singh:2019:DEO

[3554] Satya Prakash Singh and Ori Davidov. On the design of experiments with ordered treatments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(5):881– 900, November 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Greenewald:2019:TGL

[3555] Kristjan Greenewald, Shuheng Zhou, and Alfred Hero III. Tensor graphical lasso (TeraLasso). Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(5):901– 931, November 2019. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2019:IAV

[3556] Anonymous. Index of authors, volume 81, 2019. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(5):933–935, November 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2019:CV

[3557] Anonymous. Contents of volume 81, 2019. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 81(5):937–938, November 2019. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2020:IIa

[3558] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1): 1–2, February 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dunson:2020:RE

[3559] David Dunson and Simon Wood. Report of the Editors — 2019. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1): 3–4, February 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Khismatullina:2020:MIL

[3560] Marina Khismatullina and Michael Vogt. Multiscale inference and long-run variance estimation in non-parametric regression with time series errors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1):

5–37, February 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cinelli:2020:MSS

[3561] Carlos Cinelli and Chad Hazlett. Making sense of sensitivity: extending omitted variable bias. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1):39–67, February 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Luo:2020:REI

[3562] Lan Luo and Peter X.-K. Song. Renewable estimation and incremental inference in generalized linear models with streaming data sets. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1): 69–97, February 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chen:2020:TSM

[3563] Fan Chen, Yini Zhang, and Karl Rohe.

Targeted sampling from massive block model graphs with personalized PageRank. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1):99–126, February 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Christensen:2020:BHM

[3564] Jonathan Christensen and Li Ma. A Bayesian hierarchical model for related densities by using Pólya trees. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1): 127–153, February 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhao:2020:BEL

[3565] Puying Zhao, Malay Ghosh, J. N. K. Rao, and Changbao Wu. Bayesian empirical likelihood inference with complex survey data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1):155–174, February 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Berrett:2020:CPT

[3566] Thomas B. Berrett, Yi Wang, Rina Foygel Barber, and Richard J. Samworth. The conditional permutation test for independence while controlling for confounders. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1):175– 197, February 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Fulcher:2020:RIP

[3567] Isabel R. Fulcher, Ilya Shpitser, Stella Marealle, and Eric J. Tchetgen Tchetgen. Robust inference on population indirect causal effects: the generalized front door criterion. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1):199– 214, February 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bolin:2020:MTM

[3568] David Bolin and Jonas Wallin. Multivariate type G Matérn stochastic partial differential equation random fields. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1):215–239, February 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Li:2020:RRA

[3569] Xinran Li and Peng Ding. Rerandomization and regression adjustment. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1):241–268, February 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Grover:2020:CNR

[3570] Lovleen Kumar Grover and Amanpreet Kaur. Correction: ‘A new randomized response model’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(1):269–271, February 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2897] and reply [3593].

Anonymous:2020:IIb

[3571] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(2): 273–274, April 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dubey:2020:FMT

[3572] Paromita Dubey and Hans-Georg Müller. Functional models for timevarying random objects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(2):275–327, April 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See errata [3656].

Gataric:2020:SPC

[3573] Milana Gataric, Tengyao Wang, and Richard J. Samworth. Sparse principal component analysis via axisaligned random projections. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(2): 329–359, April 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Shah:2020:RSV

[3574] Rajen D. Shah, Benjamin Frot, Gian-Andrea Thanei, and Nicolai Meinshausen. Right singular vector projection graphs: fast high dimensional covariance matrix estimation under latent confounding. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(2):361–389, April 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cai:2020:SIE

[3575] T. Tony Cai and Zijian Guo. Semisupervised inference for explained variance in high dimensional linear regression and its applications. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(2): 391–419, April 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Frazier:2020:MMA

[3576] David T. Frazier, Christian P. Robert, and Judith Rousseau. Model misspecification in approximate Bayesian computation: consequences and diagnostics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(2):421–444, April 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Yang:2020:DRI

[3577] Shu Yang, Jae Kwang Kim, and Rui Song. Doubly robust inference when combining probability and nonprobability samples with high dimensional data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(2):445–465, April 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Jiang:2020:SSU

[3578] Jiming Jiang and Mahmoud Torabi.

Sumca: simple, unified, Monte-Carlo-assisted approach to second-order unbiased mean-squared prediction error estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(2):467–485, April 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Todeschini:2020:ERM

[3579] Adrien Todeschini, Xenia Miscouridou, and François Caron. Exchangeable random measures for sparse and modular graphs with overlapping communities. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82 (2):487–520, April 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Shi:2020:MRC

[3580] Xu Shi, Wang Miao, Jennifer C. Nelson, and Eric J. Tchetgen Tchetgen. Multiply robust causal inference with double-negative control adjustment for categorical unmeasured confounding. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82 (2):521–540, April 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2020:IIc

[3581] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3): 541–542, July 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Jacob:2020:UMC

[3582] Pierre E. Jacob, John O’Leary, and Yves F. Atchadé. Unbiased Markov chain Monte Carlo methods with couplings. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3):543–600, July 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Prasad:2020:RER

[3583] Adarsh Prasad, Arun Sai Suggala, Sivaraman Balakrishnan, and Pradeep Ravikumar. Robust estimation via robust gradient estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3): 601–627, July 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dette:2020:TRH

[3584] Holger Dette, Kevin Kokot, and Stanislav Volgushev. Testing relevant hypotheses in functional time series via self-normalization. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3): 629–660, July 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Diaz:2020:CMA

[3585] Iván Díaz and Nima S. Hejazi. Causal mediation analysis for stochastic interventions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3):661–683, July 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Javanmard:2020:FFH

[3586] Adel Javanmard and Jason D. Lee.

A flexible framework for hypothesis testing in high dimensions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3): 685–718, July 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Westling:2020:CIR

[3587] Ted Westling, Peter Gilbert, and Marco Carone. Causal isotonic regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3):719–747, July 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rosenblum:2020:OTS

[3588] Michael Rosenblum, Ethan X. Fang, and Han Liu. Optimal, two-stage, adaptive enrichment designs for randomized trials, using sparse linear programming. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3):749–772, July 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Jankova:2020:GFT

[3589] Jana Janková, Rajen D. Shah, Peter Bühlmann, and Richard J. Samworth. Goodness-of-fit testing in high dimensional generalized linear models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3):

773–795, July 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chauvet:2020:ITS

[3590] Guillaume Chauvet and Audrey-Anne Vallée. Inference for two-stage sampling designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3):797–815, July 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Jiang:2020:BCS

[3591] Zhenyu Jiang, Nengxiang Ling, Zudi Lu, Dag Tjøstheim, and Qiang Zhang. On bandwidth choice for spatial data density estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3):817–840, July 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Hemerik:2020:RTG

[3592] Jesse Hemerik, Jelle J. Goeman, and Livio Finos. Robust testing in generalized linear models by sign flipping score contributions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3):841–864, July 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Singh:2020:RCG

[3593] Sarjinder Singh. Reply to the correction by Grover and Kaur: a new randomized response model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(3): 865–868, July 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [2897, 3570].

Anonymous:2020:IId

[3594] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82 (4):869–870, September 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Engelke:2020:GME

[3595] Sebastian Engelke and Adrien S. Hitz.

Graphical models for extremes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(4):871– 932, September 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Liu:2020:UDA

[3596] Bin Liu, Cheng Zhou, Xinsheng Zhang, and Yufeng Liu. A unified dataadaptive framework for high dimensional change point detection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(4):933– 963, September 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rad:2020:SES

[3597] Kamiar Rahnama Rad and Arian Maleki. A scalable estimate of the outof-sample prediction error via approximate leave-one-out cross-validation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(4):965–996, September 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Taeb:2020:FDC

[3598] Armeen Taeb, Parikshit Shah, and Venkat Chandrasekaran. False discovery and its control in low rank estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(4):997–1027, September 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lane:2020:ADO

[3599] Adam Lane. Adaptive designs for optimal observed Fisher information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82 (4):1029–1058, September 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Apley:2020:VEP

[3600] Daniel W. Apley and Jingyu Zhu. Visualizing the effects of predictor variables in black box supervised learning models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(4):1059–1086, September 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Fortini:2020:QBP

[3601] Sandra Fortini and Sonia Petrone. Quasi-Bayes properties of a procedure for sequential learning in mixture models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(4):1087–1114, September 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Poss:2020:SEP

[3602] Dominik Poß, Dominik Liebl, Alois Kneip, Hedwig Eisenbarth, Tor D. Wager, and Lisa Feldman Barrett. Superconsistent estimation of points of impact in non-parametric regression with functional predictors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(4):1115– 1140, September 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Silva:2020:OAS

[3603] Ivair R. Silva, Martin Kulldorff, and W. Katherine Yih. Optimal alpha spending for sequential analysis with binomial data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(4):1141– 1164, September 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2020:IIe

[3604] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82 (5):1165–1166, December 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Pollock:2020:QSM

[3605] Murray Pollock, Paul Fearnhead, Adam M. Johansen, and Gareth O. Roberts. Quasi-stationary Monte Carlo and the ScaLE algorithm. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(5):1167– 1221, December 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Mozgunov:2020:ITA

[3606] Pavel Mozgunov and Thomas Jaki. An information theoretic approach for selecting arms in clinical trials. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82 (5):1223–1247, December 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Mukhopadhyay:2020:EDN

[3607] Minerva Mukhopadhyay, Didong Li, and David B. Dunson. Estimating densities with non-linear support by using Fisher–Gaussian kernels. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(5):1249– 1271, December 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2020:SNA

[3608] Gao Wang, Abhishek Sarkar, Peter Carbonetto, and Matthew Stephens. A simple new approach to variable selection in regression, with application to genetic fine mapping. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(5):1273– 1300, December 2020. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ye:2020:RTT

[3609] Ting Ye and Jun Shao. Robust tests for treatment effect in survival analysis under covariate-adaptive randomization. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(5):1301–1323, December 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Gorgi:2020:BNB

[3610] Paolo Gorgi. Beta-negative binomial auto-regressions for modelling integervalued time series with extreme observations. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(5):1325–1347, December 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Tang:2020:MLR

[3611] Yanbo Tang and Nancy Reid. Modified likelihood root in high dimensions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(5):1349–1369, December 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Richardson:2020:SMU

[3612] Robert Richardson, Athanasios Kottas, and Bruno Sansó. Spatiotemporal modelling using integro-difference equations with bivariate stable kernels. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(5):1371–1392, December 2020. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3469].

Anonymous:2020:IAV

[3613] Anonymous. Index of authors, volume 82, 2020. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(5):1393–1396, December 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2020:CV

[3614] Anonymous. Contents of volume 82, 2020. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 82(5):1397–1398, December 2020. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2021:IIa

[3615] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(1): 1–2, February 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Delaigle:2021:RE

[3616] Aurore Delaigle and Simon Wood. Report of the Editors — 2020. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(1): 3–4, February 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Godolphin:2021:CBF

[3617] J. D. Godolphin. Construction of blocked factorial designs to estimate main effects and selected two-factor interactions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(1):5–29, February 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Tan:2021:UMR

[3618] Linda S. L. Tan. Use of model reparametrization to improve variational Bayes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(1):30–57, February 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Xia:2021:SIL

[3619] Dong Xia and Ming Yuan. Statistical inferences of linear forms for noisy matrix completion. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(1):58–77, February 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Salvati:2021:SAE

[3620] N. Salvati, E. Fabrizi, M. G. Ranalli, and R. L. Chambers. Small area estimation with linked data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(1):78– 107, February 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kim:2021:BRS

[3621] Kwang-Rae Kim, Ian L. Dryden, Huiling Le, and Katie E. Severn. Book review: Smoothing splines on Riemannian manifolds, with applications to 3D shape space. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(1):108–132, February 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Heller:2021:OCF

[3622] Ruth Heller and Saharon Rosset. Optimal control of false discovery criteria in the two-group model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(1):133– 155, February 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See erratum [3655].

Heng:2021:GFA

[3623] Jeremy Heng, Arnaud Doucet, and Yvo Pokern. Gibbs flow for approximate transport with applications to Bayesian computation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(1):156–187, February 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Toulis:2021:PRM

[3624] Panos Toulis, Thibaut Horel, and Edoardo M. Airoldi. The proximal Robbins–Monro method. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(1):188– 212, February 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2021:IIb

[3625] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(2): 213–214, April 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rothenhausler:2021:ARH

[3626] Dominik Rothenhäusler, Nicolai Meinshausen, Peter Bühlmann, and Jonas Peters. Anchor regression: Heterogeneous data meet causality. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(2): 215–246, April 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See erratum [3665].

Yu:2021:FSC

[3627] Mengjia Yu and Xiaohui Chen. Finite sample change point inference and identification for high-dimensional mean vectors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(2):247–270, April 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Xu:2021:IAE

[3628] Sheng Xu and Zhou Fan. Iterative alpha expansion for estimating gradientsparse signals from linear measurements. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(2):271–292, April 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Noroozi:2021:ECP

[3629] Majid Noroozi, Ramchandra Rimal, and Marianna Pensky. Estimation and clustering in popularity adjusted block model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(2):293–317, April 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Pu:2021:EOT

[3630] Hongming Pu and Bo Zhang. Estimating optimal treatment rules with an instrumental variable: A partial identification learning approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(2): 318–345, April 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ferraccioli:2021:NDE

[3631] Federico Ferraccioli, Eleonora Arnone, Livio Finos, James O. Ramsay, and Laura M. Sangalli. Nonparametric density estimation over complicated domains. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(2):346–368, April 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Meng:2021:PME

[3632] Kun Meng and Ani Eloyan. Principal manifold estimation via model complexity selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(2):369–394, April 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Johansson:2021:ORD

[3633] Per Johansson, Donald B. Rubin, and Mårten Schultzberg. On optimal rerandomization designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(2): 395–403, April 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kallus:2021:ORE

[3634] Nathan Kallus. On the optimality of randomization in experimental design: How to randomize for minimax variance and design-based inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(2): 404–409, April 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2021:IIc

[3635] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(3): 411–412, July 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Reimherr:2021:PSS

[3636] Matthew Reimherr, Xiao-Li Meng, and Dan L. Nicolae. Prior sample size extensions for assessing prior impact and prior-likelihood discordance. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(3): 413–437, July 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kim:2021:VAV

[3637] Sungwook Kim, Michael P. Fay, and Michael A. Proschan. Valid and approximately valid confidence intervals for current status data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(3): 438–452, July 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Liu:2021:VSA

[3638] Yi Liu, Veronika Rocková, and Yuexi Wang. Variable selection with ABC Bayesian forests. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(3):453–481, July 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Heng:2021:IPO

[3639] Siyu Heng, Hyunseung Kang, Dylan S.

Small, and Colin B. Fogarty. Increasing power for observational studies of aberrant response: An adaptive approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83 (3):482–504, July 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Mourtada:2021:AAM

[3640] Jaouad Mourtada, Stéphane Gaïffas, and Erwan Scornet. AMF: Aggregated Mondrian forests for online learning. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83 (3):505–533, July 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Li:2021:GKF

[3641] Jinzhou Li and Marloes H. Maathuis. GGM knockoff filter: False discovery rate control for Gaussian graphical models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(3):534–558, July 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wei:2021:ECQ

[3642] Bo Wei, Limin Peng, Mei-Jie Zhang, and Jason P. Fine. Estimation of causal quantile effects with a binary instrumental variable and censored data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83 (3):559–578, July 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Stokell:2021:MHD

[3643] Benjamin G. Stokell, Rajen D. Shah, and Ryan J. Tibshirani. Modelling high-dimensional categorical data using nonconvex fusion penalties. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(3): 579–611, July 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lee:2021:IRE

[3644] Myoung jae Lee. Instrument residual estimator for any response variable with endogenous binary treatment. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(3):612–635, July 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2021:IId

[3645] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83 (4):637–638, September 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Crane:2021:IHR

[3646] Harry Crane and Min Xu. Inference on the history of a randomly growing tree. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(4):639–668, September 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cai:2021:OSI

[3647] Tianxi Cai, T. Tony Cai, and Zijian Guo. Optimal statistical inference for individualized treatment effects in high-dimensional models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83 (4):669–719, September 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ignatiadis:2021:CPC

[3648] Nikolaos Ignatiadis and Wolfgang Huber. Covariate powered cross-weighted multiple testing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(4):720–751, September 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Windmeijer:2021:CIM

[3649] Frank Windmeijer, Xiaoran Liang, Fernando P. Hartwig, and Jack Bowden. The confidence interval method for selecting valid instrumental variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(4):752–776, September 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Luo:2021:LFR

[3650] Xiaokang Luo, Tirthankar Dasgupta, Minge Xie, and Regina Y. Liu. Leveraging the Fisher randomization test using confidence distributions: Inference, combination and fusion learning. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(4):777–797, September 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lavancier:2021:SBD

[3651] Frédéric Lavancier and Ronan Le Guével. Spatial birth-death-move processes: Basic properties and estimation of their intensity functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(4):798– 825, September 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chen:2021:JQR

[3652] Xu Chen and Surya T. Tokdar. Joint quantile regression for spatial data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(4):826–852, September 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rossell:2021:ALA

[3653] David Rossell, Oriol Abril, and Anirban Bhattacharya. Approximate Laplace approximations for scalable model selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(4):853–879, September 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2021:CWL

[3654] Xiangyu Wang, Chenlei Leng, and Tom Boot. Corrigendum: Wang and Leng (2016), High-dimensional ordinary least-squares projection for screening variables, Journal of the Royal Statistical Society Series B, 78, 589–611. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(4):880–881, September 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3360].

Wang:2021:EOC

[3655] Xiangyu Wang and Chenlei Leng. Erratum: Optimal control of false discovery criteria in the two-group model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 78(3): 882, September 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3622].

Dubey:2021:EFM

[3656] Paromita Dubey and Hans-Georg Müller. Errata to “Functional models for time-varying random objects”. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(1):883, September 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3572].

Anonymous:2021:IIe

[3657] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83 (5):885–886, November 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chen:2021:ANS

[3658] Mingli Chen, Kengo Kato, and Chenlei Leng. Analysis of networks via the sparse β-model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(5):887–910, November 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lei:2021:CIC

[3659] Lihua Lei and Emmanuel J. Candès. Conformal inference of counterfactuals and individual treatment effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(5):911–938, November 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kim:2021:TSI

[3660] Byol Kim, Song Liu, and Mladen Kolar. Two-sample inference for highdimensional Markov networks. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(5):939– 962, November 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Henzi:2021:IDR

[3661] Alexander Henzi, Johanna F. Ziegel, and Tilmann Gneiting. Isotonic distributional regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(5):963– 993, November 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Su:2021:MAA

[3662] Fangzhou Su and Peng Ding. Modelassisted analyses of cluster-randomized experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(5):994–1015, November 2021. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Qiu:2021:IHT

[3663] Yumou Qiu, Jing Tao, and Xiao-Hua Zhou. Inference of heterogeneous treatment effects using observational data with high-dimensional covariates. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(5):1016–1043, November 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Tang:2021:ICN

[3664] Wenpin Tang, Lu Zhang, and Sudipto Banerjee. On identifiability and consistency of the nugget in Gaussian spatial process models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(5):1044– 1070, November 2021. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2021:EAR

[3665] Anonymous. Erratum: Anchor regression: Heterogeneous data meet causality. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(5):1071, November 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3626].

Rothenhausler:2021:CV

[3666] Dominik Rothenhäusler, Nicolai Meinshausen, Peter Bühlmann, and Jonas Peters. Contents of volume 83, 2021. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 83(5):1072–1073, November 2021. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2022:IIa

[3667] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1): 1–2, February 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dong:2022:GDP

[3668] Jinshuo Dong, Aaron Roth, and Weijie J. Su. Gaussian differential privacy. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1):3–37, February 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See discussion [3669, 3670, 3671, 3672, 3673, 3674, 3675, 3676, 3677].

Balle:2022:PVT

[3669] Borja Balle. Proposer of the vote of thanks to Dong et al. and contribution to the discussion of ‘Gaussian Differential Privacy’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1):37–38, February 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3668].

Avella-Medina:2022:SVT

[3670] Marco Avella-Medina. Seconder of the vote of thanks to Dong et al. and contribution to the discussion of ‘Gaussian Differential Privacy’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1): 39–41, February 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3668].

Krusche:2022:PKF

[3671] Peter Krusche and Frank Bretz. Peter Krusche and Frank Bretz’s contribution to the discussion of ‘Gaussian Differential Privacy’ by Dong et al. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1): 42–43, February 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3668].

Chai:2022:CPC

[3672] Christine P. Chai. Christine P. Chai’s contribution to the discussion of ‘Gaussian Differential Privacy’ by Dong et al. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1):43–44, February 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3668].

Dietz:2022:SDC

[3673] Sebastian Dietz. Sebastian Dietz’s contribution to the discussion of ‘Gaussian Differential Privacy’ by Dong et al. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1):44–45, February 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3668].

Goseling:2022:JGM

[3674] J. Goseling and M. N. M. van Lieshout. J. Goseling and M. N. M. van Lieshout’s contribution to the discussion of ‘Gaussian Differential Privacy’ by Dong et al. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1):46–47, February 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3668].

Mateu:2022:JMC

[3675] Jorge Mateu. Jorge Mateu’s contribution to the discussion of ‘Gaussian Differential Privacy’ by Dong et al. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1): 47–48, February 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3668].

Wijayatunga:2022:PWC

[3676] Priyantha Wijayatunga. Priyantha Wijayatunga’s contribution to the discussion of ‘Gaussian Differential Privacy’ by Dong et al. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1):49–50, February 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3668].

Dong:2022:ARD

[3677] Jinshuo Dong, Aaron Roth, and Weijie J. Su. Authors’ reply to the discussion of ‘Gaussian Differential Privacy’ by Dong et al. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1):50–54, February 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3668].

Jiang:2022:UPA

[3678] Jiming Jiang, Matt P. Wand, and Aishwarya Bhaskaran. Usable and precise asymptotics for generalized linear mixed model analysis and design. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1): 55–82, February 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See erratum [3785].

Chen:2022:IWG

[3679] Yao Chen, Qingyi Gao, and Xiao Wang. Inferential Wasserstein generative adversarial networks. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1):83– 113, February 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dau:2022:WFS

[3680] Hai-Dang Dau and Nicolas Chopin.

Waste-free sequential Monte Carlo. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1):114–148, February 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Li:2022:TLH

[3681] Sai Li, T. Tony Cai, and Hongzhe Li.

Transfer learning for high-dimensional linear regression: Prediction, estimation and minimax optimality. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1): 149–173, February 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Puelz:2022:GTA

[3682] David Puelz, Guillaume Basse, Avi Feller, and Panos Toulis. A graphtheoretic approach to randomization tests of causal effects under general interference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1):174–204, February 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Tan:2022:HDQ

[3683] Kean Ming Tan, Lan Wang, and Wen-Xin Zhou. High-dimensional quantile regression: Convolution smoothing and concave regularization. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1): 205–233, February 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chen:2022:HDM

[3684] Yudong Chen, Tengyao Wang, and Richard J. Samworth. High-dimensional, multiscale online changepoint detection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(1):234–266, February 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2022:IIb

[3685] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2): 267–268, April 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Luo:2022:EDR

[3686] Wei Luo. On efficient dimension reduction with respect to the interaction between two response variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2): 269–294, April 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Cohen:2022:GPF

[3687] Peter L. Cohen and Colin B. Fogarty. Gaussian prepivoting for finite population causal inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2): 295–320, April 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Syed:2022:NRP

[3688] Saifuddin Syed, Alexandre Bouchard-Côté, George Deligiannidis, and Arnaud Doucet. Non-reversible parallel tempering: a scalable highly parallel MCMC scheme. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2):321–350, April 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Ben-Michael:2022:SCS

[3689] Eli Ben-Michael, Avi Feller, and Jesse Rothstein. Synthetic controls with staggered adoption. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2):351–381, April 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhao:2022:SIE

[3690] Qingyuan Zhao, Dylan S. Small, and Ashkan Ertefaie. Selective inference for effect modification via the lasso. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2): 382–413, April 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Dunson:2022:GBG

[3691] David B. Dunson, Hau-Tieng Wu, and Nan Wu. Graph based Gaussian processes on restricted domains. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2): 414–439, April 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Mo:2022:ELO

[3692] Weibin Mo and Yufeng Liu. Efficient learning of optimal individualized treatment rules for heteroscedastic or misspecified treatment-free effect models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2):440–472, April 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

McElroy:2022:MIT

[3693] Tucker S. McElroy and Anindya Roy. Model identification via total Frobenius norm of multivariate spectra. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2): 473–495, April 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Livingstone:2022:BPC

[3694] Samuel Livingstone and Giacomo Zanella. The Barker proposal: Combining robustness and efficiency in gradient-based MCMC. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2):496–523, April 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Guan:2022:POD

[3695] Leying Guan and Robert Tibshirani.

Prediction and outlier detection in classification problems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2):524–546, April 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Sun:2022:KES

[3696] Yan Sun and Faming Liang. A kernelexpanded stochastic neural network. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (2):547–578, April 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Henckel:2022:GCE

[3697] Leonard Henckel, Emilija Perković, and Marloes H. Maathuis. Graphical criteria for efficient total effect estimation via adjustment in causal linear models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (2):579–599, April 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lee:2022:FSE

[3698] Kuang-Yao Lee and Lexin Li. Functional structural equation model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2): 600–629, April 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Fan:2022:SSI

[3699] Jianqing Fan, Yingying Fan, Xiao Han, and Jinchi Lv. SIMPLE: Statistical inference on membership profiles in large networks. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(2):630–653, April 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2022:IIc

[3700] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 655–656, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Vansteelandt:2022:ALI

[3701] Stijn Vansteelandt and Oliver Dukes. Assumption-lean inference for generalised linear model parameters. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 657–685, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See remainder of this issue for numerous discussions.

Daniel:2022:PVT

[3702] Rhian M. Daniel. Proposer of the vote of thanks and contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 686–689, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Didelez:2022:SVT

[3703] Vanessa Didelez. Seconder of the vote of thanks to Vansteelandt and Dukes and contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 689–691, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Ding:2022:PDC

[3704] Peng Ding. Peng Ding’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (3):691–693, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Stensrud:2022:MJS

[3705] Mats J. Stensrud and Aaron L.

Sarvet. Mats J. Stensrud and Aaron L. Sarvet’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 694–696, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Battey:2022:HBC

[3706] Heather Battey. Heather Battey’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 696–698, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Hennig:2022:CHC

[3707] Christian Hennig. Christian Hennig’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 698–699, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Basu:2022:PBC

[3708] Pallavi Basu. Pallavi Basu’s contribution to the Discussion of ‘Assumptionlean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3):700–701, July 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Bilodeau:2022:BBC

[3709] Blair Bilodeau. Blair Bilodeau’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 701–702, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Buja:2022:ABR

[3710] Andreas Buja, Richard A. Berk, Arun K. Kuchibhotla, Linda Zhao, and Ed George. Andreas Buja, Richard A. Berk, Arun K. Kuchibhotla, Linda Zhao and Ed George’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (3):703–705, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Choi:2022:ACW

[3711] Anna Choi and Weng Kee Wong. Anna Choi and Weng Kee Wong’s contribution to the Discussion of ‘Assumptionlean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3):705–706, July 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Dong:2022:CDJ

[3712] Chaohua Dong, Jiti Gao, and Oliver Linton. Chaohua Dong, Jiti Gao and Oliver Linton’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3):

707–708, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Hines:2022:OHK

[3713] Oliver Hines and Karla Diaz-Ordaz. Oliver Hines and Karla Diaz-Ordaz’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 709–710, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Hunt:2022:IHC

[3714] Ian Hunt. Ian Hunt’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (3):711–712, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Kumar:2022:KKC

[3715] Kuldeep Kumar. Kuldeep Kumar’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 712–713, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Lavine:2022:MLJ

[3716] Michael Lavine and James Hodges. Michael Lavine and James Hodges’ contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 713–714, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Ogburn:2022:EOJ

[3717] Elizabeth L. Ogburn, Junhui Cai, Arun K. Kuchibhotla, Richard A. Berk, and Andreas Buja. Elizabeth L Ogburn, Junhui Cai, Arun K Kuchibhotla, Richard A Berk and Andreas Buja’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 715–716, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Phillips:2022:RVP

[3718] Rachael V. Phillips and Mark J.

van der Laan. Rachael V. Phillips and Mark J. van der Laan’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (3):717–718, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Richardson:2022:TRC

[3719] Thomas S. Richardson. Thomas S.

Richardson’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 719–720, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Shpitser:2022:ISC

[3720] Ilya Shpitser. Ilya Shpitser’s contribution to the Discussion of ‘Assumptionlean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3):720–721, July 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Tang:2022:YTC

[3721] Yanbo Tang. Yanbo Tang’s contribution to the Discussion of ‘Assumptionlean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3):722–723, July 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Tchetgen:2022:EJT

[3722] Eric J. Tchetgen Tchetgen. Eric J. Tchetgen Tchetgen’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (3):723–725, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Zhao:2022:JZC

[3723] Jiwei Zhao. Jiwei Zhao’s contribution to the Discussion of ‘Assumptionlean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3):725–726, July 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Zhou:2022:NZX

[3724] Niwen Zhou and Xu Guo. Niwen Zhou and Xu Guo’s contribution to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 727–729, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Vansteelandt:2022:ARD

[3725] Stijn Vansteelandt and Oliver Dukes.

Authors’ reply to the Discussion of ‘Assumption-lean inference for generalised linear model parameters’ by Vansteelandt and Dukes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 729–739, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3701].

Chib:2022:BEC

[3726] Siddhartha Chib, Minchul Shin, and Anna Simoni. Bayesian estimation and comparison of conditional moment models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3):740–764, July 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Shi:2022:SIV

[3727] Chengchun Shi, Sheng Zhang, Wenbin Lu, and Rui Song. Statistical inference of the value function for reinforcement learning in infinite-horizon settings. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3):765–793, July 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhou:2022:SEC

[3728] Xiang Zhou. Semiparametric estimation for causal mediation analysis with multiple causally ordered mediators. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (3):794–821, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2022:FDR

[3729] Ruodu Wang and Aaditya Ramdas. False discovery rate control with evalues. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3):822–852, July 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhong:2022:EBP

[3730] Xinyi Zhong, Chang Su, and Zhou Fan. Empirical Bayes PCA in high dimensions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3):853–878, July 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Pawel:2022:SBF

[3731] Samuel Pawel and Leonhard Held. The sceptical Bayes factor for the assessment of replication success. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 879–911, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

She:2022:SML

[3732] Yiyuan She, Jiahui Shen, and Chao Zhang. Supervised multivariate learning with simultaneous feature autogrouping and dimension reduction. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (3):912–932, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Li:2022:FPM

[3733] Jialiang Li, Yaguang Li, and Tailen Hsing. On functional processes with multiple discontinuities. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3):933–972, July 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Biswas:2022:CBC

[3734] Niloy Biswas, Anirban Bhattacharya, Pierre E. Jacob, and James E. Johndrow. Coupling-based convergence assessment of some Gibbs samplers for high-dimensional Bayesian regression with shrinkage priors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 973–996, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Matsubara:2022:RGB

[3735] Takuo Matsubara, Jeremias Knoblauch, François-Xavier Briol, and Chris J. Oates. Robust generalised Bayesian inference for intractable likelihoods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (3):997–1022, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Follain:2022:HDC

[3736] Bertille Follain, Tengyao Wang, and Richard J. Samworth. Highdimensional changepoint estimation with heterogeneous missingness. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(3): 1023–1055, July 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2022:IId

[3737] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (4):1057–1058, September 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Riabiz:2022:OTM

[3738] Marina Riabiz, Wilson Ye Chen, Jon Cockayne, Pawel Swietach, Steven A. Niederer, Lester Mackey, and Chris. J. Oates. Optimal thinning of MCMC output. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1059–1081, September 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Jewell:2022:TCM

[3739] Sean Jewell, Paul Fearnhead, and Daniela Witten. Testing for a change in mean after changepoint detection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1082–1104, September 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rosset:2022:OMP

[3740] Saharon Rosset, Ruth Heller, Amichai Painsky, and Ehud Aharoni. Optimal and maximin procedures for multiple testing problems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1105– 1128, September 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See correction [3990].

Li:2022:EMA

[3741] Didong Li, Minerva Mukhopadhyay, and David B. Dunson. Efficient manifold approximation with spherelets. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1129–1149, September 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2022:BIF

[3742] Zhonglei Wang, Liuhua Peng, and Jae Kwang Kim. Bootstrap inference for the finite population mean under complex sampling designs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1150– 1174, September 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhao:2022:SLC

[3743] Wei Zhao, Limin Peng, and John Hanfelt. Semiparametric latent class analysis of recurrent event data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1175– 1197, September 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Moran:2022:FIF

[3744] Kelly R. Moran and Matthew W. Wheeler. Fast increased fidelity samplers for approximate Bayesian Gaussian process regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1198– 1228, September 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Graham:2022:MMC

[3745] Matthew M. Graham, Alexandre H. Thiery, and Alexandros Beskos. Manifold Markov chain Monte Carlo methods for Bayesian inference in diffusion models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1229–1256, September 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Tang:2022:BIR

[3746] Rong Tang and Yun Yang. Bayesian inference for risk minimization via exponentially tilted empirical likelihood. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (4):1257–1286, September 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Kontoyiannis:2022:BCT

[3747] Ioannis Kontoyiannis, Lambros Mertzanis, Athina Panotopoulou, Ioannis Papageorgiou, and Maria Skoularidou. Bayesian context trees: Modelling and exact inference for discrete time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1287–1323, September 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Feng:2022:NTF

[3748] Oliver Y. Feng, Yining Chen, Qiyang Han, Raymond J. Carroll, and Richard J. Samworth. Nonparametric, tuning-free estimation of Sshaped functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1324–1352, September 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Gronsbell:2022:EEP

[3749] Jessica Gronsbell, Molei Liu, Lu Tian, and Tianxi Cai. Efficient evaluation of prediction rules in semi-supervised settings under stratified sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (4):1353–1391, September 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

deFondeville:2022:FPT

[3750] Raphaël de Fondeville and Anthony C.

Davison. Functional peaks-over-threshold analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1392–1422, September 2022. CODEN JSTBAJ.

ISSN 1369-7412 (print), 1467-9868 (electronic).

Jiang:2022:MRE

[3751] Zhichao Jiang, Shu Yang, and Peng Ding. Multiply robust estimation of causal effects under principal ignorability. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1423–1445, September 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Rubin-Delanchy:2022:SIS

[3752] Patrick Rubin-Delanchy, Joshua Cape, Minh Tang, and Carey E. Priebe. A statistical interpretation of spectral embedding: The generalised random dot product graph. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1446– 1473, September 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Moscovich:2022:CVB

[3753] Amit Moscovich and Saharon Rosset. On the cross-validation bias due to unsupervised preprocessing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1474– 1502, September 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Wang:2022:PPP

[3754] Yudong Wang, Yanlin Tang, and Zhi-Sheng Ye. Paired or partially paired two-sample tests with unordered samples. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1503–1525, September 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Guillaumin:2022:DSW

[3755] Arthur P. Guillaumin, Adam M. Sykulski, Sofia C. Olhede, and Frederik J. Simons. The Debiased Spatial Whittle likelihood. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1526–1557, September 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Liang:2022:UPB

[3756] Tengyuan Liang. Universal prediction band via semi-definite programming. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1558–1580, September 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Bladt:2022:CSM

[3757] Mogens Bladt, Samuel Finch, and Michael Sørensen. Corrigendum to ‘Simulation of multivariate diffusion bridges’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(4):1581–1585, September 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). See [3351].

Anonymous:2022:IIe

[3758] Anonymous. Issue information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (5):1587–1588, November 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Jiang:2022:MCI

[3759] Feiyu Jiang, Zifeng Zhao, and Xiaofeng Shao. Modelling the COVID-19 infection trajectory: a piecewise linear quantile trend model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):1589– 1607, November 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Walther:2022:CSS

[3760] Guenther Walther and Andrew Perry. Calibrating the scan statistic: Finite sample performance versus asymptotics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):1608–1639, November 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Jewson:2022:GBL

[3761] Jack Jewson and David Rossell. General Bayesian loss function selection and the use of improper models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (5):1640–1665, November 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Han:2022:ECT

[3762] Rungang Han, Yuetian Luo, Miaoyan Wang, and Anru R. Zhang. Exact clustering in tensor block model: Statistical optimality and computational limit. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):1666–1698, November 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhao:2022:STS

[3763] Zifeng Zhao, Feiyu Jiang, and Xiaofeng Shao. Segmenting time series via self-normalisation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):1699– 1725, November 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Karmakar:2022:AAB

[3764] Bikram Karmakar. An approximation algorithm for blocking of an experimental design. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):1726– 1750, November 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhou:2022:DFM

[3765] Quan Zhou, Jun Yang, Dootika Vats, Gareth O. Roberts, and Jeffrey S. Rosenthal. Dimension-free mixing for high-dimensional Bayesian variable selection. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):1751–1784, November 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Sarkar:2022:CCN

[3766] Soham Sarkar and Victor M. Panaretos. CovNet: Covariance networks for functional data on multidimensional domains. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):1785–1820, November 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Lundborg:2022:CIT

[3767] Anton Rask Lundborg, Rajen D. Shah, and Jonas Peters. Conditional independence testing in Hilbert spaces with applications to functional data analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):1821–1850, November 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Le:2022:LRI

[3768] Can M. Le and Tianxi Li. Linear regression and its inference on noisy network-linked data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):1851– 1885, November 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Leung:2022:ZVA

[3769] Dennis Leung and Wenguang Sun. ZAP: Z-value adaptive procedures for false discovery rate control with side information. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):1886–1946, November 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Chang:2022:ELB

[3770] Hsin wen Chang and Ian W. McKeague. Empirical likelihood-based inference for functional means with application to wearable device data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84 (5):1947–1968, November 2022. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Papadogeorgou:2022:CIS

[3771] Georgia Papadogeorgou, Kosuke Imai, Jason Lyall, and Fan Li. Causal inference with spatio-temporal data: Estimating the effects of airstrikes on insurgent violence in Iraq. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):1969– 1999, November 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhu:2022:HDP

[3772] Ziwei Zhu, Tengyao Wang, and Richard J. Samworth. High-dimensional principal component analysis with heterogeneous missingness. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):2000– 2031, November 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

VanderWeele:2022:STR

[3773] Tyler J. VanderWeele and Stijn Vansteelandt. A statistical test to reject the structural interpretation of a latent factor model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):2032– 2054, November 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Engelke:2022:SLE

[3774] Sebastian Engelke and Stanislav Volgushev. Structure learning for extremal tree models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):2055– 2087, November 2022. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Anonymous:2022:CV

[3775] Anonymous. Contents of volume 84, 2022. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 84(5):2088–2089, November 2022. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic).

Zhao:2023:CAM

[3776] Anqi Zhao and Peng Ding. Covariate adjustment in multiarmed, possibly factorial experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(1): 1–23, February 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/1/1/7008637.

Candes:2023:CSA

[3777] Emmanuel Candès, Lihua Lei, and Zhimei Ren. Conformalized survival analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(1):24–45, February 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL https://academic.
                                        oup.com/jrsssb/article/85/1/24/
                                        7008653.

Sell:2023:TCG

[3778] Torben Sell and Sumeetpal Sidhu Singh. Trace-class Gaussian priors for Bayesian learning of neural networks with MCMC. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(1):46–66, February 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL https://academic.

oup.com/jrsssb/article/85/1/46/
7017762.

Liu:2023:BSE

[3779] Yukun Liu and Yan Fan. Biasedsample empirical likelihood weighting for missing data problems: an alternative to inverse probability weighting. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(1):67–83, February 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/1/67/7008591.

Dai:2023:BFS

[3780] Hongsheng Dai, Murray Pollock, and Gareth O. Roberts. Bayesian fusion: scalable unification of distributed statistical analyses. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(1):84–107, February 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL https://academic.
oup.com/jrsssb/article/85/1/84/
7008983.

Miao:2023:ICI

[3781] Ruizhong Miao and Tianxi Li. Informative core identification in complex networks. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(1):108–126, February 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL

https://academic.oup.com/jrsssb/
article/85/1/108/7008469.

Chang:2023:MMT

[3782] Jinyuan Chang, Jing He, Lin Yang, and Qiwei Yao. Modelling matrix time series via a tensor CP-decomposition.

Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(1):127–148, February 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/1/127/7008470.

Battey:2023:IHD

[3783] Heather S. Battey and Nancy Reid. On inference in high-dimensional regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(1):149–175, February 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL

https://academic.oup.com/jrsssb/
article/85/1/149/7018000.

Goncalves:2023:CEB

[3784] Flávio B. Gonçalves and Dani Gamerman. Corrigendum: Exact Bayesian inference in spatiotemporal Cox processes driven by multivariate Gaussian processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(1):176, February 2023. CODEN JSTBAJ. ISSN 1369-7412w (print), 1467-9868 (elec-

tronic). URL https://academic.
oup.com/jrsssb/article/85/1/176/

6998495. See [3469].

Anonymous:2023:EUP

[3785] Anonymous. Erratum: Usable and precise asymptotics for generalized linear mixed model analysis and design. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(1):177, February 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/1/177/6998488. See [3678].

Cui:2023:EHT

[3786] Yifan Cui, Michael R. Kosorok, Erik Sverdrup, Stefan Wager, and Ruoqing Zhu. Estimating heterogeneous treatment effects with right-censored data via causal survival forests. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2): 179–211, April 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/2/179/7058918.

Lahiri:2023:NER

[3787] Partha Lahiri and Nicola Salvati. A nested error regression model with high-dimensional parameter for small area estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2):212–239, April 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL https://academic.
oup.com/jrsssb/article/85/2/212/
7033452.

Lubold:2023:ILS

[3788] Shane Lubold, Arun G. Chandrasekhar, and Tyler H. McCormick. Identifying the latent space geometry of network models through analysis of curvature. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2):240–292, April 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL https://academic.
oup.com/jrsssb/article/85/2/240/
7070598.

Fraiman:2023:QHT

[3789] Ricardo Fraiman, Leonardo Moreno, and Thomas Ransford. A quantitative Heppes theorem and multivariate Bernoulli distributions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2): 293–314, April 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/2/293/7050006.

Lanteri:2023:DDH

[3790] Alessandro Lanteri, Samantha Leorato, Jesús López-Fidalgo, and Chiara Tommasi. Designing to detect heteroscedasticity in a regression model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2):315–326, April 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/2/315/7056753.

Zhang:2023:DTR

[3791] Yunyi Zhang and Dimitris N. Politis. Debiased and thresholded ridge regression for linear models with heteroskedastic and correlated errors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2):327–355, April 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/2/327/7070599.

Li:2023:EEG

[3792] Xiudi Li, Sijia Li, and Alex Luedtke. Estimating the efficiency gain of covariate-adjusted analyses in future clinical trials using external data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2):

356–377, April 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/2/356/7076962.

Lin:2023:CID

[3793] Zhenhua Lin, Dehan Kong, and Linbo Wang. Causal inference on distribution functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2):378–398, April 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL https://academic.
oup.com/jrsssb/article/85/2/378/
7082451.

Gu:2023:BPI

[3794] Yuqi Gu and David B. Dunson.

Bayesian Pyramids: identifiable multilayer discrete latent structure models for discrete data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2):399–426, April 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL https://academic.
oup.com/jrsssb/article/85/2/399/
7074362.

Chen:2023:EID

[3795] Shuxiao Chen and Bo Zhang. Estimating and improving dynamic treatment regimes with a time-varying instrumental variable. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2):427–453, April 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL https://academic.
oup.com/jrsssb/article/85/2/427/
7087213.

McElroy:2023:MIT

[3796] Tucker S. McElroy and Anindya Roy. Model identification via total Frobenius norm of multivariate spectra. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2): 454–473, April 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/2/454/7098033.

Huang:2023:NEC

[3797] Wei Huang and Zheng Zhang. Nonparametric estimation of the continuous treatment effect with measurement error. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2):474–496, April 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/2/474/7098034.

Cheung:2023:MRS

[3798] Ying Kuen Cheung and Keith M. Diaz. Monotone response surface of multi-factor condition: estimation and Bayes classifiers. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2):497–522, April 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL https://academic.
                                        oup.com/jrsssb/article/85/2/497/
                                        7083541.

Zhou:2023:SIH

[3799] Zhou Zhou and Holger Dette. Statistical inference for high-dimensional panel functional time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(2):

523–549, April 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/2/523/7099701.

Yu:2023:BLB

[3800] Weichang Yu and Howard D. Bondell. Bayesian likelihood-based regression for estimation of optimal dynamic treatment regimes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3): 551–574, July 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/3/551/7092905.

Yang:2023:EIA

[3801] Shu Yang, Chenyin Gao, Donglin Zeng, and Xiaofei Wang. Elastic integrative analysis of randomised trial and real-world data for treatment heterogeneity estimation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):575–596, July 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL https://academic.
oup.com/jrsssb/article/85/3/575/
7109328.

Thams:2023:STU

[3802] Nikolaj Thams, Sorawit Saengkyongam, Niklas Pfister, and Jonas Peters. Statistical testing under distributional shifts. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):597–663, July 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL https://academic.

                                        oup.com/jrsssb/article/85/3/597/
                                        7135919.

Vesely:2023:PBT

[3803] Anna Vesely, Livio Finos, and Jelle J. Goeman. Permutation-based true discovery guarantee by sum tests. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3): 664–683, July 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/3/664/7099913.

Ying:2023:PCI

[3804] Andrew Ying, Wang Miao, Xu Shi, and Eric J. Tchetgen Tchetgen. Proximal causal inference for complex longitudinal studies. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):684–704, July 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL https://academic.
                                        oup.com/jrsssb/article/85/3/684/
                                        7094061.

Zhu:2023:THT

[3805] Changbo Zhu and Jane-Ling Wang. Testing homogeneity: the trouble with sparse functional data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):705– 731, July 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868

                                        (electronic). URL https://academic.
                                        oup.com/jrsssb/article/85/3/705/
                                        7099972.

Goncalves:2023:EMC

[3806] Flávio B. Gonçalves, Krzysztof Łatuszyński, and Gareth O. Roberts. Exact Monte Carlo likelihood-based inference for jump-diffusion processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):732– 756, July 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868

(electronic). URL https://academic.
oup.com/jrsssb/article/85/3/732/
7101080.

Au:2023:MLS

[3807] Khai Xiang Au, Matthew M. Graham, and Alexandre H. Thiery. Manifold lifting: scaling Markov chain Monte Carlo to the vanishing noise regime. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):757–782, July 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/3/757/7137399.

Wu:2023:SKP

[3808] Sanyou Wu and Long Feng. Sparse Kronecker product decomposition: a general framework of signal region detection in image regression. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3): 783–809, July 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/3/783/7145706.

Bennett:2023:VMM

[3809] Andrew Bennett and Nathan Kallus.

The variational method of moments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):810–841, July 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/3/810/7146121.

Liebl:2023:FFS

[3810] Dominik Liebl and Matthew Reimherr. Fast and fair simultaneous confidence bands for functional parameters. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3): 842–868, July 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/3/842/7133768.

Yang:2023:OFE

[3811] Liuqing Yang, Yongdao Zhou, and Min-Qian Liu. Ordering factorial experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):869–885, July 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

                                        tronic). URL https://academic.
                                        oup.com/jrsssb/article/85/3/869/

7133803. See correction [3818].

Berenguer-Rico:2023:MWL

[3812] Vanessa Berenguer-Rico, Søren Johansen, and Bent Nielsen. A model where the least trimmed squares estimator is maximum likelihood. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3): 886–912, July 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/3/886/7162505.

Li:2023:NPI

[3813] Wei Li, Wang Miao, and Eric Tchetgen Tchetgen. Non-parametric inference about mean functionals of nonignorable non-response data without identifying the joint distribution. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3): 913–935, July 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/3/913/7156590.

Wang:2023:CED

[3814] Guanghui Wang and Long Feng.

Computationally efficient and dataadaptive changepoint inference in high dimension. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):936–958, July 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL https://academic.
oup.com/jrsssb/article/85/3/936/
7172766.

Guo:2023:CII

[3815] Zijian Guo. Causal inference with invalid instruments: post-selection problems and a solution using searching and sampling. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):959–985, July 2023. CODEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (elec-

tronic). URL https://academic.
oup.com/jrsssb/article/85/3/959/
7174915.

Kanagawa:2023:KST

[3816] Heishiro Kanagawa, Wittawat Jitkrittum, Lester Mackey, Kenji Fukumizu, and Arthur Gretton. A kernel Stein test for comparing latent variable models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):986–1011, July 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/3/986/7153413.

Zhu:2023:AOT

[3817] Changbo Zhu and Hans-Georg Müller. Autoregressive optimal transport models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):1012–1033, July 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/3/1012/7160826. See correction

[3819].

Anonymous:2023:COF

[3818] Anonymous. Correction to: Ordering factorial experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (3):1034, July 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/3/1034/7165453. See [3811].

Anonymous:2023:CAO

[3819] Anonymous. Correction to: Autoregressive optimal transport models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(3):1035, July 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/3/1035/7187238. See [3817].

Rohe:2023:VFA

[3820] Karl Rohe and Muzhe Zeng. Vintage factor analysis with Varimax performs statistical inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (4):1037–1060, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1037/7221295.

Hoff:2023:PVT

[3821] Peter Hoff. Proposer of the vote of thanks to Rohe & Zeng and contribution to the Discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (4):1061–1062, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1061/7103520.

Pensky:2023:SVT

[3822] Marianna Pensky. Seconder of the vote of thanks to Rohe & Zeng and contribution to the discussion of ‘Vintage factor analysis with varimax performs statistical inference’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (4):1062–1066, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1062/7110043.

Cape:2023:JCC

[3823] Joshua Cape. Joshua Cape’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1066–1067, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/4/1066/7107467.

Wang:2023:KWC

[3824] Kaizheng Wang. Kaizheng Wang’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1068, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/4/1068/7109146.

Han:2023:RHA

[3825] Rungang Han and Anru R. Zhang. Rungang Han and Anru R. Zhang’s contribution to the discussion of ‘Vintage factor analysis with varimax performs statistical inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1069–1070, September 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/4/1069/7107492.

VanWerde:2023:AVW

[3826] Alexander Van Werde. Alexander Van Werde’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (4):1070–1071, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1070/7179560.

He:2023:YYG

[3827] Yinqiu He, Yuqi Gu, and Zhiliang Ying. Yinqiu He, Yuqi Gu and Zhilian Ying’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (4):1071–1074, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1071/7103617.

Bickel:2023:PJB

[3828] Peter J. Bickel, Derek Bean, Aiyou Chen, and Purnamrita Sarkar. Peter J. Bickel, Derek Bean, Aiyou Chen and Purnamrita Sarkar’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1074–1075, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1074/7115882.

Cai:2023:JCD

[3829] Junhui Cai, Dan Yang, Linda Zhao, and Wu Zhu. Junhui Cai, Dan Yang, Linda Zhao and Wu Zhu’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1076–1080, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/4/1076/7109147.

Chai:2023:CPC

[3830] Christine P. Chai. Christine P. Chai’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1081– 1082, September 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/4/1081/7103623.

Chen:2023:YCG

[3831] Yunxiao Chen and Gongjun Xu. Yunxiao Chen and Gongjun Xu’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1082–1084, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/4/1082/7103502.

Kumar:2023:KKC

[3832] Kuldeep Kumar. Kuldeep Kumar’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1084, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1084/7100682.

Liu:2023:YLC

[3833] Yang Liu. Yang Liu’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (4):1085–1086, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1085/7109148.

Niu:2023:XNC

[3834] Xiaoyue Niu. Xiaoyue Niu’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1086–1087, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1086/7107493.

Pargent:2023:FPD

[3835] Florian Pargent, David Goretzko, and Timo von Oertzen. Florian Pargent, David Goretzko and Timo von Oertzen’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1087– 1088, September 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1087/7176313.

Pilling:2023:MPC

[3836] Mark Pilling. Mark Pilling’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1089, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1089/7107494.

Siroki:2023:KSK

[3837] Konstantin Siroki and Korbinian Strimmer. Konstantin Siroki and Korbinian Strimmer’s contribution to the discussion of ‘Vintage factor analysis with varimax performs statistical inference’ by Rohe and Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (4):1089–1090, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1089/7176314.

VanderWeele:2023:TJV

[3838] Tyler J. VanderWeele. Tyler J. VanderWeele’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (4):1091–1092, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1091/7110042.

Wang:2023:TWC

[3839] Tao Wang. Tao Wang’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1092–1093, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1092/7109149.

Zhou:2023:YZX

[3840] Ying Zhou and Xinyi Zhang. Ying Zhou and Xinyi Zhang’s contribution to the discussion of ‘Vintage Factor Analysis with Varimax Performs Statistical Inference’ by Rohe & Zeng. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1093–1094, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1093/7188889.

Rohe:2023:KRM

[3841] Karl Rohe and Muzhe Zeng. Karl Rohe and Muzhe Zeng’s reply to the discussion of ‘vintage factor analysis with varimax performs statistical inference’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1094–1098, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1094/7250221.

Zhu:2023:QAC

[3842] Qianqian Zhu, Songhua Tan, Yao Zheng, and Guodong Li. Quantile autoregressive conditional heteroscedasticity. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1099–1127, September 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1099/7226215.

Goeman:2023:CEI

[3843] Jelle J. Goeman, Paweł Górecki, Ramin Monajemi, Xu Chen, Thomas E. Nichols, and Wouter Weeda. Cluster extent inference revisited: quantification and localisation of brain activity. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1128–1153, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1128/7218928.

Miles:2023:CIR

[3844] Caleb H. Miles. On the causal interpretation of randomised interventional indirect effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1154– 1172, September 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1154/7209706.

Whitehouse:2023:CFI

[3845] Michael Whitehouse, Nick Whiteley, and Lorenzo Rimella. Consistent and fast inference in compartmental models of epidemics using Poisson Approximate Likelihoods. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1173– 1203, September 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1173/7218445.

Zhou:2023:TMP

[3846] Yunzhe Zhou, Chengchun Shi, Lexin Li, and Qiwei Yao. Testing for the Markov property in time series via deep conditional generative learning. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (4):1204–1222, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1204/7206744.

He:2023:REI

[3847] Xuming He, Kean Ming Tan, and Wen-Xin Zhou. Robust estimation and inference for expected shortfall regression with many regressors. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (4):1223–1246, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1223/7198207.

Beraha:2023:NLM

[3848] Mario Beraha and Jim E. Griffin. Normalised latent measure factor models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1247–1270, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1247/7206349.

Chen:2023:STS

[3849] Zengjing Chen, Xiaodong Yan, and Guodong Zhang. Strategic two-sample test via the two-armed bandit process. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1271–1298, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/4/1271/7197921.

Xue:2023:TSE

[3850] Liugen Xue. Two-stage estimation and bias-corrected empirical likelihood in a partially linear single-index varying-coefficient model. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1299– 1325, September 2023. CODEN JST-BAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/4/1299/7208993.

Wu:2023:QNU

[3851] Shuyuan Wu, Danyang Huang, and Hansheng Wang. Quasi-Newton updating for large-scale distributed learning. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1326–1354, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

                                        academic.oup.com/jrsssb/article/
                                        85/4/1326/7193744.

Anonymous:2023:CSA

[3852] Anonymous. Correction to: Sensitivity analysis for inverse probability weighting estimators via the percentile bootstrap. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(4):1355, September 2023. CO-DEN JSTBAJ. ISSN 1369-7412 (print), 1467-9868 (electronic). URL https://

academic.oup.com/jrsssb/article/
85/4/1355/7237517. See [3546].

Fong:2023:MPD

[3853] Edwin Fong, Chris Holmes, and Stephen G. Walker. Martingale posterior distributions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1357– 1391, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1357/7597700.

Dawid:2023:PVT

[3854] Philip Dawid. Proposer of the vote of thanks to Fong, Holmes and Walker and contribution to the discussion of ‘Martingale posterior distributions’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1392–1393, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1392/7252508.

Lauritzen:2023:SVT

[3855] Steffen Lauritzen. Seconder of the vote of thanks to Fong, Holmes and Walker and contribution to the discussion of ‘Martingale posterior distributions’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1394–1396, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1394/7260050.

Moya:2023:BMC

[3856] Blake Moya. Blake Moya’s contribution to the discussion of ‘Martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1396–1397, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1396/7252930.

Antoniano-Villalobos:2023:IAV

[3857] Isadora Antoniano-Villalobos. Isadora Antoniano Villalobos’s contribution to the discussion of ‘Martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1398, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1398/7257480.

Mena:2023:RMC

[3858] Ramsés H. Mena. Ramses Mena Chavez’s contribution to the discussion of ‘Martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1399–1400, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1399/7252973.

Clarke:2023:BCC

[3859] Bertrand Clarke. Bertrand Clarke’s contribution to the discussion of ‘Martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (5):1400–1401, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1400/7252509.

Draper:2023:DDE

[3860] David Draper and Erdong Guo. David Draper and Erdong Guo’s contribution to the discussion of ‘martingale posterior distributions’, by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (5):1401–1402, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1401/7260051.

Gu:2023:JGG

[3861] Jiaqi Gu and Guosheng Yin. Jiaqi Gu and Guosheng Yin’s contribution to the discussion of ‘Martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1403–1404, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1403/7252929.

Ascolani:2023:FAA

[3862] Filippo Ascolani, Antonio Lijoi, and Igor Prünster. Filippo Ascolani, Antonio Lijoi, and Igor Prünster’s contribution to the discussion of ‘Martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):

1404–1405, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1404/7252510.

Huk:2023:DHL

[3863] David Huk, Lorenzo Pacchiardi, Ritabrata Dutta, and Mark Steel. David Huk, Lorenzo Pacchiardi, Ritabrata Dutta and Mark Steel’s contribution to the discussion of ‘martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1405–1406, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1405/7274077.

Catalano:2023:MCA

[3864] Marta Catalano, Augusto Fasano, and Giovanni Rebaudo. Marta Catalano, Augusto Fasano, and Giovanni Rebaudo’s contribution to the discussion of ‘martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (5):1406–1407, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1406/7252511.

Rigo:2023:PRC

[3865] Pietro Rigo. Pietro Rigo’s contribution to the discussion of ‘Martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1407–1408, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1407/7252513.

Rossell:2023:DRC

[3866] David Rossell. David Rossell’s contribution to the discussion of ‘Martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5): 1408–1409, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1408/7258419.

Swallow:2023:BSC

[3867] Ben Swallow. Ben Swallow’s contribution to the discussion of ‘Martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (5):1410, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1410/7252512.

Ray:2023:KRB

[3868] Kolyan Ray and Botond Szabó. Kolyan Ray and Botond Szabo’s contribution to the discussion of ‘Martingale posterior distributions’ by Fong, Holmes and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1410–1411, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1410/7257482.

Wijayatunga:2023:PWC

[3869] Priyantha Wijayatunga. Priyantha Wijayatunga’s contribution to the discussion of ‘Martingale posterior distributions’ by Fong, Holmes, and Walker. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1412, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1412/7252974.

Fong:2023:ARD

[3870] Edwin Fong, Chris Holmes, and Stephen G. Walker. Authors’ reply to the discussion of ‘Martingale posterior distributions’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (5):1413–1416, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1413/7597649.

Porcu:2023:SNS

[3871] Emilio Porcu, Philip A. White, and Marc G. Genton. Stationary nonseparable space-time covariance functions on networks. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1417– 1440, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1417/7263253.

Li:2023:MCI

[3872] Shuangning Li and Molei Liu. Maxway CRT: improving the robustness of the model-X inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5): 1441–1470, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1441/7243739.

Caughey:2023:RIB

[3873] Devin Caughey, Allan Dafoe, Xinran Li, and Luke Miratrix. Randomisation inference beyond the sharp null: bounded null hypotheses and quantiles of individual treatment effects. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (5):1471–1491, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1471/7246293.

Atchade:2023:FAM

[3874] Yves Atchadé and Liwei Wang. A fast asynchronous Markov chain Monte Carlo sampler for sparse Bayesian inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1492–1516, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1492/7272760.

Yuan:2023:TWD

[3875] Chaofeng Yuan, Zhigen Gao, Xuming He, Wei Huang, and Jianhua Guo. Two-way dynamic factor models for high-dimensional matrix-valued time series. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1517–1537, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1517/7249210.

Xu:2023:BIB

[3876] Maoran Xu and Leo L. Duan. Bayesian inference with the l_{1} -ball prior: solving combinatorial problems with exact zeros. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1538–1560, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1538/7233309.

Giessing:2023:DIH

[3877] Alexander Giessing and Jingshen Wang. Debiased inference on heterogeneous quantile treatment effects with regression rank scores. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85 (5):1561–1588, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1561/7239218.

Zhang:2023:IRR

[3878] Daiwei Zhang, Lexin Li, Chandra Sripada, and Jian Kang. Image response regression via deep neural networks. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1589–1614, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1589/7230181.

Athey:2023:SPE

[3879] Susan Athey, Peter J. Bickel, Aiyou Chen, Guido W. Imbens, and Michael Pollmann. Semi-parametric estimation of treatment effects in randomised experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5): 1615–1638, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1615/7226451.

Agrawal:2023:DNC

[3880] Raj Agrawal, Chandler Squires, Neha Prasad, and Caroline Uhler. The De-CAMFounder: nonlinear causal discovery in the presence of hidden variables. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1639–1658, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1639/7223510.

Hoff:2023:CSC

[3881] Peter Hoff, Andrew McCormack, and Anru R. Zhang. Core shrinkage covariance estimation for matrix-variate data. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1659–1679, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
85/5/1659/7227061.

Qiu:2023:PSA

[3882] Hongxiang Qiu, Edgar Dobriban, and Eric Tchetgen Tchetgen. Prediction sets adaptive to unknown covariate shift. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1680–1705, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1680/7225361.

Anonymous:2023:CSS

[3883] Anonymous. Correction to: Semisupervised approaches to efficient evaluation of model prediction performance. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1706, November 2023. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1706/7274030.

Anonymous:2023:CV

[3884] Anonymous. Contents of volume 85, 2023. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 85(5):1707–1709, November 2023. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        85/5/1707/7597645.

Waudby-Smith:2024:EMB

[3885] Ian Waudby-Smith and Aaditya Ramdas. Estimating means of bounded random variables by betting. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (1):1–27, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/1/1/7043257.

Grunwald:2024:PVT

[3886] Peter Grünwald. Proposer of the vote of thanks to Waudy-Smith and Ramdas and contribution to the discussion of ‘Estimating means of bounded random variables by betting’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1): 28–30, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://academic.
oup.com/jrsssb/article/86/1/28/
7394904.

Neu:2024:SVT

[3887] Gergely Neu. Seconder of the vote of thanks to Waudby-Smith and Ramdas and contribution to the discussion of ‘Estimating means of bounded random variables by betting’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1): 30–32, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://academic.
oup.com/jrsssb/article/86/1/30/
7303214.

Wang:2024:RWC

[3888] Ruodu Wang. Ruodu Wang’s contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):32–33, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/1/32/7303976. See cor-

rection [3967].

Angelopoulos:2024:AAC

[3889] Anastasios N. Angelopoulos. Anastasios N. Angelopoulos’ contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas.

Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):33–34, February 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/1/33/7303197.

Davison:2024:ACD

[3890] Anthony C. Davison and Igor Rodionov. Anthony C. Davison and Igor Rodionov’s contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):35–36, February 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/1/35/7307584.

Howard:2024:SRH

[3891] Steven R. Howard. Steven R. Howard’s contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):36–38, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/1/36/7303978.

Jiang:2024:RJK

[3892] Rong Jiang and Keming Yu. Rong Jiang and Keming Yu’s contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):38–39, February 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/1/38/7303198.

Larsson:2024:MLJ

[3893] Martin Larsson and Johannes Ruf.

Martin Larsson and Johannes Ruf’s contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):39–40, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/1/39/7303215.

Li:2024:JLY

[3894] Jiayi Li, Yuantong Li, and Xiaowu Dai.

Jiayi Li, Yuantong Li and Xiaowu Dai’s contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):41–43, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/1/41/7303977.

Martin:2024:RMC

[3895] Ryan Martin. Ryan Martin’s contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):43–44, February 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/1/43/7301729.

Nguyen:2024:HNC

[3896] Hien Nguyen. Hien Nguyen’s contribution to the discussion of“Estimating means of bounded random variables by betting” by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):45–46, February 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/1/45/7303979.

Owen:2024:AOC

[3897] Art B. Owen. Art Owen’s contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):47–48, February 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/1/47/7319105.

Siegmund:2024:DSC

[3898] David Siegmund. David Siegmund’s contribution to the discussion of “Estimating means of bounded random variables by betting” by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):48–49, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

                                        http://academic.oup.com/jrsssb/
                                        article/86/1/48/7311100.

Stark:2024:PBS

[3899] Philip B. Stark. Philip B. Stark’s contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):49–51, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/1/49/7307585.

Thomas:2024:PTE

[3900] Philip S. Thomas, Erik Learned-Miller, and My Phan. Philip S. Thomas, Erik Learned-Miller and My Phan’s contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):51–52, February 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/1/51/7307586.

Vovk:2024:VVC

[3901] Vladimir Vovk. Vladimir Vovk’s contribution to the discussion of “Estimating means of bounded random variables by betting” by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):52–53, February 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/1/52/7301733.

Waudby-Smith:2024:ARD

[3902] Ian Waudby-Smith and Aaditya Ramdas. Authors’ reply to the discussion of ‘Estimating means of bounded random variables by betting’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1): 53–61, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://academic.
oup.com/jrsssb/article/86/1/53/
7321874.

Kosmidis:2024:EBR

[3903] Ioannis Kosmidis and Nicola Lunardon. Empirical bias-reducing adjustments to estimating functions. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1): 62–89, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://academic.
oup.com/jrsssb/article/86/1/62/
7275082.

Karwa:2024:MCG

[3904] Vishesh Karwa, Debdeep Pati, Sonja Petrović, Liam Solus, Nikita Alexeev, Mateja Raič, Dane Wilburne, Robert Williams, and Bowei Yan. Monte Carlo goodness-of-fit tests for degree corrected and related stochastic blockmodels. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):90–121, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/1/90/7274667.

Ren:2024:DKL

[3905] Zhimei Ren and Rina Foygel Barber. Derandomised knockoffs: leveraging evalues for false discovery rate control. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):122–154, February 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/1/122/7262479.

Zheng:2024:DSC

[3906] Xiangyu Zheng and Song Xi Chen.

Dynamic synthetic control method for evaluating treatment effects in autoregressive processes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):155– 176, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://academic.
oup.com/jrsssb/article/86/1/155/
7331057.

Maullin-Sapey:2024:SCR

[3907] Thomas Maullin-Sapey, Armin Schwartzman, and Thomas E. Nichols. Spatial confidence regions for combinations of excursion sets in image analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):177–193, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/1/177/7280081.

Moran:2024:HPC

[3908] Gemma E. Moran, David M. Blei, and Rajesh Ranganath. Holdout predictive checks for Bayesian model criticism. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):194–214, February 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/1/194/7274627.

Javanmard:2024:GGF

[3909] Adel Javanmard and Mohammad Mehrabi. GRASP: a goodness-of-fit test for classification learning. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (1):215–245, February 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/1/215/7281031.

Zhang:2024:ALB

[3910] Yi Zhang and Xiaofeng Shao. Another look at bandwidth-free inference: a sample splitting approach. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(1):246–272, February 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/1/246/7325862.

Li:2024:ACP

[3911] Jie Li, Paul Fearnhead, Piotr Fryzlewicz, and Tengyao Wang. Automatic change-point detection in time series via deep learning. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2): 273–285, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

                                        (electronic). URL http://academic.
                                        oup.com/jrsssb/article/86/2/273/
                                        7517020.

Benton:2024:DDD

[3912] Joe Benton, Yuyang Shi, Valentin De Bortoli, George Deligiannidis, and Arnaud Doucet. From denoising diffusions to denoising Markov models.

Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):286–301, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/286/7564909.

Wilkinson:2024:PVT

[3913] Darren Wilkinson. Proposer of the vote of thanks and contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):302–304, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/2/302/7486694.

Nemeth:2024:SVT

[3914] Christopher Nemeth. Seconder of the vote of thanks and contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):304–306, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/2/304/7504813.

vanLieshout:2024:MMV

[3915] Marie-Colette van Lieshout and Changqing Lu. M. N. M. van Lieshout and C. Lu’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):306–307, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

                                        http://academic.oup.com/jrsssb/
                                        article/86/2/306/7512678.

MacKenzie:2024:GMC

[3916] Gilbert MacKenzie. Gilbert MacKenzie’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):307–308, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

                                        http://academic.oup.com/jrsssb/
                                        article/86/2/307/7499133.

Zhang:2024:BZC

[3917] Bo Zhang. Bo Zhang’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):309–310, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/2/309/7490827.

Wang:2024:AWC

[3918] Andi Q. Wang. Andi Wang’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):310–311, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/2/310/7490811.

Mardia:2024:KMC

[3919] Kanti V. Mardia. Kanti Mardia’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):311–312, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/311/7490812.

Gavioli-Akilagun:2024:SGA

[3920] Shakeel Gavioli-Akilagun. Shakeel Gavioli-Akilagun’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):312–313, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/2/312/7490810.

Anastasiou:2024:ICA

[3921] Andreas Anastasiou and Ivor Cribben.

Ivor Cribben and Anastasiou Andreas’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):313–314, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/313/7503832.

Bodenham:2024:DBN

[3922] Dean Bodenham and Niall Adams.

Dean Bodenham and Niall Adams’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):315, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/315/7491619.

Chen:2024:YCY

[3923] Yudong Chen and Yining Chen. Yudong Chen and Yining Chen’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):316–317, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/316/7499131.

Cialfi:2024:DCC

[3924] Daniela Cialfi. Daniela Cialfi’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):318, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/318/7490828.

Gilliot:2024:PAG

[3925] Pierre-Aurelien Gilliot, Christophe Andrieu, Anthony Lee, Song Liu, and Michael Whitehouse. Pierre-aurelien gilliot, christophe andrieu, anthony Lee, song Liu, and Michael Whitehouse’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):318–320, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/2/318/7499373.

Hong:2024:YHO

[3926] Yongmiao Hong, Oliver Linton, Jiajing Sun, and Meiting Zhu. Yongmiao Hong, Oliver Linton, Jiajing Sun, and Meiting Zhu’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):320–321, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/2/320/7490813.

Jackson:2024:JJC

[3927] James Jackson. James Jackson’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):322, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/322/7511656.

Jungbluth:2024:AJJ

[3928] Ayla Jungbluth and Johannes Lederer. Ayla Jungbluth and Johannes Lederer’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):322–323, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

                                        http://academic.oup.com/jrsssb/
                                        article/86/2/322/7512680.

Kent:2024:JKC

[3929] John T. Kent. John Kent’s contribution to the discussion of the ‘discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):324, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/2/324/7486693.

Mateu:2024:JMC

[3930] Jorge Mateu. Jorge Mateu’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):325–326, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/2/325/7486691.

Ombao:2024:HOC

[3931] Hernando Ombao. Hernando Ombao’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):326, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/326/7503312.

Papp:2024:TPP

[3932] Tamás P. Papp, Paul Fearnhead, and Chris Sherlock. Tamás P. Papp, Paul Fearnhead, and Chris Sherlock’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):327–328, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/327/7564910.

Power:2024:SPC

[3933] Sam Power. Sam Power’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):328, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/328/7486692.

Schmidt-Hieber:2024:JSH

[3934] Johannes Schmidt-Hieber. Johannes Schmidt-Hieber’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):329, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/2/329/7517788.

Schoenberg:2024:FSW

[3935] Frederic Schoenberg and Weng Kee Wong. Frederic Schoenberg and Weng Kee Wong’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):330, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/2/330/7491620.

Wu:2024:YWW

[3936] Ying Nian Wu and Weng Kee Wong. Yingnian Wu and Weng Kee Wong’s contribution to the discussion of ‘the discussion meeting on probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):331, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/2/331/7502876.

Li:2024:ARD

[3937] Jie Li, Paul Fearnhead, Piotr Fryzlewicz, and Tengyao Wang. Authors’ reply to the discussion of ‘automatic change-point detection in time series via deep learning’ at the discussion meeting on ‘probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):332–334, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/332/7591224.

Benton:2024:ARD

[3938] Joe Benton, Yuyang Shi, Valentin De Bortoli, George Deligiannidis, and Arnaud Doucet. Authors’ reply to the discussion of ‘from denoising diffusions to denoising Markov models’ at the discussion meeting on ‘probabilistic and statistical aspects of machine learning’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):335–339, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/335/7600575.

Tallman:2024:BPD

[3939] Emily Tallman and Mike West.

Bayesian predictive decision synthesis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):340–363, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/340/7329254.

Tian:2024:SPE

[3940] Ye Tian and Hongquan Xu. Stratification pattern enumerator and its applications. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):364–385, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/364/7420317.

Cao:2024:CFD

[3941] Yang Cao, Xinwei Sun, and Yuan Yao. Controlling the false discovery rate in transformational sparsity: Split knockoffs. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):386–410, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/2/386/7419884.

He:2024:ABT

[3942] Yinqiu He, Peter X. K. Song, and Gongjun Xu. Adaptive bootstrap tests for composite null hypotheses in the mediation pathway analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):411–434, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/2/411/7420463.

Diaz:2024:NAI

[3943] Iván Díaz. Non-agency interventions for causal mediation in the presence of intermediate confounding. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2): 435–460, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

                                        (electronic). URL http://academic.
                                        oup.com/jrsssb/article/86/2/435/
                                        7445021.

Liu:2024:EMT

[3944] Yaowu Liu, Zhonghua Liu, and Xihong Lin. Ensemble methods for testing a global null. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):461–486, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/2/461/7456444.

Egami:2024:IEC

[3945] Naoki Egami and Eric J. Tchetgen Tchetgen. Identification and estimation of causal peer effects using double negative controls for unmeasured network confounding. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):487–511, April 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/2/487/7475696.

Mukherjee:2024:LSE

[3946] Somabha Mukherjee, Rohit K. Patra, Andrew L. Johnson, and Hiroshi Morita. Least squares estimation of a quasiconvex regression function. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(2):512–534, April 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/2/512/7470765.

Evans:2024:PSC

[3947] Robin J. Evans and Vanessa Didelez.

Parameterizing and simulating from causal models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):535–568, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/3/535/7177899. See dis-

cussion [3948, 3949, 3950, 3951, 3952, 3953, 3954, 3955, 3956] and reply

Seaman:2024:PVT

[3948] Shaun R. Seaman. Proposer of the vote of thanks to Evans and Didelez and contribution to the discussion of ‘Parameterizing and simulating from causal models’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):569–570, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

                                        http://academic.oup.com/jrsssb/
                                        article/86/3/569/7609931.  See

[3947].

Silva:2024:SVT

[3949] Ricardo Silva. Seconder of the vote of thanks to Evans and Didelez and contribution to the discussion of ‘Parameterizing and simulating from causal models’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):571–572, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

                                        http://academic.oup.com/jrsssb/
                                        article/86/3/571/7609152.  See

[3947].

Guo:2024:RGC

[3950] F. Richard Guo. Richard Guo’s contribution to the discussion of ‘Parameterizing and simulating from causal models’ by Evans and Didelez. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3): 572–574, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

                                        (electronic). URL http://academic.
                                        oup.com/jrsssb/article/86/3/572/

7608762. See [3947].

Bhansali:2024:RBC

[3951] Rajendra Bhansali. Rajendra Bhansali’s contribution to the discussion of ‘Parameterizing and simulating from causal models’ by Evans and Didelez. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):575, July 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/3/575/7609153. See [3947].

Battey:2024:HBC

[3952] Heather S. Battey. Heather Battey’s contribution to the discussion of ‘Parameterizing and simulating from causal models’ by Evans and Didelez. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):575–576, July 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/3/575/7608817. See [3947].

Dawid:2024:PDC

[3953] A. Philip Dawid. A. Philip Dawid’s contribution to the discussion of ‘Parameterizing and simulating from causal models’ by Evans and Didelez. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):576–577, July 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/3/576/7608747. See [3947].

Martinussen:2024:TMC

[3954] Torben Martinussen. Torben Martinussen’s contribution to the discussion of ‘Parameterizing and simulating from causal models’ by Evans and Didelez. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):577–578, July 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/3/577/7608748. See [3947].

Richardson:2024:TRJ

[3955] Thomas S. Richardson and James M. Robins. Thomas S. Richardson and James M. Robins’ contribution to the discussion of ‘Parameterizing and simulating from causal models’ by Evans and Didelez. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):578–580, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/3/578/7613058.  See

[3947].

Steiner:2024:GSM

[3956] Gregor Steiner and Mark F. J. Steel. Gregor Steiner and Mark Steel’s contribution to the discussion of ‘Parameterizing and simulating from causal models’ by Evans and Didelez. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3): 580–582, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://academic.
oup.com/jrsssb/article/86/3/580/

7609154. See [3947].

Evans:2024:ARD

[3957] Robin J. Evans and Vanessa Didelez. Authors’ reply to the discussion of ‘Parameterizing and simulating from causal models’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):582–585, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/3/582/7694339.  See

[3947].

Kuchibhotla:2024:HCR

[3958] Arun Kumar Kuchibhotla, Sivaraman Balakrishnan, and Larry Wasserman. The HulC: confidence regions from convex hulls. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):586–622, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/3/586/7499155.

Luo:2024:PET

[3959] Shikai Luo, Ying Yang, Chengchun Shi, Fang Yao, Jieping Ye, and Hongtu Zhu. Policy evaluation for temporal and/or spatial dependent experiments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):623–649, July 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/3/623/7511800.

Heller:2024:SDI

[3960] Ruth Heller and Aldo Solari. Simultaneous directional inference. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (3):650–670, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://academic.
oup.com/jrsssb/article/86/3/650/
7499129.

Liang:2024:ICV

[3961] Ziyi Liang, Matteo Sesia, and Wenguang Sun. Integrative conformal p-values for out-of-distribution testing with labelled outliers. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3): 671–693, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

                                        (electronic). URL http://academic.
                                        oup.com/jrsssb/article/86/3/671/
                                        7515027.

Chen:2024:GSM

[3962] Yaqing Chen, Shu-Chin Lin, Yang Zhou, Owen Carmichael, Hans-Georg / Müller, and Jane-Ling Wang. Gradient synchronization for multivariate functional data, with application to brain connectivity. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):694–713, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

                                        http://academic.oup.com/jrsssb/
                                        article/86/3/694/7584956.

Xu:2024:OIT

[3963] Qi Xu, Haoda Fu, and Annie Qu. Optimal individualized treatment rule for combination treatments under budget constraints. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):714–741, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

                                        http://academic.oup.com/jrsssb/
                                        article/86/3/714/7514594.

Zhang:2024:BMS

[3964] Yangfan Zhang and Yun Yang. Bayesian model selection via meanfield variational approximation. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3): 742–770, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://academic.
oup.com/jrsssb/article/86/3/742/
7642825.

Tran:2024:EDT

[3965] Ngoc Mai Tran, Johannes Buck, and Claudia / Klüppelberg. Estimating a directed tree for extremes. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (3):771–792, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://academic.
oup.com/jrsssb/article/86/3/771/
7606808.

Chen:2024:SPT

[3966] Elynn Y Chen, Dong Xia, Chencheng Cai, and Jianqing Fan. Semiparametric tensor factor analysis by iteratively projected singular value decomposition. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):793–823, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/3/793/7600270.

Anonymous:2024:CRW

[3967] Anonymous. Correction to: Ruodu Wang’s contribution to the discussion of ‘Estimating means of bounded random variables by betting’ by Waudby-Smith and Ramdas. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(3):824, July 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

                                        http://academic.oup.com/jrsssb/
                                        article/86/3/824/7628008.  See

[3888].

Crane:2024:RCI

[3968] Harry Crane and Min Xu. Root and community inference on the latent growth process of a network. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (4):825–865, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/4/825/7283069. See discussion

[3969, 3970, 3971, 3972, 3973, 3974, 3975, 3976, 3977, 3978, 3979, 3980] and reply [3981].

Jog:2024:PVT

[3969] Varun Jog and Po-Ling Loh. Proposers of the vote of thanks to Crane and Xu and contribution to the discussion of ‘Root and community inference on the latent growth process of a network’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):866–867, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/4/866/7687226. See [3968].

Rubin-Delanchy:2024:SVT

[3970] Patrick Rubin-Delanchy. Seconder of the vote of thanks to Crane and Xu and contribution to the discussion of ‘Root and community inference on the latent growth process of a network’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):867–870, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/867/7700654. See [3968].

Srakar:2024:ASC

[3971] Andrej Srakar. Andrej Srakar’s contribution to the discussion of ‘Root and community inference on the latent growth process of a network’ by Crane and Xu. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):870–871, September 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/870/7689243. See [3968].

Ascolani:2024:FAA

[3972] Filippo Ascolani, Antonio Lijoi, and Igor / Prünster. Filippo Ascolani, Antonio Lijoi and Igor Prünster’s contribution to the discussion of ‘Root and community inference on the latent growth process of a network’ by Crane and Xu. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):871–872, September 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/871/7687039. See [3968].

Banerjee:2024:SBC

[3973] Sayan Banerjee. Sayan Banerjee’s contribution to the discussion of ‘Root and community inference on the latent growth process of a network’ by Crane and Xu. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):873–874, September 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/873/7687041. See [3968].

Catalano:2024:MCA

[3974] Marta Catalano, Augusto Fasano, Matteo Giordano, and Giovanni Rebaudo. Marta Catalano, Augusto Fasano, Matteo Giordano, and Giovanni Rebaudo’s contribution to the discussion of ‘Root and community inference on the latent growth process of a network’ by Crane and Xu. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):874–875, September 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/874/7688442. See [3968].

Feng:2024:YFJ

[3975] Yang Feng and Jiajin Sun. Yang Feng and Jiajin Sun’s contribution to the discussion of ‘Root and community inference on the latent growth process of a network’ by Crane and Xu. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (4):875–878, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/875/7691835. See [3968].

Jiang:2024:YJZ

[3976] Yicong Jiang and Zheng Tracy Ke. Yicong Jiang and Zheng Tracy Ke’s contribution to the discussion of ‘Root and community inference on the latent growth process of a network’ by Crane and Xu. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):878–880, September 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/878/7687042. See [3968].

Li:2024:TLC

[3977] Tianxi Li. Tianxi Li’s contribution to the discussion of ‘Root and community inference on the latent growth process of a network’ by Crane and Xu. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):880–881, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/880/7687040. See [3968].

Yang:2024:QYX

[3978] Qing Yang and Xin Tong. Qing Yang and Xin Tong’s contribution to the discussion of ‘Root and community inference on the latent growth process of a network’ by Crane and Xu. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (4):881–882, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/881/7687752. See [3968].

Wang:2024:FWY

[3979] Fan Wang, Yi Yu, and Alessandro Rinaldo. Fan Wang, Yi Yu and Alessandro Rinaldo’s contribution to the discussion of ‘Root and community inference on the latent growth process of a network’ by Crane and Xu. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (4):883–884, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/4/883/7689435. See [3968].

Wyse:2024:JWJ

[3980] Jason Wyse, James Ng, Arthur White, and Michael Fop. Jason Wyse, James Ng, Arthur White and Michael Fop’s contribution to the discussion of ‘Root and community inference on the latent growth process of a network’ by Crane and Xu. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):884–885, September 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/4/884/7687043. See [3968].

Crane:2024:ARD

[3981] Harry Crane and Min Xu. Authors’ reply to the discussion of ‘Root and community inference on the latent growth process of a network’. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (4):885–895, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/4/885/7689786. See [3968].

deChaumaray:2024:FME

[3982] Marie Du Roy de Chaumaray and Matthieu Marbac. Full-model estimation for non-parametric multivariate finite mixture models. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (4):896–921, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

                                        academic.oup.com/jrsssb/article/
                                        86/4/896/7593798.

Cai:2024:THD

[3983] T Tony Cai, Zheng T Ke, and Paxton Turner. Testing high-dimensional multinomials with applications to text analysis. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):922–942, September 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (elec-

tronic).  URL http://academic.
oup.com/jrsssb/article/86/4/922/
7615734.

Yang:2024:DRC

[3984] Yachong Yang, Arun Kumar Kuchibhotla, and Eric Tchetgen Tchetgen. Doubly robust calibration of prediction sets under covariate shift. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (4):943–965, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/943/7618755.

Li:2024:SFD

[3985] Jinzhou Li, Marloes H Maathuis, and Jelle J Goeman. Simultaneous false discovery proportion bounds via knockoffs and closed testing. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):966– 986, September 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868

(electronic). URL http://academic.
oup.com/jrsssb/article/86/4/966/
7618756.

Sturma:2024:TMC

[3986] Nils Sturma, Mathias Drton, and Dennis Leung. Testing many constraints in possibly irregular models using incomplete U-statistics. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (4):987–1012, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/987/7624096.

Lila:2024:IDA

[3987] Eardi Lila, Wenbo Zhang, Swati Rane Levendovszky, Michael W. Weiner, Paul Aisen, Michael W. Weiner, Paul Aisen, Ronald Petersen, Clifford R. Jack, William Jagust, John Q. Trojanowki, Arthur W. Toga, Laurel Beckett, Robert C. Green, Andrew J. Saykin, John C. Morris, Richard J. Perrin, Leslie M. Shaw, Zaven Khachaturian, Maria Carrillo, William Potter, Lisa Barnes, Marie Bernard, Carole Ho, John K. Hsiao, Jonathan Jackson, Eliezer Masliah, Donna Masterman, Ozioma Okonkwo, Richard Perrin, Laurie Ryan, Nina Silverberg, Adam Fleisher, Michael W. Weiner, Juliet Fockler, Cat Conti, Dallas Veitch, John Neuhaus, Chengshi Jin, Rachel Nosheny, Miriam Ashford, Derek Flenniken, Adrienne Kormos, Robert C. Green, Tom Montine, Cat Conti, Ronald Petersen, Paul Aisen, Michael Rafii, Rema Raman, Gustavo Jimenez, Michael Donohue, Devon Gessert, Jennifer Salazar, Caileigh Zimmerman, Yuliana Cabrera, Sarah Walter, Garrett Miller, Godfrey Coker, Taylor Clanton, Lindsey Hergesheimer, Stephanie Smith, Olusegun Adegoke, Payam Mahboubi, Shelley Moore, Jeremy Pizzola, Elizabeth Shaffer, Brittany Sloan, Laurel Beckett, Danielle Harvey, Michael Donohue, Clifford R. Jack, Arvin Forghanian-Arani, Bret Borowski, Chad Ward, Christopher Schwarz, David Jones, Jeff Gunter, Kejal Kantarci, Matthew Senjem, Prashanthi Vemuri, Robert Reid, Nick C. Fox, Ian Malone, Paul Thompson, Sophia I. Thomopoulos, Talia M. Nir, Neda Jahanshad, Charles DeCarli, Alexander Knaack, Evan Fletcher, Danielle Harvey, Duygu Tosun-Turgut, Stephanie Rossi Chen, Mark Choe, Karen Crawford, Paul A. Yushkevich, Sandhitsu Das, William Jagust, Robert A. Koeppe, Eric M. Reiman, Kewei Chen, Chet Mathis, Susan Landau, John C. Morris, Richard Perrin, Nigel J. Cairns, Erin Householder, Erin Franklin, Haley Bernhardt, Lisa Taylor-Reinwald, Leslie M. Shaw, John Q. Trojanowki, Magdalena Korecka, Michal Figurski, Arthur W. Toga, Karen Crawford, Scott Neu, Andrew J. Saykin, Kwangsik Nho, Shannon L. Risacher, Liana G. Apostolova, Li Shen, Tatiana M. Foroud, Kelly Nudelman, Kelley Faber, Kristi Wilmes, Michael W. Weiner, Leon Thal, Zaven Khachaturian, John K. Hsiao, Lisa C. Silbert, Betty Lind, Rachel Crissey, Jeffrey A. Kaye, Raina Carter, Sara Dolen, Joseph Quinn, Lon S. Schneider, Sonia Pawluczyk, Mauricio Becerra, Liberty Teodoro, Karen Dagerman, Bryan M. Spann, James Brewer, Helen Vanderswag, Adam Fleisher, Jaimie Ziolkowski, Judith L. Heidebrink, Lisa Zbizek-Nulph, Joanne L. Lord, Lisa Zbizek-Nulph, Ronald Petersen, Sara S. Mason, Colleen S. Albers, David Knopman, Kris Johnson, Javier Villanueva-Meyer, Valory Pavlik, Nathaniel Pacini, Ashley Lamb, Joseph S. Kass, Rachelle S. Doody, Victoria Shibley, Munir Chowdhury, Susan Rountree, Mimi Dang, Yaakov Stern, Lawrence S. Honig, Akiva Mintz, Beau Ances, John C. Morris, David Winkfield, Maria Carroll, Georgia Stobbs-Cucchi, Angela Oliver, Mary L. Creech, Mark A. Mintun, Stacy Schneider, David Geldmacher, Marissa Natelson Love, Randall Griffith, David Clark, John Brockington, Daniel Marson, Hillel Grossman, Martin A. Goldstein, Jonathan Greenberg, Effie Mitsis, Raj C. Shah, Melissa Lamar, Patricia Samuels, Ranjan Duara, Maria T. Greig-Custo, Rosemarie Rodriguez, Marilyn Albert, Chiadi Onyike, Leonie Farrington, Scott Rudow, Rottislav Brichko, Stephanie Kielb, Amanda Smith, Balebail Ashok Raj, Kristin Fargher, Martin Sadowski, Thomas Wisniewski, Melanie Shulman, Arline Faustin, Julia Rao, Karen M. Castro, Anaztasia Ulysse, Shannon Chen, Mohammed O. Sheikh, Jamika Singleton-Garvin, P. Murali Doraiswamy, Jeffrey R. Petrella, Olga James, Terence Z. Wong, Salvador Borges-Neto, Jason H. Karlawish, David A. Wolk, Sanjeev Vaishnavi, Christopher M. Clark, Steven E. Arnold, Charles D. Smith, Gregory A. Jicha, Riham El Khouli, Flavius D. Raslau, Oscar L. Lopez, MaryAnn Oakley, Donna M. Simpson, Anton P. Porsteinsson, Kim Martin, Nancy Kowalski, Melanie Keltz, Bonnie S. Goldstein, Kelly M. Makino, M. Saleem Ismail, Connie Brand, Gaby Thai, Aimee Pierce, Beatriz Yanez, Elizabeth Sosa, Megan Witbracht, Brendan Kelley, Trung Nguyen, Kyle Womack, Dana Mathews, Mary Quiceno, Allan I. Levey, James J. Lah, Ihab Hajjar, Janet S. Cellar, Jeffrey M. Burns, Russell H. Swerdlow, William M. Brooks, Daniel H. S. Silverman, Sarah Kremen, Liana Apostolova, Kathleen Tingus, Po H. Lu, George Bartzokis, Ellen Woo, Edmond Teng, Neill R. Graff-Radford, Francine Parfitt, Kim Poki-Walker, Martin R. Farlow, Ann Marie Hake, Brandy R. Matthews, Jared R. Brosch, Scott Herring, Christopher H. van Dyck, Adam P. Mecca, Adam P. Mecca, Susan P. Good, Martha G. MacAvoy, Richard E. Carson, Pradeep Varma, Howard Chertkow, Susan Vaitekunis, Chris Hosein, Sandra Black, Bojana Stefanovic, Chris (Chinthaka) Heyn, Ging-Yuek Robin Hsiung, Ellen Kim, Benita Mudge, Vesna Sossi, Howard Feldman, Michele Assaly, Elizabeth Finger, Stephen Pasternak, Irina Rachinsky, Andrew Kertesz, Dick Drost, John Rogers, Ian Grant, Brittanie Muse, Emily Rogalski, Jordan Robson, M. Marsel Mesulam, Diana Kerwin, Chuang-Kuo Wu, Nancy Johnson, Kristine Lipowski, Sandra Weintraub, Borna Bonakdarpour, Nunzio Pomara, Raymundo Hernando, Antero Sarrael, Howard J. Rosen, Bruce L. Miller, David Perry, Raymond Scott Turner, Kathleen Johnson, Brigid Reynolds, Kelly MCCann, Jessica Poe, Reisa A. Sperling, Keith A. Johnson, Gad A. Marshall, Jerome Yesavage, Joy L. Taylor, Steven Chao, Jaila Coleman, Jessica D. White, Barton Lane, Allyson Rosen, Jared Tinklenberg, Christine M. Belden, Alireza Atri, Bryan M. Spann, Kelly A. Clark, Edward Zamrini, Marwan Sabbagh, Ronald Killiany, Robert Stern, Jesse Mez, Neil Kowall, Andrew E. Budson, Thomas O. Obisesan, Oyonumo E. Ntekim, Saba Wolday, Javed I. Khan, Evaristus Nwulia, Sheeba Nadarajah, Alan Lerner, Paula Ogrocki, Curtis Tatsuoka, Parianne Fatica, Evan Fletcher, Pauline Maillard, John Olichney, Charles DeCarli, Owen Carmichael, Vernice Bates, Horacio Capote, Michelle Rainka, Michael Borrie, T-Y Lee, Rob Bartha, Sterling Johnson, Sanjay Asthana, Cynthia M. Carlsson, Allison Perrin, Anna Burke, Douglas W. Scharre, Maria Kataki, Rawan Tarawneh, Brendan Kelley, David Hart, Earl A. Zimmerman, Dzintra Celmins, Delwyn D. Miller, Laura L. Boles Ponto, Karen Ekstam Smith, Hristina Koleva, Hyungsub Shim, Ki Won Nam, Susan K. Schultz, Jeff D. Williamson, Suzanne Craft, Jo Cleveland, Mia Yang, Kaycee M. Sink, Brian R. Ott, Jonathan Drake, Geoffrey Tremont, Lori A. Daiello, Jonathan D. Drake, Marwan Sabbagh, Aaron Ritter, Charles Bernick, Donna Munic, Akiva Mintz, Abigail O’Connelll, Jacobo Mintzer, Arthur Wiliams, Joseph Masdeu, Jiong Shi, Angelica Garcia, Marwan Sabbagh, Paul Newhouse, Steven Potkin, Stephen Salloway, Paul Malloy, Stephen Correia, Smita Kittur, Godfrey D. Pearlson, Karen Blank, Karen Anderson, Laura A. Flashman, Marc Seltzer, Mary L. Hynes, Robert B. Santulli, Norman Relkin, Gloria Chiang, Michael Lin, Lisa Ravdin, Athena Lee, Michael W. Weiner, Paul Aisen, Michael W. Weiner, Paul Aisen, Ronald Petersen, Robert C. Green, Danielle Harvey, Clifford R. Jack, William Jagust, John C. Morris, Andrew J. Saykin, Leslie M. Shaw, Arthur W. Toga, John Q. Trojanowki, Thomas Neylan, Jordan Grafman, Robert C. Green, Tom Montine, Michael W. Weiner, Ronald Petersen, Paul Aisen, Gustavo Jimenez, Michael Donohue, Devon Gessert, Jennifer Salazar, Caileigh Zimmerman, Sarah Walter, Olusegun Adegoke, Payam Mahboubi, Lindsey Hergesheimer, Sarah Danowski, Godfrey Coker, Taylor Clanton, Jeremy Pizzola, Elizabeth Shaffer, Catherine Nguyen-Barrera, Thomas Neylan, Jacqueline Hayes, Shannon Finley, Danielle Harvey, Michael Donohue, Clifford R. Jack, Matthew Bernstein, Bret Borowski, Jeff Gunter, Matt Senjem, Kejal Kantarci, Chad Ward, Duygu Tosun-Turgut, Stephanie Rossi Chen, Susan Landau, Robert A. Koeppe, Norm Foster, Eric M. Reiman, Kewei Chen, John C. Morris, Richard J. Perrin, Erin Franklin, Leslie M. Shaw, John Q. Trojanowki, Magdalena Korecka, Michal Figurski, Arthur W. Toga, Scott Neu, Andrew J. Saykin, Tatiana M. Foroud, Steven Potkin, Li Shen, Kelley Faber, Sungeun Kim, Kwangsik Nho, Kristi Wilmes, Lon S. Schneider, Sonia Pawluczyk, Mauricio Becerra, Liberty Teodoro, Karen Dagerman, Bryan M. Spann, James Brewer, Helen Vanderswag, Adam Fleisher, Yaakov Stern, Lawrence S. Honig, Akiva Mintz, Raj C. Shah, Ajay Sood, Kimberly S. Blanchard, Debra Fleischman, Konstantinos Arfanakis, Ranjan Duara, Daniel Varon, Maria T. Greig, P. Murali Doraiswamy, Jeffrey R. Petrella, Olga James, Salvador Borges-Neto, Terence Z. Wong, Anton P. Porsteinsson, Bonnie Goldstein, Kimberly S. Martin, Gaby Thai, Aimee Pierce, Christopher Reist, Beatriz Yanez, Elizabeth Sosa, Megan Witbracht, Carl Sadowsky, Walter Martinez, Teresa Villena, Howard Rosen, David Perry, Raymond Scott Turner, Kathleen Johnson, Brigid Reynolds, Kelly MCCann, Jessica Poe, Reisa A. Sperling, Keith A. Johnson, Gad Marshall, Christine M. Belden, Alireza Atri, Bryan M. Spann, Kelly A. Clark, Edward Zamrini, Marwan Sabbagh, Thomas O. Obisesan, Oyonumo E. Ntekim, Saba Wolday, Evaristus Nwulia, Sheeba Nadarajah, Sterling Johnson, Sanjay Asthana, Cynthia M. Carlsson, Elaine R. Peskind, Eric C. Petrie, Gail Li, Jerome Yesavage, Joy L. Taylor, Steven Chao, Jaila Coleman, Jessica D. White, Barton Lane, Allyson Rosen, Jared Tinklenberg, Michael Lin, Gloria Chiang, Lisa Ravdin, Norman Relkin, Abigail O’Connelll, Jacobo Mintzer, Arthur Wiliams, Scott Mackin, Paul Aisen, Rema Raman, Gustavo Jimenez-Maggiora, Michael Donohue, Devon Gessert, Jennifer Salazar, Caileigh Zimmerman, Sarah Walter, Olusegun Adegoke, Payam Mahboubi, Scott Mackin, Michael W. Weiner, Paul Aisen, Rema Raman, Clifford R. Jack, Susan Landau, Andrew J. Saykin, Arthur W. Toga, Charles DeCarli, Robert A. Koeppe, Robert C. Green, Erin Drake, Michael W. Weiner, Paul Aisen, Rema Raman, Mike Donohue, Scott Mackin, Craig Nelson, David Bickford, Meryl Butters, Michelle Zmuda, Clifford R. Jack, Matthew Bernstein, Bret Borowski, Jeff Gunter, Matt Senjem, Kejal Kantarci, Chad Ward, Denise Reyes, Robert A. Koeppe, Susan Landau, Arthur W. Toga, Karen Crawford, Scott Neu, Andrew J. Saykin, Tatiana M. Foroud, Kelley M. Faber, Kwangsik Nho, Kelly N. Nudelman, Scott Mackin, Howard Rosen, Craig Nelson, David Bickford, Yiu Ho Au, Kelly Scherer, Daniel Catalinotto, Samuel Stark, Elise Ong, Dariella Fernandez, Meryl Butters, Michelle Zmuda, Oscar L. Lopez, MaryAnn Oakley, and Donna M. Simpson. Interpretable discriminant analysis for functional data supported on random nonlinear domains with an application to Alzheimer’s disease. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86 (4):1013–1044, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/1013/7633881.

Ye:2024:GMR

[3988] Ting Ye, Zhonghua Liu, Baoluo Sun, and Eric Tchetgen Tchetgen. GENIUS-MAWII: for robust Mendelian randomization with many weak invalid instruments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):1045–1067, September 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

http://academic.oup.com/jrsssb/
article/86/4/1045/7629103.

Lin:2024:IVE

[3989] Yiqi Lin, Frank Windmeijer, Xinyuan Song, and Qingliang Fan. On the instrumental variable estimation with many weak and invalid instruments. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):1068–1088, September 2024. CO-DEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL http://

academic.oup.com/jrsssb/article/
86/4/1068/7629972.

Anonymous:2024:COM

[3990] Anonymous. Correction to: Optimal and maximin procedures for multiple testing problems. Journal of the Royal Statistical Society. Series B (Statistical Methodology), 86(4):1089, September 2024. CODEN ???? ISSN 1369-7412 (print), 1467-9868 (electronic). URL

                                        http://academic.oup.com/jrsssb/
                                        article/86/4/1089/7708980.  See

[3740].

Todd:1974:JNN

[3991] John Todd. John von Neumann and the National Accounting Machine. SIAM Review, 16(4):526–530, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

                                        tronic). URL https://www.jstor.
                                        org/stable/2028693. See also [42].

Moran:1981:CRD

[3992] P. A. P. Moran. Corrigenda: The random division of an interval. Journal of the Royal Statistical Society. Series A (General), 144(3):388, 1981. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981820. See [135].